A Complete Bibliography of Publications in Computers and Mathematics with Applications (2020–2029)

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: https://www.math.utah.edu/~beebe/

25 October 2025 Version 1.65

Title word cross-reference

(−Δ)^{n} [DMRV24]. (2 + 1) [KG20a, Waz20, WTQ^{+}20, LL25c]. (3 + 1) [WTSZ20]. (h − h/2) [EGP20]. 1_{σ} [Tan24]. 2 + 1 [KRO24]. 2m − 1 [ZH23b]. 3/2 [TC25]. 3D [SD25a]. 3 × 3 [XL20]. 4 [Ant21, LHSF24]. ^{1} [FKKV24, LQY22, TS21, Zhe22]. ^{2} [BEIL21, GB23, HDG23]. ^{3D} [XMQ^{+}21]. ^{∞} [CYQW21]. ^{x}3 [KGW^{+}21]. _{2} [AWA^{+}22, GO20, MAT^{+}20, ZJCG21]. _{3} [GO20, MAT^{+}20]. _{4} [ZJCG21]. _{σ} [WZ22]. α [AZHY22, CBŁP23, Tan24, WS21a, ZX24, ZQS24]. L̄1 [ZQS24]. C [BEIL21, LWL21]. C^{0} [BJyST24]. C^{1} [BLO22a, BGKV20, CHL22, GKK^{+}20, GK21, KV22, MS22, dVDR20]. C^{k} [Sch24].

C^{r} [HLW24]. chi [KN24]. d [KV20, WS25]. E [Šeb24]. G [FKKV24]. H [KCC20, LQY22, TZX23, TS21, ZJCG21, CRM22]. H(curl) [LTXZ24, WCCC25, XZT23]. H(curl, div; Ω) [KD23]. H(curl; Ω) [KD23]. H(div) [HFG23]. H(div, Ω) × H^{1}(Ω) [CRF^{+}20]. H^{1} [KPK25, RKC25, AZHY22, DS25]. H^{2} [FHM22]. H^{3} [FdOS20]. HcurlΩ [EJR24]. HdivΩ [EJR24]. hp [BHS20, DW20, CGM^{+}20]. K(x, y) = K(x − y) [LX22]. L [CWX25, GB23, KC23b, WZ22, CYQW21]. L1 [MFZX23]. L^{2} [BM24a, CDGM25, LL22a, LLXZ23, WYP23, YLW23, ZC23]. L^{∞} [LLXZ23]. L^{p} [LD21]. L_{2} [Ku24]. L_{∞} [YLZ20b]. H [AFM23]. PT [WZY23]. N [RLSC23, UR24]. n(x) = 1 [FWBY25]. O [ZJCG21]. ω [FH25]. p [BKM23, CGM^{+}20, DZ21, EAM20, ELS24, GAM22, Gun22, HGW18, JMCC24, SR20, ZR23]. p(x) [ADFP23]. P_{0} [SK21]. P_{1}(x, y) × Q_{1}(z, s) [LHSF24]. ψ [MKAM20]. q [LCX^{+}20]. Q_{1} [ZS24]. r [ABP25, OP24]. R_{0} [BDS^{+}22]. s [HUT25]. S_{N} [WJ25]. T [CJ25]. τ [HQL23]. θ [LK20, WF24, WF25]. × [MZM21]. T V − L^{2} − H^{−1} [HK20a]. T V^{p} [PMLC20]. v [MKAM20]. ε [KN20]. W^{1,p} × L^{p} [FLOQ21]. ℘ [ZR23]. Z [GCJ23].

-absolute [ZX24]. -Adaptive [OP24, ABP25, CRM22, JMCC24]. -approach [HDG23]. -based [KJ24, Šeb24]. -biharmonic [ADFP23, SR20]. -circulant [FH25]. -coercivity [CJ25]. -conforming [BLO22a]. -dimensional [KG20a, LL25c, UR24, WTSZ20, Waz20, WTQ^{+}20, RLSC23]. -DPG [LD21]. -elliptic [KD23, LTXZ24, TZX23, WCCC25]. -FEM [LSW22]. -finite [BHS20, ZS24]. -Galerkin [DS25]. -Gaussian [LCX^{+}20]. -gradient [WYP23]. -hierarchical [Gun22]. -Laplace [ELS24]. -Laplacian [DZ21, BKM23, HGW18]. -matrices [AFM23]. -matrix [KCC20]. -norm [LQY22, TS21, AZHY22, RKC25]. -norms [LLXZ23]. -optimality [KN20]. -order [HUT25]. -preconditioner [HQL23]. -refinement [EAM20]. -Robust [AZHY22, Tan24, WS21a, ZQS24, MFZX23]. -simplices [KV20]. -Stokes [GAM22]. -symmetric [WZY23]. -type [EGP20, GCJ23]. -versions [CGM^{+}20]. -water [AWA^{+}22, GO20]. -wave [WS25].

1 [Ano20-65, Ano20-66, Ano20-63, Ano20-49, Ano20-51, Ano20-60, Ano20-56, Ano20-54, Ano20-67, Ano21-52, Ano21-53, Ano21-63, Ano21-67, Ano21-57, Ano21-51, Ano21-65, Ano21-66, Ano22-52, Ano22-70, Ano22-59, Ano22-51, Ano22-56, Ano22-71, Ano22-67, Ano22-66, Ano22-68, Ano22-62, Ano22-63, Ano22-53, Ano23-57, Ano23-71, Ano23-72, Ano23-52, Ano23-51, Ano23-59, Ano23-60, Ano23-55, Ano23-53, Ano23-61, Ano23-64, Ano23-68, Ano24-66, Ano24-67, Ano24-52, Ano24-60, Ano24-58, Ano24-53, Ano24-55, Ano24-54, Ano24-47, Ano24-50, Ano24-64, Ano25-58, Ano25-60, Ano25-67, Ano25-54, Ano25-48, Ano25-55, Ano25-53, Ano25-63, Ano25-47, Ano25-61, Ano25-68, Ano25-64, HJ23, HJ25, KC23b, ŻW21a, ZKBC21]. 1-D [HJ23, HJ25, ŻW21a]. 1.0 [RHR^{+}21]. 13th [ZSH^{+}21]. 15 [Ano20-50, Ano20-58, Ano20-61, Ano20-68, Ano20-59, Ano20-64, Ano20-53, Ano20-62, Ano20-57, Ano20-55, Ano20-52, Ano21-59, Ano21-50, Ano21-64, Ano21-60, Ano21-56, Ano21-54, Ano21-49, Ano21-55, Ano21-61, Ano21-58, Ano21-62, Ano22-57, Ano22-55, Ano22-69, Ano22-58, Ano22-49, Ano22-65, Ano22-54, Ano22-61, Ano22-60, Ano22-64, Ano22-50, Ano23-56, Ano23-65, Ano23-70, Ano23-54, Ano23-50, Ano23-63, Ano23-66, Ano23-58, Ano23-49, Ano23-62, Ano23-67, Ano23-69, Ano24-49, Ano24-59, Ano24-56, Ano24-61, Ano24-63, Ano24-48, Ano24-62, Ano24-51, Ano24-68, Ano24-57, Ano24-65, Ano25-52, Ano25-56, Ano25-62, Ano25-50, Ano25-57, Ano25-51, Ano25-59, Ano25-49, Ano25-65, Ano25-69, Ano25-66]. 1D [Ale23, BGG^{+}23, FKW22, HMW21, HSJSS25, Koc22, ZXZ25]. 1D-3D [Koc22]. 1D-problems [HSJSS25]. 1D1V [RMBC24].

2 [BHD^{+}21, CWX25, CSS23, HHL^{+}21, HZA22, Vak25, ZBKC22]. 2-1 [WZ22]. 2-d [WZ24, BHD^{+}21, CSS23, HHL^{+}21, Vak25]. 2018 [ZSH^{+}21]. 2020 [Ano20-50, Ano20-58, Ano20-65, Ano20-61, Ano20-66, Ano20-68, Ano20-59, Ano20-63, Ano20-64, Ano20-49, Ano20-51, Ano20-53, Ano20-60, Ano20-62, Ano20-56, Ano20-57, Ano20-54, Ano20-55, Ano20-67, Ano20-52]. 2021 [Ano21-52, Ano21-59, Ano21-53, Ano21-50, Ano21-64, Ano21-60, Ano21-56, Ano21-63, Ano21-54, Ano21-67, Ano21-49, Ano21-57, Ano21-55, Ano21-51, Ano21-61, Ano21-65, Ano21-58, Ano21-66, Ano21-62]. 2022 [Ano22-52, Ano22-57, Ano22-70, Ano22-55, Ano22-59, Ano22-69, Ano22-51, Ano22-58, Ano22-56, Ano22-49, Ano22-71, Ano22-67, Ano22-65, Ano22-66, Ano22-54, Ano22-68, Ano22-61, Ano22-62, Ano22-60, Ano22-63, Ano22-64, Ano22-53, Ano22-50]. 2023 [Ano23-57, Ano23-56, Ano23-71, Ano23-65, Ano23-72, Ano23-70, Ano23-52, Ano23-54, Ano23-51, Ano23-50, Ano23-59, Ano23-63, Ano23-60, Ano23-66, Ano23-55, Ano23-58, Ano23-53, Ano23-49, Ano23-61, Ano23-62, Ano23-64, Ano23-67, Ano23-68, Ano23-69]. 2024 [Ano24-66, Ano24-49, Ano24-67, Ano24-59, Ano24-52, Ano24-56, Ano24-61, Ano24-60, Ano24-63, Ano24-58, Ano24-48, Ano24-53, Ano24-55, Ano24-62, Ano24-54, Ano24-51, Ano24-47, Ano24-68, Ano24-50, Ano24-57, Ano24-64, Ano24-65]. 2025 [Ano25-58, Ano25-52, Ano25-60, Ano25-56, Ano25-67, Ano25-54, Ano25-62, Ano25-48, Ano25-50, Ano25-55, Ano25-57, Ano25-53, Ano25-51, Ano25-63, Ano25-59, Ano25-47, Ano25-49, Ano25-61, Ano25-65, Ano25-68, Ano25-69, Ano25-64, Ano25-66]. 2D [AADMA23, ACLP22, Ale23, BWR21, BS23b, BZ22, BGG^{+}20, BDL20, BBM25, BDF24, Boł22, BZ21, CFP22, CWLQ22, CS25, CLL^{+}24, DN22, DFS21, DC22, DWAM25, DR24, DR25b, FKW22, FMS22, GKP25, HZ21, HJZ23, HQLJ20, HLQ24, HH23, HHR22, JWR^{+}23, JP25, JJ23, KR24, LL22a, LZY24a, LCT21, LZL^{+}21, LYL^{+}22, LSY^{+}24, LP22a, MZ23a, MKP23, NTP25, Oru20, Oru22, PS25a, PD25, QP24, RDF24, RL23, RKK25, SM22a, SN20b, SLGF20, SLYL22, VLD^{+}24, WZY23, WDYZ24, WL25, WLJF25, XZT23, XL25b, YLW^{+}22, YJZ20, ZXZ25, ZZ25b, ZS21, ZK21]. 2D-Boussinesq [BS23b]. 2D-FG [NTP25]. 2D/3D [BWR21, JWR^{+}23, RL23, WLJF25, ZS21].

3 [Pta21]. 3- [CCGV22]. 3-D [Pta21]. 30P30N [zLqLZ^{+}24]. 316L [Cao21]. 3d [DFSV22, ABL24, Ant21, BWR21, BGG^{+}23, CFG^{+}25, CAD24, CTGR24, CQG^{+}24, CZK^{+}25, CWLQ22, DLC22, EFK^{+}23, HKJ23, HLN24, JWR^{+}23, Koc22, KCC20, LZY24a, LSY^{+}24, MXLY24, MZ23b, MGA22, OKSS21, RL23, RL20, SLGF20, WNF^{+}25, WMMH25, WWG^{+}25, WPLW22, WLJF25, YL20a, YZJ^{+}22, YLQ22, YWT22, YW20, ZP24, Zha22c, ZS21, ZMS21, ZNMS23]. 3D-based [MZ23b]. 3D-heated [HKJ23].

6-simplexes [CCGV22]. 67 [JFS20].

7 [GSW21]. 75 [DZ21]. 78 [JL20]. 79

[MMAK21].

98 [LZW21a].

= [LWL21]. Ārtap [KPO^{+}21].

A-posteriori [MCW21, ZLHC24]. A-priori [ZLHC24]. A1 [Ano20-50]. A2 [Ano20-50]. ABC [HHR22]. above [EFF21]. absolute [ZX24]. absorber [HLY22]. absorbers [LXG^{+}24]. absorbing [BLX^{+}24, MBSEA24]. abstract [BBD21b, CP21]. ACA [ADD22]. accelerate [CAD24]. accelerated [DR25b, HLPW22, HC22b, KCC20, LQ24]. Accelerating [BBK24, XRS25, BGPS20, VT21]. Acceleration [DMŠ22, ADD22, CPBF21, NH24, WKC21]. accounting [GP22c]. accretive [KP20]. accuracy [CT22, CCWS20, DSF24, DR24, KWM22, KHY25, RKRRL22, SB25, SY25, SQGZ24, WH21, YZ21]. Accurate [YCCW24, ZWX24, ACXI24, AG24, AZY22, BGH20, BB23a, BBD^{+}24, FdOS20, Hag24, HJTW25, JWR^{+}23, KRO24, KJ21, LS20, Oru22, Rav20, SOPH25, SR23, SPZL21, Tan25, WZY22, WX22, WCWZ23, YTK21, ZXQQ22]. accurately [BV23]. achieved [HKJ23]. achievements [KJ22]. ACMS [GHLS25]. ACOMEN [AGNP24]. Acoustic [GCSF21, JMS^{+}24, ADD22, BFE^{+}21, CWLQ22, DMG23, FLMS24, GHL25, JYHS23, KCC20, LFGZ23, LXP^{+}24, LCE23, PSSW20, PD25, QV22, SBS24, vtWHG^{+}22, XFX^{+}24, XLCP20, XJY^{+}20, YCL20, YWLF21, ZCY22, ZCYL24, ZYC^{+}24]. acoustics [PD21]. across [AP25, CWTM20, HKJ23]. action [AHM20]. activation [ANH^{+}21, CWX22]. active [ABC^{+}21, LVDDNT25, SCD23]. active-passive [LVDDNT25]. adaptation [CHD24, DBR21, GP22c, WRMB23]. adapted [CMR22, HM24]. Adaptive [ADGG25, BFVY23, CCY24, CHY21, CWX25, DC22, GY25, GP22a, GvV22, GB21, Hak20, HSM^{+}23, KM22, Ku24, LHC23, LCC22, LC22, MB24, OP24, Shi24, Tau22, WYH20, WDGS23, WFX24, WWY25, WNF^{+}25, Alf24, ABP25, AGPP22, BBI^{+}22, BRS21, BKK^{+}23, BDGS25, Bri24, BFM^{+}25, BWMW23, CM21, CD20, CRM22, CQG^{+}24, DE22, DVA^{+}20, DMŠ22, ELS24, EGP20, FS21a, FWBY25, FP20, GYL25, GMS23, GB23, HRS24, HDG23, HWY23, HOY24, HIP23, Hop23, JMCC24, JZXR24, KWM22, KTXP21, KW20, LLSY24, LWG^{+}25b, LSGN23, LZS23, LTXZ24, LCC^{+}24, LAB23, MMC21, MJVH24, MZ23b, NC20, PMLC20, PD21, QWC^{+}23, RS21, SMDC23, SB25, SGF^{+}21, SV21, SYHN24, SMS^{+}23b, SNHH21, TL22, WWW21, WXF22, WL23, WBY23, WWCJ23, WLJF25, XWXF23, YA21, YTA23, YZW^{+}22, ZY23]. adaptivity [AMP20, AA25a, AMC23, BBK21, BX20, BDE^{+}24, CTGR24, DFHT21, HD21, TT25]. AdaPW [LWG^{+}25b]. additional [HH21, HH22]. Additive [CPS21, Par22a, AMK25, CCWC23, LAB23, dlRRG^{+}24]. additively [KJL^{+}20]. ADER [Nde25, CCGR24]. ADI [BDF24, CL21, ČPPŠ20, DZ22, DY25, DR25b, LZY24a, LZ20d, PK25, QQX21, QGQ22, RKC25, TQX23, TH23, WYHQ24, ZLPR20, ZZFL21]. ADI-based [LZ20d]. adiabatic [DBH^{+}21]. Adjoint [GS22, KFD^{+}20, OKSS21, RGS^{+}20, VSG21, WCH20, Wat22]. Adjoint-based [GS22]. adjusted [LCT22]. adjusting [BCCR20]. admittance [BHD^{+}21, BHDV22]. ADMM [TW23]. ADS [ŁWP^{+}23]. Advanced [GZG20, Lop20, JRZ21, AGNP24]. advancements [HGD21]. Advances [BDG21]. advection [BS23a, CY23, CLZW20, CSI23, DT23, DC22, DR24, FXLS25, GLÖ21, HTDT23, Hoa22, KSD^{+}25, KGK25, LDW^{+}24, LQF25, LMMMP20, ŁSSP23, MMC21, MMGS25, MMAK20, MMAK21, OTC24, RNMÖX24, RHA21, SQ23, SXLZ21, SHT^{+}24, TKMK25, TBS25, TZ21, WW24, WZ24, XSZ^{+}23, ZJJ20, ZLW23, ZHPW23]. advection-diffusion [CSI23, LDW^{+}24, MMC21, SQ23, SXLZ21, TZ21, WW24]. advection-diffusion-reaction [MMGS25, MMAK21, OTC24]. advection-dispersion [TBS25]. advection-dominated [FXLS25, ŁSSP23]. advection-reaction [HTDT23]. advective [LYZ23]. adversarial [LWG^{+}25a, MSBO23]. Aedes [dABCC25]. aegypti [dABCC25]. aeroacoustic [AAM^{+}24]. aeroacoustics [zLqLZ^{+}24]. aerodynamic [PDB22]. aerospace [ZLF^{+}22]. AFC [Jha21]. AFE [YMTS20]. AFEM [PG25]. affine [BX20, TC25]. age [DYZ22]. age-dependent [DYZ22]. agent [ZWS22]. Agglomeration [AFM^{+}24]. aggregation [AWA^{+}22, LHC^{+}22]. aggregation-based [LHC^{+}22]. aggregations [DDFZ23]. Agros [KPO^{+}21]. aided [WCZ24]. air [AADMA23, JJOG20, XLZ^{+}21]. air-vehicles [JJOG20]. airflow [TGL21]. airfoil [zLqLZ^{+}24, WS21b]. Airport [TGL21]. airway [HQCW22, LHQW25]. Al [GO20, MAT^{+}20]. ALE [NSMR25]. algae [NM23b]. Algebraic [ASV20, DDFZ23, NHCJ22, ATBT20, CAD24, DBR21, JZPG21, Loh21, SRD25]. algorithm [APK^{+}25, AHM20, ABS23, ATBT20, BF25, BMVA23, BCCR20, BCM22, CAD24, CLW^{+}24, CQG^{+}24, CSCY24, DYC^{+}22, DWY^{+}25, DYZ22, DWAM25, DR24, DR25b, FSH^{+}23, GYL25, GPRS22, GW25, GHS23, HC21, HC22a, HLCW24a, HC22b, JJZ21, JZS23, KP22d, KLRT21, KJ20, LDT24, LQLF23, LGQZ21, LHH22, LC23a, LX24a, LWL25, LC23b, LFY^{+}20, LCH23, LH23b, LQ24, LLLL24, ŁSSP23, MZHJ24, MJR20, MWRA22, Ngu20, NR25b, OEASY24, PBR24, PLVL21, PSI^{+}24, QWC^{+}23, ST23, SLYL22, SKY25, TK20, WCWZ23, WBY23, WZS24a, WZ25, WLH24, XJ21, XHL21, XHL24, YA21, YJZ20, ZP24, ZXGF22, ZZ20a, ZNH24, ZH23c]. algorithms [ABLS23, BGZ22, CDM24, FZH22, GOP25, HJTW25, JL19, JL20, KSSS23, LFF^{+}24, LT23, LZFZ22, PXLY23, Qiu21, SM23, TKMK25, TZ21, UHL^{+}24, XL25a, YK24, ZS21, ZZS24]. aliasing [LP22b]. aligned [AWA^{+}22, AAG22]. Alikhanov [MGD23a]. all-at-once [LDH24, ZS25, ZX24, ZGL23]. all-speed [SQB22]. Allen [BWMW23, CS20a, CGH24, DZ22, DLW24, DEW25, ENF24, HLJ25, JWZX25, KLK^{+}24, Lee20, LSH21, LWSZ21, LLZL24, LSGN23, LQZ23, OXY21, PL25, UHLZ24, WCC25, WWZQ25, WARX23, XHH^{+}23, YKK25, ZWMF24, ZY23, ZNH24, ZH23c]. allowing [BEIL21]. almost [YYZ22]. along [SBH^{+}21, SIAA25]. Alternating [LWL^{+}20, NSMM21, OAH^{+}24, ZX20, LYL^{+}22, MR22, QXCG20, RHA^{+}24, WC21, YQZT21]. Alternative [FSS^{+}20]. aluminium [BDJ20, SRP21]. alveolar [LHQW25]. American [ASSV20, GM22, MT25, YZL20]. among [CK23]. Ampère [SNHH21]. amplifiers [HGD21]. AMRVAC [KTXP21]. anaerobic [BAM^{+}21]. analogue [WCH23]. analyses [GLC24, JHS24, LJY^{+}25, PND^{+}24, SWZC23, TKMK25]. analysing [UuIUK21]. Analysis [AKO21, ÁCS22, AJK23, ADG23, AEF20, BDMX21, BLX^{+}24, CGO22, CLL^{+}24, DL25, EBT24, GGRBRG20b, GGRBRG20a, GSRG21, HZ24b, HW25, JZ25, KPK25, Kri20, Krz22, LHN24, LWX20, LWX21, LY23, LY25, LLLL24, MZHJ24, NP24, OSO^{+}23, OSZ22, PS25b, QL24, QS20, QFL20, RL24, RNC25, WWP^{+}25, XHL21, YHZ22, YQZL22, YYY^{+}20, ZSH^{+}21, ZZZ25, ZLL20, dABCC25, AA21, ADKH20, AHZ^{+}25, AD21, AKK^{+}25, AS23, ACLP22, AZHY22, AZ24, APS25, ASSV20, AAWJ20, BWR21, BGR23, BZDQ21, BKR24, BB23a, BLRRS20, BZ23, BBD21b, BGLZ24, BLLZ24, BGKV20, Bra20b, BRRT21, BTT20, BOSV20, CWZ21, CCGF25, CFP22, CM24, CBK24, CMY23, CHZ23, Che24, CLQL25, CWLQ22, CSS25, Dai22, DYC^{+}22, DN22, DG22a, DG22b, DSR23, DZZ^{+}24, Der23, DK23, DK24, DVA^{+}20, DKMS22, DKL24, DW20, DSF24, DBB23, FKKV24, FSH^{+}23, FK20]. analysis [FYXL21, Gal21, GP22a, GQ23, GLF25, GORBS21, GDA21, GZAR23, GCL25, HHJO20, HSWG21, HJM23a, HJ24, HLWL25, HLEF20, HH21, HH22, HC22a, HB25, HNLPP20, HYJ20, HX22, HYW^{+}24, HS20, HXZ^{+}22, HAC23, JM23a, Jaw21, JMR24, JWZ22, JCP^{+}25, JMN22, JJ23, KHD24, Kam21, KPS23a, KEUY23, KG20a, KF22, Kei21, KHH25, KSD^{+}25, KK23b, Kra23, KP24, KSJ23, KTY24, KKJ^{+}22, LL21a, LSW20, LSH21, LWSZ21, LWZ22, LL22a, LXZ23, LL23, LX24a, Li24b, Li24a, LLZ^{+}25, LML23, LYH24, LS25a, LXH24, LYZ23, LS23, LL24b, LLMC25, LZL^{+}25, LD25, LLWX25, LARGVR23, LYL25, MCD23, MFZX23, MCD24, Ma25a, MMN21, Man24, MJWD20, MLS22, MLHS23, MP23, MNB23, MS20b, NAS20, Nag24, NNH^{+}25, NTP25, NHCJ22, NERB^{+}22, PG21, PLZZ25, PCL25, PP24, PK25, QH22, QHHW20, RR22, RKC25, RGS24, SN20a, SB24b, SMV^{+}21, SMDC23, SGF^{+}21, SZ24a, SL20]. analysis [SL21, SJ21, SZ22b, SJ23, SJZ24, SLT24, SZ24b, SQ24, SS20, SJM24, SMR24b, SBY^{+}25, SRP21, SIAA25, SZCW24, SQGZ24, TT25, TA22, TSARR21, TJYL22, UHLZ24, UHL^{+}24, VBDR24, WLJ21, WM23, WLL23, WZ24, WYW^{+}24, WNF^{+}25, WWG^{+}25, WLW^{+}25, WXCL25, WQS25, WS20a, WLW^{+}22, hWCsLjH21, WCW20b, WW25, Xen20, XCLX24, XSL24a, XXLS23, XFX22, XHH^{+}23, XL25b, XJY^{+}20, YRC25, YL20a, YL22, YWL23, YJ24, Yan25b, YHL25, YLZ20a, YWL21, YTY24, ZHYL21, ZYTX22, ZJZ24, ZYD24, ZZ24b, ZZZ24, ZL25, ZHH21, ZPK22, ZLZ^{+}25a, ZL22b, ZXG^{+}24, ZHPW23, ZQS24, ZFQ24, dSMAdOW22, dlRRG^{+}24, MST25]. analysis-suitable [FKKV24]. Analysis/application [YQZL22]. Analytic [CL24b, SQC23, SD25a, ZWOM21]. Analytical [KP24, RNMÖX24, AG24, GNB24, KHY25, PZZ23, SC21a, SDJ^{+}22, SN20a, SKY25]. analyzing [HHL^{+}21, HLN24, HLY22, ZLYL20]. Anderson [NH24]. aneurysm [LH24]. angiogenesis [BGG^{+}23]. Angle [BPR22, IKST21, KV20, SS20, ZKBC21]. angle-ply [SS20]. angles [ADL^{+}21]. angular [PS25b]. Anisotropic [GP22c, AZL20, BIMS24, BCM22, CRM22, CHD24, EFFR22, HK20b, KARS23, LYP^{+}22, LAX22, LLX24, LMZ23, LP22c, LD23, MWY23, Oru25, PLZZ25, RL20, RGL21, RGS24, SCdH21, SNL23, SF25, VJH23, WBY23, WT25, WXH^{+}25, WLW^{+}22, WRMB23, XZ24, XLS22, YGZ^{+}22, YH21b, ZWCS20, ZSW23, ZWX24, ZW21b, ZZW23]. anisotropy [QP24]. annular [BPV24, HWZQ23]. anodes [FGK^{+}21]. anomalous [FC24, FYXL21, SD25a]. anomaly [KPS23a, Par24a]. antennas [SN20a]. anterior [AEF23]. Anti [LP22b, DLT20, MCZL20, MMN24]. Anti-aliasing [LP22b]. anti-plane [MCZL20]. anti-synchronization [MMN24]. any [HLW24]. aorta [ISA24]. aortic [EGF^{+}23, NDB^{+}24]. aperture [AHP20, KP22d]. Appl [DZ21, JL20, JFS20, LZW21a, MMAK21]. applicable [RZSL23]. Application [AS21, Abd25, AAU25, ABL24, BL20, CRDS20, CBLZ24, DT23, JM23a, JJOG20, KP22d, MGA22, MN23b, PN24, Par24a, QP24, SSZ20, WLX20, XXAD21, AM20, AADMA23, ABC^{+}22, Alf24, AAL24, BMH22, BFI^{+}23, BNC^{+}23, BBD21b, BHC20, CCH24, CLXZ25, DD24, RFG^{+}21, FMP^{+}24, GCSF21, GK21, HLEF20, HGD25, HKL^{+}23, JZPG21, Jaw21, Kra23, Lee21, LMBB24, LL22c, LLU24, LYLH25, MN24, NAE22, NKPA^{+}25, SRP21, SS21, WWQ20, WZC^{+}20a, hWCsLjH21, WSO^{+}23, XJ21, YY20a, Yan25a, YQZL22, Yem24, YTA23, YLWZ20, YCWS20, ZDT^{+}22, ZLF^{+}22]. Applications [GZG20, ABMS20, ABLS23, AMM23, AFM^{+}24, Axe20, Ben25, BTCV20, CSB^{+}21, CL24b, CMM23, DSS20, Din21, FSW20, GPA21, JMR24, KSB25, LNT22, LP20, MSBO23, PD21, RHR^{+}21, SD21, SBY^{+}25, VLD^{+}24, VADG22, WSS22, YKNS23, ZY21b, ZMZW24]. applied [AHOB24, BKM23, HMFM25, HNLPP20, zLqLZ^{+}24, LX24b, OKSS21, SA24, TGARGT^{+}23, VGA^{+}22, YLY20, ZK21]. approach [AI25, AKK^{+}25, AFS24, AG24, ANWW21, AWA^{+}22, AAG22, BdSC24, BZE24, BB23a, BLRRS20, BMP^{+}24, BGD22, CSB^{+}21, CTGR24, CCX21, CK23, CS23b, CZC^{+}25, CFT^{+}23, CGI^{+}22, DMŠ22, DSL24, EMSA25, ENF24, GP20, Gut20, HKJ23, HDG23, HYZZ24, IK24, KP22c, KK23a, KN20, LMP23, LLD23, Li21, LLY25, eLLAN25, MH22, MSHB22, MAT^{+}20, MKAM20, NES23, NRV22a, NE22, NSMR25, Ngo25, NP24, OKSS21, Orl23, OF20, PM23, PCL25, PSI^{+}24, SGJJ20, SFR21, SA22, SL21, SMR24b, SGSD20, SNHH21, SHXZ24, WYT23, YRC25, YZL20, YPC^{+}21, YSGD20, YOF20, YY20b, YX21, ŻW21a, ZS25, ZO21, ZTL22, dSG20]. approaches [ASB25, CO20, HUT25, HHR^{+}20, KSB25, LP22b, NSRI24]. Appropriate [CO20, SWZC23]. Approximate [AL22, BTG24, BWG25, KPJ22, TPOT20, PQSM21, SAVHB22, Tad23]. approximated [ADG23, DFK22]. approximating [JBKW23, LS21, PH23, TKT^{+}24]. Approximation [BBG23, GHAU24, HM20, Nie23, Sch21, XL24, ACDL24, AGLR21, AMN22, AMK25, ARV23, ACLP24, AGM^{+}22, BZDQ21, BKW22, BCF^{+}23, BN20, CM22, CXXZ23, CL21, CKBK24, CCWC23, CG25, CNT23, DE21, DH23, FLS23, Füh21, GHH20, GAM22, GSRG21, HRS24, HJM23b, HHR24, JM23b, KS21, KHY25, KSS^{+}20, KP23, LY23, LLW23a, LZL^{+}21, LCC22, LC22, MC24, MB24, MKM24a, Man24, MN23b, MVHÖ23, MH23, NK25, NSRI24, OCCM21, PR20, QFL20, QV22, RTA20, RKK25, SSY22, SJ21, SSD23, TCA22a, TMML23, Tür20, XXLS23, YK24, ZKBC21, ZBKC22, ZZCM20, ZSK23, Zhe22, ZH23c, ZZS24, ZMZW24, dABCC25]. approximations [APS21, GNZ24, GRR20, HSZ22, HZW21, JM20b, KKNR21, Li24a, LdO21, LYL25, SC21a, SMDC23, YZN22, YAA25, Zha22b, ZZ20b, ZX22]. April [Ano20-58, Ano21-52, Ano22-52, Ano22-57, Ano23-57, Ano23-56, Ano24-66, Ano24-49, Ano25-58, Ano25-52]. aqueous [Hoa21]. aquifers [Bou24b]. arbitrarily [MNB23]. Arbitrary [FHZ22, GNZ24, Kim21, Nde25, BHDV22, CWZ21, CM21, EH23, FJZC20, GYC25, HMW21, KTY23, LSRL24, LZG^{+}23, NFS24, SW24a, SQGZ24, VJ22, WJ25, YK20, ZLJ^{+}22, ZXXX23, ZZY^{+}24, ZAPK23]. arbitrary-order [SQGZ24]. arc [BD20, SMS23a]. arc-length [BD20, SMS23a]. architecture [EMSA25, GHSR23]. arcs [Par22c]. Argyris [GK21]. arising [APS25, BKW22, Che24, CG25, DG22b, MRP20, ZY21a]. arithmetic [AG24]. Arrhenius [ANH^{+}21]. artery [PTM^{+}23]. artificial [BDGS25, GWYR22, WDYZ24]. Asian [AG24, Rou20]. aspect [Kap24, YZ21]. aspects [GPRH^{+}24, SBS24]. assembly [CL24a]. assessing [BRR25]. Assessment [Zha21, dSMAdOW22, IBY^{+}24, KWM22, LGQZ21]. asset [CYKL23, GM22]. assimilation [ADR24, ABC^{+}22, PSDP22, SXKL24, Xu24, ZWX24]. assisted [HMW21]. associated [BB21, Hak20, QFL20, SD25a]. assumptions [CWW22]. Asymmetric [SW21, DAAEA25, SB24a]. Asymptotic [AHOB24, KP24, LGL20, LQCM20, AG25, BCT21, CGGY23, CCF^{+}21, CM24, DWY^{+}25, EP23, NS20, PV22, SSW20, WCW20b, XSL24b]. asymptotic-preserving [BCT21, SSW20]. asymptotically [WT25]. asymptotics [BD23]. asynchronicity [RLVR21]. Asynchronous [CLXZ25]. atherosclerosis [NE22]. atomic [ZTL22]. attention [Sch21]. attenuation [XJY^{+}20]. Attractors [YW20]. Aubin [FG25]. augmentation [sCpLYhG23]. augmented [ÁCS22, ACRBT23, BKK^{+}23, BGLZ24, CLZ23a, GOV20, KP22c, LZ25b, Oru25, QFL20]. August [Ano20-50, Ano20-65, Ano21-59, Ano22-70, Ano22-55, Ano23-71, Ano23-65, Ano24-67, Ano24-59, Ano25-60, Ano25-56]. AUSMPW [SQLZ21]. autoencoders [HK24]. automated [TKT^{+}24]. automated-vehicle [TKT^{+}24]. Automatic [DDFZ23, SHQL20, VBDR24, VR20, YA21, ZKTK21]. Autonomous [YMTS20, YY25, YKNS23]. Auxiliary [BHZ20, EJR24, HYZZ24]. avascular [YRC25]. average [WSS22]. averaged [PND^{+}24, YLZ^{+}24]. AX [LWL21]. axially [SN20a]. axis [KP21, SCG23]. axis-switching [KP21]. axis-symmetric [SCG23]. axisymmetric [AZL20, EP23, FL22, GNZ24, LL21a, QV22, XCFC21].

B [LYL^{+}22, AS21, BGK23, CG21, DY25, DS25, EJR24, EFK^{+}23, GP22a, GK21, HANA24, HLWL25, Kap24, KG20b, LZL^{+}25, MH25, Rou20, SRD25, SHXZ24, WLD22]. B-methods [BGK23]. B-spline [EFK^{+}23, GK21, HLWL25, KG20b, Rou20, SRD25, SHXZ24, WLD22]. B-spline-based [MH25]. B-splines [AS21, CG21, DS25, EJR24, GP22a, Kap24]. Babuska [DFHT21]. back [DLT20, HLH^{+}20a, zLqLZ^{+}24]. background [Par24a, WH25]. backtracking [Par22a]. Backward [TH21, Dai22, MKP23, SCG23, SXLZ21, TPOT20, WL22, WTSZ24, WS20b, XLLR25, ZRL^{+}22, NAE22]. Bakhvalov [ZZ23a, ZZ24a, ZZ25b]. Bakhvalov-type [ZZ23a, ZZ24a]. Balanced [LX24b, CBI^{+}23, CZLQ22, GdFFF25, JMCC24, LMD22, LZ23, RWZG24, WZZ20]. balancing [BMS22]. ball [MC24]. balls [JZ24b]. Banach [CCGI23, CG22, CP21, GORBS21, GI21]. band [CLH21]. bands [ZAPK23]. Bank [BHL^{+}23]. Barenblatt [HZ24b]. barotropic [ADR24]. barrier [AG24, RYAM21]. barycentric [LL25c, LMLZ23, Oru20, Oru22, SOPH25]. based [AS21, AHM20, Ahm24, ADD22, AG24, ACLP24, Ale23, Alf24, ASF^{+}25, ABMS20, ADL^{+}21, ABLS23, AGM^{+}22, AH25, BLLM25, BNV22, BRR25, BDP23, BGG^{+}23, BD20, BMS22, BCM22, BGD22, BSRVC20, CLLL25, CCGI23, CD20, CKFR24, CYY^{+}21, CLXZ25, CLY25b, CXJ25, CFT^{+}23, CEB24, CNT23, DDFZ23, Dai22, DYC^{+}22, DZZ^{+}24, DVA^{+}20, DTV20, DW20, DZ24, DWAM25, DNP^{+}22, DBB23, ERHL22, EJWD23, ENF24, FKKV24, FZW^{+}25, FL22, FZDS22, GS22, GORBS21, GI21, GP22c, GV21, GI20, GWYR22, GZAR23, HC21, HC23, HRS24, HZJ^{+}22, HLW23, HXW23, HLWL25, HCHM25, HLL24, HLMM20, HWL^{+}20, HLH^{+}20a, HDWL24, HHL^{+}21, HLY22, HW25, HWY25, HMX21, IBY^{+}24, IJS^{+}24, Jha21, JAHM23, JCP^{+}25, JZXR24, JJ23, KHD24, KBR20, KPO^{+}21, KM22, KLRT21, KJ24, KTFI23, Koc22, KJL^{+}20, KJ20, KSSS23, LNT22, LHC^{+}22, LZJC21, LD22, LHC23, LZF23, LT23, LZL24]. based [LpWW^{+}25, LML23, zLqLZ^{+}24, LZ20d, LYRW21, LS25b, LZL^{+}21, LCH23, LLZ^{+}23, LZG^{+}23, LLZ^{+}24, LHSF24, LLU24, LGL24, LCC^{+}24, LHTZ25, LP22b, LP22c, MWY21, MZGM25, MSHB22, Mal22, MZ23b, MB21, MH25, MNB23, MCV24, NSMM21, NW24a, NDC22, NIT^{+}20, NNT22, NNNT22, NNH^{+}25, NTP25, NNTCB25, NFFR^{+}24, NSRI24, OKSS21, Oru20, PG21, PCC21, PQSM21, Par22b, PLR21, PS25a, PL25, PBR24, PND^{+}24, PR20, PLVL21, PNT^{+}23, PTNR24, QZH23, QQX21, QSL23, QWLQ25, RNN22, RTA20, RL20, RYAM21, RLSC23, Rou20, RR22, ST20a, Šeb24, SW20a, SD21, STM21, SJ24, SFR21, SSZ20, SMS23a, SS20, SKG24, SR21, SR24, SGSD20, SPZL21, SQB22, Tad23, TCA22a, TH23, TLJ24, TCA^{+}22b, TBS25, TBO20, UuIUK21, UHLZ24, VJ22, WYS^{+}20, WWL22, WYP23, WFX24, WWZ^{+}24, WT25, WJC^{+}25, WCC25, WXH^{+}25, WH25, WL25, WQS25, WHC24, WLD22, WHH22]. based [WRMB23, WLQ22, WCY^{+}24, XKXL22, XYS^{+}24, XHY21, XLCP20, XFX22, XSZ^{+}23, XLYY25, XS25, YLQ22, YDY^{+}23, YMDW25, YLZ^{+}24, YTH^{+}21, YCL20, YAA25, YX21, Yu23, YWZ^{+}23, YZW^{+}22, ŻW21a, ZS25, ZL20b, ZZ21, ZO21, ZHZ^{+}22, ZSK23, ZZZ24, ZJ25, ZLZ^{+}25a, ZS21, Zhe22, ZS23b, ZXG^{+}24, ZWS22, ZST22, ZWW^{+}23, ZZS24, ZQ25, ZHW20, dR25a, dlRRG21, LSRL24, ZZL23]. Basic [BBD^{+}21a]. Basis [BV21, ABR20, AM22, BNB20, CBK24, DZZ^{+}24, EMSA25, GLÖ21, Gun22, HCZ25, HLX22, JKS25, JZXR24, JBKW23, LYRW21, LQF25, MFF25, MH23, Oru20, PWBS25, QWLQ25, Sch24, TO20, XCFC21, XLCP20, XFX22, Yan25a, BRÁ^{+}20a]. bat [JJOG20]. bat-inspired [JJOG20]. Battery [MJVH24, CZC^{+}25, SCD23, VGA^{+}22]. Battling [Zha22b]. Bayesian [Ben25, BCK24, HLW21, MSBO23, RBSC24]. BBM [SQ24, WLJ21]. BCD [MG25]. BDDC [KW20, TZ21]. BDF [NAE22, HZA22, MA22, TA22, WJZ^{+}24]. BDF-2 [HZA22]. BDF2 [BM21, CFG^{+}25, LLLL24, LYL25, QGQ22, SJZ22, WYHQ24, WZ25, Yan25b]. BDF2/spectral [LYL25]. BE [SP20]. beam [ADFP23, HV21, RSAG^{+}24, SJM24, WWLG24, ZWW^{+}23]. beam-deformations [ZWW^{+}23]. beams [BDS24, JSSS24, MHHR20]. Beavers [QQY^{+}23, XZ25]. beds [GORBS21]. behavior [Cao21, CCF^{+}21, EGP20, HTPV25, LXP^{+}24, LYXC21, MKM^{+}24b, MRP20, YLZL20]. behaviour [JSSS24]. Bell [GK20]. Bellman [YT22]. Belousov [KEUY23]. BEM [ADD22, AG24, EFK^{+}23, FMS22, GP22a, GvV22, HLN24, LQLF23, LXP^{+}24, MMPR20, NAS20, Pta21, SW20a, Tau22, ZYC^{+}24, ZXG^{+}24]. BEM-based [SW20a]. benchmarks [ZK21]. Bend [DFSV22]. bending [DD22, MCD23, NZ24, PCL25, SJM24, XLLR25, YLZL20]. Benjamin [CLQL25, GHAU24, KJ20, LSRL24]. Bernoulli [BLO22b, CGI^{+}22, WWLG24]. Bernoulli-type [CGI^{+}22]. Best [LCC22]. between [BM24b, DWS20, HB24, IBY^{+}24, LQLF23, Li21, MLJ^{+}20, RL24]. Beyond [BX20, XL24]. Bézier [AS21, BNC^{+}23, HL22]. BGK [LZF23]. bi [AP25, BNC^{+}23, DV22, MN23b, XGL21, EIJH21, WS25]. bi- [MN23b]. bi-cubic [BNC^{+}23]. bi-disperse [XGL21]. bi-harmonic [BZ23, DV22]. bi-material [AP25]. Bi-parametric [EIJH21]. Bi-wave [WS25]. BiCGStab [Kra20]. bicompact [BDB21]. Bidomain [BHPS24]. bifocusing [KPS23a]. bifunctional [WWY25]. Bifurcation [AKK^{+}25, WSS22]. biharmonic [ADFP23, AMV24, BLO22a, GHL24, KPO20, KPJO22, KBW23, LFY^{+}21, MCW25, Mal22, MM21, SR20, ZPK22]. bilaplacian [dlRGR20]. bilateral [LLN24a, WXCL25, ZGL23]. Bilinear [WXZW21, SZ24a, YS23, YZW^{+}22]. binarization [DH24, WH25]. binary [ADL^{+}21, XKXL22, XWXF23, ZTW22]. Bingham [BHS22]. bio [AL23, YA23]. bio-convection [YA23]. bio-heat [AL23]. bioconvection [PCC21]. biodegradable [AEF20, FGG^{+}20]. biodegradation [AM20]. biodenitrification [AM20, AO24]. biological [YYY^{+}20, ZLZ^{+}25a]. biomass [SB21]. biomass-fueled [SB21]. biomolecular [SHWC22]. biorthogonal [DLN24]. Biot [BS21, CLR23, GIRB25, HFG23, HZ24b, KKNR21, LR23, WCWZ22, WH20]. Bivariate [BDS^{+}22]. Black [RYAM21, KMP20, KN25, Ull20]. blade [HJP23, YL20b]. blanket [BAM^{+}21]. blast [BNQ21]. blasts [Abd24]. blended [MG25]. Blending [Kra23, YLY20]. Bloch [XLLA20, YLL24, ZLT^{+}21]. Bloch-Torrey [ZLT^{+}21]. Block [FK20, FH25, JL23, NSRI24, SS22, Ahm24, ATBT20, Axe20, BEG^{+}21, LY24, LZ25c, MFZX23, MC25, MZM21, WZS^{+}24b, XW24a, XL20, XW24b, ZWCS20]. Block-centered [JL23, LY24, MFZX23]. block-coupled [MC25]. block-structured [BEG^{+}21]. blocks [LDT24]. blood [AI25, ISA24, LH24, LCC^{+}24, SPC^{+}23]. Blow [DZ21, HGW18, BGK23]. Blow-up [DZ21, HGW18, BGK23]. Board [Ano20a, Ano20b, Ano20-27, Ano20-28, Ano20-29, Ano20-30, Ano20-31, Ano20-32, Ano20-33, Ano20-34, Ano20-35, Ano20-36, Ano20-37, Ano20-38, Ano20-39, Ano20-40, Ano20-41, Ano20-42, Ano20-43, Ano20-44, Ano20-45, Ano20-46, Ano20-47, Ano20-48, Ano20c, Ano20d, Ano20e, Ano20f, Ano20g, Ano20h, Ano20i, Ano20j, Ano20k, Ano20l, Ano20m, Ano20n, Ano20o, Ano20p, Ano20q, Ano20r, Ano20s, Ano20t, Ano20u, Ano20v, Ano20w, Ano20x, Ano20y, Ano20z, Ano21a, Ano21b, Ano21-27, Ano21-28, Ano21-29, Ano21-30, Ano21-31, Ano21-32, Ano21-33, Ano21-34, Ano21-35, Ano21-36, Ano21-37, Ano21-38, Ano21-39, Ano21-40, Ano21-41, Ano21-42, Ano21-43, Ano21-44, Ano21-45, Ano21-46, Ano21-47, Ano21-48, Ano21c, Ano21d, Ano21e, Ano21f, Ano21g, Ano21h, Ano21i, Ano21j, Ano21k, Ano21l, Ano21m, Ano21n, Ano21o, Ano21p]. Board [Ano21q, Ano21r, Ano21s, Ano21t, Ano21u, Ano21v, Ano21w, Ano21x, Ano21y, Ano21z, Ano22a, Ano22b, Ano22-27, Ano22-28, Ano22-29, Ano22-30, Ano22-31, Ano22-32, Ano22-33, Ano22-34, Ano22-35, Ano22-36, Ano22-37, Ano22-38, Ano22-39, Ano22-40, Ano22-41, Ano22-42, Ano22-43, Ano22-44, Ano22-45, Ano22-46, Ano22-47, Ano22-48, Ano22c, Ano22d, Ano22e, Ano22f, Ano22g, Ano22h, Ano22i, Ano22j, Ano22k, Ano22l, Ano22m, Ano22n, Ano22o, Ano22p, Ano22q, Ano22r, Ano22s, Ano22t, Ano22u, Ano22v, Ano22w, Ano22x, Ano22y, Ano22z, Ano23a, Ano23b, Ano23-27, Ano23-28, Ano23-29, Ano23-30, Ano23-31, Ano23-32, Ano23-33, Ano23-34, Ano23-35, Ano23-36, Ano23-37, Ano23-38, Ano23-39, Ano23-40, Ano23-41, Ano23-42, Ano23-43, Ano23-44, Ano23-45, Ano23-46, Ano23-47, Ano23-48, Ano23c, Ano23d, Ano23e, Ano23f]. Board [Ano23g, Ano23h, Ano23i, Ano23j, Ano23k, Ano23l, Ano23m, Ano23n, Ano23o, Ano23p, Ano23q, Ano23r, Ano23s, Ano23t, Ano23u, Ano23v, Ano23w, Ano23x, Ano23y, Ano23z, Ano24a, Ano24b, Ano24-27, Ano24-28, Ano24-29, Ano24-30, Ano24-31, Ano24-32, Ano24-33, Ano24-34, Ano24-35, Ano24-36, Ano24-37, Ano24-38, Ano24-39, Ano24-40, Ano24-41, Ano24-42, Ano24-43, Ano24-44, Ano24c, Ano24d, Ano24e, Ano24f, Ano24g, Ano24h, Ano24i, Ano24j, Ano24k, Ano24l, Ano24m, Ano24n, Ano24o, Ano24p, Ano24q, Ano24r, Ano24s, Ano24t, Ano24u, Ano24v, Ano24w, Ano24x, Ano24y, Ano24z, Ano25a, Ano25b, Ano25-27, Ano25-28, Ano25-29, Ano25-30, Ano25-31, Ano25-32, Ano25-33, Ano25-34, Ano25-35, Ano25-36, Ano25-37, Ano25-38, Ano25-39, Ano25-40, Ano25-41, Ano25-42, Ano25-43, Ano25-44, Ano25-45, Ano25-46]. Board [Ano25c, Ano25d, Ano25e, Ano25f, Ano25g, Ano25h, Ano25i, Ano25j, Ano25k, Ano25l, Ano25m, Ano25n, Ano25o, Ano25p, Ano25q, Ano25r, Ano25s, Ano25t, Ano25u, Ano25v, Ano25w, Ano25x, Ano25y, Ano25z]. bodies [BD23, CRB24, GRR20, Li21, MGBN22]. body [MIOK21, QD20, RR23, SD21, ZBCS24]. Boltzmann [GCSF21, APGK24, ABMS20, ADL^{+}21, ABLS23, ABS23, AG25, BLLY24, Bel24, BP20, BABS20, CLLL20, CCWS20, CYY^{+}21, CYC^{+}23, CRPdT20, DT23, DLT20, EFF21, EGF^{+}23, EFS24, ENF24, FGK20, FS21b, GBK23, GML^{+}20, GDŁW23, HHR^{+}20, HLH^{+}20a, HLH^{+}20b, HHY^{+}22, HQCW22, HMŁW^{+}24, HDWL24, HEDT20, HHR24, HKL^{+}23, Ima20, JHS24, KRESA21, KGB24, KFD^{+}20, KTFI23, KWK^{+}20, KKA^{+}21, Kri20, KOMK20, LMK^{+}21, LLX23, LLX24, LpWW^{+}25, zLqLZ^{+}24, LS25b, LZB25, LHQW25, MWYY22, MIOK21, MLJ^{+}20, NIT^{+}20, OF20, PGK20, PLR21, PWWJ23, PWJ^{+}24, QhCZ^{+}20, RST^{+}22, RLF21, RLSC23, RGS^{+}20, SGJJ20, SB24a, SMŁWL25, SCD25, SLX^{+}23, ST20b, SLC^{+}24, TMML23, VL21, VJ24, WWQ20, WZC^{+}20a, WZC^{+}20b, WCZ24, WWP^{+}25, WZCS20, WCY^{+}24, XL22, XLZ^{+}21, XHH^{+}23, XTZZ24, YOF20, YYY^{+}22, YCWS20, YWZ^{+}23, ZKTK21, ZCY22, ZCYL24, ZTW22, ZHHL23, ZWCS20, ZZ20b, ZYS22, ZDLL22, ZLD^{+}23]. Boltzmann-based [LpWW^{+}25]. Bona [CLQL25, GHAU24, KJ20, LSRL24]. bond [CRDS20, TLJ24, WT25, YMDW25]. bond-based [TLJ24, WT25, YMDW25]. bonds [LZ20d]. bordered [JL19, JL20]. Bose [GM21a, WTQ^{+}20, XHM20]. BoSSS [KWS21]. both [Li24a, YLQ22]. bottom [CBI^{+}23, LMD22, XDL^{+}25]. bounce [DLT20, HLH^{+}20a, zLqLZ^{+}24]. bounce-back [HLH^{+}20a]. Bound [MJZ25, DHVT20, DEW25, PL25, RWZG24, WXF22, WWZQ25, YLZ20b, ZWMF24]. Bound-preserving [MJZ25, DEW25, PL25]. boundaries [ABLS23, CCGR24, PWMC23, YX21]. Boundary [BBS21, CM20a, CRPdT20, DFS21, Gut20, HXZ^{+}22, Kar24, KCC20, LL20a, vtWHG^{+}22, Zha22e, ADGG25, Alt23, AAK24, BLX^{+}24, Bar22, BTCV20, BM24a, BBS22, CWZ21, CL24a, CLZ23a, CFP22, CRRS21, Che24, CLW^{+}24, CZ20, CWLQ22, CSS23, CGI^{+}22, DKMS22, DT24, DH23, DLT20, EFS24, ERG^{+}24, FYM23, FM25, FYXL21, GS22, GP23, GS23, GO23, HMF24, HRS24, HZ21, HTBMC20, HK22, HSR^{+}23, HYW^{+}24, HHL^{+}21, HJ23, HJ25, HYL^{+}25, JKC24, KC22a, KPJO22, KJ24, KGB24, KTFI23, LNT22, LL20b, LZ20b, Li20, LZJC21, LDW^{+}24, LpWW^{+}25, LQW22, LX24b, LC21, LLZ^{+}24, LZY24b, LSZ24a, LZLN25, LJY^{+}25, MRAC24, Man24, MM21, MZ23b, MJR20, MD21, MK20, MN23b, Nag24, NR25a, Ngu20, OEASY24, PWWJ23, PP24, PS21, QhCZ^{+}20, RLF21, ST20a, SD25b, SMŁWL25, SD21, SLX^{+}23, SXCB24, SLGF20, SPZL21, SZCW24, US20, Vak25, WWQ20, WDGS23]. boundary [WCZ24, WWLG24, WDYZ24, WSO^{+}23, WZCS20, Xen20, XZS^{+}23, YL20a, YLW^{+}22, YS23, YMTP25, YWL21, YLWZ20, ZLYL20, ZWOM21, ZTW22, ZLQ23, cZsCyC23, ZZ20b, ZHH21, ZCA23, ZQS24, BL20]. boundary-domain [CFP22]. boundary-lattice [RLF21, SLX^{+}23, WWQ20]. boundary-type [YL20a, YLW^{+}22]. bounded [ES20]. boundedness [LZ20e]. Bounds [DTW23, BMS22, GSRG21, JCJ24, Ku24, KKNR21, VLD^{+}24, Wol20, ZY21a]. Boussinesq [BS23b, AGM^{+}22, CRRS21, HC22a, HYLH22, YJ24]. box [AMPB^{+}22]. brain [PVC^{+}25]. Bratu [IZ20]. break [AS21]. breakup [ZDLL22, ZLD^{+}23]. breast [YMTP25]. breather [JXK20]. Brezzi [DFHT21]. brick [ZMS21, ZNMS23]. bridging [dR25a]. Brinkman [ALOR25, Gut20, LZBC24, MS20b, ZMS21, ZHW20]. brittle [SMS23a, SYHN24]. Brownian [LL24b, MW20, YHL25]. brush [PCM20]. bubble [DdSP21, HSM^{+}23, ZLJ^{+}22]. bubbles [WTSZ20]. buckling [AA21, KHD24, LML23, LYH24, MMP21, MNB23]. Bulkley [AI25, FNGDV23, HMŁW^{+}24]. Buongiorno [RG21]. buoyant [ZZZC24]. BURA [HLMM20]. Burgers’ [CO20, CC25, GHAU24, HCHM25, JSBZ22, LS25a, MKP23, RLSC23, SC21a, ySRC22, WLW^{+}25, ZL23c, BWG25, CLQL25, JWR^{+}23, KJ20, KMS20, LSRL24, MKM24a, jWC23, YL22]. Burton [SXCB24].

Cable [GLL^{+}20b, ZLJT20]. CADI [CC20]. Cahn [GGT24, SXKL24, AMK25, BWMW23, CS20a, CL22a, CMY23, CGH24, DZ22, DLW24, DEW25, ENF24, FRW22, GLF25, HLJ25, JWL20, JAMX24, JWZX25, KTY23, KLK^{+}24, Lee20, LKC^{+}24, LL25a, LSH21, LWSZ21, LLZL24, LSGN23, LQZ23, NR22, OXY21, PL25, SBNR23, Tan25, UHLZ24, WZY22, WCC25, WWZQ25, WHH22, WARX23, XFS23, XHH^{+}23, YKK25, YLZ20b, ZWMF24, ZZW20, ZZ21, ZO21, ZY23, ZNH24, ZH23c, ZFQ24]. calcium [JJ23]. calculation [Ant21, BDJ20, LFGZ23, LLJ23, MJR20, PWC^{+}23]. calculus [JH24]. Calderón [Li25]. cancer [AOL22, ASF^{+}25, AMC23, SAVHB22, YMTP25]. canonical [CPBF21]. CaNS [CPBF21]. cantilever [RSAG^{+}24]. canyon [WSGT21]. canyons [AADMA23]. cap [KHD24]. capacity [HNV25, NRV22b]. capillary [WYS^{+}20]. capsules [CRPdT20]. capturing [HCHM25, KK25]. Caputo [Din21, FSW20, RKK25, SD25a, SHM21]. Caputo-Riesz [SHM21]. carbon [AAS^{+}20, FGK^{+}21, HTPV25, PCM20, WCW20b]. carbon-nanotube [PCM20]. Carleman [LNT22, LLN24b, NNT22]. Carleman-based [LNT22, NNT22]. Carlo [AHM20, BW21, LLY25]. Carlo-coupled [LLY25]. carpet [HB25]. Carreau [MLHS23]. cartesian [DAAEA25, zLqLZ^{+}24, MIOK21, XHY21, ZZY^{+}24]. cartoon [RGL22]. cascaded [GML^{+}20, SCD25, XTZZ24]. cascades [HLEF20]. cascadic [SSW20]. Case [IBY^{+}24, BX20, EJR24, WJC^{+}25, WZXN21]. Casimir [SBS24]. Casson [AGA^{+}21, KC23a]. casting [YLWZ20, YWPS24]. catalytic [AGA^{+}21]. Cattaneo [ANH^{+}21, AAG22, BWZ23]. Cauchy [AAL24, BJG23, BCM22, CLXZ25, CV20, GPA21, MH23, NRV22a, SDG20, WW20a]. Cauchy-type [GPA21]. cavitating [EH23, KJ21]. cavities [KC23a, MGD23b, YZ21]. cavity [ACXI24, BRÁ^{+}20a, Fur25, GO20, HBC^{+}21, PM23, SPC^{+}23, TSARR21, WX22, ZZZC24]. CBS [YQZL22]. CBS-SEM [YQZL22]. CDNNs [LLY25]. Cell [CSI21, BRR25, Bar22, BDJ20, DAAEA25, FGK^{+}21, GHRD20, LWZ22, LLW23a, MJVH24, MWY23, VSG21, WCZ24, WKGH21, ZYM24, ZPK22]. cell-centered [LLW23a, MWY23, ZPK22]. Cell-vertex [CSI21]. cells [JJ23, KWM22, LARGVR23, MGBN22, SPC^{+}23, ZDW^{+}20]. cellular [MB21]. CEM [PFCZ23]. CEM-GMsFEM [PFCZ23]. centered [DWS20, HSY25, JL23, LY24, LLW23a, LMZ23, MFZX23, MWY23, ZPK22]. Central [Nde25, EH23, MZCK25, QD23]. Central-Upwind [Nde25, MZCK25]. centred [MC25]. cerebral [LH24]. certain [CDM24]. Certified [WARX23, BRÁ^{+}20a]. CEV [RYAM21]. CFD [JMCC24, ZvNH25, ZY21b]. CFL [TZ25]. CG [BFE^{+}21]. CH [ZJCG21]. Chaboche [RUMT^{+}24]. chain [MSBO23, ZTL22]. challenges [KJ22, LE24]. Chan [LQZ23]. Change [ZYQ^{+}24, Cha24, EBDY22, FL22, HHY^{+}22, JFS14, JFS20, MBSEA24, TWH^{+}22, ZK21]. channel [AAWJ20, BPR22, CBK24, MIOK21, MAT^{+}20, QD23, YY23]. chaos [BNSZ24]. Characteristic [KGB24, LYXC21, BRRT21, FXLS25, LZF23, QFL23, SWW22, WZ24, ZSPW20]. characteristics [JWW^{+}21b, KP21, LXG^{+}24, ZD21]. characterization [BEIL21, KJL^{+}20]. charge [LXJ^{+}21, RC25, VGA^{+}22, WMMH25, ZD22, ZZ24b, ZST22, ZZS24]. charge-conservation [ZST22, ZZS24]. charge-conservative [RC25, ZD22, ZZ24b]. Chebysev [OS24]. Chebyshev [DYC^{+}22, LVDDNT25, NML23, Pel23, PCL25, SKG24, YSB^{+}24]. checking [YJZ20]. chemical [BGG^{+}23, CM24, HEDT20, JSBZ22]. chemical-driven [CM24]. chemically [AWA^{+}22]. chemo [GGRBRG20b, GGRBRG20a, SCD23, WaZWZ23]. chemo-mechanically [SCD23]. chemo-repulsion [GGRBRG20b, GGRBRG20a]. chemo-repulsion-Navier [WaZWZ23]. chemotaxis [AO24, BGG^{+}20, BOSV20, CCGI23, HHNV23, IQS20, PCC21, ZXZF20]. chemotaxis-driven [PCC21]. chemotaxis-Navier [CCGI23]. chemotherapeutic [SAVHB22]. Chew [SBY^{+}25]. Chimera [SGSD20]. China [ZCZ^{+}21]. chip [CRDS20]. chiral [AAA24]. Choice [CKS22]. choosing [JAHM23]. Christov [ANH^{+}21, AAG22, BWZ23]. chromatography [MAT^{+}21]. Ciarlet [YJ24]. Circulant [PQSM21, FH25, LZFZ22, ZZL22]. Circulant-based [PQSM21]. circular [AMPB^{+}22, AZ24, ASV20, DBH^{+}21, DKH22, SN20a, YX21]. circumvent [DFHT21]. cities [dCG20]. clamped [KPJO22]. class [BMZ20, BLF23, CL24b, DLW24, DL25, DZ21, GLL20a, GLLZ23, HGW18, JL19, JL20, LS21, LMLZ23, LLZ^{+}24, MW20, MZM21, MD21, MZZ^{+}25, NT21, OSZ22, PQSM21, SSD23, SSdA20, hWCsLjH21]. classical [GSRG21, VGA^{+}22]. classifier [NR22]. cloak [GY25, WWY25]. cloaking [FLW20]. close [BKT^{+}23]. close-contact [BKT^{+}23]. closed [CQG^{+}24, ZL23a]. closed-loop [CQG^{+}24, ZL23a]. closest [LQF25, Yan24]. closure [AS23, MCK^{+}23]. cloud [PWBS25]. clouds [HK22]. clustered [FWBY25]. clustering [CDPL24, HK24, MN24]. CNTRC [MJWD20]. CO [ZJCG21]. coarse [BDJT21, GHS23]. coarsening [DDFZ23, HC21, He23b]. coating [AMPB^{+}22]. COCR [LWL21]. code [KKA^{+}21]. coefficient [ABFPA24, BGH20, HTDT23, KS21, KLY25, KKS25, LT23, LLZL24, LLXY23, LLXY25, MFZX23, Ma25b, PT24, WXZW21, YS23, YYSL20, ZCL20]. coefficients [AGPP22, BLO22b, CLYH22, CEL20, HLX20, HCY^{+}24, JL23, LSW22, LNNP24, Liu25a, Ngo25, PH25, PS25a, Pło22, RKC25, TLJ24, WZ22, WT25, XLLA20, YL25, YY25, ZLT^{+}21]. coercive [CJ25]. coerciveness [Zha23]. coercivity [CJ25, QN20]. cohesive [YA21]. cold [ZCZ^{+}21]. Cole [BWR21, RLSC23]. collapse [DdSP21]. Collective [CM21]. collision [ADL^{+}21, ENF24, OF20]. collocated [ADB^{+}22, NERB^{+}22]. Collocation [HWY25, BDF24, BDS^{+}22, CLZ23a, CCX21, CD20, CPS21, CFT^{+}23, DZZ^{+}24, FKW22, FS21a, FYXL21, GKHU22, JM23a, JZXR24, LZ20b, LLG25, LL25c, LZL^{+}21, LGZC23, LP22c, LN22, MHHR20, NE22, NML23, Oru25, QC24, SOPH25, SLD25, SD25a, SRD25, TjW24, TO20, TQX23, TLJ24, Tür20, WYHQ24, XLCP20, XFX22, XLYY25, YMDW25, YMTP25, YLZ^{+}24, ZDWR22, ZSL22]. Collocation-based [HWY25]. color [LP22b]. Combination [LMLZ23]. Combined [AGA^{+}21, AA20, BHD^{+}21, BSRVC20, JJ23, LLLL24, MKP23, SGJJ20, SR22, TFWX23, WYW^{+}24, WYLL25, YKNS23, Zhe22]. combustion [KRESA21]. Comments [DZ21]. common [YTY24]. Compact [CMM23, LP20, AM22, BZ22, CT22, CLL21, CLLG23, DY25, DR25b, FHM21, FXLS25, GLLZ23, HC22a, HLPW22, JZHZ22, LC23a, LZY24a, LCT22, MG25, MN23b, PM23, PK25, RWZG24, RKC25, ST20b, YL22, ZCYL24, ZAH23, ZLL23, ZQZ20]. Companion [HLL24]. Companion-based [HLL24]. comparative [LFF^{+}24, OAD24, TKT^{+}24]. Comparison [BBS22, CFTC25, KOMK20, MCK^{+}23, WZXN21, BMK21, BCF^{+}23, BDJT21, Bri24, CS20b, Kap24, LMBB24, MMGS25, RSAG^{+}24, SSK^{+}23]. Comparisons [DSS20, LAL^{+}24]. compatible [WT25]. compensation [FZW^{+}25]. competition [TSW^{+}25]. complement [Bot20]. completeness [ZLJ^{+}22]. complex [AD20, AXLW24, AAWJ20, DK23, DK24, FM25, HANA24, KJL^{+}20, LL20b, LWL21, LL21b, LFGZ23, LQW22, SMŁWL25, VJ24, WNF^{+}25, XHY21, XLCP20, XP22, YKMZ21, ZXG^{+}24]. complex-geometry [XLCP20]. complex-scaled [FM25]. complexity [BFM^{+}25, DKH22]. complicated [YWLF21]. component [Fan23, WYT23, WLH24, ZBKC22, ZGW23]. components [RGL22]. Composite [BHL24, LLG25, jWC23, Abd24, DSL24, DWY^{+}25, Hag24, HYH23, HYL^{+}25, HTPV25, LVDDNT25, MMP21, MNB23, Pta21, QP24, WWZ^{+}24, YLZL20]. composites [BNB20, MR22, Mit22]. composition [SGSD20]. comprehensive [CZC^{+}25, KWM22, PTM^{+}23]. compressed [CYY^{+}21]. compressible [ADB^{+}22, AEZ24, CYC^{+}23, CCGR24, CQS25, GPL21, HEDT20, JMCC24, JZHZ22, JZXR24, KP22b, LS25b, LZB25, PFCZ23, RZC21, RDP20, SMR^{+}24a, SJ24, XJ21, YPC^{+}21, YZN22, ZH20, ZCXG20, ZYS22, ZZY^{+}24]. compressible/incompressible [GPL21]. Compressive [BTT20, LCN20, YZW^{+}22]. comprising [AWA^{+}22, AAG22]. Comput [DZ21, JL20, JFS20, LZW21a, MMAK21]. computable [KKNR21]. Computation [NR25a, BGPS20, BRS21, CFG^{+}25, CLZ^{+}24, DWAM25, GM22, JMR24, KN20, KC23a, MRN^{+}20, SBS24, YHZ22, Yan24, YYY^{+}20, YTT^{+}22, YWLF21]. COmputational [AGNP24, GPRH^{+}24, Lop20, OAA^{+}22, BMP21, Bri24, CT22, CHJR21, DSL24, DTV20, EGF^{+}23, GKHU22, HGD21, KAKK23, KSSS23, KR24, Oru22, PTM^{+}23, RHB25, TRSM23, Xu24, YD23]. Computationally [ABP25, GP22b, HYZZ24]. Computations [GZG20, DNP^{+}22, LYP^{+}22, RG21, VBDR24, ZTL22]. computers [Fia21]. computing [AHM20, BDS^{+}22, DdSP21, EH23, HLL24, NAE22, XLS24]. concave [WL20]. concave-convex [WL20]. concentrated [HSZ22]. concentration [CML^{+}25, Hoa21, IJS^{+}24, RBSC24, SGJJ20, ZL20b]. concentrator [HB25]. concepts [BBD^{+}21a]. concrete [YA21]. Concurrency [PVC^{+}25]. condensate [WTQ^{+}20]. condensates [GM21a, XHM20]. condensation [DWAM25]. Condenser [NRV22b]. condensing [YL20b]. Condition [KC21, AAK24, BLX^{+}24, DFS21, DP25, DKMS22, DLT20, EAM20, FYM23, GV21, HMF24, HTBMC20, IKST21, KJ24, KGB24, KV20, Krz22, LSZ24b, MRAC24, NIT^{+}20, SLGF20, WSO^{+}23, YLWZ20, ZLQ23, Zha22e]. Conditioning [KF22]. conditions [Adj22, ADGG25, Alt23, BTCV20, BM24a, BGLZ24, BPR22, CSS23, DDNS21, DFHT21, EFS24, GS22, GO23, HK22, HYW^{+}24, KPJO22, LNNP24, Lee24, LZ20b, LZY24b, LJY^{+}25, MM21, MJR20, MD21, MK20, OEASY24, QQY^{+}23, SMR^{+}24a, ST20a, SD25b, SGZ^{+}23, US20, Vak25, WXZW21, WDYZ24, XZ25, XZS^{+}23, YS23, YWL21, ZCA23]. conducting [Du24, KP22d]. conduction [BCM22, CZ20, CFT^{+}23, ČPPŠ20, DD20, DWY^{+}25, FYXL21, Gra21, HZ24a, HYJ20, QX23, SD25a, SL20, TJYL22, Vab22, WQ24, XFX22, YWL21, ZLYL20, ZRL^{+}22, ZZL23, ZZ25a]. conductive [TSARR21, XKXL22]. conductivity [BD23, JM23a, MLHS23, Par22b, QP24, ZLYL20]. conductor [LSV21, LSV24]. cone [IBY^{+}24]. configuration [ZLD^{+}23]. confined [PP24]. Conformal [Rob25, HNV25, HWY23]. Conforming [NW24b, WN24, AMV20, BDMX21, BHK^{+}24, BLO22a, DDG23, EFK^{+}23, FWBY25, HFG23, LHSF24, MKM24a, MST25, NAS20, PLR21, SK25, ZQ25]. Conforming/non [SK25]. Conforming/non-conforming [SK25]. congestion [CML^{+}25]. conical [HWZQ23, LJY^{+}25, MAT^{+}20]. conical/cylindrical/spherical [HWZQ23]. conjugate [FK20, Ima20, SBH^{+}21, ZHHL23]. connection [HUT25]. consensus [ZWS22]. conservation [BCT21, BS23c, BDGS25, DTW23, GCJ23, Haj21, KG20a, Kim21, KBB^{+}21, OAD24, Pel23, SK21, ZST22, ZZS24]. Conservative [APPS25, LX24a, CWC22, CYQW21, FHZ22, FX22, FXLS25, GDA21, HC22a, HFG23, HJXW23, JZ20, JP25, KTY23, Lee20, MZCK25, QFL23, RC25, SZ22a, TH23, WZ24, WMMH25, WH20, XHH^{+}23, XF23, XLSF25, YTK21, YWL23, ZD22, ZZ24b, ZW21b, Nde25]. conserving [Jog25, LMBB24, WLC24]. considering [BLX^{+}24, FSH^{+}23, NIT^{+}20, SMR^{+}24a, SB21]. consistency [Alf24, WYP23]. consistent [GP25, HJTW25, Mit22, PDB22, SMŁWL25, WWCJ23, ZGW23, Zha22d]. consolidation [AZL20, HFG23, LR23, WCWZ22]. constant [ABFPA24, HLX20, Sau23, TSARR21]. constant-coefficient [ABFPA24]. constants [PV20]. constitution [YCZ^{+}21]. constitutive [CLZ^{+}24, PGR23, RUMT^{+}24]. Constrained [HNV25, AGM^{+}22, BV21, BGZ22, CSB^{+}21, CS23b, FWLD22, GPRS22, GS23, HLZ24, IJS^{+}24, SV21, SHWC22, TH25, YLSY24, YY20b, ZZ23b]. constraint [ERHL22, HJZ23, TS21]. constraints [BHS20, BHS25, BKK^{+}23, BJyST24, CSCY24, DDG23, ELN^{+}20, GLS25, JM23b, JW20, KW20, LDW^{+}24, LN22, SCD23]. constructed [KKNR21]. constructing [Fan23]. Construction [ABFPA24, BHDV22, DZ20, JZPG21, LLZ^{+}24, BEIL21, FKKV24, GK21, OF20]. Constructive [KMN25]. consuming [LCT22]. consumption [ZL20b]. contact [ABC^{+}21, BB24, BKT^{+}23, DFK22, DWAM25, HHJO20, HC22b, JWZX25, LSV21, LXH24, SHXZ24, VADG22, WXCL25, ZKBC21, ZNC24]. containing [KPO20]. contaminant [RBSC24, SQC23]. contamination [HLZ24]. context [Gie23]. Continuous [ADR24, LdO21, BAKR24, CRM22, CM23, CXXZ23, CG22, DO23, GQZ23, GGRBRG20a, LZFZ22, OKSS21, VSG21, YY25, YLWZ20, YWPS24, ZLL20, ZL23c]. Continuous/discontinuous [LdO21]. contour [MZWD23, XLS24]. Contraction [YKMZ21, NNT22]. contractivity [YY25]. contrast [BHD^{+}21, CCH24, GML^{+}20, vtWHG^{+}22]. contrasts [YWZ^{+}23]. Control [MRAC24, AFS24, AS23, AGM^{+}22, BHS20, BHS25, BDE^{+}24, BGZ22, BJyST24, CSB^{+}21, CRRS21, CDM24, CSCY24, CEB24, DE22, DSR23, DH23, ELN^{+}20, FLW20, FF24, GS22, GLS25, GP23, GS23, HHM21, HSY25, HLZ24, HJ23, Ila20, KUK20, Kar24, KPS23b, KN20, KVV23, LLSY24, LLSY25, LCW22, LAL^{+}24, LC23b, LC21, LSZ24a, LVDDNT25, LYL25, Man24, MP23, NE22, NM23b, RNC25, RGS^{+}20, SK25, SW20b, TS21, TKK23, WLX22, WZD24, WWLG24, XW24b, YY20a, YL20b, YTH^{+}21, YWPS24, ZCL20, ZS25, ZL20b, ZC22, ZZH23, ZZ25a, ZWS22, ZWW^{+}23]. control-net [MP23]. control-volume [HSY25]. controllability [KPJ22]. controller [ASF^{+}25, BMS22]. controller-based [ASF^{+}25]. controlling [YY20b]. Convection [JSZ24, ACXI24, AA25b, ÁCS22, AN23, ANN25, BRÁ^{+}20a, BLLZ24, CCY24, CGGY23, CZ23, CMR22, CEL20, CSS23, CS25, DK24, EBDY22, FZDS22, GO20, GO25, HHM21, HDWL24, HYJ20, HRvdZ20, Ima20, Jha21, JWR^{+}23, JMN22, KSJ23, LLD23, LD21, LL22b, LLX24, LSF25, LS23, LL24a, MZ23a, MJZ25, MN21, MN23a, PXLY23, QFL23, RL20, RST^{+}22, RR22, SG23, SB24a, SADW21, SSY22, SZ24a, SL20, SZ24b, SSD23, SBH^{+}21, SR24, TWH^{+}22, VR20, VBDR24, WX22, WJZ^{+}24, WZCS20, WCY^{+}24, WLJF25, XFLZ22, YA23, YH21a, YZ21, YZJ^{+}22, YWPS24, ZL22a, ZF22, ZZ23a, ZYQ^{+}24, ZWCS20, ZZ25b, ZSPW20, ZZB20, ZK21]. convection-cooling [HHM21]. convection-coupled [ZK21]. convection-diffusion [CSS23, GO25, Jha21, MN21, RR22, SG23, SZ24a, SR24, VBDR24, WCY^{+}24, ZZ23a]. convection-diffusion-reaction [ANN25, SSY22, YH21a]. convection-dominated [CCY24, CGGY23, FZDS22, SZ24b, VR20, XFLZ22]. convective [ALOR25, BWZ22, BAB^{+}20, GS22, GCC20, JWZX25, KUK20, LZ20a, LZS23, OAA^{+}22, TSARR21, WWZQ25]. convective-like [GS22]. convectively [AKK^{+}25]. Convergence [AEZ24, BAKR24, BLLZ24, BC20b, Che24, Cos23, DD20, DW20, FS21a, GB23, HHNV23, KHH25, KSJ23, KTY24, LXZ23, LL24b, LTXZ24, LYL25, MCD23, NT21, NCRGVR21, PG25, PFCZ23, PK25, SMV^{+}21, WW25, ZHH21, Ahm24, BHPD23, BB23a, BM24a, Bot20, Bra20b, BFM^{+}25, CLQL25, DG22a, DSL24, EGP20, FP20, GDŁW23, HH21, HH22, HC22a, He23a, HIP23, HGW21, HZA22, HJXW23, KP20, KWLW24, Ku24, LSW22, LLD23, LX24a, LZ25a, MI20, PS22, SGZ^{+}23, SQL21, SJZ24, UHLZ24, UHL^{+}24, WM23, YSB^{+}24, YHL25, YY25, YPZ^{+}21, ZvNH25, ZZ24a, ZNC24, ZHZ23, ZZ25b]. convergent [MKS^{+}21, NR25b, PS20, PSDP22, SG23, SY25, TPA24, WLL23, ZC22]. convertible [LZ20d]. Convex [SS24, ABP25, Bar21, BKW22, Boł22, CK23, FFC^{+}25, Haj21, JS25, LS22b, LLJ23, NKPA^{+}25, Par22a, RGL22, SJZ22, WCLQ25, WL20, XLLA20, ZZW20]. Convexification [KLY25, LLN24b]. conveying [ANH^{+}21]. ConvLSTM [CYY^{+}21]. Convolution [ES20, HRS24]. Convolutional [Dai22, HK24, CKFR24]. cooling [DKV21, HHM21]. coordinate [CFT^{+}23, Fan23, ZQ25]. coordinates [JRu24, LLZ^{+}24, ZZY^{+}24]. copolymer [WZY22]. core [Fia21]. Coriolis [ANH^{+}21]. corner [CKBK24, KPJO22, ZL20b]. corners [FM25]. corotational [WQS25]. Corrected [LS22a, FZDS22]. correcting [Agg20]. correction [GJS20, LYAZ20, Loh21, RL23, SLUW^{+}21, WCC25, WLJF25, YCCW24, ZSF23, ZST22]. corrections [Zhe22]. Corrector [KMS20, LS20, LZ20d]. correlations [HSZ22]. corresponding [KPJO22, STO22]. Corrigendum [LZW21a]. corrosion [SHR25]. cost [CBŁP23, LM24, TZ20, Wol20, WLJF25]. costly [YYY^{+}20]. costs [TC25]. Couette [KARS23, RGS24]. counter [SB21]. counter-flow [SB21]. counterparty [ASSV20]. couple [JSSS24, NTP25, WQS25, YY23]. Coupled [CRDS20, ALOR25, AZY22, AAK24, BML22, BBK21, BDE^{+}24, BKT^{+}23, BBDM21, BCM25, BAB^{+}20, CLZ^{+}24, CTGR24, CG21, CLR23, CO20, CFT^{+}23, CCGRL21, DFS21, DK23, FKW22, FCFW21, FCF22, GM21a, GOV20, GIRB25, GV21, HWZQ23, HKM21, HBC^{+}21, HKL^{+}23, HW25, JJZ21, JWR^{+}23, Kap24, KLY25, KWLW24, KOMK20, LD22, LH23a, LZBC24, LXG^{+}24, LaZMY25, LLY25, MZGM25, MJRM20, MC25, MA22, NSRI24, OCCM21, PL24, RC25, RLSC23, SCD23, SXCB24, SSdA20, SMR24b, SR22, SZCW24, TjW24, TA22, VJ24, WYS^{+}20, WZY22, WR23, WTSZ24, WCZ24, WH20, WCH23, WW25, ZR22, ZL23a, ZSK23, ZSF23, ZZZ24, ZSY22, ZWW^{+}23, ZK21]. Coupling [GO23, JM23b, BMP21, BGG^{+}23, BPR22, CWTM20, DNP^{+}22, EBT24, JWW21a, JM23a, KGW^{+}21, LQLF23, LYAZ20, MMPR20, RL24, SGSD20, WCZD20, ZLF^{+}22, ZST22]. CPINNs [ZL23a]. CPU [DdSP21]. CPU-GPU [DdSP21]. crack [Alf24, KPO20, MCZL20, YA21]. cracks [BNB20, BM24b, GNB24, KP22d]. Crank [BDF24, GLLZ23, HLJ25, LHN24, LSW20, Li24b, LLZL24, LS25a, RWaZ23, SX25, WZ22, ZZH23, ZFQ24]. creation [Ngu20]. criteria [BC20b]. Critical [CS22, KJ22, SR20, SSdA20, ZJCG21]. cross [AZL20, GSS24, GO20, HRS24, HH21, HH22, LZL^{+}25, Lob21, SS20, ZJZ24, dCG20]. cross-anisotropic [AZL20]. cross-diffusion [GSS24, HH21, HH22, Lob21, dCG20]. cross-field [ZJZ24]. cross-ply [SS20]. cross-section [LZL^{+}25]. Crouzeix [CS22, DGN24, Sau23]. crucial [AWA^{+}22]. cryosurgery [TK20]. crystal [AAA25, AZY22, JW20, LYP^{+}22, LAX22, LJ24, PHL21, QZH23, YWT22, ZWT23, ZHZ23, ZWX24]. crystalline [HJTW25, LQYJ23]. crystals [GHLS25]. CT [YKNS23]. cubature [ASV20, ST23, WP25]. Cubic [Kap24, BNC^{+}23, DGN24, DS25, MH25, SX25, SRD25, WLD22, Yan25b, ZC23]. Cumulant [EFF21, APGK24, FGK20, PGK20, GCSF21]. CuO [CBK24]. CuO-EG [CBK24]. curl [CWW22, FHH23, JZ24b, LD24, PV20, TZX23, WCS24]. current [AAC23, BLRRS20, BD23, EP23, OSO^{+}23, Šeb24, ZS25]. currents [CRDS20, GJM22, GDA^{+}22, HKR22]. curvature [Ahm24, HJT25, LF24a, LSH21, MHHR20, SP20, XXYL23]. curvature-based [Ahm24]. curvature-driven [LSH21]. curve [BCK24]. Curved [LWW24, WKC21, Adj22, ABLS23, APS25, BPV24, CCGR24, DD24, EJWD23, JVBM23, JCP^{+}25, MMP21, WWZ^{+}24, YY23, aKST22]. curves [JICRMG^{+}20]. curvilinear [GQZ23, JRu24, OR22]. Cut [KNBR22, DAAEA25, KWM22, KBR20, VSG21]. cut-cell [DAAEA25, VSG21]. cVEM [ABV23]. cVEM-DG [ABV23]. cycle [SAR20]. cyclic [PRN^{+}23]. cycling [SCD23]. cylinder [AMPB^{+}22, AGA^{+}21, LML23, SCG23]. cylinders [AIG^{+}21, Yem24]. cylindrical [BPV24, HWZQ23, KHD24, LL22c, YLZL20, ZZY^{+}24, ZCA23]. d [WZ24, Ant21, BHD^{+}21, CSS23, HHL^{+}21, HJ23, HJ25, LHN24, LHSF24, Pta21, Vak25, ŻW21a]. dABCD [CSCY24]. Dagn [AHOB24]. dam [AS21, YY20b]. dam-break [AS21]. damage [Abd24, BT22, WXCL25]. damped [APPS25, AKO21, Ade20, KRO24, SX25, WZS24a, YJ24]. damping [JZS23, LNNP24, WLZZ22, XZ23, YW20, ZQM20, ZW25, ZS21]. Darcy [BN20, CRF^{+}20, DFSV22, GOV20, Gut20, HKM21, HX22, Lee24, LYAZ20, LD22, QHHW20, QWC^{+}23, QC24, SR22, TKK23, WDZ21, WCH23, WW25, XZ25, XHL24, ZL20a, ZHW20]. Data [CLY25a, FWLD22, LWZ22, WZY23, ADR24, ABC^{+}22, BBS22, CCGR24, IJS^{+}24, KMZ22, KLK^{+}22, LNT22, LK20, LWX21, PWMC23, PSDP22, Pło22, PT24, SXKL24, TH25, TVKN21, UR24, UHP23, WZT23, WLW^{+}22, Xu24, ZF22, ZWX24]. Data-driven [CLY25a, FWLD22, LWZ22, WZY23]. dataset [Par24b]. Date [Waz20]. DD [CDGM25]. DDFV [CQS25, IQS20]. DDM [ZSPW20, ZLW23]. de-icing [BTCV20]. deal.II [ABD^{+}21]. deblurring [Ahm24, BML22]. debonded [AP25]. Debye [BNSZ24]. decay [DZ21, HGW18]. decaying [BWG25, LZ20c]. December [Ano20-61, Ano21-53, Ano21-50, Ano22-59, Ano22-69, Ano23-72, Ano23-70, Ano25-67, Ano24-45, Ano24-46]. decompose [RGL22]. decomposing [Zdu21]. Decomposition [LHH22, AS21, AMV21, APS21, BDB21, BMVA23, BMST20, CBLZ24, CLY25a, CDPL24, DLC22, DZ24, FLS23, FZDS22, Gun22, KCD^{+}23, LLR25, LQLF23, LXZ23, LC23a, LCE23, PV22, PRN^{+}23, PNT^{+}23, PTNR24, SM22a, SR24, SNHH21, SWC24, SS24, TjW24, VLD^{+}24, WLQ22, XHM20, ZSGT24, ZS21, Zhe22, ZS23b, ZSL22, dCLW25]. decompositions [DT24, WJC^{+}25]. deconvolution [AL22, BTG24, BWG25]. Decoupled [CXX^{+}25, TKMK25, WZY22, CXJ25, CSS25, DMX24, GLF25, HJM23b, HYLH22, JZ24b, LZL24, NR25b, RC25, SPC^{+}23, WR23, XHL24, ZD22, ZS23a, ZWT23, ZW25, Zha22d]. Decoupling [Vab25, AZY22, CLLL25, HX22, KKNR21]. dedicated [VADG22]. deduced [TSW^{+}25]. Deduction [ZDT^{+}22]. Deep [OS24, YLSY24, ZWY23, ACLP24, BB23b, CAD24, CRL^{+}25, DZZ^{+}24, GHSR23, GZAR23, HLX22, KHY25, LLY25, MHKB25, OP24, PSI^{+}24, SGR^{+}23, TH25, WTSZ24, WZXN21]. deep-GA [PSI^{+}24]. deep-genetic [PSI^{+}24]. deep-lying [CRL^{+}25]. Defect [Agg20]. deferred [WCC25, YCCW24]. defined [HK22, OR22, SPZL21]. definite [MS20a]. deflagration [BTYM^{+}21]. deflagration-to-detonation [BTYM^{+}21]. deflection [AKA^{+}21, KAKK23, YX21]. Deformable [KJ22, CRPdT20, HYL^{+}25, MHHR20, SRA20]. deformation [BBDM21, DZZ^{+}24, HTPV25, Ma25a, SCD23, SRA20, Vah22]. deformations [CCX21, EBT24, GHRD20, SSK^{+}23, ZWW^{+}23]. deforming [LR22, OXY21]. degeneracy [BCT21, FRW22]. degenerate [AGLR21, ASGGRG23, ANZZ23, AA25a, ACV21, GGT24, IQS20, LL25a]. degradation [FGG^{+}20]. degraded [DH24]. DeGVR [PWC^{+}23]. Deicing [ZCZ^{+}21]. DEIM [CS20a]. delay [HZ21, SB24b, TSW^{+}25, WSS22, YGZ^{+}22, ZWS22]. delay-reaction [HZ21]. delayed [MKM24a, XZSY23, ZW20b]. delivery [AEF23, FJP22, PTM^{+}23, YD21, YD23]. demarcation [SA24]. dendrite [ZYQ^{+}24, ZWX24]. dendritic [DMŠ22, LAX22, RST^{+}22]. denoisers [Shi24]. Denoising [GW25, BML22, CBLZ24, FFC^{+}25, HK20a, LP22c, RGL21]. dense [GBK23, HLH^{+}20a]. Density [PWC^{+}23, BDE^{+}24, GPA21, GML^{+}20, HHNV23, KMZ22, LL22a, LDM25, NM23a, SHXZ24, XTZZ24, YWZ^{+}23, ZTW22, ZDW23]. Density-extrapolation [PWC^{+}23]. density-suppressed [HHNV23]. dependence [BPR22]. dependent [AS21, AS23, AL23, AN23, BDV23, BLF23, Bar22, BKR24, BDE^{+}24, Cha24, CLH22, CMR22, CG25, DLN24, DYZ22, GDA21, GHSR23, Gra21, HJM23a, HCZ25, HYH23, JRu24, JL23, KHM25, KAKK23, LXH24, LMLZ23, LS23, ŁWP^{+}23, MH22, MMN21, PG21, Pło22, PT24, PTNR24, QX23, QHHW20, RL23, RYAM21, Šeb24, SLW22, SJM24, SC21b, THF20, UR24, Wat22, Waz20, WLJF25, WS25, YWL21, Yua23, ZA23]. depletion [Ngu20]. deposition [SADW21]. depth [GdFFF25]. depth-integrated [GdFFF25]. DERivative [Nde25, AZL20, BWZ22, CRL^{+}25, Din21, FSW20, GHAU24, HLH24, HHR22, JSZ24, MR24, QWLQ25, SD25a, SYZ23, ZGG^{+}25]. derivatives [JBKW23, LLZ20, LSPRV24, NS24, Par22b, SHM21, XL24]. derive [Zha22d]. describing [CMS23, IQS20, LARGVR23]. Design [ABD^{+}21, DKV21, HWY23, HB25, YPC^{+}21, BC20b, CWLQ22, FLW20, KPO^{+}21, OKSS21, SMR^{+}24a, WWY25]. Designing [IBY^{+}24, GY25]. detailed [HEDT20]. details [LHC23]. detection [BDGS25, KPS23a, MZ23b, MR24, Par24a, Par24b, SHQL20]. deteriorated [SM22b, ZYW20]. Determination [AL23, QX23]. determine [LNNP24, Ngu20]. determined [LNT22]. determining [DCSAM25]. deterministic [BKK^{+}23]. detonation [BTYM^{+}21]. Developing [HLZ24, HLY22]. Development [HMŁW^{+}24, KTFI23, SK23, XXAD21, XL25b, Din21, KP22b, MJRM20]. developments [BBD^{+}21a]. device [LX24a, LY24, LCH23, SJ23]. devices [AHZ^{+}25, HWY23, SMR^{+}24a]. DG [ASGGRG23, ABV23, BFE^{+}21, HFG23, Kim21, Med23, Orl23, PSSW20, SK21, WCH23, Yua23, ZPK22]. DG-DGLM [Kim21, SK21]. DGFEM [HHR24, PS20]. DGLM [Kim21, SK21]. diabatic [SB24a]. diagonal [LC23b, XW24b]. diagonalizable [LDT24]. diagonalization [YW24]. Diagonalized [YM23]. diamond [LD23, QP24]. dielectric [BHD^{+}21]. diesel [SB24b]. diffeomorphic [HJZ23]. Difference [BWR21, Bel24, BGG^{+}20, NAE22, AS20, AKO21, ANZZ23, AAL24, AJ23, AAWJ20, BLLY24, BWZ23, BZ20, BF25, BDB21, BDGS25, CT22, CKFR24, CLL21, CLLG23, CCWS20, CTZ22, DL25, Der23, DAS20, EH23, FHM21, FS21b, GDA^{+}22, GLLZ23, GCMS21, GWYR22, HLH24, HC22a, HSY25, HLCW24a, HYW^{+}24, HGW21, HQL23, HLCW24b, JM20a, JM23b, JZ20, JHG^{+}21, JL23, JBKW23, KSS22, KK23a, KC22b, KTY23, KMP20, LHN24, LZ20b, LZ22, LC23a, LZY24a, LHSF24, Lob21, LCT22, MFZX23, MC24, MG25, MLS22, MKP23, PS25b, PK25, QD23, QQX21, QSL23, QL24, RWZG24, RST^{+}22, RKC25, SG23, SQL21, SZ24b, SLGF20, ST20b, SLYL22, SQGZ24, TRL23, TGARGT^{+}23, TGDMGT^{+}25, VL21, WZZ20, WS21a, WC21, WZ22, WGeL24, WZ24, Wan25, WLW^{+}25, XYS^{+}24, XZ23, XZ24, XLLA20, YQZT21, YL22, YWT22, YLL24, Yan24, ZTL22, ZAH23, ZZ25a, ZLL23, ZLJT20, ZL22b, ZX20]. difference [ZSPW20, ZZB20, ZZFL21, ZAPK23]. difference-based [XYS^{+}24]. difference-Legendre [Der23]. difference-proper [LC23a]. difference-spectral [BWZ23]. difference/finite [XLLA20]. difference/generalized [GCMS21]. difference/Kansa [YLL24]. difference/spectral [MC24, ZLJT20]. differences [KKS25, LY24]. differencing [FX22]. different [ADL^{+}21, CQG^{+}24, CS25, HHR^{+}20, LE24, LZY24b, QWC^{+}23, WSGT21, XFX^{+}24, YZ21]. differentiable [YSB^{+}24]. differential [Abd25, AAL24, ABFPA24, AMM23, ATBT20, BLLM25, BCCR20, BHD^{+}21, BHDV22, CDLW20, Che24, CKS22, CMM23, CP21, Dai22, Din21, FAM25, FS21b, GYL25, GKHU22, GAM24, HLL24, He23a, HCY^{+}24, HM24, HZW21, HHL22, JICRMG^{+}20, Kap24, KSS^{+}20, KSSS23, LL25b, LLZ^{+}25, LMLZ23, LSY^{+}24, LZS25, MCW21, MH25, MN23b, NF25, NSRI24, OP24, PKPD^{+}24, PSI^{+}24, SMDC23, SB25, SR23, SKG24, SZ21, TQX23, TFWX23, TBO20, Ull20, WCL20, WL22, WFX24, WYHQ24, Wan24, hWCsLjH21, XS25, YK24, YTT^{+}22, YLCC21, YWPS24, ZA23, ZX20, ZWS22, ZQS24]. differentiation [MM21, ZKTK21]. diffraction [GMS23]. diffuse [KP22b, LpWW^{+}25, SWWZ23, WMMH25, ZHHL23, ZGW23]. diffuse-interface [SWWZ23]. diffusers [DAAEA25]. Diffusion [RNN22, AD20, AM23, ACV21, AHOB24, ABL24, AN23, ANN25, AN21, BHJ22, BS23a, BIMS24, Ben25, BKM23, BMK21, BHL24, Bra20b, CY23, CCY24, CS20b, CLZW20, CLYH22, CMR22, CL24b, CSI23, CL21, CEL20, CDGM25, CSS23, CS25, DSS20, DPP21, Der23, DHVT20, DWS20, DAS20, DC22, DH24, EFFR22, FKW22, FC24, FSW20, FZS22, FZDS22, FXLS25, Gal21, GSS24, GP22b, GHH20, GM22, GO25, GCMS21, GCL25, HZ21, HLW23, HLMM20, HH21, HH22, HM20, HZR23, Hoa22, Hof20, HDWL24, HRvdZ20, HS20, HQL23, HW25, Jha21, JZ24a, JWZ22, JAHM23, JHG^{+}21, JWR^{+}23, JMN22, KSS22, KSD^{+}25, KYY^{+}20, KC22b, KKS25, Krz22, KSJ23, KGK25, KP23, KC23b, LPK22, LWX20, LLW20, Li20, LD21, LL22b, LCW22, LH23a, LXY23, LLX24, LDW^{+}24, LDM25, LY21, LYL^{+}22, LMZ23, LYZ23, LLXY23, LZS23, LZ23, LL24b, LL24a, LLXY25]. diffusion [LQF25, Lob21, LMMMP20, ŁSSP23, LP22c, LZ25c, LD23, LYL25, MFZX23, MZ23a, MC24, MXLY24, MJRM20, MMC21, MMGS25, MI20, MJZ25, MMN24, MMAK20, MMAK21, MRPK21, MN21, MWY23, MN23a, MRP20, ND21, NY22, NFS24, OTC24, OAH^{+}24, PXLY23, PQSM21, PH23, PS25b, PK25, QFL23, QL24, RNMÖX24, RL20, RGL21, RR22, RKC25, SM22a, SG23, SSY22, SM23, SMV^{+}21, SHM21, SGZ^{+}23, SSW20, SZS20, SQL21, SQ23, SW24a, SZ24a, SDG20, SCG23, SZ24b, SSD23, SN20b, SR24, SIAA25, SXLZ21, SHT^{+}24, SF25, Tan24, TH23, TZ20, TLJ24, TKMK25, Top22, TH21, TZ21, TTKV20, UHL^{+}24, VR20, VBDR24, VJH23, WS21a, WML21, WC21, WLX22, WW24, WDYZ24, WZ24, WJZ^{+}24, WT25, WWYL25, Wan25, WXH^{+}25, WYLL25, WCLQ25, WLW^{+}22, WKGH21, WZCS20, WSS22, WLQ22, WCY^{+}24, Xen20, XHY21, XLS22, XSZ^{+}23, XZSY23, XL25a, XLSF25, YH21a, YZT21, YGZ^{+}22, YCCW24]. diffusion [YL25, YEO21, YLZ^{+}24, YFK20, YFK23, YAA25, YYSL20, ZM20, ZJJ20, ZHYL21, ZFL22, ZDWR22, ZYTX22, ZL22a, ZF22, ZAH23, ZZ23a, ZWCS20, ZJCG21, ZZ25b, ZSPW20, ZZB20, ZW21b, ZXQQ22, ZLW23, ZZW23, ZHPW23, ZZFL21, ZW20b, dR25a, dCG20]. diffusion-based [LP22c]. diffusion-reaction [GP22b]. diffusion-wave [Bra20b, Der23, HZR23, JHG^{+}21, KP23, KC23b, LL24b, QL24, SM23, SCG23, ZAH23]. diffusion/Burgers [JWR^{+}23]. diffusions [ZLQ23]. diffusive [ÁCS22, BT22, HB24, HSWG21, KP22a, SB21, SSW20, SW24b, SHT^{+}24, TSW^{+}25, YZ21, YJL22, ZYD24]. diffusive-viscous [HSWG21, SW24b]. diffusivity [LL22b]. digital [KLRT21]. dilatable [BBS21]. dilute [SRP21]. dimension [ČPPŠ20, HLW24, JZ24b, Koc22, LLZL24, RTA20, XFS23, XWXF23, ZHPW23]. dimensional [AHZ^{+}25, Abd24, Abd25, AM23, AMN22, AAU25, BV23, BZ20, BBK24, BGLZ24, BABS20, CBŁP23, CML^{+}25, CO20, CLL21, CLLG23, CDLW20, CDM24, CXJ25, CC25, CZ20, CYQW21, CL21, CHL21, RFG^{+}21, DHW20, DZ22, Der23, DGQT20, DY25, DS25, DAS20, EH23, FAM25, FX22, Fur25, GHAU24, GHLS25, GWN21, GCMS21, HKH^{+}24, HZ24a, HK24, HM20, HZR23, HYW^{+}24, HGW21, HJXW23, HQL23, HWY25, HHG20, HA23, Ila20, IZ20, JKS25, JRu24, JJZ21, JCJ24, JWZX25, KSS22, KK23a, KG20a, Kim21, KLK^{+}22, KGK25, KP23, LNNP24, LZ20b, LGQZ21, LL25c, LSRL24, LFY^{+}20, LFY^{+}21, LRZ23, LLJ23, LLU24, LZ25c, MFZX23, MG25, MB24, MMC21, MR22, Mit22, MN23b, NML23, Oru22, PK25, PSI^{+}24, QQX21, QGQ22, QX23, QSL23, RZSL23, RG21, RLSC23, RR22, RKC25, SB24a, SD25b, SJ24, SSD23, SBY^{+}25, ST23, SU23, SXLZ21, SWC24, SQGZ24]. dimensional [TO20, TQX23, TH23, TVKN21, Ull20, UR24, Vis24, WFH^{+}20, WTSZ20, WC21, WXF22, WZ22, WTSZ24, WQ24, WJZ^{+}24, WWYL25, WJC^{+}25, WCCC25, WYLL25, WWZX25, WQS25, WZ25, Waz20, WTQ^{+}20, WCW20b, WWH23, XLLA20, XLS22, XLSF25, YQZT21, YLL24, YMTP25, YTT^{+}22, YFK20, YLWZ20, YWPS24, ZM20, ZCYL24, ZLPR20, ZJJ20, ZLT^{+}21, ZZL22, ZZL23, ZZS23, ZF24, ZLL20, ZZ20b, ZL23c, ZSW23, ZZY^{+}24, ZLZ25b, ZLJT20, ZZB20, ZSL22, ZWY23, ZZFL21]. dimensionally [LWFS24]. dimensions [BF25, GLL^{+}20b, HZ25, HMŁW^{+}24, KRO24, MMPR20, NW24b, QXCG20, RZC21, WYHQ24, WN24, ZW21b, aKST22]. Dirac [FLOQ21, HMW21]. Direct [AD20, AHP20, APGK24, BBS21, Bou24a, GHL25, JHS24, PWWJ23, AAA24, BS23a, BS23c, CHY21, CL22b, CPBF21, GF21, GYC25, LFGZ23, NE22, WWQ20, WWC^{+}24, XL22, Zdu21, DKH22]. direction [LWL^{+}20, LYL^{+}22, LMMMP20, OAH^{+}24, QXCG20, RHA^{+}24, WC21, YQZT21, ZX20]. Directional [BIMS24, ZWX24]. directions [MAT^{+}20, QP24]. Dirichlet [DDNS21, DH23, GP23, GS23, Gut20, LC21, MM21, MK20, SM23, SGSD20, WSO^{+}23, ZLQ23]. Dirichlet-to-Neumann [MK20]. disc [Ma25b]. discharge [YTT^{+}22]. discontinuities [SHT^{+}24]. discontinuity [KK25, WZT23]. discontinuity-capturing [KK25]. Discontinuous [BDG21, CO21, DP25, GHMD22, GRR20, KC21, LZBC24, NAS20, WLX22, WW24, ZQM20, ADB^{+}22, AMV21, AFV22, AMM23, AN21, BNSZ24, BGR23, BZ23, BKB22, BS23c, BLLZ24, CY23, CHY21, CLR23, CLRT24, CRM22, CM23, CHD24, CLYH22, CZLQ22, CXXZ23, CXJ25, CMR22, CCGR24, DW20, GYC25, Haj21, HWW21, HCY^{+}24, Hop23, HYJ20, HHW21, HA23, JMCC24, KSS22, KCW22, Kar24, KPS23b, KJN21, KGJN24, KBB^{+}21, KWS21, LMD22, LaZMY25, LTXZ24, LLXY25, LD25, LHTZ25, LdO21, ND21, Nie23, NERB^{+}22, PV22, PH25, RZC21, RZSL23, RWaZ23, RHR^{+}21, SGZ^{+}23, SSW20, SCdH21, SR22, Sto20, SW24b, WLX20, WaZWZ23, WJZ^{+}24, WJ25, Wie23, WKC21, XL22, YH21a, YZN22, YEO21, YYSL20, YH21b, ZH20, Zha22b, ZQY^{+}23, ZXXX23, ZLHC24, ZZW25, ZPC20, ZPC22, ZZY^{+}24, PDB22]. Discovery [WZT23, CLY25a, WZY23]. Discrete [BV21, LLZ^{+}23, SCCS21, WZCS20, AGLR21, AHOB24, AZY22, BNV22, BGR23, BFI^{+}23, CS20b, CL22a, CHZ23, CLW^{+}24, CG22, DN22, DFHT21, DD22, DMX24, FdOS20, GGRBRG20b, GGRBRG20a, GSRG21, HYH23, JH24, JMS^{+}24, KF22, KHH25, LZJC21, LS25b, LMLZ23, LL24b, LZY24b, LSZ24b, PT24, RDP20, SSW20, SLC^{+}24, TWH^{+}22, TPOT20, TVKN21, WDYZ24, WJC^{+}25, WMMH25, WCY^{+}24, XP22, YJL22, YYY^{+}20, YTH^{+}21, ZL20b, ZLQ23, ZZZ24, VJ24, XGL21]. discrete-ordinates [SLC^{+}24]. discrete-phase [ZL20b]. discrete-time [ZLQ23]. discretely [EKB20]. discretisation [BTG24, HS20]. discretisations [BBD21b, LMBB24]. discretised [MCK^{+}23]. discretization [AMV23, BHJ22, Bal21, BLO22a, BFVY23, BBDM21, BCCR20, Bra20b, CJ25, Cos23, DSS20, Gal21, HC21, HZJ^{+}22, HLCW24b, JH24, Kam25, LSV21, LQLF23, LWX20, LQ20, LY25, LP22a, MZGM25, MN21, MWY23, MS23, PS21, Pon20, PZZ23, SQC23, SB25, SQ23, TCW23, TLJ24, WT25, Wie23, WLQ22, YS23, ZHW20, dlRGR20]. discretizations [AD21, BHPS24, Bar21, BHZ20, BBK21, BBD^{+}24, BZY20, BMvdZ21, CPS21, DD21, DS20, EFK^{+}23, Haj21, HLPW22, KCD^{+}23, LD22, NERB^{+}22, Orl23, PS22, Pup25, SW21, TZ21, ZSGT24, dlRRG21]. discretized [ADR24, AMV24, BHS20, BM20, GHLS25, HSR25, KMN25, NH24, SXKL24, VLD^{+}24]. disease [ASF^{+}25]. diseases [THF20]. disk [IBY^{+}24, MM21, MLHS23, SN20a]. disparate [MLJ^{+}20]. disperse [XGL21]. dispersed [MR22, Mit22]. Dispersion [KG20a, Ale23, HNLPP20, JFS14, JFS20, SRP21, TBS25, WCH20, WSGT21]. dispersive [BWR21, KJ24, NS20, ZAPK23]. displacement [CXXZ23, JCP^{+}25, KS25, SR21, WCH20, YZN22, ZR23, ZH20]. disrupted [TGL21]. dissimilar [MGA22]. Dissipation [WFH^{+}20, AIG^{+}21, CCF^{+}21, CLQL25, DZ22, LS25a, LCT22, PM23, SX25, WSCP24, ZWMF24]. Dissipation-preserving [WFH^{+}20, SX25]. Dissipative [Bal21, PZL23, ABV23, HLJ25, Lee20, ST22]. dissolution [SGJJ20]. distance [QWLQ25, WDGS23]. distances [MJR20]. distorted [WLQ22, ZSY22]. distortion [EAM20]. Distributed [GHK25, BHS20, BM24b, BJyST24, CRL^{+}25, CL21, Der23, GLS25, GLL^{+}20b, GF21, GCMS21, IJS^{+}24, JWZ22, JHG^{+}21, Kar24, KSS^{+}20, KSSS23, LLSY24, LLSY25, Nag24, PWMC23, QXCG20, SKG24, SIAA25, TQX23, WLX22, YCZ^{+}21, YL25, YFK23, ZM20, ZFL22, ZLJT20, ZWS22]. distributed-order [CRL^{+}25, CL21, Der23, GLL^{+}20b, GCMS21, JWZ22, JHG^{+}21, QXCG20, SKG24, SIAA25, TQX23, YCZ^{+}21, YL25, ZM20, ZFL22, ZLJT20]. distribution [BBT25, CBI^{+}23, DKH22, HRS20, LXJ^{+}21, LLZ^{+}23, XLZ^{+}21]. distributional [BLO22b]. distributions [SS21, SS24]. disturbance [WWLG24, ZWW^{+}23]. div [CWW22, HJM23a, HJ24, JZS23, LXRQ22, LD24, LH23b, PV20, QHHW20, WWZX25, WCH23]. div-curl [CWW22, LD24]. Divergence [RHA^{+}25, BHK^{+}24, DMX24, EKB20, LLX24, TLJ24]. divergence-conforming [BHK^{+}24]. divergence-free [DMX24, EKB20, LLX24]. divergent [MKS^{+}21]. DLN [CQG^{+}24]. DM [JYHS23]. DM-TS [JYHS23]. DMD [AMS24]. DMLPG [AD20, Ila20]. DMRT [SADW21]. DMRT-LBM [SADW21]. document [DH24, WH25]. does [Jog25]. DOFs [DWAM25]. Domain [AMV21, BWR21, SNHH21, Zhe22, ADD22, ADGG25, APS21, BMVA23, BMST20, Bot20, CFP22, CCX21, CLY25a, DLC22, DZ24, FLS23, FZDS22, GMS23, GHL25, HB25, JZXR24, KP22c, KPK25, KWK24, KCD^{+}23, KFD^{+}20, LLR25, Leg21, LXZ23, LZA23, LCE23, LLJ23, LYLH25, nMjWlS24, MGA22, NW24a, PG21, PWMC23, PNT^{+}23, PTNR24, PZZ23, RLVR21, RLF21, SR24, SGSD20, TjW24, VLD^{+}24, WCLQ25, WLQ22, XHM20, XL25b, Zdu21, ZYC^{+}24, ZSGT24, ZS21, ZCA23, ZS23b, ZSL22, ZY21b, ZAPK23, dCLW25]. domain-decomposition [PNT^{+}23]. domain-free-discretization [PZZ23]. domains [AD20, ABP25, APS25, BDV23, BRR25, BL20, CK23, DK23, FMS22, FAM25, FM25, FH21, GP20, GCMS21, HK22, JVBM23, JKS25, KBR20, KTY23, LLW20, LYRW21, LSRL24, MSHB22, SW24a, XHY21, XLLA20, XLCP20, YK20, YW24, ZLT^{+}21]. dominated [BLLZ24, CCY24, CGGY23, FZDS22, FXLS25, KSJ23, ŁSSP23, QFL23, SZ24b, VR20, XFLZ22]. Double [RYAM21, ÁCS22, CBLZ24, DFHT21, EJWD23, HWZQ23, KC23a, LL24a, NSMM21, SXCB24, WSS22, YZ21, ZZZC24]. double-diffusive [ÁCS22, YZ21]. double-lid-driven [ZZZC24]. doubly [AA25a, JCP^{+}25, MMP21, WTSZ24, WS20b]. DPG [BDMX21, BHPD23, DZ20, DGK20, DS22, FH21, Füh21, FHH23, FF24, HD21, LD21, LD24, MN21, PD21, RH21, RMBC24]. DPG-star [DGK20]. DPIM [CCGF25]. DPSS [CSS20]. drift [JAHM23, dR25a]. drift-diffusion [JAHM23, dR25a]. driven [ACXI24, AMK25, ACV21, AA20, BP20, BDGS25, CM24, CBLZ24, CLY25a, EGP20, FWLD22, HBC^{+}21, JWW21a, KC23a, LSH21, LWZ22, LY25, MW20, MKAM20, PCC21, WX22, WZY23, WZT23, WKGH21, YHL25, ZZZC24, ZCZ^{+}21]. DRM [Sch21]. drones [JJOG20]. drop [NM23a]. droplet [ADL^{+}21, HKL^{+}23, Vah22, WZXN21]. droplets [ENF24]. Drug [AEF20, AEF23, BFRA21, FdOS20, FJP22, FGG^{+}20, PTM^{+}23, SAVHB22, YD21, YD23, YRC25]. dry [XDL^{+}25]. DSA [PS20]. DSSR [NHCJ22]. DTC [GY25]. DtN [LLWX25, WDYZ24]. DtN-type [WDYZ24]. Dual [OTC24, ABC^{+}21, CLRT24, GDA21, GAM22, HXZ^{+}22, JZPG21, Ku24, LXRQ22, LFGZ23, LZBC24, LLY25, LYRW21, QQY^{+}23, WW20a, WZ21, ZYC^{+}24]. Dual-grid [OTC24]. dual-level [LFGZ23]. dual-mixed [GDA21, GAM22]. dual-phase-lag [LYRW21]. dual-porosity [QQY^{+}23]. dual-porosity-Brinkman [LZBC24]. dual-porosity-Stokes [CLRT24, LXRQ22, LLY25, WZ21]. Duality [PR20]. Duality-based [PR20]. duct [AA20, BPV24, HUT25, LHQW25, MKS^{+}21]. due [Alf24, KK23a]. Dufour [JSZ24]. DuMu [KGW^{+}21]. Dune [BBD^{+}21a]. Dupuit [Bou24b]. during [AOL22, BGG^{+}23, BNC^{+}23, BPID20, ISA24, PTM^{+}23, TK20]. dusty [CBK24]. Dynamic [LLMC25, Alt23, BT22, Cao21, CRPdT20, CDPL24, HLWL25, Hop23, JMCC24, Li21, LXG^{+}24, MRAC24, Nag24, SMV^{+}21, SYHN24, SZCW24, WYS^{+}20, ZDW^{+}20]. dynamic-Immersed [CRPdT20]. Dynamical [MRP20, CHJR21, LHC23, Xu24]. dynamically [LaZMY25]. Dynamics [ENF24, ADB^{+}22, ANH^{+}21, AMS24, ASF^{+}25, ERG^{+}24, GV20, HKL^{+}23, JCJ24, JM20b, Kar24, LLW20, LWZ22, PTNR24, RNMÖX24, SPC^{+}23, TSW^{+}25, WWG^{+}25, WLD22, YZJ^{+}22, ŻW21a, ZDLC22, ZZZ24, ZZW25, ZJCG21].

Easy [LST22]. economic [AGM^{+}22, dCG20]. Eddy [BD23, SGF^{+}21, BLRRS20, EP23, GJM22, HKR22, MKS^{+}21, PZZ23, Šeb24, SQB22, ZS25]. Eddy-current [BD23]. edge [AAG22, ASV20, GK21, JP25, LMZ23, PND^{+}24, RGL21, SHQL20, WCS24, ZDLL22, ZLD^{+}23]. edge-centered [LMZ23]. edge-oriented [JP25]. edge/node [PND^{+}24]. edge/node-based [PND^{+}24]. Edged [YCL20]. Edged-based [YCL20]. edges [GQZ23, KTY24, TSARR21, TKK22, WWZ^{+}24]. Editorial [Lop22, Ano20a, Ano20b, Ano20-27, Ano20-28, Ano20-29, Ano20-30, Ano20-31, Ano20-32, Ano20-33, Ano20-34, Ano20-35, Ano20-36, Ano20-37, Ano20-38, Ano20-39, Ano20-40, Ano20-41, Ano20-42, Ano20-43, Ano20-44, Ano20-45, Ano20-46, Ano20-47, Ano20-48, Ano20c, Ano20d, Ano20e, Ano20f, Ano20g, Ano20h, Ano20i, Ano20j, Ano20k, Ano20l, Ano20m, Ano20n, Ano20o, Ano20p, Ano20q, Ano20r, Ano20s, Ano20t, Ano20u, Ano20v, Ano20w, Ano20x, Ano20y, Ano20z, Ano21a, Ano21b, Ano21-27, Ano21-28, Ano21-29, Ano21-30, Ano21-31, Ano21-32, Ano21-33, Ano21-34, Ano21-35, Ano21-36, Ano21-37, Ano21-38, Ano21-39, Ano21-40, Ano21-41, Ano21-42, Ano21-43, Ano21-44, Ano21-45, Ano21-46, Ano21-47, Ano21-48, Ano21c, Ano21d, Ano21e, Ano21f, Ano21g, Ano21h, Ano21i, Ano21j, Ano21k, Ano21l, Ano21m, Ano21n, Ano21o]. Editorial [Ano21p, Ano21q, Ano21r, Ano21s, Ano21t, Ano21u, Ano21v, Ano21w, Ano21x, Ano21y, Ano21z, Ano22a, Ano22b, Ano22-27, Ano22-28, Ano22-29, Ano22-30, Ano22-31, Ano22-32, Ano22-33, Ano22-34, Ano22-35, Ano22-36, Ano22-37, Ano22-38, Ano22-39, Ano22-40, Ano22-41, Ano22-42, Ano22-43, Ano22-44, Ano22-45, Ano22-46, Ano22-47, Ano22-48, Ano22c, Ano22d, Ano22e, Ano22f, Ano22g, Ano22h, Ano22i, Ano22j, Ano22k, Ano22l, Ano22m, Ano22n, Ano22o, Ano22p, Ano22q, Ano22r, Ano22s, Ano22t, Ano22u, Ano22v, Ano22w, Ano22x, Ano22y, Ano22z, Ano23a, Ano23b, Ano23-27, Ano23-28, Ano23-29, Ano23-30, Ano23-31, Ano23-32, Ano23-33, Ano23-34, Ano23-35, Ano23-36, Ano23-37, Ano23-38, Ano23-39, Ano23-40, Ano23-41, Ano23-42, Ano23-43, Ano23-44, Ano23-45, Ano23-46, Ano23-47, Ano23-48, Ano23c, Ano23d, Ano23e]. Editorial [Ano23f, Ano23g, Ano23h, Ano23i, Ano23j, Ano23k, Ano23l, Ano23m, Ano23n, Ano23o, Ano23p, Ano23q, Ano23r, Ano23s, Ano23t, Ano23u, Ano23v, Ano23w, Ano23x, Ano23y, Ano23z, Ano24a, Ano24b, Ano24-27, Ano24-28, Ano24-29, Ano24-30, Ano24-31, Ano24-32, Ano24-33, Ano24-34, Ano24-35, Ano24-36, Ano24-37, Ano24-38, Ano24-39, Ano24-40, Ano24-41, Ano24-42, Ano24-43, Ano24-44, Ano24c, Ano24d, Ano24e, Ano24f, Ano24g, Ano24h, Ano24i, Ano24j, Ano24k, Ano24l, Ano24m, Ano24n, Ano24o, Ano24p, Ano24q, Ano24r, Ano24s, Ano24t, Ano24u, Ano24v, Ano24w, Ano24x, Ano24y, Ano24z, Ano25a, Ano25b, Ano25-27, Ano25-28, Ano25-29, Ano25-30, Ano25-31, Ano25-32, Ano25-33, Ano25-34, Ano25-35, Ano25-36, Ano25-37, Ano25-38, Ano25-39, Ano25-40, Ano25-41, Ano25-42, Ano25-43, Ano25-44, Ano25-45]. Editorial [Ano25-46, Ano25c, Ano25d, Ano25e, Ano25f, Ano25g, Ano25h, Ano25i, Ano25j, Ano25k, Ano25l, Ano25m, Ano25n, Ano25o, Ano25p, Ano25q, Ano25r, Ano25s, Ano25t, Ano25u, Ano25v, Ano25w, Ano25x, Ano25y, Ano25z]. Effect [OCCM21, PGR23, SBH^{+}21, AAC23, AAG22, BPV24, DKH22, JWW^{+}21b, LXG^{+}24, SADW21, SGAC^{+}20, SCD25, SP20, Vah22, WLZZ22, YLZ20a, YWZ21]. Effective [JCP^{+}25, MR22, Mit22, HK20a, LLN24a, Par24b, SB21, SHD23, WH21, WL25, XXYL23]. effectively [PP24]. Effects [KRESA21, MHHR20, SRA20, ZDLL22, ZLD^{+}23, HKL^{+}23, JSZ24, LS25b, NR25b, WZCS20, XXYL23, YD21, YD23]. efficiency [GDA^{+}22, KCC20, TL22]. Efficient [AJ23, BHJ22, BHK^{+}24, BM20, DE22, DBB23, EP23, FCFW21, HKS21, HC22a, JMM20, KRO24, Kam25, LF24a, LCH23, LP22c, MC24, Ma25b, NNH^{+}25, NCCD24, PCC21, VJ24, XFS23, XZ23, YW24, ZZFL21, Abd25, Ahm24, ABP25, BLLM25, BV23, BMP^{+}24, CLLL25, CDM24, CXX^{+}25, CS25, CWX22, DLC22, DY25, DWY^{+}25, DNP^{+}22, GP22b, GLLZ23, HJTW25, HKR22, HH23, JVBM23, JSSS24, JWR^{+}23, JZ24b, JP25, KTFI23, KR24, LYAZ20, LZ20b, LAX22, LC23a, LaZZ24, LaZMY25, LS20, LS23, LMST20, LST22, LAB23, MJR20, MWRA22, Ngo25, Rav20, RTA20, RGL21, SMŁWL25, SMS23a, SPZL21, SIAA25, TT25, TK20, WCL20, WXF22, WYP23, WYT23, XL22, XWXF23, XLS22, Xu24, YQZT21, YTK21, YMDW25, YA21, YHF24, ZWT23, ZSPW20, ZXQQ22, ZLW23]. Efficiently [DZ22]. efflux [JJ23]. EG [CBK24]. eigen [WJC^{+}25]. eigen-decompositions [WJC^{+}25]. Eigenanalysis [NERB^{+}22]. eigenproblem [Med23]. Eigenproblems [Hak24a]. Eigensystem [SBY^{+}25]. eigenvalue [AMV23, AMV25, AZ24, APS25, BKW22, BB21, BBG23, BZY20, CGM^{+}20, CCD25, ERG^{+}24, Hak24b, JZ24b, KMZ22, KBW23, NW24a, PN24, PP24, QV22, STO22, TCA22a, Tür20, WZA25, XXLS23, YDY^{+}23, YM23, ZY21a, ZH23a, ZLHC24, ZGG^{+}25]. eigenvalues [Ant21, DV22, FM25, FWBY25, LQZ25, RTA20, WBY23]. eighth [BZ20]. eighth-order [BZ20]. eikonal [HMF24, LQ20]. Einstein [GM21a, WTQ^{+}20, XHM20]. Elastic [NZ24, AA21, ADD22, DT24, EJWD23, Fur25, Hag24, HLQ24, HTPV24, HMX21, LLMC25, LLWX25, RL24, WF24, XLYY25, YLW^{+}22, YLZL20, YLFL20, YYY^{+}22, ZWZD24, ZH23a, ZNC24]. elastically [LLZ^{+}25]. elasticity [AHLZ24, ARV23, BHK^{+}24, BB21, CXXZ23, CS23a, CPS21, CLL^{+}24, DGQT20, DFGS24, DSL24, GENRB21, HWL^{+}20, HLW21, HWWZ24, LXZZ24, Med23, Nag24, NK25, RR23, SDJ^{+}22, TLZF20, Vis24, WNF^{+}25, WCW20b, Yu23, ZR23, Zha22c, Zha24]. elasticity-poroelasticity [BHK^{+}24]. elasto [EMSA25, KSB25, PRN^{+}23]. elasto-plasticity [KSB25, PRN^{+}23]. elasto-static [EMSA25]. elastodynamic [ABC^{+}21, CHZ23, KCC20, SQGZ24, WF25]. elastodynamics [CLZN24, FMS22]. elastoplastic [BZ21, GM21b]. elastoplasticity [MC25]. elastostatic [Oru25]. electric [CRDS20, FSH^{+}23, KTY24, XL25b, YLZL20, ZS23a]. electric/magnetic [YLZL20]. electrical [CEB24, KLRT21, WWY25]. electrically [Du24]. Electricity [FGK^{+}21]. Electrified [AAS^{+}20]. electro [AA21, ASH^{+}21, EJWD23, HTPV24, LaZZ24, LLMC25, WLZZ22, YLZL20, YLFL20]. electro-hydrodynamic [LaZZ24]. electro-hydrodynamics [ASH^{+}21]. electrochemical [CZC^{+}25]. electrochemical-thermal [CZC^{+}25]. electrode [AHZ^{+}25]. electrodes [PCM20]. electrolysis [SRP21]. electrolytes [HSZ22]. electromagnetic [AAA24, BHDV22, FLW20, GPRH^{+}24, HGD25, LSV21, TPA24]. electromagnetics [PD21]. Electromechanical [MKM^{+}24b]. electron [PP24, YCGL25]. electronegativity [LXJ^{+}21]. electroosmotic [AA20]. electroporation [SAVHB22]. electrostatic [HSZ22]. Element [BL20, BDG21, BBS21, BLMV24, CO21, DFS21, HNLPP20, HMX21, KC21, KCC20, NCRGVR21, SBMS22, VJ24, XGL21, dVDR20, ADKH20, AGLR21, APPS25, AN20, AMN22, AMV23, AMV25, Adj22, Adj25, ADGG25, ACLP22, ADFP23, ÁCS22, ACPT25, ALOR25, AZY22, AZHY22, ACRBT23, AAB^{+}21, AJK23, AMV20, APS25, ABD^{+}21, AGPP22, AN23, ANN25, AAM^{+}24, BDV23, BIMS24, BLF23, BGR23, BS23b, BM20, BHZ20, BW21, BJG23, BB23a, BM24a, BBM25, BMPP22, BER21, BS24, BNC^{+}23, BZY20, BBS22, BCF^{+}23, BMvdZ21, Bri24, CWZ21, CAD24, CL24a, CCW22, CM21, CCGI23, CGM^{+}20, CM23, CLZW20, CLYH22, CWC22, CHL22, CLH22, CLZ23b, CHZ23, CS23a, CXF24, CLZN24, CCD25, CLQL25, CC25, CKBK24, CCWC23, CSS25, CJ25, CRF^{+}20, DLN24, DS20, DFSV22, DV22, DG22b, DD20, DGN24, DE21, DGQT20, DDG23, DK23]. element [DK24, DWS20, DS25, DTV20, DH23, DZ24, Du24, DBB23, EAM20, ERG^{+}24, EBT24, ELS24, FSS^{+}20, FHM22, Fia21, FCFW21, FP20, GP25, GLS25, GLL^{+}20b, GQ23, GNZ24, GP23, GOV20, GORBS21, GIRB25, GPY22, GDA21, GNB24, GENRB21, GCC20, GSW21, GHL24, GQZ23, GS23, GGT24, Gun22, GB21, GB23, GXFT23, GCL25, GHK25, GO23, HC21, HKJ23, HRS24, Hak24b, HZ21, HZJ^{+}22, HJM23b, HDM24, HLL24, HH21, HH22, HLX20, HQLJ20, HSZ22, HWY23, HZ24b, HLQ24, HB25, HZ25, HOY24, HV21, HBC^{+}21, HYLH22, HX22, HHL^{+}21, HS20, HXZ^{+}22, HLY22, HAC23, HWC24, HYL^{+}25, HWWZ24, JVBM23, JZPG21, JM23b, JMM20, JCP^{+}25, JZS23, Jog25, KHD24, KBR20, KJ21, Kei21, KSD^{+}25, KMN25, Ku24, KSJ23, KC23a, KD23, KTY24, KS25, LLSY24, LLSY25, LS22a, LP21, LP23, LLD23, LL21b, LWSZ21, LZ21, LZJC21]. element [LD22, LL22a, LZA23, LWW24, LZL24, LaZZ24, LLZL24, Li24a, LHF24, LR23, LZ25b, LSF25, LC23b, LX24b, LFY^{+}20, LD20, LS21, LY21, LN21, LSWY21, LSGN23, LGZC23, LLXZ23, LS23, LZ23, LLW^{+}23b, LLLL24, LHSF24, LGL24, LSY^{+}24, LXZZ24, LSZ24a, LXL25, LJY^{+}25, LYLH25, LLWX25, MCD23, MGD23a, MCD24, MCW25, MKM24a, MMN21, MSHB22, Man24, MI20, Mas23, MA22, MKM^{+}24b, MT25, MZ23b, MLLL24, MN23a, MKP23, MS22, MS20b, MCV24, NR25a, NC20, Ngo25, NNNT22, NP24, NTP25, NY22, NFS24, NW24b, NML23, OAD24, OXY21, OS24, OR22, OSZ22, PS22, PG21, PZL23, PND^{+}24, QXCG20, QS20, Qiu21, QQY^{+}23, RS21, RL23, RC25, RR23, SM22a, SD25b, SR23, SPC^{+}23, SL20, SL21, SJ21, SJ23, SL24, SLT24, SQ24, SWW22, SK25, SGR^{+}23, SR21, SR22, Sto20, SS21, SWZC23, TLZF20, TA22, TCW23, Top22, TKK22, TKK23, VBDR24, VJ22]. element [WYH20, WW20b, WW20a, WWW21, WXZW21, WML21, WDZ21, WCWZ22, WWL22, WL22, WWS23, WGC23, WZD24, WWC^{+}24, WCS24, WZS24a, WSCP24, Wan24, WLC24, WYW^{+}24, WWY25, WCCC25, WYLL25, WZA25, WWZX25, WQS25, WZ25, WLW^{+}22, WWCJ23, WHC24, WH20, WHH22, WN24, WLQ22, WW25, XLS24, XSL24a, XXLS23, XLLA20, XP22, YRC25, YPC^{+}21, YA23, YL20a, YLW23, YJ24, YWL21, YWZ21, YZZ20, Yu23, YW24, YCGL25, ZLYL20, ZWOM21, ZR23, ZD21, ZH20, ZL20a, ZHYL21, Zha22b, ZD22, Zha22c, ZF22, ZS23a, ZH23b, ZQY^{+}23, ZZH23, ZZ23a, ZSF23, ZF24, ZZ24b, ZGC24, ZZ24a, ZZZ25, ZL25, ZXZF20, ZSW23, ZZZD24, ZLZ^{+}25a, ZZ25b, ZS21, ZS23b, ZCW24, ZW20a, ZW21b, ZXQQ22, ZST22, ZZW23, ZH23c, ZZS24, ZS24, ZQ25, ZHW20, ZX22, CLL^{+}24, DD24, KNBR22, PCL25]. element-free [HBC^{+}21, LL21b, MCD23, MCD24, YL20a, ZL20a]. element-meshless [EBT24]. element-projection-proximal [LC23b]. element/level [NC20]. elementary [IK24]. elements [AS21, ASV20, BNV22, BHS20, BHS25, BGLZ24, CS22, CHL21, DMRV24, DFGS24, DST25, EJR24, Fan23, FS21a, FWBY25, FK21, GK20, GLC24, HMFM25, Hak20, Hak24a, HWL^{+}20, HSM^{+}23, HHL^{+}21, JMM20, JP25, KK21, Lee21, LWW24, LLZ^{+}24, LLMC25, Liu25a, NT21, PH25, PLZZ25, Pta21, RL24, Sau23, SNHZ25, Vis24, WP25, WXH^{+}25, YHZ22, YKNS23, YZW^{+}22, ZMS21, ZNMS23, ZK21, dVDMM22]. elements-based [AS21]. Eliminating [HRvdZ20]. elimination [ST23, ZLJ^{+}22]. Elliptic [LLR25, AN20, Adj25, AA25a, BHS20, BHS25, BBI^{+}22, BMVA23, BB23b, BX20, BJyST24, CAD24, CHY21, CCW22, CD20, CKFR24, CGM^{+}20, CLZ23b, CSCY24, CCD25, CCWC23, CV20, CSS23, CS25, CNT23, DS20, DE21, EJR24, FHM22, GLS25, GP22a, GO25, GB21, GB23, GXFT23, GYC25, HLX20, HZ25, HHL22, JK22, KC22a, KSB25, KJN21, KD23, LLSY24, LNF24, LNT22, LG23, LQW22, LC23b, LS21, LC21, LC22, LTXZ24, LSZ24a, Liu25a, Ma25b, Men22, MD21, MN23b, NT21, NF25, PH25, PK23, QSL23, QWLQ25, QN20, STM21, SSD23, SSdA20, TZX23, TKK22, TBO20, WW20b, WW20a, WXZW21, WGC23, WCCC25, WZA25, WL20, XZT23, XZ24, YLY20, YEO21, ZFZ20, ZC22, Zha23, ZCW24, ZW20a, ZMZW24]. elliptical [ZH23b]. Embedded [ZBCS24, CXXZ23, CCGR24, FM25, Koc22, OAA^{+}22, WCW20b]. emergence [dCG20]. emitting [MBSEA24, PH23]. employing [ABC^{+}22, SWZC23, SQGZ24]. emulation [OS24]. enabled [RHB25]. enabling [PKPD^{+}24]. enclosed [HUT25, MSHB22]. enclosure [AA25b, DK24]. enclosures [BP20, ZSL22]. endocrinology [YKNS23]. energetic [ADD22]. Energy [GZ23, HZW21, JS25, YLFZ25, ZWMF24, Alf24, ANH^{+}21, AJ23, AIG^{+}21, BEKV21, CFG^{+}25, CBI^{+}23, CGH24, CLQL25, DG22a, DZ22, DK23, DW20, FL22, FZH22, GS22, GLS25, GGRBRG20b, GFXZ25, HKJ23, HLJ25, JMS^{+}24, JWL20, JW20, Jog25, KTY23, LLSY24, Lee20, LL25a, LWSZ21, LaZZ24, LF24b, LJ24, LS25a, MZGM25, PHL21, PLZZ25, PL25, PP24, QH22, QZH23, ySRC22, SZ22b, SQ24, SX25, SXKL24, ST22, SY25, SWZ23, SBS24, Tan25, TRL23, VJ22, WZY22, WYP23, WaZWZ23, WGeL24, WSCP24, WLC24, WHH22, XKXL22, XL25b, YWL23, YKK25, YCGL25, ZL20b, ZO21, ZD22, ZY23, ZS23a, ZDW23, ZZW25, Zha22d, ZHZ23, ZLZ25b]. energy-conservative [YWL23]. energy-conserving [WLC24]. energy-consistency [Alf24]. energy-dissipation-preserving [DZ22]. energy-dissipative [HLJ25, Lee20]. Energy-preserving [YLFZ25, AJ23, DG22a, FZH22, GFXZ25, LF24b, MZGM25, SZ22b, SQ24, TRL23, WGeL24]. Energy-production-rate [HZW21]. Energy-stable [JS25, CFG^{+}25, CBI^{+}23, GS22, GGRBRG20b, LWSZ21, LaZZ24, PLZZ25, QH22, QZH23, SY25, WaZWZ23, ZS23a, ZZW25, Zha22d, ZLZ25b]. Enforcement [KC21]. Enforcing [HK22]. ENgineering [AGNP24, JKL20, DG22b].

Enhanced [FCF22, HANA24, LLN24a, XTZZ24, AEF23, CCG^{+}23, CEB24, EMSA25, FdOS20, FJP22, KW20, LF24a, LL25b, LHH22, LYLH25, NP24, WYP23, ZD21]. Enhancement [WCZ24, MAT^{+}20, YGZ^{+}22]. Enhancing [eLLAN25, SB25]. ENO [BD20]. Enriched [DGQT20, LG23, LM24, RHA21, WF24, Yem24, YHLA22, DD24]. Enrichment [GNB24, BIMS24, HWL^{+}20, SWZC23, WHH22]. enrichments [DGN24, HSM^{+}23]. ensemble [BLLM25]. ensemble-based [BLLM25]. ensembles [MWRA22]. enstrophy [LMBB24]. enthalpy [CFTC25, ZK21]. entropic [SCD25, YOF20]. Entropy [BKB22, TSARR21, AMPB^{+}22, sCpLYhG23, CSI21, RDP20, WKC21]. entropy-stable [CSI21, RDP20]. entry [WKGH21]. envelope [HGD25]. environment [DYZ22, MJR20]. environmental [CSB^{+}21, YTH^{+}21, YTA23]. environments [LYZ23, XFX^{+}24, YT22]. epidemic [BDS^{+}22, XSL24b]. epileptic [MRP20]. EPSD [HJP23]. Epstein [AHOB24]. equal [HHW21, LHF24, PS22]. equal-order [HHW21, LHF24, PS22]. equalization [LXJ^{+}21]. equation [ADKH20, AAC23, APPS25, Ade20, AFS24, ANZZ23, ADGG25, AMK25, ACV21, ACLP22, ACLP24, ADFP23, Alt23, ALOR25, AZHY22, ADG23, ABV23, AMV24, ANN25, BS23a, BZ20, BF25, BHZ20, BLO22a, BTCV20, BDL20, BDP23, BBM25, BZ23, BHL24, BLO22b, Bou24a, BWMW23, CFP22, CCF^{+}21, CS20b, CL22a, CLH22, CMY23, Che24, CGH24, CGK24, CLXZ25, CLY25a, CZK^{+}25, CLQL25, CC25, CL20, CWX22, CDG^{+}23, RFG^{+}21, DDNS21, DZ22, DLW24, DL25, DHVT20, DS22, DZ21, DEW25, DS25, DW20, DR24, FM25, FBK^{+}24, FLMS24, FRW22, FHZ22, FX22, FZH22, GvV22, GLL^{+}20b, GHAU24, GGH25, GKP25, GHL24, GCMS21, GCL25, HMF24, HSWG21, HCHM25, HGW18, HWW21, HC22a, HSZ22, HZ24a, HNLPP20, HDWL24, HHR22, HYH23, HLJ25, HRvdZ20, HHR24, HGW21, HJXW23, HS20, HAC23, HJ23, HW25, HJ25, JZ20, JS25, JJZ21, JWZ22, JCJ24, JXK20]. equation [KRO24, Kap24, KSD^{+}25, KHM25, KK23b, KLK^{+}24, KMN25, KMS20, KPJ22, KP23, KC23b, LSW22, Lee20, LKC^{+}24, LL25a, LWL^{+}20, LWX20, LZ20b, LSW20, LWL21, LL21b, LSH21, LL22b, LCW22, LXY23, LZF23, LLX24, LY25, LL25c, LJ24, LDH24, LY21, LZL^{+}21, LYL^{+}22, LSGN23, LLXY23, LZ23, LL24b, LHSF24, LZY24b, LXL25, LLXY25, LTYZ25, Liu25a, MZWD23, MC24, Ma25b, MRAC24, MKM24a, MN21, MM21, MLLL24, MN23a, MKP23, MGA22, MS23, MH23, NAS20, Ngo25, ND21, NR25b, NR22, OTC24, OXY21, Oru20, OAH^{+}24, PXLY23, PHL21, PZL23, PL25, PSSW20, PLVL21, Pło22, PT24, QH22, QZH23, QGQ22, QX23, QL24, QXCG20, RDF24, RZSL23, RYAM21, RH21, RR22, RKK25, RKC25, SLD25, ySRC22, SJ21, SZ22b, SCG23, SZ24b, SQ24, SX25, SCdH21, SZ21, SR24, SMS^{+}23b, SJZ22, SWZ23, SW24b, SHT^{+}24, Tan24, TjW24, Tan25, TO20, TQX23]. equation [TRL23, TCW23, TGARGT^{+}23, UR24, WTSZ20, WH21, WC21, WLJ21, WLX22, WHT22, WL22, jWC23, WCWZ23, WLL23, WF24, WZD24, WSCP24, WLC24, WYW^{+}24, WCC25, WYLL25, WLW^{+}25, WJ25, WWZQ25, WZ25, Waz20, Wie23, WTQ^{+}20, WPLW22, WARX23, WLH24, XHY21, XFLZ22, XZ23, XSL24a, XP22, XHH^{+}23, XSZ^{+}23, XF23, XS25, XHL21, YQZT21, YZT21, YL22, YYZ22, YWL23, YLL24, YJ24, Yan25b, YKK25, YLFZ25, YKMZ21, YEO21, YLZ20b, YLZ^{+}24, YT22, YTT^{+}22, YH21b, Yua23, ZWOM21, ZLPR20, ZZW20, ZJJ20, ZLT^{+}21, ZHYL21, ZO21, Zha22a, ZFL22, ZDWR22, ZYTX22, ZL22a, ZF22, ZY23, ZH23b, ZAH23, ZZS23, ZSF23, ZJ25, ZLL20, ZPC20, ZL23c, ZZ25b, ZLJT20, ZNH24, ZX20, ZXQQ22, dlRGR20]. equations [Abd25, AS20, AM23, AGLR21, AKO21, AJLM24, ASB25, AXLW24, ABR20, AAL24, AJ23, ACRBT23, ABFPA24, AMM23, AN23, APS21, ATBT20, BLLM25, BV23, BWR21, BLF23, BS23b, BdSC24, BHPS24, BDB21, Ben25, BB23a, BKB22, BCCR20, BN20, BHC20, Bra20b, BPR22, BSRVC20, CHY21, CO20, CBI^{+}23, CC20, CS20a, CLZW20, CW20, CZLQ22, CLZ23b, CZ23, CKL25, CYQW21, CSI21, CHJR21, CKS22, CL21, CSS25, CEL20, CDGM25, CMM23, CG25, CP21, CTZ22, Dai22, DS20, DK23, DMX24, DKMS22, DT24, DAS20, DC22, DR25b, EMSA25, FSW20, FZS22, FAM25, FXLS25, FS21b, FK21, FH25, GP22a, GLF25, GYL25, GP22b, GI21, GIRB25, GKHU22, GJS20, GAM24, GZ23, HMW21, HZ21, HJM23a, HJ24, HTBMC20, HTDT23, HLL24, HJT25, HQLJ20, He23a, He23b, HCY^{+}24, HK24, HZR23, HM24, HZW21, Hop23, HHW21, HYLH22, HLPW22, HZA22, HHL22]. equations [HQL23, HWY25, HHG20, JKS25, JRu24, Jha21, JL19, JL20, JAHM23, JHG^{+}21, JZS23, JICRMG^{+}20, JMN22, KUK20, KPO20, KPJO22, KC22b, KC20, KWK24, KCD^{+}23, KLY25, KP20, KWLW24, KJ20, KSS^{+}20, KSJ23, KSSS23, LNT22, LHN24, LLN24b, LL21a, LPK22, LG23, LM24, LL25b, LGL20, LLW20, LQ20, LZ20c, LZ20a, LWX21, LZ22, LXZ23, LT23, LaZZ24, LDW^{+}24, LLZL24, Li24a, LHF24, LF24b, LX22, LS25a, LSRL24, LD20, LFY^{+}21, LSWY21, LZFZ22, LQY22, LMLZ23, LRZ23, LH23b, LL24a, LD25, LHTZ25, LQF25, LZS25, ŁWP^{+}23, LCT22, LZ25c, LYL25, MFZX23, MZGM25, MMC21, MCW21, MRN^{+}20, MW20, MMAK20, MMAK21, MHKB25, MRPK21, MZHJ24, MH25, MGA22, MN23b, MS22, MLL25, MZZ^{+}25, Ngu20, NNT22, NY22, NFS24, NF25, NML23, NSRI24, OP24, Oru25, PQSM21, PH25, PKPD^{+}24, PLZZ25, PS25a, PS21, Pon20, PD25, PK25, PNT^{+}23, PSI^{+}24, QSW23, QFL23, QS20]. equations [Qiu21, RZC21, RDP20, RMBC24, RLSC23, SC21a, SM22a, SM23, SMDC23, SB25, SR23, SHM21, Sch21, SCCS21, SSW20, SZS20, SQL21, SQ23, SS22, SHD23, SDG20, SZ22a, SLT24, SRD25, SWW22, SBY^{+}25, SKG24, SLGF20, SXKL24, SW21, SBNR23, ST22, SXLZ21, SF25, TH23, TFWX23, TS24, Top22, TBS25, TH21, TBO20, Ull20, Vab25, WFH^{+}20, WCL20, WZZ20, WZ22, WZY23, WZTZ23, WFX24, WYHQ24, WGeL24, WZS24a, WZ24, Wan24, WWYL25, Wan25, WS20a, WS20b, WZCS20, WL20, WKC21, WLQ22, WWH23, WCY^{+}24, WZL25, XCFC21, XCLX24, XLLA20, XLCP20, XLS22, XZSY23, XL25b, XLSF25, Yan20, YHZ22, YCCW24, YK24, YL25, YLZ20a, YWZ21, YHLA22, YTH^{+}21, YTA23, YLCC21, YFK20, YFK23, YYSL20, YWPS24, YH21b, YW20, ZM20, ZQM20, ZFZ20, ZZ21, ZD22, ZZL22, ZS23a, ZXXX23, ZDW23, ZF24, ZX24, ZZ24b, ZGC24, ZZZ25, ZW25, ZWCS20, ZLH22, ZPC22, ZGL23]. equations [ZSW23, ZA23, Zha24, ZLZ25b, ZS21, ZSPW20, ZW20a, ZZB20, ZSY22, ZST22, ZWY23, ZLW23, ZH23c, ZHPW23, ZHW20, ZZFL21, dR25a, dVDMM22]. equidistant [LLU24]. Equilibrated [AA25a, BNB20, CNT23, DFK22, Mal22]. equilibrium [ABL24, CLZ^{+}24, DK24, LLXY23, LLXY25, PZZ23]. equipped [MAT^{+}20, TSARR21]. Equivalent [FS21b, IKST21]. ergodicity [WWH23]. ERHSS [SW20b]. Erlangization [YTA23]. erodible [LMD22]. erosion [AEF20]. Erratum [MMAK21]. Error [ADKH20, BMS22, CMY23, DVA^{+}20, DSF24, Gal21, GENRB21, HSJSS25, HS20, JZ20, JMN22, LSV21, Li24b, LLW23a, MCD24, Ma25a, Man24, MA22, Pło22, QH22, STM21, TA22, ZFQ24, AA25a, ACLP22, BGR23, BS23b, BHS20, BHS22, BB24, BZE24, BKR24, BS21, BX20, BDE^{+}24, BZ23, BHL^{+}23, CY23, CC20, CHL22, CHZ23, CYQW21, CDGM25, CNT23, DN22, DFK22, DHVT20, DH23, EGP20, ES20, EFFR22, FLOQ21, GLS25, GQ23, GLF25, GWN21, GCL25, Hak24a, HLW23, HD21, HAC23, Jha21, JCJ24, Kei21, KPS23b, KMN25, KP24, KKNR21, Lee24, LZ21, LL22a, LF24b, LZ25a, LZ25b, LN21, LQY22, LS23, LXZZ24, LSZ24b, Mal22, MCW21, MS20b, MCV24, NAS20, PG21, PHL21, PR20, QZH23, RC25, RKC25, SMDC23, Šeb24, SW20a, SZ22b, SL24, SWZ23, TCA22a, TTKV20, VR20, WC21, WBY23, WS25, XZT23, XXLS23, XP22]. error [YJ24, Yan25b, ZJJ20, ZYTX22, ZC23, ZAH23, ZZH23, ZZ24b, ZLHC24, ZCA23, ZQS24, aKST22]. Error-estimate-based [DVA^{+}20]. errors [DBR21, GB23, ZAPK23]. erythrocyte [PGR23]. ESAV [CGH24, Tan25]. escape [WDGC20]. ESFR [QN20]. essential [HK22]. essentially [GCJ23, WZTZ23]. estimate [BB24, CY23, DVA^{+}20, LF24b, LLW23a, LXZZ24, OAH^{+}24, SWZ23, TTKV20, WC21, XZT23, ZJJ20]. estimates [AA25a, BS23b, BHS22, BZE24, BS21, BZ23, CC20, CLY25b, CYQW21, CDGM25, DFK22, DH23, ES20, EFFR22, FG25, FLOQ21, GLS25, GWN21, HZ25, HSJSS25, JZ20, KMN25, LSV21, Lee24, LZ21, LZ25b, LX24b, LN21, LQY22, LSZ24b, MA22, NT21, PKP21, PHL21, PR20, QZH23, RC25, STM21, SLW22, SZ22b, SL24, TCA22a, WBY23, WS25, XP22, ZC23, ZAH23, ZZH23, ZLHC24, ZCA23, aKST22]. estimating [UR24]. Estimation [YLWZ20, AOL22, BKR24, BX20, BWG25, BHL^{+}23, CHL22, CNT23, GENRB21, Hak24a, JZ24a, JJZ21, Mal22, MCV24, RBSC24, Šeb24, STO22, VR20, VGA^{+}22, WH25, YS23, YYY^{+}20]. estimations [MCW21]. estimator [HLW23, LZ25a]. estimators [EGP20, Jha21, SW20a]. estuarine [MrZQ24]. Euler [AG25, BLO22b, CZLQ22, CSI21, EH23, JZHZ22, Kam25, LL22a, MZZ^{+}25, RDP20, WWLG24, ZXXX23]. Eulerian [GNZ24, LZG^{+}23, RBSC24, SR23, WCH20, YL20b, ZXXX23, ZZY^{+}24]. European [SMR24b, TC25, TZ20]. eutectic [YZJ^{+}22]. evaluated [aKST22]. evaluating [KLRT21]. Evaluation [GOP25, SA22, AG24, GCL25, WWP^{+}25]. evaporation [NM23a]. even [JP25]. even-order [JP25]. evolution [BGK23, CMS23, LARGVR23, MW20, QQX21, ZGL23]. evolutionary [HJ24, LDH24, QWC^{+}23, Rav20, Vab25, ZL22a]. evolving [KYY^{+}20, LF24a, SR23]. Exact [DDG23, DST25, KWK24, KG20a, WJC^{+}25]. exactly [DMX24]. examination [MAT^{+}21]. examining [MGBN22]. example [RHR^{+}21]. examples [ZA23]. excavation [DZZ^{+}24]. exchange [WWP^{+}25]. excitation [HJP23, SN20a]. Existence [BMZ20, ZNC24, DHW20]. exit [HM20, SD25b]. Exner [XDL^{+}25]. exosystem [HJ25]. expansion [KG20b, XL24]. expansions [KP24, LDW^{+}24]. expected [XXY20]. expedited [JFS14, JFS20]. Experimental [EGF^{+}23, SB24b, APGK24, BAM^{+}21, BMK21, KTFI23, MAT^{+}21, NDB^{+}24, Par24b]. experiments [PV20]. Explicit [AAU25, CJ25, FX22, LGZC23, ADB^{+}22, AAL24, CZ23, ERG^{+}24, FBK^{+}24, FZDS22, HLWL25, LSW22, LLZ^{+}23, MWYY22, RDP20, Vak25, WL23, WLD22]. explicit-implicit [FZDS22]. explicitness [TSS24]. explicitness-preserving [TSS24]. Exploring [RNMÖX24, Yan25a]. explosive [Abd24]. Exponential [AFM23, BDS24, AHM20, DE22, EMSA25, FHZ22, FX22, Kam25, LCY23, PVC^{+}25, WWZQ25, XFX22, XF23, YYZ22, ZLQ23, ZZS23, ZZ24a, dR25a]. exponentially [AA25b]. exponents [CBLZ24, SR20]. exposed [Abd24]. extended [ADKH20, AA20, ATBT20, BHDV22, CCW22, CCWC23, DSR23, Ha20, HWC24, KWS21, LD23, MCW25, PZL23, Rob25, SJZ22, SWZ23, WCCC25, YTY24, ZCW24]. Extending [JRZ21]. Extension [WZC^{+}20b, YWL21, BNV22, RFG^{+}21, GHH20, Hak24a, MS20a]. extensions [Hak20]. exterior [ADD22, DFS21, JH24, TjW24]. extra [ZDT^{+}22]. extra-thick [ZDT^{+}22]. extracted [LLZ20]. Extraction [KPO20, KPJO22]. extragradient [LQ24]. extragradient-like [LQ24]. extrapolated [HZA22]. extrapolation [Agg20, HLPW22, LLZL24, PWC^{+}23, VJ22]. extrapolations [WWYL25]. extremal [Ant21]. extreme [LLR25]. eye [AEF23].

faces [GQZ23]. facesheets [AA21]. faceted [ZLJ^{+}22]. facilitating [Loh21]. factor [DCSAM25, He23b, XL25a]. factorization [FCF22]. factorized [LZS25]. factors [ADG23, ASSV20, KPO20, KPJO22]. failure [BJG23, SAR20, WLH24]. failure-informed [WLH24]. FALKSOL [Zdu21]. family [AGLR21, CYM25, HLWL25, JP25, LMZ23, Vis24, WH21, WFMD20, Zha22a, ZC23, ZW20a, ZS24]. Fast [BC20a, CMPR23, CLLG23, GCL25, JHG^{+}21, Liu25b, Lob21, LBPC20, ŁWP^{+}23, PXLY23, QGQ22, SQ23, SLYL22, WJC^{+}25, ZZ21, AAL24, CWX25, CDLW20, DR24, FSW20, FZS22, FZH22, HGW21, HHL^{+}21, HLN24, JJZ21, JWZ22, KP22d, LQ20, LGQZ21, LFGZ23, LX22, LQY22, LLJ23, LLLL24, LN22, MLLL24, Par22b, Par24b, PK25, Pta21, RKRRL22, SJZ24, SNHH21, TH23, TPA24, WWYL25, WPLW22, WW25, XL25a, XLSF25, Yan20]. faster [CL24a, GI20]. faster-than-real-time [GI20]. fatigue [SAR20]. fault [BDGS25, LBTVG22]. favorable [VLD^{+}24]. FCNN [KC22b]. FD [DMŠ22, LQYJ23, SZ21, Ull20]. FD-FV-LB [LQYJ23]. FDM [BFRA21, Jaw21]. FDTD [KJ24]. FE [BSRVC20, CSCY24, HDG23, LL22c, SP20, VR20, ŻW21a]. FE-dABCD [CSCY24]. feature [CDPL24, FZW^{+}25]. features [AWA^{+}22, AAG22, ABD^{+}21]. featuring [ZZ24a]. February [Ano20-66, Ano20-68, Ano21-64, Ano22-51, Ano22-58, Ano23-52, Ano23-54, Ano24-52, Ano24-56, Ano25-54, Ano25-62]. feed [IBY^{+}24]. feedback [CRRS21, KVV23, RNN22]. FEM [HMX21, Ade20, AWA^{+}22, AAG22, BFM24, BHS22, BBI^{+}22, BX20, BRS21, BRRT21, BFM^{+}25, CGO22, DPP21, EGP20, FMS22, HIP23, HSJSS25, Kar24, KPO^{+}21, KCD^{+}23, LFF^{+}24, LSW22, LQLF23, LSW20, LT23, Li24b, LS25a, MPVA25, MMPR20, OCRN22, SW20a, SZ22a, SLW22, SZ22b, SX25, Tan24, TK20, WBY23, XFX^{+}24, YLZ20a, YLFZ25]. FEM-BEM [FMS22]. FEM-IDS [Ade20]. FEM-PIKFNN [XFX^{+}24]. FEMs [GWN21, WLL23, WS25, ZW25]. Femurs [YKNS23]. FEniCS [BHL^{+}23, Kam21, MRS21, ZY21b]. Fermi [HSZ22]. ferrofluid [BPV24, HKL^{+}23, KHH25]. ferromagnetic [EP23]. ferronematics [MMN21]. FEs [YMTS20]. FESTUNG [RHR^{+}21]. Feynman [CYKL23, MZWD23]. FG [KHD24, MJWD20, NTP25]. FGM [HWZQ23, JCP^{+}25, LLZ^{+}25]. FGMs [JMM20, XFX22]. FHN [ZZZ24]. fiber [HGD21, HGD25]. fibre [FGK^{+}21]. fibres [SGAC^{+}20]. Fickian [ZLL23]. fictitious [KP22c, KPK25, Leg21, LZLN25]. field [ASGGRG23, ACXI24, ANWW21, AIG^{+}21, AZY22, BMH22, BWZ22, BWZ23, BPV24, Bal21, BLX^{+}24, BBS22, CLLL25, CFG^{+}25, CLH21, CZLQ22, CXX^{+}25, DMŠ22, EED^{+}25, FSH^{+}23, FCFW21, GOV20, GML^{+}20, HKH^{+}24, HXW23, HCY^{+}24, HJTW25, HMŁW^{+}24, HKL^{+}23, JWW21a, JWL20, JW20, LYP^{+}22, LWSZ21, LAX22, LQYJ23, LLX24, LaZMY25, LJ24, LLW^{+}23b, NNTCB25, PWC^{+}23, PM23, PHL21, QZH23, RST^{+}22, SMŁWL25, SYHN24, SRA20, SY25, WSGT21, WZY22, WXF22, WWCJ23, WHH22, WSO^{+}23, WYT23, XKXL22, XWXF23, XXYL23, XZS^{+}23, YK20, YZJ^{+}22, YWT22, Yan24, YYY^{+}22, YTY24, YWZ^{+}23, YZW^{+}22, ZKBC21, ZBKC22, ZWZD24, ZDT^{+}22, ZTW22, ZWT23, ZL23b, ZJZ24, ZHZ23, KEUY23]. field-finite [RST^{+}22]. Fields [BC20a, BZ21, GO20, TMML23, ZXXX23]. Fifth [GCJ23, TZ25, WZZ20, WZTZ23]. Fifth-order [GCJ23, TZ25, WZZ20, WZTZ23]. filament [WCZD20]. film [WLZZ22, ZDLL22, ZLD^{+}23]. filter [AOL22, NR22, QWC^{+}23, VGA^{+}22, ZYD24]. filtered [NW24a, Pel23]. filtering [CMPR23, Orl23]. filters [ABC^{+}22]. finance [YY20a]. financial [CWC22, SZ21, Yan25a]. Finding [WP25, Boł22]. Finite [BWR21, Bel24, BGG^{+}20, BDG21, BBS21, CO21, CLZ23b, CL21, DK23, DK24, DH23, GCMS21, HH21, HH22, HSZ22, HNLPP20, HMX21, KK23a, KNBR22, KC21, Lin24, MKM24a, MI20, NCRGVR21, PCL25, QQY^{+}23, WZD24, XLLA20, YA21, ZTL22, ZH23b, ZLJT20, AS20, AGLR21, APPS25, AKO21, Adj22, Adj25, ANZZ23, ADFP23, AJ23, ÁCS22, AEZ24, AZY22, AZHY22, ACRBT23, AAB^{+}21, AJK23, APS25, ABD^{+}21, AGPP22, AAM^{+}24, AAWJ20, AN21, BLLY24, BNV22, BDV23, BIMS24, BWZ23, BGR23, BS23b, BZ20, BRR25, BHS20, BHS25, BF25, BM20, BHZ20, BW21, BJG23, BB23a, BM24a, BER21, BS24, BNC^{+}23, BZY20, BBS22, BHL24, BCF^{+}23, BGLZ24, BDGS25, BMvdZ21, Bri24, CAD24, CCW22, CM21, CCGI23, CKFR24, CM23, CCWS20, CWTM20, CLZW20, CLYH22, CWC22, CLH22, CHZ23, CXF24, CCD25, CC25, CSI21, CHL21, CCWC23]. finite [CSS25, CJ25, CCGRL21, CRF^{+}20, Cos23, CTZ22, DLN24, DD20, DGN24, DE21, DL25, DGQT20, DDG23, DFGS24, DCCM23, DST25, DS25, DTV20, DZ24, Du24, DBB23, EAM20, ERG^{+}24, EJR24, EBT24, ELS24, EH23, FSS^{+}20, FZS22, FHM22, FS21a, FHM21, Fia21, FCFW21, FXLS25, FS21b, FP20, FK21, GP25, GLS25, GLL^{+}20b, GQ23, GNZ24, GP23, GHRD20, GOV20, GORBS21, GIRB25, GPY22, GDA21, GMS23, GHLS25, GDA^{+}22, GENRB21, GCC20, GSW21, GHL24, GQZ23, GS23, GGT24, Gun22, GB21, GWYR22, GB23, GXFT23, GHK25, GO23, HC21, HKJ23, HMFM25, Hak24b, HZ21, HJM23b, HLH24, HDM24, HLL24, HSM^{+}23, HLEF20, HLX20, HQLJ20, HWY23, HZ24b, HLQ24, HB25, HOY24, HSY25, HV21, HLCW24a, HYLH22, HX22, HYW^{+}24, HS20, HLY22, HAC23, HLCW24b, HWWZ24, JVBM23, JM20a, JZPG21, JM23b, JZ20, JCP^{+}25, JZS23]. finite [JL23, JP25, Jog25, JBKW23, KSS22, KBR20, KJ21, KMZ22, Kei21, KSD^{+}25, KC22b, KTY23, KMN25, KK21, KKS25, KMP20, Ku24, KSJ23, KC23a, KD23, KVV23, LLSY24, LLSY25, LS22a, LHN24, LP23, LLD23, LWSZ21, LZ21, LD22, LZ22, LL22a, LZA23, LWW24, LZL24, LaZZ24, LLZL24, Li24a, LZ25b, LC23b, LX24b, LFY^{+}20, LS21, LY21, LN21, LSGN23, LMZ23, LLXZ23, LS23, LZ23, LLW^{+}23b, LLLL24, LHSF24, LWFS24, LXZZ24, LLMC25, LXL25, Liu25a, LJY^{+}25, LYLH25, LLWX25, Lob21, LZB25, MFZX23, MGD23a, MC24, MMN21, MSHB22, Man24, MRN^{+}20, MC25, MA22, MRPK21, MLS22, MT25, MZ23b, MLLL24, MKP23, NT21, NC20, Ngo25, NNNT22, NTP25, NY22, NFS24, NW24b, NK25, OAD24, OXY21, OS24, OR22, OSZ22, PS22, PG21, PH25, PLZZ25, PWJ^{+}24, PL25, PK25, QSL23, QL24, QXCG20, QS20, Qiu21, RL24, RS21, RL23, RWZG24, RC25, RST^{+}22]. finite [RKC25, RR23, SG23, SR23, SPC^{+}23, SQL21, SZ24a, SL20, SL21, SJ21, SJ23, SLT24, SZ24b, SQ24, SWW22, SK25, SNHZ25, SGR^{+}23, SLGF20, SR21, SR22, Sto20, ST20b, SS21, SWZC23, SQGZ24, TA22, TCW23, TGARGT^{+}23, TGDMGT^{+}25, Top22, VBDR24, VL21, WYH20, WZZ20, WW20a, WZC^{+}20a, WZC^{+}20b, WS21a, WXZW21, WML21, WDZ21, WWL22, WL22, WGC23, WCZ24, WWC^{+}24, WCS24, WGeL24, WZS24a, WSCP24, WZ24, Wan24, WLC24, WWY25, WP25, WXH^{+}25, WCCC25, WLW^{+}25, WWZX25, WQS25, WZ25, WLW^{+}22, WH20, WHH22, WZS^{+}24b, WN24, WLQ22, XLS24, XYS^{+}24, XHY21, XSL24a, XZ24, XXLS23, XP22, XZSY23, XLSF25, YRC25, YA23, YQZT21, YHZ22, YWT22, YLW23, YLL24, YJ24, Yan24, YWL21, YWZ21, YZZ20, YKNS23, YYSL20, YCGL25, ZBCS24, ZR23, ZD21, ZCYL24, ZH20, Zha21, ZLT^{+}21, ZHYL21, Zha22b, ZD22, ZHZ^{+}22, Zha22a, Zha22c, ZS23a, ZC23, ZQY^{+}23, ZZH23]. finite [ZZ23a, ZSF23, ZZ24b, ZSGT24, ZGC24, ZZ24a, ZZZ25, ZZ25a, ZL25, ZXZF20, ZLH22, ZLL23, ZSW23, ZLZ^{+}25a, ZZ25b, ZS21, ZL22b, ZS23b, ZSPW20, ZW20a, ZW21b, ZXQQ22, ZSY22, ZST22, ZZW23, ZH23c, ZZS24, ZS24, ZHW20, ZX22, ZK21, ZAPK23]. finite-cell [BRR25]. Finite-Difference [BWR21, CCWS20, GWYR22, HLH24, ST20b, VL21, ZAPK23]. finite-element [FHM22, KMN25]. finite-size [GHLS25, PWJ^{+}24]. finite-strain [CWTM20]. finite-volume [MRPK21, Zha21]. fins [SB24a]. first [BM24a, BN20, CZ23, LL22a, OCCM21, QH22, SV21, Zha23, ZHZ23, BB23b]. first- [QH22]. first-order [LL22a, Zha23, ZHZ23, BB23b]. Fisher [ADKH20, DL25, LY25, PZL23, SLD25, SJZ22, SWZ23, TjW24]. fishery [YTY24]. fitted [FMP^{+}24, GYC25, MIOK21, dR25a]. fitting [Leg21, ZZ20a]. FitzHugh [KUK20, LZ20a, LY23, LLLL24, LTYZ25, ZZZ24, ZSY22, WPLW22]. FitzHugh-Nagumo [ZSY22]. five [Ahm24, GOV20, LLU24]. five-by-five [Ahm24]. five-field [GOV20]. five-point [LLU24]. Fixed [MMN24, dSG20]. Fixed-time [MMN24]. flame [BTYM^{+}21]. flames [SB21]. flapping [RGS^{+}20]. flat [dSMAdOW22]. flat-top} [dSMAdOW22]. flaw [MR24]. FLEXI [KBB^{+}21]. Flexible [BBK21, GSW21, ISA24, JJOG20, WCZD20, ZWW^{+}23]. flexibly [WS21b]. Flow [NDB^{+}24, YCZ^{+}21, ZSH^{+}21, AI25, ACXI24, ADR24, AMPB^{+}22, AAG22, APGK24, AGA^{+}21, AIG^{+}21, AA20, AEZ24, AFV22, AAS^{+}20, AAK24, BWZ22, BWZ23, BPV24, Bal21, BRÁ^{+}20a, BLX^{+}24, BBK21, BFI^{+}23, BV21, BDE^{+}24, BBS21, BP20, Bou24b, BLM^{+}24, BPID20, CCG^{+}23, CLZ^{+}24, CRDS20, CCWS20, CQS25, CDG^{+}23, DBH^{+}21, DAAEA25, RFG^{+}21, DKH22, DK23, DNP^{+}22, EFF21, EGF^{+}23, EFS24, FLS23, GNZ24, GPL21, GdFFF25, GAM22, GCC20, HANA24, HLEF20, HHY^{+}22, HK24, HLCW24a, HYZZ24, IBY^{+}24, ISA24, JZ25, JSBZ22, JZXR24, JWW^{+}21b, KK23a, KARS23, KFD^{+}20, KGW^{+}21, KMS20, LMP23, LH24, Lee21, LHC^{+}22, Li21, LSH21, LpWW^{+}25, LLY25, LLJ23, LCC^{+}24, Liu25b, LZL^{+}25, MIOK21, MRPK21, MD21, MZCK25, MWRA22, MS20b, NDC22, NIT^{+}20, NK25, OAA^{+}22, OSO^{+}23, OEASY24, OKSS21, OSZ22, PM23, PWMC23, PGK20, PFCZ23, QC24, RG21]. flow [RZŠ21, RGS24, SMR^{+}24a, SB21, SGJJ20, SD25b, SK23, SGAC^{+}20, SA22, SJZ24, SBH^{+}21, SRA20, SRP21, SPZL21, TKT^{+}24, TKK23, Vah22, WDGC20, WWQ20, WZC^{+}20a, WZC^{+}20b, WWW21, WSGT21, WX22, WWL22, WDGS23, WL23, WYP23, WFW^{+}24, WS21b, WZXN21, XYT^{+}21, XTZZ24, YD23, YY23, YPC^{+}21, YK20, YL20b, YQZL22, Yem24, YCWS20, ZL20b, ZO21, ZGW23, RSB^{+}21]. flowing [KC23a]. Flows [Agg20, BEKV21, AS21, ASH^{+}21, BLLY24, BBDM21, CLLL20, CBK24, CGHJ24, sCpLYhG23, CYC^{+}23, CLW^{+}24, CCGR24, CRPdT20, CCGRL21, CL22b, CPBF21, DCCM23, Du24, EH23, FNGDV23, GP25, GS22, HUT25, HLZ24, HEDT20, HBC^{+}21, KP22a, KP22b, KJ21, KHH25, LBTVG22, Lee24, LL20a, LMD22, LZF23, LS25b, LLZ^{+}23, LLW^{+}23b, LSY^{+}24, LZB25, MWYY22, MH22, MG25, MGBN22, MRN^{+}20, MCW22, MKAM20, NC20, NAE22, PL24, QD23, Rav20, SMŁWL25, SD21, SJ24, SYZ23, SNL23, SR22, ST20b, SQLZ21, TMML23, WLX20, WZC^{+}20a, WZC^{+}20b, WMMH25, WRMB23, WW25, XL22, XYS^{+}24, XRS25, YA23, YLY20, YCWS20, YWZ^{+}23, ZP24, ZKTK21, ZCXG20, ZHZ^{+}22, ZYS22, ZLL23, ZZY^{+}24, ZY21b, ZFQ24]. Fluid [ZSH^{+}21, AL22, ADB^{+}22, ANH^{+}21, AGA^{+}21, APS25, AAK24, BWZ22, BGR23, Bal21, BLX^{+}24, BP20, CCG^{+}23, CBK24, CYM25, CPBF21, Du24, FL22, GPL21, GV21, GCSF21, HC23, HANA24, HHR^{+}20, ISA24, JSBZ22, JZXR24, JJOG20, KWM22, KK23a, Kam21, KARS23, KJ21, KFD^{+}20, KMS20, KC23a, LMP23, LFF^{+}24, LWL^{+}20, LYL^{+}22, LLJ23, LL24a, LZL^{+}25, LR22, MSHB22, MLS22, MD21, MLHS23, NSMR25, OSO^{+}23, OEASY24, Oru22, PTNR24, RLVR21, RL23, RGS24, SJ24, SJZ24, SWW22, SBH^{+}21, SRA20, SNL23, SWC24, SHXZ24, Vak25, VSG21, WZC^{+}20a, WZC^{+}20b, WYP23, WR23, WS21b, WYT23, XYT^{+}21, XWXF23, YD23, YY23, YK20, YCZ^{+}21, Yan24, YCWS20, ZBCS24, ZKTK21, ZY21a, ZYD24, dCLW25, AFBM25]. fluid-fluid [AL22, CYM25]. fluid-poroelastic [WR23]. fluid-saturated [LR22]. fluid-solid [SWC24]. fluid-structure [CCG^{+}23, GV21, ISA24, KWM22, KJ21, LFF^{+}24, MSHB22, NSMR25, RLVR21, SHXZ24, dCLW25]. fluid/rigid [ZBCS24]. fluid/rigid-body [ZBCS24]. fluidized [GORBS21]. Fluids [ZSH^{+}21, BABS20, CBK24, FGK20, GML^{+}20, HLH^{+}20b, HMŁW^{+}24, JSZ24, LMBB24, LL20a, Liu25b, XKXL22, ZTW22, dSG20]. flux [AA25a, CBLZ24, CYC^{+}23, CNT23, DT23, HDWL24, JP25, Lee21, LLW23a, LS25b, LXL25, Loh21, MWYY22, RS21, SLX^{+}23, SHT^{+}24, ZCY22, ZCYL24, ZYS22]. FMM [LSPRV24]. foams [BDJ20]. Fock [YY20a]. focus [KGW^{+}21, MSHB22]. Fokas [BHDV22]. Fokker [LD25, WCW20a, HHR22, LHTZ25, PS25b, RKK25, YTH^{+}21]. folding [ZZW25]. for= [BHD^{+}21]. force [ANH^{+}21, SCD25, WYS^{+}20, WWP^{+}25, ZDT^{+}22]. Forced [RSAG^{+}24, LL24b]. forces [BLO22b, QD20]. Forchheimer [ALOR25, Gut20, HZ24c, ZL20a]. Forchheimer-Brinkman [Gut20]. Forcing [BBS21, GO25, WDGS23]. forecasts [ABC^{+}22]. form [AMV24, BDP23, Boł22, DS20, DW20, Füh21, HKM21, JM20a, KP24, SCdH21, TH23, TLJ24, TBS25, YHZ22, ZZ25a, ZK21]. formalism [Jog25]. formation [Bar22, KYY^{+}20, LH24, YLQ22]. forms [LLZ^{+}23, WSS22]. formula [CYKL23, MGD23a, Men22, WL22, XLCP20]. Formulae [SNHZ25]. formulas [CDM24, Din21, KPO20]. Formulation [LWL25, ARV23, BV21, BD23, CWZ21, CLLL25, CWTM20, CRF^{+}20, CLL^{+}24, Füh21, GCC20, HC23, HRS24, HGD25, HV21, JCP^{+}25, JJOG20, KC21, LLX23, ŁSSP23, Med23, NC20, RMBC24, Šeb24, SK25, TBO20, ZR23, ZS23a, ZF24, Zha24, ZLZ25b]. formulations [ACRBT23, BB21, CHD24, CM20a, CG22, DMRV24, GI21, Hoa22, KJ24, MHHR20, OAD24, RH21, SV21, vtWHG^{+}22]. Fort [DLW24]. Fort-Frankel [DLW24]. Fortin [DZ20]. Forward [AS23, LBTVG22, BFE^{+}21, IBY^{+}24, LWG^{+}25a, WTSZ24, ZWY23, LSPRV24].

FOSLL* [Kei21]. FOSLS [FG25, MS23]. foundation [AA21, HTPV24, YLFL20]. Foundational [NW24b]. foundations [YX21]. four [CLLL25, NW24b, WN24]. four-field [CLLL25]. Fourier [FZW^{+}25, JZ20, LDW^{+}24, LZS25, MLL25, SLD25, SWWZ23, SW24a, WCW20a, YAA25, dlRRG^{+}24]. Fourier-spectral [SWWZ23]. Fourth [CLL21, PD25, AD20, AAC23, AKO21, BZ22, CLZ23a, CLLG23, CK23, CXF24, DPP21, DLN24, HSZ22, HLPW22, HAC23, HQL23, LX24b, LS21, LY21, LdO21, MFZX23, Ma25b, Men22, PZL23, QFL23, Rou20, SJ21, TS21, TRL23, WCL20, WYW^{+}24, WWYL25, WYLL25, WL20, XZ23, YZT21, YL22, YHF24, YFK23, ZYTX22, ZAH23, ZZS23, ZL23b, ZCA23]. Fourth-order [CLL21, AD20, AAC23, AKO21, BZ22, CLZ23a, CLLG23, CK23, CXF24, DPP21, DLN24, HSZ22, HLPW22, HAC23, HQL23, LS21, LY21, LdO21, MFZX23, QFL23, SJ21, TS21, TRL23, WYW^{+}24, WWYL25, WYLL25, WL20, XZ23, YZT21, YL22, YHF24, ZYTX22, ZAH23, ZZS23, ZL23b, ZCA23]. Fractal [CBK24, QWLQ25]. Fractional [AZL20, GSS24, HCZ25, LH23a, SSZ20, AD20, AM23, AFS24, AMK25, AAD24, AAS^{+}20, BDMX21, BML22, BS23a, BFM24, BWZ22, BWZ23, Ben25, BKM23, Bra20b, BWMW23, BGD22, CCH24, CWX25, CRL^{+}25, CLL21, CLLG23, CS20b, CDLW20, CLZW20, CL22a, CLYH22, CGH24, CL21, CV20, CP21, DYC^{+}22, DY25, DCCM23, DHVT20, DAS20, DC22, DR24, DR25b, FKW22, FC24, FSW20, FZS22, FAM25, FRW22, FZH22, GLL^{+}20b, GP22b, GHH20, GKHU22, GHAU24, GWN21, GAM24, GCMS21, GCL25, HLMM20, HH23, HZR23, Hof20, HHR22, HYW^{+}24, HGW21, HJXW23, HS20, HAC23, HQL23, HW25, HWY25, HHG20, JH24, JJZ21, JWZ22, JCJ24, JHG^{+}21, JSBZ22, JWR^{+}23, JSZ24, Kam25, KG20a, KCD^{+}23, KSS^{+}20, KSSS23, KGK25, KP23, KR24, KC23b, LWX20, LWX21, LCW22, LC23a, LXY23, LT23, LLZL24, Li25, LWL25, eLLAN25, LSRL24, LFY^{+}20, LY21, LZL^{+}21, LYL^{+}22, LRZ23]. fractional [LLJ23, LZS23, LL24b, Liu25b, LD25, LZ25c, MFZX23, MC24, MZGM25, MCW21, MI20, MW20, MMAK20, MMAK21, MLS22, MZHJ24, MLLL24, ND21, Oru22, OAH^{+}24, PQSM21, Pon20, PK25, QGQ22, QL24, QXCG20, RDF24, RYAM21, RR22, RKK25, RKC25, SC21a, SLD25, SD25a, SM22a, SG23, SMDC23, SHM21, SZS20, SYZ23, SQL21, SQ23, ySRC22, SCG23, SKG24, SIAA25, SLYL22, SF25, Tan24, TQX23, TH23, Top22, TH21, UHLZ24, UHL^{+}24, WFH^{+}20, WS21a, WC21, WLX22, WHT22, WZ22, WCWZ23, WYHQ24, WZD24, WQ24, WYW^{+}24, WWYL25, Wan25, WCC25, WYLL25, WLW^{+}22, hWCsLjH21, XCFC21, XZ25, XLLA20, XLCP20, XLS22, XSZ^{+}23, XF23, Xu24, XLSF25, YQZT21, YL22, YCCW24, YLL24, YL25, YHL25, YLZ20b, YLZ^{+}24, YFK20, YAA25, YPZ^{+}21, ZM20, ZLPR20, ZZW20, ZJJ20, ZZ21, ZLT^{+}21, ZDWR22, ZZL22, ZF22, ZY23, ZAH23, ZZS23, ZLJT20, ZNH24, ZX20, ZZB20, ZXQQ22, ZHPW23]. fractional [ZZ24c, ZQS24, ZZFL21]. Fractional-order [GSS24, HCZ25, LH23a, CCH24, MI20, Xu24]. fractional-step [MZHJ24]. fractionally [AMK25]. Fracture [BV21, ANWW21, BFI^{+}23, BT22, CXJ25, Ha20, JWW21a, SMS23a, SYHN24, YZW^{+}22]. fractured [BBDM21, LWFS24, SNL23, WW25]. fractures [AFV22]. fracturing [Alf24]. fragmentation [BT22]. frame [SW24a]. Framework [ZA23, BBD^{+}21a, BEG^{+}21, BBD21b, BMvdZ21, CG21, CHD24, CKBTS24, CLXZ25, DD21, HSR25, HLH24, KJ21, KBB^{+}21, LP21, MKAM20, RZSL23, RLF21, RHB25, SD21, SMS23a, WZC^{+}20a, XHM20, Xu24, Zha22d, ZLF^{+}22]. Frankel [DLW24]. Fredholm [JKS25, LX22]. Free [KKJ^{+}22, MJWD20, ADKH20, AHLZ24, BFI^{+}23, BBM25, BLMV24, CWZ21, CLLL20, CM22, CXXZ23, CS23a, CLL^{+}24, CGI^{+}22, DMX24, DNP^{+}22, EKB20, GNB24, GHL24, HWL^{+}20, HFG23, HZ24c, HV21, HBC^{+}21, HTPV24, HTPV25, HWWZ24, JZ25, JCP^{+}25, JWR^{+}23, KPJO22, LM24, Li20, Li21, LL21b, LLX24, LR23, MCD23, MGD23a, MCD24, Men22, MWY23, MS23, MNB23, NP24, NCCD24, OCRN22, PCL25, PZZ23, SK23, SGSD20, ST22, TLZF20, WCW20b, YPC^{+}21, YL20a, YKK25, YA21, YZZ20, Yu23, ZZCM20, ZL20a, ZHH21, Zha24, ZX22, dSMAdOW22]. free-surface [CLLL20, SK23]. frequencies [BL20]. frequency [AGPP22, BDJT21, LVDDNT25, MR24, ŻW21a, ZAPK23]. Fresnel [Par24b]. friction [ZNC24]. frictional [ABC^{+}21, WXCL25]. Friedrichs [CKBTS24, LQ20, PV20]. frog [BNSZ24, YWZ21]. Front [NM23a]. Front-Tracking [NM23a]. FSDT [EJWD23, KHD24, LYXC21]. fuel [FGK^{+}21, SB24b]. fueled [SB21]. full [BFE^{+}21, BFM^{+}25, HKR22, HGD21, JZHZ22, KMN25, LSW22, WYS^{+}20]. full-discretized [KMN25]. full-implicit [WYS^{+}20]. full-space [LSW22]. full-wave [HKR22]. Fully [AGLR21, BGR23, GP25, Hoa22, HYH23, PT24, RDP20, TO20, WWC^{+}24, AZY22, BAB^{+}20, CCGI23, CL22a, CHZ23, CCGV22, DN22, DG22b, DMX24, GOV20, GORBS21, GGRBRG20b, GSRG21, JMS^{+}24, KHH25, LL24b, MSHB22, NR25b, RLF21, SYZ23, Vak25, WMMH25, XP22, YJL22, ZR22, ZS23b]. fully-discrete [DN22, JMS^{+}24]. fully-mixed [CCGI23, GOV20, GORBS21]. function [ACV21, ACLP24, AM22, DZZ^{+}24, GLÖ21, Ila20, JZXR24, JBKW23, LYRW21, LLZ^{+}23, LLU24, LQF25, MCZL20, MH23, NW24b, Oru20, PWBS25, QWLQ25, RUMT^{+}24, SA22, SZ21, SWZC23, TO20, WN24, Yan25a, ZF24]. functional [LZ25a]. Functionalized [ZO21, Tan25]. Functionally [JMM20, AKA^{+}21, BM24b, DCSAM25, EJWD23, FYXL21, HTPV24, HTPV25, LYXC21, MKM^{+}24b, Oru25, PND^{+}24, PCL25, WQS25, ZD21, ZRL^{+}22, ZZL23]. functionals [Wol20]. functions [BIMS24, BNB20, BGKV20, CS22, CM23, CWX22, EMSA25, FJZC20, GK21, HCZ25, JKS25, LCC^{+}24, MFF25, OR22, Rou20, ST20a, Sch24, WCL20, XCFC21, ZLJ^{+}22]. Fundamental [SLUW^{+}21, AMV21, Gra21, LCT21, LGL24, MCZL20, QFL20, RBSC24, dFLdSLM22]. furnace [BNQ21, YLQ22]. Further [XTZ^{+}21]. future [KJ22]. FV [LQYJ23, BSRVC20]. FVLBM [WCZ24].

GA [PSI^{+}24]. gain [XXY20]. Galerkin [KSD^{+}25, AD20, ADKH20, AS23, ADB^{+}22, AJK23, AFV22, AMM23, BS23a, BNSZ24, BGR23, BS23b, BAKR24, BDG21, BX20, BZ23, BKB22, BS23c, BLLZ24, BSRVC20, CY23, CHY21, CG21, CO21, CLR23, CLRT24, CO20, CRM22, CM23, CHD24, CM22, CZLQ22, CXXZ23, CXJ25, CC25, CMR22, CCGR24, DG22a, DP25, DS25, DW20, GM21a, GdFFF25, GHMD22, GDA21, GNB24, GRR20, GHL24, GWN21, GQZ23, GYC25, Haj21, HZ21, HWW21, Hop23, HHR22, HBC^{+}21, HYJ20, HHW21, HRvdZ20, HWWZ24, HA23, JMCC24, JJZ21, Kam25, KCW22, Kar24, KC21, KPS23b, KJN21, KGJN24, KBB^{+}21, KSJ23, KD23, KWS21, Lee21, LG23, LM24, LSW20, LL21b, LCW22, LMD22, LY23, LL23, LWW24, LZBC24, LY25, LaZMY25, LN21, LQY22, LSGN23, LTXZ24, LXZZ24, LD25, LHTZ25, LMST20, LdO21, MCD23, MGD23a, MCD24, nMjWlS24, MCW21, MA22, MLLL24, NAS20]. Galerkin [Ngo25, ND21, Nie23, NERB^{+}22, OSO^{+}23, PKP21, PK23, PS25a, PDB22, Pło22, PT24, QSW23, QC24, QXCG20, RZC21, RZSL23, RTA20, RWaZ23, RHR^{+}21, RHA21, ST20a, SCdH21, SSK^{+}23, SR22, Sto20, SW24b, TZX23, TS21, Tau22, Top22, WLX20, WFH^{+}20, WW20a, WLX22, WWL22, WGC23, WaZWZ23, WW24, WF24, WJZ^{+}24, WXH^{+}25, WJ25, WF25, WZS^{+}24b, Wie23, WKC21, XL22, XZT23, YH21a, YZN22, YWL23, Yan25b, YA21, YZZ20, YHLA22, YH21b, ZQM20, ZH20, ZL20a, ZFL22, ZR22, ZQY^{+}23, ZXXX23, ZLHC24, ZGC24, ZZW25, ZLL20, ZXZF20, ZPC20, ZPC22, ZL23c, ZSW23, ZZY^{+}24, ZXQQ22, ZX22]. Galerkin-type [ST20a]. Galerkin/least [KSD^{+}25]. Galerkin/least-square [KSD^{+}25]. Galilean [HHR^{+}20]. game [YTY24]. gap [Bar22, CLH21]. gas [Ben25, BNC^{+}23, BEKV21, CLW^{+}24, JM20b, LZF23, LpWW^{+}25, LLZ^{+}23, MZZ^{+}25, RCSM23, SMŁWL25, SCCS21, SRP21, TWH^{+}22, WLX20, WTSZ20, WLZZ22, XSL24b, ZL20b]. gas-kinetic [MZZ^{+}25]. gas-liquid-solid [LpWW^{+}25]. gas-particle [ZL20b]. gas-plasma [BNC^{+}23]. gas-surface [CLW^{+}24]. gas/HE [RCSM23]. gaseous [FL22]. gates [WDGC20]. Gauss [FSS^{+}20, HSR^{+}23, LQ20]. Gaussian [CFG^{+}23, HJP23, LCX^{+}20, LLU24, Shi24, WWH23, Yan25a]. Gelfand [IZ20]. General [ZFL22, AJLM24, ABL24, BZDQ21, BJyST24, GGH25, HZ24a, HLJ25, LX22, LSF25, MZM21, Nie23, SZ22b, TKK22, WW25, YWL21, ZQM20, Zha23, Zha22d]. Generalised [HRS24, MN24, GP25, LSY^{+}24]. Generalization [Jaw21, SZS20]. Generalized [CCD25, JBKW23, QSW23, SLGF20, AHOB24, BMZ20, CBŁP23, CSS20, CCF^{+}21, CG25, FSS^{+}20, GHAU24, GAM22, GCMS21, HLQ24, HLCW24a, HYLH22, KKS25, LQ20, LSW20, LL23, LSPRV24, MKM24a, MS20a, MLHS23, Nde25, OEASY24, Oru22, PP24, QSL23, QL24, SM22b, SS22, SJZ24, SZ21, SIAA25, SQGZ24, TGDMGT^{+}25, WZY23, WLW^{+}25, XZ24, YCWS20, ZZ25a, ZL22b, ZQ25, BGG^{+}20]. generalized- [CBŁP23]. generate [JMM20]. generated [JM20a, YOF20]. Generating [SS21, SS24, HZW21, KP21]. generation [FGK^{+}21, Ima20, LYLH25, SNHH21, Yan25a]. Generative [LWG^{+}25a, MSBO23]. genetic [PSI^{+}24, ZW20b]. genuinely [SQLZ21]. geomechanics [XMQ^{+}21]. geometric [BEIL21, HC21, HWZQ23, IKST21, Li21, LSH21, RHA^{+}24, XJY^{+}20]. geometrical [LS20]. Geometrically [PDB22, EJWD23, KNBR22, LYH24, WQS25]. geometries [BGKV20, EH23, LQW22, NKPA^{+}25, Pup25, RTA20, SMŁWL25, VJ24, YWLF21]. geometry [LH24, SA22, XLCP20]. GeoMFree [XMQ^{+}21]. geophysical [ADR24, Lee21, LMBB24]. geostrophic [MS22]. geothermal [CQG^{+}24, ZL23a]. Gevrey [CL24b]. GFDM [CCGF25, ZZ25a]. GFDM-DPIM [CCGF25]. GFEM [dSMAdOW22]. ghost [Ha20, LWL^{+}20, Oru25, WCZ24, YLCC21]. ghost-fluid [LWL^{+}20]. Gibbs [HRvdZ20, Zha22b]. Gibson [AAC23]. Gilbert [NR25b]. Ginzburg [HYH23, LSW20, LL21b, SRD25, WLL23, XSL24a, XP22, ZLPR20, ZZS23]. GJF [MCW21]. GJF-Petrov [MCW21]. glass [Ha20]. glioblastoma [PVC^{+}25]. glioma [LARGVR23]. Global [LYZ23, PWC^{+}23, WLW^{+}22, YLW23, Yua23, ZLQ23, ANWW21, LWL21, ZC22, Zhe22]. globally [YY25]. GMMP [HS20]. GMRES [Ahm24, KCC20]. GMsFEM [PFCZ23]. GNN [NBF^{+}24]. GNU [RHR^{+}21]. Goal [CNT23, DBR21, ELS24, Mal22, VR20, BBK21, BBI^{+}22, BMvdZ21, HIP23]. Goal-oriented [CNT23, DBR21, ELS24, Mal22, VR20, BBK21, BBI^{+}22, BMvdZ21, HIP23].

Godunov [XTZ^{+}21]. Godunov-type [XTZ^{+}21]. Goldberger [SBY^{+}25]. Golub [ATBT20]. Gordon [AJ23, CC20, GZ23, GFXZ25, JZ20, SZ22a, WL22, WL25, YHF24, ZW25, APPS25, FZH22, JS25, JCJ24, WH21]. governed [LCW22, LC21, LYL25, WLX22, WZD24, XCLX24, ZCL20]. governing [YTT^{+}22]. GPAV [QZH23]. GPAV-based [QZH23]. GPU [CPBF21, DdSP21, FZDS22, GM22, KLRT21, KTFI23, MJR20, WKC21, XRS25, ZCXG20]. GPU-based [FZDS22, KLRT21]. GPU-parallelized [ZCXG20]. grad [HJM23a, HJ24, JZS23, LXRQ22, LH23b, PV20, QHHW20, WCH23, LLZ^{+}23]. grad-div [HJM23a, HJ24, JZS23, LXRQ22, WCH23]. grade [Oru22, SJZ24]. Graded [JMM20, AKA^{+}21, AZ24, ABP25, BM24b, BRRT21, DCSAM25, EJWD23, FYXL21, HYW^{+}24, HTPV24, HTPV25, KC23b, LZY24a, LMLZ23, LYXC21, MKM^{+}24b, Oru25, PND^{+}24, PCL25, RR22, RKK25, SU23, Top22, WQS25, ZD21, ZRL^{+}22, ZZL23, ZXQQ22]. Gradient [BBDM21, cZsCyC23, AADMA23, BZE24, Bra20b, BCM22, CSCY24, DSL24, FK20, HYZZ24, IJS^{+}24, KN24, KK23b, LL25a, LC23b, LZY24b, MRPK21, SJM24, WZT23, WYP23, ZO21]. gradient-based [BCM22, IJS^{+}24]. gradient-flow [MRPK21]. Graetz [AA20]. Graham [RUMT^{+}24]. grains [KRESA21]. granular [FL22]. graph [AFM^{+}24, MN24]. graphene [AA21, HLY22, HTPV24, LZA23]. graphene-based [HLY22]. grating [XLS24]. gravitational [CZLQ22, ZXXX23]. gravity [HKL^{+}23]. gray [LP22b, CLLG23, GW25, HMFM25, LFY^{+}20, WWG^{+}25]. gray-scale [LP22b]. gray-scale/color/outline [LP22b]. greedy [LZ25c]. Greeks [MT25]. Green [ACLP24, MCZL20]. grey [MBSEA24]. grid [ABLS23, BZY20, CLZW20, CLH21, CLH22, CXJ25, DZ24, Du24, GWYR22, HZJ^{+}22, HX22, KTXP21, LD22, LT23, LLZL24, LCH23, LLW^{+}23b, LLLL24, MrZQ24, MIOK21, MJR20, OTC24, SPZL21, Tan24, TCW23, WCH20, Wan24, WZ25, WS20a, XHY21, YJL22, YZN22, YLZ20a, YOF20, ZFZ20, ZHYL21, ZGC24, ZL25, ZH23c]. grid-adaptive [KTXP21]. grids [AJLM24, AFM^{+}24, CLZN24, CSI21, GW25, HZ24a, HLW24, KJN21, KLK^{+}24, KWK^{+}20, LKC^{+}24, NDC22, NY22, NFS24, WJ25, ZCYL24]. Gross [FHZ22, MZGM25, WTQ^{+}20]. Ground [SSdA20, XHM20]. groundwater [BLM^{+}24, CDG^{+}23, SQC23]. group [XZSY23]. growth [Alf24, BPID20, DP25, HKJ23, JW20, LYP^{+}22, LAX22, LQYJ23, RST^{+}22, SWWZ23, YRC25, YZJ^{+}22, YYY^{+}20, ZWX24]. Grünwald [XLCP20]. Grünwald-formula [XLCP20]. Guaranteed [Ku24, KKNR21, Wol20, ZY21a]. guess [CKS22]. Guzmán [DST25]. gyroidal [APGK24].

Haar [AAU25]. half [jWC23, YLFL20].

half-space [YLFL20]. Hall [AAC23, OSO^{+}23]. Hamilton [AS20, CLXZ25, LLN24b, LZ22, YT22]. Hamiltonian [BMS22, DHW20, FZH22, Jog25, LRZ23]. Hamiltonian-preserving [LRZ23]. hand [Kra20]. handy [BRR25]. hanging [Sch24]. haptotaxis [BGG^{+}20, NCRGVR21]. hard [LDW^{+}24, Par22c]. Harmonic [Hak24a, BZ23, DV22, Hak20, LAL^{+}24, LYLH25, MN23b, SDJ^{+}22]. Harrison [AHOB24]. harvesting [NM23b, TKMK25, YTY24]. Hausdorff [QWLQ25]. HDG [CKBTS24, LL22b, TZ21]. head [TBS25]. head-based [TBS25]. Heat [BPV24, MAT^{+}20, MLHS23, ZSH^{+}21, Abd25, ANH^{+}21, AGA^{+}21, AL23, ALOR25, BWZ22, BTCV20, Bou24a, BCM22, CRDS20, Cha24, CZ20, CFT^{+}23, ČPPŠ20, DBH^{+}21, DD20, DS22, FL22, FYXL21, GvV22, Gra21, GZAR23, HANA24, HZ24a, HLCW24a, Ima20, IBY^{+}24, JSBZ22, JSZ24, JWW^{+}21b, KMN25, LWL^{+}20, LY24, LYRW21, LGZC23, LLJ23, Liu25b, MLJ^{+}20, MBSEA24, MN24, PM23, QX23, RG21, RH21, SADW21, TSARR21, TJYL22, Vab22, WQ24, XKXL22, XHL21, YLY20, YCZ^{+}21, YYZ22, YWL21, YCGL25, ZLYL20, ZWOM21, ZRL^{+}22, ZHHL23, ZZL23, ZSF23, ZX24, ZZ25a, ZSL22]. heated [AKK^{+}25, CRB24, HKJ23, KP21, Oru22]. heating [AA25b, LSV24, SADW21, SLW22, WWQ20, ZCZ^{+}21]. Heaviside [ACV21]. height [BRÁ^{+}20a]. Hele [Bar22]. Hellinger [Vis24]. Helmholtz [AXLW24, AMV21, BHPD23, BZ20, BF25, BGH20, BDJT21, BPID20, CWX22, DFS21, DT24, EMSA25, EFK^{+}23, FWBY25, HMW21, HNLPP20, KF22, LSW22, LZLN25, MGD23b, MMPR20, MS23, NAS20, Tad23, YKMZ21, YH21b, ZPC20]. Helmholtz-type [LZLN25]. hemivariational [BMZ20, FYM23, LXH24, MD21, hWCsLjH21, XCLX24]. hemivariational-variational [LXH24]. hepatic [PTM^{+}23]. Hermite [BNC^{+}23, CTZ22, GCMS21, WZA25, WCY^{+}24, ZZZD24]. Hermite-type [ZZZD24]. Hermitian [MS20a]. Herschel [AI25, FNGDV23, HMŁW^{+}24]. Heston [MT25]. heterogeneous [BNB20, CZC^{+}25, CEL20, DdSP21, ERHL22, HMW21, JMS^{+}24, LSW22, LMZ23, LYZ23, LD23, MWY23, OCRN22, OTC24, PFCZ23, SM23, SHR25, SLC^{+}24, YD21, YKMZ21]. heuristic [sCpLYhG23]. hexahedral [HWL^{+}20, JZPG21, JMM20, LZG^{+}23]. HHO [MBG24]. Hierarchical [BHL^{+}23, CRF^{+}20, CDPL24, AMP20, BGKV20, CTGR24, GP22a, Gun22, HLX22, zLqLZ^{+}24, MZ23b, WNF^{+}25]. High [BBD21b, CT22, DdSP21, FXLS25, GM22, GHLS25, GYC25, GFXZ25, HJTW25, JZHZ22, MXLY24, MGA22, MLL25, MZZ^{+}25, NDC22, NF25, QD20, SLM23, ST20b, Wan25, WKC21, XHH^{+}23, YTK21, YZ21, YLZ^{+}24, YFK20, Zha22c, ZXXX23, ZLL23, ZLZ25b, ZW21b, Adj22, AL22, AAU25, ADB^{+}22, ABMS20, AGPP22, BNV22, BF25, BEG^{+}21, BGH20, BCT21, BDJT21, BS23c, CBŁP23, CG21, CO20, CYC^{+}23, CDM24, CL21, CHL21, CCGR24, CRF^{+}20, DD21, DKV21, DR24, EAM20, FSS^{+}20, Fan23, FWBY25, FHZ22, FX22, GDA^{+}22, GML^{+}20, Ha20, Hak24b, HLX20, HBC^{+}21, HHR24, HHG20, JCJ24, JWR^{+}23, JYHS23, Kei21, KBB^{+}21, KMS20, LL25a, LQ20, LS25a, LSY^{+}24, LZLN25, LVDDNT25, MWYY22, MG25, MRN^{+}20, NM23a, Nde25, NSMR25, Ngo25, OS24, PSI^{+}24, RWZG24, RR22, SHM21, SCdH21, ST23, SU23, SAR20, TL22, TH23, Ull20, WDGS23, WCWZ23, WDYZ24, WCC25, vtWHG^{+}22, Xen20]. high [XFLZ22, XZ23, XZ24, XXY20, YSGD20, YHZ22, Yem24, ŻW21a, ZZL22, ZZY^{+}24, ZQZ20, AMV24]. High-accuracy [YZ21, DR24]. High-accurate [HJTW25]. high-contrast [vtWHG^{+}22]. high-cycle [SAR20]. high-dimensional [AAU25, CDM24, JCJ24, PSI^{+}24, ST23, SU23, Ull20, ZZL22]. high-efficient [JWR^{+}23]. high-frequency [AGPP22]. High-Order [BBD21b, QD20, CT22, FXLS25, GHLS25, GFXZ25, MLL25, MZZ^{+}25, NDC22, NF25, SLM23, ST20b, WKC21, XHH^{+}23, YTK21, YFK20, ZLZ25b, Adj22, ADB^{+}22, BNV22, BF25, BGH20, BCT21, CBŁP23, DD21, FSS^{+}20, FWBY25, FHZ22, FX22, Hak24b, HHR24, Kei21, LL25a, LQ20, LSY^{+}24, LZLN25, LVDDNT25, MWYY22, MG25, MRN^{+}20, Ngo25, OS24, RR22, SHM21, WDYZ24, Xen20, XFLZ22, XZ23, YSGD20, YHZ22, Yem24, ZZY^{+}24, ZQZ20, AMV24]. high-order/ [HHR24]. High-performance [GM22, BEG^{+}21]. high-power [DKV21]. High-precision [MXLY24, WCC25]. high-velocity [Ha20]. Higher [DSL24, DWY^{+}25, KKS25, LXL25, SMDC23, TRL23, BBK21, CCF^{+}21, CCGRL21, CDPL24, Din21, FWLD22, HHL^{+}21, IZ20, KSS22, LY25, MJR20, NAE22, PS22, Pta21, SSD23, ZR23]. higher-dimensional [IZ20]. Higher-order [DSL24, DWY^{+}25, KKS25, TRL23, BBK21, CDPL24, Din21, FWLD22, KSS22, MJR20, NAE22, PS22, Pta21]. Highly [WX22, WYP23, WYT23, ZWT23, AHM20, ACXI24, EFFR22, GHS23, PFCZ23, TGARGT^{+}23]. Hilbert [Cos23, LSZ24b, TH21]. Hilliard [GLF25, JWL20, WHH22, ZFQ24, AMK25, CL22a, CMY23, FRW22, GGT24, JAMX24, KTY23, LKC^{+}24, LL25a, NR22, SXKL24, Tan25, WZY22, XFS23, YLZ20b, ZZW20, ZZ21, ZO21]. history [LXH24]. history-dependent [LXH24]. HIV [SSZ20]. HJB [RZ21, YTH^{+}21]. HM [SLUW^{+}21]. HOC [MKAM20]. Hodge [WWZX25]. HOFEIM [DRSY20]. Hohenberg [QH22, WCWZ23, YTK21]. Hölder [CP21]. holes [XLS24]. Homoclinic [JXK20]. homogeneous [ACLP22, BL20, BM23, CCGRL21, JMS^{+}24, MRPK21, PWWJ23, PWJ^{+}24, QC24, WXZW21]. Homogenization [LR22, PCM20]. Homotopy [HHL22, ACXI24, JICRMG^{+}20, Liu20, WX22, YX21]. homotopy-based [YX21]. HomPINNs [HHL22]. Hong [TGL21]. Hopf [CDM24, RLSC23, WSS22]. Hopf-type [CDM24]. horizon [KVV23]. horizontal [CBK24]. hot [BNQ21]. hp [CKBTS24, YPC^{+}21, APS25, CM23, CHD24, HHR24, Sau23, BFM24, BHS22, BHS25, LSW22, NAS20, SGR^{+}23]. hp}-adaptation [CHD24]. hp}-Crouzeix [Sau23]. hp}-FEM [BFM24, BHS22]. hp-finite [BHS25, SGR^{+}23]. hp-HDG [CKBTS24]. hp}-mesh [CM23]. hpq [MZ23b]. hpq-adaptive [MZ23b]. HSDT [SFR21, SS20]. Hu [CLL^{+}24]. Huber [GAM22]. Huxley [MKM24a, YL22]. HWENO [ANZZ23]. Hybrid [BBD21b, CKL25, KLK^{+}24, LS22b, LVDDNT25, QD20, TCA^{+}22b, WZTZ23, ZQY^{+}23, APPS25, AKK^{+}25, AAM^{+}24, BLLY24, BMVA23, BSRVC20, CCGF25, CXJ25, DK23, DK24, GKHU22, GGH25, Hag24, IBY^{+}24, MPVA25, MMP21, MKAM20, NSMR25, OAD24, SSD23, ST22, WX22, XFLZ22, Yem24, YFK23, ZCYL24, dR25a, AMV24]. hybrid-dimensional [CXJ25]. hybrid-numerical [SSD23]. Hybridizable [CLR23, CLRT24, EFFR22]. hybridized [FCFW21, HHW21, OAD24]. hydraulic [ANWW21]. hydro [LaZMY25]. hydro-dynamically [LaZMY25]. hydroacoustics [JHS24]. hydrodynamic [BAM^{+}21, LaZZ24, NS20, RWaZ23]. hydrodynamically [AZY22, WZY22]. hydrodynamically-coupled [AZY22]. hydrodynamics [ASH^{+}21, ABL24, BKB22, CGO22, LAB23, Qiu21, SHT^{+}24, YOF20, dSG20]. hydrogels [HYH23]. hydrogeological [BKR24]. hydromagnetic [OSO^{+}23]. hydromechanical [BMP21]. hydronic [ZCZ^{+}21]. hydrostatic [CZLQ22, FNGDV23, RZSL23]. hyperbolic [BAKR24, BCT21, BCM25, CBŁP23, CMY23, DG22a, DD20, DTW23, GCJ23, Haj21, HZ24a, HA23, Kap24, Kim21, KBB^{+}21, LLSY25, LE24, LGL20, LYLH25, LBPC20, MH25, MGA22, NRV22b, NNT22, Orl23, Pac24, PS25a, SK21, WS20a, ZL25]. hyperbolic-parabolic [BAKR24]. hyperelastic [BZDQ21]. hypersingular [SXCB24]. hypersonic [sCpLYhG23, WLX20]. hyperspherical [JICRMG^{+}20]. hyperthermia [AOL22, GOP25]. hypothesis [Bou24b]. hypoxic [LARGVR23]. hysteresis [VGA^{+}22].

IB [DKH22, WCZ24, WCZD20]. IB-LBM [DKH22]. ice [BS24, LYP^{+}22]. icing [BTCV20]. ideal [WSGT21]. identical [ZDW23]. identical-relation-preserving [ZDW23]. Identification [GM21b, IJS^{+}24, KMZ22, BDJ20, CFT^{+}23, HTDT23, KP22d, KFD^{+}20, Par22c, YEO21]. identifying [AHP20]. IDS [Ade20]. IEQ [CMY23, WL25]. IFISS [BRS21]. IGA [ŁWP^{+}23, EFK^{+}23]. IGA-ADS [ŁWP^{+}23]. IgA-BEM [EFK^{+}23]. IGG [KN24]. ignition [SB24b]. iGRM [LMMMP20, ŁMMMP21]. II [BGK23, BM24a, GP22a, GZG20, GGRBRG20b, HH22, JMR24, KPJO22, Mit22, OS24, WZC^{+}20b, WN24]. II. [LC22]. Image [KJL^{+}20, PMLC20, Ahm24, BML22, BH22, BKM23, CCH24, CBLZ24, FFC^{+}25, GP22c, HK20b, HJZ23, JAMX24, KJ22, LCN20, LHC23, eLLAN25, LQZ23, LP22c, MJRM20, MB21, NES23, NZ24, NR22, RGL21, RGL22, SHQL20, SY25, TRSM23, YGZ^{+}22, ZZ21, ZZ23b]. Image-based [KJL^{+}20, MB21]. images [CCH24, DH24, LP22b]. imaging [CEB24, EGF^{+}23, KPS23a, Par24a, PLVL21, ZWZD24]. IMEX [Che24, KTY23, LYZ23, MMGS25, SMR24b, TSS24, WLL23, WJZ^{+}24]. IMEX-BDF [WJZ^{+}24]. IMEX-split [TSS24]. IMHD [RC25]. Immersed [Adj25, BBS21, CLYH22, CRPdT20, Adj22, BZ22, BBS22, CHZ23, CXF24, DG22a, DVA^{+}20, ERG^{+}24, FBK^{+}24, GBK23, GXFT23, LpWW^{+}25, PK23, RLF21, SGAC^{+}20, SD21, SLX^{+}23, SPZL21, WWQ20, WXZW21, WDGS23, WCZ24, WXH^{+}25]. immersed-boundary [LpWW^{+}25]. immersogeometric [Kam21, SGF^{+}21]. immiscible [AEZ24, CBK24, GML^{+}20, HLH^{+}20b]. immobile [BS23a, CLL21, CLZW20, DC22, QXCG20, YLZ^{+}24]. impact [Ha20, JJ23, LH24]. impacting [WZXN21]. imparity [SB24a]. impedance [CEB24]. imperfect [MKM^{+}24b, SN20a, ZHHL23]. imperfections [LS20, ZD21]. impinging [YLY20]. implant [FGG^{+}20]. Implementation [PZZ23, RUMT^{+}24, SJM24, SLUW^{+}21, BFM24, BGH20, DST25, HHM21, HSJSS25, HGW21, LS20, MPVA25, SHWC22, TMML23, XL22, YPC^{+}21, YWL21, ZKBC21, ZBKC22, dSG20]. implementations [GLLZ23]. Implicit [FBK^{+}24, TW23, AAWJ20, BZ20, BZ22, BHPS24, BZ23, CCG^{+}23, CZ23, CFTC25, FZDS22, Hoa22, HDWL24, JMS^{+}24, JHG^{+}21, JZHZ22, KN24, LF24b, LYL^{+}22, LHTZ25, MWYY22, MGA22, MLL25, NBF^{+}24, PLZZ25, PCLZ25, QXCG20, QS20, RMBC24, SYZ23, SQ23, SQB22, TBS25, WYS^{+}20, WC21, WHH22, WCZD20, WPLW22, XJ21, YPC^{+}21, YQZT21, ZDLC22, ZYS22, ZLZ25b, ZX20, ZZB20]. Implicit-explicit [FBK^{+}24, MWYY22]. implicit-in-time [RMBC24]. implicit/explicit [CZ23]. implicitly [OR22]. implicitly-defined [OR22]. imposed [Med23]. Imposing [ST20a]. Improved [EFFR22, GPRS22, GLC24, JCJ24, KHD24, KHY25, Nag24, BWZ23, CYC^{+}23, EJWD23, Gun22, HJ24, HLH^{+}20a, HBC^{+}21, LXJ^{+}21, LYZ23, LZLN25, MRN^{+}20, MKM^{+}24b, OEASY24, OCRN22, SK23, YJZ20, ZLH22]. Improved-FSDT-based [KHD24]. Improvement [SQLZ21, XJ21, DNP^{+}22]. Improving [LCY23, ZvNH25, ZKBC21, ZBKC22]. improvised [LSRL24]. impulse [RNN22, YTH^{+}21]. impulsively [WCZD20]. in-compressible [JZXR24]. In-plane [LLZ^{+}25]. In-silico [SAVHB22]. inattention [YT22]. inclined [ACXI24, OSO^{+}23, WX22]. including [APGK24, DBR21, SP20, ZZ20b]. inclusion [ZWS22]. inclusions [Par22b]. incompatible [CWTM20]. incompressible [AHLZ24, ASH^{+}21, BdSC24, CCG^{+}23, CGO22, CCWS20, CW20, CSS25, DCCM23, DMX24, DSF24, Du24, FLS23, GS22, GPL21, GI21, GCC20, GRR20, GML^{+}20, GJS20, HQLJ20, JRu24, JWL20, KP22a, KHH25, KS25, Li24b, LDM25, LH23b, LLW^{+}23b, LSY^{+}24, LL24a, LCT22, MWYY22, MH22, MG25, MGBN22, MZHJ24, MCW22, PNT^{+}23, QS20, Qiu21, QD20, RC25, SD21, SCCS21, SHD23, SR21, SPZL21, ST20b, TA22, Vak25, WWQ20, WWW21, WYP23, WS21b, WYT23, XKXL22, XYS^{+}24, XYT^{+}21, YHZ22, YQZL22, YWZ^{+}23, ZR23, ZHZ^{+}22, ZS23a, ZQY^{+}23, ZYD24, ZZZ25, ZFQ24]. incorporated [YLY20]. increased [DKH22]. indefinite [WWC^{+}24]. independent [BV23, Bot20, LQLF23, RKRRL22]. index [CHJR21, FWBY25]. index-2 [CHJR21]. indicator [SBMS22]. indifference [LLZ20, YZL20]. indirect [NE22]. indoor [SADW21]. induced [BGG^{+}23, CRL^{+}25, HHR^{+}20, SA22, SNL23, WCZD20, XFX^{+}24]. induction [BGPS20, CRB24, LSV24]. inductionless [WMMH25, ZD22, ZZ24b, ZZZ25, ZST22, ZZS24]. Industrial [SGF^{+}21]. inequalities [AMN22, BMZ20, DD21, LQ24, MD21, RZ21, ZNC24]. inequality [BHS22, BFVY23, FYM23, HJZ23, LXH24, hWCsLjH21, XCLX24]. inertia [SGAC^{+}20]. inertial [CRPdT20, ZLQ23]. Inexact [ABC^{+}21, BFVY23, BMST20, BFM^{+}25, Che24]. inextensible [GRR20]. Inf [AJLM24, CS22, DFHT21, LSZ24b, Sau23]. inf-sup [CS22, LSZ24b]. infection [SSZ20]. inference [RBSC24, UHP23]. infinite [BLX^{+}24, DHW20, YTT^{+}22]. infinite-dimensional [YTT^{+}22]. infinitely [BBS21]. Inflow [KP21]. Influence [DP20, EAM20, BPV24, Bel24, SAVHB22, SGZ^{+}23]. influences [LY24]. influx [JJ23]. information [GP22c, HM24, Par22c, Par24a]. Informed [UBP^{+}25, CGGY23, CGHJ24, CLY25a, CGI^{+}22, CDG^{+}23, DSR23, FZW^{+}25, GY25, GYL25, GZAR23, HSR25, HHL22, LDW^{+}24, LWG^{+}25b, LWG^{+}25a, LCC^{+}24, MJZ25, NES23, PL24, QWLQ25, TFWX23, WLH24, WZL25, XYS^{+}24, ZL23a, ZL23b]. Ingard [KC20, KWK24]. inhibiting [KN20]. inhomogeneities [AHP20]. Inhomogeneous [BM24a, CKL25, DG22a, EBT24, LN22, RL20, RGS24, SCG23, WZL25, YL20a, YL22, ZSW23]. initial [CL22a, CKS22, DR25b, FC24, HZ21, HTBMC20, HYW^{+}24, JAHM23, LNNP24, LS20, PS22, PK25, QX23, SB25, SGZ^{+}23, TTKV20, ZQS24]. initial-boundary [HYW^{+}24, ZQS24]. injection [BNC^{+}23]. injectivity [YJZ20]. inkjet [ZYM24]. inkjet-printed [ZYM24]. inlet [RBSC24]. inner [CFT^{+}23, Hak20]. Innovative [BBD^{+}24, WXH^{+}25]. inpainting [HK20b, JAMX24, NZ24, NR22, SY25, ZZ21]. input [KP24]. inputs [BLLM25, CSB^{+}21]. inserts [MAT^{+}20]. Insight [ANH^{+}21]. Insights [QN20, ASF^{+}25, ABD^{+}21, Par24b, RNMÖX24]. inspired [JJOG20, LZB25]. instabilities [SMS23a, XTZ^{+}21]. instability [BPID20, RCSM23, SB21]. insulation [BKW22, WWY25]. insurer [XXY20]. integer [GPRS22, LH23a]. integer-order [LH23a]. integrability [Waz20]. integrable [KCD^{+}23]. Integral [DFS21, DCSAM25, KMS20, XZS^{+}23, ASB25, Bar22, CFP22, CM20a, DT24, FM25, GP22a, GPA21, Gut20, HCZ25, JKS25, JRu24, LX22, MZWD23, ND21, PP24, PS21, RL24, SXCB24, WS20b, vtWHG^{+}22, XLS24, YTT^{+}22]. integrals [HXZ^{+}22, NR25a, Yan25a]. integrated [AA21, AMK25, BW21, GdFFF25, NC20]. Integrating [WZL25]. Integration [LP20, AMM23, DST25, DVA^{+}20, DTW23, EFK^{+}23, FBK^{+}24, HLWL25, HBC^{+}21, HLN24, KWM22, MFF25, OCCM21, TW23, VJ22, WLD22, YQZL22, Yu23, ZWOM21, ZDLC22, ZLH22]. integrator [ZZS23, dSG20]. integrators [DE22, FHZ22, PVC^{+}25]. integro [AAL24, Che24, FAM25, LMLZ23, MH25, SMDC23, SB25, SZ21, TQX23, WYHQ24, Wan24, ZX20, ZQS24]. integro-differential [AAL24, Che24, FAM25, LMLZ23, MH25, TQX23, WYHQ24, Wan24, ZX20]. integro-parabolic [SMDC23]. integro-partial [SB25]. integrodifferential [QGQ22]. intelligent [IBY^{+}24]. intense [CYC^{+}23]. intensities [ADG23]. intensity [Che24, DCSAM25, KPO20, KPJO22]. Interaction [Li21, AL22, CCG^{+}23, CLW^{+}24, CYM25, GV21, GCSF21, HC23, ISA24, JJOG20, KWM22, Kam21, MSHB22, NSMR25, QhCZ^{+}20, RLVR21, SWC24, SHXZ24, Vak25, WR23, WS21b, ZBCS24, dCLW25]. interactions [AAC23, BTYM^{+}21, BNC^{+}23, LFF^{+}24, LCE23, YRC25]. interest [BMP^{+}24, LZ20d, LSPRV24]. Interface [KC21, TWH^{+}22, Adj22, Adj25, AP25, BHK^{+}24, BZ22, BDF24, BCF^{+}23, BGLZ24, CCW22, CS23b, CHZ23, CXF24, CCWC23, CSS23, DG22a, DD20, DDG23, DNP^{+}22, FHM21, GV21, GXFT23, GYC25, GO23, HHR^{+}20, HWC24, KP22b, KP22c, Koc22, KTY24, KK25, LS22a, Lee24, LQLF23, Li21, LpWW^{+}25, LLXY25, PK23, QSL23, QQY^{+}23, RS21, RCSM23, SD21, SWWZ23, TKK22, US20, WXZW21, WGC23, WXH^{+}25, WMMH25, WCCC25, WZA25, XZ25, XZ24, Zha22c, ZHHL23, ZGW23, ZSW23, ZCW24]. interface-debonded [AP25]. interface-penalty [Zha22c]. Interface-resolved [TWH^{+}22]. interfaces [Adj22, CRB24, KJ21, LF24a, LWL^{+}20, MCW25, NML23, SGZ^{+}23, XDL^{+}25]. Interfacial [CWTM20, WZXN21]. Interior [LG23, NAS20, BJyST24, EFFR22, LTXZ24, YSGD20, ZWZD24]. interlayer [Ha20, ZNC24]. internal [MZWD23, WDGS23, ZLYL20]. International [TGL21, ZSH^{+}21]. INTERNODES [VADG22]. interpolating [ADKH20, ZL20a]. Interpolation [XMQ^{+}21, AAD24, BNV22, BBS22, GPA21, HLX20, HDWL24, KKJ^{+}22, LFF^{+}24, LL20b, zLqLZ^{+}24, LZL^{+}21, LMLZ23, LD23, MJWD20, MWY23, NP24, Oru20, Oru22, PLR21, PS25a, RSAG^{+}24, TFWX23, WLD22, YCL20, ZP24, ZDLC22]. interpolation-based [BNV22, PLR21]. interpolation-free [MWY23]. interpolations [HLW24]. interpolatory [ST20a]. intersecting [AFV22, LWFS24]. Interval [BHC20]. intervals [SS21, SS24]. interventions [ASF^{+}25]. intravitreal [FGG^{+}20]. Introduction [JYHS23, KWK^{+}20]. intrusive [AS21, LZ25a, SDM21, SC21b, ZXGF22]. invariance [HHR^{+}20]. invariant [LTYZ25, ZSY22]. invariant-set-preserving [LTYZ25]. invariants [MN24]. invasion [AMC23, NCRGVR21]. Inverse [NRV22a, STO22, AHP20, AAA24, BDJ20, BFE^{+}21, CZ20, DHVT20, FC24, GGH25, GHL25, HKS21, HLW21, Ila20, JKC24, KP22d, KHM25, KLY25, LBTVG22, LXY23, LWG^{+}25a, Liu20, MC24, MH23, MSBO23, PN24, PQSM21, Par22c, RHA^{+}24, Tad23, TH25, TTKV20, TBO20, WQ24, YDY^{+}23, ZWZD24, Zha22e, ZXZ25, ZWY23]. inversion [BCK24, BFE^{+}21, LDT24, ZDWR22]. investigate [Bar22]. Investigating [LH24]. Investigation [EFS24, LP22c, MLJ^{+}20, Par22b, SB24a, WLZZ22, YLZL20, BMH22, DKV21, DTV20, JWW^{+}21b, RCSM23, SNL23, YZ21, YZJ^{+}22, YLQ22]. investment [PR20, XXY20]. inviscid [EH23]. involved [HHR22]. involving [BKM23, HSJSS25, KP20, MCW25, PV20, SB24a, SB25, SA22, SIAA25, WLZZ22, WL20]. ion [CZC^{+}25]. ionic [BHPS24, GBK23]. IPDGFE [LSH21]. irregular [FAM25, GP20, KWK^{+}20, LYRW21, LLJ23, LGL24, MXLY24, TGARGT^{+}23, XLLA20, ZLT^{+}21, ZSW23]. irreversibilities [AGA^{+}21]. irreversible [Zha22d]. irrotational [QD20]. IS-PINN [WZL25]. Isaacs [YT22]. isoAdvector [AFBM25]. Isogeometric [AZ24, BMST20, BGKV20, GP22a, HTPV24, LMMMP20, ŁMMMP21, Pup25, Xen20, XJY^{+}20, ZLYL20, ZYC^{+}24, BLM^{+}24, BTT20, CTGR24, CPS21, DVA^{+}20, FLMS24, HLEF20, KV22, Kra23, LLG25, LMST20, LST22, LBPC20, MHHR20, PDB22, SFR21, SYHN24, TT25, WM23, WNF^{+}25, ZJZ24, ZLF^{+}22, dlRGR20, dlRRG21, MST25]. isoparametric [LS21]. isospectral [Hak24b]. isothermal [Bal21, CZLQ22, CXX^{+}25, MCW22]. isotherms [MRN^{+}20]. Isotropic [HLCW24b, CCGRL21, GRR20, HWY23, MNB23, PM23]. isotropy [SCD25]. Issue [AGNP24, ZSH^{+}21]. iterate [JAHM23]. iteration [GLL20a, KM22, LLG25, LAL^{+}24, PS20, RZ21, SQ23, SW20b, WQ24, YL25, ZZ23b, ZZZ25]. iterations [dlRRG21]. Iterative [AMV24, GHH20, HHR24, WWLG24, YK24, ZFZ20, ZWS22, BLLM25, BM20, BBK24, CLLL25, CDLW20, DTV20, HYLH22, JWW21a, KSB25, KP20, KKNR21, LQLF23, LZ20a, LZW21a, LZW21b, NSMM21, PXLY23, Par22b, Tor20, UHL^{+}24, XL25a, ZHW20].

Jacobi [YT22, AS20, CLXZ25, HH23, HZR23, LLN24b, LZ22, LXP^{+}24, LZ25c, NSRI24]. Jacobian [YPC^{+}21]. Jacobian-free [YPC^{+}21]. January [Ano20-59, Ano21-60, Ano22-56, Ano22-49, Ano23-51, Ano23-50, Ano24-60, Ano24-63, Ano25-48, Ano25-50]. Jeffrey [LLJ23]. jet [BNQ21, YLY20]. jets [KP21]. Jimbo [Waz20]. jointed [BMP21]. joints [MR24]. Joseph [XZ25, QQY^{+}23]. Joule [SLW22]. Julia [HMX21]. Julia-based [HMX21]. July [Ano20-49, Ano21-63, Ano21-54, Ano22-67, Ano22-65, Ano23-60, Ano23-66, Ano24-61, Ano24-53, Ano25-53, Ano25-51]. jump [GM22, JZ24a, Lee24, LLXY23, TZ20, WXZW21]. jump-diffusion [GM22, JZ24a, TZ20]. jumps [Che24, CDGM25, MW20, RS21]. junctions [BPR22, PV22]. June [Ano20-63, Ano20-64, Ano21-56, Ano22-71, Ano23-59, Ano23-63, Ano24-58, Ano24-48, Ano25-55, Ano25-57]. juSFEM [HMX21].

Kac [CYKL23, MZWD23]. Kahan [ATBT20]. Kalman [AOL22, VGA^{+}22]. Kansa [YLL24]. Kármán [Hop23]. Kashiwara [Waz20]. Kaskade [GSW21]. Katchalsky [Krz22]. Kedem [Krz22]. Keller [AMPB^{+}22, GSRG21, HH21, HH22, LZ20e]. Kelvin [BGR23, BPID20, SWW22]. Kernel [XLYY25, APK^{+}25, AAL24, CLY25b, LX22, Lin24, LN22, MN24, NCCD24, RKRRL22, SDM21, STM21, TH25, XJ21, ZGL23, ZX20]. Kernel-based [XLYY25, CLY25b, STM21]. kernel-smoothing [XJ21]. kernels [BTYM^{+}21, LDH24, LMLZ23, MKM24a]. Kerr [CGK24, LYLH25]. Kerr-type [CGK24]. key [GOP25]. kind [BHS22, JKS25, LX22, XL25a]. kinds [WYH20]. kinematics [SGAC^{+}20]. kinetic [CMS23, LZF23, LLZ^{+}23, MZZ^{+}25, PCLZ25, SK23, SCCS21, TWH^{+}22, XYT^{+}21, XSL24b]. kinetics [MWY21, XZSY23]. Kirchhoff [AN20, CHL22, KC23b, LZJC21, PG25, Par24b, TPOT20, WL20]. Klein [APPS25, AJ23, CC20, FZH22, GZ23, GFXZ25, JZ20, SZ22a, WL22, WL25, YHF24, ZW25]. knot [LZLN25, MHHR20, YWLF21]. knowledge [GNB24, LNNP24]. known [HJ25]. Kolmogorov [DL25, ADKH20, PZL23, SLD25, SJZ22, SWZ23]. Kong [TGL21].

Korteweg [AJK23, MRN^{+}20]. KPIK [ZS25]. KPS [ZZB20]. Kriging [WWZ^{+}24]. Kronecker [AAU25, CBŁP23]. Krylov [ASB25, JRZ21, NW24a, TS24, YPC^{+}21]. KSPHPDDM [JRZ21]. Kuramoto [DS25]. Kurchatov [YSB^{+}24]. Kutta [ADB^{+}22, BZ23, CZLQ22, FX22, KTY23, MLL25, SQ23, WWZQ25].

L [SPC^{+}23]. L-shaped [SPC^{+}23]. L1 [HYW^{+}24, JJZ21, LWX21, LXY23, YLZ^{+}24]. L1-finite [HYW^{+}24]. L2 [Tan24, ZGL23]. L2- [Tan24]. L2-type [ZGL23]. labeled [LWG^{+}25a]. labyrinth [SMR^{+}24a]. laden [PWJ^{+}24]. lag [BDS24, LYRW21]. Lagrange [LLD23, LWFS24, ZNH24, dCLW25]. Lagrangian [BKK^{+}23, GNZ24, HWL^{+}20, LZG^{+}23, LQF25, RBSC24, WCH20, YL20b, ZXXX23, ZZY^{+}24, ZWX24]. Laguerre [YW24]. Lamb [AP25]. Lamé [GI21]. lamina [LP22a]. laminar [EFS24, KWM22, SD21, WZC^{+}20a, WZC^{+}20b, WS21b]. laminated [Ha20, KKJ^{+}22, LJY^{+}25, MNB23, SP20, SS20, WWZ^{+}24, YLZL20, YLFL20]. Landau [HZA22, HYH23, LSW20, LL21b, NR25b, SRD25, WLL23, XSL24a, XP22, ZLPR20, ZZS23]. Landweber [WQ24]. Laplace [AZ24, ELS24, EMSA25, LFY^{+}21, MM21, MH23, STO22]. Laplacian [DZ21, BFM24, BKM23, HMF24, HCZ25, HGW18, HSJSS25, ND21, WWZX25, XLCP20, YPZ^{+}21, ZGG^{+}25, ZZ24c]. Larché [SBNR23]. Large [MKS^{+}21, SGF^{+}21, AKA^{+}21, AFS24, ATBT20, BMVA23, EBT24, FH21, GHRD20, KM22, LQ20, LGQZ21, LXJ^{+}21, LFGZ23, LQW22, LR22, Ma25a, PWC^{+}23, PGR23, PZZ23, QD20, SQB22, WWC^{+}24, XTZZ24, YX21, YWLF21, ZTW22, ZPC20, ZWW^{+}23]. large-scale [BMVA23, LQ20, LGQZ21, LFGZ23, LQW22, PWC^{+}23, YWLF21]. larger [KMS20]. laser [CRL^{+}25]. laser-induced [CRL^{+}25]. lateral [AHZ^{+}25, FKR21]. Lattice [ABLS23, CRPdT20, GBK23, GCSF21, KRESA21, KTFI23, KWK^{+}20, LMK^{+}21, PWJ^{+}24, QhCZ^{+}20, VL21, YOF20, ZDLL22, APGK24, ABMS20, ADL^{+}21, ABS23, BLLY24, Bel24, BP20, BABS20, CLLL20, CCWS20, CYY^{+}21, CYC^{+}23, DT23, DLT20, EFF21, EGF^{+}23, EFS24, ENF24, FGK20, FS21b, GML^{+}20, GDŁW23, HHR^{+}20, HLH^{+}20b, HHY^{+}22, HQCW22, HMŁW^{+}24, HDWL24, HKL^{+}23, JHS24, KGB24, KFD^{+}20, KKA^{+}21, Kri20, LLX23, LLX24, LpWW^{+}25, zLqLZ^{+}24, LZB25, LHQW25, MWYY22, MIOK21, MLJ^{+}20, NIT^{+}20, PGK20, PLR21, PWWJ23, RST^{+}22, RLF21, RLSC23, SGJJ20, SB24a, SMŁWL25, SK23, SCD25, SLX^{+}23, ST20b, TMML23, VJ24, WWQ20, WZC^{+}20a, WZC^{+}20b, WZY23, WCZ24, WWP^{+}25, WZCS20, WCY^{+}24, XL22, XYT^{+}21, XLZ^{+}21, XHH^{+}23, XTZZ24, YYY^{+}22, YCWS20, YWZ^{+}23, ZKTK21, ZCY22, ZCYL24, ZTW22, ZHHL23, ZWCS20, ZZ20b, ZYS22, ZLD^{+}23, HLH^{+}20a, HEDT20, Ima20, KOMK20, OF20, RGS^{+}20]. Lattice-Boltzmann [YOF20, LZB25]. lattices [ZZZ24]. lava [GdFFF25]. law [BP20, BS23c, DKMS22, PGR23]. laws [BCT21, BDGS25, DTW23, GCJ23, Haj21, KG20a, Kim21, KBB^{+}21, KVV23, Pel23, SK21]. Lax [LQ20, PCLZ25]. Layer [HM24, BM24b, CGHJ24, CMR22, Hoa21, KARS23, KTFI23, LL20a, LCC22, LC22, MZ23b, QD23, QhCZ^{+}20, RGS24, WWY25, YLQ22, YCL20, aKST22]. Layer-adapted [HM24, CMR22]. layer-wise [WWY25]. layered [AI25, ASZ24, BDB21, BDL20]. layers [BRRT21, CBK24, ZZ24a]. LB [LQYJ23, WCZD20]. LBM [DKH22, JYHS23, MJR20, SADW21]. LDA [CSI23]. leading [AAG22]. leap [BNSZ24, YWZ21]. leap-frog [BNSZ24, YWZ21]. Leapfrog [ZFQ24, CGK24]. Learning [BMP^{+}24, KC22b, ACLP24, BMvdZ21, BGD22, CAD24, CGGY23, CFT^{+}23, CGI^{+}22, FZW^{+}25, GHSR23, GOP25, GZAR23, HHL22, IBY^{+}24, KHY25, LLR25, NNH^{+}25, NFFR^{+}24, OP24, PN24, RHB25, SSK^{+}23, TH25, UHP23, VT21, WTSZ24, WWLG24, YKNS23, ZWS22, ZWY23]. learning-based [NNH^{+}25, NFFR^{+}24]. Least [BB23b, BB21, BDG21, BS24, LZ25b, UBP^{+}25, XYS^{+}24, AN21, BM24a, BBG23, Bri24, CCY24, CM21, CLY25b, FP20, FK21, JBKW23, KC21, Ku24, LZ25a, LCC22, LD23, Ma25a, MHKB25, Pac24, QC24, QFL20, SV21, STM21, SW24a, Sto20, WW20b, WCZ24, Zha23]. Least-square [XYS^{+}24, KSD^{+}25]. Least-Squares [BDG21, BB21, BS24, LZ25b, BBG23, Bri24, CCY24, CM21, CLY25b, FP20, FK21, KC21, Ku24, LZ25a, LCC22, MHKB25, Pac24, STM21, SW24a, Sto20, WW20b, Zha23]. LED [DKV21]. Legendre [Der23, FKW22, JM23a, JJZ21, KJ20, KKJ^{+}22, LYL^{+}22, NML23, QC24, SJM24, WFH^{+}20, YM23, YW24, ZDWR22]. Legendre-Laguerre [YW24]. Legendre-radial [KKJ^{+}22]. Legendre-tau [NML23]. length [BD20, NTP25, SMS23a]. Lengyel [AHOB24]. lens [LP22b]. LEoPart [MRS21]. less [LCT22]. Level [FMP^{+}24, NIT^{+}20, BS23b, CCWC23, CV20, DLC22, GPY22, GS23, HLL24, HLN24, HLW21, JMCC24, KCW22, LFGZ23, LQZ25, LL24a, MKAM20, NC20, NDC22, NNNT22, PLVL21, Qiu21, UuIUK21, YA23, YPZ^{+}21, Zdu21, ZST22, dlRRG21]. Level-set [NIT^{+}20, NNNT22]. levels [BD23, DKH22]. leveraging [XRS25]. Lévy [CWC22, DYZ22]. library [AAB^{+}21, ABD^{+}21, KWK^{+}20, MRS21]. lid [ACXI24, HBC^{+}21, KC23a, WX22, ZZZC24]. lid-driven [ACXI24, HBC^{+}21, KC23a, WX22]. Lie [WHT22]. Lifshitz [NR25b, HZA22]. lifting [WWL22]. light [GPL21]. like [CT22, GS22, LQ24, Loh21, NBF^{+}24]. Limit [YTA23, ABL24, AG25]. limitation [Zha22e]. limited [AHP20, KP22d, Par22c]. limited-aperture [AHP20, KP22d]. limited-view [Par22c]. limiter [LHTZ25, ZQZ20]. limiters [DT23]. Limiting [LLW20, Haj21, SK21]. line [SPZL21, jWC23]. line-based [SPZL21]. Linear [JW20, KG20b, KTY23, LZ25a, LJ24, RGS24, ZCY22, ZCYL24, ADKH20, Adj25, AHLZ24, Alt23, ABFPA24, ADG23, AH25, BTCV20, BB21, BZ23, BHC20, BFM^{+}25, BWMW23, CMPR23, CBŁP23, CSS20, CKFR24, Cha24, CC20, CXXZ23, CPS21, CCGRL21, CL22b, CLL^{+}24, CP21, ÇM20b, DGQT20, DMX24, EP23, Füh21, FF24, GLÖ21, GENRB21, GB21, GB23, HC21, HZJ^{+}22, HLWL25, HWL^{+}20, HLQ24, HV21, HM20, HRvdZ20, HHR24, HJXW23, HWWZ24, HA23, JS25, JL19, JWL20, JL20, JCP^{+}25, KM22, KP24, KPJ22, LZ20c, LL25c, LZW21a, LZW21b, LDH24, LXZZ24, LYL25, MGD23a, MZM21, NNH^{+}25, NR25b, Oru22, OF20, Pac24, PS25a, PD21, PK25, QH22, RL24, RC25, SD25a, SM22a, SS22, SL21, TLZF20, TKK22, UR24, WHT22, WCS24, WLW^{+}25, WZ25, WLD22, Wie23, XSZ^{+}23, YDY^{+}23, YMDW25, YM21, Yu23, Yua23, ZR23, ZHYL21, ZC22]. linear [ZFL22, ZZL22, ZH23b, ZJ25, ZL25, Zha22d, Zha24, ZXQQ22, ZMZW24]. linear-cost [CBŁP23]. Linearity [CSI23, DWS20, LD23, MKP23]. linearity-preserving [DWS20, LD23]. Linearized [ZLPR20, BZDQ21, CL20, CSS25, GWN21, HMFM25, LSW20, LL21b, LL22a, Li24b, LZY24a, LQY22, MLLL24, NS20, NK25, SZ22a, WLC24, WYW^{+}24, YWL23, Yan25b, ZW25]. linearly [LF24b, PHL21, WPLW22, ZLZ25b]. linearly-implicit [LF24b]. lines [HDG23]. Lipschitz [GQZ23, YY25]. liquid [AAA25, AAWJ20, BPID20, HHY^{+}22, HJTW25, LpWW^{+}25, SMŁWL25, SRP21, TWH^{+}22, WTSZ20, ZDLL22, ZLD^{+}23]. liquid-crystalline [HJTW25]. liquid-vapor-solid [HHY^{+}22]. liquidity [SMR24b]. liquids [LCT21]. lithium [CZC^{+}25]. lithium-ion [CZC^{+}25]. living [CM24]. LL* [Kei21]. load [JMCC24]. load-balanced [JMCC24]. loads [SS20, YLZL20]. Local [HLX20, HYW^{+}24, HLW24, LD22, MN23a, ND21, PS25a, SW24b, YWLF21, ZAH23, ZS21, ZZS24, AD20, ANWW21, AGA^{+}21, BS23a, BMH22, BZY20, BZ23, BMST20, BEKV21, CMR22, DZZ^{+}24, ERHL22, EED^{+}25, GHAU24, Hoa22, HHR22, KCW22, LE24, LDH24, MBG24, MLJ^{+}20, Oru20, PG25, PZZ23, QP24, SB25, SXKL24, WDZ21, WZ24, XMQ^{+}21, XHL21, YWL21, ZSK23, ZZ25a, ZGL23, Zhe22, dlRRG^{+}24]. localised [LE24]. localization [Par22b]. Localized [LQW22, Abd24, CCF^{+}21, CFT^{+}23, JZXR24, LFY^{+}21, QFL20, SLD25, WCH20, WWP^{+}25, XFX22]. locally [FKKV24, GDA21, JP25, LCY23, RDP20, ZW21b]. location [SADW21]. locking [CXXZ23, HWL^{+}20, HFG23, HZ24c, HV21, HWWZ24, LR23, TLZF20, YZZ20, Yu23, ZZCM20]. locking-free [CXXZ23, HWL^{+}20, HFG23, HV21, HWWZ24, LR23, TLZF20, YZZ20, Yu23]. logarithmic [sCpLYhG23, DZ21, HGW18, HSJSS25, YKK25, YLFZ25]. logistic [HH21, HH22, LZ20e]. lognormal [BN20]. long [AJK23, HD21, JCJ24, MKS^{+}21, Rav20, WLC24, WXCL25, YT22]. long-run [YT22]. long-term [WXCL25]. long-time [JCJ24, Rav20]. long-wave [AJK23]. longitudinal [Hag24, SB24a]. Looking [LP22b]. loop [CQG^{+}24, RGS^{+}20, ZL23a]. loosely [GV21]. loosely-coupled [GV21]. Lorentz [ZAPK23]. loss [GYL25, LCC^{+}24]. lossy [KJ24]. Lotka [TSW^{+}25]. Low [WLJF25, ZZ20a, ZMS21, ZMZW24, CY23, CWW22, CCH24, CL21, CRPdT20, GHH20, HK24, LM24, LLD23, MB24, MST25, RGL22, TLZF20, UHP23, ZZL22, Zhe22, SBY^{+}25]. low-contrast [CCH24]. Low-cost [WLJF25, LM24]. low-dimensional [HK24, MB24]. Low-order [ZMS21, TLZF20]. low-quality [UHP23]. Low-rank [ZZ20a, CL21, GHH20, MST25, ZZL22, Zhe22]. low-Re [CRPdT20]. Lower [GSRG21, CL20, Wol20, ZY21a]. Lowest [BBM25, ACPT25, BLMV24, DWS20, GP25, GQ23, LSF25, WGC23, Yu23]. lowest-order [ACPT25, DWS20, GP25, GQ23, LSF25, WGC23, Yu23]. LQR [YY20a]. LRBF [BBT25]. LSPIA [LLG25]. lubrication [HLH24]. luminaire [DKV21]. Lump [WTSZ20]. lump-stripe [WTSZ20]. lumped [ZH23c]. lying [CRL^{+}25].

M [BW21, CLLG23, LFY^{+}20]. MAC [He23b, WR23]. Maccari [JXK20]. Mach [CYC^{+}23, JZHZ22, JYHS23]. Machine [NFFR^{+}24, VT21, BMvdZ21, CFT^{+}23, FZW^{+}25, GOP25, IBY^{+}24, NNH^{+}25, PN24, RHB25, SSK^{+}23, YKNS23]. machine-learning [BMvdZ21, RHB25]. machines [BGPS20, LLR25]. macro [DDG23]. macroelement [LL21a]. macromolecular [HYH23]. macroscopic [QP24, TKT^{+}24, VT21]. magnesium [ZYQ^{+}24]. magnetic [ACXI24, AHP20, AMPB^{+}22, AIG^{+}21, AAWJ20, BPV24, BLX^{+}24, BHD^{+}21, EED^{+}25, EGF^{+}23, GO20, GOP25, HKL^{+}23, LSV21, LSV24, LY24, MAT^{+}21, NML23, SRA20, TMML23, WW24, YLZL20, ZS23a]. magnetic-electric [ZS23a]. magnetics [DD24]. magnetized [VJH23, WXH^{+}25]. Magneto [AAC23, CGO22, EJWD23, HTPV24, LLMC25, OAA^{+}22, Qiu21, Rav20, SDJ^{+}22, YLZL20, YLFL20]. magneto-convective [OAA^{+}22]. magneto-electro-elastic [EJWD23, HTPV24, LLMC25, YLZL20, YLFL20]. magneto-hydrodynamics [CGO22, Qiu21]. magneto-micropolar [Rav20]. magneto-thermo-elasticity [SDJ^{+}22]. Magneto-thermoelastic [AAC23]. magnetoelastic [NR25b]. magnetohydrodynamic [DSF24, HJ24, KK23a, Li24b, LLJ23, Liu25b, MWRA22, PL24, TA22, ZD22, ZST22, ZZS24]. Magnetohydrodynamics [QS20, GJS20, LL22a, LDM25, LH23b, MH22, SHD23]. magnetonanofluid [PM23]. magnitude [CS25]. Mahony [CLQL25, KJ20, LSRL24, GHAU24]. Malik [LHC23]. management [YYY^{+}20, YTH^{+}21, YT22, YTA23]. manifold [JZ25, ZZL23]. manufactured [KJL^{+}20]. manufacturing [LAB23]. map [LLWX25, ZZZ24]. mapped [LZ22, TL22]. mapping [OTC24]. MAPS [WCL20, XCFC21]. March [Ano20-51, Ano20-53, Ano21-67, Ano21-49, Ano22-66, Ano22-54, Ano23-55, Ano23-58, Ano24-55, Ano24-62, Ano25-63, Ano25-59]. marching [WYP23, WJZ^{+}24]. Market [LSPRV24, LLZ20]. marking [CM21]. Markov [MSBO23]. Markovian [LCX^{+}20, ZW20b]. marks [YLQ22]. martial [HKJ23]. martingale [WS20b]. Mass [ZSH^{+}21, ZDW23, AGA^{+}21, FXLS25, HANA24, HFG23, JSBZ22, JSZ24, KP22a, KMZ22, Lee20, LST22, OAD24, PH23, WMMH25, ZH23c]. mass-conservative [Lee20]. mass-emitting [PH23]. massive [BNC^{+}23]. massively [CPBF21, Zdu21]. massively-parallel [CPBF21]. matched [YCL20]. matching [BCF^{+}23, DDFZ23, Fur25, LFF^{+}24, ZLF^{+}22]. matching-based [DDFZ23]. material [AP25, BNC^{+}23, BHDV22, FSH^{+}23, JFS14, JFS20, KJL^{+}20, LZL24, Ma25a, NTP25, NNTCB25, NML23, SHXZ24, WR23]. Materials [JMM20, DSL24, EP23, FL22, FYXL21, HWY23, KHD24, MKM^{+}24b, MB21, Oru25, Pta21, SHR25, YSGD20, YMDW25, ZRL^{+}22, ZZL23]. Math [DZ21, JL20, JFS20, LZW21a, MMAK21].

Mathematical [AM20, AO24, ACDL24, ASSV20, BZDQ21, BDE^{+}24, BGLZ24, CM24, CMS23, FGG^{+}20, FJP22, SB21, SRP21, JJ23, LWZ22, zLqLZ^{+}24, LARGVR23, SSZ20, WZL25, dABCC25]. MATLAB [MPVA25, RHR^{+}21]. MATLAB/GNU [RHR^{+}21]. matrices [AFM23, BFRA21, JMM20, LDT24, MM21, ZZ20a]. Matrix [DSS20, AHM20, ATBT20, Axe20, BFI^{+}23, CL24a, KM22, KCC20, LWL21, LST22, Ngo25, SGSD20, Tad23, TH23, TLJ24, UHLZ24, WXZW21, WT25, YA21]. matrix-free [BFI^{+}23, SGSD20, YA21]. Matrix-oriented [DSS20]. matrix-valued [TLJ24, WT25]. maximization [HNV25]. Maximizing [XXY20]. Maximum [CC20, HKJ23, SY25, HZ25, IKST21, WXF22, WWZQ25, YKK25, YYSL20, ZWMF24, ZY23]. Maxwell [BWR21, BWZ22, BNSZ24, BLX^{+}24, BDB21, BBG23, BD23, CLW^{+}24, CGK24, CZK^{+}25, HGD21, HGD25, HZA22, Li24a, Liu25b, ŁWP^{+}23, NML23, WJC^{+}25, XL25b, YLZ20a, YWZ21, YH21b, ZLHC24, ZSW23, dVDMM22]. May [Ano20-60, Ano20-62, Ano21-57, Ano21-55, Ano22-68, Ano22-61, Ano23-53, Ano23-49, Ano24-54, Ano24-51, Ano25-47, Ano25-49, LQCM20]. MBP [HLJ25]. MBP-preserving [HLJ25]. MC [LLY25]. MC-CDNNs [LLY25]. McKean [MW20]. ME [SQLZ21]. ME-AUSMPW [SQLZ21]. Mean [WDGC20, YY25, HJT25, YTY24, ZLQ23]. Mean-square [YY25, ZLQ23]. means [JM23b]. measurement [BD20, Li25]. measurements [Ngu20]. measures [DDFZ23, FLOQ21]. mechanical [AA21, BBDM21, LML23, NDB^{+}24, SCD23, WLZZ22, YSGD20, YLZL20]. mechanically [SCD23]. mechanics [BM21, CFG^{+}23, TRSM23, VADG22, VSG21, WZS^{+}24b, ZY21a]. mechanism [LZL^{+}25, ZW25]. mechanisms [ZWW^{+}23].

Media [RSB^{+}21, AZL20, ANWW21, Ale23, ÁCS22, AEZ24, AFV22, AAA24, BB24, BDB21, BBDM21, BLM^{+}24, CGK24, CQS25, DG22a, GBK23, HANA24, Hoa21, HLZ24, JWW21a, JSZ24, KRESA21, KJ24, KGW^{+}21, KS25, LHC^{+}22, LWFS24, MLJ^{+}20, MBSEA24, MS20b, NS20, PFCZ23, SSY22, SYZ23, SLX^{+}23, SNL23, SWZC23, TPA24, VT21, WYH20, WBY23, vtWHG^{+}22, WW25, YL20a, YH21b, ZLL23, ZAPK23]. medicine [YMTS20]. medium [AGA^{+}21, BWR21, JSBZ22, NS20, OAA^{+}22, OTC24, QC24, RL20]. mediums [JMS^{+}24]. MEE [LLMC25]. melt [LAB23, RST^{+}22]. Melting [AAG22, CRB24, BKT^{+}23, GZAR23, JFS14, JFS20, ZYM24]. melts [WZY22]. membrane [KMZ22, Krz22, LaZMY25, PGR23]. membranes [AFS24]. memory [Ade20, BGH20, Fia21, JJ23, KHM25, KPJ22, Vab22, WXCL25]. memory-based [JJ23]. memristive [AHZ^{+}25]. Merton [GM22]. mesenchymal [ZDW^{+}20]. Mesh [CM22, VJH23, Alf24, BF21, Bot20, CRM22, CM23, DBR21, EAM20, GP22c, Gie23, HM24, JWR^{+}23, Kra23, KC23b, Li21, zLqLZ^{+}24, LSF25, MGD23a, MZ23a, NC20, NCCD24, RR22, RKK25, SMDC23, SHR25, SZ24a, SBMS22, WML21, WWS23, WF25, WHC24, WRMB23, ZZ24a, ZLL20, ZZ25b]. Mesh-free [CM22, JWR^{+}23, NCCD24]. mesh-independent [Bot20]. meshes [AN20, AFM23, ABL24, AZ24, ABP25, AN21, BDMX21, BCT21, BHL24, BRRT21, CWTM20, CLZ23b, CMR22, DV22, DD24, DDG23, DD22, FZS22, FGK20, FMP^{+}24, GMS23, GKK^{+}20, GQZ23, GYC25, HMF24, HWL^{+}20, HM24, HYW^{+}24, HRS20, Ima20, JZPG21, KN24, KCW22, KGJN24, KSJ23, KTY24, LFF^{+}24, LS22a, LZY24a, LHF24, LSWY21, LMLZ23, LMZ23, LZG^{+}23, LGL24, LSY^{+}24, LD23, MWY23, Nie23, PK23, PLR21, QQX21, SGF^{+}21, Sch24, SU23, ST22, SNHH21, TZ25, Top22, TKK22, WM23, WKC21, WLQ22, WW25, XLSF25, YH21a, YZZ20, ZP24, ZQM20, ZC23, ZZ23a, ZPK22, ZW20a, ZXQQ22, ZSY22, ZS24, dVDR20]. Meshfree [LLXY23, LLXY25, AADMA23, Abd25, HTPV25, LNF24, MXLY24, MJWD20, MNB23, Nag24, PS25a, SS20, WT25, XMQ^{+}21, YLW^{+}22]. Meshkov [RCSM23]. Meshless [HHR22, HHG20, LL23, MCZL20, ZSK23, AD20, BS23a, BMH22, BBT25, BDGS25, BPID20, CKL25, DMŠ22, DC22, EBT24, GM21b, HC21, HCZ25, HHNV23, JM23b, Jaw21, LYRW21, LC21, LP22c, MB21, Oru20, QSL23, QL24, SAVHB22, TO20, TGARGT^{+}23, TGDMGT^{+}25, XSZ^{+}23, XLLR25, ZCXG20, RHA^{+}25]. Mesoscopic [JKL20, ZYM24, EFS24, LMP23]. mesostructures [ERHL22]. meta [PSDP22]. meta-procedure [PSDP22]. metabolism [YD23]. metaheuristic [GM21b]. metal [BNQ21]. metallic [XLS24]. metallization [ZYM24]. metamaterial [FLW20]. metamaterials [HB25, Li24a, LYLH25, WWY25, YWZ21]. metastasis [HW25]. Method [BV21, BBS21, CO21, DFS21, HMX21, KC21, KMS20, LCT21, LMMMP20, PV22, SN20a, XMQ^{+}21, XGL21, YA21, AD20, ADKH20, AADMA23, APK^{+}25, AS21, Abd25, ABC^{+}21, AHM20, AI25, APPS25, AN20, AMV25, AAU25, ASB25, AHP20, ACLP22, ASH^{+}21, ADFP23, ÁCS22, ACPT25, ALOR25, ABMS20, AZHY22, ABLS23, ABS23, AJK23, AMV20, ABV23, AMV24, AM22, AGPP22, ANN25, AAM^{+}24, AFBM25, AH25, BLLY24, BLLM25, BMH22, BWR21, BWZ23, BNSZ24, BGR23, BS23b, BZ20, BZ22, Bar22, BAKR24, BDB21, BFVY23, BH22, BMK21, BTCV20, BDP23, BGPS20, BM24a, BBM25, BB23b, BMPP22, BS24, BDF24, BNC^{+}23, BP20, BKT^{+}23, BKK^{+}23, BM21, BHDV22, BMST20, BS23c, Bra20b, BLM^{+}24, BC20b, CCG^{+}23, CLLL25, CFG^{+}25, CBŁP23, CL24a, CLLL20, CHY21, CCW22, CWW22, CLZ23a, CCGI23, CGM^{+}20, CLRT24, CGHJ24, CCWS20, CDLW20]. method [CC20, CS20a, CLZW20, CYY^{+}21, CLH21, CLYH22, CWC22, CHL22, CLH22, CZLQ22, CLZ23b, CXXZ23, CS23a, CXF24, CGK24, CCD25, CKL25, CZK^{+}25, CLQL25, CXJ25, CC25, CZ20, CMR22, CWLQ22, CHJR21, CYKL23, CKS22, CFT^{+}23, CDGM25, CV20, CS25, CCGRL21, CWX22, DN22, DLN24, DFSV22, DV22, DT23, DG22a, DG22b, DGK20, DLC22, DZZ^{+}24, Der23, DP25, DD22, DFK22, DHVT20, DS22, DBR21, DWS20, DS25, DW20, DC22, DSF24, DZ24, Du24, DNP^{+}22, EGF^{+}23, EFS24, EBT24, FSS^{+}20, FKW22, FSW20, FWLD22, FZS22, FKR21, FHM22, FWBY25, FK20, FLMS24, FYXL21, FS21b, FH21, FHH23, FF24, GLL^{+}20b, GLL20a, GM21a, GQ23, GP23, GHRD20, GOV20, GORBS21, GPY22, GHH20, GKHU22, GDA21, GNB24, GMS23, GHLS25, GJM22, GGH25, GPA21, GENRB21, Gra21, GDŁW23, GHL24, GJS20, GQZ23, GS23, GAM24, Gun22, GB21]. method [GCMS21, GWYR22, GB23, GXFT23, GHL25, GYC25, HC21, HC23, HMW21, HZJ^{+}22, HJP23, HXW23, HJ24, HDM24, HL22, HLL24, HDG23, HLEF20, HSR^{+}23, HLH^{+}20a, HLX20, HQLJ20, HWW21, HQCW22, HFG23, HZ24b, HZ24a, HZ24c, HLQ24, HOY24, HHNV23, HH23, HMŁW^{+}24, HLCW24a, Hop23, HHR22, HBC^{+}21, HHW21, HYLH22, HYW^{+}24, HRvdZ20, HHL^{+}21, HLPW22, HS20, HLW21, HC22b, HXZ^{+}22, HLY22, HAC23, HWC24, HYL^{+}25, HTPV25, HWWZ24, HHG20, HA23, Ila20, Ima20, JVBM23, JM20a, JM23a, JRu24, JM23b, JSSS24, JZ20, JJZ21, JZ25, JZXR24, JL23, JHS24, Jog25, JMN22, KSS22, KPK25, KPS23a, Kap24, KEUY23, KBR20, KNBR22, KRESA21, KM22, KW20, Kim21, KJN21, KLK^{+}24, KGJN24, KMN25, KCD^{+}23, KTFI23, Koc22, KBW23, Kri20, KMP20, KSJ23, KTY24, KGK25, KS25, KC23b, KOMK20, KKJ^{+}22, LLSY24, LNT22, LLN24b, LNNP24, LK20, LP21, LM24, LKC^{+}24, LLR25, LL20b]. method [LZ20b, LD21, LWL21, LL21b, LSH21, LWSZ21, LLZ21, LL22b, LXRQ22, LCW22, LMD22, LXZ23, LLX23, LQYJ23, LZA23, LZBC24, LD24, LZL24, LaZZ24, LLX24, LZY24a, LLZL24, LWG^{+}25b, LY25, LLG25, LaZMY25, Li25, LL25c, LLZ^{+}25, LDM25, LQZ25, LS20, LX22, LL22c, LR23, zLqLZ^{+}24, LYH24, LSF25, LZW21a, LZW21b, LYRW21, LQW22, Lin24, LSRL24, LS25b, Liu20, LD20, LS21, LY21, LFY^{+}21, LN21, LZL^{+}21, LSWY21, LYL^{+}22, LQY22, LSGN23, LCH23, LGZC23, LLJ23, LS23, LZ23, LLW^{+}23b, LZG^{+}23, LHSF24, LGL24, LSY^{+}24, LWFS24, LXZZ24, LZY24b, LSZ24a, Liu25b, LZLN25, LD25, LQF25, LJY^{+}25, LYLH25, LP20, LP22c, LYXC21, LCT22, LN22, LZ25c, LHQW25, MCZL20, MWYY22, MCD23, MZWD23, MFZX23, MGD23a, MZ23a, MCD24, nMjWlS24, Ma25a, Ma25b, MCW25, MIOK21, MGBN22, MBG24, MGD23b, MRN^{+}20, Mas23, MS20a, Med23, MHKB25, MJWD20, Men22, MT25, MZM21, MLLL24, MH25, MN23a, MKP23, MR22, MGA22, MS22]. method [MFF25, MNB23, MS20b, Nag24, NR25a, NM23a, NZ24, Nde25, NIT^{+}20, NNNT22, NNH^{+}25, ND21, NF25, Nie23, NML23, NK25, OTC24, OXY21, OP24, Oru22, Oru25, OAH^{+}24, OSZ22, PG21, PS20, PCC21, PWC^{+}23, PWBS25, PKP21, PK23, PGK20, PZL23, Pel23, PLR21, PS25a, PWWJ23, PL25, PSSW20, PDB22, PK25, PZZ23, QSL23, QC24, QL24, QXCG20, QS20, QQY^{+}23, QFL20, QD20, RHA^{+}24, RZC21, RDF24, RNN22, RBSC24, RZSL23, RHA^{+}25, RS21, RZŠ21, RWaZ23, RL23, RST^{+}22, Rou20, RKC25, RR23, RHA21, RGS^{+}20, SC21a, SQC23, SLD25, SM22a, SDJ^{+}22, SAVHB22, SD25b, SMV^{+}21, SHR25, SMŁWL25, SR23, STM21, SJ24, SSW20, SQL21, SMS23a, SWWZ23, SW24a, SZ24a, SL20, SJ23, SXCB24, SJZ24, SLT24, SZ24b, SRD25, SK21, SS20, SCdH21, SWW22, SJM24, SLGF20, SR21, SR22, SR24, SPZL21, SY25, SIAA25]. method [SW20b, ST20b, SS21, SXLZ21, SJZ22, SQB22, SWZC23, SWC24, SLC^{+}24, SZCW24, SHXZ24, SQGZ24, Tad23, TjW24, TO20, TQX23, TH23, TZX23, TS21, TSS24, TCW23, TGDMGT^{+}25, Top22, Tür20, UuIUK21, UHLZ24, Vak25, VR20, VADG22, VSG21, WLX20, WYH20, WW20b, WYS^{+}20, WWQ20, WCW20a, WZC^{+}20a, WZC^{+}20b, WCH20, WS21a, WWW21, WXZW21, WML21, WDZ21, WLX22, WXF22, WHT22, WCWZ22, WZ22, WL22, WDGS23, WL23, jWC23, WYP23, WWS23, WGC23, WaZWZ23, WYHQ24, WTSZ24, WCZ24, WF24, WZD24, WCS24, WQ24, Wan24, WFW^{+}24, WYW^{+}24, WJZ^{+}24, WWP^{+}25, WCC25, WCCC25, WYLL25, WZA25, WWZX25, WJ25, WL25, WF25, WQS25, WZ25, WS20a, WLW^{+}22, WWCJ23, WHC24, WH20, WLD22, WHH22, WCH23, WZS^{+}24b, hWCsLjH21, WARX23, WW25, WLJF25, XLS24, XL22, XKXL22, XZ25, XJ21, XFS23, XHY21, XFLZ22, XWXF23, XXYL23, XZT23, XZ24, XLLA20, XLCP20, XLZ^{+}21, XLS22, XSZ^{+}23, XLSF25, XS25, XLLR25]. method [YD21, YRC25, Yan20, YL20a, YK20, YH21a, YZ21, YHZ22, YJL22, YLW^{+}22, YZJ^{+}22, YWT22, YZN22, YDY^{+}23, YWL23, YLL24, YJ24, YMDW25, YMTP25, YWL21, YWZ21, YZZ20, YLZ^{+}24, Yem24, YHLA22, YYY^{+}22, YLSY24, YCL20, YLCC21, YFK23, YM23, Yu23, YWZ^{+}23, YCGL25, YWLF21, ZWOM21, ZCL20, ZWZD24, ZS25, ZCYL24, ZH20, ZCXG20, ZL20a, Zha21, ZLT^{+}21, ZD22, ZHZ^{+}22, Zha22a, ZC22, ZR22, ZDWR22, ZDLC22, ZYTX22, ZRL^{+}22, ZL22a, ZF22, ZTW22, ZH23b, ZQY^{+}23, ZZL23, ZZ23b, ZZH23, ZZ23a, ZSF23, ZJZ24, ZF24, ZYD24, ZLHC24, ZGC24, ZZ24a, ZZ25a, ZJ25, ZZW25, ZLL20, ZXZF20, ZZ20b, ZLH22, ZPC22, ZPK22, ZL23c, ZSW23, ZA23, ZZY^{+}24, ZZZD24, ZZ25b, ZL22b, ZCA23, ZS23b, ZCW24, ZSL22, ZST22, ZZW23, ZZ24c, ZHW20, ZQZ20, ZMZW24, ZAPK23, dCLW25, dFLdSLM22, dlRRG21, BL20, BGG^{+}20, BLMV24, GCSF21, HNLPP20, OSO^{+}23, SBMS22, SLUW^{+}21]. Method [VJ24, YL20b, dVDR20]. method-based [KTFI23]. methodology [CO21, MC25]. Methods [AGNP24, BDG21, GZG20, JKL20, KCC20, Lop20, Adj25, AAA25, AXLW24, ADGG25, ADB^{+}22, ABR20, AMV21, AAB^{+}21, AFV22, BDMX21, BDV23, BHK^{+}24, BS23a, BLF23, BRR25, Bar21, BW21, BGK23, BER21, BZ23, BBS22, BCM25, BBD21b, BBT25, BLLZ24, BJyST24, BSRVC20, CY23, CCY24, CAD24, CM21, CG21, CLR23, CBI^{+}23, CW20, CHZ23, CLW^{+}24, CLZN24, CYM25, CLY25b, CSI21, CSI23, CSS25, CFTC25, DSS20, DD20, DZ22, DL25, DY25, DTV20, DTW23, EAM20, ERG^{+}24, ELS24, EFFR22, FMS22, FXLS25, FP20, GP25, GIRB25, GLÖ21, GM21b, GP20, GCL25, GO23, HZ21, HJM23a, HLWL25, HCZ25, HJT25, HZ25, HSY25, Hoa22, Hof20, HHR24, JZPG21, JMCC24, JS25, JK22, JRZ21, JBKW23, KRO24, KP22c, Kam25, Kei21, KPS23b, KC22b, KGB24, KFD^{+}20, Kra20, KAKK23, Ku24, KD23, KP23, KWS21]. methods [LF24a, LS22a, LG23, LP23, Leg21, LWL^{+}20, LQ20, LZ20a, LZ21, LXJ^{+}21, LD22, LWW24, LHF24, LZ25b, LAL^{+}24, LX24b, LC21, LLXY23, LTXZ24, LL24a, LXL25, LLXY25, LMST20, MXLY24, MCW21, MMGS25, MI20, MA22, MZ23b, MCW22, Mit22, MLL25, MCV24, MR24, NSMM21, NE22, NH24, NCRGVR21, NAE22, NCCD24, OAD24, OS24, Oru20, Pac24, Par22a, PD21, QSW23, QGQ22, QHHW20, QFL23, QN20, RSAG^{+}24, RKRRL22, RHR^{+}21, SOPH25, ST20a, Sch21, SZS20, SYZ23, SQ23, SLM23, SMR24b, SK25, Sto20, SW24b, TZ25, TCW23, TKT^{+}24, Tor20, TKK22, TKK23, VLD^{+}24, WFH^{+}20, WW20a, WW24, Wan24, WKC21, XZ23, XF23, XLYY25, YSB^{+}24, YA23, YTK21, YCCW24, YL25, YLZ^{+}24, YFK20, YAA25, YM21, YW24, YH21b, ZM20, ZKTK21, ZQM20, ZFZ20, ZFL22, Zha22c, ZC23, ZXXX23, ZH23a, Zha23, ZZ24b, ZSGT24, ZPC20, ZLL23, ZLZ^{+}25a, ZXZ25]. methods [ZLJT20, ZQ25, dR25a, dlRGR20, CM20a]. Metric [WRMB23]. Metric-based [WRMB23]. MEW [LL25c]. MFEM [AAB^{+}21, WLJ21]. MFPC [FZW^{+}25]. MFPC-PIML [FZW^{+}25]. MFS [AAA24, Fur25, Sch21, YLCC21]. MG [HLPW22]. mGFD [TGDMGT^{+}25]. MHD [AKK^{+}25, AAG22, CSS25, DMX24, GO20, HUT25, JSBZ22, SJZ24, Tür20, WMMH25, ZS23a, ZSF23, ZZ24b, ZZZ25]. micro [AA20, DBH^{+}21, SQC23, WLZZ22]. micro-duct [AA20]. micro-electro-mechanical [WLZZ22]. micro-organisms [SQC23]. microarchitectured [KJL^{+}20]. microbeams [NTP25]. microbial [FGK^{+}21]. microelectronic [CRDS20]. microenvironment [HW25, LWZ22]. microorganisms [AWA^{+}22, AAG22]. microplates [AA21]. micropolar [JSSS24, LL24a, Rav20, SBH^{+}21]. microshell [WQS25]. microsphere [HYH23]. microstructure [ZYQ^{+}24]. microvessels [AI25]. microwave [KPS23a]. migration [Par24b, PLVL21, SQC23]. mild [DP25]. Miller [SXCB24]. mimetic [HSY25, Lee21]. Mindlin [DD22, YZZ20]. Minimal [SW21, BMvdZ21, WP25, BMP^{+}24]. minimal-residual [BMvdZ21]. Minimization [LMMMP20, BM20, Bar21, HSM^{+}23, ŁMMMP21, RGL22]. minimum [BBK24, KV20]. miscibility [ZGW23]. miscible [BM23, HLH^{+}20b, KS25, SR21, WCH20, YZN22, ZH20]. MITC3 [PND^{+}24]. Miwa [Waz20]. Mixed [ACXI24, ADFP23, CLZN24, DMRV24, GP23, JK22, LY21, QV22, TKK23, WCCC25, WYLL25, ZK21, AGLR21, ÁCS22, AMV24, APS21, BHJ21, BMZ20, Bar21, BHZ20, BB23a, BDL20, BLMV24, BGLZ24, CGO22, CCW22, CCGI23, CO21, CW20, CWC22, CXF24, CHL21, CSS25, CRF^{+}20, DLN24, DS20, DFSV22, DG22b, DGQT20, DFGS24, DS25, FHM22, FLS23, FWBY25, FCFW21, GQ23, GPRS22, GOV20, GORBS21, GI21, GIRB25, GDA21, GAM22, GKK^{+}20, GHL24, Gut20, HZJ^{+}22, HWW21, HZ24b, HLQ24, Hoa22, HAC23, KPJO22, KK21, Koc22, KBW23, Lee21, LZ25b, LXH24, LS21, LS23, LLW^{+}23b, LTXZ24, LXL25, MHHR20, MT25, MD21, MCW22, MS20b, MCV24, OAD24, OSZ22, Qiu21, SM22a, SL20, SJ21, SK25, SLGF20, SR22, TCA22a, TCW23, WTSZ20, WDZ21, WX22, WCWZ22, WF24, WYW^{+}24, WWZX25, YJ24, YLSY24, ZR23, ZH20, ZHYL21, ZQY^{+}23, ZF24, ZL25]. mixed [Zha24, ZCA23, ZQ25, dR25a]. mixed-dimension [Koc22]. mixture [BM23, SRP21]. ML [BMvdZ21]. ML-MRes [BMvdZ21]. MLMCE [LL22b]. MLMCE-HDG [LL22b]. MLP [IBY^{+}24]. MLS [ZZL23]. MLS-based [ZZL23]. MM [TW23]. MM-ADMM [TW23]. MMALE [SJ24]. MMOC [CLZW20]. MMOCAA [ZLW23]. MMOCAA-DDM [ZLW23]. MMPDEs [TW23]. mobile [BS23a, CLL21, CLZW20, DC22, QXCG20, YLZ^{+}24]. mobile/immobile [BS23a, CLL21, CLZW20, YLZ^{+}24]. mobility [GGT24, HLJ25, LL25a]. Modal [ZLZ^{+}25a, FK20, GLC24, NERB^{+}22]. mode [AS23, CDPL24, NML23, RCSM23, SSZ20, XL25b]. Model [ABC^{+}22, HGD21, LSPRV24, WH20, YL20b, AADMA23, APK^{+}25, AI25, ASGGRG23, ADR24, AOL22, AL23, ADL^{+}21, AZY22, AGM^{+}22, AJK23, AMC23, AAK24, BWZ23, BNSZ24, BGR23, Bal21, BRÁ^{+}20a, BZDQ21, BMP21, BKR24, BDE^{+}24, BMS22, BT22, BOSV20, BM23, BABS20, CCH24, CLZ^{+}24, CRL^{+}25, CLLG23, CRM22, CM23, CBLZ24, CMS23, CLH21, CWC22, CLW^{+}24, CLZN24, CZC^{+}25, CYM25, CXX^{+}25, CHJR21, DD20, Der23, DK24, DY25, DWY^{+}25, FFC^{+}25, FGK20, FdOS20, FGG^{+}20, FL22, GORBS21, GdFFF25, GPY22, GW25, GBK23, GM22, GJM22, GHSR23, GV20, GI20, GML^{+}20, GGRBRG20b, GGRBRG20a, GZAR23, GHK25, GSRG21, Ha20, Hag24, HMFM25, HK20b, HK20a, HKH^{+}24, HJZ23, HJM23b, HXW23, HH21, HH22, HLH^{+}20a, HLH^{+}20b, HHY^{+}22, HFG23, HWZQ23, HZ24b, HZ24c, HX22, HKL^{+}23, HW25, IZ20, JWL20, JSZ24, JAMX24, JWZX25, JL23, JM20b]. model [KP22a, KP22b, KEUY23, Kar24, KTY23, KGW^{+}21, KMP20, KTY24, KR24, LMP23, LLN24a, Li20, LWSZ21, LGQZ21, LD22, LAX22, LXRQ22, LHC23, LY23, LC23a, LZF23, LQYJ23, LZBC24, LZL24, LaZMY25, LLY25, LR23, LZ20e, LFY^{+}20, LS22b, LYZ23, LQZ23, LLW^{+}23b, LLLL24, LWFS24, LCC^{+}24, LARGVR23, LP22a, MrZQ24, MJRM20, MT25, MZCK25, MS20b, NZ24, Nde25, NSMR25, NTP25, NNTCB25, NS24, NM23b, NS20, NCRGVR21, NDB^{+}24, PH23, PNT^{+}23, PTNR24, PZZ23, QHHW20, QWC^{+}23, RZSL23, RG21, RNMÖX24, RWaZ23, RWZG24, RYAM21, RGL21, RNC25, RUMT^{+}24, RLSC23, RGL22, SK23, SHWC22, SSZ20, SWWZ23, SHQL20, SJZ24, SYHN24, SRP21, SY25, SIAA25, TC25, TSW^{+}25, TJYL22, TLJ24, TKMK25, TKT^{+}24, TPOT20, VGA^{+}22, WSGT21, WZY22, WXF22, WCWZ22, WDYZ24, WT25, WMMH25, WWG^{+}25, WF25, WHH22, WS21b, WSO^{+}23, WZCS20, WCY^{+}24, XKXL22, XZ25, XWXF23, XDL^{+}25, XSL24b, XTZZ24, XHL24]. model [YLY20, YTK21, YJL22, YLQ22, YWT22, Yan24, YMDW25, YYY^{+}20, YCWS20, YCGL25, YW20, YZW^{+}22, ZKBC21, ZBKC22, ZWMF24, ZL20a, Zha21, ZXGF22, ZHHL23, ZWT23, ZL23b, ZGW23, ZYD24, ZZZ24, ZZW25, ZWCS20, ZHZ23, ZS23b, ZMS21, dCG20, dABCC25]. model-based [GZAR23]. modeled [zLqLZ^{+}24]. Modeling [FKR21, SS20, YTT^{+}22, AO24, AS21, Abd24, AKA^{+}21, AS23, AZL20, Alf24, AEF23, BDE^{+}24, BWG25, Bot20, DAAEA25, EJWD23, EKB20, FJP22, GWYR22, HTDT23, KN25, LBTVG22, LWZ22, LS20, LYRW21, LVDDNT25, MIOK21, MRP20, RST^{+}22, SQC23, SDM21, SMS23a, SC21b, US20, Vak25, WS25, YA21, ZHZ^{+}22, ZYM24, ZDLL22]. Modelling [MBSEA24, SGAC^{+}20, AM20, AFV22, DMŠ22, KAKK23, vtWHG^{+}22, XJY^{+}20, ŻW21a, ZWOM21]. models [AHZ^{+}25, AD21, AL22, AHOB24, ASSV20, BMH22, BHPS24, BC20a, BDS^{+}22, BGD22, CLL21, CDLW20, CS23b, CXJ25, CFG^{+}23, ČPPŠ20, DAAEA25, DGQT20, DFGS24, EP23, FNGDV23, GHS23, HEDT20, JZ24a, JW20, KBR20, KHY25, KK21, Kri20, KJ22, KOMK20, LF24a, LH24, LML23, LZL^{+}21, LP22b, LN22, MMGS25, MLJ^{+}20, MJVH24, MCK^{+}23, MZ23b, PN24, RHB25, SGZ^{+}23, SW24a, SLM23, SSK^{+}23, SLYL22, THF20, TZ20, UHP23, WYT23, XFS23, XHH^{+}23, XZS^{+}23, YMTP25, ZXZF20, Zha22d]. modes [Bel24, Hak20, SM22a]. modification [LZL24, LZG^{+}23, ZLL20]. Modified [ASB25, Ale23, ABLS23, LWFS24, Men22, NTP25, RZŠ21, BWZ22, CGK24, CZK^{+}25, CZ20, CSI23, CCGRL21, DK23, GW25, HSZ22, LGQZ21, LJ24, LJY^{+}25, LLWX25, LSZ24b, PHL21, RG21, RUMT^{+}24, SHQL20, SZ24b, SWW22, TZX23, Vak25, WQS25, XZT23, XP22, YM21, ZZ21, ZTW22, ZWT23, ZH23a]. Modular [PKPD^{+}24, AAB^{+}21, HJM23a, HJ24]. modulus [Hag24, ZZ23b]. moist [KTFI23]. moisture [BBT25]. MoL [BDL20]. molecular [ZJCG21]. mollification [ZDWR22]. molten [Cao21]. moment [BN20, Leg21, LLZ^{+}23, Mal22]. moments [LLZ^{+}23, WCY^{+}24]. moments-based [LLZ^{+}23]. momentum [KK23b, WWP^{+}25, ZDW23]. Monge [SNHH21]. monodomain [RNC25]. Monolithic [BAB^{+}20, Haj21, BBD21b, CWTM20, HC23, PCC21]. monotone [ABL24, DLW24, NY22, NFS24, RZ21]. monotonization [Orl23]. Monte [AHM20, BW21, LLY25]. Moore [AAC23]. Morley [DPP21, Men22]. morphable [WFW^{+}24]. Morphological [WWG^{+}25]. Morse [KC20, KWK24]. mortar [APS21, MMPR20, MFF25]. motility [HHNV23]. motion [BS24, LL24b, MW20, ZZZC24]. Motz [Rob25]. movable [XDL^{+}25]. Moving [KC21, SGZ^{+}23, BDV23, CCG^{+}23, HRS20, JBKW23, KK23a, KJ21, LS22a, LSV21, LSV24, Li21, Ma25a, MGBN22, QFL20, SHR25, SD21, SNHH21, WFW^{+}24, WHC24, XJ21, XZ24, YMTP25, ZY21b]. moving-domain [ZY21b]. MPI [KTXP21]. MPI-AMRVAC [KTXP21]. MPP [CLXZ25]. MPP-TV [CLXZ25]. MPS [Sch21, YLCC21]. MPS-MFS [YLCC21]. MR [TZ25]. MR-WENO [TZ25]. MRes [BMvdZ21]. MRT [BP20]. MRT-lattice [BP20]. mufflers [XJY^{+}20]. Multi [LLX24, LQZ23, NNTCB25, Qiu21, TT25, ZM20, ZJZ24, AHZ^{+}25, AM23, ASH^{+}21, AMC23, ASZ24, Bel24, BMST20, BHC20, CBŁP23, CLW^{+}24, CYKL23, CM20a, DSL24, DAS20, EFK^{+}23, FC24, FKKV24, Fia21, GKHU22, GCMS21, HLL24, HZR23, Hoa21, HYW^{+}24, HLN24, HAC23, HWY25, KV22, Kim21, KLK^{+}22, KP23, LH24, LHC23, LSRL24, LYL^{+}22, LLU24, nMjWlS24, MLLL24, MR24, NC20, PWMC23, PCLZ25, PLVL21, QGQ22, RCSM23, RKK25, SM22a, SYHN24, Top22, WZZ20, WFX24, WTSZ24, WNF^{+}25, WYT23, WLH24, XZSY23, XSL24b, ZP24, Zdu21, ZYTX22, ZSK23, ZGW23, ZX20, ZWS22]. multi-agent [ZWS22]. multi-asset [CYKL23]. multi-component [WYT23, ZGW23]. multi-core [Fia21]. Multi-dimensional [ZM20, AHZ^{+}25, AM23, CBŁP23, DAS20, GCMS21, HZR23, HWY25, Kim21, LSRL24, LLU24, QGQ22, WTSZ24]. multi-domain [nMjWlS24]. multi-frequency [MR24]. multi-group [XZSY23]. multi-layer [Hoa21]. multi-layered [ASZ24]. Multi-level [Qiu21, HLL24, HLN24, NC20, PLVL21, Zdu21]. multi-linear [BHC20]. Multi-material [NNTCB25]. multi-mode [RCSM23]. multi-noise [LHC23]. multi-order [HZR23]. multi-parameter [FC24]. Multi-patch [ZJZ24, BMST20, EFK^{+}23, FKKV24, KV22, SYHN24, TT25, WNF^{+}25]. Multi-phase [LQZ23, ASH^{+}21, LH24, ZP24]. Multi-relaxation-time [LLX24]. Multi-resolution [TT25, WZZ20]. multi-scale [AMC23, CLW^{+}24, DSL24, PCLZ25, WFX24, XSL24b]. multi-slice [KLK^{+}22]. multi-stage [WLH24]. multi-step [Bel24, PWMC23]. multi-symplectic [ZSK23]. multi-term [GKHU22, HYW^{+}24, HAC23, KP23, LYL^{+}22, MLLL24, RKK25, SM22a, Top22, ZYTX22, ZX20]. multi-trace [CM20a]. multiblock [ZP24]. multibody [JJOG20]. multicomponent [Bal21, HLZ24]. multiderivative [TSS24]. multidimensional [BKR24, Bou24a, Bra20b, JM20b, KJ20, KSSS23, MH25, SQLZ21]. Multigoal [ELN^{+}20, BDE^{+}24]. Multigoal-oriented [ELN^{+}20, BDE^{+}24]. Multigrid [DDFZ23, LZ25c, ZH23a, ABS23, AFM^{+}24, BHPD23, BMK21, CAD24, Gun22, HC21, KWS21, LHC^{+}22, Liu20, PD21, SSW20, TS24, WPLW22, XLS22, YPZ^{+}21, dlRGR20]. multigrid-homotopy [Liu20]. multigrid-preconditioned [TS24]. multigrid-reduction-in-time [YPZ^{+}21]. multigroup [TO20]. Multilayer [FNGDV23, ZNC24]. multilayered [AZL20]. multilevel [AHM20, BW21, CEB24, JRZ21, LNF24, LGQZ21, ZZL22]. multimodal [CLXZ25]. multinode [DDNS21]. multipatch [MST25]. multiphase [Bal21, EH23, ENF24, HQCW22, LpWW^{+}25, LVDDNT25, MRN^{+}20, NC20, NDC22, SCD25, XTZZ24, YWZ^{+}23]. Multiphysical [SP20]. multiphysics [BEG^{+}21, GPY22]. Multiple [BM24b, CLZ^{+}24, JICRMG^{+}20, ADL^{+}21, ADG23, BTYM^{+}21, BD23, BC20b, HLL24, HHL22, Kra20, LLX23, MK20, SN20b, WZ21, Waz20, WZCS20, WCY^{+}24, YWZ^{+}23]. multiple-relaxation-time [ADL^{+}21, LLX23, WZCS20, WCY^{+}24, YWZ^{+}23]. multiple-time-step [WZ21]. multiplicative [CS20a, GSS24, LH23a, LY25, LS22b, Shi24, SF25, dlRRG21]. multiplier [LWFS24, OAH^{+}24, ZNH24, dCLW25]. multipliers [RHA^{+}24]. Multipoint [Jaw21, LXL25]. multipole [HHL^{+}21, HLN24, LGQZ21, LX22, Pta21, RKRRL22]. multiquadric [SZ21]. multirate [BCCR20, JWW21a]. Multiscale [APS21, CO21, CEB24, CEL20, HLH24, KC23a, OCRN22, BdSC24, EKB20, FZW^{+}25, HLPW22, PG21, PRN^{+}23, WL23, WHH22, YD21, YD23, ZL20a, ZTL22]. multistep [LY24]. multithreaded [ÇM20b]. multivariate [SLM23]. Mumford [SHQL20]. MUSCL [CCGRL21]. MUSCL-type [CCGRL21]. MUSIC [KP22d, Par24a]. MUSIC-type [Par24a]. mutualism [LQCM20]. mutualistic [dCG20].

N [DK23]. N-S [DK23]. Nadaraya [JZ24a]. Nagumo [ZZZ24, KUK20, LZ20a, LY23, LLLL24, LTYZ25, WPLW22, ZSY22]. nano [AMPB^{+}22, CBK24, GBK23, HANA24, SJM24]. nano-fluid [HANA24]. nano-polymer [AMPB^{+}22]. nano-scanner [GBK23]. nanocomposite [AA21]. nanofluid [AKK^{+}25, AWA^{+}22, AAG22, AIG^{+}21, AAS^{+}20, DK23, DK24, GO20, IBY^{+}24, MAT^{+}20, OAA^{+}22, TSARR21, WX22]. nanoliquid [RG21]. nanoparticles [JFS14, JFS20, MAT^{+}21]. nanophotonics [NS20]. nanoshell [PCL25]. nanoshells [JCP^{+}25]. nanotube [HTPV25, MMP21, PCM20]. nanotube-reinforced [HTPV25, MMP21]. nanotubes [AAS^{+}20, WCW20b]. narrow [WDGC20]. NATriuM [KWK^{+}20]. Natural [AA25b, ÁCS22, BRÁ^{+}20a, BL20, CZ23, DK24, GO20, Ima20, LLD23, LSF25, LS23, OEASY24, SOPH25, SBH^{+}21, WLJF25]. Navier [GLF25, JWL20, WHH22, ZFQ24, Agg20, AJLM24, ASB25, ABR20, ACRBT23, BdSC24, BSRVC20, CCGI23, CW20, CL20, CG25, CTZ22, DKMS22, GOV20, GI21, GGT24, HJM23a, HDM24, HQLJ20, HX22, JRu24, Kap24, KNBR22, LHF24, LD25, ŁMMMP21, LCT22, MRN^{+}20, MZHJ24, MZZ^{+}25, PNT^{+}23, QQY^{+}23, QD20, RZC21, RDP20, SCCS21, SXKL24, WDZ21, WWS23, WaZWZ23, WFW^{+}24, WWH23, XW24b, YHZ22, YW20, ZDW23, ZF24, ZLZ25b]. NDM [RHA^{+}25]. Near [PGK20, ZWZD24, GRR20, MIOK21, US20, YLY20, ZJCG21, aKST22]. Near-field [ZWZD24]. near-incompressible [GRR20]. near-inextensible [GRR20]. Near-wall [PGK20, MIOK21, US20, YLY20]. nearly [AHLZ24, GI21, KP22a, ZR23]. nearly-incompressible [AHLZ24]. Negative [PKP21, Leg21, SR20, SV21]. negatively [ZZ23b]. Neilan [DST25]. Nektar [YPC^{+}21]. Nernst [GIRB25, GDA21, LSWY21, ST22]. nerve [JJ23]. nested [Bot20, KCC20]. nesting [MrZQ24]. net [MP23, MZCK25, LL25b].

NET-RAT [MZCK25]. Network [BV21, CGHJ24, CWX22, Dai22, DSR23, GY25, HSR25, HLW23, HCY^{+}24, HZW21, IBY^{+}24, Koc22, LCC22, LC22, MH23, NES23, Pel23, QWLQ25, SSZ20, SKG24, SGR^{+}23, WFX24, XYS^{+}24, YK24, ZL20b, ZXGF22]. network-based [SKG24]. Networks [UBP^{+}25, AFV22, AFM^{+}24, BFI^{+}23, BF21, CGGY23, CKFR24, CM24, CLY25a, CDG^{+}23, GYL25, GWYR22, HLX22, HHL22, JKC24, KC22a, LDW^{+}24, LWG^{+}25b, LWG^{+}25a, LLY25, LCC^{+}24, MB24, MJZ25, MMN24, MRP20, MVHÖ23, MSBO23, OS24, PL24, SDM21, SW24b, TFWX23, WZL25, XS25, YDY^{+}23, ZvNH25, ZL23a, ZLQ23, ZL23b, ZHH21, ZWY23, ZW20b]. Neumann [ADGG25, AH25, FM25, HYW^{+}24, LZ20b, LSZ24a, nMjWlS24, Man24, MM21, MK20, SM23, jWC23, YS23]. Neural [UBP^{+}25, YDY^{+}23, ZL20b, AFM^{+}24, BF21, BMP^{+}24, CGGY23, CKFR24, CGHJ24, CLY25a, CWX22, CDG^{+}23, Dai22, DSR23, GY25, GYL25, GWYR22, HSR25, HLW23, HLX22, HCY^{+}24, HHL22, IBY^{+}24, JKC24, KC22a, LL25b, LDW^{+}24, LWG^{+}25b, LWG^{+}25a, LLY25, LCC22, LC22, LCC^{+}24, LZS25, MB24, MJZ25, MMN24, MCK^{+}23, MRP20, MVHÖ23, MH23, MSBO23, NES23, PL24, SDM21, SSZ20, SKG24, SGR^{+}23, SW24b, TFWX23, WZL25, XYS^{+}24, XS25, YK24, ZvNH25, ZXGF22, ZL23a, ZLQ23, ZL23b, ZHH21, ZWY23]. Neural-network-based [ZL20b]. neural-weighted [BMP^{+}24]. neuron [LY23]. neurostimulation [DTV20]. neutral [HZ21]. neutrally [ZZZC24]. neutron [CDGM25, MWY21, SLC^{+}24, TO20, XZSY23]. neutrons [XZSY23]. newly [CZK^{+}25]. Newton [ASB25, BFVY23, BFE^{+}21, CKS22, EED^{+}25, HLPW22, KSB25, LQ20, Loh21, YPC^{+}21, ZC22, ZZZ25, ZHW20]. Newton-like [Loh21]. Newton-MG [HLPW22]. Newton-type [LQ20]. Newtonian [AAWJ20, CBK24, EGF^{+}23, HMŁW^{+}24, JSZ24, KC23a, OAA^{+}22, OEASY24, OSZ22]. Nicolson [BDF24, GLLZ23, HLJ25, LHN24, LSW20, Li24b, LLZL24, LS25a, RWaZ23, SX25, WZ22, ZZH23, ZFQ24]. Nicolson-Leapfrog [ZFQ24]. Nine [MN23b]. Nine-point [MN23b]. NIPG [MZ23a]. Nitsche [CCWC23, DFK22, FG25, MCW25, WCCC25]. Nitsche-type [FG25]. NLSE [FWLD22, ZXZ25]. Nodal [JRu24, KMS20, ASB25, BNSZ24, SLC^{+}24, XL22]. Node [LNF24, HC21, NF25, ST23]. node-based [PND^{+}24]. node-coarsening [HC21]. nodes [Sch24]. noise [AMK25, CS20a, DYZ22, GSS24, LHC23, LH23a, LY25, LS22b, RNN22, SDG20, Shi24, SF25, TPOT20, WWH23]. noisy [TH25, UR24]. Non [DFS21, DCCM23, EGF^{+}23, Leg21, LLX23, LZ25a, SDM21, SC21b, UHLZ24, ZQ25, ADKH20, AS21, ADG23, ABP25, APS25, AAWJ20, BDMX21, BTCV20, BZY20, BCF^{+}23, CMPR23, CLZ^{+}24, Cha24, CXX^{+}25, CCGRL21, CL22b, DKH22, DK24, EFK^{+}23, FFC^{+}25, GW25, GCC20, GO25, GCJ23, HJP23, HV21, HMŁW^{+}24, HRvdZ20, HKL^{+}23, Ima20, JKS25, JSZ24, JWZX25, KC23a, LFF^{+}24, LSV21, LE24, LSV24, LL20a, LLZ21, LXJ^{+}21, LD22, LXRQ22, LDH24, LCX^{+}20, LS22b, LMMMP20, ŁMMMP21, MIOK21, MGD23b, MLJ^{+}20, MRPK21, MCW22, MKP23, MZCK25, MBSEA24, NAS20, NNH^{+}25, NERB^{+}22, NKPA^{+}25, OAA^{+}22, PG25, Par22b, PLR21, PTNR24, PZZ23, QQX21, QC24, RZSL23, RLF21, RGL22, SB21, SD25a, ST20a, SB25, SL20, SSD23, TC25, Tan25, TMML23, TLJ24, WXZW21, WZTZ23, WH25, WCLQ25, Wat22, XLSF25, YSB^{+}24, YY25, YWLF21, ZZW20, ZJJ20, ZXGF22, ZF22]. non [ZZ23b, ZLF^{+}22, ZGL23, ZLL23]. non-affine [TC25]. non-autonomous [YY25]. non-body-fitted [MIOK21]. non-circular [DKH22]. Non-conforming [ZQ25, BDMX21, EFK^{+}23, NAS20, PLR21, SK25]. non-conservative [MZCK25]. non-contact [JWZX25]. non-convex [ABP25, FFC^{+}25, LS22b, NKPA^{+}25, RGL22, WCLQ25]. non-differentiable [YSB^{+}24]. non-divergence [TLJ24]. non-equilibrium [CLZ^{+}24, DK24, PZZ23]. non-Fickian [ZLL23]. non-Gaussian [HJP23]. non-globally [YY25]. non-grey [MBSEA24]. non-homogeneous [CCGRL21, MRPK21, QC24, WXZW21]. non-hydrostatic [RZSL23]. non-interpolatory [ST20a]. Non-intrusive [LZ25a, SC21b, AS21, ZXGF22]. non-isothermal [CXX^{+}25, MCW22]. non-isotropic [CCGRL21]. non-iterative [Par22b]. non-linear [ADKH20, ADG23, CMPR23, Cha24, CCGRL21, CL22b, HRvdZ20, NNH^{+}25, SD25a]. non-linearity [MKP23]. non-local [LE24, LDH24, MLJ^{+}20, PG25, SB25, ZGL23]. non-magnetic [LSV21, LSV24]. non-Markovian [LCX^{+}20]. non-matching [BCF^{+}23, LFF^{+}24, ZLF^{+}22]. non-modal [NERB^{+}22]. Non-negative [Leg21]. non-negatively [ZZ23b]. Non-Newtonian [EGF^{+}23, AAWJ20, HMŁW^{+}24, JSZ24, KC23a, OAA^{+}22]. Non-orthogonal [LLX23]. non-oscillatory [GCJ23, WZTZ23]. non-overlapping [BTCV20, MGD23b]. non-premixed [SB21]. non-rectangular [JKS25]. Non-relaxed [DCCM23]. non-residual [GCC20]. non-self-adjoint [Wat22]. non-selfadjoint [BZY20]. non-singular [YWLF21]. Non-smooth [UHLZ24, FFC^{+}25, GO25, LLZ21, SSD23, ZJJ20, ZF22]. non-spherical [RLF21]. non-standard [APS25]. non-stationary [HJP23, LD22, LXRQ22, LMMMP20, ŁMMMP21, MLJ^{+}20, SL20]. non-uniform [GW25, HV21, HKL^{+}23, Ima20, LL20a, LXJ^{+}21, QQX21, Tan25, TMML23, WH25, XLSF25, ZZW20]. Nonconforming [AMV23, AMV25, Bar21, CKBK24, KS25, XSL24a, BFVY23, BMPP22, HWC24, JP25, LR23, LS25a, MCW25, PZL23, PLZZ25, SL20, SL21, SZ22a, SL24, SLT24, SX25, WLW^{+}22, Yu23, ZMS21, ZNMS23]. nondivergence [Füh21]. nonequilibrium [AGA^{+}21, LS25b]. nonhomogeneous [MM21, SWZC23]. nonisothermal [AEZ24]. Nonlinear [AAA25, sCpLYhG23, DH24, SFR21, TC25, ZWW^{+}23, AKO21, AKK^{+}25, ANZZ23, AA25a, AN23, ANN25, APS21, AAK24, BMH22, BdSC24, BHPS24, BKM23, BDE^{+}24, BAB^{+}20, CLZ^{+}24, CG21, CCF^{+}21, CLL21, CLZW20, CLZ23b, CGK24, CLY25a, CYM25, CLQL25, CKS22, CFG^{+}23, DPP21, DZ22, DLW24, DBR21, DWY^{+}25, DR25b, EED^{+}25, EJWD23, ELN^{+}20, FKW22, FAM25, FHZ22, FX22, GPY22, GHAU24, GLLZ23, GCJ23, GWN21, Gut20, GO23, HJP23, HLL24, HGD25, HGW21, HJXW23, HHL22, HWY25, IQS20, JKC24, Jaw21, JJZ21, JCJ24, JZHZ22, KRO24, KEUY23, KLY25, KWLW24, KSSS23, LHN24, LHC^{+}22, LPK22, LL20b, LT23, LZY24a, LLZL24, LY25, LZ25a, LL22c, LYH24, LLW23a, Liu20, LY21, LQY22, LGZC23, LRZ23, LLXY23, LLXY25, LYLH25, Lob21, LdO21, LP22a, MCD24, MRAC24, MB24, MJRM20, MA22, MLLL24, MH25, MN23a, MN23b, MS20b, NES23, NRV22a, NH24]. nonlinear [NNT22, PG25, PKP21, PZL23, PLZZ25, PSI^{+}24, SOPH25, SLD25, SM22a, SDJ^{+}22, SDG20, SJ21, SZ22a, SZ22b, SJ23, SL24, SZ24b, SQ24, SLYL22, Tan24, TFWX23, TRL23, TCW23, UHL^{+}24, WFH^{+}20, WTSZ20, WLZZ22, WL22, WZY23, WLL23, WGeL24, WSCP24, Wan24, WLC24, WYW^{+}24, Wan25, WYLL25, WQS25, WS20a, WLD22, WLH24, XZ23, XSL24a, XSZ^{+}23, XF23, XL25b, XL25a, YGZ^{+}22, YWT22, YWL23, YMTP25, YWL21, YFK20, YFK23, YLWZ20, YX21, ZQM20, ZLPR20, ZJJ20, ZL20a, ZFL22, ZRL^{+}22, ZWT23, ZSK23, ZZS23, ZGC24, ZWCS20, ZLH22, ZWY23]. nonlinearities [RZ21, WL20]. nonlinearity [DZ21, HGW18, JCJ24, LHN24, LS25a, SX25, Yan25b, ZH23b]. Nonlocal [BKM23, JAMX24, AN20, BH22, BDP23, BS23c, BGD22, CM22, DSL24, Gal21, GLLZ23, JCP^{+}25, JW20, KCD^{+}23, KC23b, Li20, LF24b, LN22, QQX21, RZ21, ST20a, SW24a, SL24, SJM24, TSW^{+}25, TLJ24, TTKV20, WT25, WCW20b, WSS22, YWZ21, YHF24, YCGL25, ZWMF24, ZLH22, ZZ24c]. nonmonotone [GB21, GB23, GO23]. nonoverlapping [LLR25]. nonsmooth [BM20, KSB25, LK20, LWX21, Pło22, PT24]. nonstaggered [QD23]. nonstationary [HYJ20, LZL24, LDM25, QS20, SHD23, WZ21]. nonsteady [WLQ22]. nonsteady-state [WLQ22]. nonsymmetric [LZW21a, LZW21b, SM22b, XL25b, ZYW20]. nonuniform [DR25b, FZS22, KLK^{+}24, KP23, LKC^{+}24, MGD23a, ySRC22, Tan24]. nonzero [RS21]. norm [AZHY22, CC20, HZ25, KPS23b, Ku24, LX24b, LQY22, LZ23, MN21, PKP21, RKC25, SV21, STO22, TS21]. Normal [WSS22]. norms [BDMX21, LLXZ23]. notches [BNB20, GNB24]. Note [Ano20-69, FP20, XL20]. Notes [BHJ21]. notice [JL20, JFS20]. Novel [BB23a, GMS23, GP20, HUT25, AAL24, AZY22, AN21, BGPS20, BLLZ24, BT22, CRL^{+}25, CBLZ24, DR24, DR25b, FC24, FSH^{+}23, GLL^{+}20b, GLC24, HKH^{+}24, HJXW23, HQL23, HTPV25, JWL20, KGK25, KN25, LL22c, LS25a, eLLAN25, LSRL24, MFF25, MNB23, Par24b, PCL25, SJ24, SA22, SWZC23, TL22, TLJ24, UHLZ24, WDZ21, WLC24, WZA25, WLD22, XHM20, XFX22, XSZ^{+}23, XF23, YL20a, YCZ^{+}21, ZCXG20, ZX24, ZCA23, ZS24, ZZ24c].

November

[Ano20-56, Ano20-57, Ano21-51, Ano21-61, Ano22-62, Ano22-60, Ano23-61, Ano23-62, Ano24-47, Ano24-68, Ano25-61, Ano25-65]. Noyes [KEUY23]. nozzles [BEKV21]. ntkFoam [MWY21]. nuclear [MWY21]. nuclei [BM24b]. null [Yan20]. null-space [Yan20]. number [CYC^{+}23, EAM20, KMS20, NSMR25, TZ25, ZPC20]. numbers [AL22, ABMS20, CO20, HMW21, HBC^{+}21, JZHZ22, WDGS23]. Numerical [AHZ^{+}25, AKA^{+}21, ARV23, ASH^{+}21, Ant21, ANN25, AFBM25, BS23a, BWR21, BTYM^{+}21, BNQ21, BKW22, BLRRS20, BT22, BOSV20, CML^{+}25, CRRS21, Cha24, CSS23, DMG23, DKH22, DKMS22, DKL24, EJWD23, FAM25, FL22, GZG20, GDA21, GCC20, HHJO20, HSWG21, HJM23a, HJ24, HLMM20, HQCW22, HYJ20, HX22, ISA24, JMR24, JMM20, JWW^{+}21b, KP22b, KK23b, LE24, LSH21, LWSZ21, LQYJ23, LL25c, LSRL24, LXH24, LFY^{+}21, LLJ23, LD25, LARGVR23, LHQW25, MAT^{+}21, MLS22, MH25, MZCK25, NE22, NS24, NDB^{+}24, QHHW20, RG21, RYAM21, SC21a, SGZ^{+}23, SV21, SN20b, SMR24b, SBS24, TSARR21, Vab22, WZC^{+}20a, WSGT21, WGeL24, WLW^{+}25, WXCL25, WKGH21, WSO^{+}23, WWH23, XCLX24, XXAD21, XGL21, XLZ^{+}21, YZ21, YL22, YZJ^{+}22, YYZ22, YL20b, YLFL20, YTY24, ZSH^{+}21, ZDW^{+}20, ZDWR22, ZYD24, ZYQ^{+}24, ZJCG21, ZL23c, ZXZ25, ACDL24, AI25, AFS24]. numerical [ACV21, AAL24, ABMS20, AFV22, ASSV20, AEF20, BRR25, BZDQ21, BGK23, BDL20, BDP23, BGPS20, BDE^{+}24, BKT^{+}23, BCM25, BGLZ24, BPID20, Cao21, CWW22, CK23, CMS23, CMY23, CLXZ25, CZC^{+}25, CXX^{+}25, CS25, CCGRL21, CPBF21, DD21, DN22, Din21, DWY^{+}25, DP20, DC22, DYZ22, EFK^{+}23, FMS22, FC24, FRW22, GF21, GKP25, GGRBRG20a, HLEF20, HH22, HHM21, HZ24a, HZ24c, HJTW25, HD21, HSJSS25, HHNV23, HZR23, Hof20, HZW21, Hop23, HLN24, HWY25, HHG20, IZ20, JSSS24, JWZ22, JZ25, JM20b, KRO24, KRESA21, KHH25, KLK^{+}24, KN20, KJ20, KGK25, KN25, LF24a, LNT22, LSV24, Lee20, LL25a, LX24a, Li25, LYZ23, LLU24, LP22b, MRP20, NM23a, NF25, OAD24, OCCM21, PV20, PS25a, PWWJ23, RHA^{+}24, RNMÖX24, Rou20, RR22, RKK25, SLD25, SSY22, SB24b, SHWC22, SP20, SSD23, SKG24, SSK^{+}23]. numerical [SNL23, SIAA25, Tad23, TGARGT^{+}23, TKT^{+}24, VJ22, WYS^{+}20, WCWZ23, WDYZ24, WZ24, WCC25, WL25, hWCsLjH21, WS21b, XL22, XFS23, XTZ^{+}21, YRC25, YZL20, YTK21, YLQ22, YCCW24, YMTP25, YLZ^{+}24, ZZCM20, ZO21, cZsCyC23, ZWT23, ZZL23, ZNC24, ZHZ23, dR25a, dABCC25]. Numerically [LZ20d, JL19, JL20]. numerics [CFP22]. NURBS [RHA^{+}25, SFR21, WNF^{+}25]. NURBS-based [SFR21]. NURBS-Divergence-Meshless [RHA^{+}25]. NVD [YL20b]. Nyström [HSR^{+}23, TPA24].

O [GO20, MAT^{+}20]. object [ASZ24, MN24, Par24b, PH23]. objects [BHDV22, FGK20, JWZX25]. oblique [BWZ22, BWZ23, TSARR21]. observation [AOL22, YYY^{+}20, YTH^{+}21]. observations [HTBMC20, Wat22, Xu24]. observed [SXKL24]. observer [SSZ20, ZWW^{+}23]. observer-based [ZWW^{+}23]. obstacle [KPS23b]. obstacles [BBS21, XLYY25]. occurring [BNC^{+}23]. ocean [MS22, XFX^{+}24, APK^{+}25]. Octave [RHR^{+}21]. octic [JVBM23]. October [Ano20-54, Ano20-55, Ano21-65, Ano21-58, Ano22-63, Ano22-64, Ano23-64, Ano23-67, Ano24-50, Ano24-57, Ano25-68, Ano25-69]. octree [SGF^{+}21]. Off [CCGR24, BABS20]. off-lattice [BABS20]. Off-site [CCGR24]. ohmic [AHZ^{+}25]. oil [CCGRL21, DP20, LZL24, WYS^{+}20]. oil-water [WYS^{+}20]. Oldroyd [DY25, HANA24, LYL^{+}22, LZL^{+}25]. Oldroyd-B [DY25, HANA24, LZL^{+}25, LYL^{+}22]. once [LDH24, ZS25, ZX24, ZGL23]. One [LL24a, GHAU24, HUT25, HM20, HA23, MMC21, PLVL21, SBY^{+}25, Tan25, TO20, YMTP25]. one-dimensional [HM20, HA23, MMC21, SBY^{+}25, TO20, YMTP25]. One-level [LL24a]. one-parameter [Tan25]. one-step [HUT25]. one-way [PLVL21]. Online [RBSC24, LHH22, PKPD^{+}24]. onset [KRESA21]. onto [HSM^{+}23, WZXN21]. Open [Kam21, KKA^{+}21, RGS^{+}20, XXAD21, ZK21, BDJ20, GS22, HMX21, KGW^{+}21, Par22c, QD23, Vak25, ZLF^{+}22, RSB^{+}21]. Open-loop [RGS^{+}20]. Open-Source [XXAD21, Kam21, HMX21, KGW^{+}21, ZLF^{+}22]. OpenCAL [RFG^{+}21]. OpenFOAM [MWY21]. OpenLB [KKA^{+}21]. operational [UHLZ24]. Operator [ACLP24, UHP23, AAU25, ADFP23, BHD^{+}21, BHDV22, CBLZ24, CYQW21, CYKL23, DMRV24, EIJH21, FM25, HJXW23, JYHS23, KEUY23, LKC^{+}24, LL25b, LL24a, LZS25, LP22c, NZ24, OF20, PL25, PS25b, Sch21, SZ22b, SIAA25, VLD^{+}24, WXF22, WHT22, WWL22, WSCP24, YWL23, YKMZ21, YTT^{+}22, ZWX24]. operator-splitting [WXF22]. operators [CV20, DDNS21, DZ20, DE21, GPA21, IK24, KSB25, KP20, KKS25, LQW22, MGD23b, PV20, Pon20, STO22, SW21, YSB^{+}24]. optical [WZY23]. optics [HWY23, HGD25]. Optimal [ACLP22, BM24a, Boł22, CY23, CLQL25, CYQW21, DG22a, EGP20, FLW20, GWYR22, HHM21, He23b, HZ25, HAC23, KC20, KN20, KVV23, LL22a, NM23b, NKPA^{+}25, RC25, SL24, YLQ22, YWPS24, AFS24, BHS20, BHS25, BKW22, BBI^{+}22, BF21, BGZ22, BJyST24, BFM^{+}25, BC20b, CSB^{+}21, CM21, CM23, CDM24, CSCY24, DSR23, DO23, DHVT20, DH23, ELN^{+}20, FF24, GS22, GLS25, GLF25, GWN21, HLZ24, HYZZ24, KUK20, Kar24, LLSY24, LLSY25, LCW22, LX24a, LWL25, LAL^{+}24, LC23b, LC21, LQY22, LSZ24a, LYL25, MI20, MN21, NE22, PWBS25, PR20, RNC25, RR23, RGS^{+}20, SZ22b, SK25, SGR^{+}23, SW20b, TZ25, TS21, TCA^{+}22b, TKK23, WLX22, WM23, WZD24, XP22, YY20a, ZCL20, ZS25, ZL20b, ZJJ20, ZC22, ZZH23]. Optimal-order [KC20]. Optimality [Axe20, BC20b, GB23, KN20]. optimisation [AMPB^{+}22, MJVH24, PNT^{+}23, PTNR24, ZKTK21]. optimisation-based [PNT^{+}23]. Optimised [CM20a]. Optimistic [BGZ22]. Optimization [PBR24, WWZ^{+}24, AAA25, BV21, BGG^{+}23, BKK^{+}23, BLM^{+}24, BGD22, CK23, CLH21, CEB24, ERHL22, FYM23, FMP^{+}24, GPRS22, IJS^{+}24, KPO^{+}21, LQZ25, LL22c, LML23, LHTZ25, NIT^{+}20, NNNT22, NNTCB25, OKSS21, Par22a, PWMC23, PDB22, RHA^{+}24, SMR^{+}24a, UuIUK21, VSG21, WFW^{+}24, WSO^{+}23, Wol20, XYT^{+}21, XXYL23, YS23, YYY^{+}22, YLSY24, ZYW20]. Optimization-based [PBR24, BGD22, LHTZ25]. optimizations [HLCW24b]. Optimized [Ahm24, AXLW24, ADB^{+}22, BTCV20, LCY23, WFMD20]. Optimizing [UBP^{+}25]. option [Che24, KHY25, LLU24, NS24, SLM23, TZ20, Yan25a]. options [AG24, ASSV20, CDLW20, CYKL23, GM22, MT25, RYAM21, Rou20, SMR24b, TC25, YZL20]. orbit [GM21a]. orbit-coupled [GM21a]. Orbital [LZ20c]. Order [BBD21b, KNBR22, QD20, AD20, AADMA23, AS21, AAC23, AKO21, Ade20, Adj22, AFS24, ANZZ23, ADB^{+}22, ACPT25, AZY22, AMM23, BNV22, BS23a, BZ20, BZ22, BF25, BBK21, BGH20, Ben25, BB23a, BM24a, BBM25, BCT21, BM21, BLMV24, BS23c, BGLZ24, BWMW23, BSRVC20, CT22, CBŁP23, CLZ23a, CCH24, CRL^{+}25, CM21, CG21, CCF^{+}21, CLL21, CLLG23, CK23, CCWS20, CS20a, CLZW20, CMY23, CXF24, CQG^{+}24, CYM25, CL21, CCGR24, CFG^{+}23, CCGRL21, CRF^{+}20, CDPL24, DD21, DPP21, DLN24, DG22a, DE21, Der23, Din21, DMX24, DWS20, DSL24, DWY^{+}25, DW20, DAS20, DC22, DSF24, EAM20, FSS^{+}20, FSW20, FWLD22, FHM21, FWBY25, FHZ22, FX22, FXLS25, GP25, GLL^{+}20b, GQ23, GSS24, GP23, GNB24, GHLS25, GDA^{+}22, GDŁW23, GCJ23, GQZ23, GCMS21, GYC25, GFXZ25, GHK25, Hak24b, HSR25, HUT25, HJM23b, HCZ25, HLX20]. order [HSZ22, HH23, HZR23, HHW21, HHL^{+}21, HLPW22, HAC23, HQL23, HHG20, JVBM23, JWZ22, JHG^{+}21, JZHZ22, JW20, JP25, KSS22, KUK20, KBR20, Kei21, Kim21, KC20, KWLW24, KBB^{+}21, KKS25, KAKK23, KSS^{+}20, KSSS23, KGK25, KC23b, LG23, LCY23, LL25a, LLD23, LQ20, LSH21, LWSZ21, LL22a, LXRQ22, LH23a, LHF24, LY25, LS20, LML23, LYH24, LSF25, LS25a, LX24b, LS21, LY21, LSY^{+}24, LWFS24, LXL25, LZLN25, LTYZ25, Liu25a, LdO21, LN22, LVDDNT25, MWYY22, MFZX23, Ma25b, MG25, MI20, MRN^{+}20, MW20, MJVH24, MA22, Men22, MJR20, MKP23, MN23b, MLL25, MZZ^{+}25, Nde25, NSMR25, NDC22, Ngo25, NF25, NAE22, OS24, OCCM21, PS22, PHL21, PZL23, PD25, PNT^{+}23, PTNR24, Pta21, QH22, QQX21, QFL23, QWC^{+}23, QXCG20, Rav20, RWZG24, Rou20, RR22, SDM21, SMDC23, SHM21, SV21, STM21, ySRC22, SHD23, SJ21, SLM23, SSD23, SCdH21]. order [SKG24, SSK^{+}23, SR21, SR24, SvV20, SY25, SIAA25, ST20b, SC21b, SWZ23, SQGZ24, TZ25, Tan25, TLZF20, TA22, TQX23, TS21, TRL23, TBS25, WCL20, WW20b, WZZ20, WH21, WZY22, WZ22, WZTZ23, WCWZ23, WGC23, WLL23, WGeL24, WDYZ24, WYW^{+}24, WWYL25, Wan25, WYLL25, WWZQ25, WF25, WL20, WKC21, WARX23, Xen20, XFLZ22, XZ23, XZ24, XHH^{+}23, XSZ^{+}23, Xu24, XHL24, YSGD20, YCZ^{+}21, YTK21, YZT21, YHZ22, YL22, YS23, YL25, YLZ^{+}24, Yem24, YHF24, YFK20, YFK23, YM23, Yu23, ZM20, ZR23, ZCYL24, ZLT^{+}21, ZO21, ZFL22, Zha22c, ZXGF22, ZYTX22, ZLJ^{+}22, ZXXX23, ZAH23, ZZS23, ZL23b, ZSF23, Zha23, ZL25, ZZ20b, ZLL23, ZHZ23, ZZY^{+}24, ZZZD24, ZLZ25b, ZLJT20, ZCA23, ZW21b, ZMS21, ZQZ20, ZZFL21, AMV24, BB23b]. order/ [HHR24]. ordered [NNTCB25]. orders [CZ23]. ordinate [SSW20]. ordinates [SLC^{+}24]. organisms [CM24, SQC23]. orientation [SGAC^{+}20]. oriented [BBK21, BBI^{+}22, BDE^{+}24, BMvdZ21, CNT23, DSS20, DBR21, ELN^{+}20, ELS24, HIP23, JP25, Mal22, VR20]. orifice [JHS24]. original [HLJ25]. Orthogonal [LHH22, AS21, BDF24, Fur25, GCC20, LC23a, LLX23, LYXC21, PLR21, SM22a, TQX23, WYHQ24]. orthopedic [YKNS23]. orthotropic [AAC23, TPA24]. OSC [YZT21, ZYTX22]. oscillating [GHS23, NIT^{+}20]. oscillatory [AMS24, GCJ23, KP24, LZL^{+}25, WZTZ23]. Oseen [Agg20, BLF23, MBG24, PG21, SK25]. osmosis [CMPR23, LLN24a, eLLAN25]. outline [LP22b]. outlook [KEUY23]. Output [HJ23, HJ25]. outs [UHP23]. oval [MAT^{+}20]. over-determined [LNT22]. over-integration [YQZL22]. over-penalized [LG23]. overcoming [Zha22e]. Overlapping [CPS21, ZSGT24, BTCV20, DLC22, DZ24, GLC24, JRZ21, LLMC25, MGD23b, ZS23b]. overlying [KARS23, RGS24]. overset [ZP24]. Overview [RHR^{+}21]. oxide [YLQ22].

p [WXF22, XWXF23]. p-adaptive [WXF22, XWXF23]. Pacific [APK^{+}25]. package [HMX21, KWS21, XMQ^{+}21]. packaging [CRDS20]. Padé [YAA25]. Pages [Ano20-50, Ano20-58, Ano20-65, Ano20-61, Ano20-66, Ano20-68, Ano20-59, Ano20-63, Ano20-64, Ano20-49, Ano20-51, Ano20-53, Ano20-60, Ano20-62, Ano20-56, Ano20-57, Ano20-54, Ano20-55, Ano20-67, Ano20-52, Ano21-52, Ano21-59, Ano21-53, Ano21-50, Ano21-64, Ano21-60, Ano21-56, Ano21-63, Ano21-54, Ano21-67, Ano21-49, Ano21-57, Ano21-55, Ano21-51, Ano21-61, Ano21-65, Ano21-58, Ano21-66, Ano21-62, Ano22-52, Ano22-57, Ano22-70, Ano22-55, Ano22-59, Ano22-69, Ano22-51, Ano22-58, Ano22-56, Ano22-49, Ano22-71, Ano22-67, Ano22-65, Ano22-66, Ano22-54, Ano22-68, Ano22-61, Ano22-62, Ano22-60, Ano22-63, Ano22-64, Ano22-53, Ano22-50, Ano23-57, Ano23-56, Ano23-71, Ano23-65, Ano23-72, Ano23-70, Ano23-52, Ano23-54, Ano23-51, Ano23-50, Ano23-59, Ano23-63, Ano23-60, Ano23-66, Ano23-55, Ano23-58, Ano23-53, Ano23-49, Ano23-61, Ano23-62, Ano23-64, Ano23-67, Ano23-68, Ano23-69, Ano24-66, Ano24-49]. Pages [Ano24-67, Ano24-59, Ano24-52, Ano24-56, Ano24-61, Ano24-60, Ano24-63, Ano24-58, Ano24-48, Ano24-53, Ano24-55, Ano24-62, Ano24-54, Ano24-51, Ano24-47, Ano24-68, Ano24-50, Ano24-57, Ano24-64, Ano24-65, Ano25-58, Ano25-52, Ano25-60, Ano25-56, Ano25-67, Ano25-54, Ano25-62, Ano25-48, Ano25-50, Ano25-55, Ano25-57, Ano25-53, Ano25-51, Ano25-63, Ano25-59, Ano25-47, Ano25-49, Ano25-61, Ano25-65, Ano25-68, Ano25-69, Ano25-64, Ano25-66]. Pagosa [RCSM23]. Painlevé [Waz20]. pairs [AJLM24]. Palabos [LMK^{+}21]. pancreatic [ASF^{+}25]. panel [LL22c, SP20]. panels [LYH24, MMP21]. papers [ZSH^{+}21]. parabolic [AD21, ACDL24, AGLR21, AMN22, AAU25, ANZZ23, APS21, BAKR24, BB23a, BDF24, BCT21, BCM25, CWX25, CCX21, CKFR24, CBLZ24, CM22, CKBK24, CNT23, DN22, DLC22, DP25, DZ21, ELS24, FK21, FG25, FH25, GQ23, GLLZ23, HZJ^{+}22, HTBMC20, HGW18, HA23, IQS20, Ila20, JS25, KPK25, KLY25, KBB^{+}21, KPJ22, KSJ23, KK25, LK20, LXZ23, LAL^{+}24, LZ20e, LN21, LXL25, LMST20, Man24, Ngu20, PH25, PKP21, PSI^{+}24, QGQ22, RDF24, RS21, SMDC23, SL21, SL24, SU23, SW21, SNHH21, TPOT20, UR24, WYHQ24, Wan24, Wol20, YS23, ZCL20, ZXGF22, ZZH23, ZGC24, ZJ25, ZX22]. parabolic-hyperbolic [BCM25]. parabolic-parabolic [LZ20e]. Paradigm [HC22a, DFHT21]. Parallel [BFI^{+}23, DS20, Fia21, LMK^{+}21, WLQ22, AHM20, BF25, BHPS24, BW21, BGH20, CGHJ24, CPBF21, ÇM20b, DZ24, GHMD22, GDA^{+}22, HDM24, HMX21, JZS23, KTXP21, LDT24, LD22, ŁWP^{+}23, PLVL21, TW23, WDZ21, WWC^{+}24, WZS24a, XHL21, XHL24, ZWOM21, Zdu21, ZS21, ZS23b, ZWX24, ZZS24]. parallelization [LHH22, XRS25]. Parallelized [AGM^{+}22, ZCXG20]. Parameter [AHLZ24, BDJ20, MMN21, BKR24, BGD22, CGGY23, FC24, HHW21, JJZ21, KM22, LCY23, LM24, MJVH24, MVHÖ23, PTNR24, SH22, Tan25, WZY23, YEO21, ZDWR22, Zha24, ZWS22]. parameter-adaptive [MJVH24]. Parameter-free [AHLZ24, LM24, Zha24]. parameterization [MHHR20, ZJZ24]. parameterized [AXLW24, KV22, MWRA22, SC21b, YL25, YX21, ZXGF22, ZZ20b, ZYW20]. parameters [BHDV22, CLY25a, CS25, FLW20, IJS^{+}24, NTP25, RYAM21, RHB25, RR23, SB21, SAVHB22, SLUW^{+}21, VT21, XW24a]. parametric [AH25, BX20, BMP^{+}24, CL24b, DSR23, EIJH21, GHSR23, JVBM23, Kri20, LL25b, PKPD^{+}24]. parametrization [HK24]. parametrized [ABR20, CSB^{+}21, KBR20, KNBR22, SDM21, UuIUK21, WARX23]. Parareal [ZJ25, GHS23, MZHJ24]. Parareal-based [ZJ25]. parasitic [Bel24]. Part [JMR24, BGK23, BM24a, GP22a, NW24b, WN24, GGRBRG20b, GGRBRG20a, HH21, HH22, MKM24a, MR22, Mit22, WZC^{+}20a, WZC^{+}20b, ZKBC21, ZBKC22]. Partial [NSRI24, Abd25, ABFPA24, BLLM25, BCCR20, CMM23, Dai22, FS21b, GYL25, GO20, GKHU22, He23a, HCY^{+}24, HM24, HZW21, LL25b, LZ25b, LMLZ23, LZS25, LZ25c, MH25, MN23b, NF25, OP24, PV22, PKPD^{+}24, PSI^{+}24, SMDC23, SB25, SR23, SZ21, TFWX23, TBO20, Ull20, WCL20, WFX24, XS25, YK24, YOF20, YLCC21, YWPS24, ZGW23, ZA23]. partial-integro [SZ21]. Partially [ADD22, ČPPŠ20, JMS^{+}24, MGBN22, NTP25, WL23]. Particle [ABC^{+}22, APK^{+}25, ASH^{+}21, CCG^{+}23, JWW^{+}21b, LGQZ21, Lin24, LAB23, MRS21, RLF21, SADW21, SGAC^{+}20, SHT^{+}24, XJ21, YD21, ZZZC24, ZL20b, ZHZ^{+}22, dSG20]. particles [ANH^{+}21, DKH22, HKJ23, HLH^{+}20a, PWJ^{+}24, SGJJ20, SCD23, SLX^{+}23, VJ24, XGL21, ZYM24]. particular [ABFPA24, SAVHB22]. particulate [ENF24]. partition [AM22, Du24, FAM25, LQF25, RDF24, HNLPP20]. Partitioned [XZ25, GPL21, JJOG20, LXRQ22, WZ21, YK24]. partitioning [GY25, YD21]. parts [NERB^{+}22]. passing [WP25]. passive [DT23, Kri20, KOMK20, LVDDNT25, PBR24, YL20b]. patch [BMST20, BGKV20, EFK^{+}23, FKKV24, KV22, SS20, SYHN24, TT25, WNF^{+}25, WP25, YH21a, ZJZ24]. patches [YJZ20]. path [JICRMG^{+}20, MZCK25, OKSS21, Nde25]. path-conservative [MZCK25, Nde25]. path-tracking [JICRMG^{+}20]. Pattern [KYY^{+}20, ZZZ24, Bar22]. patterns [CML^{+}25, CBK24, SOPH25, SKY25]. PCA [CEB24]. PCA-based [CEB24]. PCCU [Nde25]. PCCU-ADER [Nde25]. PCHPDDM [JRZ21]. PCM [JFS20, JFS14]. PDE [ASSV20, BV21, CSB^{+}21, CKBTS24, CKBK24, ELN^{+}20, GP22a, GPRS22, HK20b, HK20a, HLZ24, IJS^{+}24, Kar24, KTXP21, LNF24, LP22b, NES23, SW20b, TC25, WH25, YLSY24, YLWZ20]. PDE-based [LP22b]. PDE-constrained [BV21, CSB^{+}21, HLZ24, YLSY24]. PDEs [AAU25, BBI^{+}22, BMVA23, BB23b, BBD^{+}24, BX20, BMP^{+}24, CAD24, CCX21, CD20, CSS23, DSS20, DN22, Füh21, FG25, HOY24, HA23, Ila20, KP24, KN25, eLLAN25, LC21, LC22, LZLN25, LSPRV24, Lop20, LP22c, MCW25, MB24, MCK^{+}23, NH24, SOPH25, SDM21, STM21, SU23, WZT23, XXAD21, Zha22b, ZXGF22, Zha23, ZZ24c, ZMZW24]. Penalized [BBS21, KPK25, LG23]. Penalty [NAS20, BLRRS20, BJyST24, DSF24, EFFR22, GPRS22, HBC^{+}21, HHW21, LM24, LTXZ24, MH22, RZ21, Zha22c]. penalty-projection [DSF24, MH22]. pencil [GF21]. penetration [Abd24]. pentatope [WN24]. pentatopes [WFMD20]. perfectly [KP22d, YCL20]. Performance [HJ25, MMGS25, MC25, ZDLC22, BEG^{+}21, DdSP21, Fan23, GM22, IBY^{+}24, Kra20]. performances [XJY^{+}20, ZCZ^{+}21]. perfusion [BGG^{+}23]. Peri [AP25]. Peri-ultrasound [AP25]. peridynamic [Abd24, Ale23, BDP23, Ha20, LP22a, SGZ^{+}23, SMS23a, WDYZ24, YMDW25]. perimeter [NRV22b]. periodic [DHW20, KMN25, PWWJ23, Vak25, Wol20]. periodicity [ZLQ23]. permeability [Cha24]. permeable [LL20a]. permeation [ZBKC22]. permutation [PXLY23, XL25a]. Perona [LHC23]. personalized [YMTS20]. perspective [AMP20, HLX22]. perturbation [LL22c, Men22]. perturbed [BHJ22, CGGY23, CMR22, CSS23, CS25, CG22, DN22, GO25, HM24, KSS22, LHN24, LY23, LX24b, LZ23, MZ23a, SG23, STO22, SZ24a, SSD23, SN20b, WML21, WS25, Xen20, ZZ23a, ZZ24a, ZZ25b, ZMZW24]. Peterlin [HJM23b, YJL22, ZYD24]. petroleum [DG22b]. Petrov [AD20, BS23a, BDG21, CO21, CRM22, CM23, CHD24, HHR22, HRvdZ20, Lee21, LCW22, MCW21, Nie23, PS25a, RZC21, Sto20, ZXZF20, ZSW23]. Petrovsky [DL25]. PETSc [JRZ21]. PFASST [BMK21]. PFEM [LFF^{+}24]. PFEM-FEM [LFF^{+}24]. PFM [Alf24]. PGD [PKPD^{+}24]. PGHW [HJP23]. phantom [EGF^{+}23]. Phase [BC20a, CFG^{+}25, HCY^{+}24, LYP^{+}22, SMŁWL25, ASGGRG23, ANWW21, ASH^{+}21, AEZ24, AZY22, BLLY24, BMH22, Bal21, BAM^{+}21, BBDM21, BDS24, BM23, CLLL20, Cha24, CLH21, CXX^{+}25, CCGRL21, CQS25, DBH^{+}21, DMŠ22, EBDY22, FZW^{+}25, FL22, GNZ24, GML^{+}20, HKH^{+}24, HXW23, HHY^{+}22, HMŁW^{+}24, JWW21a, JWL20, JW20, JFS14, JFS20, KP22b, KHH25, LH24, LHC^{+}22, LWSZ21, LAX22, LQYJ23, LaZMY25, LJ24, LYRW21, LQZ23, LLW^{+}23b, LLMC25, MAT^{+}20, MBSEA24, NRV22a, NNTCB25, PHL21, PL24, QZH23, QP24, RST^{+}22, SK23, SYHN24, SY25, TK20, TWH^{+}22, Vak25, WYS^{+}20, WZY22, WXF22, WYP23, WMMH25, WHH22, WSO^{+}23, WYT23, XKXL22, XWXF23, XXYL23, XZS^{+}23, YK20, YZJ^{+}22, YWT22, Yan24, YCWS20, YWZ^{+}23, YZW^{+}22, ZP24, ZKBC21, ZBKC22, ZL20b, ZTW22, ZWT23, ZL23b, ZGW23, ZHZ23, ZFQ24, ZK21, CBLZ24]. phase-change [ZK21]. Phase-field [CFG^{+}25, LYP^{+}22, ASGGRG23, ANWW21, BMH22, CLH21, DMŠ22, GML^{+}20, HKH^{+}24, HXW23, HMŁW^{+}24, JWW21a, JWL20, LAX22, LaZMY25, SYHN24, SY25, WZY22, WHH22, WYT23, YK20, YZJ^{+}22, YWT22, Yan24, YZW^{+}22, ZKBC21, ZBKC22, ZTW22, ZWT23, ZHZ23]. phase-field-based [YWZ^{+}23]. phase-wise [NRV22a]. phases [BM23]. phenomena [HRvdZ20]. phenomenon [KP21, Zha22b]. phonon [PCLZ25]. photoacoustic [AOL22]. photonic [CLH21, GHLS25]. photonics [Bot20]. photovoltage [FKR21]. Physical [LDW^{+}24, FCF22, PL24, XZ23]. physical-informed [PL24]. Physics [CGGY23, FZW^{+}25, GYL25, GI20, GZAR23, HSR25, QWLQ25, TFWX23, TH25, UBP^{+}25, CGHJ24, CLY25a, CGI^{+}22, CDG^{+}23, DSR23, GY25, HHL22, JAHM23, LWG^{+}25b, LWG^{+}25a, LCC^{+}24, MJZ25, MSBO23, NES23, WZL25, XYS^{+}24, ZL23a, ZL23b]. Physics-based [GI20, JAHM23]. Physics-constrained [TH25]. Physics-Informed [UBP^{+}25, CGGY23, FZW^{+}25, GYL25, GZAR23, QWLQ25, TFWX23, CLY25a, CGI^{+}22, CDG^{+}23, GY25, HHL22, LWG^{+}25b, LWG^{+}25a, LCC^{+}24, MJZ25, NES23, WZL25, XYS^{+}24, ZL23a, ZL23b]. PI [CCGV22]. PI-type [CCGV22]. Picard [ZZZ25, ZST22]. PID [ASF^{+}25]. PIDE [LZY24a]. Piecewise [AMS24, BL20, CCH24, HCHM25]. PIES [BZ21]. piezoelectric [AA21, BM24b, DCSAM25, LYXC21, SS22]. piezoelectric-viscoelastic [DCSAM25]. piezolaminated [MKM^{+}24b]. PIKFNN [XFX^{+}24]. PIML [FZW^{+}25]. pin [DBH^{+}21]. PINN [CLXZ25, EMSA25, FWLD22, GY25, HCY^{+}24, KK25, WZL25]. PINN-DTC [GY25]. PINNs [HCHM25]. pipe [HLCW24a, JHS24, TMML23]. pipelined [WWC^{+}24]. Piscounov [DL25]. Pitaevskii [FHZ22, MZGM25, WTQ^{+}20]. pitcher [HA23]. pitching [PSSW20]. pitting [SHR25]. pivoted [ADD22]. placement [Gra21, MHHR20]. Plain [HIP23, CBK24, FP20, LZ25a]. Planar [BCK24, BGKV20, GKK^{+}20, LD24]. Planck [GIRB25, GDA21, HHR22, LSWY21, LD25, LHTZ25, PS25b, RKK25, ST22, WCW20a, YTH^{+}21]. Plane [YH21b, CGHJ24, CS23a, CWX22, LLZ^{+}25, MCZL20, Nag24, Oru25]. planetary [CMS23]. planning [BGZ22]. plasma [BNC^{+}23, WXH^{+}25]. plasmas [VJH23]. plasmon [LZA23]. plastic [BZ21]. plasticity [KSB25, PRN^{+}23]. plate [BLX^{+}24, CCGF25, DYC^{+}22, DD22, HYL^{+}25, HTPV24, KHD24, LZJC21, LC23a, LXG^{+}24, MCD23, MCD24, MZ23b, NNNT22, NP24, XLLR25, YZZ20, YX21, ZDT^{+}22, ZXG^{+}24]. Plateau [HL22]. platelets [AA21, HTPV24]. plates [Abd24, AP25, CWZ21, CHL22, HWZQ23, HTPV25, KAKK23, KKJ^{+}22, LLZ^{+}25, LYXC21, MNB23, NP24, SFR21, SS20, WWZ^{+}24, YLFL20, ZD21]. platform [AEF20, DdSP21]. plicRDF [AFBM25]. plicRDF-isoAdvector [AFBM25]. ply [SS20]. PMCHWT [BD23]. PMLs [Li24a]. POD [AGM^{+}22, BSRVC20, CSB^{+}21, CS20a, HK24, JMN22, SDM21, SR21, SR24]. POD-based [AGM^{+}22]. POD-reduction [CSB^{+}21]. POD-ROM [JMN22]. Poincaré [FM25, PV20, VLD^{+}24]. Point [XMQ^{+}21, AGA^{+}21, BHJ21, BWZ22, BWZ23, CSS20, CG22, DWAM25, Fur25, GLL20a, GdFFF25, HK22, KKJ^{+}22, LL20b, LWG^{+}25b, LZW21a, LZW21b, LLW23a, LZL^{+}21, LLU24, LQF25, Ma25a, MS20a, MJWD20, MN23b, NK25, PWBS25, PS25a, PH23, RSAG^{+}24, SM22b, SS22, SH22, SHXZ24, Xen20, XW24a, XL20, XHY21, Yan24, YCL20, YLCC21, ZDLC22, ZJCG21, ZYW20, dSG20]. point-weighting [LWG^{+}25b]. points [LZLN25, Oru25, WFX24]. Pointwise [KWLW24, WC21, Adj22, BJyST24, HAC23, dFLdSLM22]. pointwise-in-time [HAC23]. Poiseuille [HLCW24a, KARS23, RGS24]. Poisson [AG25, WCW20a, BZ22, BJG23, BBM25, DDNS21, FHM21, GIRB25, GDA21, GHS23, HSZ22, HLPW22, LSWY21, LHSF24, MW20, NBF^{+}24, OAD24, RZSL23, RMBC24, ST22]. Poisson-like [NBF^{+}24]. Poisson-type [RZSL23]. Polar [MM21, CFT^{+}23, LMP23, RTA20]. polaritons [LZA23]. poles [ZXG^{+}24]. policy [RZ21]. pollutant [CML^{+}25, RWZG24, WSGT21]. polluted [DYZ22]. pollution [AADMA23, HD21, MS23]. pollution-free [MS23]. polygon [GYC25, Kra23]. polygonal [AN20, AFM^{+}24, ASV20, BHDV22, CLZN24, DD22, GENRB21, KSJ23, KTY24, LZJC21, LHF24, LSWY21, LGL24, MWY23, NP24, NFS24, PK23, ST22, TKK22, WWS23, WJ25, WF25, WW25, YZZ20, ZPK22, ZZW23]. polygons [DWS20, OR22, ZLJ^{+}22]. polyharmonic [AMV20, CKL25, DMRV24, ZCL20]. Polyhedral [MP23, SBMS22, BDMX21, DV22, DFGS24, HMF24, LD23, ST22, ZP24, dVDR20]. polyhedrons [ZLJ^{+}22]. polymer [AMPB^{+}22, DP20]. polymeric [AEF20, BFRA21, WWCJ23]. polymers [HJTW25, LQYJ23]. polynomial [ABFPA24, BNSZ24, BFI^{+}23, CLZ23a, DGN24, GK21, JZPG21, KVV23, LCY23, Oru25, TFWX23, XCFC21, ZH23b]. polynomial-augmented [CLZ23a, Oru25]. Polynomially [ES20]. polynomials [HZR23, LYXC21]. polynomials-Ritz [LYXC21]. polytopal [DDG23, FMP^{+}24, GQZ23, SNHZ25]. polytopes [ST23]. Polytopic [AFV22, YH21a]. pool [Cao21, LAB23, WZXN21]. population [DYZ22, Kar24, TSW^{+}25]. populations [dABCC25]. pore [ZDW^{+}20]. pores [ZD21]. poro [BZDQ21, CLZN24, JM23a, NK25]. poro- [NK25]. poro-elastodynamics [CLZN24]. poro-hyperelastic [BZDQ21]. poro-thermoelastic [JM23a]. poroelastic [ANWW21, HZ24b, WR23, ZR22]. poroelasticity [BHK^{+}24, BBD21b, BAB^{+}20, CLLL25, GPY22, HZ24c, LR23, VT21, ZZCM20]. poroelasticity-Forchheimer [HZ24c]. poromechanics [BER21, FCFW21, FCF22]. porosity [CLRT24, LXRQ22, LZBC24, LLY25, NKPA^{+}25, QQY^{+}23, SRA20, WZ21].

Porous

[RSB^{+}21, AKA^{+}21, AZL20, AGA^{+}21, ÁCS22, AEZ24, AFV22, BB24, BBDM21, BLM^{+}24, CQS25, DK24, HANA24, Hoa21, HLZ24, HTPV24, JWW21a, JSBZ22, JSZ24, KHD24, KARS23, KRESA21, KGW^{+}21, KS25, LHC^{+}22, LWFS24, LR22, MKM^{+}24b, MS20b, OAA^{+}22, PND^{+}24, PCL25, RGS24, SSY22, SYZ23, SLX^{+}23, SRA20, SNL23, WW25, YY23, YSGD20, ZLL23]. port [BMS22]. port-Hamiltonian [BMS22]. posedness [CCF^{+}21, CG22, FRW22, GI21, LLWX25, MKM24a, SSY22, YHL25, dCLW25]. Positive [IQS20, BDGS25, MS20a, SvV20]. positive-definite [MS20a]. Positivity [CQS25, QD23, QFL23, RWZG24, DYZ22, LLW23a, LHTZ25, NCRGVR21, ST22, TZ25, XZSY23, YYZ22]. Positivity-preserving [CQS25, QD23, DYZ22, LLW23a, LHTZ25, ST22, XZSY23]. post [GQ23, LZ23, WZS24a]. post-processing [GQ23, LZ23, WZS24a]. Posteriori [HLW23, AA25a, APS25, BHS20, BHS22, BB24, BZE24, BKR24, BS21, BX20, BDE^{+}24, BHL^{+}23, CHL22, CDGM25, DFK22, DH23, FLOQ21, Jha21, KPS23b, Lee24, LZ21, LZ25a, LZ25b, LN21, Mal22, MCW21, MZ23b, MS20b, MCV24, Šeb24, WBY23, XZT23, XXLS23, ZLHC24, ZPK22, HM24, PR20, LXZZ24, MMC21]. postprocessing [HLX20, ZX22]. potential [CNT23, HHL^{+}21, HLN24, LMBB24, LZ20c, LGQZ21, Li24b, Li25, MC24, NNT22, PS21, RZŠ21, WTQ^{+}20, YWT22, ZWT23]. potentials [ARV23, ZXZ25, aKST22]. Powell [BEIL21]. Power [BP20, DKV21, DKMS22, YDY^{+}23]. Power-law [BP20]. powers [CV20]. PPM [ZHPW23]. practical [YWT22, Yan24]. Prandtl [HKJ23]. pre [CLXZ25]. pre-train [CLXZ25]. precise [Bar21]. precision [MXLY24, MGA22, WCC25]. Preconditioned [LZ20a, LZW21a, LZW21b, ZZS23, ASB25, CDLW20, EH23, FK20, PS20, TS24, XLS22, XLSF25]. preconditioner [AM23, ACPT25, Axe20, CL20, CCWC23, FCF22, HQL23, Krz22, NHCJ22, SM22b, SH22, XW24a, ZX24, Zhe22, ZZB20, ZYW20]. preconditioners [BFI^{+}23, CSS20, CPS21, FH25, JRZ21, KC20, KCC20, PQSM21, SS22, SvV20, XW24b, ZZL22]. Preconditioning [CG25, KK21, LP23, YL25, AHLZ24, Ahm24, BHJ22, BHZ20, BW21, DTV20, EJR24, EIJH21, HKM21, LZ21, LDH24, LST22, MS20a, TH23, VADG22, Wat22, XL20, ZGL23]. predator [Li20]. predefined [ZWW^{+}23]. predict [QP24]. predicting [BBT25, CLW^{+}24, YSGD20]. prediction [FWLD22, GOP25, Hag24, NKPA^{+}25, SGR^{+}23, VT21]. predictions [CM24]. predictive [AGM^{+}22]. Predictor [KMS20, LS20, LZ20d, SAR20]. Predictor-Corrector [KMS20]. Preface [Ano21-68, GZG20, GRYX24, Lop20, PRS24, RLS24, Yua24, ZSH^{+}21]. prefilmer [ZDLL22, ZLD^{+}23]. Preisach [VGA^{+}22]. Preliminary [BMH22, DSR23]. premixed [SB21]. prescribed [HJT25]. presence [GO20, LYH24, RST^{+}22, SQC23, TMML23]. preservation [LHC23, RGL21]. preservative [KS21, LZS23]. preserving [AD21, ASGGRG23, AJ23, AG25, BCT21, CG21, CXX^{+}25, CSI23, CQS25, DG22a, DZ22, DLW24, DL25, DWS20, DEW25, DYZ22, DTW23, FZH22, GGT24, GFXZ25, HZW21, HLJ25, HGW21, JH24, JMS^{+}24, KWLW24, LF24b, LLW23a, LRZ23, LHTZ25, LTYZ25, LD23, MZGM25, MJZ25, PL25, PBR24, QD23, QFL23, QS20, RWZG24, SSW20, SZ22b, SQ24, SX25, ST22, SY25, SLYL22, TZ25, TSS24, TRL23, WFH^{+}20, WH21, WXF22, WGeL24, WWZQ25, XZ23, XDL^{+}25, XZSY23, XSL24b, YTK21, YKK25, YLFZ25, YHF24, YYSL20, ZWMF24, ZY23, ZS23a, ZXXX23, ZDW23, ZW25, ZNH24, ZSY22]. Pressure [RL23, AA20, BWZ22, BWZ23, sCpLYhG23, HHW21, JWW21a, KK23b, LM24, LYAZ20, LSF25, PS22, RWaZ23, SK25, TBS25, WWS23, WLJF25, YHZ22, ZR23, ZZ20b, ZPC22]. pressure-correction [LYAZ20, WLJF25]. pressure-driven [AA20]. pressure-projection [RWaZ23]. pressure-robust [LM24, LSF25, WWS23, YHZ22]. pressureless [JM20b]. prey [Li20]. price [LLZ20]. Pricing [LLZ20, CDLW20, Che24, GM22, KHY25, KN25, LZ20d, LLU24, LSPRV24, MT25, NS24, RYAM21, Rou20, SLM23, SU23, TC25, TZ20, YZL20, Yan25a]. Primal [GIRB25, OAD24, WW20a, ABC^{+}21, APPS25, HKM21, KW20, MPVA25, MHHR20, Nie23]. Primal-dual [WW20a]. Primal-mixed [GIRB25]. primary [ZDLL22, ZLD^{+}23]. principle [Alf24, CS20b, NNT22, SY25, WXF22, WWZQ25, YKK25, YYSL20, ZWMF24, ZY23]. principle-preserving [SY25, WWZQ25]. principles [NW24b, WZL25]. printed [ZYM24]. prior [KW20, SHQL20]. priori [BHS22, DH23, Kei21, LZ25b, LN21, PG21, Par22c, ZZH23, ZLHC24, ZPK22, GNB24, BS23b, GSRG21, NAS20]. prism [WN24]. probabilistic [BMVA23]. Probability [CWC22, SS21, SS24]. Probability-conservative [CWC22]. problem [AD20, AN20, AMV23, AMV25, AL22, AFS24, AHP20, ARV23, AAL24, AA20, ALOR25, AZ24, APS25, BHJ22, BHK^{+}24, BKW22, BJG23, BMK21, BLRRS20, BS21, BBG23, BZY20, Boł22, BN20, BCM25, BLLZ24, BRRT21, BJyST24, BCM22, CGGY23, CCGI23, CLR23, CLRT24, CM22, Cha24, CLXZ25, CZ20, CFT^{+}23, CJ25, CV20, CRF^{+}20, CG25, DD22, DO23, DHVT20, DKL24, DH23, DZ24, FC24, FWBY25, FHH23, GP23, GOV20, GGH25, GWN21, GS23, GGRBRG20a, GYC25, Gut20, GO23, HHJO20, HZJ^{+}22, HJM23b, HDM24, HL22, HTDT23, HZ25, HKM21, HYJ20, HYW^{+}24, JZ24b, KSS22, KS21, KP22d, KPS23b, KGJN24, KLY25, KBW23, Krz22, KKNR21, KGK25, KS25, LS22a, LSV21, LSV24, LL20b, LYAZ20, LD21, LCW22, LXY23, LL23, LY24, LWG^{+}25a, Li25, LSF25, LC23b, LS21, LC21, LCH23, LZS23, LS23, LSZ24a, LLWX25, LdO21, ŁSSP23]. problem [MCD23, MGD23a, MZ23a, MCD24, MXLY24, nMjWlS24, Mal22, MBG24, MMPR20, MLS22, NR22, OAD24, Oru22, PG25, Par22c, PP24, Pup25, QQX21, QV22, QD20, RL23, Rob25, RR23, SD25a, SG23, Šeb24, SL20, SL21, SLW22, SCG23, SJ23, SSD23, SK25, SR21, SW20b, SWC24, SA24, Tad23, TZX23, TS21, TPOT20, TH21, TVKN21, TTKV20, Tür20, Vab22, VT21, WS21a, WML21, WDZ21, WZ21, WLX22, jWC23, WWS23, WZD24, WCS24, WQ24, WXH^{+}25, WXCL25, WWZX25, WF25, WLW^{+}22, WCH23, WS25, XL20, XZ24, XXLS23, XL25a, YY20a, YA23, YS23, YEO21, Yem24, ZCL20, ZWZD24, ZH20, ZC22, ZR22, ZQY^{+}23, ZH23a, ZZ23a, ZLHC24, ZZ24a, ZGG^{+}25, ZL25, ZHH21, Zha22e, ZPK22, ZZZD24, ZXZ25, ZNMS23, ZZS24]. Problems [XXAD21, AS21, Abd25, ABC^{+}21, Adj25, AAA25, AA25a, ADD22, AMV21, AMV20, AMM23, AGPP22, AAM^{+}24, AAA24, BHJ21, BHPD23, BDV23, BMH22, BZ22, BHS20, BB24, BHS25, BM20, Bar21, BBK21, BGH20, BGK23, BB21, BDF24, BCF^{+}23, BGZ22, BZ21, BM21, BDJT21, BMST20, BGLZ24, BT22, BC20b, CY23, CCY24, CCW22, CLZ23a, CWX25, CSB^{+}21, CKFR24, CGM^{+}20, CK23, CLYH22, CHZ23, CDM24, CXF24, CSCY24, CCD25, CMR22, CL24b, CSI23, CHL21, CCWC23, CFTC25, CNT23, CGI^{+}22, DPP21, DLN24, DFSV22, DV22, DG22a, DSR23, DLC22, DFS21, DP25, DFK22, DBR21, DT24, DWS20, DSL24, DWY^{+}25, EFS24, EJR24, EBDY22, EP23, ELN^{+}20, ELS24, EFK^{+}23, EFFR22, FMS22, FHM22, FHM21, FSH^{+}23, FZDS22, FF24, GLS25, GLL20a, GQ23, GPRS22, GLÖ21, GLLZ23, GHSR23, GAM22, GCC20, Gra21, GO25]. problems [GQZ23, GB21, GB23, GXFT23, GHL25, HC23, Hak24b, HZ21, HLW23, HLMM20, HSR^{+}23, HLX20, HLQ24, HSJSS25, HM24, Hoa22, HKM21, HHL^{+}21, HLN24, HLW21, HC22b, HWC24, HMX21, HWWZ24, JKC24, JSSS24, JWR^{+}23, JK22, KWM22, KPK25, KC22a, KNBR22, KF22, KHM25, KJN21, KFD^{+}20, KGW^{+}21, KD23, KK25, KVV23, LLSY24, LLSY25, LK20, Lee21, LLR25, LLD23, LD24, LZ25a, LAL^{+}24, LQW22, LXH24, LX24b, Liu20, LN21, LMZ23, LTXZ24, LXZZ24, LMST20, LBPC20, LMMMP20, ŁMMMP21, LD23, LYL25, MCZL20, Ma25a, MSHB22, Man24, MI20, MGD23b, MS20a, MJZ25, Men22, MWY23, MN23b, MH23, MSBO23, Nag24, NT21, NRV22a, NBF^{+}24, OEASY24, Orl23, OCCM21, Pac24, PG21, PN24, PCC21, PKP21, PK23, PR20, PTNR24, QSL23, QWLQ25, QN20, RHA^{+}24, RL24, RS21, RL20, RHA21, SSY22, SD25b, STO22, SD21, SM22b, SA22, SMS23a]. problems [SZ24a, SL24, SBY^{+}25, SH22, SvV20, SC21b, SLC^{+}24, TCA22a, TjW24, TLZF20, TH25, Tor20, TZ21, TKK22, UuIUK21, UHLZ24, UHL^{+}24, VR20, VBDR24, Vis24, WYH20, WW20b, WW20a, WCH20, WXZW21, WL23, WGC23, WW24, WWC^{+}24, WJZ^{+}24, WCCC25, WZA25, WCLQ25, Wat22, Wol20, WLJF25, Xen20, XW24a, XZT23, XW24b, XLLR25, YH21a, YLW^{+}22, YZN22, YDY^{+}23, YHL25, YLSY24, YM23, Yu23, YW24, YWLF21, YPZ^{+}21, ZBCS24, ZLYL20, ZS25, ZZCM20, ZY21a, Zha22c, ZRL^{+}22, ZZL23, ZZH23, ZYC^{+}24, ZZ25a, ZCA23, ZCW24, ZW21b, ZWY23, ZZW23, ZQS24, ZYW20, ZX22]. procedure [PSDP22]. process [AO24, Cha24, CFG^{+}23, LCX^{+}20, SLUW^{+}21, ZYM24]. processes [DP20, Gie23, HM20]. processing [GQ23, LZ23, TRSM23, WZS24a]. produced [BNQ21]. product [AAU25, CBŁP23, CHL21, DS25, PH25]. production [GGRBRG20b, GGRBRG20a, HZW21, TSARR21]. profile [SGJJ20]. profiles [CS23b]. program [Boł22]. progress [Ano24-45, Ano24-46]. progression [PSDP22]. Project [BHL^{+}23]. projected [ABC^{+}22]. Projection [KBR20, Koc22, DSF24, EBDY22, FK20, GJS20, HC23, HJM23a, HJ24, HDWL24, LC23b, MH22, MBG24, MN23a, PG21, PCC21, RWaZ23, RL23, SWW22, WLJF25, ZSF23]. Projection-based [KBR20, Koc22, HDWL24, PCC21]. proof [Zha23]. propagation [AS21, AP25, AMM23, DG22b, DCSAM25, GCSF21, GI20, HLY22, JMS^{+}24, JWW21a, KG20b, LBPC20, SDJ^{+}22, XFX^{+}24, YWZ21, YA21, ZCY22, ZCYL24, ZDLC22, cZsCyC23, ZQY^{+}23]. propagations [SWZC23]. propelling [AAWJ20]. Proper [LHH22, AS21, LC23a, SM22a]. properties [AD21, ACDL24, Ale23, AZ24, Axe20, DP20, GM21b, MR22, Mit22, NS20, Nie23, YSGD20, YWL21]. property [CLQL25, LCX^{+}20, LCT22, NERB^{+}22, PBR24, XZ23, ZSY22]. property-preserving [PBR24, XZ23]. proposal [FSS^{+}20]. prostate [TK20]. prosthetic [NDB^{+}24]. protein [ZZW25]. provable [KTY23]. Proving [CLY25b]. proximal [LC23b]. Pseudo [AA20, AD21, ACDL24, DZ21, HGW18, MKS^{+}21, Wan24, ZvNH25]. pseudo-parabolic [AD21, ACDL24, DZ21, HGW18, Wan24]. pseudo-time [ZvNH25]. pseudoparabolic [KHM25]. pseudoplastic [AMPB^{+}22]. pseudopotential [ADL^{+}21]. pseudospectral [CBK24, JZ20, MLL25, WCW20a]. PSS [SM22b, ZYW20]. PT [ZXZ25].

PT-symmetric [ZXZ25]. Publisher

[Ano20-69]. pulmonary

[HQCW22, LHQW25]. pulsating [DKH22]. Pure [TZ25, WSO^{+}23]. Pure-positivity-preserving [TZ25]. purposes [SGSD20]. pursuit [Fur25].

quad [FHH23, JZ24b, WCS24]. quad-curl [FHH23, JZ24b, WCS24]. quadratic [DGN24, FF24, GGRBRG20b, GGRBRG20a, KPS23b, LLXZ23, YLW23, YLZ^{+}24, ZDT^{+}22, Zha22a, ZW20a]. Quadrature [JMM20, OR22, aKST22, CWZ21, CCGV22, FSS^{+}20, FJZC20, HRS24, Kap24, Leg21, LLZ^{+}25, LSY^{+}24, LP20, MH25, Tau22, WFMD20, ZA23, ES20]. quadratures [MB24]. quadrilateral [CLZ23b, CLL^{+}24, GKK^{+}20, GLC24, HWL^{+}20, JZPG21, JMM20, LLZ^{+}24, MrZQ24, Nie23, VJ22, WM23, WWZ^{+}24, XSL24a, ZS24]. quadrilaterals [nMjWlS24]. qualitative [KPS23a]. Quality [PH23, DDFZ23, Gie23, HTDT23, eLLAN25, SBMS22, UHP23]. Quantification [KAKK23, BLLM25, CL24b, CFG^{+}23, NNH^{+}25, SC21b]. quantifying [RHB25]. quantitative [LWZ22]. quantities [BMP^{+}24, GOP25]. quantum [PP24, YY20a]. quartic [BEIL21, Kap24]. Quasi [KSB25, SGR^{+}23, WS25, Adj25, BAB^{+}20, CZ20, EED^{+}25, GLL20a, GB21, GB23, Hop23, JSSS24, MN21, MD21, MS22, OSZ22, SMS23a, TMML23, ZR22, ZXQQ22]. quasi-boundary [CZ20]. quasi-brittle [SMS23a]. quasi-geostrophic [MS22]. quasi-linear [Adj25, GB21, GB23, ZXQQ22]. Quasi-Newton [KSB25, EED^{+}25]. quasi-Newtonian [OSZ22]. Quasi-optimal [SGR^{+}23, MN21]. quasi-optimality [GB23]. quasi-static [BAB^{+}20, Hop23, JSSS24, TMML23, ZR22]. Quasi-uniform [WS25]. quasi-variational-hemivariational [MD21]. quasicrystal [SKY25]. quasilinear [LNT22, LGL20, LZ20e, MGA22, NT21, SSdA20, ZZCM20]. quasineutral [AG25]. quasistatic [WXCL25]. quenching [KS21, LZS23]. quintic [SRD25].

Radial [LZL^{+}21, MH23, AM22, CBK24, DZZ^{+}24, GLÖ21, HCZ25, JKS25, JZXR24, JBKW23, KKJ^{+}22, LYRW21, LQF25, MJWD20, MFF25, Oru20, PWBS25, PS25a, QWLQ25, RSAG^{+}24, SRA20, TO20, XLCP20, XFX22, Yan25a, ZDLC22, XMQ^{+}21]. radially [FC24, WS21a]. radiation [AKK^{+}25, AAG22, ABL24, JSBZ22, LLXY23, LLXY25, MXLY24, PWC^{+}23, SP20, SBH^{+}21, XL25a, YSGD20, YCL20]. radiative [MBSEA24, PS20, SSW20, ZSL22]. radioembolization [PTM^{+}23]. rainfall [NS24]. Raman [LYLH25]. random [BLLM25, CSB^{+}21, CCX21, LL22b, TPOT20, TVKN21, YTH^{+}21, ZLQ23]. randomized [NCCD24, SW24b]. randomly [WDGC20]. range [SAR20]. rank [CL21, GHH20, MST25, RGL22, ZZL22, ZZ20a, Zhe22, ZMZW24]. RANS [TGL21]. Rapid [JMM20, LFGZ23, LGQZ21]. rapidly [TPA24]. rarefied [CLW^{+}24, LLZ^{+}23, WLZZ22]. RAT [MZCK25]. Rate [BBI^{+}22, He23a, HZW21, IJS^{+}24, LZ20d, LSPRV24, YY25]. Rate-optimal [BBI^{+}22]. Rates [KP20, BM24a, CM21]. ratio [JYHS23, XTZZ24, YZ21]. Rational [HLQ24, WM23, AM23, LL25c, LMLZ23, Oru20, YT22, ZWY23]. ratios [NM23a, ZTW22]. Raviart [AAD24, CS22, DGN24, GQ23, Sau23, YJ24]. Rayleigh [GWN21, Oru22]. RBF [CLZ23a, CD20, DMŠ22, FAM25, JM20a, JKC24, KC22a, LSRL24, LP22c, MVHÖ23, Oru25, RDF24, RL20, SSZ20, SS20, SZ21, Ull20, XSZ^{+}23, ZSK23]. RBF-based [LSRL24, RL20]. RBF-FD [DMŠ22, SZ21].

RBF-generated [JM20a]. RBFI [YLW^{+}22]. RBFNN [MH23]. RBFs [HHG20]. RBFWENO [AS20]. Re [CRPdT20, MGBN22]. Re-examining [MGBN22]. reacting [HEDT20]. reaction [AD20, AHOB24, AN23, ANN25, BHJ22, BKM23, BMK21, CY23, CCY24, CL24b, DSS20, DPP21, DR24, GP22b, GDŁW23, GCMS21, HZ21, HTDT23, HLMM20, HRvdZ20, HW25, IJS^{+}24, JSBZ22, JMN22, KSS22, KEUY23, KYY^{+}20, KC22b, KR24, LPK22, LLW20, Li20, LY21, LYZ23, LZ23, LYL25, MFZX23, MGD23a, MMGS25, MJZ25, MMN24, MMAK20, MMAK21, MN23a, MRP20, OTC24, PK25, RNMÖX24, RL20, RKC25, SSY22, SN20b, SIAA25, TKMK25, WML21, WLW^{+}22, WSS22, Xen20, XHY21, XSZ^{+}23, YH21a, YS23, YFK20, YFK23, YAA25, ZHYL21, ZDWR22, ZL22a, ZF22, ZW20b]. reaction-diffusion [BHJ22, BMK21, CL24b, DPP21, GCMS21, KC22b, LPK22, MFZX23, PK25, SIAA25, WLW^{+}22, XHY21]. reaction-subdiffusion [KR24, MGD23a]. reaction-type [GDŁW23]. reactions [AGA^{+}21]. reactive [XJY^{+}20]. reactor [BAM^{+}21, MWY21]. real [BEKV21, GI20, KPS23a, WSGT21]. real-time [KPS23a]. reality [BBD^{+}24]. receptor [WKGH21]. reciprocity [ZYC^{+}24]. recirculation [HJ23, HJ25]. reconstruct [BBS22, NNT22]. reconstructing [Li25]. Reconstruction [CCGR24, PWMC23, BS21, Bou24a, CFG^{+}25, CNT23, HTBMC20, HLW21, JP25, KLK^{+}22, LCN20, Li21, MWYY22, MRN^{+}20, MZZ^{+}25, XDL^{+}25, YH21a, dFLdSLM22]. reconstructions [DFK22]. Recovering [Hoa21, BZE24, DLC22]. Recovery [SW20a, Ben25, DP20, HLW23]. Recovery-based [SW20a]. rectangular [Fur25, JKS25, KMZ22, KP21, KJN21, KGJN24, KC23a, KKJ^{+}22, MGD23b, YZ21]. rectification [CLZ^{+}24]. Recurrent [BF21]. Recursive [VL21, BN20]. red [SPC^{+}23]. redefined [MH25]. Reduced [AADMA23, BRÁ^{+}20a, BV21, KUK20, SM22a, AS21, ABR20, BGLZ24, BSRVC20, CS20a, CFG^{+}23, ČPPŠ20, GHS23, HSR25, HBC^{+}21, KBR20, KAKK23, LWSZ21, LLZL24, LS20, LML23, LYH24, LWFS24, LN22, MJVH24, NSMR25, PNT^{+}23, PTNR24, SDM21, SSK^{+}23, SR21, SR24, SC21b, UHP23, WF25, WARX23, Yu23, ZAPK23, KNBR22]. reduced-dimension [LLZL24]. reduced-order [AS21, CS20a, LWSZ21, LS20, LML23, LYH24, LN22, NSMR25, SR21, SR24, SC21b, WF25]. Reduction [PWC^{+}23, ABC^{+}22, BMS22, CSB^{+}21, CHJR21, CEB24, GHSR23, GHK25, JZ24b, KC23b, LNNP24, RTA20, YPZ^{+}21, ZXGF22]. reductions [KG20a]. reference [FSS^{+}20, Hak20]. refilling [WWP^{+}25]. refinement [BF21, EAM20, NC20, VJH23, YWL21]. refinements [SGR^{+}23]. Reflecting [DFS21]. reformulation [BHJ21]. refraction [FWBY25]. Regime [YY20b, BDJT21, WZC^{+}20a, WZC^{+}20b]. Regime-switching [YY20b]. region [GP22c, ZSY22]. region-based [GP22c]. regions [TGARGT^{+}23, ZCZ^{+}21, ZSW23]. registration [BCK24, HJZ23]. regression [CFG^{+}23, PN24]. Regular [HYL^{+}25, AJLM24, FAM25, YL25]. regularised [AJK23]. Regularity [AZ24, BN20, MMAK20, MMAK21, CY23, CWW22, CL24b, CP21, MKM24a]. Regularization [GLS25, TTKV20, UR24, BCM22, GPA21, GAM22, GO23, LLSY24, PMLC20, SN20a, Shi24]. Regularized [CSS20, CLLL20, ELS24, HMF24, RZŠ21, WLC24, YLFZ25]. regularizers [LS22b]. Regularizing [WQ24]. regulatory [ZW20b]. reheating [YLQ22]. reinforced [AA21, HTPV24, HTPV25, MMP21]. reinitialization [KCW22, NDC22]. reinsurance [XXY20]. Reissner [DD22, Vis24, YZZ20]. Relation [HB24, CLZ^{+}24, Sto20, ZDW23]. relations [Bar21, JCP^{+}25]. Relationship [DWS20, YCZ^{+}21]. relativistic [BKB22]. relaxation [ADL^{+}21, BM23, CMY23, JM20b, LLX23, LLX24, SM23, WZCS20, WCY^{+}24, XF23, XL25a, YWZ^{+}23, ZJ25, ZWCS20]. relaxed [CL20, DFGS24, DCCM23, FCF22, HYZZ24]. release [AEF20, BFRA21, FdOS20, FGG^{+}20]. Reliability [TJYL22, NNH^{+}25]. reliable [BHK^{+}24, BER21]. ReLU [HLX22, LCC22, LC22, OS24]. remeshing [GHRD20]. removal [BH22, GSS24, LLN24a, LHC23, LH23a, RNN22, SDG20, Shi24, SF25]. removing [LS22b]. reopening [HQCW22, LHQW25]. reparameterization [WM23]. representation [BBD^{+}24, CDM24, RL24, ZZ20a]. representations [GK20, JZPG21]. reproducing [APK^{+}25, Lin24, LN22]. repulsion [GGRBRG20b, GGRBRG20a, WaZWZ23]. require [Jog25]. RERK [YLFZ25]. RERK-FEM [YLFZ25]. reservoir [RSB^{+}21, WYS^{+}20, YY20b]. reservoirs [CCGRL21]. Residual [LMMMP20, MCV24, OSO^{+}23, Šeb24, TBO20, BBK24, BMvdZ21, CBI^{+}23, GCC20, HSM^{+}23, HRS20, Jha21, ŁMMMP21, RNN22, SW21, WFX24, YLSY24, BMP^{+}24]. Residual-based [MCV24, Šeb24, TBO20]. resistance [PWMC23]. resistivity [KLRT21]. ResNet [CYY^{+}21]. resolution [MZ23b, TT25, WZZ20]. resolved [Koc22, RLF21, TWH^{+}22]. resolvent [Pon20]. resonance [EGF^{+}23]. resonances [FM25, XLS24]. resonator [SJM24]. resource [YTY24]. resources [YYY^{+}20]. Response [HJP23, AKA^{+}21, Alf24, CCGF25, LL22c, LVDDNT25, PGR23]. responses [EJWD23, SP20]. resting [AA21, HTPV24, YLFL20]. restoration [BKM23, DH24, LCN20, MJRM20, NES23, PMLC20, ZZ23b]. restrained [LLZ^{+}25]. restricted [YAA25]. result [LZ20e]. resulting [AA25b, OAD24]. results [AHOB24, BMZ20, PKP21, SV21, WL22]. retarded [PS21]. retention [MAT^{+}21]. Retraction [JL20, JFS20]. reversible [Zha22d]. reversible-irreversible [Zha22d]. Review [Bri24, KJ22]. Revised [ZLJ^{+}22]. Revision [Mit22]. revisited [DZ20]. Revisiting [DE21, GDŁW23, Kra20]. Reweighted [JZ24a]. Reynolds [AL22, ABMS20, CO20, HBC^{+}21, JYHS23, KMS20, NSMR25, WDGS23]. rezoning [LZG^{+}23]. RH [HCHM25]. RH-piecewise [HCHM25]. Rham [DD22]. rheological [AGA^{+}21]. Riccati [ATBT20, YTT^{+}22]. Richards [BHZ20, BDL20, BDP23, RFG^{+}21, SMS^{+}23b, TBS25]. Richardson [VJ22]. Richtmyer [RCSM23]. Riemann [sCpLYhG23, HB24, ZHZ^{+}22]. Riesz [AM23, Der23, HQL23, JHG^{+}21, SHM21, SQ23, Wan25, YL25, YFK20, ZZL22, ZZFL21]. right [Kra20, LZL^{+}25]. right-hand [Kra20]. rigid [ZWW^{+}23]. rigid-body [ZBCS24]. rigid-flexible [ZWW^{+}23]. rings [CMS23]. Ripa [RWZG24]. risk [ADG23, ASSV20, LLZ20]. Ritz [LXP^{+}24, LC22, LYXC21]. Ritz-spectral [LXP^{+}24]. river [YTH^{+}21, YT22]. RKDG [ZQZ20]. RKPM [APK^{+}25]. RLW [SJ21]. RN [WL20]. road [ZCZ^{+}21]. Robin [BM24a, CSS23, FLS23, GV21, Gut20, Lee24, MM21]. Robin-type [CSS23, FLS23, Lee24]. robotic [ZWW^{+}23]. Robust [AZHY22, BNV22, BHPS24, BER21, DT24, HKM21, LLSY25, LZ25b, UHL^{+}24, WWL22, ZW20b, BCCR20, FH21, GCL25, JRZ21, KPO^{+}21, KJ21, LM24, LSF25, MFZX23, NK25, QD20, RHA^{+}24, SG23, SR23, SD21, SJ24, SBNR23, Tan24, WS21a, WWS23, YHZ22, ZYTX22, ZL23b, ZPC22, ZXQQ22, ZWW^{+}23, ZQS24, dlRGR20]. Robustness [HNLPP20]. rocks [BMP21, KLRT21]. rods [GM21b]. Rogue [WTQ^{+}20, JXK20, WTSZ20]. role [Mas23]. roll [UHP23]. rolling [ZDT^{+}22]. ROM [JMN22]. root [BBT25]. Rosenau [AJK23, SJ21]. Rosenau-regularised [AJK23]. Rosenau-RLW [SJ21]. Rot [WWZX25]. Rot-div [WWZX25]. Rotated [XW24b, CYC^{+}23]. rotating [ANH^{+}21, AIG^{+}21, Bar22, FGK20, MLHS23]. rotation [CWZ21, YHZ22]. rotation-free [CWZ21]. rotational [GJS20, LPK22, LYAZ20, ZSF23]. rotationally [BKW22]. Rothe [hWCsLjH21]. rotor [HJP23, OKSS21]. rotor-blade [HJP23]. Rough [ZCL20, EFF21, He23a, SLM23]. roughness [JWW^{+}21b]. rows [MAT^{+}20]. rp} [LX24b]. rp}-finite [LX24b]. RPIM [XMQ^{+}21, ZDLC22]. Rule [ES20]. Rules [LP20, CCGV22, FJZC20, HSR^{+}23, Leg21, LQ24, Tau22, WP25, WFMD20]. run [YT22]. Runge [ADB^{+}22, BZ23, CZLQ22, FX22, KTY23, MLL25, SQ23, WWZQ25]. runner [BNQ21].

S [DK23, HMX21, JCP^{+}25, ZSPW20]. S-DDM [ZSPW20]. S-FEM [HMX21]. S-FGM [JCP^{+}25]. Sabin [BEIL21]. saddle [BHJ21, CSS20, CG22, DWAM25, GLL20a, LZW21a, LZW21b, MS20a, SM22b, SS22, SH22, XW24a, XL20, ZYW20]. saddle-point [CG22, XW24a]. Saffman [CBK24, XZ25]. Saint [XDL^{+}25]. Saint-Venant [XDL^{+}25]. samples [LWG^{+}25a]. sampling [AHP20, BKK^{+}23, GHL25, WFX24]. sandwich [AA21, NTP25, SP20]. saturable [LHN24]. saturated [LR22, MGBN22]. saturated-cells [MGBN22]. SAV [HC22a, LSGN23, SWWZ23, WYT23, ZO21]. SBC [KJ24]. SBDF [MMGS25]. SBFEM [ZLJ^{+}22]. Scalable [BHPD23, SYZ23, GdFFF25, KCD^{+}23]. scalar [BDGS25, DT23, HYZZ24, Kri20, KOMK20]. scale [AMC23, BMP21, BMVA23, BKT^{+}23, Cao21, CLW^{+}24, CHL21, DFGS24, DSL24, DWY^{+}25, ERHL22, LQ20, LGQZ21, LXJ^{+}21, LFGZ23, LQW22, NTP25, PWC^{+}23, PH25, PCLZ25, SGF^{+}21, WFX24, XSL24b, YSGD20, YLQ22, YWLF21]. scale-coupled [BKT^{+}23]. scale/color/outline [LP22b]. scaled [FM25, LZJC21, LLZ^{+}24, LJY^{+}25]. scales [SN20b, TH21]. scaling [JYHS23]. scalings [SSW20]. scanner [GBK23]. scanning [FKR21]. scans [YKNS23]. scarce [UHP23]. scattering [AHP20, AGPP22, AAA24, ASZ24, BHDV22, DT24, Fur25, GMS23, GHL25, HLW21, KP22d, LFGZ23, LCE23, LLWX25, MK20, Par22c, RL24, SWC24, SBS24, TPA24, XLYY25, ZWZD24]. scheme [AS20, ASGGRG23, ANZZ23, AAL24, AEZ24, AZY22, ABL24, AMM23, AG25, AN21, BF25, BDB21, BB23a, BZY20, BDGS25, CWX25, CD20, CL22a, Che24, CYQW21, CXX^{+}25, CL21, ČPPŠ20, CRPdT20, CQS25, CTZ22, DMX24, DWS20, DEW25, DR24, DR25b, EBDY22, FZDS22, GLF25, GdFFF25, GV21, GKP25, GGRBRG20b, HMFM25, HUT25, HJM23b, HLH24, HC22a, He23a, He23b, HDWL24, HX22, HYH23, HLJ25, HGW21, HZA22, HJXW23, HQL23, IQS20, JMS^{+}24, JWW21a, JWL20, JWZ22, JZHZ22, JWR^{+}23, JZ24b, KN24, KHH25, KWLW24, KSS^{+}20, KN25, KC23b, LHN24, LSV24, LL25a, LWX21, LZ22, LL22a, LXY23, LZF23, LpWW^{+}25, LZ20d, LYZ23, LZS23, LL24b, LWFS24, LTYZ25, Loh21, LD23, LAB23, MrZQ24, MZGM25, MG25, MRPK21, MLS22, MKP23, MZZ^{+}25, NY22, NFS24, NFFR^{+}24, OCCM21, PHL21, PLZZ25, PL25, PCLZ25, Pło22, PT24, QZH23, QQX21, QP24, RHA^{+}24, Rav20]. scheme [RZ21, RTA20, RWZG24, RC25, RR22, SG23, SK23, SHM21, SCD25, SPC^{+}23, SCCS21, ySRC22, SHD23, SQ24, SKG24, SXKL24, ST22, SQLZ21, TCA22a, TA22, TL22, TWH^{+}22, TRL23, TGARGT^{+}23, TBS25, Ull20, WZZ20, WC21, WZ21, WLJ21, WZY22, WL23, WZTZ23, WR23, WF24, WDYZ24, WSCP24, WZ24, WYW^{+}24, WWP^{+}25, WWYL25, WXH^{+}25, WMMH25, WLW^{+}25, WWZQ25, WF25, WCZD20, WPLW22, XYT^{+}21, XDL^{+}25, XP22, XZSY23, XSL24b, YQZT21, YL22, Yan25b, YKK25, YLZ^{+}24, YHF24, YYSL20, ZWMF24, ZZW20, ZDLC22, ZY23, ZS23a, ZAH23, ZDW23, ZSK23, ZHZ23, ZX20, ZSPW20, ZZB20, ZXQQ22, ZSY22, ZHPW23, ZQS24, ZFQ24]. schemes [AKO21, ADB^{+}22, AJ23, Alt23, Bel24, BD20, BKB22, BCT21, BHL24, BC20a, Bri24, BAB^{+}20, BWMW23, CT22, CMPR23, CLL21, CLLG23, CRM22, CM23, CMY23, CZ23, CGH24, CV20, CMM23, DLW24, DCCM23, DP20, DAS20, DLT20, FHM21, FHZ22, FX22, Gal21, GLLZ23, GCJ23, GGRBRG20a, GGT24, GZ23, GFXZ25, HYJ20, Jha21, JHG^{+}21, JW20, JYHS23, Kar24, KTY23, LPK22, LCY23, LMBB24, LYAZ20, LAX22, LY23, LF24b, LJ24, LLW23a, LMZ23, LRZ23, LHTZ25, Lob21, MMC21, MZCK25, NSRI24, PR20, PS25b, QH22, QD23, RDP20, SQB22, TZ25, Tan25, VL21, WH21, WGeL24, WLC24, Wan25, WYLL25, WL25, WZCS20, WLQ22, XTZ^{+}21, XL25b, YY25, Yua23, ZLPR20, ZO21, ZTW22, ZWT23, Zha22d, ZLZ25b, ZW20a, ZS24, ZZFL21]. Scholes [KMP20, KN25, RYAM21, Ull20]. Schottky [AHZ^{+}25]. Schrödinger [AMV25, AJ23, BV23, CG21, CCF^{+}21, CC20, CLH22, CLY25a, CYQW21, DR25b, FX22, HGW21, HJXW23, HWY25, JZ20, JJZ21, KWLW24, LHN24, LZ20b, LZ20c, LRZ23, MA22, SZ22a, SZ22b, TCW23, WHT22, WL22, WZY23, WSCP24, WL25, WZ25, WLH24, XF23, YWL23, Yan25b, YWZ21, YLFZ25, YHF24, ZSK23, ZW25, ZLH22].

Schulz [LLG25]. Schur [Bot20, NSRI24]. Schur-based [NSRI24]. Schwarz [CM20a, ACPT25, BTCV20, CPS21, CCWC23, GJM22, JRZ21, MGD23b, MR22, Par22a, SR24, dlRRG21, dlRRG^{+}24]. Science [JKL20]. sciences [CSB^{+}21, SA24]. Scientific [GZG20, JMR24]. Scott [AJLM24, CLLG23, GW25, HMFM25, LFY^{+}20, WWG^{+}25]. SDEs [WTSZ24, YY25]. SDFEM [BRRT21]. sea [BS24]. sea-ice [BS24]. seal [SMR^{+}24a]. seal-type [SMR^{+}24a]. Second [BWMW23, DMX24, MKP23, YS23, ZSF23, Ade20, AZY22, AMM23, BHS22, BB23a, BM21, CCWS20, CMY23, CZ23, CQG^{+}24, CYM25, DG22a, DE21, DW20, DAS20, DSF24, GP23, GDŁW23, GQZ23, HJM23b, JKS25, JHG^{+}21, JW20, KWLW24, LG23, LSH21, LXRQ22, LX22, LWFS24, LTYZ25, Liu25a, MW20, MA22, Oru22, PHL21, QH22, QQX21, QFL23, QWC^{+}23, Rav20, STM21, ySRC22, SHD23, SJZ24, SY25, Tan25, TA22, TBS25, WW20b, WH21, WZY22, WZ22, WLL23, WYW^{+}24, WYLL25, WWZQ25, XHL24, YM23, ZO21, ZLJ^{+}22, Zha23, ZL25, ZZ20b, ZZZD24, ZZFL21]. second-grade [SJZ24]. Second-order [BWMW23, YS23, ZSF23, AZY22, CCWS20, CMY23, CQG^{+}24, CYM25, DG22a, DE21, DW20, GDŁW23, HJM23b, JHG^{+}21, LSH21, LWFS24, LTYZ25, Liu25a, MW20, MA22, PHL21, QH22, QQX21, QFL23, QWC^{+}23, Rav20, ySRC22, SHD23, SY25, Tan25, TA22, TBS25, WW20b, WH21, WZY22, WZ22, WLL23, WYW^{+}24, WYLL25, WWZQ25, XHL24, YM23, ZO21, Zha23, ZZ20b, ZZFL21]. section [LZL^{+}25]. sectorial [CWZ21]. sectors [AZ24]. securities [KN25]. sediment [Nde25, RWaZ23]. sedimentation [SLX^{+}23]. seeding [ZDW^{+}20]. seepage [JZ25]. Segel [GSRG21, HH21, HH22, LZ20e]. segment [AEF23]. segmentation [CCH24, GP22c, KJ22, LQZ23, YGZ^{+}22, YKNS23]. segregated [CFP22, HQLJ20, NSMR25]. segregation [XGL21]. Seidel [LQ20]. seismic [AMM23, PLVL21]. seizure [MRP20]. selected [KW20, ZSH^{+}21]. selection [CDPL24, Fur25, SHQL20]. selective [NP24]. self [BCCR20, HJTW25, Mit22, Wat22, WWCJ23]. self-consistent [HJTW25, Mit22, WWCJ23]. selfadjoint [BZY20]. SEM [YQZL22]. Semi [CHZ23, HC23, PCLZ25, SDJ^{+}22, SKY25, WHH22, ZWOM21, AG24, Alt23, ÁCS22, BLX^{+}24, CCG^{+}23, CP21, HZJ^{+}22, JZHZ22, LQYJ23, LZS23, LHTZ25, LQF25, PK25, QS20, SC21a, SQC23, SD25a, SL21, TBS25, UHLZ24, WR23, XJ21, ZHYL21, ZC22, ZH23b, ZJ25, ZL25]. semi-adaptive [LZS23]. Semi-analytic [ZWOM21, SQC23, SD25a]. Semi-analytical [SDJ^{+}22, SKY25, AG24, SC21a]. semi-augmented [ÁCS22]. semi-crystalline [LQYJ23]. semi-decoupled [WR23]. Semi-implicit [PCLZ25, WHH22, CCG^{+}23, JZHZ22, LHTZ25, QS20, TBS25, XJ21]. semi-infinite [BLX^{+}24]. semi-Lagrangian [LQF25]. semi-linear [Alt23, CP21, HZJ^{+}22, PK25, SL21, ZHYL21, ZC22, ZH23b, ZJ25, ZL25]. Semi-monolithic [HC23]. semi-smooth [ZC22]. semi-spectral [UHLZ24]. semiclassical [WHT22]. semiconductor [LX24a, LY24, LCH23, PP24, SJ23]. semilinear [BBI^{+}22, CL24b, DP25, HLPW22, KSJ23, LWX20, LZY24b, LSZ24a, Pło22, PT24, SMR24b, Tor20, ZFZ20]. Semilocal [YSB^{+}24]. sensitivity [AKK^{+}25, AS23, CWLQ22, CL22b, RG21]. separated [LCC^{+}24, YLY20]. separations [SA22]. September [Ano20-67, Ano20-52, Ano21-66, Ano21-62, Ano22-53, Ano22-50, Ano23-68, Ano23-69, Ano24-64, Ano24-65, Ano25-64, Ano25-66]. sequence [LL22c, WWZ^{+}24]. sequences [DDG23]. series [AH25]. set [ABC^{+}21, FMP^{+}24, GO23, HLW21, KCW22, KP20, LQZ25, LTYZ25, MKAM20, NC20, NDC22, NIT^{+}20, NNNT22, UuIUK21]. set-fitted [FMP^{+}24]. set-valued [GO23, KP20]. sets [NF25]. setting [BBG23, RHA^{+}24]. settling [RLF21]. Seven [ZZ20b]. Seven-velocity [ZZ20b]. several [BBS22]. severely [ZCZ^{+}21]. sewage [NM23b]. SFB [CLY25a]. shadow [BH22, LLN24a]. Shah [SHQL20]. shale [Ben25, LZL24]. shallow [ABC^{+}22, Bou24b, BPR22, CBI^{+}23, EJWD23, LMD22, WZZ20, WZTZ23, WKC21]. Shape [FYM23, Par22c, ZGG^{+}25, AAA25, AJLM24, BLM^{+}24, CK23, HKH^{+}24, HXW23, HLW21, JWZX25, KLK^{+}22, LCY23, MVHÖ23, RHA^{+}24, SHQL20, WXF22, ZKTK21]. shape-regular [AJLM24]. shaped [DWS20, LZS25, MNB23, SPC^{+}23]. shapes [BKW22, KRESA21]. shared [BGH20, Fia21]. shared-memory [BGH20]. sharp [SD21, WZT23]. Shaw [Bar22]. shear [BNQ21, HYL^{+}25, HTPV25, MHHR20, PWWJ23, PWJ^{+}24, SGJJ20, SGAC^{+}20, Vah22, YLY20, YZW^{+}22]. shear-deformable [MHHR20]. sheath [WXH^{+}25]. sheet [OAA^{+}22, OSO^{+}23, YHL25]. sheets [LZA23]. shell [CFG^{+}25, HWZQ23, HYL^{+}25, KHD24, LXP^{+}24, LL22c, MJWD20, MKM^{+}24b, MZ23b, PND^{+}24, SP20, ZD21]. shells [AKA^{+}21, EJWD23, JZ24b, LJY^{+}25, LVDDNT25, SRA20, YLZL20, ZD21, ZLF^{+}22]. Shepard [DDNS21]. shift [GLL20a, SH22]. shift-splitting [GLL20a, SH22]. shifted [DYC^{+}22]. Shishkin [MZ23a, SZ24a, WML21]. Shock [NR22, BTYM^{+}21, CYC^{+}23, HCHM25, MKS^{+}21, QhCZ^{+}20, SB24b]. shocks [SA22, SMR24b, XTZ^{+}21]. shooting [NE22]. shooting/collocation [NE22]. short [FP20]. shrinkage [NKPA^{+}25]. shrinking [OAA^{+}22]. sided [JSZ24, LCW22, LDH24, SQL21, ZFL22, ZHPW23]. sides [Kra20]. sideways [TVKN21]. Signorini [LL20b]. silico [SAVHB22]. silicon [ZYM24]. silver [ZYM24]. similarity [KG20a]. SIMP [NNTCB25]. SIMP-based [NNTCB25]. simple [ASH^{+}21, Cos23, EGP20, HXW23, KN24, NK25, SGAC^{+}20, Vah22, WWQ20, YWT22, Zha23, ZZ24c, ZQZ20]. simplexes [CCGV22]. simplices [KV20]. simplicial [AJLM24, HLW24, SNHZ25]. simplified [ABS23, KP22a, SSK^{+}23, TMML23, WZC^{+}20a, WZC^{+}20b, WF25]. simply [KPJO22]. simulate [KP22a, NC20, Ull20]. Simulating [SOPH25, WCZD20, HC23, HLY22, LZA23, LX24a, LCT21, LYL^{+}22, LSY^{+}24, SPC^{+}23, ZCY22, ZXZF20]. Simulation [ADL^{+}21, BM23, DAAEA25, FGK20, Gie23, HLCW24a, SGJJ20, SCD23, SHXZ24, XLLR25, AADMA23, AHZ^{+}25, AMPB^{+}22, ABMS20, AJK23, AEF20, AEF23, BTYM^{+}21, BNQ21, BAM^{+}21, BS24, BPID20, CCG^{+}23, CRDS20, CML^{+}25, Cha24, CWC22, CYC^{+}23, CCGRL21, Dai22, DMG23, FAM25, FKR21, FJP22, GBK23, GHLS25, GF21, GML^{+}20, HLH24, HH22, HQCW22, HWY23, HWY25, KRESA21, KJ20, KMS20, LQYJ23, zLqLZ^{+}24, LQW22, LHQW25, MCZL20, MWY21, MJR20, MB21, MRP20, MWRA22, MKS^{+}21, MBSEA24, PGK20, PSDP22, PWWJ23, PZZ23, QhCZ^{+}20, RL24, RNC25, SB24b, SMR24b, SQB22, TK20, TWH^{+}22, TSARR21, WYS^{+}20, WZC^{+}20a, WZC^{+}20b, WDGS23, WXCL25, WKGH21, WS21b, XL22, XFS23, YMTP25, YL20b, ZDW^{+}20, cZsCyC23, ZJCG21, ZYS22, ZWX24, dSG20, DKH22]. Simulations [HKJ23, SGF^{+}21, APGK24, ASH^{+}21, AFBM25, BRR25, BEG^{+}21, BFI^{+}23, BV21, BT22, Cao21, CPBF21, DTV20, EFF21, GDA^{+}22, GI20, HD21, HEDT20, JMCC24, JYHS23, KK23b, KWK^{+}20, LHH22, LAB23, MIOK21, NDC22, PVC^{+}25, PWJ^{+}24, RLF21, SPZL21, TGL21, VJ24, WZC^{+}20a, WZC^{+}20b, XTZZ24, XRS25, YCCW24, ZvNH25, ZCYL24, ZY21b]. simulator [KTFI23, KGW^{+}21, RSB^{+}21]. Simultaneous [QX23, BML22, DH24]. simultaneously [LDT24]. sine [JS25, JCJ24, WH21]. sine-Gordon [JS25, JCJ24]. Single [KP23, CLLL20, JWW^{+}21b, Li25, LDH24, SK23, Vah22, WCW20b, WZXN21, YOF20]. single-phase [CLLL20, SK23]. single-sided [LDH24]. single-square-grid [YOF20]. Single-term [KP23]. single-walled [WCW20b]. Singular [BNB20, CGHJ24, CWLQ22, BZ21, DST25, FHM21, GP22a, GENRB21, HXZ^{+}22, LQW22, LDH24, LMLZ23, LD25, MKM24a, Men22, NR25a, SZCW24, YZT21, YEO21, YM21, YWLF21, ZJZ24, ZGL23, ZX20]. singularities [KPO20, KPJO22, SB25, SGR^{+}23]. singularity [AZHY22, CL22a, CKBK24, DR25b, FC24, GNB24, PK25, RR23]. Singularly [GO25, BHJ22, CGGY23, CMR22, CSS23, CS25, DN22, HM24, KSS22, LX24b, LZ23, MZ23a, SG23, SZ24a, SSD23, SN20b, WML21, WS25, Xen20, ZZ23a, ZZ24a, ZZ25b]. singularly-perturbed [SZ24a]. sinks [DBH^{+}21]. sintering [CXX^{+}25]. SIR [XSL24b]. SIRS [GV20]. SIS [LYZ23]. site [CCGR24]. situ} [AM20]. Sivashinsky [DS25]. Sixth [FHM21, ANZZ23, MN23b, WGeL24]. sixth-order [ANZZ23, MN23b, WGeL24]. size [DE22, GHLS25, KAKK23, NM23b, PWJ^{+}24, SJM24, YD21]. size-dependent [KAKK23]. size-structured [NM23b]. sized [CTZ22, LZ22, WZTZ23]. skew [LLZ^{+}25, LZFZ22, MS20a]. skew-circulant [LZFZ22]. skew-Hermitian [MS20a]. skid [YLQ22]. skin [LYH24]. slice [KLK^{+}22]. sliding [FGK20, SSZ20]. slip [AAK24, DKMS22, FYM23, OAA^{+}22]. SLM [Cao21]. sloshing [LCT21]. slowly [LZ20c]. sludge [BAM^{+}21]. slug [BPID20]. Smagorinsky [BRÁ^{+}20a, ZS23b]. small [CKFR24, EBT24, HHW21, KP22d, KTY24, LF24a, LWG^{+}25a, MC25, Par22b, TKK22, VJ24]. small-strain [MC25]. smart [EJWD23, MMP21]. smeared [LML23]. smooth [CCH24, FFC^{+}25, GENRB21, GO25, GKK^{+}20, LSW22, LLZ21, MB24, SSD23, UHLZ24, UHL^{+}24, ZJJ20, ZC22, ZF22]. Smoothed [SHT^{+}24, ASH^{+}21, ACPT25, LP21, LAB23, PND^{+}24, SGJJ20, VJ22, YCL20, ZLZ^{+}25a, dSG20, HMX21]. smoother [Gun22, LZ25c, XLS22]. smoothers [dlRRG^{+}24]. smoothing [AADMA23, BFVY23, He23b, HCY^{+}24, LP23, SHXZ24, XJ21]. smoothing-PINN [HCY^{+}24]. smoothly [AGPP22]. smoothness [BD20]. snap [LL22c]. snap-through [LL22c]. snapping [SMS23a]. Sobolev [AAD24, BDMX21, HHG20, KJ20, LLW20, LS25a, MKP23, Ngo25, PLZZ25, XFLZ22]. soft [CCD25, LDW^{+}24]. SoftFEM [DE21]. Software [XXAD21, ZK21]. soil [BBT25, KTFI23, XLZ^{+}21]. soils [BDL20]. solar [ZYM24]. solid [AAC23, APS25, CFG^{+}23, HHR^{+}20, HHY^{+}22, KHD24, LpWW^{+}25, MKM^{+}24b, SRA20, SWC24, TWH^{+}22, WZS^{+}24b]. solid-liquid [TWH^{+}22]. solid-shell [KHD24, MKM^{+}24b]. solidification [DMŠ22, RST^{+}22, WSO^{+}23]. solids [Alf24, OCRN22]. soliton [FWLD22, JMS^{+}24, Waz20]. solitons [WZY23, ZWY23]. soluble [BLLY24]. solute [CS23b, HKJ23]. solute-solvent [CS23b]. Solution [SD25b, AAU25, AA20, AMV24, ASSV20, ANN25, AH25, BM20, BGH20, BGK23, BDB21, BDL20, BLRRS20, BAB^{+}20, CSS23, DHW20, DT24, EMSA25, FMS22, GGH25, GHK25, Hak24b, HLMM20, HZR23, Hop23, HHR24, KS21, KC22a, KSB25, KMN25, LZF23, LLZL24, LGL24, LZY24b, Loh21, NE22, PKPD^{+}24, PS25a, SLD25, SM22a, SP20, SN20b, SZ21, SMS^{+}23b, SBNR23, TO20, TGARGT^{+}23, TPOT20, TBO20, Vab22, XHM20, ZNC24, dlRGR20]. Solutions [CKFR24, WL20, ACXI24, AMV21, AZHY22, ABFPA24, BIMS24, BGK23, CLZ^{+}24, CLXZ25, CLY25a, CL22b, FWLD22, FdOS20, FRW22, Gra21, HLL24, HHL22, HHG20, JXK20, JICRMG^{+}20, KG20a, KP20, LE24, Lee20, LGL20, LL25c, LCT21, LFY^{+}21, LD25, MCZL20, MMN21, MH25, NW24a, NS24, QFL20, RBSC24, Rob25, SAVHB22, TKT^{+}24, UHLZ24, UHL^{+}24, WTSZ20, WX22, WGeL24, Waz20, WS20b, YZT21, YLZ20b, YLFL20, YY20b, ZJJ20, Zha22b, ZW21b, dFLdSLM22, SLUW^{+}21]. solvable [CFG^{+}25]. solvation [CS23b, SHWC22]. solve [ADKH20, Abd25, BZ20, CKS22, HL22, JRu24, SKG24, TFWX23, WZL25]. solved [MLS22]. solvent [CS23b]. Solver [LMK^{+}21, ŁSSP23, BHPD23, BFE^{+}21, Bot20, BFM^{+}25, BSRVC20, CBŁP23, CCGF25, sCpLYhG23, CYC^{+}23, CL21, ÇM20b, DNP^{+}22, EH23, Fia21, GPL21, GHMD22, GF21, GDA^{+}22, Gun22, HB24, HWL^{+}20, HCY^{+}24, HKR22, HDWL24, LHC^{+}22, LFGZ23, LS25b, LLZ^{+}23, LMST20, ŁWP^{+}23, LZB25, MWY21, MST25, NW24a, NBF^{+}24, PD21, SB24a, SLX^{+}23, SU23, TS24, TPA24, WWC^{+}24, WPLW22, XFX22, YPC^{+}21, YZT21, YPZ^{+}21, Zdu21, ZCY22, ZCYL24, ZHZ^{+}22, ZYS22, ZY21b]. solvers [AFM^{+}24, BHPS24, BBK24, BMST20, CAD24, DS20, EED^{+}25, FCFW21, GHH20, HKS21, LLSY25, LNF24, LBPC20, NSMR25, WWL22, WJC^{+}25, WZXN21]. Solving [BdSC24, BGG^{+}20, CDG^{+}23, DDNS21, HSR^{+}23, HBC^{+}21, JKS25, KFD^{+}20, PWBS25, XLCP20, YY20a, APK^{+}25, ABC^{+}21, AFS24, AXLW24, BV23, BF25, BL20, BB23b, BHC20, CWX25, CD20, CO20, CK23, CZC^{+}25, CKL25, CV20, CWX22, DZ22, DC22, DWAM25, GYL25, GHS23, GAM24, GHL25, HZ21, HJM23b, HZ24a, HYZZ24, JKC24, JL19, JL20, JAHM23, KGW^{+}21, Kra20, KGK25, KR24, LSV24, LL25b, LWL^{+}20, LQ20, LCW22, LY23, LXY23, LDW^{+}24, LWG^{+}25a, LMLZ23, LXZZ24, LZLN25, LHTZ25, LZS25, LCT22, LZ25c, MXLY24, MGD23b, MSBO23, NF25, OP24, Oru20, PXLY23, PP24, QD23, QWLQ25, RHA^{+}24, SD25a, Sch21, SLGF20, SU23, SIAA25, TBS25, WCL20, WZZ20, WXZW21, WFX24, WTSZ24, WCC25, WCCC25, WJ25, WL25, XCFC21, XZ24, XSZ^{+}23, XS25, YDY^{+}23, YLSY24, YLCC21, YM21, ZWMF24, ZCXG20, Zha22a, ZYC^{+}24, ZHH21, ZSW23, ZA23, ZXZ25, ZLW23, ZHPW23]. Some [WL25, Wat22, Gal21, Hof20, LP20, PR20, THF20, WZCS20, ZNMS23]. Soner [HMF24]. Soret [JSZ24]. sound [Par22c, SP20, WYH20]. sound-hard [Par22c]. Source [Fur25, XXAD21, AL23, AMV21, Cha24, DHVT20, GdFFF25, GDŁW23, HANA24, HH21, HH22, HMX21, KSS22, Kam21, KHM25, KGW^{+}21, KKA^{+}21, KSSS23, LZ20e, OAH^{+}24, PS20, PH23, QX23, SSD23, TSARR21, UR24, YL20a, ZLYL20, Zha22e, ZLF^{+}22, dFLdSLM22]. source-term-discontinuous [KSS22]. sources [DLC22, FHM21, Gra21, LGL20, Nag24, RZŠ21]. Space [AMC23, BDV23, FK21, HH21, HRS20, KSD^{+}25, THF20, ZRL^{+}22, AL23, ABV23, BHZ20, BBK21, BMK21, CLLG23, CCWS20, CEB24, CCGR24, DD24, Der23, DS22, DMŠ22, DBB23, EJR24, ELS24, FLMS24, FG25, GvV22, GKK^{+}20, Gun22, HLCW24a, Hop23, HGW21, HQL23, JCJ24, JHG^{+}21, JWR^{+}23, JSZ24, KV22, KR24, LSW22, LCY23, LLW20, LXZ23, LFY^{+}20, LMST20, LP22a, ŁSSP23, LZ25c, MSHB22, NH24, PS21, QX23, RKK25, SOPH25, SC21a, SHM21, SQL21, SQ23, SK21, SW21, SC21b, SXLZ21, SLYL22, SF25, TH23, TGARGT^{+}23, TH21, UR24, WL23, WCWZ23, WWYL25, Wan25, Wie23, XLLA20, XLS22, XF23, XLSF25, XHL21, YY20a, Yan20, YCZ^{+}21, YLL24, YL25, YHL25, YLFL20, YFK23, YAA25, Yua23, ZWOM21, ZLPR20, ZZ21, ZLT^{+}21, ZFL22, ZY23, ZZZ24, ZGL23, ZZB20, ZHPW23, ZZFL21]. space-dependent [AL23, QX23, UR24]. space-fractional [HGW21, HQL23, JWR^{+}23, LZ25c, SQL21, SLYL22, WCWZ23, XLS22, XF23, XLSF25, YL25, ZLPR20, ZZ21, ZHPW23, ZZFL21]. Space-time [AMC23, BDV23, HRS20, KSD^{+}25, ZRL^{+}22, ABV23, BMK21, CCGR24, DS22, DMŠ22, ELS24, FG25, GvV22, Hop23, LXZ23, LP22a, ŁSSP23, MSHB22, PS21, SOPH25, SW21, SXLZ21, Wie23, XHL21, YHL25, Yua23, ZWOM21]. spaces [AMP20, Adj22, AAD24, BDJT21, CCGI23, CRF^{+}20, CG22, CP21, DDG23, FLOQ21, GORBS21, GI21, MXLY24, NW24b, WN24]. spaces-based [CCGI23, GORBS21, GI21]. spacetime [HA23]. Sparse [CHL21, PH25, BN20, ÇM20b, KM22, LH23b, RNC25, SU23, Ull20, WWC^{+}24, Xu24, ZZ20a]. spatial [Bal21, CL21, FZS22, FSH^{+}23, GP22c, GV20, KMN25, LQLF23, LLZL24, MHHR20, PLZZ25, QFL23, SZS20, SW21, WSS22, YMTP25, ZLQ23]. spatially [IJS^{+}24, ZLYL20]. spatio [AMS24]. spatio-temporal [AMS24]. Spatiotemporal [TSW^{+}25, YMTP25]. SPDEs [Kam25]. Special [AGNP24, BKB22, Sch21, ZSH^{+}21]. species [BOSV20]. Specific [WZXN21]. specifying [Ila20]. Spectral [AD21, HJT25, TCA22a, TjW24, AKK^{+}25, ACLP22, AA20, AAM^{+}24, BWZ23, BDB21, CT22, CK23, CL22a, CKBK24, DE21, Der23, GP20, GCMS21, HUT25, HQLJ20, HWZQ23, HH23, JJZ21, KMN25, Lee21, LZ20b, LCW22, LY23, LXP^{+}24, LY25, LYL^{+}22, LGZC23, LYL25, MC24, nMjWlS24, Ma25b, NH24, NHCJ22, NML23, Pel23, QC24, RTA20, SM22a, SWWZ23, TS21, Tür20, UHLZ24, UHL^{+}24, WFH^{+}20, WHT22, jWC23, WCC25, WYLL25, WZA25, YPC^{+}21, YCCW24, YMTP25, YAA25, YM23, YW24, ZM20, ZY21a, ZFL22, ZDWR22, ZLJT20, ZCA23, ZSL22, ZZ24c]. spectral-collocation [LZ20b]. spectral-Galerkin [nMjWlS24, RTA20]. spectral-geometric [HWZQ23]. spectral-like [CT22]. spectral/hp [YPC^{+}21]. spectrum [VLD^{+}24]. speed [SQB22]. SPH [LAB23, Vak25]. sphere [JWW^{+}21b]. spheres [SS21, SS24]. spherical [HWZQ23, JZ24b, KHD24, LYP^{+}22, LXP^{+}24, RLF21, XGL21]. spheroidal [BRR25]. spheroids [SGAC^{+}20]. spherulite [LQYJ23]. spin [GM21a]. Spline [MSHB22, AMP20, BDF24, EFK^{+}23, FKKV24, GKK^{+}20, GK21, HLWL25, HSR^{+}23, KV22, KG20b, Kra23, LLZ^{+}24, MH25, Rou20, SRD25, SHXZ24, TQX23, WYHQ24, WLD22, YLZ^{+}24, ŻW21a]. Spline-based [MSHB22, ŻW21a]. splines [AS21, BEIL21, CG21, CM22, DS25, EJR24, GP22a, GK21, Kap24, MP23, ZCL20]. Split [SRD25, TSS24]. Split-step [SRD25]. Splitting [Alt23, BGK23, NSRI24, BAB^{+}20, CL20, CYKL23, GLL20a, GM21a, He23a, JS25, KS21, KEUY23, KM22, LKC^{+}24, LZFZ22, LL24a, LMMMP20, LP22c, MS20a, NZ24, OEASY24, PL25, SZS20, SH22, SJZ22, TH23, WCW20a, WXF22, WHT22, WL23, XFS23, XWXF23, YL25, ZS25, ZZW20, ZWX24]. splitting-based [ZS25]. splitting-up [He23a]. splittings [IK24, NSMM21]. spraying [SLUW^{+}21]. spread [THF20]. square [ACXI24, AA25b, Axe20, HBC^{+}21, KSD^{+}25, LD23, MC24, Ma25a, PM23, WX22, WCZ24, XYS^{+}24, YY25, YOF20, ZZZC24, ZLQ23, ZHZ23]. Squares [BDG21, UBP^{+}25, AN21, BM24a, BB21, BBG23, BS24, Bri24, CCY24, CM21, CLY25b, FP20, FK21, JBKW23, KC21, Ku24, LZ25a, LZ25b, LCC22, MHKB25, Pac24, QC24, QFL20, SV21, STM21, SW24a, Sto20, WW20b, Zha23, BB23b]. squeeze [WLZZ22]. squeeze-film [WLZZ22]. squeezed [AKK^{+}25]. stabilisation [YOF20]. stabilised [MBG24]. Stability [BZ23, DN22, Der23, NS20, QZH23, SQL21, AD21, AHOB24, BDS24, CS22, CYC^{+}23, CZ23, CLY25b, GV21, KTY23, LGL20, LZ20c, LQCM20, Pac24, RGS24, SGZ^{+}23, STM21, SIAA25, YYZ22, ZL22a, ZY23, ZLQ23, ZHZ23, ZW20b]. Stabilization [BMPP22, CS23a, AN23, BBM25, BLMV24, CRRS21, CEL20, CLL^{+}24, ERG^{+}24, HJM23a, HJ24, KCW22, KSD^{+}25, LXRQ22, Mas23, MN23a, QHHW20, SCD25, WCH23, YEO21, YQZL22]. Stabilization-free [CS23a, BLMV24, CLL^{+}24]. Stabilized [ABR20, BNC^{+}23, AJLM24, ANN25, BLF23, CO20, CXXZ23, CRF^{+}20, DL25, DZ24, FLS23, GCC20, HDM24, HSM^{+}23, HOY24, JWW21a, JZS23, JMN22, KC23a, LZL24, LHF24, LH23b, ŁSSP23, SLT24, SR22, TBO20, WHH22, ZYD24]. stabilizer [GHL24, MGD23a, ZX22]. stabilizer-free [GHL24, MGD23a, ZX22]. Stable [HTBMC20, JYHS23, LPK22, Bel24, BKB22, BFRA21, BC20a, CFG^{+}25, CMPR23, CBI^{+}23, sCpLYhG23, CGH24, CGK24, CSI21, DMX24, FLMS24, GS22, GLF25, GHMD22, GLÖ21, GKP25, GK20, GGRBRG20b, GZ23, HYJ20, JS25, JWL20, JW20, LL25a, LWSZ21, LaZZ24, LJ24, Lob21, OF20, PHL21, PLZZ25, PL25, QH22, QZH23, RDP20, Rav20, RC25, ySRC22, SU23, SXKL24, SY25, SWZ23, Tan25, VR20, VBDR24, WLJ21, WZY22, WaZWZ23, WDYZ24, WHH22, WKC21, XKXL22, XL25b, YCGL25, ZO21, ZD22, ZS23a, ZWT23, ZZW25, Zha22d, ZLZ25b, dSMAdOW22]. stacking [LL22c, WWZ^{+}24]. stage [CD20, DWY^{+}25, GHSR23, WLH24]. Staggered [ZPC20, BHPS24, GWYR22, KJN21, KGJN24, WJ25, ZPC22, ZPK22]. staggered-grid [GWYR22]. stagnation [AGA^{+}21, BWZ22, BWZ23]. stagnation-point [AGA^{+}21, BWZ22]. stainless [Cao21]. staircase [ZKBC21, ZBKC22]. standard [APS25]. standing [DMG23, LZ20c]. star [DGK20, DWS20]. star-shaped [DWS20]. start [CSCY24]. started [WCZD20]. state [Ale23, AGM^{+}22, BGPS20, BGZ22, BJyST24, BCM22, BC20b, CZLQ22, CSCY24, GLS25, GORBS21, HBC^{+}21, LSWY21, MJZ25, SD25b, TS21, VGA^{+}22, WLQ22, XHM20, ZLYL20, ZZL23, ZS23b]. state-based [Ale23]. state-constrained [AGM^{+}22, BGZ22]. states [ABS23, MZWD23, PP24, SSdA20]. Static [LJY^{+}25, PND^{+}24, ZKBC21, BAB^{+}20, EMSA25, Hop23, JSSS24, MC25, TMML23, XZ24, ZR22]. stationary [BDE^{+}24, CGO22, CCGI23, HJP23, LD22, LXRQ22, LMMMP20, ŁMMMP21, MGBN22, MLJ^{+}20, MS22, NSMM21, PTNR24, Qiu21, QQY^{+}23, SL20, SLGF20, ZZ24b, ZZZ25, ZXZF20, ZST22, ZZS24, ZHW20]. statistical [Kap24]. steady [ABR20, ABS23, BGPS20, BCM22, GORBS21, HBC^{+}21, JICRMG^{+}20, KNBR22, LSWY21, LLZ^{+}23, MG25, MJZ25, SSY22, SD25b, WDGC20, WFW^{+}24, XYS^{+}24, YA23, ZLYL20, ZZL23, ZS23b]. steady-state [BGPS20, BCM22, GORBS21, HBC^{+}21, LSWY21, MJZ25, ZLYL20, ZZL23, ZS23b]. steam [HLEF20]. steel [Cao21]. Stefan [CFTC25, DKL24, NRV22a]. Stekloff [BZY20]. Steklov [Ant21, VLD^{+}24, XXLS23, ZPK22]. stem [ZDW^{+}20]. stencils [CT22, LLU24]. stenotic [ISA24]. step [BS23b, Bel24, BCM25, DE22, DCCM23, HUT25, LLLL24, LYL25, MZHJ24, PWMC23, SRD25, WZ21, WLD22, YLCC21]. stepping [BWMW23, CYM25, DP25, DR25b, HMFM25, HJM23b, Hoa22, HLCW24a, KP23, LAX22, PD25, Rav20, RH21, ySRC22, WYW^{+}24, WYLL25, XZ25, YFK20, YFK23, ZvNH25, ZZW20]. steps [CQG^{+}24, CGH24, KMS20, PLZZ25, QWC^{+}23, SJZ22]. stepsize [LQ24, MMGS25]. stepsizes [Tan25]. stiff [DdSP21]. stiffened [LXG^{+}24, LML23]. stiffener [LML23]. Stochastic [YYZ22, ADG23, ASSV20, BX20, BKK^{+}23, CCGF25, CCX21, CS20a, Che24, Dai22, DKL24, DYZ22, FS21a, FSH^{+}23, GAM24, HJP23, HZR23, JZ24a, KHY25, LLW20, LZL24, LY25, LLY25, LZ20d, LQCM20, LD20, LY21, LQ24, MW20, TC25, TZ20, VT21, WS20b, WWH23, YY20a, YHL25, YY25, YTH^{+}21, YTT^{+}22, ZLQ23, ZMZW24]. stocking [TKMK25]. Stokes [GAM22, GGT24, LD25, SXKL24, WDZ21, Agg20, AJLM24, ASB25, ABR20, ACRBT23, BdSC24, BSRVC20, CCGI23, CLR23, CLRT24, CW20, CL20, CHJR21, CJ25, CG25, CTZ22, DKMS22, DZ24, FYM23, FLOQ21, GP25, GLF25, GOV20, GI21, GWN21, GS23, HJM23a, HDM24, HQLJ20, He23b, HKM21, HHW21, HX22, HWC24, JRu24, JWL20, JZS23, Kap24, KNBR22, KGJN24, LS22a, LL21a, LM24, LYAZ20, LD22, LXRQ22, LL23, LHF24, LLY25, LQZ25, LD20, ŁMMMP21, LCT22, MRN^{+}20, MHKB25, MZHJ24, MZZ^{+}25, NK25, Oru22, OSZ22, PS22, PNT^{+}23, Pup25, QSW23, QHHW20, QWC^{+}23, QQY^{+}23, QD20, RZC21, RDP20, SD25b, SCCS21, SLT24, SLGF20, SR22, TS24, Tür20, WZ21, WWL22, WWS23, WaZWZ23, WZS24a, WFW^{+}24, WH20, WHH22, WCH23, WWH23, XZ25, XCLX24, XW24b, XHL24, Yan20, YHZ22, Yem24, YHLA22, YW20, ZQM20, Zha22a, ZDW23, ZF24, ZPC22, Zha24, ZLZ25b]. Stokes [ZS21, ZNMS23, ZFQ24].

Stokes-Korteweg [MRN^{+}20].

Stokes/Darcy [GOV20, HX22, LYAZ20, QHHW20, QWC^{+}23, XHL24]. storage [JFS14, JFS20]. storages [FL22]. strain [BZ21, BM24b, CWTM20, JCP^{+}25, LP21, LP23, MC25, SJM24, TRL23, VJ22]. strain-smoothed [LP21]. strains [PGR23]. Strands [BV23]. strategies [ASF^{+}25, BF21, DTV20, LZG^{+}23, Loh21]. strategy [BGG^{+}23, CL24a, CGGY23, CRM22, DBB23, GHRD20, GP22b, HLN24, JAHM23, MMC21, MVHÖ23, SMS^{+}23b, SBNR23, ZZ21, dR25a]. stream [ZF24]. stream-function [ZF24]. streamflow [YTT^{+}22]. streamfunction [LLX23]. streaming [DMG23]. streamline [LDM25, SSW20]. streamlined [SZ24b]. street [AADMA23, WSGT21]. strength [MMP21]. Stress [DFGS24, BNQ21, BS21, CXXZ23, DGQT20, DFK22, DCSAM25, JSSS24, KPO20, KPJO22, Med23, NNNT22, NTP25, NK25, WQS25, YY23, YLY20, Zha24]. stress-based [NNNT22, NTP25]. stretching [OAA^{+}22, OSO^{+}23]. stretching/shrinking [OAA^{+}22]. stride [YMTS20]. stringer [LML23]. strip [MAT^{+}20]. stripe [WTSZ20]. Strong [AMK25, AD21, CXXZ23, XTZ^{+}21, YWT22, YHL25, ZWT23]. strongly [AKO21, Ade20, FWLD22, KRO24, Med23, PKP21, WH20, Zha24]. strongly-constrained [FWLD22]. Structural [CL22b, BM21, FMP^{+}24, FK20, HLWL25, XFX^{+}24]. Structure [DZZ^{+}24, GGT24, JH24, ASGGRG23, AZY22, CCG^{+}23, CG21, CXX^{+}25, DLW24, DL25, DCSAM25, FZH22, GPL21, GV21, GCSF21, HC23, HGW21, ISA24, JJOG20, KWM22, Kam21, KJ21, KWLW24, LFF^{+}24, MSHB22, MZM21, MKS^{+}21, NSMR25, QS20, RLVR21, SLYL22, SHXZ24, TT25, Vak25, WH21, XDL^{+}25, YTK21, YGZ^{+}22, YHF24, ZDW^{+}20, ZS23a, ZXXX23, ZJZ24, ZW25, ZNH24, dCLW25]. Structure-preserving [JH24, ASGGRG23, CG21, CXX^{+}25, DLW24, DL25, KWLW24, QS20, SLYL22, WH21, XDL^{+}25, YTK21, YHF24, ZS23a, ZXXX23, ZW25, ZNH24]. structured [BEG^{+}21, NM23b, OF20, TZ25]. structures [APGK24, BDS^{+}22, CCGF25, DWY^{+}25, DBB23, EBT24, Gie23, Ha20, HXZ^{+}22, KJL^{+}20, LLMC25, LR22, MJWD20, MKM^{+}24b, NNNT22, NNH^{+}25, NNTCB25, PND^{+}24, PP24, WQS25, XLS24, YA21, ŻW21a, ZLF^{+}22, ZLZ^{+}25a, ZXG^{+}24]. studies [XTZ^{+}21]. Study [GGRBRG20b, GGRBRG20a, RLVR21, SLX^{+}23, Adj22, ACV21, BHPD23, BS23a, CMS23, CS23b, DBH^{+}21, EGF^{+}23, GKHU22, HCHM25, HHR^{+}20, HD21, IZ20, ISA24, KAKK23, LFF^{+}24, LXG^{+}24, Liu25b, MRP20, MZCK25, NM23a, NDB^{+}24, OAA^{+}22, OAD24, PTM^{+}23, RSAG^{+}24, SB21, SAVHB22, SIAA25, TKT^{+}24, WSGT21, XGL21, XLZ^{+}21, YD21, YD23, YY23, YZL20, ZYQ^{+}24, ZJCG21, ZL23c, ZLD^{+}23]. studying [FL22, HGD21, THF20]. style [MB21]. sub [DAS20, JVBM23, MJR20, SM23, WCLQ25, WLD22, YZT21, ZYTX22]. sub-diffusion [DAS20, SM23, WCLQ25, YZT21, ZYTX22]. sub-grid [MJR20]. sub-parametric [JVBM23]. sub-step [WLD22]. subcell [KCW22, RHA21, SK21]. subcell-enriched [RHA21]. subdiffusion [DR24, JL23, KR24, LQY22, MGD23a, Pło22, PT24, Pon20]. subdiffusive [KMP20]. subdomain [CQG^{+}24, QWC^{+}23]. subgrid [Zha21]. subject [ACXI24, ANH^{+}21, LXY23, PWWJ23]. subjected [YLZL20]. sublimation [Cha24]. submerged [Yem24]. suboptimal [AGM^{+}22]. subsampling [LNF24]. subscales [EKB20]. subsonic [SMR^{+}24a, SQLZ21]. Subspace [LLZ21, XS25, AH25, FK20, TS24]. substance [WSO^{+}23]. substitution [SXLZ21, XLLR25, ZRL^{+}22]. substrate [BM24b]. subsurface [LBTVG22, SYZ23]. subwavelength [XLS24]. successive [PXLY23, XL25a]. sufficient [KN20, TZ25]. suitable [FKKV24]. suited [BMVA23]. summation [NERB^{+}22]. summation-by-parts [NERB^{+}22]. sup [AJLM24, CS22, DFHT21, LSZ24b, Sau23]. super [GKK^{+}20]. super-smooth [GKK^{+}20]. Supercloseness [LZ23, MZ23a, ZZ23a, ZX22]. superconductors [WS25]. Superconvergence [BRRT21, CSS25, Li24a, LS23, Liu25a, SJ21, SJ23, SLT24, WLJ21, WL22, WGC23, WSCP24, WS20a, YWL23, YJ24, YLZ20a, ZNMS23, LSW20, LS25a, LLXZ23, PKP21, SL21, SLW22, SZ22b, SZ24b, SQ24, WLW^{+}22, XXLS23, YLW23]. Superconvergent [SL20, WLL23, WS25]. Superconvergnce [PLZZ25]. supercritical [JWW^{+}21b]. superdiffusion [AZHY22]. superhydrophobic [LZL24]. superlinear [SSdA20]. Supermodeling [PSDP22]. superposition [BL20]. supersonic [sCpLYhG23, SA22, WLX20]. SUPG [JMN22, ZL22a]. SUPG-stabilized [JMN22]. supply [BEKV21]. support [JZXR24]. supported [KPJO22, NTP25, WS21b]. supports [CWZ21]. suppressed [HHNV23]. suppression [ZWW^{+}23]. supraconvergent [BFRA21]. Supremum [KPS23b]. Supremum-norm [KPS23b]. surely [YYZ22]. Surface [KJ24, Yan24, ANH^{+}21, AIG^{+}21, AEF20, BHD^{+}21, BHDV22, CLLL20, CLW^{+}24, CFT^{+}23, DNP^{+}22, HKJ23, HLH24, HOY24, JZ25, JWW^{+}21b, JBKW23, KK23a, LYP^{+}22, LL20a, Li21, LLZ21, LZA23, MKAM20, NC20, OXY21, SK23, SBH^{+}21, WWG^{+}25, XDL^{+}25, YSGD20, YJZ20, ZDLL22]. surfaces [CXF24, EFF21, FKKV24, GXFT23, Hak24b, KYY^{+}20, Kra23, LQF25, SR23, YWT22, ZXZF20, aKST22]. surfactant [XWXF23, Yan24]. surfactants [BLLY24]. survey [Mas23]. suspended [AAS^{+}20, HLH^{+}20a]. Svärd [KP22a]. SVD [NCCD24]. sweeping [LQ20]. Swift [QH22, WCWZ23, YTK21]. swimmers [AAWJ20]. swimming [CM24]. switching [KP21, WDGC20, YY20b, ZW20b]. Sylvester [LWL21, LZFZ22]. Symmetric [FJZC20, KC23b, BKW22, BS21, CXXZ23, CCGV22, FC24, HZJ^{+}22, KMZ22, LPK22, LWL21, SN20a, SS22, SCG23, SCdH21, TS24, WS21a, WZY23, WWC^{+}24, WFMD20, Zha24, ZXZ25]. symmetric-triangular [SS22]. Symmetrical [WS20b]. symmetries [LE24]. symmetry [DGQT20, DFGS24, Med23, ZYQ^{+}24]. symplectic [ZSK23]. Symposium [ZSH^{+}21]. synapse [Gie23]. synchronization [MMN24, MRP20]. System [BB23b, AHOB24, AN23, AG25, BML22, BS23b, BW21, BKM23, BGG^{+}20, BM24a, BMS22, CRRS21, CQG^{+}24, CSI21, ÇM20b, DN22, RFG^{+}21, DKV21, DYZ22, DSF24, DH24, FKW22, Fan23, FLOQ21, GSS24, GDA21, GHS23, GGT24, GFXZ25, Gut20, HSR25, HJP23, HTDT23, HHNV23, HH23, IQS20, JZHZ22, JM20b, KG20a, KLY25, LGQZ21, LXJ^{+}21, LL22a, Li24b, LDM25, LDH24, LQCM20, MA22, MZM21, PS22, PTM^{+}23, QQY^{+}23, RC25, SADW21, SR20, SV21, SN20b, SMR24b, TA22, WCW20a, WYP23, WaZWZ23, WH25, WL25, WSS22, Xu24, YGZ^{+}22, YHF24, YCWS20, ZL23a, ZSK23, Zha23, ZCZ^{+}21, ZGL23, ZSY22, ZQ25]. systems [ACDL24, AS23, AH25, BAKR24, BCT21, BHC20, CBŁP23, CSS20, CWW22, CFP22, CG21, CKBTS24, CHJR21, CS25, DLN24, DHW20, DWAM25, GP22a, HYZZ24, JL19, JL20, KM22, KYY^{+}20, Kim21, Koc22, Kra20, LE24, LZW21a, LZW21b, SS22, SK21, SSdA20, Vab25, WLZZ22, XZ24, XSZ^{+}23, YY20b, YAA25, YM21, ZCL20, ZZL22, ZX24, ZNC24, ZWS22, ZMZW24].

T [BRS21]. T-IFISS [BRS21]. tactoids [AAA25]. tailored [EFK^{+}23, XHY21]. tangent [PWBS25]. tank [LCT21]. Targeted [PTM^{+}23]. targets [LFGZ23]. tau [NML23, ZM20]. tax [XXY20]. Taylor [GdFFF25, XL24]. TBC [XLYY25]. TD [LQLF23]. TD-BEM [LQLF23]. TD-FEM [LQLF23]. technique [Agg20, Alf24, BV23, BL20, BGH20, BBK24, DAAEA25, DCSAM25, ERG^{+}24, FC24, GQ23, HLX20, KR24, LCE23, LMLZ23, LP22c, Par22b, RL20, RKK25, SD25a, SWZ23, WZ21, YQZL22, Yu23, YW24, ZBCS24]. Techniques [DO23, BCF^{+}23, KHY25, MHHR20, TRSM23, UR24, VADG22, WCZ24, YL25]. technology [Vab25]. telegraph [MJRM20, YQZT21, ZLL20]. telegraph-diffusion [MJRM20]. telegraphic [MGA22]. temperature [AOL22, BDE^{+}24, BABS20, Cha24, FdOS20, TSARR21, YWL21, ZTL22]. temperature-dependent [Cha24, YWL21]. tempered [DR24, DR25b, JL23, LL24b, QGQ22]. tempered-fractional [DR25b]. templates [SNHZ25]. Temporal [DAS20, HZA22, AMS24, DTW23, HS20, LWX20, Pac24, QFL23, SQ23, Top22, TBS25, YMTP25, ZHZ23]. temporal-spatial [YMTP25]. TENO [MZZ^{+}25]. tensile [YZW^{+}22]. tensile-compressive-shear [YZW^{+}22]. tension [MKAM20, NC20, SRD25]. tensor [CHL21, DS25, Mal22, PH25, YGZ^{+}22]. tensorial [SNHZ25]. Tensorized [ZZL22]. tensors [BHC20]. Tent [HA23, PSSW20]. Tent-pitcher [HA23]. term [GKHU22, GF21, GCC20, GO25, HYW^{+}24, HAC23, KSS22, KSSS23, KP23, LF24a, LYL^{+}22, MLLL24, MN23a, Ngu20, OAH^{+}24, RKK25, SM22a, Top22, WCH20, WXCL25, XFLZ22, YWPS24, ZQM20, ZYTX22, ZX20]. term-by-term [GCC20]. terminal [XXY20]. terms [AMV21, CS25, GDŁW23, HB24, KP24, SSD23, ZW20b]. ternary [HLH^{+}20b, IBY^{+}24, WZY22, ZWMF24]. terrain [TGL21]. terrain-disrupted [TGL21]. tertiary [DP20]. tesseract [NW24b]. test [CM23, GM21b, MN21, SBY^{+}25, WP25]. Tests [Zdu21]. tetrahedral [Gie23, LLXZ23, WN24, YLW23]. tetrahedron [BRR25]. tetrahedron-based [BRR25]. tetrahedrons [IKST21]. textile [FGK^{+}21]. texture [RGL22]. their [BHC20, Din21, GLLZ23, Kra23, LMBB24, MRP20, OAD24, YTA23]. Theoretical [WZ24, ZL22b, Liu25b, XL25b]. Theorist [Sch21]. Theory [PV20, AP25, BDP23, Ha20, HJTW25, HTPV25, LZJC21, MMAK20, MMAK21, SJM24, WQS25, WWCJ23, WCW20b]. therapeutic [ASF^{+}25, GOP25]. therapy [CRL^{+}25]. Thermal [LML23, MMP21, ZXG^{+}24, Alf24, AAG22, AGA^{+}21, AL23, CRL^{+}25, CZC^{+}25, DK24, DWY^{+}25, FL22, GY25, HKS21, HHY^{+}22, HB25, JM23a, JFS14, JFS20, KGB24, MLHS23, QP24, SB21, SD25a, SBH^{+}21, SLUW^{+}21, TWH^{+}22, UuIUK21, WWQ20, WWY25, XFX22, YL20a, YLZ20a]. thermal-diffusive [SB21]. Thermal-mechanical [LML23]. thermal-wave [AL23]. thermally [OCCM21, RC25, TA22]. Thermo [AA21, BAB^{+}20, CLLL25, HEDT20, SDJ^{+}22, YSGD20, YLW^{+}22, ZR22]. thermo-chemical [HEDT20]. thermo-elastic [YLW^{+}22]. Thermo-electro-mechanical [AA21]. thermo-mechanical [YSGD20]. thermo-poroelastic [ZR22]. thermo-poroelasticity [BAB^{+}20, CLLL25]. thermodynamic [AGA^{+}21]. thermodynamically [ZGW23, Zha22d]. Thermoelastic [ASZ24, AAC23, JM23a, NNTCB25, SZCW24]. thermoelasticity [BDS24]. thermoelectrical [LX24a]. Thermogravitational [PM23]. thermomechanical [CWTM20]. thermomechanics [CTGR24]. thermomicropolar [RL23]. thermoviscoelasticity [Alf24]. theta [YY25]. thick [ZDT^{+}22]. thickness [KKJ^{+}22]. thin [BBS21, CWZ21, CRDS20, HXZ^{+}22, LLW20, LZJC21, MNB23, NP24]. thin-walled [NP24]. third [LCY23, LYLH25, OCCM21, SWZ23, ZCYL24]. third-harmonic [LYLH25]. third-order [LCY23, SWZ23, ZCYL24]. thixotropy [Vah22]. Thomas [AAD24, GQ23]. Thompson [AAC23]. thoracic [ISA24]. Three [Abd24, BAM^{+}21, Fan23, KLK^{+}22, SQGZ24, WQS25, WCW20b, ZBKC22, Abd25, BS23b, BZ20, BF25, BDS24, BOSV20, BM23, RFG^{+}21, DWY^{+}25, EH23, FCFW21, HKH^{+}24, He23b, HZ24a, HZ25, HMŁW^{+}24, JWZX25, LGQZ21, LLMC25, MMPR20, QQX21, QSL23, QP24, RZC21, RZSL23, RG21, SD25b, SXLZ21, SWC24, Vis24, WXF22, WYHQ24, WCCC25, WLH24, YSGD20, YQZT21, YWPS24, ZCYL24, ZZL23, ZZ20b, ZSW23, ZHPW23, ZZFL21, aKST22]. Three-component [Fan23, ZBKC22, WLH24]. Three-dimensional [Abd24, KLK^{+}22, SQGZ24, WQS25, WCW20b, Abd25, BZ20, RFG^{+}21, EH23, HKH^{+}24, HZ24a, JWZX25, LGQZ21, QQX21, QSL23, RZSL23, RG21, SD25b, SXLZ21, Vis24, WXF22, WCCC25, YQZT21, YWPS24, ZCYL24, ZZL23, ZZ20b, ZSW23, ZZFL21]. three-field [FCFW21]. Three-phase [BAM^{+}21, BM23, LLMC25, QP24]. three-phase-lag [BDS24]. three-scale [DWY^{+}25, YSGD20]. three-species [BOSV20]. three-step [BS23b]. threshold [LHC23, SB21]. tidal [ZY21b]. tide [KK21]. tIGAr [Kam21]. Tikhonov [DHVT20]. Time [BWR21, FRW22, GM21a, HH22, KG20a, KHM25, Pon20, RH21, SQC23, SF25, AD20, AS21, AD21, ADD22, ADGG25, AMK25, AHOB24, ADL^{+}21, AZY22, ABV23, AMM23, AN23, AMC23, BDV23, BV23, BS23a, BLF23, BHPS24, Bar22, BBK21, BKR24, Ben25, BMK21, BB23a, BCCR20, Bot20, Bra20b, BWMW23, CWX25, CRL^{+}25, CS20b, CCWS20, CLZW20, CL22a, CLYH22, CLH22, CQG^{+}24, CGH24, CLY25a, CYM25, CMR22, CCGR24, CMM23, CG25, DLN24, DP25, DHVT20, DS22, DMŠ22, DAS20, DC22, DR25b, DBB23, ELS24, FKW22, FC24, FSW20, FBK^{+}24, FLMS24, FX22, FK21, FG25, GvV22, GLL^{+}20b, GP22b, GKHU22, GDA21, GMS23, GHSR23, Gra21, GI20, GAM24, GGRBRG20a, GCMS21, GHL25, GCL25, HMFM25, HJM23a, HJM23b, HLWL25, HDG23, HB25, HH23, HZR23, Hoa22, Hop23, HYH23, HYW^{+}24, HS20, HAC23, HWY25, HRS20]. time [JRu24, JJZ21, JWZ22, JCJ24, JHG^{+}21, JL23, Jog25, KPS23a, Kar24, KSD^{+}25, KMN25, KMS20, KGK25, KP23, KR24, KC23b, LSV21, LNNP24, LAX22, LXZ23, LLX23, LZA23, LT23, LLX24, LY25, LWL25, LAL^{+}24, LSRL24, LY21, LYL^{+}22, LMLZ23, LS23, LLLL24, LD25, LYLH25, LMST20, LP22a, ŁWP^{+}23, ŁSSP23, LCT22, LYL25, MH22, MFZX23, MC24, MSHB22, MMC21, MMN24, MMAK20, MMAK21, MJVH24, MLLL24, NW24a, NH24, OCCM21, OAH^{+}24, PG21, PRN^{+}23, PLZZ25, Pło22, PT24, PS21, PD25, PK25, QHHW20, QWC^{+}23, QL24, QXCG20, RDF24, Rav20, RL23, RYAM21, RMBC24, RR22, RKK25, RKC25, SOPH25, SC21a, SM22a, SG23, SB24b, SB25, SHM21, Šeb24, ySRC22, SLW22, SCG23, SK21, SW21, SC21b, SXLZ21, SJZ22, Tan24, Tan25, Top22, TH21, UHLZ24, UHL^{+}24, WDGC20, WS21a, WC21, WZ21, WLX22, WZY22, WZ22, WL23, WYP23, WZD24, WWLG24]. time [WQ24, WYW^{+}24, WJZ^{+}24, WCC25, WYLL25, Wat22, Waz20, WLW^{+}22, WLD22, Wie23, Wol20, WZCS20, WTQ^{+}20, WCY^{+}24, WLJF25, WS25, XCFC21, XZ25, XLLA20, XSZ^{+}23, XL25b, XHL21, XHL24, YQZT21, YTK21, YGZ^{+}22, YL22, YS23, YLL24, YHL25, YLZ^{+}24, YFK20, YFK23, Yua23, YWZ^{+}23, YPZ^{+}21, ZvNH25, ZWOM21, ZZW20, ZJJ20, ZLT^{+}21, ZDWR22, ZDLC22, ZRL^{+}22, ZF22, ZLQ23, ZAH23, ZYC^{+}24, ZYD24, ZZZ24, ZWCS20, ZGL23, ZA23, ZLJT20, ZNH24, ZX20, ZZB20, ZXQQ22, ZWS22, ZWW^{+}23, ZQS24, ZAPK23, dSG20]. time- [MJVH24]. time-accurate [YTK21]. time-adaptive [HDG23]. time-consuming [LCT22]. time-delay [YGZ^{+}22, ZWS22]. Time-dependent [KHM25, AS21, AN23, BDV23, BLF23, CLH22, CMR22, CG25, DLN24, GDA21, GHSR23, Gra21, HJM23a, HYH23, JRu24, JL23, LMLZ23, LS23, ŁWP^{+}23, MH22, PG21, Pło22, PT24, QHHW20, RL23, RYAM21, Šeb24, SLW22, SC21b, Wat22, Waz20, WLJF25, WS25, Yua23, ZA23]. time-discrete [GGRBRG20a]. Time-Domain [BWR21, ADD22, ADGG25, GMS23, HB25, LZA23, LYLH25, XL25b, ZYC^{+}24, ZAPK23]. Time-fractional [FRW22, AD20, AMK25, BS23a, Ben25, CS20b, CLZW20, DHVT20, FKW22, FSW20, GP22b, GAM24, GCMS21, HYW^{+}24, HS20, HAC23, HWY25, JWZ22, KP23, KR24, LT23, LWL25, LSRL24, LY21, LYL^{+}22, LD25, MFZX23, MC24, MMAK20, MMAK21, OAH^{+}24, PK25, QL24, QXCG20, RDF24, RR22, RKK25, RKC25, SG23, SCG23, Tan24, UHLZ24, UHL^{+}24, WS21a, WC21, WZ22, WQ24, WCC25, XCFC21, XZ25, YQZT21, YL22, ZZW20, ZDWR22, ZAH23, ZLJT20, ZNH24, ZX20, ZXQQ22, ZQS24]. time-harmonic [LAL^{+}24]. time-independent [BV23]. time-marching [WYP23]. time-periodic [Wol20]. Time-space [SF25, SHM21, TH21, ZGL23]. Time-splitting [GM21a]. Time-stepping [RH21, BWMW23, DP25, DR25b, HJM23b, Hoa22, KP23, LAX22, PD25, Rav20, ySRC22, WYW^{+}24, YFK20, ZZW20]. time-steps [SJZ22]. time-varying [Ben25, WWLG24, WTQ^{+}20]. time/space [SK21]. times [HM20]. Timoshenko [BDS24, HV21, JSSS24]. tiny [ANH^{+}21]. TiO [AWA^{+}22]. tissue [BGG^{+}23]. tissues [SAVHB22]. TMoL [BDL20]. Toeplitz [ZZL22]. tokamaks [BNC^{+}23]. tomography [CEB24, GPRH^{+}24, HKS21]. tool [BRR25]. toolbox [BRS21, GSW21, RHR^{+}21]. toolkit [KTXP21]. tools [NES23]. top} [dSMAdOW22]. topography [CBI^{+}23, XDL^{+}25]. topological [MR24, Par22b]. Topology [LQZ25, LZG^{+}23, OKSS21, SMR^{+}24a, VSG21, WFW^{+}24, XYT^{+}21, YYY^{+}22, ERHL22, FMP^{+}24, NIT^{+}20, NNNT22, NNTCB25, UuIUK21, XXYL23]. toric [YJZ20]. Torrey [XLLA20, YLL24, ZLT^{+}21]. torsion [GM21b, MHHR20]. total [CS23b, LLN24a, LCN20, NZ24, Shi24]. TPFA [AEZ24]. TR [BM21]. TR-BDF2 [BM21]. trace [CXXZ23, CM20a, HOY24]. tracer [PBR24]. tracking [BJyST24, HJ23, HJ25, JICRMG^{+}20, LLZ21, ZWW^{+}23, NM23a]. traffic [CML^{+}25, MZCK25, TKT^{+}24]. train [CLXZ25]. Training [JKC24, KC22a, LWG^{+}25b, WLH24]. transaction [TC25, TZ20]. Transfer [ZSH^{+}21, Abd25, ANH^{+}21, AGA^{+}21, AL23, BWZ22, BPV24, DK23, FL22, GZAR23, HANA24, HLCW24a, IBY^{+}24, JSBZ22, JSZ24, JWW^{+}21b, LYRW21, LGZC23, LLJ23, Liu25b, MAT^{+}20, MLJ^{+}20, MLHS23, MBSEA24, PS20, RG21, SADW21, SSW20, TSARR21, YLY20, YCZ^{+}21, ZHHL23, ZSL22]. transform [DCSAM25, SZS20]. transformation [Cos23, HKH^{+}24, HXW23, HWY23, HXZ^{+}22, KLK^{+}22, LSZ24b, Rob25, RLSC23, WXF22, YKMZ21, ZXG^{+}24]. transformations [JWZX25, SNHZ25]. transformed [ZA23]. Transient [AIG^{+}21, SWZC23, VBDR24, Abd25, BMH22, BIMS24, BLRRS20, BKT^{+}23, CRDS20, DBH^{+}21, DWY^{+}25, FMS22, HQLJ20, JZ25, LGZC23, MLJ^{+}20, NIT^{+}20, SFR21, SZCW24, ZDLC22, ZRL^{+}22]. transition [BTYM^{+}21, TK20, WS21b]. transitional [APGK24]. translations [RKRRL22]. transmissibility [LBTVG22]. Transmission [MGD23b, AMV23, BCM25, FWBY25, GO23, Krz22, RTA20, TCA22a, WBY23, ZH23a]. transport [BBK21, BGG^{+}23, BAB^{+}20, BSRVC20, CLL21, CRPdT20, DO23, GGH25, GKP25, HHR24, KTFI23, KGW^{+}21, Kri20, KOMK20, LHC^{+}22, LWL25, LWFS24, MWY21, MMC21, Nde25, NKPA^{+}25, PWBS25, PCLZ25, PBR24, RWaZ23, RWZG24, SAVHB22, SR22, SLC^{+}24, TO20, TCA^{+}22b, WJ25, Wie23, YLY20, YCGL25]. transportation [HKJ23]. transverse [NML23, XL25b]. transversely [GRR20]. TransVNet [CLXZ25]. Trapezoidal [ES20]. trapping [WTQ^{+}20]. treatment [AOL22, ASF^{+}25, CCGR24, LSRL24, NM23b, PGK20, SMŁWL25]. treatments [GF21, ZZ20b]. tree [ABLS23, HMW21, WSGT21]. Trefftz [LFY^{+}21, MNB23, PSSW20, Sch21, XFX22, Yua23]. Trefftz-based [MNB23]. Tresca [AAK24]. tri [MN23b]. tri-harmonic [MN23b]. triangle [FSS^{+}20, FJZC20, GKK^{+}20]. triangles [GK21]. triangular [AA20, CL20, ÇM20b, HL22, LLMC25, LZL^{+}25, NY22, Sau23, SS22, WKC21, XW24a, ZC23, ZZ23a, ZZ24a, ZW20a]. triangulations [BEIL21]. triblock [WZY22]. tridiagonal [JL19, JL20]. tridirectional [PCL25]. trigonometric [SZS20, SRD25, WCL20]. trilinear [WXH^{+}25]. trilinearly [KV22]. trimmed [Pup25]. triple [KWLW24, ZWCS20]. triple-coupled [KWLW24]. triple-relaxation-time [ZWCS20]. triquadratic [Liu25a]. tropical [APK^{+}25]. Trotter [WHT22]. truncated [NCCD24, WNF^{+}25]. truncation [KWK24]. truss [NNH^{+}25]. TS [JYHS23]. TSDT [JCP^{+}25]. TT-M [CLLG23, LFY^{+}20]. TTTS [SQ23]. tube [LZL^{+}25, QhCZ^{+}20, SB24b]. tubular [Koc22]. tumbler [XGL21]. tumor [ASGGRG23, BGG^{+}23, CRL^{+}25, HW25, KN20, LWZ22, PVC^{+}25, PSDP22, SWWZ23, YRC25]. tumor-growth [SWWZ23]. tumor-induced [BGG^{+}23]. tumoral [NCRGVR21]. tumors [YMTP25]. turbidity [GDA^{+}22]. turbine [HLEF20, YL20b, ZY21b]. turbulence [BWG25, DAAEA25, PWWJ23, PWJ^{+}24, US20, WS21b, YLY20, YOF20]. Turbulent [MIOK21, DAAEA25, EFF21, EGF^{+}23, NAE22, PGK20, SMR^{+}24a, TGL21, WZC^{+}20a, WZC^{+}20b, XL22, XRS25, ZCXG20, ZY21b]. Turing [AMS24]. TV [CLXZ25, ZZ23b]. twins [TCA^{+}22b]. twisted [MAT^{+}20]. Two [BZY20, BCM25, BABS20, CLH22, CV20, DY25, DZ24, ERHL22, FMS22, Fur25, JMCC24, LXJ^{+}21, LT23, LC21, NK25, Oru20, Wan24, WLC24, WZ25, XF23, YA23, YLCC21, ZGC24, ZW25, ZL25, ZST22, ZH23c, AI25, AMN22, AIG^{+}21, AEZ24, ASSV20, BLLY24, BV23, BS23b, BF25, BMP21, BBK24, BBDM21, BGKV20, BDS^{+}22, CD20, CML^{+}25, CO20, CLL21, CLLG23, CDLW20, CLZW20, CLH21, CXJ25, CC25, CZ20, CYQW21, CHL21, CCWC23, CCGRL21, CQS25, DBH^{+}21, DZ22, DLC22, Der23, DGQT20, DFGS24, DWY^{+}25, DS25, Du24, EH23, FAM25, FX22, GLL^{+}20b, GNZ24, GPY22, GHAU24, GM22, GHLS25, GHSR23, GK21, GWN21, GS23, HZJ^{+}22, HX22, HYW^{+}24, HGW21, HJXW23, HQL23, HHG20, Ila20, JKS25, JRu24, JJZ21, JSZ24, JWZX25, KSS22, KP22b, KK23a, KHH25, KLY25, KGK25, KP23, LHC^{+}22, LZ20b, LD22, LCW22, LY23, LLZL24, LAL^{+}24]. two [LLW23a, LFY^{+}20, LFY^{+}21, LCC22, LC22, LCH23, LRZ23, LLJ23, LLW^{+}23b, LLLL24, LL24a, LP22c, LZ25c, MZWD23, MFZX23, MrZQ24, MG25, MAT^{+}20, MR22, Mit22, MN23b, NRV22a, NML23, Oru22, PH25, PL24, PK25, QD23, QX23, QHHW20, QXCG20, RR22, RKC25, SB24a, SJ24, SQL21, SSD23, SH22, SWC24, Tan24, TQX23, TCW23, TVKN21, Vak25, WYH20, WFH^{+}20, WYS^{+}20, WCH20, WC21, WZ22, WYP23, WQ24, WJZ^{+}24, WWYL25, WMMH25, WYLL25, WWZX25, WS20a, WZXN21, WWH23, Xen20, XW24a, XLLA20, XLS22, XLSF25, YK20, YJL22, YZN22, YLL24, YLZ20a, YFK20, YLWZ20, YCWS20, YPZ^{+}21, ZFZ20, ZLPR20, ZJJ20, ZLT^{+}21, ZY21a, ZHYL21, ZFL22, ZZS23, ZGW23, ZF24, ZLL20, ZZ20b, ZL23c, ZZY^{+}24, ZLZ25b, ZLJT20, ZZB20, ZW21b, ZSL22, ZWY23, ZHPW23, ZFQ24, dlRRG21, WJC^{+}25]. two- [EH23, JWZX25]. two-asset [GM22]. Two-dimensional [BABS20, Fur25, AMN22, BV23, BBK24, CML^{+}25, CO20, CLL21, CLLG23, CDLW20, CC25, CZ20, CYQW21, DZ22, Der23, DY25, DS25, FAM25, FX22, GHLS25, GWN21, HYW^{+}24, HGW21, HJXW23, HQL23, HHG20, Ila20, JKS25, JRu24, JJZ21, KSS22, KK23a, KGK25, KP23, LZ20b, LFY^{+}20, LFY^{+}21, LRZ23, LLJ23, LZ25c, MFZX23, MG25, MR22, Mit22, MN23b, Oru22, PK25, QX23, RR22, RKC25, SB24a, SJ24, SSD23, TQX23, TVKN21, WFH^{+}20, WC21, WZ22, WQ24, WWYL25, WYLL25, WZ25, XLLA20, XLS22, XLSF25, YLL24, YFK20, YLWZ20, ZLPR20, ZJJ20, ZLT^{+}21, ZZS23, ZF24, ZLL20, ZL23c, ZZY^{+}24, ZLZ25b, ZLJT20, ZZB20, ZSL22, ZWY23, WJC^{+}25]. Two-grid [BZY20, CLH22, DZ24, LT23, Wan24, WZ25, ZGC24, ZL25, ZH23c, CLZW20, CLH21, CXJ25, HZJ^{+}22, HX22, LD22, LLZL24, LCH23, LLW^{+}23b, LLLL24, Tan24, TCW23, WCH20, WS20a, YJL22, YZN22, YLZ20a, ZFZ20, ZHYL21]. two-layer [LCC22, LC22, QD23]. two-layered [AI25]. Two-level [CV20, JMCC24, YA23, ZST22, BS23b, CCWC23, DLC22, GPY22, LL24a, YPZ^{+}21, dlRRG21]. two-parameter [SH22]. two-parameterized [ZZ20b]. two-patch [BGKV20]. two-phase [AEZ24, BLLY24, BBDM21, CCGRL21, CQS25, GNZ24, KP22b, KHH25, LHC^{+}22, LLW^{+}23b, MAT^{+}20, NRV22a, Vak25, WYS^{+}20, WYP23, WMMH25, YK20, YCWS20, ZGW23, ZFQ24]. Two-point [NK25, LLW23a, Xen20]. Two-scale [ERHL22, BMP21, CHL21, DFGS24]. two-sided [JSZ24, LCW22, SQL21, ZFL22, ZHPW23]. two-stage [CD20, DWY^{+}25, GHSR23]. Two-step [BCM25, YLCC21]. two-way [MrZQ24]. type [AN20, BHS22, BHL^{+}23, BGD22, CDM24, CGK24, CSS23, CCGRL21, CGI^{+}22, EGP20, FLS23, FG25, GPA21, GK21, GDŁW23, GCJ23, HLW23, JKS25, JAMX24, KC22b, KCD^{+}23, KC23b, Lee24, LQ20, LZ20c, LY25, LS25a, LYL^{+}22, LZLN25, LYLH25, MZM21, MKP23, PG25, Par24a, PV20, QGQ22, RZSL23, SMR^{+}24a, ST20a, SX25, TPOT20, WYHQ24, WL20, Xen20, XFS23, XTZ^{+}21, YSB^{+}24, YL20a, YLW^{+}22, Yan24, ZZ23a, ZZ24a, ZGL23, ZZZD24, ZWS22, ZMZW24, CCGV22, WDYZ24]. types [ZZ20b]. typical [ZL23b].

U [LL25b, LZS25]. U-Net [LL25b]. U-shaped [LZS25]. U-WNO [LL25b]. Ultra [WJZ^{+}24, MS23]. Ultra-weak [WJZ^{+}24, MS23]. ultraconvergence [HLX20]. ultrasound [AP25, AEF23, FJP22]. Ultraweak [Füh21, CHD24]. unbalanced [PWBS25]. unbounded [AAC23, FMS22, GCMS21, IK24, KK23b, RLF21, YW24]. uncertain [IJS^{+}24, TJYL22]. uncertainties [FSH^{+}23, RHB25]. Uncertainty [CFG^{+}23, KG20b, AS21, BLLM25, CL24b, NNH^{+}25, SC21b, YT22]. Unconditional [CZ23, LSW20, LF24b, LS25a, SZ22b, SZ24b, SQ24, WYW^{+}24, XP22, ZHZ23, JWL20, LX24a, LaZZ24, SXKL24, WS25, ZY23]. Unconditionally [CGH24, GHMD22, GKP25, GWN21, HLJ25, LL25a, LQY22, Tan25, WLL23, WaZWZ23, YKK25, ZO21, ZZW25, CGK24, DMX24, FLMS24, GLF25, GGRBRG20b, HYJ20, LWSZ21, LJ24, LTYZ25, PHL21, QH22, QZH23, RC25, SY25, WZY22, WHH22, XKXL22, ZD22, ZS23a, ZWT23]. undamped [WH21]. Understanding [YRC25, LE24]. underwater [XFX^{+}24]. undulation [BP20]. unequal [BM24b, CTZ22, LZ22, WZTZ23]. unequal-sized [CTZ22, LZ22, WZTZ23]. unfitted [BNV22, BDV23, LS22a, NF25, OSZ22]. unicellular [NM23b]. unified [CG21, CKBTS24, DD21, DLN24, DG22b, HLH^{+}20b, HWZQ23, Hof20, LXP^{+}24, SCCS21, TWH^{+}22, XSL24b]. Uniform [SvV20, ZL22a, ZZ24a, ZZ25b, GW25, HV21, HKL^{+}23, Ima20, JCJ24, LL20a, LXJ^{+}21, QQX21, SRD25, Tan25, TMML23, WH25, WS25, XLSF25, ZZW20]. uniformly [SG23]. uniquely [CFG^{+}25]. Uniquenes [BJG23]. uniqueness [BMZ20, ZNC24]. unit [JFS14, JFS20, MC24, Ma25b, ZCZ^{+}21]. united [CFP22]. Unity [HNLPP20, AM22, Du24, FAM25, LQF25, RDF24]. unknown [LLZL24, OAH^{+}24, YEO21]. unmanned [JJOG20]. unsaturated [RFG^{+}21, XLZ^{+}21]. unstable [Bel24]. Unsteady [BWZ22, BWZ23, US20, BEKV21, CSI23, CSS25, DLC22, DCCM23, GJS20, HSR25, KNBR22, LHH22, LLZ^{+}23, LL24a, LCT22, OEASY24, QhCZ^{+}20, Yan20, YWPS24, YPZ^{+}21, ZSF23]. unstructured [AFBM25, AN21, BCT21, BHL24, CSI21, EH23, GMS23, HZ24a, KN24, KCW22, LMZ23, LZG^{+}23, LSY^{+}24, NDC22, WM23, YA21, ZCYL24]. unsupervised [CKFR24]. Unveiling [Par24b]. updates [EED^{+}25]. upflow [BAM^{+}21]. upper [CL20, Ku24, LYH24, XW24a]. Upwind [LY24, Nde25, ASGGRG23, LWFS24, LCT22, MZCK25, ST20b, ZHPW23]. upwinding [Lee21, LMBB24]. urban [AADMA23]. usage [RHR^{+}21]. use [BNB20, EMSA25, Loh21, NES23, Oru25]. used [Kri20, PSDP22]. Using [TRSM23, UBP^{+}25, YL20b, AS21, AI25, ASB25, ASH^{+}21, AIG^{+}21, AFM^{+}24, ANN25, AFBM25, BDMX21, BS23a, BGG^{+}20, BP20, BD20, BDJ20, BCK24, CRM22, CRRS21, CM22, CHZ23, CFG^{+}23, CL22b, CDG^{+}23, DLN24, DAAEA25, DS25, EED^{+}25, EGF^{+}23, EBT24, EFK^{+}23, EFFR22, Fur25, GCSF21, GM21b, GLC24, Gun22, GHK25, HKJ23, HLH24, HSR^{+}23, HXZ^{+}22, HTPV25, IJS^{+}24, JH24, JMS^{+}24, JZ25, JCP^{+}25, JSZ24, JHS24, Kam21, KEUY23, KG20b, KRESA21, KSD^{+}25, KLK^{+}22, KHY25, KMS20, KKJ^{+}22, LH24, LCY23, LQYJ23, LDW^{+}24, LLG25, LL25c, LLZ^{+}25, LML23, LJY^{+}25, LP22c, LYXC21, MCZL20, MIOK21, MB24, MRN^{+}20, MJVH24, MJWD20, MKM^{+}24b, MH25, MVHÖ23, MFF25, NM23a, NE22, NIT^{+}20, NP24, NNH^{+}25, NR22, OSO^{+}23, PDB22, PND^{+}24, RNN22, RSAG^{+}24, RBSC24, RG21, RCSM23, RGS^{+}20, SB24a, SAVHB22, SB24b]. using [SGF^{+}21, SS20, SHXZ24, TWH^{+}22, UR24, VSG21, WCL20, WCZ24, WFW^{+}24, WWG^{+}25, WQS25, WS21b, XYT^{+}21, XLCP20, XLZ^{+}21, YPC^{+}21, YCCW24, YA21, YYY^{+}22, ZP24, ZKTK21, ZD21, ZJZ24, ZL23c, ZXG^{+}24, dlRGR20]. utility [LLZ20, XXY20, YZL20]. utility-indifference [YZL20]. utilizing [DDFZ23, SZ21, TT25]. Uzawa [AXLW24, HC22b, MZM21, YM21]. Uzawa-type [MZM21].

V [GV20]. vacancy [YWT22, ZWT23]. vaccination [GV20]. vaginal [SSK^{+}23]. Validation [SAR20, APGK24, BAM^{+}21, HMŁW^{+}24, KP22b, KTFI23, XL22]. valuation [LLZ20]. value [ADG23, CLZ23a, CZ20, Gut20, HZ21, HSR^{+}23, HYW^{+}24, JKC24, KC22a, LX24b, MN23b, QX23, Xen20, ZX24, ZQS24]. valued [GO23, KP20, TLJ24, WT25]. valve [NDB^{+}24]. vapor [HHY^{+}22, KTFI23]. Variable [LCC^{+}24, ACRBT23, BS23a, BRÁ^{+}20a, BGH20, BABS20, CBLZ24, CLZW20, CQG^{+}24, CGH24, DYC^{+}22, DAS20, DC22, EFFR22, FSW20, GP25, HMW21, HH23, HYZZ24, JM23a, KS21, KKS25, Ku24, KGK25, LLSY24, LK20, LGL20, LL20b, LL22a, LT23, LDM25, LLLL24, Liu25a, LYL25, MFZX23, Ma25b, MMGS25, MVHÖ23, Ngo25, NTP25, PLZZ25, PS25a, QWC^{+}23, RKC25, SK25, SJZ22, WZ22, WT25, Wan25, XLLA20, XSZ^{+}23, YL20a, YL25, YLZ^{+}24, ZLT^{+}21, ZDW23]. variable- [LK20]. variable-coefficient [MFZX23]. variable-order [BS23a, CLZW20, DAS20, FSW20, HH23, KGK25, Wan25, XSZ^{+}23, YLZ^{+}24, ZLT^{+}21]. variable-stepsize [MMGS25]. variable-time-step [LLLL24, LYL25]. variably [JWZ22]. Variance [PWC^{+}23]. variant [LWL21, LZY24b, SM22b, WYT23]. variation [CS23b, LLN24a, LCN20, NZ24]. Variational [EKB20, UBP^{+}25, ZL20a, ZNC24, AMN22, BHJ21, BMZ20, BHS22, BLO22a, BFVY23, CLY25b, DD21, DMRV24, JJOG20, LP21, LMBB24, LLZ^{+}25, LXH24, LQ24, MD21, NR25a, PG21, PS21, RZ21, STM21, SHWC22, Shi24, SLC^{+}24, Tor20, XCLX24, YWL21, dSG20]. variational-hemivariational [BMZ20, XCLX24]. variationally [VR20, VBDR24]. variations [XFX22]. various [DAAEA25, KAKK23, LJY^{+}25, MAT^{+}20, ZZ20b]. varying [AA25b, AGPP22, Ben25, CBI^{+}23, KKJ^{+}22, MHHR20, WWLG24, WTQ^{+}20, ZLYL20]. vault [Boł22]. vector [AHM20, DSF24, JH24, Li24b, MH22]. vectorial [HGD21, HGD25, ZZ20b]. vectorization [AAU25]. Vectorized [MPVA25]. vectors [LLZL24]. vehicle [TKT^{+}24]. vehicles [JJOG20]. velocities [WSGT21]. velocity [BWG25, CLW^{+}24, CEL20, GJS20, Ha20, HJ23, HJ25, LLX24, PWMC23, SK25, WHC24, YW20, ZDT^{+}22, ZSF23, ZZ20b, ZLZ25b]. velocity-based [WHC24]. velocity-correction [GJS20]. velocity-vorticity-Voigt [YW20]. VEM [CGH24, PD25, SW20a]. Venant [XDL^{+}25]. vents [GdFFF25]. verification [GCC20, WWY25]. verified [ZK21]. Versatile [CW20, MCW22, GF21]. version [DW20]. versions [CGM^{+}20]. versus [LZ25b, ZR23]. Vertex [HSY25, CSI21, DWS20, MC25]. Vertex-centered [HSY25, DWS20]. vertex-centred [MC25]. vertical [HKJ23, HDG23, SBH^{+}21]. Very [BCT21, CCGR24, BMVA23, SJ24]. Vese [LQZ23]. vesicle [LaZMY25]. vessel [LH24]. via [AADMA23, AJK23, BTYM^{+}21, BBK24, CTGR24, CKFR24, CLY25a, CEB24, DFK22, DT24, FWLD22, GPA21, Hak20, HJP23, HSM^{+}23, HSR^{+}23, HM24, HLCW24a, HBC^{+}21, HJ23, JKS25, JICRMG^{+}20, Kam25, Kap24, MR24, NES23, Oru22, PMLC20, PWMC23, PP24, Pon20, SOPH25, SC21a, SQC23, SNHZ25, SGR^{+}23, UHLZ24, WZL25, YMTP25]. vibrating [DBB23]. Vibration [DYC^{+}22, ZD21, AFS24, ARV23, CWZ21, HTPV24, HTPV25, JCP^{+}25, KKJ^{+}22, LC23a, LXG^{+}24, LLZ^{+}25, LJY^{+}25, LYXC21, MJWD20, MNB23, NP24, PND^{+}24, PCL25, SFR21, WCZD20, WCW20b, XLLR25, ZWW^{+}23, dSMAdOW22]. vibrations [APS25, HHR^{+}20, OCRN22, RSAG^{+}24, XFX^{+}24]. vibro [LXP^{+}24]. vibro-acoustic [LXP^{+}24]. vicinity [BBS22]. view [Hof20, Par22c, Sch21]. viral [WKGH21]. Virtual [AN20, AMN22, AN23, DV22, DD24, DFS21, LSZ24a, TKK22, dVDMM22, AMV23, AMV25, ACPT25, ALOR25, ABLS23, AMV20, ANN25, BLF23, BBM25, BMPP22, CGM^{+}20, CHL22, CS23a, CLZN24, CLQL25, DS20, DFSV22, DMRV24, DG22b, DWS20, GCL25, HZ25, HWC24, JK22, KTY24, KS25, LHF24, LR23, LSF25, LD20, LSWY21, MCW25, Mas23, MN23a, MS22, MS20b, MCV24, PZL23, TLZF20, TKK23, Vis24, WW20b, WWW21, WCWZ22, WWS23, WWCJ23, WHC24, WW25, Yu23, ZF22, cZsCyC23, ZF24, ZZZD24, ZCW24, BLMV24, CLL^{+}24, SBMS22, dVDR20]. Visco [BFE^{+}21, XLCP20]. Visco-acoustic [BFE^{+}21, XLCP20]. viscoelastic [AAC23, AZL20, AEF20, BGR23, BFRA21, DYC^{+}22, DCSAM25, HJM23b, HV21, JSSS24, LL20a, LC23a, MLS22, Oru20, SFR21, SCdH21, SWW22, WRMB23, YCZ^{+}21, YJL22, YMDW25, cZsCyC23, ZYD24]. viscoplastic [FNGDV23, GAM22, RUMT^{+}24]. viscosity [ACRBT23, BDGS25, GP25, RUMT^{+}24, SK25, YY20b, YWZ^{+}23]. viscous [CMY23, GF21, HSWG21, LZF23, MWYY22, MG25, PM23, ST20b, SW24b, XYT^{+}21]. Vlasov [GHS23, MW20, RMBC24, WCW20a]. VMS [BRÁ^{+}20a]. VOF [AFBM25, AI25, SJ24]. Vogelius [AJLM24]. void [WFW^{+}24]. Voigt [BGR23, SWW22, YW20]. volatility [JZ24a, KHY25, LZ20d, SLM23, TC25, TZ20]. Volterra [JKS25, TSW^{+}25, WS20b]. Volume [YA21, AEZ24, AAM^{+}24, AN21, BHL24, CLH22, CLZ23b, CSI21, CL21, CCGRL21, DCCM23, ERHL22, EH23, Fan23, FZS22, FXLS25, GMS23, HLEF20, HSY25, JK22, KLK^{+}22, Lin24, LMZ23, LWFS24, LN22, LZB25, MRN^{+}20, MC25, MRPK21, NY22, NFS24, NK25, PL25, SZ24a, WZC^{+}20a, WZC^{+}20b, WXZW21, WML21, WCZ24, WLQ22, XZSY23, XLSF25, YYSL20, ZBCS24, ZCYL24, Zha21, ZLT^{+}21, Zha22a, ZC23, ZSGT24, ZZ25a, ZW20a, ZSY22, ZZW23, ZS24, ZQ25, AFBM25]. Volume-of-Fluid [AFBM25]. volumes [KV22]. volumetric [Cha24]. vortex [HHR^{+}20, WZY23, WCZD20]. vortex-induced [HHR^{+}20]. vorticity [ADR24, ACRBT23, LLX23, SK25, YW20, ZLZ25b]. vorticity-streamfunction [LLX23]. vorticity-velocity [ZLZ25b]. Vries [AJK23]. vs [JBKW23]. Vulanović [ZZ25b].

Waals [BABS20]. wafer [ZYM24]. waLBerla [BEG^{+}21]. wall [AA25b, zLqLZ^{+}24, MIOK21, PGK20, PZZ23, SA22, US20, YLY20]. walled [NP24, WCW20b]. Walles [RUMT^{+}24]. warm [CSCY24]. warm-start [CSCY24]. Washizu [CLL^{+}24]. Water [Bou24b, MAT^{+}20, ABC^{+}22, AWA^{+}22, BPR22, CBK24, CBI^{+}23, GO20, HTDT23, JWW^{+}21b, LMD22, WZZ20, WYS^{+}20, WZTZ23, WKC21, XLZ^{+}21, ZJCG21]. Watson [JZ24a]. wave [AKO21, Ade20, ADGG25, ACLP22, ACLP24, Ale23, AP25, AL23, Alt23, AJK23, ABV23, AMM23, ASZ24, BLO22a, Bra20b, BPID20, CYQW21, CWX22, Der23, DCSAM25, DW20, FBK^{+}24, FLMS24, Fur25, GMS23, GCSF21, GWYR22, GHL25, GCL25, HMW21, HSWG21, HWW21, HKR22, HZR23, HJXW23, HLY22, HJ23, HJ25, JHG^{+}21, JXK20, KRO24, KP23, KC23b, LWX21, LT23, LF24b, LCE23, LRZ23, LL24b, LZY24b, LBPC20, MRAC24, MLLL24, MLL25, Oru20, PSSW20, PLVL21, PD25, QL24, QhCZ^{+}20, SM22a, SM23, SZ22b, SCG23, SX25, SCdH21, SNL23, SLYL22, SWZC23, SW24b, TRL23, TGARGT^{+}23, WYH20, WFH^{+}20, WTSZ20, WZ22, WF24, WZD24, WGeL24, WSCP24, WLC24, WYW^{+}24, WZL25, WS25, XZ23, XLCP20, XLYY25, YWL23, YHL25, YWZ21, YH21b, Yua23, ZDLC22, cZsCyC23, ZAH23, ZYC^{+}24, ZPC20, ZWY23]. wave-induced [SNL23]. wave-stripe [WTSZ20]. wave/boundary [QhCZ^{+}20]. waveform [BFE^{+}21, SM23, ZJ25]. waveguide [HD21]. waveguides [HGD25, MK20]. Wavelet [KSS^{+}20, AAU25, ACXI24, FKW22, JM23a, KG20b, LL25b, SJM24, WX22, YX21]. wavelet-homotopy [ACXI24, WX22]. Wavelets [KSSS23, DYC^{+}22, KJ20]. Wavenumber [LSW22]. Wavenumber-explicit [LSW22]. waves [ADD22, CYC^{+}23, DMG23, LZ20c, RL24, SDJ^{+}22, WTQ^{+}20, XRS25]. wavy [AIG^{+}21, AAWJ20]. way [MrZQ24, PLVL21]. WCSPH [HB24]. Weak [GQZ23, KD23, AZHY22, CC25, DG22a, DGQT20, DR25b, GDA21, GHL24, HWWZ24, JM20a, JCJ24, KSJ23, LWW24, LN21, LSGN23, LXZZ24, MGD23a, MS23, Ngo25, PK23, PK25, QSW23, TZX23, Top22, WW20a, WWL22, WGC23, WJZ^{+}24, WF25, XZT23, YZZ20, ZGC24, ZZ25a, ZXQQ22, ZX22]. weak-form [JM20a]. Weakly [HEDT20, APPS25, Bel24, BS21, GP22a, HXZ^{+}22, KP22b, LDH24, LMLZ23, LD25, MKM24a, XJ21, YZT21, ZGL23, ZX20]. weakly-singular [GP22a]. wear [HHJO20, WXCL25]. weather [LLZ20, NS24]. web [CM22]. web-splines [CM22]. wedge [RG21]. weight [Yan25a]. Weighted [OSO^{+}23, RR23, AAD24, BMP^{+}24, CSB^{+}21, CM24, FFC^{+}25, GCJ23, JZHZ22, KMP20, LCC^{+}24, WZTZ23, WLL23]. weighting [GYL25, LWG^{+}25b]. weights [GCJ23]. Weiser [BHL^{+}23]. welding [MR24]. Well [CCF^{+}21, CBI^{+}23, FRW22, LLWX25, MKM24a, SSY22, YHL25, dCLW25, CZLQ22, CG22, GI21, GdFFF25, LMD22, RWZG24, SPZL21, WZZ20]. Well-balanced [CBI^{+}23, CZLQ22, GdFFF25, LMD22, RWZG24, WZZ20]. well-defined [SPZL21]. Well-posedness [CCF^{+}21, FRW22, LLWX25, MKM24a, SSY22, YHL25, dCLW25, CG22, GI21]. Wendroff [PCLZ25]. WENO [BD20, CTZ22, LCY23, LZ22, TZ25, TL22, WZZ20, ZQZ20]. WENO5 [NFFR^{+}24]. wet [HLEF20, XDL^{+}25]. wet-dry [XDL^{+}25]. wet-steam [HLEF20]. wettability [ZKBC21, ZBKC22, ZDLL22]. Wetting [ZTW22, AFBM25, HKL^{+}23, SMŁWL25]. wide [CT22]. Wiener [KG20b]. wildfire [GI20, RNMÖX24]. Williamson [OSO^{+}23]. Wilson [SL21, SL24]. wind [WSGT21]. wing [LYH24, RGS^{+}20]. wing-upper-skin [LYH24]. Winkler [YLFL20]. wires [CRDS20]. wise [NRV22a, WWY25]. within [HDG23, RUMT^{+}24, WSGT21, ZAPK23]. without [Bou24b, GNB24, LNNP24, Par22c, Par24a, SXCB24, XXYL23]. WNO [LL25b]. Workpiece [Cao21]. Workpiece-scale [Cao21]. wormhole [DG22b, ZQY^{+}23]. WSK [LZY24a]. Wu [HH23, KG20a].

X [TK20]. X-FEM [TK20]. XB [LWL21]. XFEM [WCLQ25].

yield [LLZ21].

Zakharov [GZ23, GFXZ25]. Zhabotinsky [KEUY23]. Zhang [HH23, KG20a]. Zienkiewicz [DST25]. ZK [LL25c]. ZK-MEW [LL25c]. zone [BBT25].

References

Alshammari:2020:PSS

[AA20] Fehaid Salem Alshammari and F. Talay Akyildiz. Pseudo spectral solution of extended Graetz problem for combined pressure-driven and electroosmotic flow in a triangular micro-duct. Computers and Mathematics with Applications, 80(5):990– 1008, September 1, 2020. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                                      /www.sciencedirect.com/
                                                      science/article/pii/S0898122120301723.

Abbaspour:2021:TEM

[AA21] Fatemeh Abbaspour and Hadi Arvin. Thermo-electro-mechanical buckling analysis of sandwich nanocomposite microplates reinforced with graphene platelets integrated with piezoelectric facesheets resting on elastic foundation. Computers and Mathematics with Applications, 101(??): 38–50, November 1, 2021. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                                      /www.sciencedirect.com/
                                                      science/article/pii/S0898122121003369.

Ahmed:2025:EFP

[AA25a] Elyes Ahmed and Saber Amdouni. Equilibrated flux a posteriori error estimates and adaptivity for nonlinear and doubly degenerate elliptic problems. Computers and Mathematics with Applications, 195(??):239– 264, October 1, 2025. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122125003086.

Alsoy-Akgun:2025:NCR

[AA25b] Nagehan Alsoy-Akgün. Natural convection resulting from exponentially varying wall heating in a square enclosure. Computers and Mathematics with Applications, 197(??):235–258, November 1, 2025. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122125003578.

Athanasiadis:2024:MDI

[AAA24] C. E. Athanasiadis, E. S.

Athanasiadou, and I. Arkoudis. The MFS for direct and inverse electromagnetic scattering problems in chiral media. Computers and Mathematics with Applications, 163(??):201– 225, June 1, 2024. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122124001330.

Adler:2025:NMS

[AAA25] J. H. Adler, A. S. Andrei, and T. J. Atherton. Nonlinear methods for shape optimization problems in liquid crystal tactoids. Computers and Mathematics with Applications, 187(??):231– 248, June 1, 2025. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122125001506.

Anderson:2021:MMF

[AAB^{+}21] Robert Anderson, Julian Andrej, Andrew Barker, Jamie Bramwell, Jean-Sylvain Camier, Jakub Cerveny, Veselin Dobrev, Yohann Dudouit, Aaron Fisher, Tzanio Kolev, Will Pazner, Mark Stowell, Vladimir Tomov, Ido Akkerman, Johann Dahm, David Medina, and Stefano Zampini. MFEM: a modular finite element methods library. Computers and Mathematics with Applications, 81(??):42–74, January 1, 2021. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122120302583.

Abouelregal:2023:MTI

[AAC23] Ahmed E. Abouelregal, Bekir Akgöz, and Ömer Civalek. Magneto-thermoelastic interactions in an unbounded orthotropic viscoelastic solid under the Hall current effect by the fourth-order Moore– Gibson–Thompson equation. Computers and Mathematics with Applications, 141 (??):102–115, July 1, 2023. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S089812212300130X.

Alvarez:2024:RTI

[AAD24] María Luz Alvarez, María Gabriela Armentano, and Ricardo G. Durán. Raviart–Thomas interpolation in fractional weighted Sobolev spaces. Computers and Mathematics with Applications, 168 (??):39–45, August 15, 2024. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122124002463.

Abbaszadeh:2023:ROM

[AADMA23] Mostafa Abbaszadeh, Mohammad Ivan Azis, Mehdi Dehghan, and Reza Mohammadi-Arani. Reduced order model for simulation of air pollution model and application in 2D urban street canyons via the meshfree gradient smoothing method. Computers and Mathematics with Applications, 140 (??):195–210, June 15, 2023. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122123001049.

Ali:2022:MEC

[AAG22] Liaqat Ali, Bagh Ali, and Muhammad Bilal Ghori. Melting effect on Cattaneo– Christov and thermal radiation features for aligned MHD nanofluid flow comprising microorganisms to leading edge: FEM approach. Computers and Mathematics with Applications, 109(??):260–269, March 1, 2022. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122122000104.

Ati:2024:CMF

[AAK24] Dania Ati, Rahma Agroum, and Jonas Koko. A coupled model of the fluid flow with nonlinear slip Tresca boundary condition. Computers and Mathematics with Applications, 153(??):130– 158, January 1, 2024. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122123004972.

Alikhanov:2024:NEF

[AAL24] Anatoly A. Alikhanov, Mohammad Shahbazi Asl, and Dongfang Li. A novel explicit fast numerical scheme for the Cauchy problem for integro-differential equations with a difference kernel and its application. Computers and Mathematics with Applications, 175(??):330–344, December 1, 2024. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122124004590.

Artoni:2024:HFV

[AAM^{+}24] Alberto Artoni, Paola F.

Antonietti, Ilario Mazzieri, Nicola Parolini, and Daniele Rocchi. A hybrid finite volume — spectral element method for aeroacoustic problems. Computers and Mathematics with Applications, 155(??):193–208, February 1, 2024. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122123005540.

Anwar:2020:EFN

[AAS^{+}20] Muhammad Shoaib Anwar, Rana Tariq Mehmood Ahmad, Tahir Shahzad, Muhammad Irfan, and Muhammad Zeeshan Ashraf. Electrified fractional nanofluid flow with suspended carbon nanotubes. Computers and Mathematics with Applications, 80(5):1375–1386, September 1, 2020. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122120302716.

Ahmad:2025:ESH

[AAU25] Masood Ahmad, Muhammad Ahsan, and Zaheer Uddin. Explicit solution of high-dimensional parabolic PDEs: Application of Kronecker product and vectorization operator in the Haar wavelet method. Computers and Mathematics with Applications, 186(??):1–15, May 15, 2025. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122125000884.

Asghar:2020:IFD

[AAWJ20] Z. Asghar, N. Ali, M. Waqas, and M. A. Javed. An implicit finite difference analysis of magnetic swimmers propelling through non-Newtonian liquid in a complex wavy channel. Computers and Mathematics with Applications, 79(8):2189– 2202, April 15, 2020. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122119305206.

Abide:2021:IPD

[ABC^{+}21] Stéphane Abide, Mikaël Barboteu, Soufiane Cherkaoui, David Danan, and Serge Dumont. Inexact primal–dual active set method for solving elastodynamic frictional contact problems. Computers and Mathematics with Applications, 82(??):36–59, January 15, 2021. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122120304454.

Albarakati:2022:MDR

[ABC^{+}22] Aishah Albarakati, Marko Budišić, Rose Crocker, Juniper Glass-Klaiber, Sarah Iams, John Maclean, Noah Marshall, Colin Roberts, and Erik S. Van Vleck. Model and data reduction for data assimilation: Particle filters employing projected forecasts and data with application to a shallow water model. Computers and Mathematics with Applications, 116(??):194– 211, June 15, 2022. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122121002121.

Arndt:2021:DIF

[ABD^{+}21] Daniel Arndt, Wolfgang Bangerth, Denis Davydov, Timo Heister, Luca Heltai, Martin Kronbichler, Matthias Maier, Jean-Paul Pelteret, Bruno Turcksin, and David Wells. The deal.II finite element library: Design, features, and insights. Computers and Mathematics with Applications, 81(??): 407–422, January 1, 2021. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122120300894.

Abdoh:2024:TDP

[Abd24]  D. A. Abdoh.  Three-

dimensional peridynamic modeling of damage and penetration in composite plates exposed to localized explosive blasts. Computers and Mathematics with Applications, 175(??):399–415, December 1, 2024. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122124004632.

Abdoh:2025:NEM

[Abd25] Daud Ali Abdoh. A new and efficient meshfree method to solve partial differential equations: Application to three-dimensional transient heat transfer problems. Computers and Mathematics with Applications, 187 (??):181–202, June 1, 2025. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122125001294.

Anderson:2024:CPP

[ABFPA24] Thomas G. Anderson, Marc Bonnet, Luiz M. Faria, and Carlos Pérez-Arancibia. Construction of polynomial particular solutions of linear constant-coefficient partial differential equations. Computers and Mathematics with Applications, 162 (??):94–103, May 15, 2024. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122124000907.

Anguill:2024:MDS

[ABL24] Pierre Anguill, Xavier Blanc, and Emmanuel Labourasse. A monotone diffusion scheme for 3D general meshes: Application to radiation hydrodynamics in the equilibrium diffusion limit. Computers and Mathematics with Applications, 158(??):56–73, March 15, 2024. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122124000051.

An:2023:MAC

[ABLS23] Bo An, J. M. Bergadà, D. Li, and W. M. Sang. Modified algorithms for curved and virtual boundaries in lattice Boltzmann method applications based on tree grid. Computers and Mathematics with Applications, 146(??):

71–83, September 15, 2023. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122123002729.

An:2020:NAN

[ABMS20] Bo An, J. M. Bergadà, F. Mellibovsky, and W. M. Sang. New applications of numerical simulation based on lattice Boltzmann method at high Reynolds numbers. Computers and Mathematics with Applications, 79(6): 1718–1741, March 15, 2020. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122119304857.

Appella:2025:CEA

[ABP25] Simone Appella, Chris Budd, and Tristan Pryer. Computationally efficient r-adaptive graded meshes over nonconvex domains. Computers and Mathematics with Applications, 192(??):240– 258, August 15, 2025. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122125002226.

Ali:2020:SRB

[ABR20] Shafqat Ali, Francesco Ballarin, and Gianluigi Rozza. Stabilized reduced basis methods for parametrized steady Stokes and Navier– Stokes equations. Computers and Mathematics with Applications, 80(11):2399– 2416, December 2020. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122120301231.

An:2023:SNM

[ABS23] Bo An, J. M. Bergadà, and W. M. Sang. A simplified new multigrid algorithm of lattice Boltzmann method for steady states. Computers and Mathematics with Applications, 135 (??):102–110, April 1, 2023. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122123000135.

Antonietti:2023:CDS

[ABV23] Paola F. Antonietti, Francesca Bonizzoni, and Marco Verani. A cVEM-DG spacetime method for the dissipative wave equation. Computers and Mathematics with Applications, 152(??):341– 354, December 15, 2023. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122123004698.

Abreu:2024:MPN

[ACDL24] Eduardo Abreu, Eduardo Cuesta, Angel Durán, and Wanderson Lambert. Mathematical properties and numerical approximation of pseudo-parabolic systems. Computers and Mathematics with Applications, 165 (??):163–179, July 1, 2024. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122124001743.

Aldirany:2022:OEA

[ACLP22] Ziad Aldirany, Régis Cottereau, Marc Laforest, and Serge Prudhomme. Optimal error analysis of the spectral element method for the 2D homogeneous wave equation. Computers and Mathematics with Applications, 119(??): 241–256, August 1, 2022. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122122002310.

Aldirany:2024:OAW

[ACLP24] Ziad Aldirany, Régis Cottereau, Marc Laforest, and Serge Prudhomme. Operator approximation of the wave equation based on deep learning of Green’s function. Computers and Mathematics with Applications, 159 (??):21–30, April 1, 2024. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122124000324.

Amey:2025:SPS

[ACPT25] Luis F. Amey, Juan G.

Calvo, and Josué Padilla-Torres. A Schwarz preconditioner for the smoothed lowest-order virtual element method. Computers and Mathematics with Applications, 197(??):71–87, November 1, 2025. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122125003335.

Anaya:2023:AFE

[ACRBT23] Verónica Anaya, Rubén Caraballo, Ricardo Ruiz-Baier, and Héctor Torres. On augmented finite element formulations for the Navier– Stokes equations with vorticity and variable viscosity. Computers and Mathematics with Applications, 143(??): 397–416, August 1, 2023. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122123002134.

Alvarez:2022:ASA

[ÁCS22] Mario Álvarez, Eligio Colmenares, and Filánder A. Sequeira. Analysis of a semiaugmented mixed finite element method for doublediffusive natural convection in porous media. Computers and Mathematics with Applications, 114(??):112– 131, May 15, 2022. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122122001225.

Alberini:2021:NSD

[ACV21] C. Alberini, R. Capitanelli, and S. Finzi Vita. A numerical study of a degenerate diffusion equation driven by a Heaviside function. Computers and Mathematics with Applications, 89 (??):139–149, May 1, 2021. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122121000705.

Ahmed:2024:MCF

[ACXI24] Sohail Ahmed, Zhi-Min Chen, Hang Xu, and Muhammad Ishaq. Mixed convection flow in a square liddriven cavity subject to inclined magnetic field with highly accurate wavelethomotopy solutions. Computers and Mathematics with Applications, 162(??):33– 51, May 15, 2024. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S089812212400107X.

Abbaszadeh:2020:DML

[AD20] Mostafa Abbaszadeh and Mehdi Dehghan. Direct meshless local Petrov– Galerkin (DMLPG) method for time-fractional fourthorder reaction–diffusion problem on complex domains. Computers and Mathematics with Applications, 79(3): 876–888, February 1, 2020. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122119303918.

Abreu:2021:SDA

[AD21] Eduardo Abreu and Angel Durán. Spectral discretizations analysis with time strong stability preserving properties for pseudoparabolic models. Computers and Mathematics with Applications, 102(??):15–44, November 15, 2021. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122121003497.

AlJahdali:2022:OER

[ADB^{+}22] R. Al Jahdali, L. Dalcin, R. Boukharfane, I. R. Nolasco, D. E. Keyes, and M. Parsani. Optimized explicit Runge–Kutta schemes for high-order collocated discontinuous Galerkin methods for compressible fluid dynamics. Computers and Mathematics with Applications, 118(??):1–17, July 15, 2022. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122122001936.

Aimi:2022:PPA

[ADD22] A. Aimi, L. Desiderio, and G. Di Credico. Partially pivoted ACA based acceleration of the energetic BEM for time-domain acoustic and elastic waves exterior problems. Computers and Mathematics with Applications, 119(??):351– 370, August 1, 2022. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122122002188.

Adewole:2020:FIS

[Ade20] Matthew O. Adewole. FEM-IDS for a second order strongly damped wave equation with memory. Computers and Mathematics with Applications, 80(12):2896– 2914, December 15, 2020. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S089812212030417X.

Almeida:2023:MFE

[ADFP23] Rui M. P. Almeida, José C. M. Duque, Jorge Ferreira, and Willian S. Panni. Mixed finite element method for a beam equation with the p(x)-biharmonic operator. Computers and Mathematics with Applications, 139(??):57–67, June 1, 2023. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122123000937.

Antonelli:2023:ANL

[ADG23] Fabio Antonelli, Raffaele D’Ambrosio, and Ivan Gallo. Analysis of non-linear approximated value equation under multiple risk factors and stochastic intensities. Computers and Mathematics with Applications, 140 (??):24–32, June 15, 2023. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122123001086.

Aimi:2025:ATD

[ADGG25] A. Aimi, G. Di Credico, H. Gimperlein, and C. Guardasoni. Adaptive time-domain boundary element methods for the wave equation with Neumann boundary conditions. Computers and Mathematics with Applications, 198(??):196– 213, November 15, 2025. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122125003530.

Adjerid:2022:SHO

[Adj22] Slimane Adjerid. A study of high-order immersed finite element spaces by pointwise interface conditions on curved interfaces. Computers and Mathematics with Applications, 128(??):331– 353, December 15, 2022. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122122004576.

Adjerid:2025:IFE

[Adj25] Slimane Adjerid. Immersed finite element methods for linear and quasi-linear elliptic interface problems. Computers and Mathematics with Applications, 197(??):19–42, November 1, 2025. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122125003281.

Abbaszadeh:2020:EAI

[ADKH20] Mostafa Abbaszadeh, Mehdi Dehghan, Amirreza Khodadadian, and Clemens Heitzinger. Error analysis of interpolating element free Galerkin method to solve non-linear extended Fisher–Kolmogorov equation. Computers and Mathematics with Applications, 80 (1):247–262, July 1, 2020. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122120301188.

An:2021:SBD

[ADL^{+}21] Xiang An, Bo Dong, Weizhong Li, Xun Zhou, and Tao Sun. Simulation of binary droplet collision with different angles based on a pseudopotential multiple-relaxation-time lattice Boltzmann model. Computers and Mathematics with Applications, 92 (??):76–87, June 15, 2021. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122121001255.

Akbas:2024:CDA

[ADR24] Mine Akbas, Amanda E.

Diegel, and Leo G. Rebholz. Continuous data assimilation of a discretized barotropic vorticity model of geophysical flow. Computers and Mathematics with Applications, 160(??):30– 45, April 15, 2024. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122124000439.

Azhdari:2020:DRS

[AEF20] E. Azhdari, A. Emami, and J. A. Ferreira. Drug release from a surface erosion biodegradable viscoelastic polymeric platform: Analysis and numerical simulation. Computers and Mathematics with Applications, 80(12):3004–3026, December 15, 2020. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122120304090.

Azhdari:2023:DDA

[AEF23] E. Azhdari, A. Emami, and J. A. Ferreira. Drug delivery to the anterior segment of the eye enhanced by ultrasound

— modeling and simulation. Computers and Mathematics with Applications, 144 (??):51–61, August 15, 2023. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122123002092.

Amaziane:2024:CTF

[AEZ24] Brahim Amaziane, Mustapha El Ossmani, and Youssef Zahraoui. Convergence of a TPFA finite volume scheme for nonisothermal immiscible compressible two-phase flow in porous media. Computers and Mathematics with Applications, 165(??):118– 149, July 1, 2024. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S089812212400169X.

Asghar:2025:NWS

[AFBM25] Muhammad Hassan Asghar, Mathis Fricke, Dieter Bothe, and Tomislav Marić. Numerical wetting simulations using the plicRDF-isoAdvector unstructured Volume-of-Fluid (VOF) method. Computers and Mathematics with Applications, 180(??):191–213, February 15, 2025. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122124005595.

Angleitner:2023:EMM

[AFM23] Niklas Angleitner, Markus Faustmann, and Jens Markus Melenk. Exponential meshes and H-matrices. Computers and Mathematics with Applications, 130(??):21–40, January 15, 2023. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122122004746.

Antonietti:2024:APG

[AFM^{+}24] P. F. Antonietti, N. Farenga, E. Manuzzi, G. Martinelli, and L. Saverio. Agglomeration of polygonal grids using graph neural networks with applications to multigrid solvers. Computers and Mathematics with Applications, 154(??):45–57, January 15, 2024. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122123005059.

Aghchi:2024:NAS

[AFS24] Sima Aghchi, Hossein Fazli, and HongGunag Sun. A numerical approach for solving optimal control problem of fractional order vibration equation of large membranes. Computers and Mathematics with Applications, 165 (??):19–27, July 1, 2024. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122124001512.

Antonietti:2022:PDG

[AFV22] Paola F. Antonietti, Chiara Facciolà, and Marco Verani. Polytopic discontinuous Galerkin methods for the numerical modelling of flow in porous media with networks of intersecting fractures. Computers and Mathematics with Applications, 116 (??):116–139, June 15, 2022. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122121002996.

Aimi:2024:BBS

[AG24] A. Aimi and C. Guardasoni. BEM based semianalytical approach for accurate evaluation of arithmetic Asian barrier options. Computers and Mathematics with Applications, 167 (??):74–91, August 1, 2024. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122124002104.

Arun:2025:APS

[AG25] K. R. Arun and R. Ghorai. An asymptotic preserving scheme for the Euler– Poisson–Boltzmann system in the quasineutral limit. Computers and Mathematics with Applications, 185(??): 1–28, May 1, 2025. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S089812212500080X.

Alizadeh:2021:CHM

[AGA^{+}21] Rasool Alizadeh, Sina Rezaei Gomari, Ahmad Alizadeh, Nader Karimi, and Larry K. B. Li. Combined heat and mass transfer and thermodynamic irreversibilities in the stagnation-point flow of Casson rheological fluid over a cylinder with catalytic reactions and inside a porous medium under local thermal nonequilibrium.

Computers and Mathematics with Applications, 81(??): 786–810, January 1, 2021. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122119305103.

Aggul:2020:DCE

[Agg20] Mustafa Aggul. Defect correcting extrapolation technique for Oseen and Navier– Stokes flows. Computers and Mathematics with Applications, 80(5):1029–1049, September 1, 2020. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122120302431.

Acevedo:2021:FDF

[AGLR21] Ramiro Acevedo, Christian Gómez, and Bibiana López-Rodríguez. Fully discrete finite element approximation for a family of degenerate parabolic mixed equations. Computers and Mathematics with Applications, 96(??): 155–177, August 15, 2021. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122121001905.

Andrej:2022:PPB

[AGM^{+}22] Julian Andrej, Lars Grüne, Luca Mechelli, Thomas Meurer, Simon Pirkelmann, and Stefan Volkwein. Parallelized POD-based suboptimal economic model predictive control of a stateconstrained Boussinesq approximation. Computers and Mathematics with Applications, 116(??):71–81, June 15, 2022. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122121003242.

Arnst:2024:SIA

[AGNP24] Maarten Arnst, Christophe Geuzaine, Ludovic Noels, and Jean-Philippe Ponthot. Special issue: Advanced COmputational Methods in ENgineering (ACOMEN 2022). Computers and Mathematics with Applications, 176(??):189, December 15, 2024. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122124004553.

Arnold:2022:AFE

[AGPP22] Anton Arnold, Sjoerd Geevers, Ilaria Perugia, and Dmitry Ponomarev. An adaptive finite element method for high-frequency scattering problems with smoothly varying coefficients. Computers and Mathematics with Applications, 109(??):1–14, March 1, 2022. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122122000098.

Autio:2025:SMB

[AH25] Antti Autio and Antti Hannukainen. A subspace method based on the Neumann series for the solution of parametric linear systems. Computers and Mathematics with Applications, 178 (??):1–18, January 15, 2025. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122124005169.

Adler:2024:PFP

[AHLZ24] James H. Adler, Xiaozhe Hu, Yuwen Li, and Ludmil T. Zikatanov. Parameter-free preconditioning for nearlyincompressible linear elasticity. Computers and Mathematics with Applications, 154 (??):39–44, January 15, 2024. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122123005084.

Acebron:2020:HPA

[AHM20] Juan A. Acebrón, José R. Herrero, and José Monteiro. A highly parallel algorithm for computing the action of a matrix exponential on a vector based on a multilevel Monte Carlo method. Computers and Mathematics with Applications, 79(12): 3495–3515, June 15, 2020. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122120300808.

Ahmad:2024:OFF

[Ahm24] Shahbaz Ahmad. Optimized five-by-five block preconditioning for efficient GMRES convergence in curvaturebased image deblurring. Computers and Mathematics with Applications, 175(??): 174–183, December 1, 2024. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122124004358.

Almatroud:2024:ASR

[AHOB24] Othman Abdullah Almatroud, Amel Hioual, Adel Ouannas, and Iqbal M. Batiha. Asymptotic stability results of generalized discrete time reaction diffusion system applied to Lengyel–Epstein and Dagn Harrison models. Computers and Mathematics with Applications, 170(??):25–32, September 15, 2024. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122124002967.

Ahn:2020:DSM

[AHP20] Chi Young Ahn, Taeyoung Ha, and Won-Kwang Park. Direct sampling method for identifying magnetic inhomogeneities in limitedaperture inverse scattering problem. Computers and Mathematics with Applications, 80(12):2811–2829, December 15, 2020. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122120304089.

Abdel:2025:NAS

[AHZ^{+}25] Dilara Abdel, Maxime Herda, Martin Ziegler, Claire Chainais-Hillairet, Benjamin Spetzler, and Patricio Farrell. Numerical analysis and simulation of lateral memristive devices: Schottky, ohmic, and multi-dimensional electrode models. Computers and Mathematics with Applications, 199(??):286–308, December 1, 2025. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122125004158.

Achab:2025:NNA

[AI25] Louiza Achab and Farida Iachachene. A new numerical approach using the VOF method to model the twolayered Herschel–Bulkley blood flow in microvessels.

Computers and Mathematics with Applications, 190 (??):154–169, July 15, 2025. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122125001786.

Alsabery:2021:TNF

[AIG^{+}21] Ammar I. Alsabery, Muneer A.

Ismael, Engin Gedik, Ali J. Chamkha, and Ishak Hashim. Transient nanofluid flow and energy dissipation from wavy surface using magnetic field and two rotating cylinders. Computers and Mathematics with Applications, 97(??): 329–343, September 1, 2021. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122121002224.

Almushaira:2023:EEP

[AJ23] Mustafa Almushaira and Yan-Fei Jing. Efficient energy-preserving finite difference schemes for the Klein–Gordon–Schrödinger equations. Computers and Mathematics with Applications, 149(??):150–170, November 1, 2023. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122123003802.

Ankur:2023:ASK

[AJK23] Ankur, Ram Jiwari, and Naresh Kumar. Analysis and simulation of Korteweg–de Vries–Rosenauregularised long-wave model via Galerkin finite element method. Computers and Mathematics with Applications, 135(??):134– 148, April 1, 2023. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122123000330.

Ahmed:2024:ISS

[AJLM24] Naveed Ahmed, Volker John, Xu Li, and Christian Merdon. Inf–sup stabilized Scott–Vogelius pairs on general shape-regular simplicial grids for Navier–Stokes equations. Computers and Mathematics with Applications, 168(??):148–161, August 15, 2024. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122124002608.

Abid:2021:NMP

[AKA^{+}21] M. Abid, S. Koubaa, A. Abdelkefi, A. Frikha, and F. Dammak. Numerical modeling of porous functionally graded shells response in large deflection. Computers and Mathematics with Applications, 104(??):

59–70, December 15, 2021. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122121003965.

Adeosun:2025:BSA

[AKK^{+}25] Adeshina Taofeeq Adeosun, Abdul Rahman Mohd Kasim, Ilyas Khan, Yusuf Olatunji Tijani, and Mojeed Taiwo Akolade. Bifurcation and sensitivity analysis of a convectively heated squeezed MHD hybrid nanofluid with nonlinear radiation: a spectral approach. Computers and Mathematics with Applications, 195(??):28–41, October 1, 2025. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122125002792.

Achouri:2021:AFD

[AKO21] Talha Achouri, Tlili Kadri, and Khaled Omrani. Analysis of finite difference schemes for a fourth-order strongly damped nonlinear wave equations. Computers and Mathematics with Applications, 82(??):74–96, January 15, 2021. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122120304405.

afKlinteberg:2022:QEE

[aKST22] Ludvig af Klinteberg, Chiara Sorgentone, and Anna-Karin Tornberg. Quadrature error estimates for layer potentials evaluated near curved surfaces in three dimensions. Computers and Mathematics with Applications, 111 (??):1–19, April 1, 2022. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122122000517.

Aggul:2022:ADM

[AL22] Mustafa Aggul and Alexander E. Labovsky. Approximate deconvolution models for a fluid-fluid interaction problem with high Reynolds numbers. Computers and Mathematics with Applications, 117(??):113– 126, July 1, 2022. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122122001675.

Alosaimi:2023:DSD

[AL23] M. Alosaimi and D. Lesnic. Determination of a spacedependent source in the thermal-wave model of bioheat transfer. Computers and Mathematics with Applications, 129(??):34–49, January 1, 2023. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122122004564.

Alebrahim:2023:MWD

[Ale23]  Reza Alebrahim.  Modi-

fied wave dispersion properties in 1D and 2D statebased peridynamic media. Computers and Mathematics with Applications, 151(??): 21–35, December 1, 2023. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122123003929.

Alfat:2024:ECP

[Alf24] Sayahdin Alfat. On energyconsistency principle of PFM for thermal fracturing in thermoviscoelasticity solids and its application for modeling thermal response due to crack growth based on adaptive mesh technique. Computers and Mathematics with Applications, 175(??): 107–118, December 1, 2024. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122124004255.

Amigo:2025:VEM

[ALOR25] Danilo Amigo, Felipe Lepe, Enrique Otárola, and Gonzalo Rivera. A virtual element method for a convective Brinkman–Forchheimer problem coupled with a heat equation. Computers and Mathematics with Applications, 191(??):1–23, August 1, 2025. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122125001609.

Altmann:2023:SSS

[Alt23]  R. Altmann.  Splitting

schemes for the semi-linear wave equation with dynamic boundary conditions. Computers and Mathematics with Applications, 151(??):12–20, December 1, 2023. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122123004042.

Abaali:2020:MMB

[AM20] Mostafa Abaali and Zoubida Mghazli. Mathematical modelling of biodegradation in situ application to biodenitrification. Computers and Mathematics with Applications, 79(6):1833–1844, March 15, 2020. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122119304936.

Arefian:2022:CRB

[AM22] Sara Arefian and Davoud Mirzaei. A compact radial basis function partition of unity method. Computers and Mathematics with Applications, 127(??):1–11, December 1, 2022. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122122004151.

Aceto:2023:RPM

[AM23] L. Aceto and M. Mazza. A rational preconditioner for multi-dimensional Riesz fractional diffusion equations. Computers and Mathematics with Applications, 143(??): 372–382, August 1, 2023. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122123002201.

Aswin:2023:STA

[AMC23] V. S. Aswin, J. Manimaran, and Nagaiah Chamakuri. Space-time adaptivity for a multi-scale cancer invasion model. Computers and Mathematics with Applications, 146(??):309–322, September 15, 2023. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122123002997.

Al-Maskari:2025:SAT

[AMK25] Mariam Al-Maskari and Samir Karaa. Strong approximation of the timefractional Cahn–Hilliard equation driven by a fractionally integrated additive noise. Computers and Mathematics with Applications, 180(??): 28–45, February 15, 2025. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122124005522.

Antonietti:2023:DGT

[AMM23] Paola F. Antonietti, Ilario Mazzieri, and Francesco Migliorini. A discontinuous Galerkin time integration scheme for second order differential equations with applications to seismic wave propagation problems. Computers and Mathematics with Applications, 134 (??):87–100, March 15, 2023. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122123000159.

Adak:2022:VEA

[AMN22] D. Adak, G. Manzini, and S. Natarajan. Virtual element approximation of twodimensional parabolic variational inequalities. Computers and Mathematics with Applications, 116(??):48– 70, June 15, 2022. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122121003345.

Actis:2020:NPH

[AMP20] Marcelo Actis, Pedro Morin, and M. Sebastian Pauletti. A new perspective on hierarchical spline spaces for adaptivity. Computers and Mathematics with Applications, 79(8):2276–2303, April 15, 2020. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122119305231.

Al-Mdallal:2022:KBS

[AMPB^{+}22] Qasem Al-Mdallal, V. Ramachandra Prasad, H. Thameem Basha, Ioannis Sarris, and Nevzat Akkurt. Keller box simulation of magnetic pseudoplastic nano-polymer coating flow over a circular cylinder with entropy optimisation. Computers and Mathematics with Applications, 118 (??):132–158, July 15, 2022. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122122002000.

Alla:2024:PDO

[AMS24] Alessandro Alla, Angela Monti, and Ivonne Sgura. Piecewise DMD for oscillatory and Turing spatiotemporal dynamics. Computers and Mathematics with Applications, 160(??):108– 124, April 15, 2024. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122124000671.

Antonietti:2020:CVE

[AMV20] P. F. Antonietti, G. Manzini, and M. Verani. The conforming virtual element method for polyharmonic problems. Computers and Mathematics with Applications, 79 (7):2021–2034, April 2020. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S089812211930478X.

Alves:2021:DDM

[AMV21] Carlos J. S. Alves, Nuno F. M. Martins, and Svilen S. Valtchev. Domain decomposition methods with fundamental solutions for Helmholtz problems with discontinuous source terms. Computers and Mathematics with Applications, 88 (??):16–32, April 15, 2021. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122118307065.

Adak:2023:NVE

[AMV23] Dibyendu Adak, David Mora, and Iván Velásquez. Nonconforming virtual element discretization for the transmission eigenvalue problem. Computers and Mathematics with Applications, 152(??):250–267, December 15, 2023. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122123004790.

Antonietti:2024:ISB

[AMV24] P. F. Antonietti, P. Matalon, and M. Verani. Iterative solution to the biharmonic equation in mixed form discretized by the Hybrid High-Order method. Computers and Mathematics with Applications, 171(??): 154–163, October 1, 2024. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S089812212400316X.

Adak:2025:NVE

[AMV25] Dibyendu Adak, Gianmarco Manzini, and Jesus Vellojin. Nonconforming virtual element method for the Schrödinger eigenvalue problem. Computers and Mathematics with Applications, 182(??):213–235, March 15, 2025. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122125000471.

Adak:2020:VEM

[AN20] D. Adak and S. Natarajan. Virtual element method for a nonlocal elliptic problem of Kirchhoff type on polygonal meshes. Computers and Mathematics with Applications, 79(10):2856– 2871, May 15, 2020. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122119305863.

Assam:2021:NLS

[AN21] Ashwani Assam and Ganesh Natarajan. A novel least squares finite volume scheme for discontinuous diffusion on unstructured meshes. Computers and Mathematics with Applications, 96(??): 120–130, August 15, 2021. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122121001930.

Arrutselvi:2023:VES

[AN23] M. Arrutselvi and E. Natarajan. Virtual element stabilization for the system of time-dependent nonlinear convection–diffusion– reaction equations. Computers and Mathematics with Applications, 142(??):121– 139, July 15, 2023. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122123001554.

Ali:2021:IDF

[ANH^{+}21] Bagh Ali, Yufeng Nie, Sajjad Hussain, Danial Habib, and Sohaib Abdal. Insight into the dynamics of fluid conveying tiny particles over a rotating surface subject to Cattaneo–Christov heat transfer, Coriolis force, and Arrhenius activation energy. Computers and Mathematics with Applications, 93 (??):130–143, July 1, 2021. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122121001383.

Arrutselvi:2025:NSN

[ANN25] M. Arrutselvi, E. Natarajan, and S. Natarajan. Numerical solution of nonlinear convection-diffusion-reaction equation using a stabilized virtual element method. Computers and Mathematics with Applications, 183 (??):46–70, April 1, 2025. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122125000409.

Anonymous:2020:EBa

[Ano20a]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 79(1):i, January 2020. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122119305760.

Anonymous:2020:EBb

[Ano20b]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 79(1):ii, January 2020. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122119305772.

Anonymous:2020:EBc

[Ano20c]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 79(2):i, January 15, 2020. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0898122119305917.

Anonymous:2020:EBd

[Ano20d]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 79(2):ii, January 15, 2020. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0898122119305929.

Anonymous:2020:EBe

[Ano20e]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 79(3):i, February 1, 2020. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122120300055.

Anonymous:2020:EBf

[Ano20f]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 79(3):ii, February 1, 2020. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122120300067.

Anonymous:2020:EBg

[Ano20g]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 79(4):i, February 15, 2020. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122120300171.

Anonymous:2020:EBh

[Ano20h]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 79(4):ii, February 15, 2020. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122120300183.

Anonymous:2020:EBi

[Ano20i]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 79(5):i, March 1, 2020. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122120300432.

Anonymous:2020:EBj

[Ano20j]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 79(5):ii, March 1, 2020. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122120300444.

Anonymous:2020:EBk

[Ano20k]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 79(6):i, March 15, 2020. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0898122120300584.

Anonymous:2020:EBl

[Ano20l]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 79(6):ii, March 15, 2020. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122120300596.

Anonymous:2020:EBm

[Ano20m]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 79(7):i, April 2020. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122120300729.

Anonymous:2020:EBn

[Ano20n]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 79(7):ii, April 2020. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122120300730.

Anonymous:2020:EBo

[Ano20o]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 79(8):i, April 15, 2020. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122120300948.

Anonymous:2020:EBp

[Ano20p]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 79(8):ii, April 15, 2020. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S089812212030095X.

Anonymous:2020:EBq

[Ano20q]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 79(9):i, May 1, 2020. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122120301103.

Anonymous:2020:EBr

[Ano20r]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 79(9):ii, May 1, 2020. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122120301115.

Anonymous:2020:EBs

[Ano20s]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 79(10):i, May 15, 2020. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122120301346.

Anonymous:2020:EBt

[Ano20t]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 79(10):ii, May 15, 2020. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122120301358.

Anonymous:2020:EBu

[Ano20u]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 79(11):i, June 1, 2020. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122120301504.

Anonymous:2020:EBv

[Ano20v]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 79(11):ii, June 1, 2020. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122120301516.

Anonymous:2020:EBw

[Ano20w]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 79(12):i, June 15, 2020. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0898122120301796.

Anonymous:2020:EBx

[Ano20x]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 79(12):ii, June 15, 2020. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0898122120301802.

Anonymous:2020:EBy

[Ano20y]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 80(1):i, July 1, 2020. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122120301954.

Anonymous:2020:EBz

[Ano20z]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 80(1):ii, July 1, 2020. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122120301966.

Anonymous:2020:EBba

[Ano20-27]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 80(2):i, July 2020.

CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S089812212030211X.

Anonymous:2020:EBbb

[Ano20-28]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 80(2):ii, July 2020. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122120302121.

Anonymous:2020:EBbc

[Ano20-29]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 80(3):i, August 1, 2020. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122120302273.

Anonymous:2020:EBbd

[Ano20-30]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 80(3):ii, August 1, 2020. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122120302285.

Anonymous:2020:EBbe

[Ano20-31]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 80(4):i, August 15, 2020. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122120302376.

Anonymous:2020:EBbf

[Ano20-32]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 80(4):ii, August 15, 2020. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122120302388.

Anonymous:2020:EBbg

[Ano20-33]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 80(5):i, September 1, 2020. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122120302984.

Anonymous:2020:EBbh

[Ano20-34]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 80(5):ii, September 1, 2020. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122120302996.

Anonymous:2020:EBbi

[Ano20-35]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 80(6):i, September 15, 2020. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122120303357.

Anonymous:2020:EBbj

[Ano20-36]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 80(6):ii, September 15, 2020. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122120303369.

Anonymous:2020:EBbk

[Ano20-37]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 80(7):i, October 1, 2020. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S089812212030345X.

Anonymous:2020:EBbl

[Ano20-38]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 80(7):ii, October 1, 2020. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122120303461.

Anonymous:2020:EBbm

[Ano20-39]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 80(8):i, October 15, 2020. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122120303576.

Anonymous:2020:EBbn

[Ano20-40]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 80(8):ii, October 15, 2020. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122120303588.

Anonymous:2020:EBbo

[Ano20-41]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 80(9):i, November 1, 2020. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122120303746.

Anonymous:2020:EBbp

[Ano20-42]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 80(9):ii, November 1, 2020. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122120303758.

Anonymous:2020:EBbq

[Ano20-43]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 80(10):i, November 15, 2020. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122120303928.

Anonymous:2020:EBbr

[Ano20-44]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 80(10):ii, November 15, 2020. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S089812212030393X.

Anonymous:2020:EBbs

[Ano20-45]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 80(11):i, December 2020. CODEN CMAPDK.

ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S089812212030403X.

Anonymous:2020:EBbt

[Ano20-46]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 80(11):ii, December 2020. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122120304041.

Anonymous:2020:EBbu

[Ano20-47]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 80(12):i, December 15, 2020. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S089812212030448X.

Anonymous:2020:EBbv

[Ano20-48]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 80(12):ii, December 15, 2020. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122120304491.

Anonymous:2020:PJd

[Ano20-49] Anonymous. Pages 1–284 (1 July 2020). Computers and Mathematics with Applications, 80(1):??, July 1, 2020. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic).

Anonymous:2020:PAAa

[Ano20-50] Anonymous. Pages 1–92, A1–A2 (15 August 2020). Computers and Mathematics with Applications, 80(4): ??, August 15, 2020. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic).

Anonymous:2020:PMa

[Ano20-51] Anonymous. Pages 1241– 1618 (1 March 2020). Computers and Mathematics with Applications, 79(5):??, March 1, 2020. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic).

Anonymous:2020:PSb

[Ano20-52] Anonymous. Pages 1479– 1744 (15 September 2020). Computers and Mathematics with Applications, 80(6):??, September 15, 2020. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic).

Anonymous:2020:PMb

[Ano20-53] Anonymous. Pages 1619– 1888 (15 March 2020).

Computers and Mathematics with Applications, 79(6): ??, March 15, 2020. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic).

Anonymous:2020:POa

[Ano20-54] Anonymous. Pages 1745– 1872 (1 October 2020). Computers and Mathematics with Applications, 80(7):??, October 1, 2020. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic).

Anonymous:2020:POb

[Ano20-55] Anonymous. Pages 1873– 2018 (15 October 2020). Computers and Mathematics with Applications, 80(8): ??, October 15, 2020. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic).

Anonymous:2020:PNb

[Ano20-56] Anonymous. Pages 2019– 2104 (1 November 2020). Computers and Mathematics with Applications, 80(9):??, November 1, 2020. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic).

Anonymous:2020:PNc

[Ano20-57] Anonymous. Pages 2105– 2258 (15 November 2020). Computers and Mathematics with Applications, 80(10):??, November 15, 2020. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic).

Anonymous:2020:PAa

[Ano20-58] Anonymous. Pages 2119– 2472 (15 April 2020). Computers and Mathematics with Applications, 79(8):??, April 15, 2020. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic).

Anonymous:2020:PJa

[Ano20-59] Anonymous. Pages 213–538 (15 January 2020). Computers and Mathematics with Applications, 79(2):??, January 15, 2020. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic).

Anonymous:2020:PMc

[Ano20-60] Anonymous. Pages 2473– 2794 (1 May 2020). Computers and Mathematics with Applications, 79(9):??, May 1, 2020. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic).

Anonymous:2020:PD

[Ano20-61] Anonymous. Pages 2623– 3194 (15 December 2020). Computers and Mathematics with Applications, 80(12):??, December 15, 2020. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic).

Anonymous:2020:PMd

[Ano20-62] Anonymous. Pages 2795– 3026 (15 May 2020). Computers and Mathematics with Applications, 79(10): ??, May 15, 2020. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic).

Anonymous:2020:PJb

[Ano20-63] Anonymous. Pages 3027– 3256 (1 June 2020). Computers and Mathematics with Applications, 79(11):??, June 1, 2020. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic).

Anonymous:2020:PJc

[Ano20-64] Anonymous. Pages 3257– 3612 (15 June 2020). Computers and Mathematics with Applications, 79(12): ??, June 15, 2020. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic).

Anonymous:2020:PAb

[Ano20-65] Anonymous. Pages 417–568 (1 August 2020). Computers and Mathematics with Applications, 80(3):??, August 1, 2020. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic).

Anonymous:2020:PFa

[Ano20-66] Anonymous. Pages 539– 888 (1 February 2020).

Computers and Mathematics with Applications, 79(3): ??, February 1, 2020. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic).

Anonymous:2020:PSa

[Ano20-67] Anonymous. Pages 569– 1478 (1 September 2020). Computers and Mathematics with Applications, 80(5):??, September 1, 2020. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic).

Anonymous:2020:PFb

[Ano20-68] Anonymous. Pages 889– 1240 (15 February 2020). Computers and Mathematics with Applications, 79(4):??, February 15, 2020. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic).

Anonymous:2020:PNa

[Ano20-69]  Anonymous.  Publisher’s

note. Computers and Mathematics with Applications, 80(4):A1, August 15, 2020. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122120302418.

Anonymous:2021:EBa

[Ano21a]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 81(??):i, January 1, 2021. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122120304582.

Anonymous:2021:EBb

[Ano21b]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 81(??):ii, January 1, 2021. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122120304594.

Anonymous:2021:EBc

[Ano21c]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 82(??):i, January 15, 2021. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122120304764.

Anonymous:2021:EBd

[Ano21d]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 82(??):ii, January 15, 2021. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122120304776.

Anonymous:2021:EBe

[Ano21e]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 83(??):i, February 1, 2021. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122121000110.

Anonymous:2021:EBf

[Ano21f]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 83(??):ii, February 1, 2021. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122121000122.

Anonymous:2021:EBg

[Ano21g]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 84(??):i, February 15, 2021. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122121000304.

Anonymous:2021:EBh

[Ano21h]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 84(??):ii, February 15, 2021. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122121000316.

Anonymous:2021:EBi

[Ano21i]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 85(??):i, March 1, 2021. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122121000432.

Anonymous:2021:EBj

[Ano21j]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 85(??):ii, March 1, 2021. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122121000444.

Anonymous:2021:EBk

[Ano21k]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 86(??):i, March 15, 2021. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122121000584.

Anonymous:2021:EBl

[Ano21l]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 86(??):ii, March 15, 2021. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122121000596.

Anonymous:2021:EBm

[Ano21m]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 87(??):i, April 1, 2021. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S089812212100081X.

Anonymous:2021:EBn

[Ano21n]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 87(??):ii, April 1, 2021. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122121000821.

Anonymous:2021:EBo

[Ano21o]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 88(??):i, April 15, 2021. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0898122121001061.

Anonymous:2021:EBp

[Ano21p]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 88(??):ii, April 15, 2021. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0898122121001073.

Anonymous:2021:EBq

[Ano21q]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 89(??):i, May 1, 2021. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122121001188.

Anonymous:2021:EBr

[Ano21r]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 89(??):ii, May 1, 2021. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S089812212100119X.

Anonymous:2021:EBs

[Ano21s]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 90(??):i, May 15, 2021.

CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122121001322.

Anonymous:2021:EBt

[Ano21t]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 90(??):ii, May 15, 2021. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122121001334.

Anonymous:2021:EBu

[Ano21u]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 91(??):i, June 1, 2021. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122121001474.

Anonymous:2021:EBv

[Ano21v]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 91(??):ii, June 1, 2021. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122121001486.

Anonymous:2021:EBw

[Ano21w]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 92(??):i, June 15, 2021. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122121001668.

Anonymous:2021:EBx

[Ano21x]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 92(??):ii, June 15, 2021. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S089812212100167X.

Anonymous:2021:EBy

[Ano21y]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 93(??):i, July 1, 2021. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122121001796.

Anonymous:2021:EBz

[Ano21z]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 93(??):ii, July 1, 2021. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122121001802.

Anonymous:2021:EBba

[Ano21-27]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 94(??):i, July 15, 2021. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122121002042.

Anonymous:2021:EBbb

[Ano21-28]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 94(??):ii, July 15, 2021. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122121002054.

Anonymous:2021:EBbc

[Ano21-29]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 95(??):i, August 1, 2021. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122121002182.

Anonymous:2021:EBbd

[Ano21-30]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 95(??):ii, August 1, 2021. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0898122121002194.

Anonymous:2021:EBbe

[Ano21-31]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 96(??):i, August 15, 2021. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0898122121002273.

Anonymous:2021:EBbf

[Ano21-32]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 96(??):ii, August 15, 2021. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0898122121002285.

Anonymous:2021:EBbg

[Ano21-33]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 97(??):i, September 1, 2021. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122121002637.

Anonymous:2021:EBbh

[Ano21-34]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 97(??):ii, September 1, 2021. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122121002649.

Anonymous:2021:EBbi

[Ano21-35]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 98(??):i, September 15, 2021. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122121003023.

Anonymous:2021:EBbj

[Ano21-36]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 98(??):ii, September 15, 2021. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122121003035.

Anonymous:2021:EBbk

[Ano21-37]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 99(??):i, October 1, 2021. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122121003291.

Anonymous:2021:EBbl

[Ano21-38]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 99(??):ii, October 1, 2021. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122121003308.

Anonymous:2021:EBbm

[Ano21-39]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 100(??):i, October 15, 2021. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122121003436.

Anonymous:2021:EBbn

[Ano21-40]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 100(??):ii, October 15, 2021. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122121003448.

Anonymous:2021:EBbo

[Ano21-41]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 101(??):i, November 1, 2021. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122121003667.

Anonymous:2021:EBbp

[Ano21-42]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 101(??):ii, November 1, 2021. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122121003679.

Anonymous:2021:EBbq

[Ano21-43]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 102(??):i, November 15, 2021. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122121003874.

Anonymous:2021:EBbr

[Ano21-44]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 102(??):ii, November 15, 2021. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122121003886.

Anonymous:2021:EBbs

[Ano21-45]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 103(??):i, December 1, 2021. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122121004004.

Anonymous:2021:EBbt

[Ano21-46]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 103(??):ii, December 1, 2021. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122121004016.

Anonymous:2021:EBbu

[Ano21-47]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 104(??):i, December 15, 2021. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S089812212100420X.

Anonymous:2021:EBbv

[Ano21-48]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 104(??):ii, December 15, 2021. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122121004211.

Anonymous:2021:PMb

[Ano21-49] Anonymous. Pages 1–106 (15 March 2021). Computers and Mathematics with Applications, 86(??):1–106, March 15, 2021. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic).

Anonymous:2021:PDb

[Ano21-50] Anonymous. Pages 1–112 (15 December 2021). Computers and Mathematics with Applications, 104(??):1–112, December 15, 2021. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic).

Anonymous:2021:PNa

[Ano21-51] Anonymous. Pages 1–134 (1 November 2021). Computers and Mathematics with Applications, 101(??):1–134, November 1, 2021. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic).

Anonymous:2021:PAa

[Ano21-52] Anonymous. Pages 1–138 (1 April 2021). Computers and Mathematics with Applications, 87(??):1–138, April 1, 2021. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic).

Anonymous:2021:PDa

[Ano21-53] Anonymous. Pages 1–170 (1 December 2021). Computers and Mathematics with Applications, 103(??):1–170, December 1, 2021. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic).

Anonymous:2021:PJd

[Ano21-54] Anonymous. Pages 1–170 (15 July 2021). Computers and Mathematics with Applications, 94(??):1–170, July 15, 2021. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic).

Anonymous:2021:PMd

[Ano21-55] Anonymous. Pages 1–170 (15 May 2021). Computers and Mathematics with Applications, 90(??):1–170, May 15, 2021. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic).

Anonymous:2021:PJb

[Ano21-56] Anonymous. Pages 1–180 (15 June 2021). Computers and Mathematics with Applications, 92(??):1–180, June 15, 2021. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic).

Anonymous:2021:PMc

[Ano21-57] Anonymous. Pages 1–184 (1 May 2021). Computers and Mathematics with Applications, 89(??):1–184, May 1, 2021. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic).

Anonymous:2021:POb

[Ano21-58] Anonymous. Pages 1–196 (15 October 2021). Computers and Mathematics with Applications, 100(??):1–196, October 15, 2021. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic).

Anonymous:2021:PAb

[Ano21-59] Anonymous. Pages 1–228 (15 August 2021). Computers and Mathematics with Applications, 96(??):1–228, August 15, 2021. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic).

Anonymous:2021:PJa

[Ano21-60] Anonymous. Pages 1–228 (15 January 2021). Computers and Mathematics with Applications, 82(??):1–228, January 15, 2021. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic).

Anonymous:2021:PNb

[Ano21-61] Anonymous. Pages 1–276 (15 November 2021). Computers and Mathematics with Applications, 102(??):1–276, November 15, 2021. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic).

Anonymous:2021:PSb

[Ano21-62] Anonymous. Pages 1–280 (15 September 2021). Computers and Mathematics with Applications, 98(??):1–280, September 15, 2021. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic).

Anonymous:2021:PJc

[Ano21-63] Anonymous. Pages 1–288 (1 July 2021). Computers and Mathematics with Applications, 93(??):1–288, July 1, 2021. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic).

Anonymous:2021:PF

[Ano21-64] Anonymous. Pages 1–314 (15 February 2021). Computers and Mathematics with Applications, 84(??):1–314, February 15, 2021. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic).

Anonymous:2021:POa

[Ano21-65] Anonymous. Pages 1–360 (1 October 2021). Computers and Mathematics with Applications, 99(??):1–360, October 1, 2021. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic).

Anonymous:2021:PSa

[Ano21-66] Anonymous. Pages 1–438 (1 September 2021). Computers and Mathematics with Applications, 97(??):1–438, September 1, 2021. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic).

Anonymous:2021:PMa

[Ano21-67] Anonymous. Pages 1–98 (1 March 2021). Computers and Mathematics with Applications, 85(??):1–98, March 1, 2021. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic).

Anonymous:2021:P

[Ano21-68] Anonymous. Preface. Computers and Mathematics with Applications, 88(??): 1, April 15, 2021. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122121001036.

Anonymous:2022:EBa

[Ano22a]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 105(??):i, January 1, 2022. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122121004351.

Anonymous:2022:EBb

[Ano22b]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 105(??):ii, January 1, 2022. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122121004363.

Anonymous:2022:EBc

[Ano22c]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 106(??):i, January 15, 2022. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S089812212100451X.

Anonymous:2022:EBd

[Ano22d]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 106(??):ii, January 15, 2022. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122121004521.

Anonymous:2022:EBe

[Ano22e]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 107(??):i, February 1, 2022. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122122000189.

Anonymous:2022:EBf

[Ano22f]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 107(??):ii, February 1, 2022. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122122000190.

Anonymous:2022:EBg

[Ano22g]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 108(??):i, February 15, 2022. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122122000414.

Anonymous:2022:EBh

[Ano22h]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 108(??):ii, February 15, 2022. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122122000426.

Anonymous:2022:EBi

[Ano22i]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 109(??):i, March 1, 2022. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122122000578.

Anonymous:2022:EBj

[Ano22j]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 109(??):ii, March 1, 2022. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S089812212200058X.

Anonymous:2022:EBk

[Ano22k]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 110(??):i, March 15, 2022. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122122000724.

Anonymous:2022:EBl

[Ano22l]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 110(??):ii, March 15, 2022. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122122000736.

Anonymous:2022:EBm

[Ano22m]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 111(??):i, April 1, 2022. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122122001080.

Anonymous:2022:EBn

[Ano22n]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 111(??):ii, April 1, 2022. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122122001092.

Anonymous:2022:EBo

[Ano22o]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 112(??):i, April 15, 2022. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0898122122001316.

Anonymous:2022:EBp

[Ano22p]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 112(??):ii, April 15, 2022. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0898122122001328.

Anonymous:2022:EBq

[Ano22q]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 113(??):i, May 1, 2022. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122122001456.

Anonymous:2022:EBr

[Ano22r]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 113(??):ii, May 1, 2022. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0898122122001468.

Anonymous:2022:EBs

[Ano22s]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 114(??):i, May 15, 2022. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122122001596.

Anonymous:2022:EBt

[Ano22t]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 114(??):ii, May 15, 2022. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122122001602.

Anonymous:2022:EBu

[Ano22u]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 115(??):i, June 1, 2022. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122122001857.

Anonymous:2022:EBv

[Ano22v]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 115(??):ii, June 1, 2022. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122122001869.

Anonymous:2022:EBw

[Ano22w]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 116(??):i, June 15, 2022. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S089812212200236X.

Anonymous:2022:EBx

[Ano22x]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 116(??):ii, June 15, 2022. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0898122122002371.

Anonymous:2022:EBy

[Ano22y]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 117(??):i, July 1, 2022. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122122002073.

Anonymous:2022:EBz

[Ano22z]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 117(??):ii, July 1, 2022. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122122002085.

Anonymous:2022:EBba

[Ano22-27]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 118(??):i, July 15, 2022. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122122002449.

Anonymous:2022:EBbb

[Ano22-28]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 118(??):ii, July 15, 2022. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122122002450.

Anonymous:2022:EBbc

[Ano22-29]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 119(??):i, August 1, 2022. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122122002802.

Anonymous:2022:EBbd

[Ano22-30]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 119(??):ii, August 1, 2022. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122122002814.

Anonymous:2022:EBbe

[Ano22-31]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 120(??):i, August 15, 2022. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122122003005.

Anonymous:2022:EBbf

[Ano22-32]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 120(??):ii, August 15, 2022. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122122003017.

Anonymous:2022:EBbg

[Ano22-33]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 121(??):i, September 1, 2022. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122122003182.

Anonymous:2022:EBbh

[Ano22-34]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 121(??):ii, September 1, 2022. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122122003194.

Anonymous:2022:EBbi

[Ano22-35]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 122(??):i, September 15, 2022. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122122003571.

Anonymous:2022:EBbj

[Ano22-36]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 122(??):ii, September 15, 2022. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122122003583.

Anonymous:2022:EBbk

[Ano22-37]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 123(??):i, October 1, 2022. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122122003832.

Anonymous:2022:EBbl

[Ano22-38]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 123(??):ii, October 1, 2022. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122122003844.

Anonymous:2022:EBbm

[Ano22-39]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 124(??):i, October 15, 2022. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122122004096.

Anonymous:2022:EBbn

[Ano22-40]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 124(??):ii, October 15, 2022. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122122004102.

Anonymous:2022:EBbo

[Ano22-41]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 125(??):i, November 1, 2022. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122122004217.

Anonymous:2022:EBbp

[Ano22-42]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 125(??):ii, November 1, 2022. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122122004229.

Anonymous:2022:EBbq

[Ano22-43]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 126(??):i, November 15, 2022. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S089812212200445X.

Anonymous:2022:EBbr

[Ano22-44]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 126(??):ii, November 15, 2022. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122122004461.

Anonymous:2022:EBbs

[Ano22-45]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 127(??):i, December 1, 2022. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122122004667.

Anonymous:2022:EBbt

[Ano22-46]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 127(??):ii, December 1, 2022. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122122004679.

Anonymous:2022:EBbu

[Ano22-47]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 128(??):i, December 15, 2022. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122122004850.

Anonymous:2022:EBbv

[Ano22-48]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 128(??):ii, December 15, 2022. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122122004862.

Anonymous:2022:PJb

[Ano22-49] Anonymous. Pages 1–106 (15 January 2022). Computers and Mathematics with Applications, 106(??):1–106, January 15, 2022. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic).

Anonymous:2022:PSb

[Ano22-50] Anonymous. Pages 1–126 (15 September 2022). Computers and Mathematics with Applications, 122(??):1–126, September 15, 2022. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic).

Anonymous:2022:PFa

[Ano22-51] Anonymous. Pages 1–132 (1 February 2022). Computers and Mathematics with Applications, 107(??):1–132, February 1, 2022. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic).

Anonymous:2022:PAa

[Ano22-52] Anonymous. Pages 1–134 (1 April 2022). Computers and Mathematics with Applications, 111(??):1–134, April 1, 2022. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic).

Anonymous:2022:PSa

[Ano22-53] Anonymous. Pages 1–144 (1 September 2022). Computers and Mathematics with Applications, 121(??):1–144, September 1, 2022. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic).

Anonymous:2022:PMb

[Ano22-54] Anonymous. Pages 1–146 (15 March 2022). Computers and Mathematics with Applications, 110(??):1–146, March 15, 2022. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic).

Anonymous:2022:PAd

[Ano22-55] Anonymous. Pages 1–150 (15 August 2022). Computers and Mathematics with Applications, 120(??):1–150, August 15, 2022. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic).

Anonymous:2022:PJa

[Ano22-56] Anonymous. Pages 1–172 (1 January 2022). Computers and Mathematics with Applications, 105(??):1–172, January 1, 2022. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic).

Anonymous:2022:PAb

[Ano22-57] Anonymous. Pages 1–180 (15 April 2022). Computers and Mathematics with Applications, 112(??):1–180, April 15, 2022. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic).

Anonymous:2022:PFb

[Ano22-58] Anonymous. Pages 1–216 (15 February 2022). Computers and Mathematics with Applications, 108(??):1–216, February 15, 2022. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic).

Anonymous:2022:PDa

[Ano22-59] Anonymous. Pages 1–224 (1 December 2022). Computers and Mathematics with Applications, 127(??):1–224, December 1, 2022. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic).

Anonymous:2022:PNb

[Ano22-60] Anonymous. Pages 1–226 (15 November 2022). Computers and Mathematics with Applications, 126(??):1–226, November 15, 2022. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic).

Anonymous:2022:PMd

[Ano22-61] Anonymous. Pages 1–244 (15 May 2022). Computers and Mathematics with Applications, 114(??):1–244, May 15, 2022. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic).

Anonymous:2022:PNa

[Ano22-62] Anonymous. Pages 1–248 (1 November 2022). Computers and Mathematics with Applications, 125(??):1–248, November 1, 2022. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic).

Anonymous:2022:POa

[Ano22-63] Anonymous. Pages 1–252 (1 October 2022). Computers and Mathematics with Applications, 123(??):1–252, October 1, 2022. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic).

Anonymous:2022:POb

[Ano22-64] Anonymous. Pages 1–258 (15 October 2022). Computers and Mathematics with Applications, 124(??):1–258, October 15, 2022. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic).

Anonymous:2022:PJf

[Ano22-65] Anonymous. Pages 1–264 (15 July 2022). Computers and Mathematics with Applications, 118(??):1–264, July 15, 2022. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic).

Anonymous:2022:PMa

[Ano22-66] Anonymous. Pages 1–324 (1 March 2022). Computers and Mathematics with Applications, 109(??):1–324, March 1, 2022. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic).

Anonymous:2022:PJe

[Ano22-67] Anonymous. Pages 1–326 (1 July 2022). Computers and Mathematics with Applications, 117(??):1–326, July 1, 2022. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic).

Anonymous:2022:PMc

[Ano22-68] Anonymous. Pages 1–352 (1 May 2022). Computers and Mathematics with Applications, 113(??):1–352, May 1, 2022. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic).

Anonymous:2022:PDb

[Ano22-69] Anonymous. Pages 1–382 (15 December 2022). Computers and Mathematics with Applications, 128(??):1–382, December 15, 2022. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic).

Anonymous:2022:PAc

[Ano22-70] Anonymous. Pages 1–386 (1 August 2022). Computers and Mathematics with Applications, 119(??):1–386, August 1, 2022. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic).

Anonymous:2022:PJc

[Ano22-71] Anonymous. Pages 1–80 (1 June 2022). Computers and Mathematics with Applications, 115(??):1–80, June 1, 2022. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic).

Anonymous:2023:EBa

[Ano23a]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 129(??):i, January 1, 2023. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122122005156.

Anonymous:2023:EBb

[Ano23b]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 129(??):ii, January 1, 2023. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122122005168.

Anonymous:2023:EBc

[Ano23c]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 130(??):i, January 15, 2023. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122122005314.

Anonymous:2023:EBd

[Ano23d]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 130(??):ii, January 15, 2023. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122122005326.

Anonymous:2023:EBe

[Ano23e]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 131(??):i, February 1, 2023. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122123000226.

Anonymous:2023:EBf

[Ano23f]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 131(??):ii, February 1, 2023. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122123000238.

Anonymous:2023:EBg

[Ano23g]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 132(??):i, February 15, 2023. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122123000421.

Anonymous:2023:EBh

[Ano23h]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 132(??):ii, February 15, 2023. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122123000433.

Anonymous:2023:EBi

[Ano23i]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 133(??):i, March 1, 2023. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122123000561.

Anonymous:2023:EBj

[Ano23j]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 133(??):ii, March 1, 2023. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122123000573.

Anonymous:2023:EBk

[Ano23k]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 134(??):i, March 15, 2023. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122123000846.

Anonymous:2023:EBl

[Ano23l]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 134(??):ii, March 15, 2023. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0898122123000858.

Anonymous:2023:EBm

[Ano23m]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 135(??):i, April 1, 2023. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0898122123000974.

Anonymous:2023:EBn

[Ano23n]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 135(??):ii, April 1, 2023. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0898122123000986.

Anonymous:2023:EBo

[Ano23o]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 136(??):i, April 15, 2023. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0898122123001207.

Anonymous:2023:EBp

[Ano23p]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 136(??):ii, April 15, 2023. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122123001219.

Anonymous:2023:EBq

[Ano23q]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 137(??):i, May 1, 2023. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122123001347.

Anonymous:2023:EBr

[Ano23r]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 137(??):ii, May 1, 2023. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122123001359.

Anonymous:2023:EBs

[Ano23s]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 138(??):i, May 15, 2023. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122123001633.

Anonymous:2023:EBt

[Ano23t]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 138(??):ii, May 15, 2023. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0898122123001645.

Anonymous:2023:EBu

[Ano23u]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 139(??):i, June 1, 2023. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S089812212300189X.

Anonymous:2023:EBv

[Ano23v]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 139(??):ii, June 1, 2023. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0898122123001906.

Anonymous:2023:EBw

[Ano23w]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 140(??):i, June 15, 2023. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122123001967.

Anonymous:2023:EBx

[Ano23x]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 140(??):ii, June 15, 2023. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122123001979.

Anonymous:2023:EBy

[Ano23y]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 141(??):i, July 1, 2023. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S089812212300216X.

Anonymous:2023:EBz

[Ano23z]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 141(??):ii, July 1, 2023. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122123002171.

Anonymous:2023:EBba

[Ano23-27]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 142(??):i, July 15, 2023. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122123002407.

Anonymous:2023:EBbb

[Ano23-28]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 142(??):ii, July 15, 2023. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122123002419.

Anonymous:2023:EBbc

[Ano23-29]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 143(??):i, August 1, 2023. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122123002511.

Anonymous:2023:EBbd

[Ano23-30]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 143(??):ii, August 1, 2023. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122123002523.

Anonymous:2023:EBbe

[Ano23-31]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 144(??):i, August 15, 2023. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122123003243.

Anonymous:2023:EBbf

[Ano23-32]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 144(??):ii, August 15, 2023. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122123003255.

Anonymous:2023:EBbg

[Ano23-33]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 145(??):i, September 1, 2023. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122123003395.

Anonymous:2023:EBbh

[Ano23-34]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 145(??):ii, September 1, 2023. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122123003401.

Anonymous:2023:EBbi

[Ano23-35]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 146(??):i, September 15, 2023. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122123003644.

Anonymous:2023:EBbj

[Ano23-36]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 146(??):ii, September 15, 2023. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122123003656.

Anonymous:2023:EBbk

[Ano23-37]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 147(??):i, October 1, 2023. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122123003838.

Anonymous:2023:EBbl

[Ano23-38]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 147(??):ii, October 1, 2023. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S089812212300384X.

Anonymous:2023:EBbm

[Ano23-39]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 148(??):i, October 15, 2023. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122123004339.

Anonymous:2023:EBbn

[Ano23-40]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 148(??):ii, October 15, 2023. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122123004340.

Anonymous:2023:EBbo

[Ano23-41]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 149(??):i, November 1, 2023. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122123004479.

Anonymous:2023:EBbp

[Ano23-42]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 149(??):ii, November 1, 2023. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122123004480.

Anonymous:2023:EBbq

[Ano23-43]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 150(??):i, November 15, 2023. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122123004832.

Anonymous:2023:EBbr

[Ano23-44]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 150(??):ii, November 15, 2023. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122123004844.

Anonymous:2023:EBbs

[Ano23-45]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 151(??):i, December 1, 2023. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122123005151.

Anonymous:2023:EBbt

[Ano23-46]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 151(??):ii, December 1, 2023. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122123005163.

Anonymous:2023:EBbu

[Ano23-47]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 152(??):i, December 15, 2023. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122123005308.

Anonymous:2023:EBbv

[Ano23-48]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 152(??):ii, December 15, 2023. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S089812212300531X.

Anonymous:2023:PMd

[Ano23-49] Anonymous. Pages 1–120 (15 May 2023). Computers and Mathematics with Applications, 138(??):1–120, May 15, 2023. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic).

Anonymous:2023:PJb

[Ano23-50] Anonymous. Pages 1–148 (15 January 2023). Computers and Mathematics with Applications, 130(??):1–148, January 15, 2023. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic).

Anonymous:2023:PJa

[Ano23-51] Anonymous. Pages 1–150 (1 January 2023). Computers and Mathematics with Applications, 129(??):1–150, January 1, 2023. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic).

Anonymous:2023:PFa

[Ano23-52] Anonymous. Pages 1–158 (1 February 2023). Computers and Mathematics with Applications, 131(??):1–158, February 1, 2023. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic).

Anonymous:2023:PMc

[Ano23-53] Anonymous. Pages 1–172 (1 May 2023). Computers and Mathematics with Applications, 137(??):1–172, May 1, 2023. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic).

Anonymous:2023:PFb

[Ano23-54] Anonymous. Pages 1–180 (15 February 2023). Computers and Mathematics with Applications, 132(??):1–180, February 15, 2023. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic).

Anonymous:2023:PMa

[Ano23-55] Anonymous. Pages 1–188 (1 March 2023). Computers and Mathematics with Applications, 133(??):1–188, March 1, 2023. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic).

Anonymous:2023:PAb

[Ano23-56] Anonymous. Pages 1–190 (15 April 2023). Computers and Mathematics with Applications, 136(??):1–190, April 15, 2023. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic).

Anonymous:2023:PAa

[Ano23-57] Anonymous. Pages 1–206 (1 April 2023). Computers and Mathematics with Applications, 135(??):1–206, April 1, 2023. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic).

Anonymous:2023:PMb

[Ano23-58] Anonymous. Pages 1–224 (15 March 2023). Computers and Mathematics with Applications, 134(??):1–224, March 15, 2023. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic).

Anonymous:2023:PJc

[Ano23-59] Anonymous. Pages 1–230 (1 June 2023). Computers and Mathematics with Applications, 139(??):1–230, June 1, 2023. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic).

Anonymous:2023:PJe

[Ano23-60] Anonymous. Pages 1–254 (1 July 2023). Computers and Mathematics with Applications, 141(??):1–254, July 1, 2023. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic).

Anonymous:2023:PNa

[Ano23-61] Anonymous. Pages 1–258 (1 November 2023). Computers and Mathematics with Applications, 149(??):1–258, November 1, 2023. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic).

Anonymous:2023:PNb

[Ano23-62] Anonymous. Pages 1–266 (15 November 2023). Computers and Mathematics with Applications, 150(??):1–266, November 15, 2023. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic).

Anonymous:2023:PJd

[Ano23-63] Anonymous. Pages 1–290 (15 June 2023). Computers and Mathematics with Applications, 140(??):1–290, June 15, 2023. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic).

Anonymous:2023:POa

[Ano23-64] Anonymous. Pages 1–300 (1 October 2023). Computers and Mathematics with Applications, 147(??):1–300, October 1, 2023. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic).

Anonymous:2023:PAd

[Ano23-65] Anonymous. Pages 1–322 (15 August 2023). Computers and Mathematics with Applications, 144(??):1–322, August 15, 2023. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic).

Anonymous:2023:PJf

[Ano23-66] Anonymous. Pages 1–322 (15 July 2023). Computers and Mathematics with Applications, 142(??):1–322, July 15, 2023. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic).

Anonymous:2023:POb

[Ano23-67] Anonymous. Pages 1–338 (15 October 2023). Computers and Mathematics with Applications, 148(??):1–338, October 15, 2023. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic).

Anonymous:2023:PSa

[Ano23-68] Anonymous. Pages 1–340 (1 September 2023). Computers and Mathematics with Applications, 145(??):1–340, September 1, 2023. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic).

Anonymous:2023:PSb

[Ano23-69] Anonymous. Pages 1–360 (15 September 2023). Computers and Mathematics with Applications, 146(??):1–360, September 15, 2023. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic).

Anonymous:2023:PDb

[Ano23-70] Anonymous. Pages 1–378 (15 December 2023). Computers and Mathematics with Applications, 152(??):1–378, December 15, 2023. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic).

Anonymous:2023:PAc

[Ano23-71] Anonymous. Pages 1–434 (1 August 2023). Computers and Mathematics with Applications, 143(??):1–434, August 1, 2023. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic).

Anonymous:2023:PDa

[Ano23-72] Anonymous. Pages 1–508 (1 December 2023). Computers and Mathematics with Applications, 151(??):1–508, December 1, 2023. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic).

Anonymous:2024:EBa

[Ano24a]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 153(??):i, January 1, 2024. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122123005710.

Anonymous:2024:EBb

[Ano24b]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 153(??):ii, January 1, 2024. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122123005722.

Anonymous:2024:EBc

[Ano24c]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 154(??):i, January 15, 2024. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122123005862.

Anonymous:2024:EBd

[Ano24d]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 154(??):ii, January 15, 2024. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122123005874.

Anonymous:2024:EBe

[Ano24e]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 155(??):i, February 1, 2024. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122124000154.

Anonymous:2024:EBf

[Ano24f]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 155(??):ii, February 1, 2024. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122124000166.

Anonymous:2024:EBg

[Ano24g]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 156(??):i, February 15, 2024. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122124000282.

Anonymous:2024:EBh

[Ano24h]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 156(??):ii, February 15, 2024. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122124000294.

Anonymous:2024:EBi

[Ano24i]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 157(??):i, March 1, 2024. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0898122124000609.

Anonymous:2024:EBj

[Ano24j]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 157(??):ii, March 1, 2024. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0898122124000610.

Anonymous:2024:EBk

[Ano24k]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 158(??):i, March 15, 2024. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0898122124000956.

Anonymous:2024:EBl

[Ano24l]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 158(??):ii, March 15, 2024. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122124000968.

Anonymous:2024:EBm

[Ano24m]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 159(??):i, April 1, 2024. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122124001159.

Anonymous:2024:EBn

[Ano24n]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 159(??):ii, April 1, 2024. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122124001160.

Anonymous:2024:EBo

[Ano24o]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 160(??):i, April 15, 2024. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S089812212400138X.

Anonymous:2024:EBp

[Ano24p]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 160(??):ii, April 15, 2024. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0898122124001391.

Anonymous:2024:EBq

[Ano24q]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 161(??):i, May 1, 2024. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S089812212400155X.

Anonymous:2024:EBr

[Ano24r]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 161(??):ii, May 1, 2024. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0898122124001561.

Anonymous:2024:EBs

[Ano24s]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 162(??):i, May 15, 2024. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122124001834.

Anonymous:2024:EBt

[Ano24t]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 162(??):ii, May 15, 2024. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122124001846.

Anonymous:2024:EBu

[Ano24u]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 163(??):i, June 1, 2024. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122124002013.

Anonymous:2024:EBv

[Ano24v]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 163(??):ii, June 1, 2024. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122124002025.

Anonymous:2024:EBw

[Ano24w]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 164(??):i, June 15, 2024. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0898122124002190.

Anonymous:2024:EBx

[Ano24x]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 164(??):ii, June 15, 2024. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0898122124002207.

Anonymous:2024:EBy

[Ano24y]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 165(??):i, July 1, 2024. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122124002359.

Anonymous:2024:EBz

[Ano24z]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 165(??):ii, July 1, 2024. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0898122124002360.

Anonymous:2024:EBba

[Ano24-27]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 166(??):i, July 15, 2024. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122124002530.

Anonymous:2024:EBbb

[Ano24-28]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 166(??):ii, July 15, 2024. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122124002542.

Anonymous:2024:EBbc

[Ano24-29]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 167(??):i, August 1, 2024. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122124002669.

Anonymous:2024:EBbd

[Ano24-30]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 167(??):ii, August 1, 2024. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122124002670.

Anonymous:2024:EBbe

[Ano24-31]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 168(??):i, August 15, 2024. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0898122124003286.

Anonymous:2024:EBbf

[Ano24-32]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 168(??):ii, August 15, 2024. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122124003298.

Anonymous:2024:EBbg

[Ano24-33]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 169(??):i, September 1, 2024. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122124003535.

Anonymous:2024:EBbh

[Ano24-34]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 169(??):ii, September 1, 2024. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122124003547.

Anonymous:2024:EBbi

[Ano24-35]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 170(??):i, September 15, 2024. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122124003602.

Anonymous:2024:EBbj

[Ano24-36]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 170(??):ii, September 15, 2024. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122124003614.

Anonymous:2024:EBbk

[Ano24-37]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 171(??):i, October 1, 2024. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122124003778.

Anonymous:2024:EBbl

[Ano24-38]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 171(??):ii, October 1, 2024. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S089812212400378X.

Anonymous:2024:EBbm

[Ano24-39]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 172(??):i, October 15, 2024. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122124003900.

Anonymous:2024:EBbn

[Ano24-40]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 172(??):ii, October 15, 2024. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122124003912.

Anonymous:2024:EBbo

[Ano24-41]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 173(??):i, November 1, 2024. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122124004085.

Anonymous:2024:EBbp

[Ano24-42]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 173(??):ii, November 1, 2024. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122124004097.

Anonymous:2024:EBbq

[Ano24-43]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 174(??):i, November 15, 2024. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122124004796.

Anonymous:2024:EBbr

[Ano24-44]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 174(??):ii, November 15, 2024. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122124004802.

Anonymous:2024:PDa

[Ano24-45] Anonymous. In progress (1 December 2024). Computers and Mathematics with Applications, 175(??):??, December 1, 2024. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic).

Anonymous:2024:PDb

[Ano24-46] Anonymous. In progress (15 December 2024). Computers and Mathematics with Applications, 176(??):??, December 15, 2024. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic).

Anonymous:2024:PNa

[Ano24-47] Anonymous. Pages 1–150 (1 November 2024). Computers and Mathematics with Applications, 173(??):1–150, November 1, 2024. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic).

Anonymous:2024:PJd

[Ano24-48] Anonymous. Pages 1–166 (15 June 2024). Computers and Mathematics with Applications, 164(??):1–166, June 15, 2024. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic).

Anonymous:2024:PAb

[Ano24-49] Anonymous. Pages 1–208 (15 April 2024). Computers and Mathematics with Applications, 160(??):1–208, April 15, 2024. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic).

Anonymous:2024:POa

[Ano24-50] Anonymous. Pages 1–214 (1 October 2024). Computers and Mathematics with Applications, 171(??):1–214, October 1, 2024. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic).

Anonymous:2024:PMd

[Ano24-51] Anonymous. Pages 1–214 (15 May 2024). Computers and Mathematics with Applications, 162(??):1–214, May 15, 2024. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic).

Anonymous:2024:PFa

[Ano24-52] Anonymous. Pages 1–224 (1 February 2024). Computers and Mathematics with Applications, 155(??):1–224, February 1, 2024. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic).

Anonymous:2024:PJe

[Ano24-53] Anonymous. Pages 1–224 (1 July 2024). Computers and Mathematics with Applications, 165(??):1–224, July 1, 2024. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic).

Anonymous:2024:PMc

[Ano24-54] Anonymous. Pages 1–224 (1 May 2024). Computers and Mathematics with Applications, 161(??):1–224, May 1, 2024. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic).

Anonymous:2024:PMa

[Ano24-55] Anonymous. Pages 1–234 (1 March 2024). Computers and Mathematics with Applications, 157(??):1–234, March 1, 2024. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic).

Anonymous:2024:PFb

[Ano24-56] Anonymous. Pages 1–238 (15 February 2024). Computers and Mathematics with Applications, 156(??):1–238, February 15, 2024. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic).

Anonymous:2024:POb

[Ano24-57] Anonymous. Pages 1–240 (15 October 2024). Computers and Mathematics with Applications, 172(??):1–240, October 15, 2024. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic).

Anonymous:2024:PJc

[Ano24-58] Anonymous. Pages 1–242 (1 June 2024). Computers and Mathematics with Applications, 163(??):1–242, June 1, 2024. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic).

Anonymous:2024:PAd

[Ano24-59] Anonymous. Pages 1–254 (15 August 2024). Computers and Mathematics with Applications, 168(??):1–254, August 15, 2024. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic).

Anonymous:2024:PJa

[Ano24-60] Anonymous. Pages 1–256 (1 January 2024). Computers and Mathematics with Applications, 153(??):1–256, January 1, 2024. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic).

Anonymous:2024:PJ

[Ano24-61] Anonymous. Pages 1–268 (15 July 2024). Computers and Mathematics with Applications, 166(??):1–268, July 15, 2024. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic).

Anonymous:2024:PMb

[Ano24-62] Anonymous. Pages 1–274 (15 March 2024). Computers and Mathematics with Applications, 158(??):1–274, March 15, 2024. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic).

Anonymous:2024:PJb

[Ano24-63] Anonymous. Pages 1–286 (15 January 2024). Computers and Mathematics with Applications, 154(??):1–286, January 15, 2024. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic).

Anonymous:2024:PSa

[Ano24-64] Anonymous. Pages 1–298 (1 September 2024). Computers and Mathematics with Applications, 169(??):1–298, September 1, 2024. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic).

Anonymous:2024:PSb

[Ano24-65] Anonymous. Pages 1–300 (15 September 2024). Computers and Mathematics with Applications, 170(??):1–300, September 15, 2024. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic).

Anonymous:2024:PAa

[Ano24-66] Anonymous. Pages 1–318 (1 April 2024). Computers and Mathematics with Applications, 159(??):1–318, April 1, 2024. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic).

Anonymous:2024:PAc

[Ano24-67] Anonymous. Pages 1–328 (1 August 2024). Computers and Mathematics with Applications, 167(??):1–328, August 1, 2024. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic).

Anonymous:2024:PNb

[Ano24-68] Anonymous. Pages 1–552 (15 November 2024). Computers and Mathematics with Applications, 174(??):1–552, November 15, 2024. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic).

Anonymous:2025:EBa

[Ano25a]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 177(??):i, January 1, 2025. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122124005455.

Anonymous:2025:EBb

[Ano25b]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 177(??):ii, January 1, 2025. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122124005467.

Anonymous:2025:EBc

[Ano25c]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 178(??):i, January 15, 2025. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122124005662.

Anonymous:2025:EBd

[Ano25d]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 178(??):ii, January 15, 2025. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122124005674.

Anonymous:2025:EBe

[Ano25e]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 179(??):i, February 1, 2025. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122125000197.

Anonymous:2025:EBf

[Ano25f]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 179(??):ii, February 1, 2025. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122125000203.

Anonymous:2025:EBg

[Ano25g]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 180(??):i, February 15, 2025. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122125000434.

Anonymous:2025:EBh

[Ano25h]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 180(??):ii, February 15, 2025. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122125000446.

Anonymous:2025:EBi

[Ano25i]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 181(??):i, March 1, 2025. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0898122125000537.

Anonymous:2025:EBj

[Ano25j]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 181(??):ii, March 1, 2025. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0898122125000549.

Anonymous:2025:EBk

[Ano25k]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 182(??):i, March 15, 2025. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0898122125000756.

Anonymous:2025:EBl

[Ano25l]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 182(??):ii, March 15, 2025. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122125000768.

Anonymous:2025:EBm

[Ano25m]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 183(??):i, April 1, 2025. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122125000847.

Anonymous:2025:EBn

[Ano25n]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 183(??):ii, April 1, 2025. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122125000859.

Anonymous:2025:EBo

[Ano25o]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 184(??):i, April 15, 2025. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122125000914.

Anonymous:2025:EBp

[Ano25p]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 184(??):ii, April 15, 2025. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122125000926.

Anonymous:2025:EBq

[Ano25q]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 185(??):i, May 1, 2025. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122125001336.

Anonymous:2025:EBr

[Ano25r]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 185(??):ii, May 1, 2025. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122125001348.

Anonymous:2025:EBs

[Ano25s]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 186(??):i, May 15, 2025. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122125001439.

Anonymous:2025:EBt

[Ano25t]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 186(??):ii, May 15, 2025. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122125001440.

Anonymous:2025:EBu

[Ano25u]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 187(??):i, June 1, 2025. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122125001671.

Anonymous:2025:EBv

[Ano25v]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 187(??):ii, June 1, 2025. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122125001683.

Anonymous:2025:EBw

[Ano25w]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 188(??):i, June 15, 2025. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0898122125001865.

Anonymous:2025:EBx

[Ano25x]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 188(??):ii, June 15, 2025. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0898122125001877.

Anonymous:2025:EBy

[Ano25y]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 189(??):i, July 1, 2025. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122125001981.

Anonymous:2025:EBz

[Ano25z]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 189(??):ii, July 1, 2025. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0898122125001993.

Anonymous:2025:EBba

[Ano25-27]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 190(??):i, July 15, 2025. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122125002172.

Anonymous:2025:EBbb

[Ano25-28]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 190(??):ii, July 15, 2025. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122125002184.

Anonymous:2025:EBbc

[Ano25-29]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 191(??):i, August 1, 2025. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122125002330.

Anonymous:2025:EBbd

[Ano25-30]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 191(??):ii, August 1, 2025. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122125002342.

Anonymous:2025:EBbe

[Ano25-31]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 192(??):i, August 15, 2025. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0898122125002433.

Anonymous:2025:EBbf

[Ano25-32]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 192(??):ii, August 15, 2025. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122125002445.

Anonymous:2025:EBbg

[Ano25-33]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 193(??):i, September 1, 2025. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122125002901.

Anonymous:2025:EBbh

[Ano25-34]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 193(??):ii, September 1, 2025. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122125002913.

Anonymous:2025:EBbi

[Ano25-35]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 194(??):i, September 15, 2025. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122125003451.

Anonymous:2025:EBbj

[Ano25-36]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 194(??):ii, September 15, 2025. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122125003463.

Anonymous:2025:EBbk

[Ano25-37]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 195(??):i, October 1, 2025. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S089812212500375X.

Anonymous:2025:EBbl

[Ano25-38]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 195(??):ii, October 1, 2025. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122125003761.

Anonymous:2025:EBbm

[Ano25-39]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 196(??):i, October 15, 2025. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122125003876.

Anonymous:2025:EBbn

[Ano25-40]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 196(??):ii, October 15, 2025. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122125003888.

Anonymous:2025:EBbo

[Ano25-41]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 197(??):i, November 1, 2025. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122125004201.

Anonymous:2025:EBbp

[Ano25-42]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 197(??):ii, November 1, 2025. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122125004213.

Anonymous:2025:EBbq

[Ano25-43]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 198(??):i, November 15, 2025. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122125004390.

Anonymous:2025:EBbr

[Ano25-44]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 198(??):ii, November 15, 2025. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122125004407.

Anonymous:2025:EBbs

[Ano25-45]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 199(??):i, December 1, 2025. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122125004493.

Anonymous:2025:EBbt

[Ano25-46]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 199(??):ii, December 1, 2025. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S089812212500450X.

Anonymous:2025:PMc

[Ano25-47] Anonymous. Pages 1–110 (1 May 2025). Computers and Mathematics with Applications, 185(??):1–110, May 1, 2025. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic).

Anonymous:2025:PJa

[Ano25-48] Anonymous. Pages 1–166 (1 January 2025). Computers and Mathematics with Applications, 177(??):1–166, January 1, 2025. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic).

Anonymous:2025:PMd

[Ano25-49] Anonymous. Pages 1–170 (15 May 2025). Computers and Mathematics with Applications, 186(??):1–170, May 15, 2025. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic).

Anonymous:2025:PJb

[Ano25-50] Anonymous. Pages 1–214 (15 January 2025). Computers and Mathematics with Applications, 178(??):1–214, January 15, 2025. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic).

Anonymous:2025:PJf

[Ano25-51] Anonymous. Pages 1–218 (15 July 2025). Computers and Mathematics with Applications, 190(??):1–218, July 15, 2025. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic).

Anonymous:2025:PAb

[Ano25-52] Anonymous. Pages 1–220 (15 April 2025). Computers and Mathematics with Applications, 184(??):1–220, April 15, 2025. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic).

Anonymous:2025:PJe

[Ano25-53] Anonymous. Pages 1–224 (1 July 2025). Computers and Mathematics with Applications, 189(??):1–224, July 1, 2025. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic).

Anonymous:2025:PFa

[Ano25-54] Anonymous. Pages 1–242 (1 February 2025). Computers and Mathematics with Applications, 179(??):1–242, February 1, 2025. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic).

Anonymous:2025:PJc

[Ano25-55] Anonymous. Pages 1–248 (1 June 2025). Computers and Mathematics with Applications, 187(??):1–248, June 1, 2025. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic).

Anonymous:2025:PAd

[Ano25-56] Anonymous. Pages 1–258 (15 August 2025). Computers and Mathematics with Applications, 192(??):1–258, August 15, 2025. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic).

Anonymous:2025:PJd

[Ano25-57] Anonymous. Pages 1–264 (15 June 2025). Computers and Mathematics with Applications, 188(??):1–264, June 15, 2025. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic).

Anonymous:2025:PAa

[Ano25-58] Anonymous. Pages 1–286 (1 April 2025). Computers and Mathematics with Applications, 183(??):1–286, April 1, 2025. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic).

Anonymous:2025:PMb

[Ano25-59] Anonymous. Pages 1–290 (15 March 2025). Computers and Mathematics with Applications, 182(??):1–290, March 15, 2025. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic).

Anonymous:2025:PAc

[Ano25-60] Anonymous. Pages 1–296 (1 August 2025). Computers and Mathematics with Applications, 191(??):1–296, August 1, 2025. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic).

Anonymous:2025:PNa

[Ano25-61] Anonymous. Pages 1–316 (1 November 2025). Computers and Mathematics with Applications, 197(??):1–316, November 1, 2025. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic).

Anonymous:2025:PFb

[Ano25-62] Anonymous. Pages 1–322 (15 February 2025). Computers and Mathematics with Applications, 180(??):1–322, February 15, 2025. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic).

Anonymous:2025:PMa

[Ano25-63] Anonymous. Pages 1–350 (1 March 2025). Computers and Mathematics with Applications, 181(??):1–350, March 1, 2025. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic).

Anonymous:2025:PSa

[Ano25-64] Anonymous. Pages 1–364 (1 September 2025). Computers and Mathematics with Applications, 193(??):1–364, September 1, 2025. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic).

Anonymous:2025:PNb

[Ano25-65] Anonymous. Pages 1–376 (15 November 2025). Computers and Mathematics with Applications, 198(??):1–376, November 15, 2025. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic).

Anonymous:2025:PSb

[Ano25-66] Anonymous. Pages 1–378 (15 September 2025). Computers and Mathematics with Applications, 194(??):1–378, September 15, 2025. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic).

Anonymous:2025:PD

[Ano25-67] Anonymous. Pages 1–422 (1 December 2025). Computers and Mathematics with Applications, 199(??):1–422, December 1, 2025. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic).

Anonymous:2025:POa

[Ano25-68] Anonymous. Pages 1–432 (1 October 2025). Computers and Mathematics with Applications, 195(??):1–432, October 1, 2025. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic).

Anonymous:2025:POb

[Ano25-69] Anonymous. Pages 1–476 (15 October 2025). Computers and Mathematics with Applications, 196(??):1–476, October 15, 2025. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic).

Antunes:2021:NCE

[Ant21] Pedro R. S. Antunes. Numerical calculation of extremal Steklov eigenvalues in 3D and 4D. Computers and Mathematics with Applications, 104(??):50–58, December 15, 2021. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122121004053.

Aldakheel:2021:GLA

[ANWW21] Fadi Aldakheel, Nima Noii, Thomas Wick, and Peter Wriggers. A global–local approach for hydraulic phasefield fracture in poroelastic media. Computers and Mathematics with Applications, 91 (??):99–121, June 1, 2021. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122120302790.

Ahmat:2023:SOF

[ANZZ23] Muyassar Ahmat, SuYuan Ni, Min Zhang, and Zhuang Zhao. A sixth-order finite difference HWENO scheme for nonlinear degenerate parabolic equation. Computers and Mathematics with Applications, 150(??):196– 210, November 15, 2023. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122123003942.

Abaali:2024:MMC

[AO24] M. Abaali and S. Ouchtout. Mathematical modeling of the chemotaxis–biodenitrification process. Computers and Mathematics with Applications, 169(??):56–67, September 1, 2024. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122124002748.

Alaeian:2022:KFT

[AOL22] Mohsen Alaeian, Helcio R. B. Orlande, and Bernard Lamien. Kalman filter temperature estimation with a photoacoustic observation model during the hyperthermia treatment of cancer. Computers and Mathematics with Applications, 119(??): 193–207, August 1, 2022. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122122002528.

Alebrahim:2025:PUT

[AP25] Reza Alebrahim and Riccardo Panciroli. Periultrasound theory for Lamb wave propagation across an interface-debonded bimaterial plates. Computers and Mathematics with Applications, 194(??):18–30, September 15, 2025. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122125002470.

Alihussein:2024:DCL

[APGK24] Hussein Alihussein, Adarsh Prasannakumar, Martin Geier, and Manfred Krafczyk. Direct cumulant lattice Boltzmann simulations of transitional flow in gyroidal structures including experimental validation. Computers and Mathematics with Applications, 157(??):159– 172, March 1, 2024. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122123005904.

Abbaszadeh:2025:RKP

[APK^{+}25] Mostafa Abbaszadeh, Maryam Parvizi, Amirreza Khodadadian, Thomas Wick, and Mehdi Dehghan. A reproducing kernel particle method (RKPM) algorithm for solving the tropical Pacific Ocean model. Computers and Mathematics with Applications, 179(??):197– 211, February 1, 2025. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122124005571.

Acharya:2025:CPH

[APPS25] Sanjib K. Acharya, Amiya K. Pani, Ajit Patel, and Ravina Shokeen. Conservative primal hybrid finite element method for weakly damped Klein–Gordon equation. Computers and Mathematics with Applications, 186 (??):16–36, May 15, 2025. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122125000963.

Arshad:2021:MMM

[APS21] Muhammad Arshad, Eun-Jae Park, and Dongwook Shin. Multiscale mortar mixed domain decomposition approximations of nonlinear parabolic equations. Computers and Mathematics with Applications, 97(??): 375–385, September 1, 2021. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122121002479.

Armentano:2025:HFE

[APS25] María Gabriela Armentano, Claudio Padra, and Mario Scheble. An hp finite element a posteriori analysis of a nonstandard eigenvalue problem arising in fluid–solid vibrations on curved domains. Computers and Mathematics with Applications, 184 (??):65–85, April 15, 2025. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S089812212500063X.

Albella:2023:NAP

[ARV23] J. Albella, R. Rodríguez, and P. Venegas. Numerical approximation of a potentials formulation for the elasticity vibration problem. Computers and Mathematics with Applications, 137 (??):61–72, May 1, 2023. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122123000639.

Abedian:2020:RFD

[AS20] Rooholah Abedian and Rezvan Salehi. A RBFWENO finite difference scheme for Hamilton–Jacobi equations. Computers and Mathematics with Applications, 79 (7):2002–2020, April 2020. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122119304833.

Abdedou:2021:NIR

[AS21] Azzedine Abdedou and Azzeddine Soulaïmani. A non-intrusive reduced-order modeling for uncertainty propagation of time-dependent problems using a B-splines Bézier elements-based method and proper orthogonal decomposition: Application to dam-break flows. Computers and Mathematics with Applications, 102(??):187– 205, November 15, 2021.

CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122121003540.

Ahmed:2023:FSA

[AS23] Shady E. Ahmed and Omer San. Forward sensitivity analysis and mode dependent control for closure modeling of Galerkin systems. Computers and Mathematics with Applications, 145(??): 289–302, September 1, 2023. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122123002900.

Ahmed:2025:MPN

[ASB25] Nadeem Ahmed, Suneet Singh, and Ram Prakash Bharti. Modified preconditioned Newton–Krylov approaches for Navier–Stokes equations using nodal integral method. Computers and Mathematics with Applications, 181(??):163– 192, March 1, 2025. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122124005777.

Amilo:2025:NID

[ASF^{+}25] David Amilo, Khadijeh Sadri, Muhammad Farman, Evren Hincal, and Kottakkaran Sooppy Nisar. New insights into disease dynamics and treatment interventions with PID controllerbased therapeutic strategies for pancreatic cancer. Computers and Mathematics with Applications, 181 (??):133–162, March 1, 2025. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122125000070.

Acosta-Soba:2023:SPU

[ASGGRG23] Daniel Acosta-Soba, Francisco Guillén-González, and J. Rafael Rodríguez-Galván. A structure-preserving upwind DG scheme for a degenerate phase-field tumor model. Computers and Mathematics with Applications, 152(??):317–333, December 15, 2023. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122123004741.

Almasi:2021:NSM

[ASH^{+}21] F. Almasi, M. S. Shadloo, A. Hadjadj, M. Ozbulut, N. Tofighi, and M. Yildiz. Numerical simulations of multi-phase electro-hydrodynamics flows using a simple incompressible smoothed particle hydrodynamics method. Computers and Mathematics with Applications, 81(??): 772–785, January 1, 2021.

CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122119305243.

Arregui:2020:PMA

[ASSV20] Iñigo Arregui, Beatriz Salvador, Daniel Sevcovic, and Carlos Vázquez. PDE models for American options with counterparty risk and two stochastic factors: Mathematical analysis and numerical solution. Computers and Mathematics with Applications, 79(5):1525– 1542, March 1, 2020. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S089812211930464X.

Artioli:2020:ACP

[ASV20] E. Artioli, A. Sommariva, and M. Vianello. Algebraic cubature on polygonal elements with a circular edge. Computers and Mathematics with Applications, 79 (7):2057–2066, April 2020. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122119305115.

Athanasiadou:2024:TWS

[ASZ24] E. S. Athanasiadou, V. Sevroglou, and S. Zoi. Thermoelastic wave scattering by a multilayered object. Computers and Mathematics with Applications, 163(??):186– 200, June 1, 2024. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122124001354.

Asgari:2020:EBG

[ATBT20] Z. Asgari, F. Toutounian, E. Babolian, and E. Tohidi. An extended block Golub–Kahan algorithm for large algebraic and differential matrix Riccati equations. Computers and Mathematics with Applications, 79(8): 2447–2457, April 15, 2020. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122119305437.

Ali:2022:CFA

[AWA^{+}22] Liaqat Ali, Ya-Jie Wu, Bagh Ali, Sohaib Abdal, and Sajjad Hussain. The crucial features of aggregation in TiO_{2}-water nanofluid aligned of chemically comprising microorganisms: a FEM approach. Computers and Mathematics with Applications, 123(??):241– 251, October 1, 2022. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122122003492.

Axelsson:2020:OPS

[Axe20] Owe Axelsson. Optimality properties of a square block matrix preconditioner with applications. Computers and Mathematics with Applications, 80(2): 286–294, July 2020. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122119304808.

Ai:2024:OPU

[AXLW24] Xia Ai, Wei Xu, Li-Dan Liao, and Xiang Wang. Optimized parameterized Uzawa methods for solving complex Helmholtz equations. Computers and Mathematics with Applications, 164 (??):34–44, June 15, 2024. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122124001457.

Apel:2024:IAL

[AZ24] Thomas Apel and Philipp Zilk. Isogeometric analysis of the Laplace eigenvalue problem on circular sectors: Regularity properties and graded meshes. Computers and Mathematics with Applications, 175(??): 236–254, December 1, 2024. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122124004280.

An:2022:RNA

[AZHY22] Na An, Guoye Zhao, Chaobao Huang, and Xijun Yu. αrobust H^{1}-norm analysis of a finite element method for the superdiffusion equation with weak singularity solutions. Computers and Mathematics with Applications, 118 (??):159–170, July 15, 2022. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122122002036.

Ai:2020:FDM

[AZL20] Zhi Yong Ai, Yong Zhi Zhao, and Wen Jie Liu. Fractional derivative modeling for axisymmetric consolidation of multilayered cross-anisotropic viscoelastic porous media. Computers and Mathematics with Applications, 79(5):1321– 1334, March 1, 2020. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122119304365.

An:2022:NSO

[AZY22] Jing An, Jun Zhang, and Xiaofeng Yang. A novel secondorder time accurate fully discrete finite element scheme with decoupling structure for the hydrodynamicallycoupled phase field crystal model. Computers and Mathematics with Applications, 113(??):70–85, May 1, 2022. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122122000360.

Brun:2020:MSS

[BAB^{+}20] Mats Kirkesæther Brun, Elyes Ahmed, Inga Berre, Jan Martin Nordbotten, and Florin Adrian Radu. Monolithic and splitting solution schemes for fully coupled quasi-static thermoporoelasticity with nonlinear convective transport. Computers and Mathematics with Applications, 80(8): 1964–1984, October 15, 2020. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122120303308.

Busuioc:2020:TDL

[BABS20] Sergiu Busuioc, Victor E.

Ambrus, Tonino Biciusca, and Victor Sofonea. Twodimensional off-lattice Boltzmann model for van der Waals fluids with variable temperature. Computers and Mathematics with Applications, 79(1):111–140, January 2020. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122118307077.

Bause:2024:CCG

[BAKR24] Markus Bause, Mathias Anselmann, Uwe Köcher, and Florin A. Radu. Convergence of a continuous Galerkin method for hyperbolic-parabolic systems. Computers and Mathematics with Applications, 158(??):118–138, March 15, 2024. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122124000208.

Balashov:2021:DSD

[Bal21]  V. A. Balashov.  Dis-

sipative spatial discretization of a phase field model of multiphase multicomponent isothermal fluid flow. Computers and Mathematics with Applications, 90 (??):112–124, May 15, 2021. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122121000870.

Bastiani:2021:TPH

[BAM^{+}21] Camila D’ Bastiani, Jéferson Luis Alba, Gabriel Tomazzoni Mazzarotto, Severino Rodrigues de Farias Neto, Anthony Reynolds, David Kennedy, and Lademir Luiz Beal. Three-phase hydrodynamic simulation and experimental validation of an upflow anaerobic sludge blanket reactor. Computers and Mathematics with Applications, 83(??):95–110, February 1, 2021. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122120300845.

Bartels:2021:NDC

[Bar21]  Sören Bartels.  Noncon-

forming discretizations of convex minimization problems and precise relations to mixed methods. Computers and Mathematics with Applications, 93(??):214– 229, July 1, 2021. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122121001541.

Barua:2022:BIM

[Bar22] Amlan K. Barua. A boundary integral method to investigate pattern formation in a rotating Hele–Shaw cell with time dependent gap. Computers and Mathematics with Applications, 118 (??):36–45, July 15, 2022. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S089812212200181X.

Bertrand:2021:LSF

[BB21] Fleurianne Bertrand and Daniele Boffi. Least-squares formulations for eigenvalue problems associated with linear elasticity. Computers and Mathematics with Applications, 95(??):19–27, August 1, 2021. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122120304831.

Benkhaldoun:2023:NAA

[BB23a] Fayssal Benkhaldoun and Abdallah Bradji. Novel analysis approach for the convergence of a second order time accurate mixed finite element scheme for parabolic equations. Computers and Mathematics with Applications, 133 (??):85–103, March 1, 2023. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122123000019.

Bersetche:2023:DFO

[BB23b] Francisco M. Bersetche and Juan Pablo Borthagaray. A deep First-Order System Least Squares method for solving elliptic PDEs. Computers and Mathematics with Applications, 129(??): 136–150, January 1, 2023. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122122004771.

Banz:2024:PEE

[BB24] L. Banz and F. Bertrand.

A posteriori error estimate for contact problems in porous media. Computers and Mathematics with Applications, 174(??):219–229, November 15, 2024. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122124003572.

Bastian:2021:DFB

[BBD^{+}21a] Peter Bastian, Markus Blatt, Andreas Dedner, Nils-Arne Dreier, Christian Engwer, René Fritze, Carsten Gräser, Christoph Grüninger, Dominic Kempf, Robert Klöfkorn, Mario Ohlberger, and Oliver Sander. The Dune framework: Basic concepts and recent developments. Computers and Mathematics with Applications, 81(??):75–112, January 1, 2021. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S089812212030256X.

Botti:2021:AAF

[BBD21b] Lorenzo Botti, Michele Botti, and Daniele A. Di Pietro. An abstract analysis framework for monolithic discretisations of poroelasticity with application to hybrid high-order methods. Computers and Mathematics with Applications, 91 (??):150–175, June 1, 2021. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122120302534.

Bertrand:2024:IDP

[BBD^{+}24] Fleurianne Bertrand, Daniele Boffi, Alexander Düster, Jean-Luc Guermond, Norbert Heuer, Jichun Li, and Waldemar Rachowicz. Innovative discretizations of PDEs: Towards an accurate representation of the reality. Computers and Mathematics with Applications, 176(??): 221–223, December 15, 2024. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122124004565.

Bonaldi:2021:GDT

[BBDM21] Francesco Bonaldi, Konstantin Brenner, Jérôme Droniou, and Roland Masson. Gradient discretization of two-phase flows coupled with mechanical deformation in fractured porous media. Computers and Mathematics with Applications, 98(??): 40–68, September 15, 2021. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122121002558.

Bertrand:2023:AME

[BBG23] Fleurianne Bertrand, Daniele Boffi, and Lucia Gastaldi. Approximation of the Maxwell eigenvalue problem in a leastsquares setting. Computers and Mathematics with Applications, 148(??):302–312, October 15, 2023. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122123003486.

Becker:2022:ROG

[BBI^{+}22] Roland Becker, Maximilian Brunner, Michael Innerberger, Jens Markus Melenk, and Dirk Praetorius. Rate-optimal goal-oriented adaptive FEM for semilinear elliptic PDEs. Computers and Mathematics with Applications, 118(??):18– 35, July 15, 2022. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S089812212200195X.

Bause:2021:FGO

[BBK21] Markus Bause, Marius Paul Bruchhäuser, and Uwe Köcher. Flexible goal-oriented adaptivity for higher-order space– time discretizations of transport problems with coupled flow. Computers and Mathematics with Applications, 91(??):17–35, June 1, 2021. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122120303424.

Beik:2024:AIS

[BBK24] Fatemeh P. A. Beik, Michele Benzi, and Mehdi Najafi Kalyani. Accelerating iterative solvers via a twodimensional minimum residual technique. Computers and Mathematics with Applications, 166(??):224– 236, July 15, 2024. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122124002128.

Berrone:2025:LOS

[BBM25] Stefano Berrone, Andrea Borio, and Francesca Marcon. Lowest order stabilization free virtual element method for the 2D Poisson equation. Computers and Mathematics with Applications, 177 (??):78–99, January 1, 2025. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122124005145.

Billo:2021:FEP

[BBS21] Georis Billo, Michel Belliard, and Pierre Sagaut. A finite element penalized direct forcing immersed boundary method for infinitely thin obstacles in a dilatable flow. Computers and Mathematics with Applications, 99(??): 292–304, October 1, 2021. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122121002893.

Billo:2022:CSI

[BBS22] Georis Billo, Michel Belliard, and Pierre Sagaut. Comparison of several interpolation methods to reconstruct field data in the vicinity of a finite element immersed boundary. Computers and Mathematics with Applications, 123(??): 123–135, October 1, 2022. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122122003157.

Boujoudar:2025:LMM

[BBT25] Mohamed Boujoudar, Abdelaziz Beljadid, and Ahmed Taik. LRBF meshless methods for predicting soil moisture distribution in root zone. Computers and Mathematics with Applications, 179(??):178–196, February 1, 2025. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S089812212400525X.

Brachet:2020:FSS

[BC20a] Matthieu Brachet and Jean-Paul Chehab. Fast and stable schemes for phase fields models. Computers and Mathematics with Applications, 80(6):1683–1713, September 15, 2020. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122120302819.

Burazin:2020:COC

[BC20b] Kresimir Burazin and Ivana Crnjac. Convergence of the optimality criteria method for multiple state optimal design problems. Computers and Mathematics with Applications, 79(5):1382– 1392, March 1, 2020. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122119304432.

Bonaventura:2020:SAM

[BCCR20] L. Bonaventura, F. Casella, L. Delpopolo Carciopolo, and A. Ranade. A self adjusting multirate algorithm for robust time discretization of partial differential equations. Computers and Mathematics with Applications, 79 (7):2086–2098, April 2020. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122119305565.

Boffi:2023:CNM

[BCF^{+}23] Daniele Boffi, Andrea Cangiani, Marco Feder, Lucia Gastaldi, and Luca Heltai. A comparison of non-matching techniques for the finite element approximation of interface problems. Computers and Mathematics with Applications, 151(??):101–115, December 1, 2023. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122123004029.

Bock:2024:PCR

[BCK24] Andreas Bock, Colin J. Cotter, and Robert C. Kirby. Planar curve registration using Bayesian inversion. Computers and Mathematics with Applications, 159(??):155– 172, April 1, 2024. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122124000427.

Bucataru:2022:GBR

[BCM22] Mihai Bucataru, Iulian Cîmpean, and Liviu Marin. A gradient-based regularization algorithm for the Cauchy problem in steadystate anisotropic heat conduction. Computers and Mathematics with Applications, 119(??):220–240, August 1, 2022. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122122002590.

Borachok:2025:TSN

[BCM25] Ihor Borachok, Roman Chapko, and Leonidas Mindrinos. Two-step numerical methods for a coupled parabolic-hyperbolic transmission problem. Computers and Mathematics with Applications, 177(??):115– 129, January 1, 2025. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122124005121.

Blachere:2021:VHO

[BCT21] Florian Blachère, Christophe Chalons, and Rodolphe Turpault. Very highorder asymptotic-preserving schemes for hyperbolic systems of conservation laws with parabolic degeneracy on unstructured meshes. Computers and Mathematics with Applications, 87 (??):41–49, April 1, 2021. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122121000390.

Biswas:2020:EWS

[BD20] Biswarup Biswas and Ritesh Kumar Dubey. ENO and WENO schemes using arclength based smoothness measurement. Computers and Mathematics with Applications, 80(12):2780–2795, December 15, 2020. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122120303989.

Bonnet:2023:ECA

[BD23] Marc Bonnet and Edouard Demaldent. Eddy-current asymptotics of the Maxwell PMCHWT formulation for multiple bodies and conductivity levels. Computers and Mathematics with Applications, 141(??):80– 101, July 1, 2023. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S089812212300127X.

Belov:2021:BSS

[BDB21] A. A. Belov, Zh. O. Dombrovskaya, and A. N. Bogolyubov. A bicompact scheme and spectral decomposition method for difference solution of Maxwell’s equations in layered media. Computers and Mathematics with Applications, 96(??): 178–187, August 15, 2021. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122121002017.

Beuchler:2024:MMN

[BDE^{+}24] S. Beuchler, A. Demircan, B. Endtmayer, U. Morgner, and T. Wick. Mathematical modeling and numerical multigoal-oriented a posteriori error control and adaptivity for a stationary, nonlinear, coupled flow temperature model with temperature dependent density. Computers and Mathematics with Applications, 175(??): 138–151, December 1, 2024. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122124004243.

Bhal:2024:ACN

[BDF24] Santosh Kumar Bhal, P. Danumjaya, and G. Fairweather. An ADI Crank–Nicolson orthogonal spline collocation method for 2D parabolic problems with an interface. Computers and Mathematics with Applications, 160 (??):142–147, April 15, 2024. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122124000658.

Bertrand:2021:RAL

[BDG21] Fleurianne Bertrand, Leszek Demkowicz, and Jay Gopalakrishnan. Recent advances in least-squares and discontinuous Petrov–Galerkin finite element methods. Computers and Mathematics with Applications, 95(??):1– 3, August 1, 2021. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122121002157.

Bracco:2025:PMF

[BDGS25] Cesare Bracco, Oleg Davydov, Carlotta Giannelli, and Alessandra Sestini. A positive meshless finite difference scheme for scalar conservation laws with adaptive artificial viscosity driven by fault detection. Computers and Mathematics with Applications, 190(??):103– 121, July 15, 2025. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122125001518.

Bleistein:2020:PIO

[BDJ20] T. Bleistein, S. Diebels, and A. Jung. Parameter identification for open cell aluminium foams using inverse calculation. Computers and Mathematics with Applications, 79(9):2644– 2654, May 1, 2020. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122119305553.

Bootland:2021:CCS

[BDJT21] Niall Bootland, Victorita Dolean, Pierre Jolivet, and Pierre-Henri Tournier. A comparison of coarse spaces for Helmholtz problems in the high frequency regime. Computers and Mathematics with Applications, 98 (??):239–253, September 15, 2021. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0898122121002741.

Berardi:2020:MMT

[BDL20] Marco Berardi, Fabio Difonzo, and Luciano Lopez. A mixed MoL–TMoL for the numerical solution of the 2D Richards’ equation in layered soils. Computers and Mathematics with Applications, 79 (7):1990–2001, April 2020. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122119303785.

Bacuta:2021:ANC

[BDMX21] Constantin Bacuta, Leszek Demkowicz, Jaime Mora, and Christos Xenophontos. Analysis of non-conforming DPG methods on polyhedral meshes using fractional Sobolev norms. Computers and Mathematics with Applications, 95(??):215–241, August 1, 2021. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122120303801.

Berardi:2023:NMN

[BDP23] Marco Berardi, Fabio V.

Difonzo, and Sabrina F. Pellegrino. A numerical method for a nonlocal form of Richards’ equation based on peridynamic theory. Computers and Mathematics with Applications, 143(??):23–32, August 1, 2023. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122123001761.

Breda:2022:BCC

[BDS^{+}22] Dimitri Breda, Simone De Reggi, Francesca Scarabel, Rossana Vermiglio, and Jianhong Wu. Bivariate collocation for computing R_{0} in epidemic models with two structures. Computers and Mathematics with Applications, 116(??):15– 24, June 15, 2022. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122121003801.

Bouraoui:2024:EST

[BDS24] Hamed Abderrahmane Bouraoui, Abdelhak Djebabla, and Abdourazek Souahi. Exponential stability of Timoshenko beams with threephase-lag thermoelasticity. Computers and Mathematics with Applications, 168 (??):58–83, August 15, 2024. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122124002499.

Baptista:2024:SIN

[BdSC24] Riedson Baptista, Isaac P. dos Santos, and Lucia Catabriga. Solving incompressible Navier–Stokes equations: a nonlinear multiscale approach. Computers and Mathematics with Applications, 175(??):366–384, December 1, 2024. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S089812212400453X.

Badia:2023:STU

[BDV23] Santiago Badia, Hridya Dilip, and Francesc Verdugo. Space-time unfitted finite element methods for time-dependent problems on moving domains. Computers and Mathematics with Applications, 135 (??):60–76, April 1, 2023. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S089812212300038X.

Bauer:2021:WBS

[BEG^{+}21] Martin Bauer, Sebastian Eibl, Christian Godenschwager, Nils Kohl, Michael Kuron, Christoph Rettinger, Florian Schornbaum, Christoph Schwarzmeier, Dominik Thönnes, Harald Köstler, and Ulrich Rüde. waLBerla: a blockstructured high-performance framework for multiphysics simulations. Computers and Mathematics with Applications, 81(??):478–501, January 1, 2021. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122120300146.

Barrera:2021:GCP

[BEIL21] D. Barrera, S. Eddargani, M. J. Ibáñez, and A. Lamnii. A geometric characterization of Powell–sabin triangulations allowing the construction of C ^{2} quartic splines. Computers and Mathematics with Applications, 100 (??):30–40, October 15, 2021. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122121003205.

Brykov:2021:FRG

[BEKV21] N. A. Brykov, V. N. Emelyanov, A. G. Karpenko, and K. N. Volkov. Flows of real gas in nozzles with unsteady local energy supply. Computers and Mathematics with Applications, 81(??): 702–724, January 1, 2021. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122119305814.

Bellotti:2024:IPM

[Bel24] Thomas Bellotti. The influence of parasitic modes on stable lattice Boltzmann schemes and weakly unstable multi-step finite difference schemes. Computers and Mathematics with Applications, 174(??):397–416, November 15, 2024. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122124004371.

BenSalah:2025:BRT

[Ben25] Mohamed BenSalah. Bayesian recovery of time-varying fractional order in timefractional diffusion equations for shale gas applications. Computers and Mathematics with Applications, 191(??): 228–244, August 1, 2025. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122125002123.

Bertrand:2021:RRF

[BER21] Fleurianne Bertrand, Alexandre Ern, and Florin Adrian Radu. Robust and reliable finite element methods in poromechanics. Computers and Mathematics with Applications, 91(??): 1–2, June 1, 2021. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122121001449.

Bohn:2021:RNN

[BF21] Jan Bohn and Michael Feischl. Recurrent neural networks as optimal mesh refinement strategies. Computers and Mathematics with Applications, 97(??):61–76, September 1, 2021. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S089812212100198X.

Bao:2025:NPA

[BF25] Tiantian Bao and Xiufang Feng. A new parallel algorithm with high-order finite difference scheme for solving the Helmholtz equation in two and three dimensions. Computers and Mathematics with Applications, 183 (??):71–97, April 1, 2025.

CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122125000276.

Bohlen:2021:VAF

[BFE^{+}21] Thomas Bohlen, Mario Ruben Fernandez, Johannes Ernesti, Christian Rheinbay, Andreas Rieder, and Christian Wieners. Visco-acoustic full waveform inversion: From a DG forward solver to a Newton–CG inverse solver. Computers and Mathematics with Applications, 100(??): 126–140, October 15, 2021. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122121003217.

Bergamaschi:2023:PMF

[BFI^{+}23] L. Bergamaschi, M. Ferronato, G. Isotton, C. Janna, and A. Martínez. Parallel matrix-free polynomial preconditioners with application to flow simulations in discrete fracture networks. Computers and Mathematics with Applications, 146(??): 60–70, September 15, 2023. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122123002845.

Bahr:2024:IHF

[BFM24] Björn Bahr, Markus Faustmann, and Jens Markus Melenk. An implementation of hp-FEM for the fractional Laplacian. Computers and Mathematics with Applications, 176(??):324– 348, December 15, 2024. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122124004486.

Bringmann:2025:FLC

[BFM^{+}25] Philipp Bringmann, Michael Feischl, Ani Miraçi, Dirk Praetorius, and Julian Streitberger. On full linear convergence and optimal complexity of adaptive FEM with inexact solver. Computers and Mathematics with Applications, 180(??):102–129, February 15, 2025. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122124005583.

Borges:2021:DRV

[BFRA21] J. S. Borges, J. A. Ferreira, G. Romanazzi, and E. Abreu. Drug release from viscoelastic polymeric matrices — a stable and supraconvergent FDM. Computers and Mathematics with Applications, 99(??):257–269, October 1, 2021. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122121002911.

BenGharbia:2023:AIS

[BFVY23] Ibtihel Ben Gharbia, Joëlle Ferzly, Martin Vohralík, and Soleiman Yousef. Adaptive inexact smoothing Newton method for a nonconforming discretization of a variational inequality. Computers and Mathematics with Applications, 133(??):12– 29, March 1, 2023. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122122005004.

Burkovska:2022:OBA

[BGD22] Olena Burkovska, Christian Glusa, and Marta D’Elia. An optimization-based approach to parameter learning for fractional type nonlocal models. Computers and Mathematics with Applications, 116(??):229– 244, June 15, 2022. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122121001851.

Benito:2020:SCH

[BGG^{+}20] J. J. Benito, A. García, L. Gavete, M. Negreanu, F. Ureña, and A. M. Vargas. Solving a chemotaxis– haptotaxis system in 2D using Generalized Finite Difference Method. Computers and Mathematics with Applications, 80(5):762–777, September 1, 2020. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122120302066.

Berrone:2023:OBC

[BGG^{+}23] Stefano Berrone, Chiara Giverso, Denise Grappein, Luigi Preziosi, and Stefano Scialò. An optimization based 3D–1D coupling strategy for tissue perfusion and chemical transport during tumor-induced angiogenesis. Computers and Mathematics with Applications, 151(??): 252–270, December 1, 2023. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122123004376.

Beams:2020:PSM

[BGH20] Natalie N. Beams, Adrianna Gillman, and Russell J. Hewett. A parallel sharedmemory implementation of a high-order accurate solution technique for variable coefficient Helmholtz problems. Computers and Mathematics with Applications, 79(4): 996–1011, February 15, 2020. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S089812211930416X.

Beck:2023:BMN

[BGK23] Mélanie Beck, Martin J. Gander, and Felix Kwok. Bmethods for the numerical solution of evolution problems with blow-up solutions part II: Splitting methods. Computers and Mathematics with Applications, 152(??): 143–154, December 15, 2023. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122123004595.

Bracco:2020:IAH

[BGKV20] Cesare Bracco, Carlotta Giannelli, Mario Kapl, and Rafael Vázquez. Isogeometric analysis with C^{1} hierarchical functions on planar two-patch geometries. Computers and Mathematics with Applications, 80(11): 2538–2562, December 2020. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S089812212030122X.

Boulakia:2024:MNA

[BGLZ24] Muriel Boulakia, Céline Grandmont, Fabien Lespagnol, and Paolo Zunino. Mathematical and numerical analysis of reduced order interface conditions and augmented finite elements for mixed dimensional problems. Computers and Mathematics with Applications, 175(??): 536–569, December 1, 2024. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122124004723.

Bermudez:2020:NNM

[BGPS20] A. Bermúdez, D. Gómez, M. Piñeiro, and P. Salgado. A novel numerical method for accelerating the computation of the steadystate in induction machines. Computers and Mathematics with Applications, 79(2): 274–292, January 15, 2020. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122119303359.

Bajpai:2023:FDF

[BGR23] Saumya Bajpai, Deepjyoti Goswami, and Kallol Ray. Fully discrete finite element error analysis of a discontinuous Galerkin method for the Kelvin–Voigt viscoelastic fluid model. Computers and Mathematics with Applications, 130(??):69–97, January 15, 2023. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S089812212200476X.

Bokanowski:2022:OPA

[BGZ22] Olivier Bokanowski, Nidhal Gammoudi, and Hasnaa Zidani. Optimistic planning algorithms for stateconstrained optimal control problems. Computers and Mathematics with Applications, 109(??):158– 179, March 1, 2022. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122122000232.

Benalia:2022:NMI

[BH22] S. Benalia and M. Hachama. A nonlocal method for image shadow removal. Computers and Mathematics with Applications, 107(??):95–103, February 1, 2022. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122121004636.

Bozorgmanesh:2020:ITT

[BHC20] Hassan Bozorgmanesh, Masoud Hajarian, and Anthony Theodore Chronopoulos. Interval tensors and their application in solving multilinear systems of equations. Computers and Mathematics with Applications, 79(3): 697–715, February 1, 2020. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122119303761.

Bosman:2021:DDS

[BHD^{+}21] Dries Bosman, Martijn Huynen, Daniël De Zutter, Hendrik Rogier, and Dries Vande Ginste. A 2-D differential surface admittance operator for= combined magnetic and dielectric contrast. Computers and Mathematics with Applications, 102(??): 175–186, November 15, 2021. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S089812212100359X.

Bosman:2022:CDS

[BHDV22] Dries Bosman, Martijn Huynen, Daniël De Zutter, and Dries Vande Ginste. Construction of the differential surface admittance operator with an extended Fokas method for electromagnetic scattering at polygonal objects with arbitrary material parameters. Computers and Mathematics with Applications, 128(??):44–54, December 15, 2022. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122122004345.

Bacuta:2021:NSP

[BHJ21] Constantin Bacuta, Daniel Hayes, and Jacob Jacavage. Notes on a saddle point reformulation of mixed variational problems. Computers and Mathematics with Applications, 95 (??):4–18, August 1, 2021. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122120302820.

Bacuta:2022:EDP

[BHJ22] Constantin Bacuta, Daniel Hayes, and Jacob Jacavage. Efficient discretization and preconditioning of the singularly perturbed reactiondiffusion problem. Computers and Mathematics with Applications, 109(??):270– 279, March 1, 2022. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122122000372.

Badia:2024:ERD

[BHK^{+}24] Santiago Badia, Martin Hornkjøl, Arbaz Khan, Kent-André Mardal, Alberto F. Martín, and Ricardo Ruiz-Baier. Efficient and reliable divergence-conforming methods for an elasticityporoelasticity interface problem. Computers and Mathematics with Applications, 157 (??):173–194, March 1, 2024. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122123006004.

Bulle:2023:HPE

[BHL^{+}23] Raphaël Bulle, Jack S. Hale, Alexei Lozinski, Stéphane P. A. Bordas, and Franz Chouly. Hierarchical a posteriori error estimation of Bank–Weiser type in the FEniCS Project. Computers and Mathematics with Applications, 131(??):103– 123, February 1, 2023. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122122004722.

Blanc:2024:CFV

[BHL24] Xavier Blanc, Philippe Hoch, and Clément Lasuen. Composite finite volume schemes for the diffusion equation on unstructured meshes. Computers and Mathematics with Applications, 156(??):207– 217, February 15, 2024. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122123005965.

Badger:2023:SDM

[BHPD23] Jacob Badger, Stefan Henneking, Socratis Petrides, and Leszek Demkowicz. Scalable DPG multigrid solver for Helmholtz problems: a study on convergence. Computers and Mathematics with Applications, 148(??):81–92, October 15, 2023. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122123003012.

Barnafi:2024:RPN

[BHPS24] Nicolás A. Barnafi, Ngoc Mai Monica Huynh, Luca F. Pavarino, and Simone Scacchi. Robust parallel nonlinear solvers for implicit time discretizations of the bidomain equations with staggered ionic models. Computers and Mathematics with Applications, 167(??):134– 149, August 1, 2024. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122124001731.

Banz:2020:PEC

[BHS20] Lothar Banz, Michael Hintermüller, and Andreas Schröder. A posteriori error control for distributed elliptic optimal control problems with control constraints discretized by hpfinite elements. Computers and Mathematics with Applications, 80(11):2433– 2450, December 2020. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S089812212030314X.

Banz:2022:PPE

[BHS22] Lothar Banz, Orlando Hernández, and Ernst P. Stephan. A priori and a posteriori error estimates for hp-FEM for a Bingham type variational inequality of the second kind. Computers and Mathematics with Applications, 126(??): 14–30, November 15, 2022. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122122003777.

Banz:2025:HFE

[BHS25] Lothar Banz, Michael Hintermüller, and Andreas Schröder. hp-finite elements for elliptic optimal control problems with control constraints. Computers and Mathematics with Applications, 196(??):288–311, October 15, 2025. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122125003190.

Batista:2020:ASP

[BHZ20] Juan Batista, Xiaozhe Hu, and Ludmil T. Zikatanov. Auxiliary space preconditioning for mixed finite element discretizations of Richards’ equation. Computers and Mathematics with Applications, 80(2): 405–416, July 2020. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122119304596.

Bahssini:2024:DEF

[BIMS24] Abderrahim Bahssini, Nouh Izem, M. Shadi Mohamed, and Mohammed Seaid. Directional enrichment functions for finite element solutions of transient anisotropic diffusion. Computers and Mathematics with Applications, 163(??):42–55, June 1, 2024. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0898122124001238.

Belgacem:2023:UFF

[BJG23] F. Ben Belgacem, F. Jelassi, and V. Girault. Uniqueness’ failure for the finite element Cauchy–Poisson’s problem. Computers and Mathematics with Applications, 135 (??):77–92, April 1, 2023. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S089812212300010X.

Brenner:2024:IPM

[BJyST24] Susanne C. Brenner, SeongHee Jeong, Li yeng Sung, and Zhiyu Tan. C^{0} interior penalty methods for an elliptic distributed optimal control problem with general tracking and pointwise state constraints. Computers and Mathematics with Applications, 155(??):80–90, February 1, 2024. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122123005369.

Biswas:2022:ESD

[BKB22] Biswarup Biswas, Harish Kumar, and Deepak Bhoriya. Entropy stable discontinuous Galerkin schemes for the special relativistic hydrodynamics equations. Computers and Mathematics with Applications, 112(??):55– 75, April 15, 2022. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122122000827.

Bollapragada:2023:ASA

[BKK^{+}23] Raghu Bollapragada, Cem Karamanli, Brendan Keith, Boyan Lazarov, Socratis Petrides, and Jingyi Wang. An adaptive sampling augmented Lagrangian method for stochastic optimization with deterministic constraints. Computers and Mathematics with Applications, 149(??):239–258, November 1, 2023. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122123003991.

Bendaida:2023:NLI

[BKM23] Fatiha Bendaida, Fahd Karami, and Driss Meskine. Nonlocal p-Laplacian involving a nonlinear fractional reaction–diffusion system applied to image restoration. Computers and Mathematics with Applications, 152(??): 56–66, December 15, 2023. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S089812212300442X.

BenAmeur:2024:PEA

[BKR24] Hend Ben Ameur, Nizar Kharrat, and Mohamed Hedi Riahi. A posteriori error analysis for a multidimensional hydrogeological parameter estimation in a time dependent model. Computers and Mathematics with Applications, 169(??):237– 252, September 1, 2024. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S089812212400292X.

Boledi:2023:SCN

[BKT^{+}23] Leonardo Boledi, Fabian Key, Benjamin Terschanski, Stefanie Elgeti, and Julia Kowalski. A scale-coupled numerical method for transient close-contact melting.

Computers and Mathematics with Applications, 143(??): 277–288, August 1, 2023. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122123002055.

Bartels:2022:NAO

[BKW22] Sören Bartels, Hedwig Keller, and Gerd Wachsmuth. Numerical approximation of optimal convex and rotationally symmetric shapes for an eigenvalue problem arising in optimal insulation. Computers and Mathematics with Applications, 119(??): 327–339, August 1, 2022. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S089812212200219X.

Barbosa:2020:ABE

[BL20] João Paulo Barbosa and Carlos Friedrich Loeffler. Application of Boundary Element Method superposition technique for solving natural frequencies in piecewise homogeneous domains. Computers and Mathematics with Applications, 79(4):1131–1144, February 15, 2020. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122119304298.

Bai:2023:NCS

[BLF23] Yanhong Bai, Yang Li, and Minfu Feng. A new class of stabilized virtual element methods for the timedependent Oseen equations. Computers and Mathematics with Applications, 145(??): 303–317, September 1, 2023. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122123002936.

Ba:2025:EBE

[BLLM25] Yuming Ba, Qiuqi Li, Zehua Li, and Lingling Ma. An ensemble-based efficient iterative method for uncertainty quantification of partial differential equations with random inputs. Computers and Mathematics with Applications, 182(??):256– 274, March 15, 2025. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122125000483.

Ba:2024:HLB

[BLLY24] Yan Ba, Haihu Liu, Wenqiang Li, and Wenjing Yang. A hybrid lattice Boltzmann and finite difference method for two-phase flows with soluble surfactants. Computers and Mathematics with Applications, 174(??):325– 339, November 15, 2024. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122124004310.

Boyana:2024:CAN

[BLLZ24] Satyajith Bommana Boyana, Thomas Lewis, Sijing Liu, and Yi Zhang. Convergence analysis of novel discontinuous Galerkin methods for a convection dominated problem. Computers and Mathematics with Applications, 175(??):224–235, December 1, 2024. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S089812212400436X.

Bressan:2024:ISO

[BLM^{+}24] Andrea Bressan, Gabriele Loli, Sauro Manenti, Alessandro Reali, and Giancarlo Sangalli. An isogeometric shape optimization method for groundwater flow in porous media. Computers and Mathematics with Applications, 162(??):104– 119, May 15, 2024. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122124000889.

Borio:2024:LOS

[BLMV24] Andrea Borio, Carlo Lovadina, Francesca Marcon, and Michele Visinoni. A lowest order stabilizationfree mixed Virtual Element Method. Computers and Mathematics with Applications, 160(??):161– 170, April 15, 2024. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122124000695.

Bause:2022:CVD

[BLO22a] Markus Bause, Maria Lymbery, and Kevin Osthues. C^{1}-conforming variational discretization of the biharmonic wave equation. Computers and Mathematics with Applications, 119(??):208– 219, August 1, 2022. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122122002498.

Blommaert:2022:EBE

[BLO22b] Robin Blommaert, Srdan Lazendić, and Ljubica Oparnica. The Euler–Bernoulli equation with distributional coefficients and forces. Computers and Mathematics with Applications, 123(??):171– 183, October 1, 2022. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122122003303.

Bermudez:2020:NAP

[BLRRS20] Alfredo Bermúdez, Bibiana López-Rodríguez, Rodolfo Rodríguez, and Pilar Salgado. Numerical analysis of a penalty approach for the solution of a transient eddy current problem. Computers and Mathematics with Applications, 79(9): 2503–2526, May 1, 2020. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122119305486.

Bao:2024:AAB

[BLX^{+}24] Chunxu Bao, Lin Liu, Chiyu Xie, Jing Zhu, Yufeng Quan, Siyu Chen, Libo Feng, and Liancun Zheng. Analysis of the absorbing boundary condition for the Maxwell fluid flow over a semi-infinite plate with considering the magnetic field. Computers and Mathematics with Applications, 154(??):212–223, January 15, 2024. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122123005461.

Bartels:2020:EIS

[BM20] Sören Bartels and Marijo Milicevic. Efficient iterative solution of finite element discretized nonsmooth minimization problems. Computers and Mathematics with Applications, 80(5):588–603, September 1, 2020. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122120301735.

Bonaventura:2021:TBM

[BM21] Luca Bonaventura and Macarena Gómez Mármol. The TR-BDF2 method for second order problems in structural mechanics. Computers and Mathematics with Applications, 92(??):13–26, June 15, 2021. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122121001267.

Bussac:2023:SHR

[BM23] J. Bussac and H. Mathis.

Simulation of a homogeneous relaxation model for a three-phase mixture with miscible phases. Computers and Mathematics with Applications, 143(??):327– 347, August 1, 2023. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122123002110.

Bernkopf:2024:OCR

[BM24a] M. Bernkopf and J. M. Melenk. Optimal convergence rates in L^{2} for a first order system least squares finite element method — part II: Inhomogeneous Robin boundary conditions. Computers and Mathematics with Applications, 173(??):1–18, November 1, 2024. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122124003419.

Boroujerdi:2024:MUC

[BM24b] R. Boroujerdi and M. M.

Monfared. Multiple unequal cracks between a functionally graded piezoelectric layer and a piezoelectric substrate by distributed strain nuclei. Computers and Mathematics with Applications, 170(??):116–141, September 15, 2024. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122124002992.

Bahramifar:2022:LMM

[BMH22] Saeed Bahramifar, Farshid Mossaiby, and Hamed Haftbaradaran. A local meshless method for transient nonlinear problems: Preliminary investigation and application to phase-field models. Computers and Mathematics with Applications, 124(??): 163–187, October 15, 2022. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122122003479.

Benedusi:2021:ECS

[BMK21] Pietro Benedusi, Michael L. Minion, and Rolf Krause. An experimental comparison of a space-time multigrid method with PFASST for a reaction-diffusion problem. Computers and Mathematics with Applications, 99(??): 162–170, October 1, 2021. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122121002716.

Bahador:2022:FCS

[BML22] F. Gholami Bahador, P. Mokhtary, and M. Lakestani. A fractional coupled system for simultaneous image denoising and deblurring. Computers and Mathematics with Applications, 128(??):285– 299, December 15, 2022. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122122004552.

Barroso:2021:NTS

[BMP21] Josue S. Barroso, Marcio A. Murad, and Patricia A. Pereira. A new two-scale computational model for hydromechanical coupling in jointed rocks. Computers and Mathematics with Applications, 91(??):67–98, June 1, 2021. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122120303850.

Brevis:2024:LQI

[BMP^{+}24] Ignacio Brevis, Ignacio Muga, David Pardo, Oscar Rodriguez, and Kristoffer G. van der Zee. Learning quantities of interest from parametric PDEs: an efficient neural-weighted Minimal Residual approach. Computers and Mathematics with Applications, 164 (??):139–149, June 15, 2024. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122124001640.

Bertoluzza:2022:SNV

[BMPP22] S. Bertoluzza, G. Manzini, M. Pennacchio, and D. Prada. Stabilization of the nonconforming virtual element method. Computers and Mathematics with Applications, 116(??):25–47, June 15, 2022. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122121003576.

Breiten:2022:EBP

[BMS22] Tobias Breiten, Riccardo Morandin, and Philipp Schulze. Error bounds for port-Hamiltonian model and controller reduction based on system balancing. Computers and Mathematics with Applications, 116(??):100– 115, June 15, 2022. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122121003102.

Bosy:2020:DDM

[BMST20] Michał Bosy, Monica Montardini, Giancarlo Sangalli, and Mattia Tani. A domain decomposition method for isogeometric multi-patch problems with inexact local solvers. Computers and Mathematics with Applications, 80(11):2604–2621, December 2020. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122120303321.

Bernal:2023:HPD

[BMVA23] Francisco Bernal, Jorge Morón-Vidal, and Juan A. Acebrón. A hybrid probabilistic domain decomposition algorithm suited for very large-scale elliptic PDEs. Computers and Mathematics with Applications, 146(??):294–308, September 15, 2023. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122123002985.

Brevis:2021:MLM

[BMvdZ21] Ignacio Brevis, Ignacio Muga, and Kristoffer G. van der Zee. A machinelearning minimal-residual (ML-MRes) framework for goal-oriented finite element discretizations. Computers and Mathematics with Applications, 95(??):186–199, August 1, 2021. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122120303199.

Bai:2020:CGM

[BMZ20] Yunru Bai, Stanisław Migórski, and Shengda Zeng. A class of generalized mixed variational-hemivariational inequalities I: Existence and uniqueness results. Computers and Mathematics with Applications, 79(10):2897– 2911, May 15, 2020. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122119305991.

Bonizzoni:2020:RSA

[BN20] Francesca Bonizzoni and Fabio Nobile. Regularity and sparse approximation of the recursive first moment equations for the lognormal Darcy problem. Computers and Mathematics with Applications, 80(12):2925– 2947, December 15, 2020. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122120304168.

Bateniparvar:2020:SFH

[BNB20] O. Bateniparvar, N. Noormohammadi, and B. Boroomand. Singular functions for heterogeneous composites with cracks and notches; the use of equilibrated singular basis functions. Computers and Mathematics with Applications, 79(5):1461– 1482, March 1, 2020. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122119304560.

Bhole:2023:SBC

[BNC^{+}23] Ashish Bhole, Boniface Nkonga, José Costa, Guido Huijsmans, Stanislas Pamela, and Matthias Hoelzl. Stabilized bi-cubic Hermite Bézier finite element method with application to gas-plasma interactions occurring during massive material injection in tokamaks. Computers and Mathematics with Applications, 142(??):225– 256, July 15, 2023. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122123001773.

Barral:2021:NSS

[BNQ21] P. Barral, B. Nicolás, and P. Quintela. Numerical simulation of the shear stress produced by the hot metal jet on the blast furnace runner. Computers and Mathematics with Applications, 102(??): 146–159, November 15, 2021. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122121003618.

Bai:2024:LFN

[BNSZ24] Xixian Bai, Chunyan Niu, Dongyang Shi, and Yongwei Zhang. A leap-frog nodal discontinuous Galerkin method for Maxwell polynomial chaos Debye model. Computers and Mathematics with Applications, 162 (??):120–132, May 15, 2024. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122124001056.

Badia:2022:RHO

[BNV22] Santiago Badia, Eric Neiva, and Francesc Verdugo. Robust high-order unfitted finite elements by interpolationbased discrete extension. Computers and Mathematics with Applications, 127(??):

105–126, December 1, 2022. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122122004138.

Bolbotowski:2022:OVP

[Boł22] Karol Bołbotowski. Optimal vault problem — form finding through 2D convex program. Computers and Mathematics with Applications, 109(??):280– 324, March 1, 2022. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122122000268.

Burger:2020:NAT

[BOSV20] Raimund Bürger, Rafael Ordoñez, Mauricio Sepúlveda, and Luis Miguel Villada. Numerical analysis of a threespecies chemotaxis model. Computers and Mathematics with Applications, 80 (1):183–203, July 1, 2020. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122120301061.

Botchev:2020:NSC

[Bot20] M. A. Botchev. A nested Schur complement solver with mesh-independent convergence for the time domain photonics modeling. Computers and Mathematics with Applications, 80(2): 295–304, July 2020. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122119304079.

Boumenir:2024:DRM

[Bou24a] A. Boumenir. Direct reconstruction of a multidimensional heat equation. Computers and Mathematics with Applications, 176(??):91–98, December 15, 2024. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122124004176.

Bourel:2024:WFS

[Bou24b] Christophe Bourel. Water flow in shallow aquifers without the Dupuit hypothesis. Computers and Mathematics with Applications, 163 (??):165–185, June 1, 2024. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122124001342.

Bisht:2020:PLF

[BP20] Manju Bisht and Dhiraj V.

Patil. Power-law fluid flow in driven enclosures with undulation using MRTlattice Boltzmann method. Computers and Mathematics with Applications, 79 (1):100–110, January 2020.

CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122117305540.

Budiana:2020:MNS

[BPID20] Eko Prasetya Budiana, Pranowo, Indarto, and Deendarlianto. The meshless numerical simulation of Kelvin– Helmholtz instability during the wave growth of liquid– liquid slug flow. Computers and Mathematics with Applications, 80(7):1810–1838, October 1, 2020. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122120303138.

Briani:2022:ADC

[BPR22] M. Briani, G. Puppo, and M. Ribot. Angle dependence in coupling conditions for shallow water equations at channel junctions. Computers and Mathematics with Applications, 108(??): 49–65, February 15, 2022. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122121004612.

Bakalis:2024:HTE

[BPV24] Panteleimon A. Bakalis, Polycarpos K. Papadopoulos, and Panayiotis Vafeas. Heat transfer effect on the ferrofluid flow in a curved cylindrical annular duct under the influence of a magnetic field. Computers and Mathematics with Applications, 170(??):172–185, September 15, 2024. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122124002943.

Ballarin:2020:CRB

[BRÁ^{+}20a] Francesco Ballarin, Tomás Chacón Rebollo, Enrique Delgado Ávila, Macarena Gómez Mármol, and Gianluigi Rozza. Certified Reduced Basis VMS–Smagorinsky model for natural convection flow in a cavity with variable height. Computers and Mathematics with Applications, 80(5):973–989, September 1, 2020. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122120302170.

Bradji:2020:NAC

[Bra20b] Abdallah Bradji. A new analysis for the convergence of the gradient discretization method for multidimensional time fractional diffusion and diffusion-wave equations. Computers and Mathematics with Applications, 79(2): 500–520, January 15, 2020. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122119305255.

Bringmann:2024:RCC

[Bri24] Philipp Bringmann. Review and computational comparison of adaptive least-squares finite element schemes. Computers and Mathematics with Applications, 172(??):1–15, October 15, 2024. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122124003225.

Baldelli:2025:HTA

[BRR25] Andrés León Baldelli, Vitoriano Ruas, and Marco Antonio Silva Ramos. A handy tool for assessing tetrahedron-based finite-cell methods and for numerical simulations in spheroidal domains. Computers and Mathematics with Applications, 181(??):323–339, March 1, 2025. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122125000239.

Brdar:2021:SAF

[BRRT21] M. Brdar, G. Radojev, H.- G. Roos, and Lj. Teofanov. Superconvergence analysis of FEM and SDFEM on graded meshes for a problem with characteristic layers. Computers and Mathematics with Applications, 93(??):50–57, July 1, 2021. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122121001413.

Bespalov:2021:ITA

[BRS21] Alex Bespalov, Leonardo Rocchi, and David Silvester. T-IFISS: a toolbox for adaptive FEM computation. Computers and Mathematics with Applications, 81(??): 373–390, January 1, 2021. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122120301036.

Bertrand:2021:PEE

[BS21] Fleurianne Bertrand and Gerhard Starke. A posteriori error estimates by weakly symmetric stress reconstruction for the Biot problem. Computers and Mathematics with Applications, 91(??): 3–16, June 1, 2021. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122120304107.

Bahmani:2023:NSV

[BS23a] Erfan Bahmani and Ali Shokri. Numerical study of the variable-order timefractional mobile/immobile advection–diffusion equation using direct meshless local Petrov–Galerkin methods. Computers and Mathematics with Applications, 135 (??):111–123, April 1, 2023. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122123000317.

Bajpai:2023:PEE

[BS23b] Saumya Bajpai and Debendra Kumar Swain. A priori error estimates of a three-step two-level finite element Galerkin method for a 2D-boussinesq system of equations. Computers and Mathematics with Applications, 146(??):137–164, September 15, 2023. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122123002778.

Bouharguane:2023:DDG

[BS23c] Afaf Bouharguane and Nour Seloula. A direct discontinuous Galerkin method for a high order nonlocal conservation law. Computers and Mathematics with Applications, 141(??):1–14, July 1, 2023. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122123001177.

Bertrand:2024:LSF

[BS24] Fleurianne Bertrand and Henrik Schneider. Leastsquares finite element method for the simulation of seaice motion. Computers and Mathematics with Applications, 172(??):38–46, October 15, 2024. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122124003237.

Busto:2020:PGR

[BSRVC20] S. Busto, G. Stabile, G. Rozza, and M. E. Vázquez-Cendón. POD–Galerkin reduced order methods for combined Navier–Stokes transport equations based on a hybrid FV–FE solver. Computers and Mathematics with Applications, 79(2):256–273, January 15, 2020. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122119303293.

Bui:2022:NSD

[BT22] Tinh Quoc Bui and Hung Thanh Tran. Numerical simulations of dynamic fracture and fragmentation problems by a novel diffusive damage model. Computers and Mathematics with Applications, 125(??):193–212, November 1, 2022. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122122003637.

Bennani:2020:NOO

[BTCV20] Lokman Bennani, Pierre Trontin, Rémi Chauvin, and Philippe Villedieu. A non-overlapping optimized Schwarz method for the heat equation with non linear boundary conditions and with applications to deicing. Computers and Mathematics with Applications, 80(6):1500–1522, September 15, 2020. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122120302832.

Boguslawski:2024:ADD

[BTG24] A. Boguslawski, A. Tyliszczak, and B. J. Geurts. Approximate deconvolution discretisation. Computers and Mathematics with Applications, 154(??):175–198, January 15, 2024. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122123005424.

Brugiapaglia:2020:CIA

[BTT20] Simone Brugiapaglia, Lorenzo Tamellini, and Mattia Tani. Compressive isogeometric analysis. Computers and Mathematics with Applications, 80(12):3137–3155, December 15, 2020. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122120304296.

Bakalis:2021:NSD

[BTYM^{+}21] Georgios Bakalis, Kelsey C.

Tang-Yuk, Xiaocheng Mi, Nikos Nikiforakis, and Hoi Dick Ng. Numerical simulation of deflagration-to-detonation transition via shock–multiple flame kernels interactions. Computers and Mathematics with Applications, 83(??): 111–126, February 1, 2021. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122120302200.

Berrone:2021:RBM

[BV21] Stefano Berrone and Fabio Vicini. A reduced basis method for a PDEconstrained optimization formulation in discrete fracture network flow simulations. Computers and Mathematics with Applications, 99(??): 182–194, October 1, 2021. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S089812212100290X.

Baeyens:2023:SNE

[BV23] Toon Baeyens and Marnix Van Daele. Strands: New efficient technique for solving time-independent two-dimensional Schrödinger equations accurately. Computers and Mathematics with Applications, 152(??):334– 340, December 15, 2023. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122123004108.

Baumgarten:2021:PFE

[BW21] Niklas Baumgarten and Christian Wieners. The parallel finite element system M++ with integrated multilevel preconditioning and multilevel Monte Carlo methods. Computers and Mathematics with Applications, 81(??):391–406, January 1, 2021. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122120301024.

Boguslawski:2025:ADV

[BWG25] A. Boguslawski, K. Wawrzak, and B. J. Geurts. Approximate deconvolution and velocity estimation modeling of decaying Burgers turbulence. Computers and Mathematics with Applications, 193(??): 279–296, September 1, 2025. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122125002676.

Bu:2023:SOL

[BWMW23] Linlin Bu, Jianhua Wu, Liquan Mei, and Ying Wang. Second-order linear adaptive time-stepping schemes for the fractional Allen– Cahn equation. Computers and Mathematics with Applications, 145(??):260–274, September 1, 2023. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122123002808.

Bai:2021:NAF

[BWR21] Xixian Bai, Shuang Wang, and Hongxing Rui. Numerical analysis of finitedifference time-domain method for 2D/3D Maxwell’s equations in a Cole–Cole dispersive medium. Computers and Mathematics with Applications, 93(??):230– 252, July 1, 2021. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122121001553.

Bai:2022:UOS

[BWZ22] Yu Bai, Xin Wang, and Yan Zhang. Unsteady oblique stagnation-point flow and heat transfer of fractional Maxwell fluid with convective derivative under modified pressure field. Computers and Mathematics with Applications, 123(??):13–25, October 1, 2022. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122122002966.

Bai:2023:UOS

[BWZ23] Yu Bai, Xin Wang, and Yan Zhang. Unsteady oblique stagnation point flow with improved pressure field and fractional Cattaneo– Christov model by finite difference-spectral method. Computers and Mathematics with Applications, 147 (??):38–52, October 1, 2023. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122123003024.

Bespalov:2020:PEE

[BX20] Alex Bespalov and Feng Xu. A posteriori error estimation and adaptivity in stochastic Galerkin FEM for parametric elliptic PDEs: Beyond the affine case. Computers and Mathematics with Applications, 80(5):1084–1103, September 1, 2020. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122120302455.

Balam:2020:NEO

[BZ20] Reymundo Itzá Balam and Miguel Uh Zapata. A new eighth-order implicit finite difference method to solve the three-dimensional Helmholtz equation. Computers and Mathematics with Applications, 80(5):1176– 1200, September 1, 2020. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122120302601.

Boltuc:2021:PEP

[BZ21] Agnieszka Bołtuć and Eugeniusz Zieniuk. PIES for 2D elastoplastic problems with singular plastic strain fields. Computers and Mathematics with Applications, 103(??): 53–64, December 1, 2021. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122121003849.

Balam:2022:FOC

[BZ22] Reymundo Itza Balam and Miguel Uh Zapata. A fourth-order compact implicit immersed interface method for 2D Poisson interface problems. Computers and Mathematics with Applications, 119(??):257– 277, August 1, 2022. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122122002565.

Bi:2023:SAE

[BZ23] Hui Bi and Mengyuan Zhang. Stability analysis and error estimates of implicit Runge–Kutta local discontinuous Galerkin methods for linear bi-harmonic equation. Computers and Mathematics with Applications, 149(??): 211–220, November 1, 2023. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122123004078.

Barnafi:2021:MAN

[BZDQ21] Nicolás Barnafi, Paolo Zunino, Luca Dedè, and Alfio Quarteroni. Mathematical analysis and numerical approximation of a general linearized poro-hyperelastic model. Computers and Mathematics with Applications, 91 (??):202–228, June 1, 2021. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122120303060.

Barakat:2024:NAR

[BZE24] Mohamed Barakat, Waheed Zahra, and Ahmed Elsaid.

A new approach for recovering the gradient and a posteriori error estimates. Computers and Mathematics with Applications, 159 (??):202–218, April 1, 2024. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122124000488.

Bi:2020:TGD

[BZY20] Hai Bi, Yu Zhang, and Yidu Yang. Two-grid discretizations and a local finite element scheme for a nonselfadjoint Stekloff eigenvalue problem. Computers and Mathematics with Applications, 79(7):1895– 1913, April 2020. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122118304735.

Caldana:2024:DLA

[CAD24] Matteo Caldana, Paola F.

Antonietti, and Luca Dede’. A deep learning algorithm to accelerate algebraic multigrid methods in finite element solvers of 3D elliptic PDEs. Computers and Mathematics with Applications, 167(??):217–231, August 1, 2024. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122124002256.

Cao:2021:WSN

[Cao21] Liu Cao. Workpiece-scale numerical simulations of SLM molten pool dynamic behavior of 316L stainless steel. Computers and Mathematics with Applications, 96(??): 209–228, August 15, 2021. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S089812212030167X.

Chang:2023:WBE

[CBI^{+}23] Wei Shyang Chang, Muna Mohammed Bazuhair, Farzad Ismail, Hossain Chizari, and Mohammad Hafifi Hafiz Ishak. Well-balanced energystable residual distribution methods for the shallow water equations with varying bottom topography. Computers and Mathematics with Applications, 148(??):188– 199, October 15, 2023. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122123003449.

Chandrawat:2024:FPF

[CBK24] Rajesh Kumar Chandrawat, Gurpreet Singh Bhatia, and Deepak Kumar. Fractal patterns in fluid flows: a radial basis pseudospectral analysis of CuO-EG water nano, Saffman dusty, and plain Newtonian fluids in immiscible layers of a horizontal channel. Computers and Mathematics with Applications, 156(??):167–179, February 15, 2024. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S089812212300593X.

Calo:2023:KPL

[CBŁP23] V. M. Calo, P. Behnoudfar, M. Łoś, and M. Paszyński. A Kronecker product linearcost solver for the highorder generalized-α method for multi-dimensional hyperbolic systems. Computers and Mathematics with Applications, 142(??):257– 267, July 15, 2023. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122123001852.

Charkaoui:2024:NPM

[CBLZ24] Abderrahim Charkaoui, Anouar Ben-Loghfyry, and Shengda Zeng. A novel parabolic model driven by double phase flux operator with variable exponents: Application to image decomposition and denoising. Computers and Mathematics with Applications, 174(??):97–141, November 15, 2024. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122124003821.

Chen:2020:MNE

[CC20] Juan Chen and Fangqi Chen. Maximum norm error estimates of a linear CADI method for the Klein–Gordon–Schrödinger equations. Computers and Mathematics with Applications, 80(5):1327–1342, September 1, 2020. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122120302649.

Chen:2025:WGF

[CC25] Xiaoxiao Chen and Yanli Chen. A weak Galerkin finite element method for the two-dimensional Burgers’ equation. Computers and Mathematics with Applications, 196(??):84–103, October 15, 2025. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122125002986.

Chen:2025:GSF

[CCD25] Jipei Chen, Victor M. Calo, and Quanling Deng. Generalized soft finite element method for elliptic eigenvalue problems. Computers and Mathematics with Applications, 184(??):168– 184, April 15, 2025. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122125000665.

Cavalcanti:2021:WPA

[CCF^{+}21] Marcelo M. Cavalcanti, Wellington J. Corrêa, Andrei V. Faminskii, Mauricio A. Sepúlveda C., and Rodrigo Véjar-Asem. Wellposedness and asymptotic behavior of a generalized higher order nonlinear Schrödinger equation with localized dissipation. Computers and Mathematics with Applications, 96(??): 188–208, August 15, 2021. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122121001747.

Cai:2023:EMP

[CCG^{+}23] Qinghang Cai, Ronghua Chen, Kailun Guo, Wenxi Tian, Suizheng Qiu, and G. H. Su. An enhanced moving particle semi-implicit method for simulation of incompressible fluid flow and fluid-structure interaction. Computers and Mathematics with Applications, 145(??): 41–57, September 1, 2023. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122123002596.

Cao:2025:HGD

[CCGF25] Yangfan Cao, Hanshu Chen, Jakub Krzysztof Grabski, and Zhuojia Fu. A hybrid GFDM-DPIM solver for the stochastic response analysis of plate structures. Computers and Mathematics with Applications, 194(??):316– 334, September 15, 2025. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122125002639.

Caucao:2023:BSB

[CCGI23] Sergio Caucao, Eligio Colmenares, Gabriel N. Gatica, and Cristian Inzunza. A Banach spaces-based fully-mixed finite element method for the stationary chemotaxis-Navier–Stokes problem. Computers and Mathematics with Applications, 145(??):65–89, September 1, 2023. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0898122123002572.

Ciallella:2024:VHO

[CCGR24] Mirco Ciallella, Stephane Clain, Elena Gaburro, and Mario Ricchiuto. Very high order treatment of embedded curved boundaries in compressible flows: ADER discontinuous Galerkin with a space-time reconstruction for off-site data. Computers and Mathematics with Applications, 175(??):1–18, December 1, 2024. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S089812212400395X.

Contreras:2021:NLF

[CCGRL21] Fernando R. L. Contreras, Darlan K. E. Carvalho, Gustavo Galindez-Ramirez, and Paulo R. M. Lyra. A nonlinear finite volume method coupled with a modified higher order MUSCL-type method for the numerical simulation of two-phase flows in non-homogeneous and non-isotropic oil reservoirs. Computers and Mathematics with Applications, 92 (??):120–133, June 15, 2021. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122121000985.

Chuluunbaatar:2022:PTF

[CCGV22] G. Chuluunbaatar, O. Chuluunbaatar, A. A. Gusev, and S. I. Vinitsky. PItype fully symmetric quadrature rules on the 3-, …, 6simplexes. Computers and Mathematics with Applications, 124(??):89–97, October 15, 2022. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122122003364.

Cao:2024:FOI

[CCH24] Junfeng Cao, Ke Chen, and Huan Han. A fractionalorder image segmentation model with application to low-contrast and piecewise smooth images. Computers and Mathematics with Applications, 153(??):159– 171, January 1, 2024. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122123004996.

Cao:2022:EMF

[CCW22] Pei Cao, Jinru Chen, and Feng Wang. An extended mixed finite element method for elliptic interface problems. Computers and Mathematics with Applications, 113 (??):148–159, May 1, 2022. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122122000955.

Chu:2023:TLA

[CCWC23] Hanyu Chu, Ying Cai, Feng Wang, and Jinru Chen. A two-level additive Schwarz preconditioner for the Nitsche extended finite element approximation of elliptic interface problems. Computers and Mathematics with Applications, 147(??): 111–120, October 1, 2023. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122123003127.

Chen:2020:FDL

[CCWS20] Xinmeng Chen, Zhenhua Chai, Huili Wang, and Baochang Shi. A finitedifference lattice Boltzmann method with secondorder accuracy of time and space for incompressible flow. Computers and Mathematics with Applications, 80(12):3066–3081, December 15, 2020. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122120304247.

Castrillon-Candas:2021:SCA

[CCX21] Julio E. Castrillón-Candás and Jie Xu. A stochastic collocation approach for parabolic PDEs with random domain deformations. Computers and Mathematics with Applications, 93 (??):32–49, July 1, 2021. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122121001371.

Cai:2024:ALS

[CCY24] Zhiqiang Cai, Binghe Chen, and Jing Yang. Adaptive least-squares methods for convection-dominated diffusion–reaction problems. Computers and Mathematics with Applications, 173(??): 141–150, November 1, 2024. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122124003675.

Cavoretto:2020:TSA

[CD20] R. Cavoretto and A. De Rossi. A two-stage adaptive scheme based on RBF collocation for solving elliptic PDEs. Computers and Mathematics with Applications, 79(11):3206– 3222, June 1, 2020. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122120300341.

Cuomo:2023:SGF

[CDG^{+}23] Salvatore Cuomo, Mariapia De Rosa, Fabio Giampaolo, Stefano Izzo, and Vincenzo Schiano Di Cola. Solving groundwater flow equation using physics-informed neural networks. Computers and Mathematics with Applications, 145(??):106–123, September 1, 2023. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122123002602.

Ciarlet:2025:PEE

[CDGM25] Patrick Ciarlet, Minh Hieu Do, Mario Gervais, and François Madiot. A posteriori error estimates for the DD+ L^{2} jumps method on the neutron diffusion equations. Computers and Mathematics with Applications, 195(??):349–365, October 1, 2025. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122125003165.

Chen:2020:FPI

[CDLW20] Xu Chen, Deng Ding, Siu-Long Lei, and Wenfei Wang. A fast preconditioned iterative method for twodimensional options pricing under fractional differential models. Computers and Mathematics with Applications, 79(2):440–456, January 15, 2020. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122119303487.

Chen:2024:HTR

[CDM24] Paula Chen, Jérôme Darbon, and Tingwei Meng. Hopftype representation formulas and efficient algorithms for certain high-dimensional optimal control problems. Computers and Mathematics with Applications, 161 (??):90–120, May 1, 2024. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122124000828.

Corrochano:2024:HHO

[CDPL24] Adrián Corrochano, Giuseppe D’Alessio, Alessandro Parente, and Soledad Le Clainche. Hierarchical higher-order dynamic mode decomposition for clustering and feature selection. Computers and Mathematics with Applications, 158(??):36–45, March 15, 2024. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122124000038.

Chun:2024:MOE

[CEB24] Maria M. F. M. Chun, Briana L. Edwards, and Vladislav Bukshtynov. Multiscale optimization via enhanced multilevel PCAbased control space reduction for electrical impedance tomography imaging. Computers and Mathematics with Applications, 157(??):215– 234, March 1, 2024. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122124000075.

Chung:2020:MSC

[CEL20] Eric T. Chung, Yalchin Efendiev, and Wing Tat Leung. Multiscale stabilization for convection–diffusion equations with heterogeneous velocity and diffusion coefficients. Computers and Mathematics with Applications, 79(8):2336– 2349, April 15, 2020. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122119305334.

Cicci:2023:UQN

[CFG^{+}23] Ludovica Cicci, Stefania Fresca, Mengwu Guo, Andrea Manzoni, and Paolo Zunino. Uncertainty quantification for nonlinear solid mechanics using reduced order models with Gaussian process regression. Computers and Mathematics with Applications, 149(??):1–23, November 1, 2023. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122123003541.

Cai:2025:PFC

[CFG^{+}25] Dongting Cai, Boyi Fu, Renjun Gao, Xiangjie Kong, and Junxiang Yang. Phase-field computation for 3D shell reconstruction with an energystable and uniquely solvable BDF2 method. Computers and Mathematics with Applications, 189(??): 1–23, July 1, 2025. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122125001166.

Caruso:2022:ANU

[CFP22] N. Caruso and C. Fresneda-Portillo. On the analysis and numerics of united and segregated boundary-domain integral equation systems in 2D. Computers and Mathematics with Applications, 118 (??):46–55, July 15, 2022. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122122001973.

Chu:2023:MLA

[CFT^{+}23] Wen-Hui Chu, Zhuo-Jia Fu, Zhuo-Chao Tang, Wen-Zhi Xu, and Xiao-Ying Zhuang. A machine learning approach coupled with polar coordinate based localized collocation method for inner surface identification in heat conduction problem. Computers and Mathematics with Applications, 148(??):41–61, October 15, 2023. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122123003310.

Colera:2025:CIE

[CFTC25] Manuel Colera, Mario Freire-Torres, and Jaime Carpio. Comparison of implicit enthalpy methods for Stefan problems. Computers and Mathematics with Applications, 198(??):93–105, November 15, 2025. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122125003426.

Castillo:2021:UFH

[CG21] Paul Castillo and Sergio Gómez. A unified framework of high order structure-preserving Bsplines Galerkin methods for coupled nonlinear Schrödinger systems. Computers and Mathematics with Applications, 102(??):45–53, November 15, 2021. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122121003552.

Correa:2022:CDW

[CG22] Claudio I. Correa and Gabriel N. Gatica. On the continuous and discrete well-posedness of perturbed saddle-point formulations in Banach spaces. Computers and Mathematics with Applications, 117 (??):14–23, July 1, 2022. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122122001651.

Creff:2025:PGS

[CG25] Melvin Creff and Jean-Luc Guermond. Preconditioning of the generalized Stokes problem arising from the approximation of the time-dependent Navier– Stokes equations. Computers and Mathematics with Applications, 191(??):255– 274, August 1, 2025. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S089812212500183X.

Cao:2023:PIN

[CGGY23] Fujun Cao, Fei Gao, Xiaobin Guo, and Dongfang Yuan. Physics-informed neural networks with parameter asymptotic strategy for learning singularly perturbed convectiondominated problem. Computers and Mathematics with Applications, 150(??):229– 242, November 15, 2023. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122123004157.

Chen:2024:UES

[CGH24] Yanping Chen, Qiling Gu, and Jian Huang. Unconditionally energy stable ESAV– VEM schemes with variable time steps for the time fractional Allen–Cahn equation. Computers and Mathematics with Applications, 175(??): 276–288, December 1, 2024. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122124004450.

Chang:2024:SLP

[CGHJ24] Teng-Yuan Chang, Gung-Min Gie, Youngjoon Hong, and Chang-Yeol Jung. Singular layer physics informed neural network method for plane parallel flows. Computers and Mathematics with Applications, 166(??):91– 105, July 15, 2024. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122124001901.

Cuomo:2022:PIL

[CGI^{+}22] Salvatore Cuomo, Fabio Giampaolo, Stefano Izzo, Carlo Nitsch, Francesco Piccialli, and Cristina Trombetti. A physics-informed learning approach to Bernoullitype free boundary problems. Computers and Mathematics with Applications, 128(??): 34–43, December 15, 2022.

CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122122004266.

Chen:2024:USM

[CGK24] Meng Chen, Rong Gao, and Linghua Kong. An unconditionally stable modified leapfrog method for Maxwell’s equation in Kerrtype nonlinear media. Computers and Mathematics with Applications, 159(??):186– 201, April 1, 2024. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122124000531.

Certik:2020:HVV

[CGM^{+}20] O. Certík, F. Gardini, G. Manzini, L. Mascotto, and G. Vacca. The pand hp-versions of the virtual element method for elliptic eigenvalue problems. Computers and Mathematics with Applications, 79 (7):2035–2056, April 2020. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122119305073.

Camano:2022:ANM

[CGO22] Jessika Camaño, Carlos García, and Ricardo Oyarzúa. Analysis of a new mixed FEM for stationary incompressible magneto-hydrodynamics. Computers and Mathematics with Applications, 127(??): 65–79, December 1, 2022. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122122003972.

Chaurasiya:2024:NSN

[Cha24] Vikas Chaurasiya. Numerical simulation of a nonlinear sublimation process with temperature-dependent permeability and volumetric heat source: a phase change problem. Computers and Mathematics with Applications, 176(??):55–66, December 15, 2024. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122124004140.

Chakraborty:2024:AHA

[CHD24] Ankit Chakraborty, Stefan Henneking, and Leszek Demkowicz. An anisotropic hp-adaptation framework for ultraweak discontinuous Petrov–Galerkin formulations. Computers and Mathematics with Applications, 167(??):315–327, August 1, 2024. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S089812212400244X.

Chen:2024:CAI

[Che24] Yong Chen. Convergence analysis of an IMEX scheme for an integro-differential equation with inexact boundary arising in option pricing with stochastic intensity jumps. Computers and Mathematics with Applications, 161(??):63–77, May 1, 2024. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122124000853.

Chkifa:2021:CMM

[CHJR21] A. Chkifa, M. A. Hamadi, K. Jbilou, and A. Ratnani. A computational method for model reduction in index-2 dynamical systems for Stokes equations. Computers and Mathematics with Applications, 99(??): 171–181, October 1, 2021. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122121002935.

Chu:2021:STP

[CHL21] Van Tiep Chu, Viet Ha Hoang, and Roktaek Lim. Sparse tensor product high dimensional finite elements for two-scale mixed problems. Computers and Mathematics with Applications, 85 (??):42–56, March 1, 2021. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122120304855.

Chen:2022:PEE

[CHL22] Mingqing Chen, Jianguo Huang, and Sen Lin. A posteriori error estimation for a C^{1} virtual element method of Kirchhoff plates. Computers and Mathematics with Applications, 120(??): 132–150, August 15, 2022. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122122001808.

Cao:2021:ADD

[CHY21] Huihui Cao, Yunqing Huang, and Nianyu Yi. Adaptive direct discontinuous Galerkin method for elliptic equations. Computers and Mathematics with Applications, 97(??): 394–415, September 1, 2021. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122121002522.

Chen:2023:SFD

[CHZ23] Yuan Chen, Songming Hou, and Xu Zhang. Semi and fully discrete error analysis for elastodynamic interface problems using immersed finite element methods. Computers and Mathematics with Applications, 147 (??):92–110, October 1, 2023. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122123003073.

Ciarlet:2025:ECS

[CJ25] Patrick Ciarlet and Erell Jamelot. Explicit T - coercivity for the Stokes problem: a coercive finite element discretization. Computers and Mathematics with Applications, 188(??):137– 159, June 15, 2025. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122125001233.

Chakib:2023:NNA

[CK23] Abdelkrim Chakib and Ibrahim Khalil. A new numerical approach for solving shape optimization fourth-order spectral problems among convex domains. Computers and Mathematics with Applications, 149(??):171–189, November 1, 2023. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122123003784.

Choudhury:2024:NSE

[CKBK24] Sanuwar Ahmed Choudhury, N. Kishore Kumar, Pankaj Biswas, and Arbaz Khan. Nonconforming spectral element approximation for parabolic PDE with corner singularity. Computers and Mathematics with Applications, 167 (??):54–73, August 1, 2024. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122124001949.

Chen:2024:UHH

[CKBTS24] Jau-Uei Chen, Shinhoo Kang, Tan Bui-Thanh, and John N. Shadid. A unified hp-HDG framework for Friedrichs’ PDE systems. Computers and Mathematics with Applications, 154(??): 236–266, January 15, 2024. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122123005588.

Celaya:2024:SEP

[CKFR24] Adrian Celaya, Keegan Kirk, David Fuentes, and Beatrice Riviere. Solutions to elliptic and parabolic problems via finite difference based unsupervised small linear convolutional neural networks. Computers and Mathematics with Applications, 174(??): 31–42, November 15, 2024. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122124003663.

Chen:2025:HMM

[CKL25] C. S. Chen, Andreas Karageorghis, and Q. G. Liu. Hybrid meshless method for solving inhomogeneous polyharmonic equations. Computers and Mathematics with Applications, 196(??):218– 232, October 15, 2025. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122125003116.

Choi:2022:CIG

[CKS22] Hayoung Choi, Sang Dong Kim, and Byeong-Chun Shin. Choice of an initial guess for Newton’s method to solve nonlinear differential equations. Computers and Mathematics with Applications, 117(??):69– 73, July 1, 2022. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122122001699.

Cheng:2020:RUL

[CL20] Guo Cheng and Ji-Cheng Li. A relaxed upper and lower triangular splitting preconditioner for the linearized Navier–Stokes equation. Computers and Mathematics with Applications, 80(1):43–60, July 1, 2020. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122120301012.

Chou:2021:FVA

[CL21] Lot-Kei Chou and Siu-Long Lei. Finite volume approximation with ADI scheme and low-rank solver for high dimensional spatial distributed-order fractional diffusion equations. Computers and Mathematics with Applications, 89 (??):116–126, May 1, 2021. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122121000626.

Chen:2022:FDS

[CL22a] Li Chen and Shujuan Lü. A fully discrete spectral scheme for time fractional Cahn– Hilliard equation with initial singularity. Computers and Mathematics with Applications, 127(??):213–224, December 1, 2022. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122122004382.

Corrochano:2022:SSN

[CL22b] Adrián Corrochano and Soledad Le Clainche. Structural sensitivity in non-linear flows using direct solutions. Computers and Mathematics with Applications, 128(??):

69–78, December 15, 2022. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122122004291.

Campos:2024:NSF

[CL24a] Lucas Silveira Campos and Carlos Friedrich Loeffler. A new strategy for a faster matrix assembly in the boundary element method. Computers and Mathematics with Applications, 174(??):449– 453, November 15, 2024. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122124004449.

Chernov:2024:AGC

[CL24b] Alexey Chernov and Tùng Lê. Analytic and Gevrey class regularity for parametric semilinear reactiondiffusion problems and applications in uncertainty quantification. Computers and Mathematics with Applications, 164(??):116– 130, June 15, 2024. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122124001652.

Chen:2021:TGM

[CLH21] Jinyue Chen, Yixin Li, and Xianliang Hu. A twogrid method for the phasefield model of photonic band gap optimization. Computers and Mathematics with Applications, 96(??):44–54, August 15, 2021. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122121001838.

Chen:2022:TGF

[CLH22] Chuanjun Chen, Yuzhi Lou, and Hanzhang Hu. Twogrid finite volume element method for the timedependent Schrödinger equation. Computers and Mathematics with Applications, 108(??):185–195, February 15, 2022. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122122000086.

Chai:2021:FOC

[CLL21] Li Chai, Yang Liu, and Hong Li. Fourth-order compact difference schemes for the two-dimensional nonlinear fractional mobile/immobile transport models. Computers and Mathematics with Applications, 100(??):1–10, October 15, 2021. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122121003199.

Cremonesi:2024:ASF

[CLL^{+}24]  Massimiliano  Cremonesi,

Andrea Lamperti, Carlo Lovadina, Umberto Perego, and Alessandro Russo. Analysis of a stabilization-free quadrilateral Virtual Element for 2D linear elasticity in the Hu–Washizu formulation. Computers and Mathematics with Applications, 155(??): 142–149, February 1, 2024. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122123005503.

Chai:2023:FTM

[CLLG23] Li Chai, Yang Liu, Hong Li, and Wei Gao. Fast TT-M fourth-order compact difference schemes for a twodimensional space fractional Gray–Scott model. Computers and Mathematics with Applications, 141(??):191– 206, July 1, 2023. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122123001827.

Cao:2020:RSP

[CLLL20] Wenjin Cao, Zhe Li, Xuhui Li, and David Le Touzé. A regularized single-phase lattice Boltzmann method for free-surface flows. Computers and Mathematics with Applications, 80(10):2194– 2211, November 15, 2020.

CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122120303710.

Cai:2025:EID

[CLLL25] Mingchao Cai, Jingzhi Li, Ziliang Li, and Qiang Liu. An efficient iterative decoupling method for thermoporoelasticity based on a four-field formulation. Computers and Mathematics with Applications, 195(??):139– 160, October 1, 2025. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122125003128.

Chen:2025:OCA

[CLQL25] Yanping Chen, Wanxiang Liu, Fangfang Qin, and Qin Liang. Optimal convergence analysis of an energy dissipation property virtual element method for the nonlinear Benjamin–Bona– Mahony–Burgers equation. Computers and Mathematics with Applications, 192 (??):37–53, August 15, 2025. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122125001932.

Cesmelioglu:2023:HDG

[CLR23] Aycil Cesmelioglu, Jeonghun J.

Lee, and Sander Rhebergen.

Hybridizable discontinuous Galerkin methods for the coupled Stokes–Biot problem. Computers and Mathematics with Applications, 144 (??):12–33, August 15, 2023. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122123002286.

Cesmelioglu:2024:HDG

[CLRT24] Aycil Cesmelioglu, Jeonghun J. Lee, Sander Rhebergen, and Dorisa Tabaku. A hybridizable discontinuous Galerkin method for the dual-porosity-Stokes problem. Computers and Mathematics with Applications, 165 (??):180–195, July 1, 2024. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122124001627.

Chen:2024:GSI

[CLW^{+}24] Jianfeng Chen, Sha Liu, Yong Wang, Congshan Zhuo, Yanguang Yang, and Chengwen Zhong. A gassurface interaction algorithm for discrete velocity methods in predicting rarefied and multi-scale flows: For Maxwell boundary model. Computers and Mathematics with Applications, 175(??): 570–601, December 1, 2024. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122124004838.

Chen:2025:AMP

[CLXZ25] Tianhao Chen, Zeyu Li, Pengbo Xu, and Haibiao Zheng. Asynchronous multimodal PINN pre-train framework based on TransVNet(MPP-TV) and its application in numerical solutions of the Cauchy problem for the Hamilton–Jacobi equation. Computers and Mathematics with Applications, 187 (??):203–230, June 1, 2025. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122125001221.

Chen:2025:DDS

[CLY25a] Jiaxin Chen, Biao Li, and Manwai Yuen. Data-driven SFB solutions and parameters discovery for nonlinear Schrödinger equation via time domain decomposition physics-informed neural networks. Computers and Mathematics with Applications, 199(??):45–63, December 1, 2025. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122125003724.

Chen:2025:PSE

[CLY25b] Meng Chen, Leevan Ling, and Dongfang Yun. Proving the stability estimates of variational least-squares kernel-based methods. Computers and Mathematics with Applications, 180(??):46–60, February 15, 2025. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122124005534.

Chen:2022:IFE

[CLYH22] Yanping Chen, Qingfeng Li, Huaming Yi, and Yunqing Huang. Immersed finite element method for time fractional diffusion problems with discontinuous coefficients. Computers and Mathematics with Applications, 128(??):121–129, December 15, 2022. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122122004011.

Cao:2023:PAR

[CLZ23a] Dingding Cao, Xinxiang Li, and Huiqing Zhu. A polynomial-augmented RBF collocation method for fourthorder boundary value problems. Computers and Mathematics with Applications, 133 (??):1–11, March 1, 2023. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122122005211.

Chen:2023:FVE

[CLZ23b] Guofang Chen, Junliang Lv, and Xinye Zhang. Finite volume element method for nonlinear elliptic equations on quadrilateral meshes. Computers and Mathematics with Applications, 140(??):154– 168, June 15, 2023. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122123001463.

Cao:2024:MSN

[CLZ^{+}24] Junzhe Cao, Sha Liu, Chengwen Zhong, Congshan Zhuo, and Kun Xu. Multiple solutions of nonlinear coupled constitutive relation model and its rectification in non-equilibrium flow computation. Computers and Mathematics with Applications, 174(??):1–17, November 15, 2024. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122124003705.

Chen:2024:MVE

[CLZN24] Yanli Chen, Xin Liu, Wenhui Zhang, and Yufeng Nie. Mixed virtual element methods for the poroelastodynamics model on polygonal grids. Computers and Mathematics with Applications, 174(??):431– 448, November 15, 2024. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122124004334.

Chen:2020:TGM

[CLZW20] Chuanjun Chen, Huan Liu, Xiangcheng Zheng, and Hong Wang. A twogrid MMOC finite element method for nonlinear variable-order timefractional mobile/immobile advection–diffusion equations. Computers and Mathematics with Applications, 79 (9):2771–2783, May 1, 2020. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122119305668.

Claeys:2020:BIM

[CM20a] X. Claeys and P. Marchand. Boundary integral multi-trace formulations and Optimised Schwarz Methods. Computers and Mathematics with Applications, 79(11): 3241–3256, June 1, 2020. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122120300328.

Cugu:2020:PMS

[ÇM20b] Ilke Çugu and Murat Manguoglu. A parallel multithreaded sparse triangular linear system solver. Computers and Mathematics with Applications, 80(2): 371–385, July 2020. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122119304602.

Carstensen:2021:CMA

[CM21] Carsten Carstensen and Rui Ma. Collective marking for arbitrary order adaptive least-squares finite element methods with optimal rates. Computers and Mathematics with Applications, 95(??): 271–281, August 1, 2021. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122120304673.

Chaudhary:2022:MFG

[CM22] Sudhakar Chaudhary and Jitesh P. Mandaliya. Meshfree Galerkin approximation for parabolic nonlocal problem using web-splines. Computers and Mathematics with Applications, 128(??):180– 187, December 15, 2022. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S089812212200431X.

Chakraborty:2023:CHM

[CM23] Ankit Chakraborty and Georg May. A continuous hp-mesh model for discontinuous Petrov–Galerkin finite element schemes with optimal test functions. Computers and Mathematics with Applications, 147(??):64–75, October 1, 2023. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122123003115.

Chamoun:2024:MAA

[CM24] Georges Chamoun and Nahia Mourad. Mathematical analysis and asymptotic predictions of chemical-driven swimming living organisms in weighted networks. Computers and Mathematics with Applications, 172(??):16–37, October 15, 2024. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122124003377.

Chaari:2025:NSP

[CML^{+}25] Anis Chaari, Waleed Mouhali, Mohammed Louaked, Nacer Sellila, and Houari Mechkour. Numerical simulation of pollutant concentration patterns of a twodimensional congestion traffic. Computers and Mathematics with Applications, 188 (??):97–114, June 15, 2025.

CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122125001154.

Clain:2023:CST

[CMM23] Stéphane Clain, Gaspar J.

Machado, and M. T. Malheiro. Compact schemes in time with applications to partial differential equations. Computers and Mathematics with Applications, 140 (??):107–125, June 15, 2023. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122123001062.

Calatroni:2023:FSS

[CMPR23] L. Calatroni, S. Morigi, S. Parisotto, and G. A. Recupero. Fast and stable schemes for non-linear osmosis filtering. Computers and Mathematics with Applications, 133(??):30– 47, March 1, 2023. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122122005223.

Cheng:2022:LDG

[CMR22] Yao Cheng, Yanjie Mei, and Hans-Görg Roos. The local discontinuous Galerkin method on layer-adapted meshes for time-dependent singularly perturbed convection– diffusion problems. Computers and Mathematics with Applications, 117(??):245– 256, July 1, 2022. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122122001912.

Charles:2023:MNS

[CMS23] Frédérique Charles, Annamaria Massimini, and Francesco Salvarani. Mathematical and numerical study of a kinetic model describing the evolution of planetary rings. Computers and Mathematics with Applications, 143 (??):48–56, August 1, 2023. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122123001724.

Chen:2023:EAS

[CMY23] Xiangling Chen, Lina Ma, and Xiaofeng Yang. Error analysis of second-order IEQ numerical schemes for the viscous Cahn–Hilliard equation with hyperbolic relaxation. Computers and Mathematics with Applications, 152(??):112–128, December 15, 2023. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122123004443.

Creuse:2023:GOE

[CNT23] Emmanuel Creusé, Serge Nicaise, and Zuqi Tang. Goal-oriented error estimation based on equilibrated flux and potential reconstruction for the approximation of elliptic and parabolic problems. Computers and Mathematics with Applications, 146(??):323– 338, September 15, 2023. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122123003000.

Chai:2020:ASG

[CO20] Yong Chai and Jie Ouyang.

Appropriate stabilized Galerkin approaches for solving twodimensional coupled Burgers’ equations at high Reynolds numbers. Computers and Mathematics with Applications, 79(5):1287–1301, March 1, 2020. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122119304407.

Cecot:2021:DPG

[CO21] Witold Cecot and Marta Oleksy. The discontinuous Petrov–Galerkin methodology for the mixed multiscale finite element method. Computers and Mathematics with Applications, 95 (??):28–40, August 1, 2021.

CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122120303692.

Costabel:2023:CSD

[Cos23] Martin Costabel. Convergence of a simple discretization of the finite Hilbert transformation. Computers and Mathematics with Applications, 150(??):15–21, November 15, 2023. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122123003899.

Cuesta:2021:HRA

[CP21] Eduardo Cuesta and Rodrigo Ponce. Hölder regularity for abstract semilinear fractional differential equations in Banach spaces. Computers and Mathematics with Applications, 85 (??):57–68, March 1, 2021. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122121000201.

Costa:2021:GAC

[CPBF21] Pedro Costa, Everett Phillips, Luca Brandt, and Massimiliano Fatica. GPU acceleration of CaNS for massively-parallel direct numerical simulations of canonical fluid flows. Computers and Mathematics with Applications, 81(??):502–511, January 1, 2021. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122120300092.

Ciegis:2020:ASP

[ČPPŠ20] R. Čiegis, G. Panasenko, K. Pileckas, and V. Šumskas. ADI scheme for partially dimension reduced heat conduction models. Computers and Mathematics with Applications, 80(5):1275–1286, September 1, 2020. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122120302613.

Cho:2021:OAS

[CPS21] D. Cho, L. F. Pavarino, and S. Scacchi. Overlapping additive Schwarz preconditioners for isogeometric collocation discretizations of linear elasticity. Computers and Mathematics with Applications, 93(??):66–77, July 1, 2021. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122121001395.

Chen:2024:SOA

[CQG^{+}24] Lele Chen, Yi Qin, Xinyue Gao, Yang Wang, Yi Li, and Jian Li. A second-order adaptive DLN algorithm with different subdomain variable time steps for the 3D closed-loop geothermal system. Computers and Mathematics with Applications, 165(??):1–18, July 1, 2024. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122124001494.

Crozon:2025:PPD

[CQS25] Thomas Crozon, El Houssaine Quenjel, and Mazen Saad. Positivity-preserving DDFV scheme for compressible two-phase flow in porous media. Computers and Mathematics with Applications, 194(??):110– 134, September 15, 2025. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122125002524.

Canelas:2024:MII

[CRB24] Alfredo Canelas, Jean R. Roche, and Jean P. Brancher. Melting interfaces in induction heated bodies. Computers and Mathematics with Applications, 153(??):213– 224, January 1, 2024. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122123005266.

Casper:2020:CST

[CRDS20] Thorben Casper, Ulrich Römer, Herbert De Gersem, and Sebastian Schöps. Coupled simulation of transient heat flow and electric currents in thin wires: Application to bond wires in microelectronic chip packaging. Computers and Mathematics with Applications, 79(6): 1781–1801, March 15, 2020. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122119304985.

Correa:2020:HHO

[CRF^{+}20] Maicon R. Correa, Juan C.

Rodriguez, Agnaldo M. Farias, Denise de Siqueira, and Philippe R. B. Devloo. Hierarchical high order finite element spaces in H(div, Ω) × H^{1}(Ω) for a stabilized mixed formulation of Darcy problem. Computers and Mathematics with Applications, 80(5):1117–1141, September 1, 2020. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122120302522.

Cao:2025:NDO

[CRL^{+}25] Xiran Cao, Zhengze Rong, Ping Lin, Liancun Zheng, and Xuelan Zhang. A novel distributed-order time fractional derivative model of laser-induced thermal therapy for deep-lying tumor. Computers and Mathematics with Applications, 184 (??):107–115, April 15, 2025. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122125000689.

Chakraborty:2022:AAS

[CRM22] Ankit Chakraborty, Ajay Mandyam Rangarajan, and Georg May. An anisotropic h-adaptive strategy for discontinuous Petrov–Galerkin schemes using a continuous mesh model. Computers and Mathematics with Applications, 106 (??):1–17, January 15, 2022. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122121004272.

Coclite:2020:LBD

[CRPdT20] Alessandro Coclite, Sergio Ranaldo, Giuseppe Pascazio, and Marco D. de Tullio. A lattice Boltzmann dynamicimmersed boundary scheme for the transport of deformable inertial capsules in low-re flows. Computers and Mathematics with Applications, 80(12):2860– 2876, December 15, 2020. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122120303795.

Chandrashekar:2021:NSB

[CRRS21] Praveen Chandrashekar, Mythily Ramaswamy, Jean-Pierre Raymond, and Ruchi Sandilya. Numerical stabilization of the Boussinesq system using boundary feedback control. Computers and Mathematics with Applications, 89 (??):163–183, May 1, 2021. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122121000699.

Chen:2020:PDR

[CS20a] Dongqin Chen and Huailing Song. The POD– DEIM reduced-order method for stochastic Allen–Cahn equations with multiplicative noise. Computers and Mathematics with Applications, 80(12):2691–2706, December 15, 2020. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122120303825.

Chen:2020:DCP

[CS20b] Hu Chen and Martin Stynes.

A discrete comparison principle for the time-fractional diffusion equation. Computers and Mathematics with Applications, 80(5):917–922, September 1, 2020. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122120301656.

Carstensen:2022:CFI

[CS22] C. Carstensen and S. Sauter. Critical functions and infsup stability of Crouzeix– Raviart elements. Computers and Mathematics with Applications, 108(??):12–23, February 15, 2022. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122121004442.

Chen:2023:SFV

[CS23a] Alvin Chen and N. Sukumar. Stabilization-free virtual element method for plane elasticity. Computers and Mathematics with Applications, 138(??):88– 105, May 15, 2023. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122123000913.

Chen:2023:NAC

[CS23b] Zhan Chen and Yuanzhen Shao. A new approach to constrained total variation solvation models and the study of solute-solvent interface profiles. Computers and Mathematics with Applications, 130(??):119–136, January 15, 2023. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S089812212200503X.

Clavero:2025:ENM

[CS25] Carmelo Clavero and Ram Shiromani. An efficient numerical method for 2D elliptic singularly perturbed systems with different magnitude parameters in the diffusion and the convection terms. Computers and Mathematics with Applications, 181(??):287–322, March 1, 2025. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122125000112.

Carere:2021:WPR

[CSB^{+}21] Giuseppe Carere, Maria Strazzullo, Francesco Ballarin, Gianluigi Rozza, and Rob Stevenson. A weighted POD-reduction approach for parametrized PDE-constrained optimal control problems with random inputs and applications to environmental sciences. Computers and Mathematics with Applications, 102(??):261–276, November 15, 2021. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122121003746.

Chen:2024:WSF

[CSCY24] Zixuan Chen, Xiaoliang Song, Xiaotong Chen, and Bo Yu. A warm-start FEdABCD algorithm for elliptic optimal control problems with constraints on the control and the gradient of the state. Computers and Mathematics with Applications, 161(??):1–12, May 1, 2024. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S089812212400083X.

Chizari:2021:CVE

[CSI21] Hossain Chizari, Vishal Singh, and Farzad Ismail. Cell-vertex entropy-stable finite volume methods for the system of Euler equations on unstructured grids. Computers and Mathematics with Applications, 98(??):261– 279, September 15, 2021. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122121002753.

Chizari:2023:LPM

[CSI23] Hossain Chizari, Vishal Singh, and Farzad Ismail. Linearity preserving modified LDA methods for unsteady advectiondiffusion problems. Computers and Mathematics with Applications, 146(??):12–21, September 15, 2023. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122123002791.

Cao:2020:RDP

[CSS20] Yang Cao, Zhen-Quan Shi, and Quan Shi. Regularized DPSS preconditioners for generalized saddle point linear systems. Computers and Mathematics with Applications, 80(5):956–972, September 1, 2020. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122120302340.

Clavero:2023:NSS

[CSS23] Carmelo Clavero, Ram Shiromani, and Vembu Shanthi. Numerical solution of singularly perturbed 2-D convection-diffusion elliptic interface PDEs with Robintype boundary conditions. Computers and Mathematics with Applications, 140 (??):1–16, June 15, 2023. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122123001050.

Chu:2025:SAD

[CSS25] Xiaochen Chu, Xiangyu Shi, and Dongyang Shi. Superconvergence analysis of the decoupled and linearized mixed finite element methods for unsteady incompressible MHD equations. Computers and Mathematics with Applications, 188 (??):160–182, June 15, 2025. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122125001282.

Caban:2022:HOC

[CT22] Lena Caban and Artur Tyliszczak. High-order compact difference schemes on wide computational stencils with a spectral-like accuracy. Computers and Mathematics with Applications, 108(??): 123–140, February 15, 2022. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122122000062.

Carraturo:2024:IAC

[CTGR24] Massimo Carraturo, Michele Torre, Carlotta Giannelli, and Alessandro Reali. An isogeometric approach to coupled thermomechanics in 3D via hierarchical adaptivity. Computers and Mathematics with Applications, 162 (??):133–144, May 15, 2024. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122124001020.

Cui:2022:NFD

[CTZ22] Shengzhu Cui, Zhanjing Tao, and Jun Zhu. New finite difference unequal-sized Hermite WENO scheme for Navier–Stokes equations. Computers and Mathematics with Applications, 128(??): 273–284, December 15, 2022. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122122004400.

Ciegis:2020:TLS

[CV20] Raimondas Ciegis and Petr N.

Vabishchevich. Two-level schemes of Cauchy problem method for solving fractional powers of elliptic operators. Computers and Mathematics with Applications, 80(2):305–315, July 2020. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122119304092.

Chen:2020:VMM

[CW20] Xi Chen and David M.

Williams. Versatile mixed methods for the incompressible Navier–Stokes equations. Computers and Mathematics with Applications, 80(6):1555–1577, September 15, 2020. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122120303059.

Chen:2022:PCS

[CWC22] Jing Chen, Feng Wang, and Huanzhen Chen. Probabilityconservative simulation for Lévy financial model by a mixed finite element method. Computers and Mathematics with Applications, 106(??): 92–105, January 15, 2022. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122121004405.

Cheng:2022:SBM

[CWLQ22] Suifu Cheng, Fajie Wang, Po-Wei Li, and Wenzhen Qu. Singular boundary method for 2D and 3D acoustic design sensitivity analysis. Computers and Mathematics with Applications, 119(??): 371–386, August 1, 2022. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S089812212200253X.

Chen:2020:ICA

[CWTM20] Pinlei Chen, Ignasius P. A. Wijaya, Ian Tuttle, and Arif Masud. Interfacial coupling across incompatible meshes in a monolithic finite-strain thermomechanical formulation. Computers and Mathematics with Applications, 79(11):3068– 3091, June 1, 2020. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122120300213.

Cao:2022:NNM

[CWW22] Shuhao Cao, Chunmei Wang, and Junping Wang. A new numerical method for divcurl systems with low regularity assumptions. Computers and Mathematics with Applications, 114(??):47– 59, May 15, 2022. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122122000992.

Cui:2022:ENN

[CWX22] Tao Cui, Ziming Wang, and Xueshuang Xiang. An efficient neural network method with plane wave activation functions for solving Helmholtz equation. Computers and Mathematics with Applications, 111(??):34– 49, April 1, 2022. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122122000542.

Cao:2025:AFS

[CWX25] Jiliang Cao, Wansheng Wang, and Aiguo Xiao.

Adaptive fast L1–2 scheme for solving time fractional parabolic problems. Computers and Mathematics with Applications, 179(??):59–76, February 1, 2025. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122124005406.

Cai:2021:RFQ

[CWZ21] Deng’an Cai, Xinwei Wang, and Guangming Zhou. A rotation-free quadrature element formulation for free vibration analysis of thin sectorial plates with arbitrary boundary supports. Computers and Mathematics with Applications, 99 (??):84–98, October 1, 2021. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122121002923.

Chen:2024:MIF

[CXF24] Jiaqi Chen, Xufeng Xiao, and Xinlong Feng. A mixed immersed finite element method for fourthorder interface problems on surfaces. Computers and Mathematics with Applications, 176(??):122–139, December 15, 2024. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S089812212400422X.

Chen:2025:TGD

[CXJ25] Shuangshuang Chen, Yuna Xu, and Longchao Jin. The two-grid decoupled method for hybrid-dimensional fracture models based on the discontinuous Galerkin method. Computers and Mathematics with Applications, 195(??): 296–321, October 1, 2025. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122125003189.

Cheng:2025:DES

[CXX^{+}25] Jingjie Cheng, Qing Xia, Binhu Xia, Junseok Kim, and Yibao Li. Decoupled, efficient and structurepreserving numerical scheme for a non-isothermal phase field sintering model. Computers and Mathematics with Applications, 196(??):49–63, October 15, 2025. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122125002998.

Chen:2023:LFS

[CXXZ23] Gang Chen, Xiaoping Xie, Youcai Xu, and Yangwen Zhang. A locking-free stabilized embedded discontinuous Galerkin method for linear elasticity with strong symmetric stress and continuous displacement trace approximation. Computers and Mathematics with Applications, 134(??):66–86, March 15, 2023. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122122005259.

Cai:2023:OEE

[CY23] Zhiqiang Cai and Jing Yang. Optimal error estimate of discontinuous Galerkin methods for advection– diffusion–reaction problems with low regularity. Computers and Mathematics with Applications, 140(??):282– 290, June 15, 2023. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122122004618.

Chen:2023:RLB

[CYC^{+}23] Jiabao Chen, Dangguo Yang, Qing Chen, Jianhong Sun, and Yan Wang. A rotated lattice Boltzmann flux solver with improved stability for the simulation of compressible flows with intense shock waves at high Mach number. Computers and Mathematics with Applications, 132(??): 18–31, February 15, 2023. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122122005041.

Cho:2023:OSM

[CYKL23] Junhyun Cho, Donghee Yang, Yejin Kim, and Sungchul Lee. An operator splitting method for multi-asset options with the Feynman–Kac formula. Computers and Mathematics with Applications, 135 (??):93–101, April 1, 2023. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122123000196.

Chen:2025:FSO

[CYM25] Yiru Chen, Yun-Bo Yang, and Lijie Mei. A family of second-order time stepping methods for a nonlinear fluid-fluid interaction model. Computers and Mathematics with Applications, 182 (??):1–23, March 15, 2025. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122125000100.

Cheng:2021:OEE

[CYQW21] Xiujun Cheng, Xiaoqiang Yan, Hongyu Qin, and Huiru Wang. Optimal l ^{∞} error estimates of the conservative scheme for two-dimensional Schrödinger equations with wave operator. Computers and Mathematics with Applications, 100(??):74–82, October 15, 2021. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122121003187.

Chen:2021:CLB

[CYY^{+}21] Xinyang Chen, Gengchao Yang, Qinghe Yao, Zisen Nie, and Zichao Jiang. A compressed lattice Boltzmann method based on ConvLSTM and ResNet. Computers and Mathematics with Applications, 97(??):162–174, September 1, 2021. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122121002418.

Cheng:2020:MQB

[CZ20] Wei Cheng and Qi Zhao. A modified quasi-boundary value method for a twodimensional inverse heat conduction problem. Computers and Mathematics with Applications, 79(2):293–302, January 15, 2020. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122119303347.

Chen:2023:USF

[CZ23] Chuanjun Chen and Tong Zhang. Unconditional stability of first and second orders implicit/explicit schemes for the natural convection equations. Computers and Mathematics with Applications, 139 (??):152–172, June 1, 2023. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122122002668.

Chen:2025:CNA

[CZC^{+}25] Qiyu Chen, Lance Zhao, Susan Chen, Ke Ge, Zhe Lyu, and Zhe Li. A comprehensive numerical approach for solving heterogeneous electrochemicalthermal model of lithiumion battery. Computers and Mathematics with Applications, 197(??):88–111, November 1, 2025. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122125003360.

Chen:2025:NMM

[CZK^{+}25] Meng Chen, Juan Zhang, Linghua Kong, Peng Zhang, and Bin He. A newly modified method for 3D Maxwell’s equation. Computers and Mathematics with Applications, 193(??):77– 102, September 1, 2025. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122125002299.

Chen:2022:WBR

[CZLQ22] Ziming Chen, Yingjuan Zhang, Gang Li, and Shouguo Qian. A wellbalanced Runge–Kutta discontinuous Galerkin method for the Euler equations in isothermal hydrostatic state under gravitational field. Computers and Mathematics with Applications, 119(??): 340–350, August 1, 2022. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122122002176.

Zhang:2023:GVV

[cZsCyC23] Te chao Zhang, Xiao shan Cao, and Si yuan Chen. Gradient viscoelastic virtual boundary for numerical simulation of wave propagation. Computers and Mathematics with Applications, 147(??): 202–209, October 1, 2023. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122123003218.

Dawood:2025:SCC

[DAAEA25] A. S. Dawood, A. S. Amer, R. M. Abumandour, and W. A. El-Askary. Simulation of cartesian cut-cell technique for modeling turbulent flow in asymmetric diffusers using various turbulence models. Computers and Mathematics with Applications, 182(??):84–101, March 15, 2025. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S089812212500015X.

deAraujo:2025:ANA

[dABCC25] Anderson L. A. de Araujo, Jose L. Boldrini, Bianca M. R. Calsavara, and Maicon R. Correa. Analysis and numerical approximation of a mathematical model for Aedes aegypti populations. Computers and Mathematics with Applications, 180(??):214–241, February 15, 2025. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122125000021.

Dai:2022:CNN

[Dai22] Wanyang Dai. Convolutional neural network based simulation and analysis for backward stochastic partial differential equations. Computers and Mathematics with Applications, 119 (??):21–58, August 1, 2022. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122122002115.

Du:2020:TSO

[DAS20] Ruilian Du, Anatoly A. Alikhanov, and Zhi-Zhong Sun. Temporal second order difference schemes for the multi-dimensional variableorder time fractional subdiffusion equations. Computers and Mathematics with Applications, 79(10):2952– 2972, May 15, 2020. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122120300109.

Dyniewicz:2023:ESS

[DBB23] Bartłomiej Dyniewicz, Jacek M. Bajkowski, and Czesław I. Bajer. Efficient strategy for space–time based finite element analysis of vibrating structures. Computers and Mathematics with Applications, 148(??):70–80, October 15, 2023. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122123003346.

DaqiqShirazi:2021:TST

[DBH^{+}21] Mohammadreza DaqiqShirazi, Azeez A. Barzinjy, Samir M. Hamad, Rezvan Alamian, and Mostafa Safdari Shadloo. A transient study on two phase adiabatic flow over micro circular pin heat sinks. Computers and Mathematics with Applications, 81(??):811–822, January 1, 2021. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122119305085.

Dolejsi:2021:GOM

[DBR21] Vít Dolejší, Ondřej Bartoš, and Filip Roskovec. Goaloriented mesh adaptation method for nonlinear problems including algebraic errors. Computers and Mathematics with Applications, 93 (??):178–198, July 1, 2021. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S089812212100136X.

Du:2022:AMN

[DC22] Hong Du and Zhong Chen.

Adaptive meshless numerical method of solving 2D variable order time fractional mobile–immobile advection– diffusion equations. Computers and Mathematics with Applications, 124(??):42–51, October 15, 2022. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122122003236.

Diaz:2023:NRF

[DCCM23] Felipe A. Díaz, Ernesto Castillo, Roberto C. Cabrales, and Nelson O. Moraga. Nonrelaxed finite volume fractional step schemes for unsteady incompressible flows. Computers and Mathematics with Applications, 146 (??):241–252, September 15, 2023. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122123002961.

de-Cordoba:2020:ECD

[dCG20] Gonzalo F. de Córdoba and Gonzalo Galiano. An economic cross-diffusion mutualistic model for cities emergence. Computers and Mathematics with Applications, 79(3):643–655, February 1, 2020. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S089812211930361X.

deCastro:2025:LMM

[dCLW25] Amy de Castro, Hyesuk Lee, and Margaret M. Wiecek. A Lagrange multiplier method for fluid-structure interaction: Well-posedness and domain decomposition. Computers and Mathematics with Applications, 181(??):193– 215, March 1, 2025. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122124005716.

Diksha:2025:ITT

[DCSAM25] Diksha, Soniya Chaudhary, Pawan Kumar Sharma, and Qasem M. Al-Mdallal. Integral transform technique for determining stress intensity factor in wave propagation through functionally graded piezoelectricviscoelastic structure. Computers and Mathematics with Applications, 186(??):130– 154, May 15, 2025. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122125001178.

Deka:2020:CFE

[DD20] Bhupen Deka and Jogen Dutta. Convergence of finite element methods for hyperbolic heat conduction model with an interface. Computers and Mathematics with Applications, 79(11): 3139–3159, June 1, 2020. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122120300298.

Dabaghi:2021:UFH

[DD21] Jad Dabaghi and Guillaume Delay. A unified framework for high-order numerical discretizations of variational inequalities. Computers and Mathematics with Applications, 92(??):62– 75, June 15, 2021. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122121000857.

DiPietro:2022:DRM

[DD22] Daniele A. Di Pietro and Jérôme Droniou. A discrete de Rham method for the Reissner–Mindlin plate bending problem on polygonal meshes. Computers and Mathematics with Applications, 125(??):136–149, November 1, 2022. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122122003650.

Dassi:2024:EVE

[DD24] F. Dassi and P. Di Barba. Enriched Virtual Element space on curved meshes with an application in magnetics. Computers and Mathematics with Applications, 161(??):43–50, May 1, 2024. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122124000816.

DAmbra:2023:ACA

[DDFZ23] Pasqua D’Ambra, Fabio Durastante, Salvatore Filippone, and Ludmil Zikatanov. Automatic coarsening in algebraic multigrid utilizing quality measures for matching-based aggregations. Computers and Mathematics with Applications, 144(??):290–305, August 15, 2023. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S089812212300278X.

Devloo:2023:ESC

[DDG23] Philippe R. B. Devloo, Omar Durán, and Sônia M. Gomes. Exact sequences of conforming finite element spaces with interface constraints for macro polytopal meshes. Computers and Mathematics with Applications, 134(??): 124–139, March 15, 2023. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122123000056.

DellAccio:2021:SPE

[DDNS21] Francesco Dell’Accio, Filomena Di Tommaso, Otheman Nouisser, and Najoua Siar. Solving Poisson equation with Dirichlet conditions through multinode Shepard operators. Computers and Mathematics with Applications, 98(??):254–260, September 15, 2021. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122121002844.

Dubois:2021:HPC

[DdSP21] Remy Dubois, Eric Goncalves da Silva, and Philippe Parnaudeau. High performance computing of stiff bubble collapse on CPU-GPU heterogeneous platform. Computers and Mathematics with Applications, 99(??):246– 256, October 1, 2021. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S089812212100273X.

Deng:2021:SRS

[DE21] Quanling Deng and Alexandre Ern. SoftFEM: Revisiting the spectral finite element approximation of second-order elliptic operators. Computers and Mathematics with Applications, 101(??):119–133, November 1, 2021. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122121003382.

Deka:2022:EAS

[DE22] Pranab Jyoti Deka and Lukas Einkemmer. Efficient adaptive step size control for exponential integrators. Computers and Mathematics with Applications, 123 (??):59–74, October 1, 2022. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122122002954.

Derakhshan:2023:SAD

[Der23] Mohammad Hossein Derakhshan. Stability analysis of difference-Legendre spectral method for two-dimensional Riesz space distributed-order diffusion-wave model. Computers and Mathematics with Applications, 144(??):150– 163, August 15, 2023. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122123002456.

Dong:2025:BPS

[DEW25] Zhaonan Dong, Alexandre Ern, and Zuodong Wang. A bound-preserving scheme for the Allen–Cahn equation. Computers and Mathematics with Applications, 199(??): 225–241, December 1, 2025. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122125004067.

Devloo:2024:SMP

[DFGS24] Philippe R. B. Devloo, Jeferson W. D. Fernandes, Sônia M. Gomes, and Nathan Shauer. Stress mixed polyhedral finite elements for two-scale elasticity models with relaxed symmetry. Computers and Mathematics with Applications, 159 (??):302–318, April 1, 2024. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122124000750.

Demkowicz:2021:DAP

[DFHT21] Leszek Demkowicz, Thomas Führer, Norbert Heuer, and Xiaochuan Tian. The double adaptivity paradigm: (how to circumvent the discrete inf–sup conditions of Babuška and Brezzi). Computers and Mathematics with Applications, 95(??):41–66, August 1, 2021. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122120303898.

DiPietro:2022:PEE

[DFK22] Daniele A. Di Pietro, Ilaria Fontana, and Kyrylo Kazymyrenko. A posteriori error estimates via equilibrated stress reconstructions for contact problems approximated by Nitsche’s method. Computers and Mathematics with Applications, 111(??):61– 80, April 1, 2022. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122122000645.

deFaria:2022:MFS

[dFLdSLM22] Jairo Rocha de Faria, Daniel Lesnic, Rômulo da Silva Lima, and Thiago José Machado. The method of fundamental solutions for pointwise source reconstruction. Computers and Mathematics with Applications, 114(??):171– 179, May 15, 2022. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122122001389.

Desiderio:2021:VEM

[DFS21] L. Desiderio, S. Falletta, and L. Scuderi. A virtual element method coupled with a boundary integral non reflecting condition for 2D exterior Helmholtz problems. Computers and Mathematics with Applications, 84(??): 296–313, February 15, 2021. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122121000043.

Dassi:2022:BMV

[DFSV22] Franco Dassi, Alessio Fumagalli, Anna Scotti, and Giuseppe Vacca. Bend 3d mixed virtual element method for Darcy problems. Computers and Mathematics with Applications, 119 (??):1–12, August 1, 2022. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122122002140.

Dehghan:2022:OCA

[DG22a] Mehdi Dehghan and Zeinab Gharibi. Optimal convergence analysis of the energy-preserving immersed weak Galerkin method for second-order hyperbolic interface problems in inhomogeneous media. Computers and Mathematics with Applications, 105(??):150– 171, January 1, 2022. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122121004119.

Dehghan:2022:UAF

[DG22b] Mehdi Dehghan and Zeinab Gharibi. A unified analysis of fully mixed virtual element method for wormhole propagation arising in the petroleum engineering. Computers and Mathematics with Applications, 121(??): 30–51, September 1, 2022. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122122002486.

Demkowicz:2020:DSM

[DGK20] Leszek Demkowicz, Jay Gopalakrishnan, and Brendan Keith. The DPGstar method. Computers and Mathematics with Applications, 79(11):3092– 3116, June 1, 2020. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122120300250.

DellAccio:2024:NQC

[DGN24] Francesco Dell’Accio, Allal Guessab, and Federico Nudo. New quadratic and cubic polynomial enrichments of the Crouzeix–Raviart finite element. Computers and Mathematics with Applications, 170(??):204–212, September 15, 2024. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122124002876.

Devloo:2020:ETD

[DGQT20] Philippe R. B. Devloo, Sônia M. Gomes, Thiago O. Quinelato, and Shudan Tian. Enriched two dimensional mixed finite element models for linear elasticity with weak stress symmetry. Computers and Mathematics with Applications, 79(9):2678– 2700, May 1, 2020. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122119305620.

Du:2023:FEA

[DH23] Shaohong Du and Xiaoxia He. Finite element approximation to optimal Dirichlet boundary control problem: a priori and a posteriori error estimates. Computers and Mathematics with Applications, 131(??):14–25, February 1, 2023. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122122004916.

Du:2024:NDS

[DH24] Zhongjie Du and Chuanjiang He. Nonlinear diffusion system for simultaneous restoration and binarization of degraded document images. Computers and Mathematics with Applications, 153(??): 237–248, January 1, 2024. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122123005345.

Dien:2020:TMO

[DHVT20] Nguyen Minh Dien, Dinh Nguyen Duy Hai, Tran Quoc Viet, and Dang Duc Trong. On Tikhonov’s method and optimal error bound for inverse source problem for a time-fractional diffusion equation. Computers and Mathematics with Applications, 80(1):61–81, July 1, 2020. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122120300912.

Deng:2020:EPS

[DHW20] Weibing Deng, Wunming Han, and Qi Wang. The existence of periodic solution for infinite dimensional Hamiltonian systems. Computers and Mathematics with Applications, 79(2):354–362, January 15, 2020. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122119303384.

Ding:2021:DHO

[Din21] Hengfei Ding. The development of higher-order numerical differential formulas of Caputo derivative and their applications (I). Computers and Mathematics with Applications, 84(??): 203–223, February 15, 2021. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S089812212100002X.

Dey:2023:FEA

[DK23] Sangita Dey and B. V. Rathish Kumar. Finite element analysis of modified n-s equations coupled with energy transfer for hybrid nanofluid flow in complex domains.

Computers and Mathematics with Applications, 150(??): 37–53, November 15, 2023. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122123003796.

Dey:2024:FEA

[DK24] Sangita Dey and B. V. Rathish Kumar. Finite element analysis of a thermal nonequilibrium hybrid nanofluid model for natural convection in a complex porous enclosure. Computers and Mathematics with Applications, 169(??):177–194, September 1, 2024. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122124002840.

Delouei:2022:DNS

[DKH22] Amin Amiri Delouei, Sajjad Karimnejad, and Fuli He. Direct Numerical Simulation of pulsating flow effect on the distribution of non-circular particles with increased levels of complexity: IB-LBM. Computers and Mathematics with Applications, 121(??): 115–130, September 1, 2022. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122122002899.

Droniou:2024:NAS

[DKL24] Jérôme Droniou, Muhammad Awais Khan, and Kim-Ngan Le. Numerical analysis of the stochastic Stefan problem. Computers and Mathematics with Applications, 173(??):114–140, November 1, 2024. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122124003493.

Djoko:2022:SNS

[DKMS22] J. K. Djoko, J. Koko, M. Mbehou, and Toni Sayah. Stokes and Navier–Stokes equations under power law slip boundary condition: Numerical analysis. Computers and Mathematics with Applications, 128(??):198– 213, December 15, 2022. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122122004394.

Delendik:2021:DIC

[DKV21] K. Delendik, N. Kolyago, and O. Voitik. Design and investigation of cooling system for high-power LED luminaire. Computers and Mathematics with Applications, 83 (??):84–94, February 1, 2021. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122120300511.

Deng:2025:ACS

[DL25] Dingwen Deng and Yuxin Liang. Analysis of a class of stabilized and structurepreserving finite difference methods for Fisher–Kolmogorov– Petrovsky–Piscounov equation. Computers and Mathematics with Applications, 184 (??):86–106, April 15, 2025. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122125000628.

Deng:2022:ETL

[DLC22] Xiaomao Deng, Zi-Ju Liao, and Xiao-Chuan Cai. An efficient two-level overlapping domain decomposition method for recovering unsteady sources of 3D parabolic problems. Computers and Mathematics with Applications, 111(??):98– 108, April 1, 2022. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122122000670.

Das:2024:UMF

[DLN24] Avijit Das, Bishnu P. Lamichhane, and Neela Nataraj. A unified mixed finite element method for fourth-order time-dependent problems using biorthogonal systems. Computers and Mathematics with Applications, 165(??):52– 69, July 1, 2024. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122124001718.

delaRiva:2020:RSI

[dlRGR20] A. Pé de la Riva, F. J.

Gaspar, and C. Rodrigo. On the robust solution of an isogeometric discretization of bilaplacian equation by using multigrid methods. Computers and Mathematics with Applications, 80(2):386–394, July 2020. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122119304080.

delaRiva:2021:TLM

[dlRRG21] Álvaro Pé de la Riva, Carmen Rodrigo, and Francisco J. Gaspar. A twolevel method for isogeometric discretizations based on multiplicative Schwarz iterations. Computers and Mathematics with Applications, 100 (??):41–50, October 15, 2021. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122121003126.

delaRiva:2024:LFA

[dlRRG^{+}24] Álvaro Pé de la Riva, Carmen Rodrigo, Francisco J. Gaspar, James H. Adler, Xiaozhe Hu, and Ludmil Zikatanov. A local Fourier analysis for additive Schwarz smoothers. Computers and Mathematics with Applications, 158(??):13–20, March 15, 2024. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122123006016.

Dubois:2020:ABB

[DLT20] François Dubois, Pierre Lallemand, and Mohamed Mahdi Tekitek. On anti bounce back boundary condition for lattice Boltzmann schemes. Computers and Mathematics with Applications, 79(3): 555–575, February 1, 2020. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122119301646.

Deng:2024:CMS

[DLW24] Dingwen Deng, Shuhua Lin, and Qihong Wang. A class of monotone and structurepreserving du fort-Frankel schemes for nonlinear Allen– Cahn equation. Computers and Mathematics with Applications, 170(??):1–24, September 15, 2024. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122124002918.

Delis:2023:NSA

[DMG23] A. I. Delis, V. Mandikas, and H. Guillard. Numerical simulation of acoustic streaming in standing waves. Computers and Mathematics with Applications, 152(??): 199–220, December 15, 2023. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122123004753.

Dassi:2024:MVF

[DMRV24] Franco Dassi, David Mora, Carlos Reales, and Iván Velásquez. Mixed variational formulations of virtual elements for the polyharmonic operator (−Δ)^{n}. Computers and Mathematics with Applications, 158(??): 150–166, March 15, 2024. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122124000221.

Dobravec:2022:ARF

[DMŠ22] Tadej Dobravec, Boštjan Mavrič, and Božidar Šarler. Acceleration of RBF-FD meshless phase-field modelling of dendritic solidification by space-time adaptive approach. Computers and Mathematics with Applications, 126(??):77–99, November 15, 2022. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122122003881.

Ding:2024:SOF

[DMX24] Qianqian Ding, Shipeng Mao, and Ruijie Xi. Second order, fully decoupled, linear, exactly divergence-free and unconditionally stable discrete scheme for incompressible MHD equations. Computers and Mathematics with Applications, 169(??): 195–204, September 1, 2024. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122124002864.

Das:2022:SEA

[DN22] Avijit Das and Srinivasan Natesan. Stability and error analysis of a fully-discrete numerical method for system of 2D singularly perturbed parabolic PDEs. Computers and Mathematics with Applications, 110(??):135– 145, March 15, 2022. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122122000530.

Duy:2022:FSF

[DNP^{+}22] Trong-Nguyen Duy, Van-Tu Nguyen, Thanh-Hoang Phan, Dong-Hyun Kim, and Warn-Gyu Park. A free surface flow solver based on an efficient improvement to a coupling method for interface computations. Computers and Mathematics with Applications, 124(??):21–41, October 15, 2022. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122122003406.

Dieci:2023:TCO

[DO23] Luca Dieci and Daniyar Omarov. Techniques for continuous optimal transport problem. Computers and Mathematics with Applications, 146(??):176–191, September 15, 2023. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122123002882.

Druetta:2020:IPP

[DP20] P. Druetta and F. Picchioni. Influence of the polymer properties and numerical schemes on tertiary oil recovery processes. Computers and Mathematics with Applications, 79(4):1094–1110, February 15, 2020. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122119304316.

Devi:2025:DGT

[DP25] Raksha Devi and Dwijendra Narain Pandey. Discontinuous Galerkin timestepping method for semilinear parabolic problems with mild growth condition. Computers and Mathematics with Applications, 198(??): 274–292, November 15, 2025. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S089812212500416X.

Danumjaya:2021:MFF

[DPP21] P. Danumjaya, Ambit Kumar Pany, and Amiya K. Pani. Morley FEM for the fourth-order nonlinear reaction-diffusion problems. Computers and Mathematics with Applications, 99(??): 229–245, October 1, 2021. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122121002947.

Dwivedi:2024:NFT

[DR24] Himanshu Kumar Dwivedi and Rajeev. A novel fast tempered algorithm with high-accuracy scheme for 2D tempered fractional reaction–advection– subdiffusion equation. Computers and Mathematics with Applications, 176(??):371– 397, December 15, 2024. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122124004954.

daRocha:2025:NNS

[dR25a] Aline C. da Rocha. A new numerical strategy for the drift-diffusion equations based on bridging the hybrid mixed and exponential fitted methods. Computers and Mathematics with Applications, 190(??):185– 205, July 15, 2025. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122125001828.

Dwivedi:2025:NAT

[DR25b] Himanshu Kumar Dwivedi and Rajeev. A novel accelerated tempered algorithm with nonuniform time-stepping compact ADI scheme for 2D temperedfractional nonlinear Schrödinger equations with weak initial singularity. Computers and Mathematics with Applications, 196(??):312–337, October 15, 2025. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122125003256.

Duster:2020:H

[DRSY20] Alexander Düster, Alessandro Reali, Giancarlo Sangalli, and Zohar Yosibash. HOFEIM 2019. Computers and Mathematics with Applications, 80(11):2259– 2260, December 2020. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122120304004.

Dassi:2020:PSV

[DS20] F. Dassi and S. Scacchi. Parallel solvers for virtual element discretizations of elliptic equations in mixed form. Computers and Mathematics with Applications, 79 (7):1972–1989, April 2020. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122119303797.

Diening:2022:STD

[DS22] Lars Diening and Johannes Storn. A space-time DPG method for the heat equation. Computers and Mathematics with Applications, 105 (??):41–53, January 1, 2022. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122121004107.

Doss:2025:GMF

[DS25] L. Jones Tarcius Doss and V. Sindhujarani. H^{1}- Galerkin mixed finite element method using tensor product of cubic Bsplines for two-dimensional Kuramoto–Sivashinsky equation. Computers and Mathematics with Applications, 188 (??):19–39, June 15, 2025. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S089812212500104X.

Du:2024:EAV

[DSF24] Zijun Du, Haiyan Su, and Xinlong Feng. Error analysis of vector penalty-projection method with second order accuracy for incompressible magnetohydrodynamic system. Computers and Mathematics with Applications, 160 (??):46–59, April 15, 2024. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S089812212400052X.

daSilva:2020:FPI

[dSG20] Leandro Tavares da Silva and Gilson Antonio Giraldi. Fixed point implementation of a variational time integrator approach for smoothed particle hydrodynamics simulation of fluids. Computers and Mathematics with Applications, 79(4):1111–1130, February 15, 2020. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122119304304.

Dong:2024:HOM

[DSL24] Hao Dong, Jie Shi, and Jiale Linghu. Higher-order multi-scale computational approach and its convergence for nonlocal gradient elasticity problems of composite materials. Computers and Mathematics with Applications, 163(??):66– 83, June 1, 2024. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122124001214.

daSilva:2022:AFT

[dSMAdOW22] Ivan Assing da Silva, Roberto Dalledone Machado, Marcos Arndt, and Paulo de Oliveira Weinhardt. Assessment of the flattop stable GFEM for free vibration analysis. Computers and Mathematics with Applications, 117(??):271– 283, July 1, 2022. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122122001687.

Demo:2023:EPI

[DSR23] Nicola Demo, Maria Strazzullo, and Gianluigi Rozza.

An extended physics informed neural network for preliminary analysis of parametric optimal control problems. Computers and Mathematics with Applications, 143(??):383–396, August 1, 2023. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122123002018.

DAutilia:2020:MOD

[DSS20] Maria Chiara D’Autilia, Ivonne Sgura, and Valeria Simoncini. Matrix-oriented discretization methods for reaction–diffusion PDEs: Comparisons and applications. Computers and Mathematics with Applications, 79 (7):2067–2085, April 2020. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122119305097.

Diening:2025:EIS

[DST25] Lars Diening, Johannes Storn, and Tabea Tscherpel. Exact integration for singular Zienkiewicz and Guzmán–Neilan finite elements with implementation. Computers and Mathematics with Applications, 191 (??):60–85, August 1, 2025. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122125001634.

DeVincentis:2023:AFL

[DT23] Brian DeVincentis and John Thomas. Application of flux limiters to passive scalar advection for the lattice Boltzmann method. Computers and Mathematics with Applications, 144(??):1–11, August 15, 2023. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122123002365.

Dominguez:2024:RBI

[DT24] Víctor Domínguez and Catalin Turc. Robust boundary integral equations for the solution of elastic scattering problems via Helmholtz decompositions. Computers and Mathematics with Applications, 175(??):152–173, December 1, 2024. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122124004218.

Dougherty:2020:CIP

[DTV20] E. T. Dougherty, J. C. Turner, and F. Vogel. A computational investigation of preconditioning strategies and iterative methods for finite element based neurostimulation simulations. Computers and Mathematics with Applications, 79(12): 3531–3546, June 15, 2020. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122120300882.

Dzanic:2023:BPT

[DTW23] T. Dzanic, W. Trojak, and F. D. Witherden. Bounds preserving temporal integration methods for hyperbolic conservation laws. Computers and Mathematics with Applications, 135 (??):6–18, April 1, 2023. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122123000299.

Du:2024:TGPb

[Du24] Guangzhi Du. A two grid partition of unity finite element method for electrically conducting incompressible fluid flows. Computers and Mathematics with Applications, 156(??):121–128, February 15, 2024. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122123005990.

Dassi:2022:VEM

[DV22] Franco Dassi and Iván Velásquez. Virtual element method on polyhedral meshes for bi-harmonic eigenvalues problems. Computers and Mathematics with Applications, 121(??):85– 101, September 1, 2022. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S089812212200284X.

Divi:2020:EEB

[DVA^{+}20] Sai C. Divi, Clemens V. Verhoosel, Ferdinando Auricchio, Alessandro Reali, and E. Harald van Brummelen. Error-estimate-based adaptive integration for immersed isogeometric analysis. Computers and Mathematics with Applications, 80(11):2481– 2516, December 2020. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122120301590.

daVeiga:2022:VEM

[dVDMM22] L. Beirão da Veiga, F. Dassi, G. Manzini, and L. Mascotto. Virtual elements for Maxwell’s equations. Computers and Mathematics with Applications, 116 (??):82–99, June 15, 2022. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122121003114.

daVeiga:2020:CVE

[dVDR20] L. Beirão da Veiga, F. Dassi, and A. Russo. A C^{1} Virtual Element Method on polyhedral meshes. Computers and Mathematics with Applications, 79(7):1936– 1955, April 2020. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122119303220.

Du:2020:CAE

[DW20] Yu Du and Jiangxing Wang.

Convergence analysis of an energy based discontinuous Galerkin method for the wave equation in secondorder form: hp version. Computers and Mathematics with Applications, 79(11): 3223–3240, June 1, 2020. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S089812212030033X.

Duan:2025:DCB

[DWAM25] Xiaoyu Duan, Zihan Wang, Hengbin An, and Zeyao Mo. A DOFs condensation based algorithm for solving saddle point systems in 2D contact computation. Computers and Mathematics with Applications, 199(??):64–79, December 1, 2025. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122125003657.

Dong:2020:RBV

[DWS20] Qiannan Dong, Jiming Wu, and Shuai Su. Relationship between the vertexcentered linearity-preserving scheme and the lowest-order virtual element method for diffusion problems on starshaped polygons. Computers and Mathematics with Applications, 79(11):3117– 3138, June 1, 2020. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122120300225.

Dong:2025:HOT

[DWY^{+}25] Hao Dong, Yanqi Wang, Changqing Ye, Yihan Nie, and Puyang Gao. Higherorder three-scale asymptotic model and efficient two-stage numerical algorithm for transient nonlinear thermal conduction problems of composite structures. Computers and Mathematics with Applications, 192(??):72–103, August 15, 2025. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S089812212500207X.

Diao:2025:TEC

[DY25] Xinyu Diao and Bo Yu. Two efficient compact ADI methods for the two-dimensional fractional Oldroyd-B model.

Computers and Mathematics with Applications, 190 (??):72–89, July 15, 2025. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122125001531.

Dang:2022:VAV

[DYC^{+}22] Rongqi Dang, Aiming Yang, Yiming Chen, Yanqiao Wei, and Chunxiao Yu. Vibration analysis of variable fractional viscoelastic plate based on shifted Chebyshev wavelets algorithm. Computers and Mathematics with Applications, 119(??):149– 158, August 1, 2022. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122122002577.

Du:2022:PPN

[DYZ22] Yanyan Du, Ming Ye, and Qimin Zhang. A positivitypreserving numerical algorithm for stochastic agedependent population system with Lévy noise in a polluted environment. Computers and Mathematics with Applications, 125(??):51–79, November 1, 2022. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122122003662.

Demkowicz:2020:CDF

[DZ20] Leszek Demkowicz and Pietro Zanotti. Construction of DPG Fortin operators revisited. Computers and Mathematics with Applications, 80(11):2261–2271, December 2020. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122120302868.

Ding:2021:CBD

[DZ21] Hang Ding and Jun Zhou. Comments on “Blow-up and decay for a class of pseudoparabolic p-Laplacian equation with logarithmic nonlinearity” [Comput. Math. Appl. 75(2) (2018) 459–469]. Computers and Mathematics with Applications, 84(??): 144–147, February 15, 2021. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122120304703.

See [HGW18].

Deng:2022:EED

[DZ22] Dingwen Deng and Zilin Zhao. Efficiently energydissipation-preserving ADI methods for solving twodimensional nonlinear Allen– Cahn equation. Computers and Mathematics with Applications, 128(??):249– 272, December 15, 2022. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122122004539.

Du:2024:TGPa

[DZ24] Guangzhi Du and Liyun Zuo. Two-grid parallel stabilized finite element method based on overlapping domain decomposition for the Stokes problem. Computers and Mathematics with Applications, 155(??):91–96, February 1, 2024. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122123005412.

Deng:2024:SDA

[DZZ^{+}24] Cheng Deng, Hui Zheng, Rongping Zhang, Liangyong Gong, and Xiangcou Zheng. Structure deformation analysis of the deep excavation based on the local radial basis function collocation method. Computers and Mathematics with Applications, 174(??):495– 509, November 15, 2024. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122124004577.

Eisentrager:2020:CNH

[EAM20] S. Eisenträger, E. Atroshchenko, and R. Makvandi. On the condition number of high order finite element methods: Influence of p-refinement and mesh distortion. Computers and Mathematics with Applications, 80(11):2289– 2339, December 2020. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122120302169.

ElHaddad:2022:PSP

[EBDY22] M. El Haddad, Y. Belhamadia, J. Deteix, and D. Yakoubi. A projection scheme for phase change problems with convection. Computers and Mathematics with Applications, 108(??):109–122, February 15, 2022. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122122000013.

ElKadmiri:2024:AIS

[EBT24] Redouane El Kadmiri, Youssef Belaasilia, and Abdelaziz Timesli. Analysis of inhomogeneous structures in small and large deformations using the finite element-meshless coupling method. Computers and Mathematics with Applications, 169(??):273– 297, September 1, 2024. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122124003183.

Egger:2025:NMF

[EED^{+}25] H. Egger, F. Engertsberger, L. Domenig, K. Roppert, and M. Kaltenbacher. On nonlinear magnetic field solvers using local quasi-Newton updates. Computers and Mathematics with Applications, 183(??):20–31, April 1, 2025. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122125000392.

Eichler:2021:CLB

[EFF21] Pavel Eichler, Vladimír Fuka, and Radek Fučík. Cumulant lattice Boltzmann simulations of turbulent flow above rough surfaces. Computers and Mathematics with Applications, 92 (??):37–47, June 15, 2021. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122121000912.

Etangsale:2022:IEE

[EFFR22] Grégory Etangsale, Marwan Fahs, Vincent Fontaine, and Nalitiana Rajaonison. Improved error estimates of hybridizable interior penalty methods using a variable penalty for highly anisotropic diffusion problems. Computers and Mathematics with Applications, 119 (??):89–99, August 1, 2022. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S089812212200222X.

Esposti:2023:IBH

[EFK^{+}23] Bruno Degli Esposti, Antonella Falini, Tadej Kanduč, Maria Lucia Sampoli, and Alessandra Sestini. IgA-BEM for 3D Helmholtz problems using conforming and non-conforming multipatch discretizations and B-spline tailored numerical integration. Computers and Mathematics with Applications, 147(??):164– 184, October 1, 2023. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122123003061.

Eichler:2024:IMB

[EFS24] Pavel Eichler, Radek Fučík, and Pavel Strachota. Investigation of mesoscopic boundary conditions for lattice Boltzmann method in laminar flow problems. Computers and Mathematics with Applications, 173(??):87– 101, November 1, 2024. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S089812212400350X.

Eichler:2023:NNT

[EGF^{+}23] Pavel Eichler, Radek Galabov, Radek Fučík, Kateřina Škardová, Tomáš Oberhuber, Petr Pauš, Jaroslav Tintěra, and Radomír Chabiniok. Non-Newtonian turbulent flow through aortic phantom: Experimental and computational study using magnetic resonance imaging and lattice Boltzmann method. Computers and Mathematics with Applications, 136 (??):80–94, April 15, 2023. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122123000378.

Erath:2020:OCB

[EGP20] Christoph Erath, Gregor Gantner, and Dirk Praetorius. Optimal convergence behavior of adaptive FEM driven by simple (h − h/2)-type error estimators. Computers and Mathematics with Applications, 79(3): 623–642, February 1, 2020. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122119303529.

Ezzatneshan:2023:UPC

[EH23] Eslam Ezzatneshan and Kazem Hejranfar. An unstructured preconditioned central difference finite volume multiphase Euler solver for computing inviscid cavitating flows over arbitrary two- and three-dimensional geometries. Computers and Mathematics with Applications, 143(??):234–248, August 1, 2023. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122123002109.

Escapil-Inchauspe:2021:BPO

[EIJH21] Paul Escapil-Inchauspé and Carlos Jerez-Hanckes. Biparametric operator preconditioning. Computers and Mathematics with Applications, 102(??):220–232, November 15, 2021. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122121003606.

ElAkri:2024:ASS

[EJR24] A. El Akri, K. Jbilou, and A. Ratnani. Auxiliary splines space preconditioning for Bsplines finite elements: the case of HcurlΩ and HdivΩ elliptic problems. Computers and Mathematics with Applications, 159(??):102– 121, April 1, 2024. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122124000397.

Ellouz:2023:NMG

[EJWD23] Hajer Ellouz, Hanen Jrad, Mondher Wali, and Fakhreddine Dammak. Numerical modeling of geometrically nonlinear responses of smart magneto-electro-elastic functionally graded double curved shallow shells based on improved FSDT. Computers and Mathematics with Applications, 151(??): 271–287, December 1, 2023. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S089812212300425X.

Evans:2020:VMM

[EKB20] John A. Evans, David Kamensky, and Yuri Bazilevs. Variational multiscale modeling with discretely divergencefree subscales. Computers and Mathematics with Applications, 80(11):2517– 2537, December 2020. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122120301152.

Limami:2025:EIQ

[eLLAN25] Fakhr eddine Limami, Amine Laghrib, Lekbir Afraites, and Mourad Nachaoui. Enhancing image quality through fractional PDEs: a novel approach to image osmosis. Computers and Mathematics with Applications, 193(??): 54–76, September 1, 2025. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122125002263.

Endtmayer:2020:MOO

[ELN^{+}20] B. Endtmayer, U. Langer, I. Neitzel, T. Wick, and W. Wollner. Multigoaloriented optimal control problems with nonlinear PDE constraints. Computers and Mathematics with Applications, 79(10):3001– 3026, May 15, 2020. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122120300122.

Endtmayer:2024:GOA

[ELS24] B. Endtmayer, U. Langer, and A. Schafelner. Goaloriented adaptive spacetime finite element methods for regularized parabolic p-Laplace problems. Computers and Mathematics with Applications, 167(??):286– 297, August 1, 2024. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122124002293.

Esfahani:2025:UEB

[EMSA25] Sadegh Ghamsari Esfahani, Bashir Movahedian, Saeid Sarrami, and Mojtaba Azhari. On the use of exponential basis functions in the PINN architecture: an enhanced solution approach for the Laplace, Helmholtz, and elasto-static equations. Computers and Mathematics with Applications, 183 (??):234–255, April 1, 2025. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122125000574.

Ezzatneshan:2024:DPD

[ENF24] Eslam Ezzatneshan, Kian Nakhaei, and Ayoub Fattahi. Dynamics of particulate droplets collision: an Allen– Cahn based multiphase lattice Boltzmann approach. Computers and Mathematics with Applications, 174(??): 167–182, November 15, 2024. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122124003961.

ElYafi:2023:EAM

[EP23] Dima Abou El Nasser El Yafi and Victor Péron. Efficient asymptotic models for axisymmetric eddy current problems in linear ferromagnetic materials. Computers and Mathematics with Applications, 151(??):335–345, December 1, 2023. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122123004273.

Eisentrager:2024:EST

[ERG^{+}24] S. Eisenträger, L. Radtke, W. Garhuom, S. Löhnert, A. Düster, D. Juhre, and D. Schillinger. An eigenvalue stabilization technique for immersed boundary finite element methods in explicit dynamics. Computers and Mathematics with Applications, 166(??):129– 168, July 15, 2024. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122124001664.

Ebeling-Rump:2022:TST

[ERHL22] Moritz Ebeling-Rump, Dietmar Hömberg, and Robert Lasarzik. Two-scale topology optimization with heterogeneous mesostructures based on a local volume constraint. Computers and Mathematics with Applications, 126(??): 100–114, November 15, 2022. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122122003765.

Eruslu:2020:PBE

[ES20] Hasan Eruslu and Francisco-Javier Sayas. Polynomially bounded error estimates for Trapezoidal Rule Convolution Quadrature. Computers and Mathematics with Applications, 79(6):1634–1643, March 15, 2020. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122119304766.

Fardi:2025:NSN

[FAM25] M. Fardi, B. Azarnavid, and S. Mohammadi. Numerical simulation of nonlinear fractional integrodifferential equations on twodimensional regular and irregular domains: RBF partition of unity. Computers and Mathematics with Applications, 181(??):21– 43, March 1, 2025. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122124005704.

Fan:2023:TCV

[Fan23] Huo Fan. Three-component volume coordinate system for constructing high performance elements. Computers and Mathematics with Applications, 144(??):62–89, August 15, 2023. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122123002250.

Fassbender:2024:IET

[FBK^{+}24] Christian Faßbender, Tim Bürchner, Philipp Kopp, Ernst Rank, and Stefan Kollmannsberger. Implicitexplicit time integration for the immersed wave equation. Computers and Mathematics with Applications, 163(??):1–13, June 1, 2024. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122124001032.

Fan:2024:NNI

[FC24] Wenping Fan and Hao Cheng. A novel numerical inverse technique for multi-parameter time fractional radially symmetric anomalous diffusion problem with initial singularity. Computers and Mathematics with Applications, 158 (??):95–101, March 15, 2024. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122124000105.

Frigo:2022:ERP

[FCF22] Matteo Frigo, Nicola Castelletto, and Massimiliano Ferronato. Enhanced relaxed physical factorization preconditioner for coupled poromechanics. Computers and Mathematics with Applications, 106(??):27–39, January 15, 2022. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122121004120.

Frigo:2021:ESH

[FCFW21] Matteo Frigo, Nicola Castelletto, Massimiliano Ferronato, and Joshua A. White. Efficient solvers for hybridized three-field mixed finite element coupled poromechanics. Computers and Mathematics with Applications, 91(??):36– 52, June 1, 2021. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122120302765.

Ferreira:2020:DRE

[FdOS20] J. A. Ferreira, P. de Oliveira, and E. Silveira. Drug release enhanced by temperature: an accurate discrete model for solutions in H^{3}. Computers and Mathematics with Applications, 79(3): 852–875, February 1, 2020. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S089812211930392X.

Fuhrer:2024:DML

[FF24] Thomas Führer and Francisco Fuica. A DPG method for linear quadratic optimal control problems. Computers and Mathematics with Applications, 166(??):106– 117, July 15, 2024. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122124001925.

Fan:2025:NCN

[FFC^{+}25] Huayu Fan, Qiqi Feng, Rui Chen, Xiangyang Cao, and Zhi-Feng Pang. A non-convex and non-smooth weighted image denoising model. Computers and Mathematics with Applications, 187(??):85–105, June 1, 2025. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122125001051.

Fuhrer:2025:ANT

[FG25] Thomas Führer and Gregor Gantner. Aubin–Nitschetype estimates for space-time FOSLS for parabolic PDEs. Computers and Mathematics with Applications, 186 (??):155–170, May 15, 2025. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122125001129.

Ferreira:2020:MMD

[FGG^{+}20] J. A. Ferreira, M. B. Gonçalves, E. Gudiño, M. Maia, and C. M. Oishi. Mathematical model for degradation and drug release from an intravitreal biodegradable implant. Computers and Mathematics with Applications, 80(10):2212–2240, November 15, 2020. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122120303631.

Far:2020:SRO

[FGK20] Ehsan Kian Far, Martin Geier, and Manfred Krafczyk. Simulation of rotating objects in fluids with the cumulant lattice Boltzmann model on sliding meshes. Computers and Mathematics with Applications, 79(1):3– 16, January 2020. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122118304814.

Farber:2021:EGM

[FGK^{+}21] Peter Farber, Jens Gräbel, Norman Kroppen, Liesa Pötschke, Dirk Roos, Miriam Rosenbaum, Georg Stegschuster, and Peer Ueberholz. Electricity generation in a microbial fuel cell with textile carbon fibre anodes. Computers and Mathematics with Applications, 83 (??):4–23, February 1, 2021. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122119305528.

Fuhrer:2021:RDM

[FH21] Thomas Führer and Norbert Heuer. A robust DPG method for large domains. Computers and Mathematics with Applications, 94 (??):15–27, July 15, 2021. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122121001619.

Fung:2025:BOC

[FH25] Po Yin Fung and Sean Y. Hon. Block ω-circulant preconditioners for parabolic equations. Computers and Mathematics with Applications, 182(??):122–138, March 15, 2025. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122125000240.

Fuhrer:2023:DMQ

[FHH23] Thomas Führer, Pablo Herrera, and Norbert Heuer. A DPG method for the quadcurl problem. Computers and Mathematics with Applications, 149(??):221–238, November 1, 2023. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S089812212300411X.

Feng:2021:SOC

[FHM21] Qiwei Feng, Bin Han, and Peter Minev. Sixth order compact finite difference schemes for Poisson interface problems with singular sources. Computers and Mathematics with Applications, 99(??):2–25, October 1, 2021. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122121002832.

Farrell:2022:NMF

[FHM22] Patrick E. Farrell, Abdalaziz Hamdan, and Scott P. MacLachlan. A new mixed finite-element method for H^{2} elliptic problems. Computers and Mathematics with Applications, 128(??):300– 319, December 15, 2022. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122122004540.

Fu:2022:AHO

[FHZ22] Yayun Fu, Dongdong Hu, and Gengen Zhang. Arbitrary high-order exponential integrators conservative schemes for the nonlinear Gross–Pitaevskii equation. Computers and Mathematics with Applications, 121(??): 102–114, September 1, 2022. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122122002875.

Fialko:2021:PFE

[Fia21] Sergiy Fialko. Parallel finite element solver for multi-core computers with shared memory. Computers and Mathematics with Applications, 94(??):1–14, July 15, 2021. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S089812212100153X.

Ferreira:2022:DDE

[FJP22] J. A. Ferreira, D. Jordão, and L. Pinto. Drug delivery enhanced by ultrasound: Mathematical modeling and simulation. Computers and Mathematics with Applications, 107(??):57–69, February 1, 2022. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122121004429.

Freno:2020:STQ

[FJZC20] Brian A. Freno, William A. Johnson, Brian F. Zinser, and Salvatore Campione. Symmetric triangle quadrature rules for arbitrary functions. Computers and Mathematics with Applications, 79(10):2885–2896, May 15, 2020. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122119305954.

Fialko:2020:BSP

[FK20] Sergiy Fialko and Viktor Karpilovskyi. Block subspace projection preconditioned conjugate gradient method in modal structural analysis. Computers and Mathematics with Applications, 79(12): 3410–3428, June 15, 2020. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S089812212030064X.

Fuhrer:2021:STL

[FK21] Thomas Führer and Michael Karkulik. Space–time leastsquares finite elements for parabolic equations. Computers and Mathematics with Applications, 92(??):27– 36, June 15, 2021. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122121000729.

Farahat:2024:LBC

[FKKV24] Andrea Farahat, Mario Kapl, Aljaž Kosmač, and Vito Vitrih. A locally based construction of analysis-suitable G ^{1} multi-patch spline surfaces. Computers and Mathematics with Applications, 168 (??):46–57, August 15, 2024.

CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122124002487.

Farrell:2021:MSL

[FKR21] Patricio Farrell, Stefan Kayser, and Nella Rotundo. Modeling and simulation of the lateral photovoltage scanning method. Computers and Mathematics with Applications, 102(??):248– 260, November 15, 2021. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122121003710.

Faheem:2022:LWC

[FKW22] Mo Faheem, Arshad Khan, and Patricia J. Y. Wong. A Legendre wavelet collocation method for 1D and 2D coupled time-fractional nonlinear diffusion system. Computers and Mathematics with Applications, 128(??): 214–238, December 15, 2022. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122122004370.

Fetsov:2022:NMS

[FL22] Sergey S. Fetsov and Nickolay A. Lutsenko. Numerical model for studying axisymmetric thermal energy storages based on granular phase change materials with gaseous heat transfer fluid. Computers and Mathematics with Applications, 112 (??):138–153, April 15, 2022. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122122000852.

Fraschini:2024:USS

[FLMS24] S. Fraschini, G. Loli, A. Moiola, and G. Sangalli. An unconditionally stable space– time isogeometric method for the acoustic wave equation. Computers and Mathematics with Applications, 169(??): 205–222, September 1, 2024. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122124002773.

Fuica:2021:PEE

[FLOQ21] Francisco Fuica, Felipe Lepe, Enrique Otárola, and Daniel Quero. A posteriori error estimates in W^{1,p} × L^{p} spaces for the Stokes system with Dirac measures. Computers and Mathematics with Applications, 94 (??):47–59, July 15, 2021. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122121001577.

Feng:2023:RTD

[FLS23] Yani Feng, Qifeng Liao, and David Silvester. Robin-type domain decomposition with stabilized mixed approximation for incompressible flow. Computers and Mathematics with Applications, 147 (??):53–63, October 1, 2023. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122123003103.

Fang:2020:OCE

[FLW20] Zhiwei Fang, Jichun Li, and Xiang Wang. Optimal control for electromagnetic cloaking metamaterial parameters design. Computers and Mathematics with Applications, 79(4):1165–1176, February 15, 2020. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122119304262.

Faria:2025:CSB

[FM25] Luiz M. Faria and Florian Monteghetti. A complexscaled boundary integral equation for the embedded eigenvalues and complex resonances of the Neumann– Poincaré operator on domains with corners. Computers and Mathematics with Applications, 197(??):135– 166, November 1, 2025. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122125003402.

Ferro:2024:LSF

[FMP^{+}24] Nicola Ferro, Stefano Micheletti, Nicola Parolini, Simona Perotto, Marco Verani, and Paola Francesca Antonietti. Level set-fitted polytopal meshes with application to structural topology optimization. Computers and Mathematics with Applications, 169(??):99–111, September 1, 2024. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122124002797.

Falletta:2022:TFB

[FMS22] S. Falletta, G. Monegato, and L. Scuderi. Two FEM-BEM methods for the numerical solution of 2D transient elastodynamics problems in unbounded domains. Computers and Mathematics with Applications, 114 (??):132–150, May 15, 2022. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122122001377.

Fernandez-Nieto:2023:MMH

[FNGDV23] E. D. Fernández-Nieto, J. Garres-Díaz, and P. Vigneaux. Multilayer models for hydrostatic Herschel–Bulkley viscoplastic flows. Computers and Mathematics with Applications, 139(??):99– 117, June 1, 2023. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122123001141.

Fuhrer:2020:SNP

[FP20] Thomas Führer and Dirk Praetorius. A short note on plain convergence of adaptive least-squares finite element methods. Computers and Mathematics with Applications, 80(6):1619–1632, September 15, 2020. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122120302881.

Fritz:2022:TFC

[FRW22] Marvin Fritz, Mabel L. Rajendran, and Barbara Wohlmuth. Time-fractional Cahn–Hilliard equation: Wellposedness, degeneracy, and numerical solutions. Computers and Mathematics with Applications, 108(??):66–87, February 15, 2022. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122122000025.

Feischl:2021:CAS

[FS21a] Michael Feischl and Andrea Scaglioni. Convergence of adaptive stochastic collocation with finite elements. Computers and Mathematics with Applications, 98(??):139–156, September 15, 2021. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122121002571.

Fucik:2021:EFD

[FS21b] Radek Fučík and Robert Straka. Equivalent finite difference and partial differential equations for the lattice Boltzmann method. Computers and Mathematics with Applications, 90 (??):96–103, May 15, 2021. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122121000882.

Feng:2023:NAS

[FSH^{+}23] S. Z. Feng, Q. J. Sun, X. Han, Atilla Incecik, and Z. X. Li. A novel algorithm for the stochastic analysis of electric field problems considering material and spatial uncertainties. Computers and Mathematics with Applications, 133(??):124– 141, March 1, 2023. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122123000263.

Facco:2020:APH

[FSS^{+}20] W. G. Facco, R. C. Silva, K. F. O. Santos, E. J. Silva, A. S. Moura, and R. R. Saldanha. Alternative proposal of the high-order Gauss quadrature for reference triangle in the generalized finite element method. Computers and Mathematics with Applications, 80(10):2162– 2175, November 15, 2020. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122120303679.

Fang:2020:FMV

[FSW20] Zhi-Wei Fang, Hai-Wei Sun, and Hong Wang. A fast method for variable-order Caputo fractional derivative with applications to timefractional diffusion equations. Computers and Mathematics with Applications, 80(5):1443–1458, September 1, 2020. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122120302753.

Fuhrer:2021:UFL

[Füh21]  Thomas Führer.  Ultra-

weak formulation of linear PDEs in nondivergence form and DPG approximation. Computers and Mathematics with Applications, 95 (??):67–84, August 1, 2021. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S089812212030273X.

Furukawa:2025:SPS

[Fur25] Akira Furukawa. Source point selection in the MFS using orthogonal matching pursuit: Two-dimensional elastic wave scattering by a rectangular cavity. Computers and Mathematics with Applications, 199(??):80– 105, December 1, 2025. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122125003670.

Feng:2025:ACM

[FWBY25] Jinhua Feng, Shixi Wang, Hai Bi, and Yidu Yang. An adaptive conforming mixed method with high-order elements for clustered eigenvalues of the Helmholtz transmission problem with index of refraction n(x) = 1. Computers and Mathematics with Applications, 196(??): 201–217, October 15, 2025. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122125003049.

Fang:2022:DDP

[FWLD22] Yin Fang, Yue-Yue Wang, Wei Liu, and Chao-Qing Dai. Data-driven prediction of soliton solutions of the higher-order NLSE via the strongly-constrained PINN method. Computers and Mathematics with Applications, 127(??):144–153, December 1, 2022. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122122004059.

Fu:2022:EHO

[FX22] Yayun Fu and Zhuangzhi Xu. Explicit high-order conservative exponential time differencing Runge–Kutta schemes for the two-dimensional nonlinear Schrödinger equation. Computers and Mathematics with Applications, 119(??):141–148, August 1, 2022. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122122002152.

Fu:2025:HOM

[FXLS25] Kai Fu, Shusen Xie, Dong Liang, and Yilei Shi. Highorder mass conservative compact characteristic finite volume methods for advection-dominated diffusion equations. Computers and Mathematics with Applications, 197(??):112–134, November 1, 2025. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122125003323.

Fang:2023:SOS

[FYM23] Changjie Fang, Meifang Yang, and Stanisław Migórski. Shape optimization for the Stokes hemivariational inequality with slip boundary condition. Computers and Mathematics with Applications, 146(??):213–224, September 15, 2023. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122123002912.

Fu:2021:BCM

[FYXL21] Zhuo-Jia Fu, Li-Wen Yang, Qiang Xi, and Chein-Shan Liu. A boundary collocation method for anomalous heat conduction analysis in functionally graded materials. Computers and Mathematics with Applications, 88 (??):91–109, April 15, 2021. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122120300900.

Foadaddini:2022:NGB

[FZDS22] Ali Foadaddini, Seyed Alireza Zolfaghari, Hossein Mahmoodi Darian, and Hamid Saadatfar. A new GPUbased corrected explicitimplicit domain decomposition scheme for convectiondominated diffusion problems. Computers and Mathematics with Applications, 123(??):184–203, October 1, 2022. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122122003352.

Fu:2022:HSF

[FZH22] Yayun Fu, Yanmin Zhao, and Dongdong Hu. The Hamiltonian structure and fast energy-preserving algorithms for the fractional Klein– Gordon equation. Computers and Mathematics with Applications, 113(??):86– 102, May 1, 2022. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122122001122.

Fang:2022:FFV

[FZS22] Zhi-Wei Fang, Jia-Li Zhang, and Hai-Wei Sun. A fast finite volume method for spatial fractional diffusion equations on nonuniform meshes. Computers and Mathematics with Applications, 108(??): 175–184, February 15, 2022. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122122000153.

Feng:2025:MPP

[FZW^{+}25] Naixing Feng, Shuiqing Zeng, Xianpeng Wang, Jinfeng Zhu, and Atef Z. Elsherbeni. MFPC-PIML: Physicsinformed machine learning based on multiscale Fourier feature phase compensation. Computers and Mathematics with Applications, 187 (??):166–180, June 1, 2025. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122125001191.

Galiano:2021:EAS

[Gal21] Gonzalo Galiano. Error analysis of some nonlocal diffusion discretization schemes. Computers and Mathematics with Applications, 103(??): 40–52, December 1, 2021. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122121003771.

Gonzalez-Andrade:2022:DMA

[GAM22] Sergio González-Andrade and Paul E. Méndez. A dual-mixed approximation for a Huber regularization of generalized p-stokes viscoplastic flow problems. Computers and Mathematics with Applications, 112 (??):76–96, April 15, 2022. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122122000839.

Guidotti:2024:SMS

[GAM24] Nicolas L. Guidotti, Juan A. Acebrón, and José Monteiro. A stochastic method for solving time-fractional differential equations. Computers and Mathematics with Applications, 159(??):240– 253, April 1, 2024. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122124000646.

Guo:2021:AFE

[GB21] Liming Guo and Chunjia Bi. Adaptive finite element method for nonmonotone quasi-linear elliptic problems. Computers and Mathematics with Applications, 93 (??):94–105, July 1, 2021. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122121001152.

Guo:2023:CQO

[GB23] Liming Guo and Chunjia Bi. Convergence and quasioptimality of an adaptive finite element method for nonmonotone quasi-linear elliptic problems on L ^{2} errors. Computers and Mathematics with Applications, 139 (??):38–56, June 1, 2023. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122123000949.

Gharib:2023:LBM

[GBK23] Mohammad Reza Gharib, Seyed Ehsan Davarpanah Baygi, and Ali Koochi. Lattice Boltzmann model for simulation of a nano-scanner immersed in ionic dense media. Computers and Mathematics with Applications, 140 (??):237–249, June 15, 2023. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122123001578.

Gonzalez:2020:NVN

[GCC20] A. González, E. Castillo, and M. A. Cruchaga. Numerical verification of a non-residual orthogonal term-by-term stabilized finite element formulation for incompressible convective flow problems. Computers and Mathematics with Applications, 80(5):1009– 1028, September 1, 2020. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122120302479.

Gu:2023:FOW

[GCJ23] Jiaxi Gu, Xinjuan Chen, and Jae-Hun Jung. Fifth-order weighted essentially nonoscillatory schemes with new Z-type nonlinear weights for hyperbolic conservation laws. Computers and Mathematics with Applications, 134(??): 140–166, March 15, 2023. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122123000093.

Guo:2025:FER

[GCL25] Jixiao Guo, Yanping Chen, and Qin Liang. Fast evaluation and robust error analysis of the virtual element methods for time fractional diffusion wave equation. Computers and Mathematics with Applications, 177 (??):41–57, January 1, 2025. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122124004929.

Guo:2021:FDG

[GCMS21] Shimin Guo, Yaping Chen, Liquan Mei, and Yining Song. Finite difference/ generalized Hermite spectral method for the distributedorder time-fractional reactiondiffusion equation on multidimensional unbounded domains. Computers and Mathematics with Applications, 93(??):1–19, July 1, 2021. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122121001280.

Gorakifard:2021:AWP

[GCSF21] Mohsen Gorakifard, Ildefonso Cuesta, Clara Salueña, and Ehsan Kian Far. Acoustic wave propagation and its application to fluid structure interaction using the Cumulant Lattice Boltzmann Method. Computers and Mathematics with Applications, 87(??):91–106, April 1, 2021. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122121000535.

Gharibi:2021:NAL

[GDA21] Zeinab Gharibi, Mehdi Dehghan, and Mostafa Abbaszadeh. Numerical analysis of locally conservative weak Galerkin dual-mixed finite element method for the time-dependent Poisson– Nernst–Planck system. Computers and Mathematics with Applications, 92(??):88– 108, June 15, 2021. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122121000766.

Gong:2022:HOF

[GDA^{+}22] Zheng Gong, Gefei Deng, Chenge An, Zi Wu, and Xudong Fu. A high order finite difference solver for simulations of turbidity currents with high parallel efficiency. Computers and Mathematics with Applications, 128(??): 21–33, December 15, 2022. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122122004047.

Gatti:2025:SWB

[GdFFF25] Federico Gatti, Carlo de Falco, Marco Fois, and Luca Formaggia. A scalable wellbalanced Taylor–Galerkin scheme for a lava flow depthintegrated model with point source vents. Computers and Mathematics with Applications, 184(??):153– 167, April 15, 2025. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122125000677.

Gruszczynski:2023:RSO

[GDŁW23] Grzegorz Gruszczyński, Michał Dzikowski, and Łukasz Łaniewski-Wołłk. Revisiting the secondorder convergence of the lattice Boltzmann method with reaction-type source terms. Computers and Mathematics with Applications, 144 (??):34–50, August 15, 2023.

CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122123002237.

Gonzalez-Estrada:2021:EEP

[GENRB21] Octavio A. González-Estrada, Sundararajan Natarajan, Juan José Ródenas, and Stéphane P. A. Bordas. Error estimation for the polygonal finite element method for smooth and singular linear elasticity. Computers and Mathematics with Applications, 92(??):109– 119, June 15, 2021. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122121000924.

Gong:2021:PDD

[GF21] Zheng Gong and Xudong Fu.

A pencil distributed direct numerical simulation solver with versatile treatments for viscous term. Computers and Mathematics with Applications, 100(??):141–151, October 15, 2021. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122121003230.

Guo:2025:HOE

[GFXZ25] Hongji Guo, Yayun Fu, Xi Xi, and Gengen Zhang. High-order energy-preserving schemes for the Klein– Gordon–Zakharov system. Computers and Mathematics with Applications, 197(??): 217–234, November 1, 2025. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122125003633.

Golgeleyen:2025:HSM

[GGH25] Fikret Gölgeleyen, Ismet Gölgeleyen, and Muhammed Hasdemir. A hybrid solution method for an inverse problem for the general transport equation. Computers and Mathematics with Applications, 192(??):172– 188, August 15, 2025. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122125002135.

Guillen-Gonzalez:2020:SCRb

[GGRBRG20a] F. Guillén-González, M. A. Rodríguez-Bellido, and D. A. Rueda-Gómez. Study of a chemo-repulsion model with quadratic production. Part I: Analysis of the continuous problem and time-discrete numerical schemes. Computers and Mathematics with Applications, 80(5):692–713, September 1, 2020. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122120301553.

Guillen-Gonzalez:2020:SCRa

[GGRBRG20b] F. Guillén-González, M. A.

Rodríguez-Bellido, and D. A. Rueda-Gómez. Study of a chemo-repulsion model with quadratic production. Part II: Analysis of an unconditionally energy-stable fully discrete scheme. Computers and Mathematics with Applications, 80(5):636–652, September 1, 2020. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122120301565.

Guillen-Gonzalez:2024:SPF

[GGT24] Francisco Guillén-González and Giordano Tierra. Structure preserving finite element schemes for the Navier– Stokes–Cahn–Hilliard system with degenerate mobility. Computers and Mathematics with Applications, 172(??):181–201, October 15, 2024. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122124003444.

Ghafoor:2024:AOT

[GHAU24] Abdul Ghafoor, Manzoor Hussain, Danyal Ahmad, and Shams Ul Arifeen. Approximation of one and two dimensional nonlinear generalized Benjamin–Bona– Mahony Burgers’ equation with local fractional derivative. Computers and Mathematics with Applications, 172(??):125–133, October 15, 2024. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122124003390.

Georgieva:2020:ILR

[GHH20] Irina Georgieva, Stanislav Harizanov, and Clemens Hofreither. Iterative lowrank approximation solvers for the extension method for fractional diffusion. Computers and Mathematics with Applications, 80(2): 351–366, July 2020. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122119303542.

Gustafsson:2025:DFE

[GHK25] Tom Gustafsson, Antti Hannukainen, and Vili Kohonen. Distributed finite element solution using model order reduction. Computers and Mathematics with Applications, 199(??):403–422, December 1, 2025. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122125004122.

Gu:2024:SFW

[GHL24] Shanshan Gu, Fuchang Huo, and Shicheng Liu. A stabilizer-free weak Galerkin mixed finite element method for the biharmonic equation. Computers and Mathematics with Applications, 176(??): 109–121, December 15, 2024. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122124004231.

Guo:2025:DSM

[GHL25] Hong Guo, Jin Huang, and Zhaoxing Li. Direct sampling method for solving the inverse acoustic wave scattering problems in the time domain. Computers and Mathematics with Applications, 179(??):229–242, February 1, 2025. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122124005637.

Giammatteo:2025:HOD

[GHLS25] Elena Giammatteo, Alexander Heinlein, Philip L. Lederer, and Matthias Schlottbom. High-order discretized ACMS method for the simulation of finitesize two-dimensional photonic crystals. Computers and Mathematics with Applications, 196(??):376–394, October 15, 2025. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122125003396.

Gerhard:2022:USP

[GHMD22] Pierre Gerhard, Philippe Helluy, and Victor Michel-Dansac. Unconditionally stable and parallel discontinuous Galerkin solver. Computers and Mathematics with Applications, 112(??):116– 137, April 15, 2022. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122122000785.

Garhuom:2020:RSL

[GHRD20] Wadhah Garhuom, Simeon Hubrich, Lars Radtke, and Alexander Düster. A remeshing strategy for large deformations in the finite cell method. Computers and Mathematics with Applications, 80(11):2379– 2398, December 2020. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122120301243.

Grigori:2023:PAH

[GHS23] Laura Grigori, Sever A. Hirstoaga, and Julien Salomon. A parareal algorithm for a highly oscillating Vlasov–Poisson system with reduced models for the coarse solving. Computers and Mathematics with Applications, 130(??):137–148, January 15, 2023. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122122005065.

Gonnella:2023:TSD

[GHSR23] Isabella Carla Gonnella, Martin W. Hess, Giovanni Stabile, and Gianluigi Rozza. A two-stage deep learning architecture for model reduction of parametric timedependent problems. Computers and Mathematics with Applications, 149(??):115– 127, November 1, 2023. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122123003711.

Grasso:2020:PBM

[GI20] Paolo Grasso and Mauro S.

Innocente. Physics-based model of wildfire propagation towards faster-than-real-time simulations. Computers and Mathematics with Applications, 80(5):790–808, September 1, 2020. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122120302078.

Gatica:2021:WPB

[GI21] Gabriel N. Gatica and Cristian Inzunza. On the well-posedness of Banach spaces-based mixed formulations for the nearly incompressible Navier–Lamé and Stokes equations. Computers and Mathematics with Applications, 102(??):87–94, November 15, 2021. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122121003527.

Gierdziewicz:2023:SPS

[Gie23] Maciej Gierdziewicz. Simulation of processes and structures in the synapse in the context of tetrahedral mesh quality. Computers and Mathematics with Applications, 145(??):58–64, September 1, 2023. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122123002559.

Gatica:2025:PMF

[GIRB25] Gabriel N. Gatica, Cristian Inzunza, and Ricardo Ruiz-Baier. Primal-mixed finite element methods for the coupled Biot and Poisson– Nernst–Planck equations. Computers and Mathematics with Applications, 186 (??):53–83, May 15, 2025. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122125000975.

Gmati:2022:SME

[GJM22] Nabil Gmati, Faten Jelassi, and Anis Mohamed. On the Schwarz method for the eddy currents model. Computers and Mathematics with Applications, 113(??):174– 188, May 1, 2022. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122122000803.

Guan:2020:RVC

[GJS20] Jixiang Guan, Shujie Jing, and Zhiyong Si. A rotational velocity-correction projection method for unsteady incompressible magnetohydrodynamics equations. Computers and Mathematics with Applications, 80(5):809–821, September 1, 2020. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122120301644.

Groselj:2020:SRB

[GK20] Jan Groselj and Marjeta Knez. On stable representations of Bell elements. Computers and Mathematics with Applications, 79(10): 2924–2941, May 15, 2020. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122119305978.

Groselj:2021:CEB

[GK21] Jan Grošelj and Marjeta Knez. A construction of edge B-spline functions for a C^{1} polynomial spline on two triangles and its application to Argyris type splines. Computers and Mathematics with Applications, 99(??): 329–344, October 1, 2021. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122121003072.

Ghafoor:2022:HCM

[GKHU22] Abdul Ghafoor, Nazish Khan, Manzoor Hussain, and Rahman Ullah. A hybrid collocation method for the computational study of multiterm time fractional partial differential equations. Computers and Mathematics with Applications, 128(??): 130–144, December 15, 2022. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122122004278.

Groselj:2020:SSS

[GKK^{+}20] Jan Grošelj, Mario Kapl, Marjeta Knez, Thomas Takacs, and Vito Vitrih. A super-smooth C^{1} spline space over planar mixed triangle and quadrilateral meshes. Computers and Mathematics with Applications, 80(12):2623–2643, December 15, 2020. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122120303977.

Grec:2025:USN

[GKP25] Bérénice Grec, Davor Kumozec, and Yohan Penel. Unconditionally stable numerical scheme for the 2D transport equation. Computers and Mathematics with Applications, 182(??):275– 290, March 15, 2025. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122125000501.

Gui:2024:IMA

[GLC24] Qiang Gui, Wei Li, and Yingbin Chai. Improved modal analyses using the novel quadrilateral overlapping elements. Computers and Mathematics with Applications, 154(??):138–152, January 15, 2024. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122123005242.

Gao:2025:NDU

[GLF25] Haijun Gao, Xi Li, and Minfu Feng. A new decoupled unconditionally stable scheme and its optimal error analysis for the Cahn–Hilliard–Navier– Stokes equations. Computers and Mathematics with Applications, 194(??):53–85, September 15, 2025. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122125002494.

Gao:2020:QSS

[GLL20a] Wen-Li Gao, Xi-An Li, and Xin-Ming Lu. On quasi shift-splitting iteration method for a class of saddle point problems. Computers and Mathematics with Applications, 79(10):2912– 2923, May 15, 2020. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S089812211930598X.

Gao:2020:NFE

[GLL^{+}20b] Xinghua Gao, Fawang Liu, Hong Li, Yang Liu, Ian Turner, and Baoli Yin. A novel finite element method for the distributed-order time fractional cable equation in two dimensions. Computers and Mathematics with Applications, 80(5):923–939, September 1, 2020. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122120301668.

Gong:2023:CNC

[GLLZ23] Chunye Gong, Dongfang Li, Lili Li, and Dan Zhao. Crank–Nicolson compact difference schemes and their efficient implementations for a class of nonlocal nonlinear parabolic problems. Computers and Mathematics with Applications, 132 (??):1–17, February 15, 2023. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122122004588.

Glaubitz:2021:TSR

[GLÖ21] Jan Glaubitz, Elise Le Meledo, and Philipp Öffner. Towards stable radial basis function methods for linear advection problems. Computers and Mathematics with Applications, 85 (??):84–97, March 1, 2021. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122121000225.

Gangl:2025:RFE

[GLS25] Peter Gangl, Richard Löscher, and Olaf Steinbach. Regularization and finite element error estimates for elliptic distributed optimal control problems with energy regularization and state or control constraints. Computers and Mathematics with Applications, 180(??):242–260, February 15, 2025. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122124005728.

Gao:2021:TSG

[GM21a] Yali Gao and Liquan Mei. Time-splitting Galerkin method for spin–orbit-coupled Bose– Einstein condensates. Computers and Mathematics with Applications, 87(??):77– 90, April 1, 2021. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122121000511.

Grabski:2021:IEP

[GM21b] Jakub Krzysztof Grabski and Agata Mrozek. Identification of elastoplastic properties of rods from torsion test using meshless methods and a metaheuristic. Computers and Mathematics with Applications, 92 (??):149–158, June 15, 2021. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122121000997.

Ghosh:2022:HPC

[GM22] Abhijit Ghosh and Chittaranjan Mishra. Highperformance computation of pricing two-asset American options under the Merton jump-diffusion model on a GPU. Computers and Mathematics with Applications, 105(??):29–40, January 1, 2022. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122121004090.

Gruszczynski:2020:CPF

[GML^{+}20] G. Gruszczyński, T. Mitchell, C. Leonardi, L. Laniewski-Wołłk, and T. Barber. A cascaded phase-field lattice Boltzmann model for the simulation of incompressible, immiscible fluids with high density contrast. Computers and Mathematics with Applications, 79(4):1049–1071, February 15, 2020. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122119304158.

Ghoudi:2023:NAF

[GMS23] Tarek Ghoudi, M. Shadi Mohamed, and Mohammed Seaid. Novel adaptive finite volume method on unstructured meshes for time-domain wave scattering and diffraction. Computers and Mathematics with Applications, 141(??):54– 66, July 1, 2023. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122123001268.

Ghorbani:2024:EEF

[GNB24] Mehran Ghorbani, Nima Noormohammadi, and Bijan Boroomand. Enrichment of the element free Galerkin method for cracks and notches without a priori knowledge of the analytical singularity order. Computers and Mathematics with Applications, 162 (??):155–179, May 15, 2024. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122124001081.

Garcke:2024:ALE

[GNZ24] Harald Garcke, Robert Nürnberg, and Quan Zhao. Arbitrary Lagrangian–Eulerian finite element approximations for axisymmetric twophase flow. Computers and Mathematics with Applications, 155(??):209–223, February 1, 2024. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122123005618.

Geridonmez:2020:MNC

[GO20] B. Pekmen Geridonmez and H. F. Oztop. MHD natural convection in a cavity in the presence of cross partial magnetic fields and Al_{2}O_{3}- water nanofluid. Computers and Mathematics with Applications, 80(12):2796– 2810, December 15, 2020. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122120303965.

Gwinner:2023:CFE

[GO23] J. Gwinner and N. Ovcharova.

Coupling of finite element and boundary element methods with regularization for a nonlinear interface problem with nonmonotone setvalued transmission conditions. Computers and Mathematics with Applications, 134 (??):45–54, March 15, 2023. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122122005284.

Gracia:2025:SPC

[GO25] José Luis Gracia and Eugene O’Riordan. Singularly perturbed convection-diffusion elliptic problems with a nonsmooth forcing term. Computers and Mathematics with Applications, 196(??):458– 476, October 15, 2025. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S089812212500358X.

Gontijo:2025:EML

[GOP25] R. G. Gontijo, F. E. L. Ossege, and J. L. J. Pereira. Evaluation of machine learning algorithms in the prediction of key therapeutic quantities in magnetic hyperthermia. Computers and Mathematics with Applications, 194(??):362–378, September 15, 2025. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S089812212500272X.

Gatica:2021:BSB

[GORBS21] Gabriel N. Gatica, Ricardo Oyarzúa, Ricardo Ruiz-Baier, and Yuri D. Sobral. Banach spaces-based analysis of a fully-mixed finite element method for the steadystate model of fluidized beds. Computers and Mathematics with Applications, 84(??): 244–276, February 15, 2021. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122121000031.

Gatica:2020:FFA

[GOV20] Gabriel N. Gatica, Ricardo Oyarzúa, and Nathalie Valenzuela. A five-field augmented fully-mixed finite element method for the Navier–Stokes/Darcy coupled problem. Computers and Mathematics with Applications, 80(8):1944–1963, October 15, 2020. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122120303242.

Guimaraes:2020:NAS

[GP20] Osvaldo Guimarães and José Roberto C. Piqueira. Novel approach to spectral methods for irregular domains. Computers and Mathematics with Applications, 80(1): 1–12, July 1, 2020. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122120300833.

Gantner:2022:ABE

[GP22a] Gregor Gantner and Dirk Praetorius. Adaptive BEM for elliptic PDE systems, part II: Isogeometric analysis with hierarchical Bsplines for weakly-singular integral equations. Computers and Mathematics with Applications, 117(??):74– 96, July 1, 2022. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122122001559.

Garrappa:2022:CES

[GP22b] Roberto Garrappa and Marina Popolizio. A computationally efficient strategy for time-fractional diffusionreaction equations. Computers and Mathematics with Applications, 116(??):181– 193, June 15, 2022. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122121002133.

Giacomini:2022:AMA

[GP22c] Matteo Giacomini and Simona Perotto. Anisotropic mesh adaptation for regionbased segmentation accounting for image spatial information. Computers and Mathematics with Applications, 121(??):1–17, September 1, 2022. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122122002711.

Garg:2023:MFE

[GP23] Divay Garg and Kamana Porwal. Mixed finite element method for a second order Dirichlet boundary control problem. Computers and Mathematics with Applications, 135(??):31– 59, April 1, 2023. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122123000111.

Galarce:2025:FCL

[GP25] Felipe Galarce and Douglas R. Q. Pacheco. Fully consistent lowest-order finite element methods for generalised Stokes flows with variable viscosity. Computers and Mathematics with Applications, 188(??):40– 49, June 15, 2025. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122125001087.

Gomez:2021:RCT

[GPA21] Vicente Gómez and Carlos Pérez-Arancibia. On the regularization of Cauchytype integral operators via the density interpolation method and applications. Computers and Mathematics with Applications, 87 (??):107–119, April 1, 2021. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122121000389.

Garg:2021:PSC

[GPL21] Deepak Garg, Paolo Papale, and Antonella Longo. A partitioned solver for compressible/incompressible fluid flow and light structure. Computers and Mathematics with Applications, 100(??):182– 195, October 15, 2021. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122121003254.

Gatto:2024:CAE

[GPRH^{+}24] Paolo Gatto, Ramon Planas, Francisco Romero-Hinrichsen, Tommy Henriksson, Christian El-Salloum, Abdolrahman Khavari, and Stephen Pearce. Computational aspects of electromagnetic tomography. Computers and Mathematics with Applications, 153(??):249–255, January 1, 2024. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122123005382.

Garmatter:2022:IPA

[GPRS22] Dominik Garmatter, Margherita Porcelli, Francesco Rinaldi, and Martin Stoll. Improved penalty algorithm for mixed integer PDE constrained optimization problems. Computers and Mathematics with Applications, 116(??):2–14, June 15, 2022. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122121003953.

Ge:2022:TLM

[GPY22] Zhihao Ge, Shuaichao Pei, and Yinyin Yuan. A twolevel multiphysics finite element method for a nonlinear poroelasticity model. Computers and Mathematics with Applications, 124 (??):63–73, October 15, 2022. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122122003418.

Gao:2023:NEA

[GQ23] Huadong Gao and Weifeng Qiu. A new error analysis and post-processing technique of the lowest-order Raviart–Thomas mixed finite element method for parabolic problems. Computers and Mathematics with Applications, 133(??):174– 188, March 1, 2023. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122123000275.

Guan:2023:WGF

[GQZ23] Qingguang Guan, Gillian Queisser, and Wenju Zhao. Weak Galerkin finite element method for second order problems on curvilinear polytopal meshes with Lipschitz continuous edges or faces.

Computers and Mathematics with Applications, 148(??): 282–292, October 15, 2023. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122123003565.

Grabski:2021:SPM

[Gra21] Jakub Krzysztof Grabski. On the sources placement in the method of fundamental solutions for time-dependent heat conduction problems. Computers and Mathematics with Applications, 88 (??):33–51, April 15, 2021. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122119302330.

Grieshaber:2020:DGA

[GRR20] B. J. Grieshaber, F. Rasolofoson, and B. D. Reddy. Discontinuous Galerkin approximations for near-incompressible and near-inextensible transversely isotropic bodies. Computers and Mathematics with Applications, 79 (7):1914–1935, April 2020. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122119302147.

Georgiou:2024:P

[GRYX24] Georgios Georgiou, Ernst Rank, Zohar Yosibash, and Christos Xenophontos. Preface. Computers and Mathematics with Applications, 176 (??):491, December 15, 2024. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122124004759.

Galecki:2022:ABO

[GS22] Jakub Gałecki and Jacek Szumbarski. Adjoint-based optimal control of incompressible flows with convectivelike energy-stable open boundary conditions. Computers and Mathematics with Applications, 106(??):40–56, January 15, 2022. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122121004302.

Gudi:2023:TLF

[GS23] Thirupathi Gudi and Ramesh Ch.

Sau. A two level finite element method for Stokes constrained Dirichlet boundary control problem. Computers and Mathematics with Applications, 129(??):126– 135, January 1, 2023. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122122004965.

Gutierrez-Santacreu:2021:AFD

[GSRG21] Juan Vicente Gutiérrez-Santacreu and José Rafael Rodríguez-Galván. Analysis of a fully discrete approximation for the classical Keller–Segel model: Lower and a priori bounds. Computers and Mathematics with Applications, 85 (??):69–81, March 1, 2021. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122121000195.

Gao:2024:FOC

[GSS24] Juanjuan Gao, Jiebao Sun, and Shengzhu Shi. Fractionalorder cross-diffusion system for multiplicative noise removal. Computers and Mathematics with Applications, 164(??):1–11, June 15, 2024. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122124001469.

Gotschel:2021:KFF

[GSW21] Sebastian Götschel, Anton Schiela, and Martin Weiser. Kaskade 7 — a flexible finite element toolbox. Computers and Mathematics with Applications, 81(??):444–458, January 1, 2021. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S089812212030078X.

Gunatilake:2022:IMS

[Gun22]  Janitha Gunatilake.  An

improved multigrid solver for the p-hierarchical basis finite element method using a space decomposition smoother. Computers and Mathematics with Applications, 124(??):52–62, October 15, 2022. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S089812212200339X.

Gutt:2020:BIA

[Gut20] Robert Gutt. Boundary integral approach for the mixed Dirichlet–Robin boundary value problem for the nonlinear Darcy–Forchheimer-Brinkman system. Computers and Mathematics with Applications, 79(10):2805– 2818, May 15, 2020. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122119305723.

Gonzalez:2020:SDV

[GV20] Eduardo González and Marcelo J.

Villena. On the spatial dynamics of vaccination: a spatial SIRS–V model. Computers and Mathematics with Applications, 80(5):733–743, September 1, 2020. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122120301620.

Gigante:2021:SLC

[GV21] Giacomo Gigante and Christian Vergara. On the stability of a loosely-coupled scheme based on a Robin interface condition for fluidstructure interaction. Computers and Mathematics with Applications, 96(??):109– 119, August 15, 2021. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122121001929.

Gantner:2022:AST

[GvV22] Gregor Gantner and Raymond van Venetië. Adaptive space-time BEM for the heat equation. Computers and Mathematics with Applications, 107(??):117–131, February 1, 2022. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122121004624.

Ge:2025:DAM

[GW25] Shanshan Ge and Jian Wang. Denoising algorithm of the modified Gray–Scott model on non-uniform grids. Computers and Mathematics with Applications, 195(??):191– 211, October 1, 2025. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122125003074.

Guan:2021:UOE

[GWN21] Zhen Guan, Jungang Wang, and Yufeng Nie. Unconditionally optimal error estimates of two linearized Galerkin FEMs for the twodimensional nonlinear fractional Rayleigh–Stokes problem. Computers and Mathematics with Applications, 93(??):78–93, July 1, 2021. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122121001401.

Guo:2022:OSG

[GWYR22] Xu Guo, Jiansen Wang, Senlin Yang, and Yuxiao Ren. Optimal staggeredgrid finite-difference method for wave modeling based on artificial neural networks. Computers and Mathematics with Applications, 108(??): 141–158, February 15, 2022. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122122000116.

Guo:2023:IFE

[GXFT23] Changyin Guo, Xufeng Xiao, Xinlong Feng, and Zhijun Tan. An immersed finite element method for elliptic interface problems on surfaces. Computers and Mathematics with Applications, 131 (??):54–67, February 1, 2023. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122122004795.

Gan:2025:PDA

[GY25] Ziyu Gan and Bo Yu. PINN-DTC: Adaptive partitioning of physics-informed neural network for designing thermal cloak. Computers and Mathematics with Applications, 195(??):419– 431, October 1, 2025. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122125003268.

Guo:2025:HOD

[GYC25] Hanlin Guo, Li Yin, and Xia Cui. High order direct discontinuous Galerkin method for elliptic interface problem on arbitrary polygon fitted meshes. Computers and Mathematics with Applications, 191(??):144– 166, August 1, 2025. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122125001816.

Gao:2025:PIN

[GYL25] Bo Gao, Ruoxia Yao, and Yan Li. Physics-informed neural networks with adaptive loss weighting algorithm for solving partial differential equations. Computers and Mathematics with Applications, 181(??):216– 227, March 1, 2025. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122125000069.

Guo:2023:ESS

[GZ23] Jiaojiao Guo and Qingqu Zhuang. Energy stable schemes for the Klein– Gordon–Zakharov equations. Computers and Mathematics with Applications, 147(??): 150–163, October 1, 2023. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S089812212300305X.

Guo:2023:PID

[GZAR23] Hongwei Guo, Xiaoying Zhuang, Naif Alajlan, and Timon Rabczuk. Physicsinformed deep learning for melting heat transfer analysis with model-based transfer learning. Computers and Mathematics with Applications, 143(??):303–317, August 1, 2023. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122123002122.

Georgiev:2020:NMS

[GZG20] Krassimir Georgiev, Zahari Zlatev, and Ivan Georgiev. Numerical methods for scientific computations and advanced applications II: Preface. Computers and Mathematics with Applications, 80(2):285, July 2020. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S089812212030208X.

Ha:2020:EGI

[Ha20] Youn Doh Ha. An extended ghost interlayer model in peridynamic theory for highvelocity impact fracture of laminated glass structures. Computers and Mathematics with Applications, 80(5): 744–761, September 1, 2020. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122120301991.

Huynh:2023:TPS

[HA23] Giang D. Huynh and Reza Abedi. Tent-pitcher spacetime discontinuous Galerkin method for one-dimensional linear hyperbolic and parabolic PDEs. Computers and Mathematics with Applications, 148(??):26–40, October 15, 2023. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122123003164.

Huang:2023:OPT

[HAC23] Chaobao Huang, Na An, and Hu Chen. Optimal pointwise-in-time error analysis of a mixed finite element method for a multiterm time-fractional fourthorder equation. Computers and Mathematics with Applications, 135(??):149– 156, April 1, 2023. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122123000342.

Hage:2024:HMA

[Hag24] Ilige S. Hage. A hybrid model for accurate prediction of composite longitudinal elastic modulus. Computers and Mathematics with Applications, 174(??):287– 297, November 15, 2024. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122124004279.

Hajduk:2021:MCL

[Haj21] Hennes Hajduk. Monolithic convex limiting in discontinuous Galerkin discretizations of hyperbolic conservation laws. Computers and Mathematics with Applications, 87 (??):120–138, April 1, 2021. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122121000547.

Hakula:2020:ARE

[Hak20]  Harri Hakula.  Adaptive

reference elements via harmonic extensions and associated inner modes. Computers and Mathematics with Applications, 80(11):2272– 2288, December 2020. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122120302856.

Hakula:2024:HEE

[Hak24a] Harri Hakula. Harmonic extension elements: Eigenproblems and error estimation. Computers and Mathematics with Applications, 163 (??):27–41, June 1, 2024. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122124001202.

Hakula:2024:HOF

[Hak24b] Harri Hakula. On highorder finite element solution of eigenvalue problems on isospectral surfaces. Computers and Mathematics with Applications, 168(??):22–32, August 15, 2024. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122124002414.

Haider:2024:EHM

[HANA24] Ali Haider, M. S. Anwar, Yufeng Nie, and M. S. Alqarni. Enhanced heat and mass transfer in porous media with Oldroyd-B complex nano-fluid flow and heat source. Computers and Mathematics with Applications, 170(??):282–300, September 15, 2024. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122124003146.

Halada:2024:RBD

[HB24] Tomáš Halada and Luděk Beneš. Relation between diffusive terms and Riemann solver in WCSPH. Computers and Mathematics with Applications, 167(??):239– 248, August 1, 2024. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122124002281.

He:2025:DTD

[HB25] Bin He and Shouzhu Bao.

Design and time-domain finite element analysis of a carpet thermal concentrator in metamaterials. Computers and Mathematics with Applications, 185(??):94– 109, May 1, 2025. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122125001117.

Hostos:2021:SSS

[HBC^{+}21] Juan C. Álvarez Hostos, Joselynne C. Salazar Bove, Marcela A. Cruchaga, Víctor D. Fachinotti, and Rafael A. Mujica Agelvis. Solving steadystate lid-driven square cavity flows at high Reynolds numbers via a coupled improved element-free Galerkin–reduced integration penalty method. Computers and Mathematics with Applications, 99(??): 211–228, October 1, 2021. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122121002972.

Ha:2021:MGM

[HC21] Sang Truong Ha and Hyoung Gwon Choi. A meshless geometric multigrid method based on a node-coarsening algorithm for the linear finite element discretization. Computers and Mathematics with Applications, 96 (??):31–43, August 15, 2021. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122121001899.

He:2022:EAC

[HC22a] Yuyu He and Hongtao Chen.

Efficient algorithm and convergence analysis of conservative SAV compact difference scheme for Boussinesq paradigm equation. Computers and Mathematics with Applications, 125(??):34–50, November 1, 2022. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122122003625.

Huang:2022:AMU

[HC22b] Zhenjie Huang and Xiaoliang Cheng. An accelerated method of Uzawa algorithm in contact problems. Computers and Mathematics with Applications, 127(??): 97–104, December 1, 2022. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122122004163.

Ha:2023:SMF

[HC23] Sang Truong Ha and Hyoung Gwon Choi. Semimonolithic formulation based on a projection method for simulating fluid–structure interaction problems. Computers and Mathematics with Applications, 134(??):207– 224, March 15, 2023. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122123000354.

Han:2025:SSC

[HCHM25] Jiarong Han, Junxi Chen, Fei Hu, and Liangcai Mei. A study on shock capturing in the Burgers’ equation based on RH-piecewise PINNs. Computers and Mathematics with Applications, 199(??): 309–324, December 1, 2025. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122125004298.

He:2024:PFS

[HCY^{+}24] Rui He, Yanfu Chen, Zihao Yang, Jizu Huang, and Xiaofei Guan. Phase field smoothing-PINN: a neural network solver for partial differential equations with discontinuous coefficients. Computers and Mathematics with Applications, 171(??):188– 203, October 1, 2024. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122124003249.

Hao:2025:FOD

[HCZ25] Zhaopeng Hao, Zhiqiang Cai, and Zhongqiang Zhang. Fractional-order dependent radial basis functions meshless methods for the integral fractional Laplacian. Computers and Mathematics with Applications, 178(??): 197–213, January 15, 2025. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122124005248.

Henneking:2021:NSP

[HD21] S. Henneking and L. Demkowicz. A numerical study of the pollution error and DPG adaptivity for long waveguide simulations. Computers and Mathematics with Applications, 95(??):85–100, August 1, 2021. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122120301280.

Hartmann:2023:TAF

[HDG23] Stefan Hartmann, Pranav Kumar Dileep, and Matthias Grafenhorst. A timeadaptive FE ^{2}-approach within the method of vertical lines. Computers and Mathematics with Applications, 151(??):222–243, December 1, 2023. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122123004212.

Han:2024:PSF

[HDM24] Jing Han, Guangzhi Du, and Shilin Mi. A parallel stabilized finite element method for the Navier– Stokes problem. Computers and Mathematics with Applications, 170(??):33–41, September 15, 2024. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122124003043.

Hong:2024:ILB

[HDWL24] Peng Hong, Chuanshan Dai, Guiling Wang, and Haiyan Lei. An implicit lattice Boltzmann flux solver with a projection-based interpolation scheme for the convection–diffusion equation. Computers and Mathematics with Applications, 159 (??):287–301, April 1, 2024. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122124000701.

He:2023:CRS

[He23a] Yuchen He. On the convergence rate of the splittingup scheme for rough partial differential equations. Computers and Mathematics with Applications, 139 (??):1–6, June 1, 2023. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122123000901.

He:2023:OSF

[He23b] Yunhui He. Optimal smoothing factor with coarsening by a factor of three for the MAC scheme for the Stokes equations. Computers and Mathematics with Applications, 132(??):63–72, February 15, 2023. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122122005089.

Hosseini:2020:WCL

[HEDT20] S. A. Hosseini, A. Eshghinejadfard, N. Darabiha, and D. Thévenin. Weakly compressible Lattice Boltzmann simulations of reacting flows with detailed thermo-chemical models. Computers and Mathematics with Applications, 79 (1):141–158, January 2020. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122117305473.

He:2023:LFM

[HFG23] Linshuang He, Minfu Feng, and Jun Guo. A lockingfree and mass conservative H(div) conforming DG method for the Biot’s consolidation model. Computers and Mathematics with Applications, 136(??):151– 164, April 15, 2023. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122123000470.

Henneking:2021:MCA

[HGD21] Stefan Henneking, Jacob Grosek, and Leszek Demkowicz. Model and computational advancements to full vectorial Maxwell model for studying fiber amplifiers. Computers and Mathematics with Applications, 85 (??):30–41, March 1, 2021. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S089812212100016X.

Henneking:2025:VEM

[HGD25] Stefan Henneking, Jacob Grosek, and Leszek Demkowicz. A vectorial envelope Maxwell formulation for electromagnetic waveguides with application to nonlinear fiber optics. Computers and Mathematics with Applications, 193(??):34–53, September 1, 2025. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122125002305.

He:2018:BDC

[HGW18] Yijun He, Huaihong Gao, and Hua Wang. Blow-up and decay for a class of pseudoparabolic p-Laplacian equation with logarithmic nonlinearity. Computers and Mathematics with Applications, 75(2):459–469, January 15, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122117305904.

See comments [DZ21].

Hu:2021:CSP

[HGW21] Dongdong Hu, Yuezheng Gong, and Yushun Wang. On convergence of a structure preserving difference scheme for two-dimensional space-fractional nonlinear Schrödinger equation and its fast implementation. Computers and Mathematics with Applications, 98(??):10–23, September 15, 2021. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S089812212100256X.

Hashim:2021:FEA

[HH21] Mohammed H. Hashim and Akil J. Harfash. Finite element analysis of a Keller– Segel model with additional cross-diffusion and logistic source. Part I: Space convergence. Computers and Mathematics with Applications, 89(??):44–56, May 1, 2021. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122121000481.

Hashim:2022:FEA

[HH22] Mohammed H. Hashim and Akil J. Harfash. Finite element analysis of a Keller– Segel model with additional cross-diffusion and logistic source. Part II: Time convergence and numerical simulation. Computers and Mathematics with Applications, 109 (??):216–234, March 1, 2022. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122122000384.

Heydari:2023:EJS

[HH23] M. H. Heydari and M. Hosseininia. An efficient Jacobi spectral method for variable-order time fractional 2D Wu–Zhang system. Computers and Mathematics with Applications, 140 (??):89–106, June 15, 2023. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122123001128.

Hussain:2020:MRM

[HHG20] Manzoor Hussain, Sirajul Haq, and Abdul Ghafoor.

Meshless RBFs method for numerical solutions of twodimensional high order fractional Sobolev equations. Computers and Mathematics with Applications, 79(3): 802–816, February 1, 2020. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S089812211930389X.

Han:2020:NAC

[HHJO20] Danfu Han, Weimin Han, Michal Jureczka, and Anna Ochal. Numerical analysis of a contact problem with wear. Computers and Mathematics with Applications, 79(10): 2942–2951, May 15, 2020. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122120300018.

Hu:2021:FMB

[HHL^{+}21] Bin Hu, Zongjun Hu, Cong Li, Zhongrong Niu, and Xiaobao Li. A fast multipole boundary element method based on higher order elements for analyzing 2-D potential problems. Computers and Mathematics with Applications, 87 (??):65–76, April 1, 2021. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S089812212100050X.

Huang:2022:HHP

[HHL22] Yao Huang, Wenrui Hao, and Guang Lin. HomPINNs: Homotopy physics-informed neural networks for learning multiple solutions of nonlinear elliptic differential equations. Computers and Mathematics with Applications, 121(??):62–73, September 1, 2022. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122122002851.

He:2021:OCC

[HHM21] Cuiyu He, Weiwei Hu, and Lin Mu. Optimal control of convection-cooling and numerical implementation. Computers and Mathematics with Applications, 92 (??):48–61, June 15, 2021. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S089812212100095X.

Herrero-Hervas:2023:CMN

[HHNV23] F. Herrero-Hervás, M. Negreanu, and A. M. Vargas. Convergence of a meshless numerical method for a chemotaxis system with density-suppressed motility. Computers and Mathematics with Applications, 148(??): 293–301, October 15, 2023. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122123003607.

Haussmann:2020:GIS

[HHR^{+}20] Marc Haussmann, Nicolas Hafen, Florian Raichle, Robin Trunk, Hermann Nirschl, and Mathias J. Krause. Galilean invariance study on different lattice Boltzmann fluid–solid interface approaches for vortexinduced vibrations. Computers and Mathematics with Applications, 80(5):671–691, September 1, 2020. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122120301693.

Hosseininia:2022:MLP

[HHR22] M. Hosseininia, M. H.

Heydari, and M. Razzaghi. Meshless local Petrov– Galerkin method for 2D fractional Fokker–Planck equation involved with the ABC fractional derivative. Computers and Mathematics with Applications, 125(??):176– 192, November 1, 2022. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122122003674.

Houston:2024:ISM

[HHR24] Paul Houston, Matthew E.

Hubbard, and Thomas J.

Radley. Iterative solution methods for high-order/hp– DGFEM approximation of the linear Boltzmann transport equation. Computers and Mathematics with Applications, 166(??):37– 49, July 15, 2024. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122124001676.

Hou:2021:EOH

[HHW21] Yanren Hou, Yongbin Han, and Jing Wen. An equalorder hybridized discontinuous Galerkin method with a small pressure penalty parameter for the Stokes equations. Computers and Mathematics with Applications, 93(??):58–65, July 1, 2021. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122121001127.

He:2022:LBM

[HHY^{+}22] Qiang He, Weifeng Huang, Yuan Yin, Decai Li, and Yuming Wang. A lattice Boltzmann model for liquidvapor-solid flow with thermal phase change. Computers and Mathematics with Applications, 114(??):60– 72, May 15, 2022. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122122001183.

Helml:2023:PCG

[HIP23] Valentin Helml, Michael Innerberger, and Dirk Praetorius. Plain convergence of goal-oriented adaptive FEM. Computers and Mathematics with Applications, 147(??): 130–149, October 1, 2023. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122123003152.

Huang:2023:OTD

[HJ23] Shuangxi Huang and Feng-Fei Jin. Output tracking for a 1-D wave equation with velocity recirculation via boundary control. Computers and Mathematics with Applications, 148(??): 228–240, October 15, 2023. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122123003498.

Han:2024:NAI

[HJ24] Wei-Wei Han and Yao-Lin Jiang. Numerical analysis of an improved projection method for the evolutionary magnetohydrodynamic equations with modular grad-div stabilization. Computers and Mathematics with Applications, 167(??): 298–314, August 1, 2024.

CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122124002438.

Huang:2025:PBO

[HJ25] Shuangxi Huang and Feng-Fei Jin. Performance boundary output tracking for 1-D wave equation with velocity recirculation and known exosystem. Computers and Mathematics with Applications, 199(??):273–285, December 1, 2025. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122125004262.

Han:2023:NAP

[HJM23a] Wei-Wei Han, Yao-Lin Jiang, and Zhen Miao. Numerical analysis of projection methods for the timedependent Navier–Stokes equations with modular grad-div stabilization. Computers and Mathematics with Applications, 141(??):145– 158, July 1, 2023. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122123001530.

Han:2023:SOD

[HJM23b] Wei-Wei Han, Yao-Lin Jiang, and Zhen Miao. On a second-order decoupled timestepping scheme for solving a finite element problem for the approximation of Peterlin viscoelastic model. Computers and Mathematics with Applications, 142 (??):48–63, July 15, 2023. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122123001438.

Han:2023:RER

[HJP23] Xiaojing Han, Oscar Sánchez Jiménez, and Emmanuel Pagnacco. Response and EPSD of rotor-blade nonlinear system with nonstationary non-Gaussian stochastic excitation via PGHW method. Computers and Mathematics with Applications, 142(??):140–156, July 15, 2023. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122123001566.

Haug:2025:SMP

[HJT25] Jonas Haug, Rachel Jewell, and Ray Treinen. Spectral methods for prescribed mean curvature equations. Computers and Mathematics with Applications, 197 (??):1–18, November 1, 2025. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122125003220.

He:2025:HAE

[HJTW25] Zhijuan He, Kai Jiang, Liwei Tan, and Xin Wang. High-accurate and efficient numerical algorithms for the self-consistent field theory of liquid-crystalline polymers. Computers and Mathematics with Applications, 194(??): 31–52, September 15, 2025. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122125002500.

Hu:2023:CNL

[HJXW23] Dongdong Hu, Huiling Jiang, Zhuangzhi Xu, and Yushun Wang. On convergence of a novel linear conservative scheme for the twodimensional fractional nonlinear Schrödinger equation with wave operator. Computers and Mathematics with Applications, 150(??):254– 266, November 15, 2023. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122123004017.

Han:2023:DIR

[HJZ23] Huan Han, Qi Jin, and Yimin Zhang. A 2D diffeomorphic image registration model with inequality constraint. Computers and Mathematics with Applications, 132(??):135–144, February 15, 2023. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122122005193.

Halim:2020:PME

[HK20a] Abdul Halim and B. V. Rathish Kumar. A T V − L^{2} − H^{−1} PDE model for effective denoising. Computers and Mathematics with Applications, 80(10):2176– 2193, November 15, 2020. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122120303655.

Halim:2020:APM

[HK20b] Abdul Halim and B. V. Rathish Kumar. An anisotropic PDE model for image inpainting. Computers and Mathematics with Applications, 79(9): 2701–2721, May 1, 2020. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122119305607.

Hartmann:2022:EEB

[HK22] Frank Hartmann and Stefan Kollmannsberger. Enforcing essential boundary conditions on domains defined by point clouds. Computers and Mathematics with Applications, 113(??):13– 23, May 1, 2022. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S089812212200092X.

Heiland:2024:CAC

[HK24] Jan Heiland and Yongho Kim. Convolutional autoencoders, clustering, and POD for low-dimensional parametrization of flow equations. Computers and Mathematics with Applications, 175(??):49–61, December 1, 2024. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122124003997.

Ham:2024:NPF

[HKH^{+}24] Seokjun Ham, Hyundong Kim, Youngjin Hwang, Soobin Kwak, Jyoti, Jian Wang, Heming Xu, Wenjing Jiang, and Junseok Kim. A novel phase-field model for three-dimensional shape transformation. Computers and Mathematics with Applications, 176(??):67–76, December 15, 2024. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122124004164.

Hafeez:2023:MTG

[HKJ23] Muhammad Bilal Hafeez, Marek Krawczuk, and Wasim Jamshed. Maximum transportation growth in energy and solute particles in Prandtl martial across a vertical 3D-heated surface: Simulations achieved using by finite element approach. Computers and Mathematics with Applications, 133 (??):48–60, March 1, 2023. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122122004953.

Huang:2023:NUM

[HKL^{+}23] Yichen Huang, Zhaoqing Ke, Zhihao Li, Yan Gao, Zhaoxuan Tang, and Ying Zhang. A non-uniform magnetic field coupled lattice Boltzmann model and its application on the wetting dynamics of a ferrofluid droplet under gravity effects. Computers and Mathematics with Applications, 143 (??):73–93, August 1, 2023. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122123001797.

Holter:2021:RPC

[HKM21] Karl Erik Holter, Miroslav Kuchta, and Kent-Andre Mardal. Robust preconditioning for coupled Stokes– Darcy problems with the Darcy problem in primal form. Computers and Mathematics with Applications, 91(??):53–66, June 1, 2021. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122120303291.

Helsing:2022:EFW

[HKR22] Johan Helsing, Anders Karlsson, and Andreas Rosén. An efficient full-wave solver for eddy currents. Computers and Mathematics with Applications, 128(??):145– 162, December 15, 2022. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122122004412.

Havelka:2021:EIS

[HKS21] Jan Havelka, Anna Kučerová, and Jan Sýkora. Efficient inverse solvers for thermal tomography. Computers and Mathematics with Applications, 97(??):314–328, September 1, 2021. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122121002431.

Hao:2022:NMS

[HL22] Yong-Xia Hao and Ting Li. A new method to solve the triangular Plateau– Bézier problem. Computers and Mathematics with Applications, 125(??):101–115, November 1, 2022. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122122003649.

Hong:2024:SHT

[HLCW24a] Yongxing Hong, Ji Lin, Alexander H. D. Cheng, and Yanjie Wang. Simulation of heat transfer in Poiseuille pipe flow via generalized finite difference method with a space stepping algorithm. Computers and Mathematics with Applications, 163 (??):136–145, June 1, 2024. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122124001299.

Huang:2024:IOF

[HLCW24b] Yuhao Huang, Qilin Liu, Zhenhua Chai, and Binghai Wen. Isotropic optimizations of finite difference discretization. Computers and Mathematics with Applications, 168(??):190– 206, August 15, 2024. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122124002591.

Hashemian:2020:AIF

[HLEF20] Ali Hashemian, Esmail Lakzian, and Amir Ebrahimi-Fizik. On the application of isogeometric finite volume method in numerical analysis of wet-steam flow through turbine cascades. Computers and Mathematics with Applications, 79(6): 1687–1705, March 15, 2020. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S089812211930481X.

He:2020:LBM

[HLH^{+}20a] Qiang He, Yongjian Li, Weifeng Huang, Yang Hu, Decai Li, and Yuming Wang. Lattice Boltzmann model for dense suspended particles based on improved bounce-back method. Computers and Mathematics with Applications, 80(3):552–567, August 1, 2020. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122120301462.

He:2020:ULB

[HLH^{+}20b] Qiang He, Yongjian Li, Weifeng Huang, Yang Hu, Decai Li, and Yuming Wang. A unified lattice Boltzmann model for immiscible and miscible ternary fluids. Computers and Mathematics with Applications, 80(12):2830– 2859, December 15, 2020. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122120304077.

Han:2024:MSS

[HLH24] Yanxiang Han, Zeyu Lu, and Juan Huang. Multiscale simulation scheme by using a finite-difference derivative framework for lubrication surface. Computers and Mathematics with Applications, 155(??):176–192, February 1, 2024. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122123005606.

Hou:2025:UOE

[HLJ25] Dianming Hou, Hui Liu, and Lili Ju. Unconditionally original energy-dissipative and MBP-preserving Crank– Nicolson scheme for the Allen–Cahn equation with general mobility. Computers and Mathematics with Applications, 191(??):86–104, August 1, 2025. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122125001749.

Hao:2024:CBM

[HLL24] Wenrui Hao, Sun Lee, and Young Ju Lee. Companionbased multi-level finite element method for computing multiple solutions of nonlinear differential equations. Computers and Mathematics with Applications, 168(??): 162–173, August 15, 2024. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S089812212400261X.

Harizanov:2020:NSF

[HLMM20] Stanislav Harizanov, Raytcho Lazarov, Svetozar Margenov, and Pencho Marinov. Numerical solution of fractional diffusion–reaction problems based on BURA. Computers and Mathematics with Applications, 80(2): 316–331, July 2020. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122119303402.

Hu:2024:NML

[HLN24] Bin Hu, Cong Li, and Zhongrong Niu. A new multilevel strategy of numerical integration in the fast multipole BEM for analyzing 3D potential problems. Computers and Mathematics with Applications, 161 (??):174–189, May 1, 2024. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122124000920.

Hu:2022:EAM

[HLPW22] Hongling Hu, Ming Li, Kejia Pan, and Pinxia Wu. An extrapolation accelerated multiscale Newton-MG method for fourth-order compact discretizations of semilinear Poisson equations. Computers and Mathematics with Applications, 113(??):189– 197, May 1, 2022. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122122000888.

He:2024:RGM

[HLQ24] Lu He, Yanhong Liu, and Guanghui Qing. Rational generalized mixed finite element method for 2D linear elastic problems. Computers and Mathematics with Applications, 163(??):103– 116, June 1, 2024. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122124001196.

Huang:2021:BLS

[HLW21] Jiangfeng Huang, Zhaoxing Li, and Bo Wang. A Bayesian level set method for the shape reconstruction of inverse scattering problems in elasticity. Computers and Mathematics with Applications, 97(??):18–27, September 1, 2021. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122121002108.

Han:2023:PEN

[HLW23] Feng Han, Yu Liu, and Jianguo Wang. Posteriori error neural network: a recovery type posteriori error estimator based on neural network for diffusion problems. Computers and Mathematics with Applications, 143(??): 169–188, August 1, 2023. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122123002067.

Hu:2024:LIS

[HLW24] Jun Hu, Ting Lin, and Qingyu Wu. Local C^{r} interpolations on simplicial grids in any dimension. Computers and Mathematics with Applications, 154(??):153–161, January 15, 2024. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122123005229.

Han:2025:NFB

[HLWL25] Yanqun Han, Tianhao Liu, Weibin Wen, and Xiaomin Liu. A new family of Bspline based explicit time integration methods for linear structural dynamic analysis. Computers and Mathematics with Applications, 185(??):29–51, May 1, 2025. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122125000707.

He:2020:LUH

[HLX20] Wenming He, Xiong Liu, and Jin Xiao. Local ultraconvergence of high order finite element method by interpolation postprocessing technique for elliptic problems with constant coefficients. Computers and Mathematics with Applications, 79 (9):2492–2502, May 1, 2020. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122119305498.

He:2022:RDN

[HLX22] Juncai He, Lin Li, and Jinchao Xu. ReLU deep neural networks from the hierarchical basis perspective. Computers and Mathematics with Applications, 120(??): 105–114, August 15, 2022. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122122002516.

Huang:2022:DAF

[HLY22] Yunqing Huang, Jichun Li, and Wei Yang. Developing and analyzing a finite element method for simulating wave propagation in graphene-based absorber. Computers and Mathematics with Applications, 122(??): 76–92, September 15, 2022. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122122003078.

Hossain:2024:DPC

[HLZ24] Khan Enaet Hossain, Dong Liang, and Hongmei Zhu. Developing PDE-constrained optimal control of multicomponent contamination flows in porous media. Computers and Mathematics with Applications, 176(??):224– 243, December 15, 2024. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122124004760.

Herrmann:2020:AET

[HM20] Samuel Herrmann and Nicolas Massin. Approximation of exit times for onedimensional linear diffusion processes. Computers and Mathematics with Applications, 80(6):1668–1682, September 15, 2020. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122120302893.

Hill:2024:LAM

[HM24] Róisín Hill and Niall Madden. Layer-adapted meshes for singularly perturbed problems via mesh partial differential equations and a posteriori information. Computers and Mathematics with Applications, 168 (??):1–9, August 15, 2024. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122124002311.

Hahn:2024:LRE

[HMF24] Jooyoung Hahn, Karol Mikula, and Peter Frolkovič. Laplacian regularized eikonal equation with Soner boundary condition on polyhedral meshes. Computers and Mathematics with Applications, 156(??):74–86, February 15, 2024. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122123005655.

Haggar:2025:LTS

[HMFM25] M. S. Daoussa Haggar, K. Mahamat Malloum, J. M. Fokam, and M. Mbehou. A linearized time stepping scheme for finite elements applied to Gray–Scott model.

Computers and Mathematics with Applications, 191(??): 129–143, August 1, 2025. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122125001804.

Hill:2024:DVP

[HMŁW^{+}24] B. M. Hill, T. R. Mitchell, Ł. Łaniewski-Wołłk, S. M. Aminossadati, and C. R. Leonardi. Development and validation of a phase-field lattice Boltzmann method for non-Newtonian Herschel– Bulkley fluids in three dimensions. Computers and Mathematics with Applications, 176(??):398–414, December 15, 2024. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S089812212400484X.

Han:2021:DAT

[HMW21] Bin Han, Michelle Michelle, and Yau Shu Wong. Dirac assisted tree method for 1D heterogeneous Helmholtz equations with arbitrary variable wave numbers. Computers and Mathematics with Applications, 97(??): 416–438, September 1, 2021. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122121002534.

Huo:2021:JJB

[HMX21] Zenan Huo, Gang Mei, and Nengxiong Xu. juSFEM: a Julia-based open-source package of parallel Smoothed Finite Element Method (S-FEM) for elastic problems. Computers and Mathematics with Applications, 81(??): 459–477, January 1, 2021. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122120300523.

Hervella-Nieto:2020:RDA

[HNLPP20] Luis Hervella-Nieto, Paula M.

López-Pérez, and Andrés Prieto. Robustness and dispersion analysis of the Partition of Unity Finite Element Method applied to the Helmholtz equation. Computers and Mathematics with Applications, 79(8): 2426–2446, April 15, 2020. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122119305425.

Hakula:2025:CMC

[HNV25] Harri Hakula, Mohamed M. S. Nasser, and Matti Vuorinen. Constrained maximization of conformal capacity. Computers and Mathematics with Applications, 193(??):132–156, September 1, 2025. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122125002251.

Hoang:2021:RAC

[Hoa21] Quan Pham Hoang. Recovering the aqueous concentration in a multi-layer porous media. Computers and Mathematics with Applications, 100(??):83–98, October 15, 2021. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122121003175.

Hoang:2022:FIL

[Hoa22] Thi-Thao-Phuong Hoang. Fully implicit local timestepping methods for advection– diffusion problems in mixed formulations. Computers and Mathematics with Applications, 118(??):248– 264, July 15, 2022. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122122002164.

Hofreither:2020:UVS

[Hof20] Clemens Hofreither. A unified view of some numerical methods for fractional diffusion. Computers and Mathematics with Applications, 80(2):332–350, July 2020. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122119303773.

Hoppe:2023:STA

[Hop23] R. H. W. Hoppe. A spacetime adaptive discontinuous Galerkin method for the numerical solution of the dynamic quasi-static von Kármán equations. Computers and Mathematics with Applications, 132(??):32–47, February 15, 2023. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122122005053.

Heister:2024:AST

[HOY24] Timo Heister, Maxim A.

Olshanskii, and Vladimir Yushutin. An adaptive stabilized trace finite element method for surface PDEs. Computers and Mathematics with Applications, 171(??): 164–174, October 1, 2024. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122124003171.

He:2022:NSP

[HQCW22] Bing He, Chunyan Qin, Wenbo Chen, and Binghai Wen. Numerical simulation of pulmonary airway reopening by the multiphase lattice Boltzmann method. Computers and Mathematics with Applications, 108(??): 196–205, February 15, 2022. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S089812212200013X.

Huang:2023:PNF

[HQL23] Yuan-Yuan Huang, Wei Qu, and Siu-Long Lei. On τ-preconditioner for a novel fourth-order difference scheme of two-dimensional Riesz space-fractional diffusion equations. Computers and Mathematics with Applications, 145(??):124–140, September 1, 2023. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122123002675.

He:2020:SSF

[HQLJ20] Wenqiang He, Guoliang Qin, Jingxiang Lin, and Cheng Jia. A segregated spectral finite element method for the 2D transient incompressible Navier–Stokes equations. Computers and Mathematics with Applications, 79(2): 521–537, January 15, 2020. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S089812211930536X.

Hubbard:2020:STR

[HRS20] M. E. Hubbard, M. Ricchiuto, and D. Sármány. Space-time residual distribution on moving meshes. Computers and Mathematics with Applications, 79(5): 1561–1589, March 1, 2020. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122119304699.

Haider:2024:GAC

[HRS24] A. M. Haider, S. Rjasanow, and M. Schanz. Generalised adaptive cross approximation for convolution quadrature based boundary element formulation. Computers and Mathematics with Applications, 175(??):470– 486, December 1, 2024. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122124004681.

Houston:2020:EGP

[HRvdZ20] Paul Houston, Sarah Roggendorf, and Kristoffer G. van der Zee. Eliminating Gibbs phenomena: a non-linear Petrov–Galerkin method for the convection– diffusion–reaction equation. Computers and Mathematics with Applications, 80(5): 851–873, September 1, 2020. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122120301292.

Huang:2020:EAF

[HS20] Chaobao Huang and Martin Stynes. Error analysis of a finite element method with GMMP temporal discretisation for a timefractional diffusion equation. Computers and Mathematics with Applications, 79(9): 2784–2794, May 1, 2020. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122119305644.

Hernandez-Santamaria:2025:FPI

[HSJSS25] Víctor Hernández-Santamaría, Sven Jarohs, Alberto Saldaña, and Leonard Sinsch. FEM for 1D-problems involving the logarithmic Laplacian: Error estimates and numerical implementation. Computers and Mathematics with Applications, 192(??):189– 211, August 15, 2025. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122125002093.

Hasbani:2023:ASF

[HSM^{+}23] Jose G. Hasbani, Paulina Sepúlveda, Ignacio Muga, Victor M. Calo, and Sergio Rojas. Adaptive stabilized finite elements via residual minimization onto bubble enrichments. Computers and Mathematics with Applications, 151(??):1–11, December 1, 2023. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122123003954.

Hashemian:2023:SBV

[HSR^{+}23] Ali Hashemian, Hanna Sliusarenko, Sara Remogna, Domingo Barrera, and Michael Bartoň. Solving boundary value problems via the Nyström method using spline Gauss rules. Computers and Mathematics with Applications, 143 (??):33–47, August 1, 2023. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122123001785.

Halder:2025:PIN

[HSR25] Rahul Halder, Giovanni Stabile, and Gianluigi Rozza. Physics informed neural network framework for unsteady discretized reduced order system. Computers and Mathematics with Applications, 191(??):167– 187, August 1, 2025. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122125001944.

Han:2021:NAD

[HSWG21] Weimin Han, Chenghang Song, Fei Wang, and Jinghuai Gao. Numerical analysis of the diffusive-viscous wave equation. Computers and Mathematics with Applications, 102(??):54–64, November 15, 2021. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122121003515.

Helmig:2025:VCC

[HSY25] Rainer Helmig, Martin Schneider, and Ivan Yotov. Vertex-centered control-volume mimetic finite difference methods. Computers and Mathematics with Applications, 196(??):104–126, October 15, 2025. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122125003062.

He:2022:FEA

[HSZ22] Mingyan He, Pengtao Sun, and Hui Zhao. Finite element approximations to a fourthorder modified Poisson– Fermi equation for electrostatic correlations in concentrated electrolytes. Computers and Mathematics with Applications, 117(??):229– 244, July 1, 2022. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122122001924.

Hao:2020:SRI

[HTBMC20] Dinh Nho Hào, Phan Xuan Thanh, Bandar Bin-Mohsin, and Nguyen Huu Cong. Stable reconstruction of the initial condition in parabolic equations from boundary observations. Computers and Mathematics with Applications, 79(12):3570– 3587, June 15, 2020. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122120300869.

Hao:2023:CIP

[HTDT23] Dinh Nho Hào, Nguyen Trung Thành, Nguyen Van Duc, and Nguyen Van Thang. A coefficient identification problem for a system of advection-reaction equations in water quality modeling. Computers and Mathematics with Applications, 148(??): 126–150, October 15, 2023. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S089812212300336X.

Hung:2024:IFV

[HTPV24] P. T. Hung, Chien H. Thai, and P. Phung-Van. Isogeometric free vibration of functionally graded porous magneto-electro-elastic plate reinforced with graphene platelets resting on an elastic foundation. Computers and Mathematics with Applications, 169(??):68–87, September 1, 2024. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122124002736.

Hung:2025:NMM

[HTPV25] P. T. Hung, Chien H. Thai, and P. Phung-Van. A novel meshfree method for free vibration behavior of the functionally graded carbon nanotube-reinforced composite plates using a new shear deformation theory. Computers and Mathematics with Applications, 189(??):208– 224, July 1, 2025. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122125001762.

Hamid:2025:NCS

[HUT25] Muhammad Hamid, Muhammad Usman, and Zhenfu Tian. Novel connection of spectral scheme and onestep of s-order approaches for MHD flows enclosed a duct. Computers and Mathematics with Applications, 184 (??):185–220, April 15, 2025. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122125000690.

Hernandez:2021:LFF

[HV21] Erwin Hernández and Jesus Vellojin. A locking-free finite element formulation for a non-uniform linear viscoelastic Timoshenko beam. Computers and Mathematics with Applications, 99(??): 305–322, October 1, 2021. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122121002984.

Huang:2025:ATM

[HW25] Yating Huang and Zhenyou Wang. Analysis of tumor metastasis model in microenvironment based on coupled fractional reaction diffusion equation. Computers and Mathematics with Applications, 182(??):24–45, March 15, 2025. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122125000136.

Huang:2024:NEV

[HWC24] Yuxiang Huang, Feng Wang, and Jinru Chen. A nonconforming extended virtual element method for Stokes interface problems. Computers and Mathematics with Applications, 175(??):509–535, December 1, 2024. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S089812212400470X.

Weng:2021:RMN

[hWCsLjH21] Yun hua Weng, Tao Chen, Xue song Li, and Nan jing Huang. Rothe method and numerical analysis for a new class of fractional differential hemivariational inequality with an application. Computers and Mathematics with Applications, 98(??):118– 138, September 15, 2021. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122121002595.

Harper:2020:LFS

[HWL^{+}20] Graham Harper, Ruishu Wang, Jiangguo Liu, Simon Tavener, and Ran Zhang. A locking-free solver for linear elasticity on quadrilateral and hexahedral meshes based on enrichment of Lagrangian elements. Computers and Mathematics with Applications, 80(6):1578– 1595, September 15, 2020. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122120302807.

He:2021:MDG

[HWW21] Limin He, Fei Wang, and Jing Wen. A mixed discontinuous Galerkin method for the wave equation. Computers and Mathematics with Applications, 82(??):60–73, January 15, 2021. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122120304521.

Huo:2024:LFW

[HWWZ24] Fuchang Huo, Ruishu Wang, Yanqiu Wang, and Ran Zhang. A locking-free weak Galerkin finite element method for linear elasticity problems. Computers and Mathematics with Applications, 160(??):181– 190, April 15, 2024. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122124000774.

He:2023:DAF

[HWY23] Bin He, Hao Wang, and Wei Yang. Design and adaptive finite element simulation of conformal transformation optics devices with isotropic materials. Computers and Mathematics with Applications, 144(??):198– 209, August 15, 2023. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122123002304.

Huang:2025:CBN

[HWY25] Rong Huang, Zhifeng Weng, and Jianhua Yuan. Collocationbased numerical simulation of multi-dimensional nonlinear time-fractional Schrödinger equations. Computers and Mathematics with Applications, 183(??):214– 233, April 1, 2025. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122125000495.

He:2023:USG

[HWZQ23] Dongze He, Qingshan Wang, Rui Zhong, and Bin Qin. A unified spectral-geometric model of FGM double conical/cylindrical/spherical shell coupled with annular plates. Computers and Mathematics with Applications, 143(??):348–371, August 1, 2023. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122123001931.

Hou:2022:NAT

[HX22] Yanren Hou and Dandan Xue. Numerical analysis of two-grid decoupling finite element scheme for Navier–Stokes/Darcy model.

Computers and Mathematics with Applications, 113 (??):45–51, May 1, 2022. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122122000943.

Han:2023:SST

[HXW23] Ziwei Han, Heming Xu, and Jian Wang. A simple shape transformation method based on phasefield model. Computers and Mathematics with Applications, 147(??):121–129, October 1, 2023. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122123003140.

Huang:2022:BEA

[HXZ^{+}22] Rongjie Huang, Guizhong Xie, Yudong Zhong, Hongrui Geng, Hao Li, and Liangwen Wang. Boundary element analysis of thin structures using a dual transformation method for weakly singular boundary integrals. Computers and Mathematics with Applications, 113 (??):198–213, May 1, 2022. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122122000979.

Hou:2023:FDS

[HYH23] Bingrui Hou, Maoqin Yuan, and Pengzhan Huang. Fully discrete scheme for a timedependent Ginzburg–Landau equation in macromolecular microsphere composite hydrogels. Computers and Mathematics with Applications, 151(??):127–133, December 1, 2023. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122123004236.

Hou:2020:NAU

[HYJ20] Yuanyuan Hou, Wenjing Yan, and Feifei Jing. Numerical analysis of the unconditionally stable discontinuous Galerkin schemes for the nonstationary conduction– convection problem. Computers and Mathematics with Applications, 80(6):1479– 1499, September 15, 2020. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122120302741.

Huang:2025:RBE

[HYL^{+}25] W. Huang, X. B. Yan, J. X. Liao, L. K. Feng, and P. H. Wen. Regular boundary element method for composite shear deformable plate and shell. Computers and Mathematics with Applications, 199(??):106–126, December 1, 2025. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122125003967.

Hou:2022:DIF

[HYLH22] Yuanyuan Hou, Wenjing Yan, Maojun Li, and Xiaoming He. A decoupled and iterative finite element method for generalized Boussinesq equations. Computers and Mathematics with Applications, 115 (??):14–25, June 1, 2022. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122122001523.

Hou:2024:LAL

[HYW^{+}24] Jian Hou, Yongguang Yu, Jingjia Wang, Hongpeng Ren, and Xiangyun Meng. Local analysis of L1-finite difference method on graded meshes for multi-term twodimensional time-fractional initial-boundary value problem with Neumann boundary conditions. Computers and Mathematics with Applications, 157(??):209– 214, March 1, 2024. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122124000026.

Huang:2024:COR

[HYZZ24] Qiong-Ao Huang, Cheng Yuan, Gengen Zhang, and Lian Zhang. A computationally optimal relaxed scalar auxiliary variable approach for solving gradient flow systems. Computers and Mathematics with Applications, 156(??):64–73, February 15, 2024. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122123005667.

Han:2021:GFE

[HZ21] Hao Han and Chengjian Zhang. Galerkin finite element methods solving 2D initial–boundary value problems of neutral delayreaction–diffusion equations. Computers and Mathematics with Applications, 92 (??):159–171, June 15, 2021. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122121001115.

He:2024:GNM

[HZ24a] Huizhi He and Xiaobing Zhang. A general numerical method for solving the three-dimensional hyperbolic heat conduction equation on unstructured grids. Computers and Mathematics with Applications, 158 (??):85–94, March 15, 2024.

CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122124000129.

He:2024:ANM

[HZ24b] Wenlong He and Jiwei Zhang. Analysis of new mixed finite element method for a Barenblatt–Biot poroelastic model. Computers and Mathematics with Applications, 171(??):34–59, October 1, 2024. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122124003109.

He:2024:LFN

[HZ24c] Wenlong He and Jiwei Zhang. A locking free numerical method for the poroelasticity-Forchheimer model. Computers and Mathematics with Applications, 174(??):183–203, November 15, 2024. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122124003869.

He:2025:OMN

[HZ25] Wenming He and Ren Zhao.

Optimal maximum norm estimates of virtual element methods for elliptic problem in three dimensions. Computers and Mathematics with Applications, 197(??): 167–182, November 1, 2025. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122125003384.

Hu:2022:TCE

[HZA22] Shuaifei Hu, Guomei Zhao, and Rong An. Temporal convergence of extrapolated BDF-2 scheme for the Maxwell–Landau– Lifshitz equations. Computers and Mathematics with Applications, 119(??):278– 287, August 1, 2022. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122122002309.

Han:2022:NSM

[HZJ^{+}22] Huiran Han, Jiansong Zhang, Bingjie Ji, Yue Yu, and Yun Yu. A new symmetric mixed element method for semi-linear parabolic problem based on two-grid discretization. Computers and Mathematics with Applications, 108(??):206–215, February 15, 2022. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122122000256.

Heydari:2023:JPN

[HZR23] M. H. Heydari, Sh. Zhagharian, and M. Razzaghi. Jacobi polynomials for the numerical solution of multidimensional stochastic multiorder time fractional diffusionwave equations. Computers and Mathematics with Applications, 152(??):91–101, December 15, 2023. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S089812212300456X.

Hong:2021:EPR

[HZW21] Qi Hong, Jia Zhao, and Qi Wang. Energy-production-rate preserving numerical approximations to network generating partial differential equations. Computers and Mathematics with Applications, 84(??):148–165, February 15, 2021. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122120304429.

Ismail:2024:DML

[IBY^{+}24] Ismail, B. S. Bhadauria, Moh Yaseen, Sawan Kumar Rawat, and Manish Pant. Designing machine learning based intelligent network for assessment of heat transfer performance of ternary hybrid nanofluid flow between a cone and a disk: Case of MLP feed forward neural network. Computers and Mathematics with Applications, 169(??):17–38, September 1, 2024. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122124002712.

Ito:2024:IRR

[IJS^{+}24] Shota Ito, Julius Jeßberger, Stephan Simonis, Fedor Bukreev, Adrian Kummerländer, Alexander Zimmermann, Gudrun Thäter, Georg R. Pesch, Jorg Thöming, and Mathias J. Krause. Identification of reaction rate parameters from uncertain spatially distributed concentration data using gradientbased PDE constrained optimization. Computers and Mathematics with Applications, 167(??):249–263, August 1, 2024. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122124002451.

Iserles:2024:EAS

[IK24] Arieh Iserles and Karolina Kropielnicka. An elementary approach to splittings of unbounded operators. Computers and Mathematics with Applications, 176(??):29–34, December 15, 2024. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122124003985.

Ishizaka:2021:NGC

[IKST21] Hiroki Ishizaka, Kenta Kobayashi, Ryo Suzuki, and Takuya Tsuchiya. A new geometric condition equivalent to the maximum angle condition for tetrahedrons. Computers and Mathematics with Applications, 99(??):323–328, October 1, 2021. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122121003084.

Ilati:2020:DMS

[Ila20] Mohammad Ilati. DMLPG method for specifying a control function in twodimensional parabolic inverse PDEs. Computers and Mathematics with Applications, 80(5):604–616, September 1, 2020. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122120301541.

Imani:2020:LBM

[Ima20] Gholamreza Imani. Lattice Boltzmann method for conjugate natural convection with heat generation on nonuniform meshes. Computers and Mathematics with Applications, 79(4):1188–1207, February 15, 2020. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122119304249.

Ibrahim:2020:PND

[IQS20] Moustafa Ibrahim, El Houssaine Quenjel, and Mazen Saad. Positive nonlinear DDFV scheme for a degenerate parabolic system describing chemotaxis. Computers and Mathematics with Applications, 80(12):2972– 3003, December 15, 2020. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S089812212030420X.

Issakhov:2024:FSI

[ISA24] Alibek Issakhov, Aidana Sabyrkulova, and Aizhan Abylkassymova. The fluidstructure interaction during blood flow in a flexible stenotic thoracic aorta: Numerical study. Computers and Mathematics with Applications, 165(??):39– 51, July 1, 2024. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122124001500.

Iqbal:2020:NSH

[IZ20] Sehar Iqbal and Paul Andries Zegeling. A numerical study of the higher-dimensional Gelfand–Bratu model. Computers and Mathematics with Applications, 79(6):1619– 1633, March 15, 2020. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122119304687.

Jia:2023:PBS

[JAHM23] Xiaowei Jia, Hengbin An, Yi Hu, and Zeyao Mo. A physics-based strategy for choosing initial iterate for solving drift-diffusion equations. Computers and Mathematics with Applications, 131 (??):1–13, February 1, 2023. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122122004989.

Jiang:2024:NCH

[JAMX24] Dandan Jiang, Mejdi Azaiez, Alain Miranville, and Chuanju Xu. Nonlocal Cahn– Hilliard type model for image inpainting. Computers and Mathematics with Applications, 159(??):76– 91, April 1, 2024. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122124000336.

Jaworska:2021:GMM

[Jaw21] Irena Jaworska. Generalization of the multipoint meshless FDM application to the nonlinear analysis. Computers and Mathematics with Applications, 87 (??):1–11, April 1, 2021. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122121000365.

Jones:2023:GML

[JBKW23] Andrew M. Jones, Peter A. Bosler, Paul A. Kuberry, and Grady B. Wright. Generalized moving least squares vs. radial basis function finite difference methods for approximating surface derivatives. Computers and Mathematics with Applications, 147 (??):1–13, October 1, 2023. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122123003097.

Jia:2024:IUE

[JCJ24] Junqing Jia, Xiaoqing Chi, and Xiaoyun Jiang. Improved uniform error bounds for long-time dynamics of the high-dimensional nonlinear space fractional sine-Gordon equation with weak nonlinearity. Computers and Mathematics with Applications, 175(??):62–86, December 1, 2024. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122124004048.

Jiammeepreecha:2025:ENF

[JCP^{+}25] Weeraphan Jiammeepreecha, Komkorn Chaidachatorn, Boonchai Phungpaingam, Karun Klaycham, and Somchai Chucheepsakul. Effective nonlocal finite element formulation for free vibration analysis of S-FGM doubly curved nanoshells based on linear strain–displacement relations using TSDT. Computers and Mathematics with Applications, 179(??):77– 102, February 1, 2025. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122124005170.

Jourabian:2014:EMP

[JFS14] Mahmoud Jourabian, Mousa Farhadi, and Kurosh Sedighi. On the expedited melting of phase change material (PCM) through dispersion of nanoparticles in the thermal storage unit. Computers and Mathematics with Applications, 67 (7):1358–1372, April 2014. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122114000625.

See retraction notice [JFS20].

Jourabian:2020:RNE

[JFS20] Mahmoud Jourabian, Mousa Farhadi, and Kurosh Sedighi. Retraction notice to “On the expedited melting of phase change material (PCM) through dispersion of nanoparticles in the thermal storage unit” [Comput. Math. Appl. 67(7) (2014) 1358–1372]. Computers and Mathematics with Applications, 80(12): 3193, December 15, 2020. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122120304144.

See [JFS14].

Jacobson:2024:SPD

[JH24] Alon Jacobson and Xiaozhe Hu. Structure-preserving discretization of fractional vector calculus using discrete exterior calculus. Computers and Mathematics with Applications, 153(??):186– 196, January 1, 2024. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122123005035.

Jha:2021:RBP

[Jha21]  Abhinav Jha.  A resid-

ual based a posteriori error estimators for AFC schemes for convectiondiffusion equations. Computers and Mathematics with Applications, 97(??):86–99, September 1, 2021. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122121002236.

Jian:2021:FSO

[JHG^{+}21] Huan-Yan Jian, Ting-Zhu Huang, Xian-Ming Gu, Xi-Le Zhao, and Yong-Liang Zhao. Fast second-order implicit difference schemes for time distributed-order and Riesz space fractional diffusionwave equations. Computers and Mathematics with Applications, 94(??):136– 154, July 15, 2021. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122121001760.

Joe:2024:DHA

[JHS24] Beom-Jin Joe, Suk-Yoon Hong, and Jee-Hun Song. Direct hydroacoustics analyses of pipe orifice using lattice Boltzmann method. Computers and Mathematics with Applications, 174(??): 142–150, November 15, 2024. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122124003857.

Jimenez-Islas:2020:MSS

[JICRMG^{+}20] Hugo Jiménez-Islas, Mario Calderón-Ramírez, Gloria María Martínez-González, Martha Patricia Calderón-Álvarado, and Juan Manuel Oliveros-Muñoz. Multiple solutions for steady differential equations via hyperspherical path-tracking of homotopy curves. Computers and Mathematics with Applications, 79(8):2216– 2239, April 15, 2020. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122119305188.

Joshi:2023:MBM

[JJ23] Hardik Joshi and Brajesh Kumar Jha. 2D memory-based mathematical analysis for the combined impact of calcium influx and efflux on nerve cells. Computers and Mathematics with Applications, 134 (??):33–44, March 15, 2023. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122122005235.

Joshi:2020:VFM

[JJOG20] Vaibhav Joshi, Rajeev K. Jaiman, and Carl Ollivier-Gooch. A variational flexible multibody formulation for partitioned fluid–structure interaction: Application to bat-inspired drones and unmanned air-vehicles. Computers and Mathematics with Applications, 80(12):2707– 2737, December 15, 2020. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122120303837.

Jia:2021:LLG

[JJZ21] Junqing Jia, Xiaoyun Jiang, and Hui Zhang. An L1 Legendre–Galerkin spectral method with fast algorithm for the two-dimensional nonlinear coupled time fractional Schrödinger equation and its parameter estimation. Computers and Mathematics with Applications, 82 (??):13–35, January 15, 2021. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122120304442.

Jo:2022:MVV

[JK22] Gwanghyun Jo and Do Y.

Kwak. Mixed virtual volume methods for elliptic problems. Computers and Mathematics with Applications, 113 (??):345–352, May 1, 2022. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122122001286.

Jankowska:2024:TRN

[JKC24] Malgorzata A. Jankowska, Andreas Karageorghis, and C. S. Chen. Training RBF neural networks for solving nonlinear and inverse boundary value problems. Computers and Mathematics with Applications, 165 (??):205–216, July 1, 2024. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122124001937.

Janssen:2020:MME

[JKL20] Christian Janßen, Manfred Krafczyk, and Li-Shi Luo. Mesoscopic methods in engineering and science. Computers and Mathematics with Applications, 79(1): 1–2, January 2020. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122119305802.

Jalalian:2025:SSK

[JKS25] Mohsen Jalalian, Manochehr Kazemi, and Mohammad Esmael Samei. Solving the second kind Volterra–Fredholm type of two-dimensional integral equations on nonrectangular domains via radial basis functions. Computers and Mathematics with Applications, 195(??):265– 279, October 1, 2025. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122125003153.

Jia:2019:NAN

[JL19] Jiteng Jia and Sumei Li.

New algorithms for numerically solving a class of bordered tridiagonal systems of linear equations. Computers and Mathematics with Applications, 78 (1):144–151, July 1, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122119301026.

See retraction notice [JL20].

Jia:2020:RNN

[JL20] Jiteng Jia and Sumei Li.

Retraction notice to “New algorithms for numerically solving a class of bordered tridiagonal systems of linear equations” [Comput. Math. Appl. 78 (1) (2019) 144–151]. Computers and Mathematics with Applications, 80(6): 1744, September 15, 2020. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122120303278.

See [JL19].

Jing:2023:BCF

[JL23] Yinlong Jing and Can Li.

Block-centered finite difference method for a tempered subdiffusion model with time-dependent coefficients. Computers and Mathematics with Applications, 145(??): 202–223, September 1, 2023.

CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122123002663.

Jabalameli:2020:WFR

[JM20a] Mozhgan Jabalameli and Davoud Mirzaei. A weakform RBF-generated finite difference method. Computers and Mathematics with Applications, 79(9):2624– 2643, May 1, 2020. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122119305577.

Jung:2020:RMN

[JM20b] SungKi Jung and R. S. Myong. A relaxation model for numerical approximations of the multidimensional pressureless gas dynamics system. Computers and Mathematics with Applications, 80(5):1073–1083, September 1, 2020. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122120302510.

Jangid:2023:ALW

[JM23a] Komal Jangid and Santwana Mukhopadhyay. Application of Legendre wavelet collocation method to the analysis of poro-thermoelastic coupling with variable thermal conductivity. Computers and Mathematics with Applications, 146(??):1–11, September 15, 2023. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122123002626.

Jaskowiec:2023:CFE

[JM23b] Jan Jaśkowiec and Sławomir Milewski. Coupling finite element method with meshless finite difference method by means of approximation constraints. Computers and Mathematics with Applications, 142(??):208– 224, July 15, 2023. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122123001803.

Jang:2024:TLD

[JMCC24] Yongseok Jang, Emeric Martin, Jean-Baptiste Chapelier, and Vincent Couaillier. Two-level dynamic loadbalanced p-adaptive discontinuous Galerkin methods for compressible CFD simulations. Computers and Mathematics with Applications, 176(??):165–178, December 15, 2024. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122124004516.

Jeyakarthikeyan:2020:ERN

[JMM20] P. V. Jeyakarthikeyan, Radha Jeyakarthikeyan M., and Rohith Narendiran M. An efficient and rapid numerical quadrature to generate element matrices for quadrilateral and hexahedral elements in functionally graded materials (FGMs). Computers and Mathematics with Applications, 80 (4):1–35, August 15, 2020. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122120301577.

John:2022:EAS

[JMN22] Volker John, Baptiste Moreau, and Julia Novo. Error analysis of a SUPG-stabilized POD-ROM method for convection–diffusion–reaction equations. Computers and Mathematics with Applications, 122(??):48–60, September 15, 2022. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S089812212200308X.

Jbilou:2024:NAS

[JMR24] Khalide Jbilou, Marilena Mitrouli, and Ahmed Ratnani. Numerical analysis and scientific computation with applications, part II. Computers and Mathematics with Applications, 176 (??):564, December 15, 2024.

CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122124004917.

Jaglan:2024:ASP

[JMS^{+}24] Jyoti Jaglan, Vikas Maurya, Ankit Singh, Vivek S. Yadav, and Manoj K. Rajpoot. Acoustic and soliton propagation using fullydiscrete energy preserving partially implicit scheme in homogeneous and heterogeneous mediums. Computers and Mathematics with Applications, 174(??):379– 396, November 15, 2024. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122124004425.

Jog:2025:ECT

[Jog25] C. S. Jog. An energy conserving time finite element method that does not require a Hamiltonian formalism. Computers and Mathematics with Applications, 195 (??):42–51, October 1, 2025. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122125002834.

Jo:2025:FEO

[JP25] Gwanghyun Jo and Hyeokjoo Park. A family of evenorder edge-oriented nonconforming finite elements in 2D with efficient locally conservative flux reconstruction. Computers and Mathematics with Applications, 196(??): 127–134, October 15, 2025. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S089812212500313X.

Jarrah:2024:NIM

[JRu24] Ibrahim Jarrah and Rizwanuddin. Nodal integral method to solve the twodimensional, time-dependent, incompressible Navier–Stokes equations in curvilinear coordinates. Computers and Mathematics with Applications, 158(??):219–243, March 15, 2024. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S089812212400049X.

Jolivet:2021:KPE

[JRZ21] Pierre Jolivet, Jose E. Roman, and Stefano Zampini. KSPHPDDM and PCH-PDDM: Extending PETSc with advanced Krylov methods and robust multilevel overlapping Schwarz preconditioners. Computers and Mathematics with Applications, 84(??):277–295, February 15, 2021. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122121000055.

Ji:2025:ESL

[JS25] Minhwan Ji and Jaemin Shin. Energy-stable linear convex splitting methods for the parabolic sine-Gordon equation. Computers and Mathematics with Applications, 198(??):24–37, November 15, 2025. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122125003347.

Jiang:2022:MFR

[JSBZ22] Yuehua Jiang, HongGuang Sun, Yu Bai, and Yan Zhang. MHD flow, radiation heat and mass transfer of fractional Burgers’ fluid in porous medium with chemical reaction. Computers and Mathematics with Applications, 115(??):68– 79, June 1, 2022. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122122000141.

Javadpour:2024:ENM

[JSSS24] Seyed Alireza Javadpour, Manouchehr Salehi, and Saeed Saber-Samandari. An efficient numerical method for the quasi-static behaviour of micropolar viscoelastic Timoshenko beams for couple stress problems. Computers and Mathematics with Applications, 155(??):15–34, February 1, 2024. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122123005400.

Jiang:2024:CHM

[JSZ24] Yuehua Jiang, HongGuang Sun, and Yong Zhang. Convection heat and mass transfer of non-Newtonian fluids in porous media with Soret and Dufour effects using a two-sided space fractional derivative model. Computers and Mathematics with Applications, 173(??):74–86, November 1, 2024. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122124003468.

J:2023:EOO

[JVBM23] Sasikala J., Kesavulu Naidu V., Venkatesh B., and S. M. Mallikarjunaiah. On an efficient octic order subparametric finite element method on curved domains. Computers and Mathematics with Applications, 143(??): 249–268, August 1, 2023. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122123002043.

Jing:2020:LSO

[JW20] Xiaobo Jing and Qi Wang.

Linear second order energy stable schemes for phase field crystal growth models with nonlocal constraints. Computers and Mathematics with Applications, 79(3): 764–788, February 1, 2020. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122119303827.

Wang:2023:CSM

[jWC23] Tian jun Wang and Guo Chai. Composite spectral method for the Neumann problem of the Burgers equation on the half line. Computers and Mathematics with Applications, 134(??):194– 206, March 15, 2023. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122123000184.

Jia:2020:NLU

[JWL20] Hongen Jia, Xue Wang, and Kaitai Li. A novel linear, unconditional energy stable scheme for the incompressible Cahn–Hilliard– Navier–Stokes phase-field model. Computers and Mathematics with Applications, 80(12):2948–2971, December 15, 2020. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122120303990.

Jiang:2023:HEA

[JWR^{+}23] Tao Jiang, Xing-Chi Wang, Jin-Lian Ren, Jin-Jing Huang, and Jin-Yun Yuan. A highefficient accurate coupled mesh-free scheme for 2D/3D space-fractional convection– diffusion/Burgers’ problems. Computers and Mathematics with Applications, 140 (??):260–281, June 15, 2023. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122122004503.

Jammoul:2021:PFM

[JWW21a] Mohamad Jammoul, Mary F. Wheeler, and Thomas Wick. A phase-field multirate scheme with stabilized iterative coupling for pressure driven fracture propagation in porous media. Computers and Mathematics with Applications, 91 (??):176–191, June 1, 2021. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S089812212030434X.

Jin:2021:NIE

[JWW^{+}21b] Hui Jin, Huibo Wang, Zhenqun Wu, Zhiwei Ge, and Yunan Chen. Numerical investigation of the effect of surface roughness on flow and heat transfer characteristics of single sphere particle in supercritical water. Computers and Mathematics with Applications, 81(??): 562–572, January 1, 2021. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122119305000.

Jia:2022:FNS

[JWZ22] Jinhong Jia, Hong Wang, and Xiangcheng Zheng. A fast numerical scheme for a variably distributedorder time-fractional diffusion equation and its analysis. Computers and Mathematics with Applications, 108(??):24–32, February 15, 2022. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122121004557.

Jiang:2025:CAC

[JWZX25] Anwen Jiang, Yan Wang, Fenglian Zheng, and Xufeng Xiao. A convective Allen– Cahn model for the twoand three-dimensional shape transformations of noncontact objects. Computers and Mathematics with Applications, 188(??):72– 82, June 15, 2025. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122125001142.

Jiang:2020:HBR

[JXK20] Ying Jiang, Da-Quan Xian, and Xiao-Rong Kang. Homoclinic breather and rogue wave solutions to Maccari equation. Computers and Mathematics with Applications, 79(7):1890– 1894, April 2020. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122118302694.

Joe:2023:SLS

[JYHS23] Beom-Jin Joe, Sang-Jae Yeo, Suk-Yoon Hong, and Jee-Hun Song. Stable LBM schemes for acoustic scaling simulations under high Reynolds to Mach ratio: Introduction to the DM-TS operator. Computers and Mathematics with Applications, 145(??):1–12, September 1, 2023. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S089812212300247X.

Ji:2020:EEC

[JZ20] Bingquan Ji and Luming Zhang. Error estimates of a conservative finite difference Fourier pseudospectral method for the Klein– Gordon–Schrödinger equation. Computers and Mathematics with Applications, 79 (7):1956–1971, April 2020. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122119303311.

Ji:2024:RNW

[JZ24a] Shaolin Ji and Linlin Zhu.

Reweighted Nadaraya–Watson estimation of stochastic volatility jump-diffusion models. Computers and Mathematics with Applications, 174(??):352–360, November 15, 2024. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122124004383.

Jiang:2024:EDD

[JZ24b] Jiantao Jiang and Zhimin Zhang. An efficient decoupled and dimension reduction scheme for quadcurl eigenvalue problem in balls and spherical shells. Computers and Mathematics with Applications, 174(??): 454–480, November 15, 2024. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122124004528.

Jia:2025:ATF

[JZ25] Zhen Jia and Hong Zheng. Analysis of transient free surface seepage flow using numerical manifold method. Computers and Mathematics with Applications, 189 (??):129–143, July 1, 2025. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122125001567.

Jiang:2022:HOS

[JZHZ22] Yan-Qun Jiang, Shu-Guang Zhou, Ying-Gang Hu, and Xu Zhang. High order semiimplicit weighted compact nonlinear scheme for the full compressible Euler system at all Mach numbers. Computers and Mathematics with Applications, 109 (??):125–139, March 1, 2022. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122122000281.

Jain:2021:CAA

[JZPG21] V. Jain, Y. Zhang, A. Palha, and M. Gerritsma. Construction and application of algebraic dual polynomial representations for finite element methods on quadrilateral and hexahedral meshes. Computers and Mathematics with Applications, 95(??): 101–142, August 1, 2021. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122120303862.

Jiang:2023:PGD

[JZS23] Ye Jiang, Bo Zheng, and Yueqiang Shang. A parallel grad-div stabilized finite element algorithm for the Stokes equations with damping. Computers and Mathematics with Applications, 135 (??):171–192, April 1, 2023. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122123000391.

Jiang:2024:ASD

[JZXR24] Pengfei Jiang, Hui Zheng, Jingang Xiong, and Timon Rabczuk. An adaptive support domain for the incompressible fluid flow based on the localized radial basis function collocation method. Computers and Mathematics with Applications, 156(??): 29–41, February 15, 2024. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122123005631.

Krysko:2023:QVR

[KAKK23] V. A. Krysko, J. Awrejcewicz, L. A. Kalutsky, and V. A. Krysko. Quantification of various reduced order modelling computational methods to study deflection of size-dependent plates. Computers and Mathematics with Applications, 133 (??):61–84, March 1, 2023. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122123000044.

Kamensky:2021:OSI

[Kam21] David Kamensky. Opensource immersogeometric analysis of fluid–structure interaction using FEniCS and tIGAr. Computers and Mathematics with Applications, 81(??):634–648, January 1, 2021. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122120300481.

Kamrani:2025:EDF

[Kam25] Minoo Kamrani. Efficient discretization of fractional SPDEs via Galerkin and exponential Euler methods. Computers and Mathematics with Applications, 193(??): 205–221, September 1, 2025. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S089812212500255X.

Kapoor:2024:CQH

[Kap24]  Mamta Kapoor.  Cubic

and quartic hyperbolic Bsplines comparison for coupled Navier Stokes equation via differential quadrature method — a statistical aspect. Computers and Mathematics with Applications, 171 (??):6–33, October 1, 2024. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122124003067.

Karatzas:2024:BDO

[Kar24] Efthymios N. Karatzas.

Boundary and distributed optimal control for a population dynamics PDE model with discontinuous in time Galerkin FEM schemes. Computers and Mathematics with Applications, 165 (??):70–87, July 1, 2024. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122124001688.

Karmakar:2023:CPF

[KARS23] Timir Karmakar, Meraj Alam, Motahar Reza, and G. P. Raja Sekhar. Couette– Poiseuille flow in a fluid overlying an anisotropic porous layer. Computers and Mathematics with Applications, 151(??):346–358, December 1, 2023. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122123004546.

Krais:2021:FHO

[KBB^{+}21] Nico Krais, Andrea Beck, Thomas Bolemann, Hannes Frank, David Flad, Gregor Gassner, Florian Hindenlang, Malte Hoffmann, Thomas Kuhn, Matthias Sonntag, and Claus-Dieter Munz. FLEXI: a high order discontinuous Galerkin framework for hyperbolic– parabolic conservation laws. Computers and Mathematics with Applications, 81(??): 186–219, January 1, 2021. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122120302029.

Karatzas:2020:PBR

[KBR20] Efthymios N. Karatzas, Francesco Ballarin, and Gianluigi Rozza. Projectionbased reduced order models for a cut finite element method in parametrized domains. Computers and Mathematics with Applications, 79(3):833–851, February 1, 2020. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122119303931.

Kosin:2023:NMM

[KBW23] V. Kosin, S. Beuchler, and T. Wick. A new mixed method for the biharmonic eigenvalue problem. Computers and Mathematics with Applications, 136(??):44– 53, April 15, 2023. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122123000512.

Kirby:2020:OOP

[KC20] Robert C. Kirby and Peter Coogan. Optimal-order preconditioners for the Morse– Ingard equations. Computers and Mathematics with Applications, 79(8):2458– 2471, April 15, 2020. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122119305449.

Kercher:2021:LSF

[KC21] Andrew D. Kercher and Andrew Corrigan. A least-squares formulation of the moving discontinuous Galerkin finite element method with interface condition enforcement. Computers and Mathematics with Applications, 95(??): 143–171, August 1, 2021. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122120303680.

Karageorghis:2022:TRN

[KC22a] Andreas Karageorghis and C. S. Chen. Training RBF neural networks for the solution of elliptic boundary value problems. Computers and Mathematics with Applications, 126(??):196– 211, November 15, 2022. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122122003480.

Kim:2022:LFD

[KC22b] Yongho Kim and Yongho Choi. Learning finite difference methods for reactiondiffusion type equations with FCNN. Computers and Mathematics with Applications, 123(??):115–122, October 1, 2022. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122122003273.

Kumar:2023:MSF

[KC23a] B. V. Rathish Kumar and Manisha Chowdhury. Multiscale stabilized finite element computation of the non-Newtonian Casson fluid flowing in double lid-driven rectangular cavities. Computers and Mathematics with Applications, 143(??):57–72, August 1, 2023. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122123001864.

Kundaliya:2023:SFO

[KC23b] Pari J. Kundaliya and Sudhakar Chaudhary. Symmetric fractional order reduction method with L1 scheme on graded mesh for time fractional nonlocal diffusion-wave equation of Kirchhoff type. Computers and Mathematics with Applications, 149(??):128–134, November 1, 2023. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122123003760.

Kpadonou:2020:ENG

[KCC20] Félix Kpadonou, Stéphanie Chaillat, and Patrick Ciarlet. On the efficiency of nested GMRES preconditioners for 3D acoustic and elastodynamic H-matrix accelerated boundary element methods. Computers and Mathematics with Applications, 80(3): 471–489, August 1, 2020. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122120301255.

Klar:2023:SDD

[KCD^{+}23] Manuel Klar, Giacomo Capodaglio, Marta D’Elia, Christian Glusa, Max Gunzburger, and Christian Vollmann. A scalable domain decomposition method for FEM discretizations of nonlocal equations of integrable and fractional type. Computers and Mathematics with Applications, 151(??):434– 448, December 1, 2023. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122123004662.

Karakus:2022:LDG

[KCW22] A. Karakus, N. Chalmers, and T. Warburton. A local discontinuous Galerkin level set reinitialization with subcell stabilization on unstructured meshes. Computers and Mathematics with Applications, 123(??):160– 170, October 1, 2022. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122122003315.

Kumar:2023:WGF

[KD23] Raman Kumar and Bhupen Deka. Weak Galerkin finite element methods for H(curl; Ω) and H(curl, div; Ω)- elliptic problems. Computers and Mathematics with Applications, 147(??):210– 221, October 1, 2023. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122123003139.

Keith:2021:PEA

[Kei21] Brendan Keith. A priori error analysis of high-order LL* (FOSLL*) finite element methods. Computers and Mathematics with Applications, 103(??):12–18, December 1, 2021. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122121003631.

Karaagac:2023:NOA

[KEUY23] Berat Karaagac, Alaattin Esen, Yusuf Ucar, and Nuri Murat Yagmurlu. A new outlook for analysis of Noyes—Field model for the nonlinear Belousov– Zhabotinsky reaction using operator splitting method. Computers and Mathematics with Applications, 136 (??):127–135, April 15, 2023. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S089812212300069X.

Kaya:2022:CAD

[KF22] Adem Kaya and Melina A.

Freitag. Conditioning analysis for discrete Helmholtz problems. Computers and Mathematics with Applications, 118(??):171–182, July 15, 2022. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122122002048.

Klemens:2020:SFF

[KFD^{+}20] Fabian Klemens, Benjamin Förster, Márcio Dorn, Gudrun Thäter, and Mathias J. Krause. Solving fluid flow domain identification problems with adjoint lattice Boltzmann methods. Computers and Mathematics with Applications, 79(1): 17–33, January 2020. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122118303754.

Kaur:2020:TFD

[KG20a] Bikramjeet Kaur and R. K. Gupta. Time fractional (2 + 1)-dimensional Wu–Zhang system: Dispersion analysis, similarity reductions, conservation laws, and exact solutions. Computers and Mathematics with Applications, 79(4):1031–1048, February 15, 2020. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122119304110.

Kaur:2020:UPU

[KG20b] Navjot Kaur and Kavita Goyal. Uncertainty propagation using Wiener–linear B-spline wavelet expansion. Computers and Mathematics with Applications, 79(9): 2598–2623, May 1, 2020. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122119305541.

Klass:2024:CBC

[KGB24] Friedemann Klass, Alessandro Gabbana, and Andreas Bartel. Characteristic boundary condition for thermal lattice Boltzmann methods. Computers and Mathematics with Applications, 157 (??):195–208, March 1, 2024. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122123005953.

Kim:2024:SDG

[KGJN24] Hyea Hyun Kim, Gung-Min Gie, Chang-Yeol Jung, and Thien Binh Nguyen. A staggered discontinuous Galerkin method for the Stokes problem on rectangular meshes. Computers and Mathematics with Applications, 162 (??):180–195, May 15, 2024. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122124001123.

Kumar:2025:NNM

[KGK25] Saurabh Kumar, Vikas Gupta, and Ajay Kumar. A novel numerical method for solving a two-dimensional variable-order time fractional advection–diffusion problem. Computers and Mathematics with Applications, 195(??): 177–190, October 1, 2025. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122125003141.

Koch:2021:DOS

[KGW^{+}21] Timo Koch, Dennis Gläser, Kilian Weishaupt, Sina Ackermann, Martin Beck, Beatrix Becker, Samuel Burbulla, Holger Class, Edward Coltman, Simon Emmert, Thomas Fetzer, Christoph Grüninger, Katharina Heck, Johannes Hommel, Theresa Kurz, Melanie Lipp, Farid Mohammadi, Samuel Scherrer, Martin Schneider, Gabriele Seitz, Leopold Stadler, Martin Utz, Felix Weinhardt, and Bernd Flemisch. DuMu^{x}3

— an open-source simulator for solving flow and transport problems in porous media with a focus on model coupling. Computers and Mathematics with Applications, 81(??):423–443, January 1, 2021. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122120300791.

Kalleli:2024:IFB

[KHD24] S. Kalleli, A. Hajlaoui, and F. Dammak. Improved-FSDT-based solid-shell element for buckling analysis of plate, spherical cap, and cylindrical shell of FG porous materials. Computers and Mathematics with Applications, 157(??):1–14, March 1, 2024. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122123005771.

Keram:2025:CAF

[KHH25] Aytura Keram, Pengzhan Huang, and Yinnian He. Convergence analysis of a fully discrete numerical scheme for the two-phase incompressible ferrofluid flows. Computers and Mathematics with Applications, 199(??): 343–377, December 1, 2025. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122125004328.

Khompysh:2025:TDI

[KHM25] Kh. Khompysh, M. J.

Huntul, and M. Mukhambetkaliyev. Time-dependent inverse source problems for a pseudoparabolic equation with memory. Computers and Mathematics with Applications, 198(??):239–254, November 15, 2025. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S089812212500361X.

Kim:2025:IAA

[KHY25] Donghyun Kim, Jeonggyu Huh, and Ji-Hun Yoon. Improved accuracy of an analytical approximation for option pricing under stochastic volatility models using deep learning techniques. Computers and Mathematics with Applications, 187 (??):150–165, June 1, 2025. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122125001245.

Kim:2021:AOD

[Kim21] Mi-Young Kim. Arbitrary order DG-DGLM method for hyperbolic systems of multi-dimensional conservation laws. Computers and Mathematics with Applications, 97(??):100–121, September 1, 2021. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122121002340.

Kumar:2020:LWB

[KJ20] K. Harish Kumar and Ram Jiwari. Legendre wavelets based numerical algorithm for simulation of multidimensional Benjamin– Bona–Mahony–Burgers and Sobolev equations. Computers and Mathematics with Applications, 80(3):417–433, August 1, 2020. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122120301206.

Kashyap:2021:RAF

[KJ21] Suraj R. Kashyap and Rajeev K. Jaiman. A robust and accurate finite element framework for cavitating flows with moving fluidstructure interfaces. Computers and Mathematics with Applications, 103(??):19–39, December 1, 2021. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122121003783.

Kumar:2022:DMI

[KJ22] Ankit Kumar and Subit Kumar Jain. Deformable models for image segmentation: a critical review of achievements and future challenges. Computers and Mathematics with Applications, 119(??): 288–311, August 1, 2022. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122122002279.

Kim:2024:SBC

[KJ24] Yong-Jin Kim and Kyung-Young Jung. Surface boundary condition (SBC)-based FDTD formulations for lossy dispersive media. Computers and Mathematics with Applications, 171(??):204– 213, October 1, 2024. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122124003250.

Korshunova:2020:IBM

[KJL^{+}20] N. Korshunova, J. Jomo, G. Lékó, D. Reznik, P. Balázs, and S. Kollmannsberger. Image-based material characterization of complex microarchitectured additively manufactured structures. Computers and Mathematics with Applications, 80(11): 2462–2480, December 2020. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122120302844.

Kim:2021:SDG

[KJN21] H. H. Kim, C.-Y. Jung, and T. B. Nguyen. A staggered discontinuous Galerkin method for elliptic problems on rectangular grids. Computers and Mathematics with Applications, 99(??):

133–154, October 1, 2021. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122121002959.

Kirby:2021:PMF

[KK21] Robert C. Kirby and Tate Kernell. Preconditioning mixed finite elements for tide models. Computers and Mathematics with Applications, 82(??):212–227, January 15, 2021. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122120304272.

Kalpana:2023:FDA

[KK23a] G. Kalpana and Ramesh B.

Kudenatti. Finite difference approach to two-dimensional magnetohydrodynamic fluid flow due to moving surface. Computers and Mathematics with Applications, 144(??): 257–263, August 15, 2023. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122123002614.

Kim:2023:AME

[KK23b] Ji Hoo Kim and Jae Ryong Kweon. An analysis for the momentum equation with unbounded pressure gradient: Numerical simulations. Computers and Mathematics with Applications, 140 (??):126–133, June 15, 2023. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122123001293.

Kumar:2025:DCP

[KK25] Rajendra Kumar and B. V. Rathish Kumar. A discontinuitycapturing PINN for parabolic interface problems. Computers and Mathematics with Applications, 195(??):93– 108, October 1, 2025. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122125002809.

Krause:2021:OOS

[KKA^{+}21] Mathias J. Krause, Adrian Kummerländer, Samuel J. Avis, Halim Kusumaatmaja, Davide Dapelo, Fabian Klemens, Maximilian Gaedtke, Nicolas Hafen, Albert Mink, Robin Trunk, Jan E. Marquardt, Marie-Luise Maier, Marc Haussmann, and Stephan Simonis. OpenLB — open source lattice Boltzmann code. Computers and Mathematics with Applications, 81(??):258–288, January 1, 2021. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122120301875.

Kwak:2022:FVA

[KKJ^{+}22] Songhun Kwak, Kwanghun Kim, Sunjong Jon, Jongguk Yun, and Chanil Pak. Free vibration analysis of laminated rectangular plates with varying thickness using Legendre-radial point interpolation method. Computers and Mathematics with Applications, 117(??):187– 205, July 1, 2022. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122122001766.

Kumar:2021:GCE

[KKNR21] Kundan Kumar, Svetlana Kyas, Jan Martin Nordbotten, and Sergey Repin. Guaranteed and computable error bounds for approximations constructed by an iterative decoupling of the Biot problem. Computers and Mathematics with Applications, 91 (??):122–149, June 1, 2021. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122120302030.

Kraus:2025:HOG

[KKS25] Heinrich Kraus, Jörg Kuhnert, and Pratik Suchde. Higher-order generalized finite differences for variable coefficient diffusion operators. Computers and Mathematics with Applications, 194(??):257–271, September 15, 2025. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122125002640.

Kim:2022:TDV

[KLK^{+}22] Hyundong Kim, Chaeyoung Lee, Soobin Kwak, Youngjin Hwang, Sangkwon Kim, Yongho Choi, and Junseok Kim. Three-dimensional volume reconstruction from multi-slice data using a shape transformation. Computers and Mathematics with Applications, 113(??):52– 58, May 1, 2022. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122122001031.

Kim:2024:HNM

[KLK^{+}24] Hyundong Kim, Gyeonggyu Lee, Seungyoon Kang, Seokjun Ham, Youngjin Hwang, and Junseok Kim. Hybrid numerical method for the Allen–Cahn equation on nonuniform grids. Computers and Mathematics with Applications, 158(??): 167–178, March 15, 2024. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122124000245.

Khachkova:2021:GBA

[KLRT21] Tatyana Khachkova, Vadim Lisitsa, Galina Reshetova, and Vladimir Tcheverda. GPU-based algorithm for evaluating the electrical resistivity of digital rocks. Computers and Mathematics with Applications, 82(??): 200–211, January 15, 2021. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122120304302.

Klibanov:2025:CCI

[KLY25] Michael V. Klibanov, Jingzhi Li, and Zhipeng Yang. Convexification for a coefficient inverse problem for a system of two coupled nonlinear parabolic equations. Computers and Mathematics with Applications, 179 (??):41–58, February 1, 2025. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S089812212400542X.

Ke:2022:APB

[KM22] Yifen Ke and Changfeng Ma. Adaptive parameter based matrix splitting iteration method for the large and sparse linear systems. Computers and Mathematics with Applications, 122(??): 19–27, September 15, 2022. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122122002942.

Kimura:2025:CEE

[KMN25] Takuma Kimura, Teruya Minamoto, and Mitsuhiro T. Nakao. Constructive error estimates for a fulldiscretized periodic solution of heat equation by spatial finite-element and time spectral method. Computers and Mathematics with Applications, 181(??):248– 258, March 1, 2025. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122125000082.

Krzyzanowski:2020:WFD

[KMP20] Grzegorz Krzyżanowski, Marcin Magdziarz, and Łukasz Płociniczak. A weighted finite difference method for subdiffusive Black–Scholes model. Computers and Mathematics with Applications, 80(5):653–670, September 1, 2020. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122120301760.

Kumar:2020:PCN

[KMS20] Niteen Kumar, Rudrodip Majumdar, and Suneet Singh. Predictor-corrector nodal integral method for simulation of high Reynolds number fluid flow using larger time steps in Burgers ’ equation. Computers and Mathematics with Applications, 79(5):1362– 1381, March 1, 2020. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122119304377.

Kawano:2022:ISM

[KMZ22] A. Kawano, A. Morassi, and R. Zaera. Identification of a symmetric mass density in a rectangular membrane from finite eigenvalue data. Computers and Mathematics with Applications, 125(??): 222–237, November 1, 2022. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122122003753.

Krawczyk:2020:OCI

[KN20] Anita Krawczyk and Andrzej Nowakowski. Optimal control of inhibiting tumor, new approach to sufficient ε-optimality and numerical computation. Computers and Mathematics with Applications, 80(5):778–789, September 1, 2020. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122120301589.

K:2024:ICN

[KN24] Vishnu Prakash K. and Ganesh Natarajan. IGG(chi): a new and simple implicit gradient scheme on unstructured meshes. Computers and Mathematics with Applications, 175(??):385–398, December 1, 2024. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122124004656.

Kumar:2025:NNS

[KN25] Sachin Kumar and Srinivasan Natesan. A novel numerical scheme for Black– Scholes PDEs modeling pricing securities. Computers and Mathematics with Applications, 190(??):57– 71, July 15, 2025. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122125001488.

Karatzas:2022:ROC

[KNBR22] Efthymios N. Karatzas, Monica Nonino, Francesco Ballarin, and Gianluigi Rozza. A Reduced Order Cut Finite Element method for geometrically parametrized steady and unsteady Navier– Stokes problems. Computers and Mathematics with Applications, 116(??):140– 160, June 15, 2022. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122121002790.

Koch:2022:PBR

[Koc22] Timo Koch. Projectionbased resolved interface 1D-3D mixed-dimension method for embedded tubular network systems. Computers and Mathematics with Applications, 109(??):15– 29, March 1, 2022. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S089812212200027X.

Kung:2020:CPS

[KOMK20] V. E. Küng, F. Osmanlic, M. Markl, and C. Körner. Comparison of passive scalar transport models coupled with the Lattice Boltzmann method. Computers and Mathematics with Applications, 79(1):55–65, January 2020. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122118300336.

Kohlenbach:2020:RCI

[KP20] Ulrich Kohlenbach and Thomas Powell. Rates of convergence for iterative solutions of equations involving setvalued accretive operators. Computers and Mathematics with Applications, 80(3): 490–503, August 1, 2020. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122120301310.

Kim:2021:ICG

[KP21] Won Hyun Kim and Tae Seon Park. Inflow characteristics for generating axis-switching phenomenon of heated rectangular jets. Computers and Mathematics with Applications, 86(??):73–89, March 15, 2021. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122121000377.

Kajzer:2022:MDM

[KP22a] Adam Kajzer and Jacek Pozorski. The mass diffusive model of Svärd simplified to simulate nearly incompressible flows. Computers and Mathematics with Applications, 121(??):18–29, September 1, 2022. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122122002619.

Kajzer:2022:WCD

[KP22b] Adam Kajzer and Jacek Pozorski. A weakly compressible, diffuse interface model of two-phase flows: Numerical development and validation. Computers and Mathematics with Applications, 106 (??):74–91, January 15, 2022. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122121004260.

Kale:2022:AIA

[KP22c] Swapnil Kale and Debasish Pradhan. An augmented interface approach in fictitious domain methods. Computers and Mathematics with Applications, 125(??):238–247, November 1, 2022. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122122003741.

Kang:2022:AMA

[KP22d] Sangwoo Kang and Won-Kwang Park. Application of MUSIC algorithm for a fast identification of small perfectly conducting cracks in limited-aperture inverse scattering problem. Computers and Mathematics with Applications, 117 (??):97–112, July 1, 2022. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122122001717.

Kumari:2023:STM

[KP23] Sarita Kumari and Rajesh K. Pandey. Single-term and multi-term nonuniform time-stepping approximation methods for two-dimensional time-fractional diffusionwave equation. Computers and Mathematics with Applications, 151(??):359–383, December 1, 2023. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122123004558.

Kropielnicka:2024:AEL

[KP24] Karolina Kropielnicka and Rafał Perczyński. Asymptotic expansions for the linear PDEs with oscillatory input terms; analytical form and error analysis. Computers and Mathematics with Applications, 156(??):16–27, February 15, 2024. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S089812212300562X.

Kumar:2022:ACL

[KPJ22] Anil Kumar, Amiya K. Pani, and Mohan C. Joshi. Approximate controllability of linear parabolic equation with memory. Computers and Mathematics with Applications, 128(??):320–330, December 15, 2022. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S089812212200459X.

Kim:2022:ESI

[KPJO22] Seokchan Kim, Birce Palta, Jaewoo Jeong, and Hae-Soo Oh. Extraction of stress intensity factors of biharmonic equations with corner singularities corresponding to mixed boundary conditions of clamped, simply supported, and free (II). Computers and Mathematics with Applications, 109 (??):235–259, March 1, 2022. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122122000359.

Kale:2025:APF

[KPK25] Swapnil Kale, Debasish Pradhan, and Sarvesh Kumar. Analysis of H^{1} penalized fictitious domain method for parabolic problems. Computers and Mathematics with Applications, 196(??):183–200, October 15, 2025. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122125002962.

Kim:2020:EFS

[KPO20] Seokchan Kim, Birce Palta, and Hae-Soo Oh. Extraction formulas of stress intensity factors for the biharmonic equations containing crack singularities. Computers and Mathematics with Applications, 80(5):1142–1163, September 1, 2020. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122120302480.

Karban:2021:FBR

[KPO^{+}21] Pavel Karban, David Pánek, Tamás Orosz, Iveta Petrášová, and Ivo Doležel. FEM based robust design optimization with Agros and Ārtap. Computers and Mathematics with Applications, 81(??): 618–633, January 1, 2021. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122120300778.

Kang:2023:QAB

[KPS23a] Sangwoo Kang, Won-Kwang Park, and Seong-Ho Son. A qualitative analysis of bifocusing method for a real-time anomaly detection in microwave imaging. Computers and Mathematics with Applications, 137 (??):93–101, May 1, 2023. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122123000767.

Khandelwal:2023:SNP

[KPS23b] Rohit Khandelwal, Kamana Porwal, and Ritesh Singla. Supremum-norm a posteriori error control of quadratic discontinuous Galerkin methods for the obstacle problem. Computers and Mathematics with Applications, 137 (??):147–171, May 1, 2023. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122123000615.

Kumari:2024:ECT

[KR24] Trishna Kumari and Pradip Roul. An efficient computational technique for solving a time-fractional reaction-subdiffusion model in 2D space. Computers and Mathematics with Applications, 160(??):191– 208, April 15, 2024. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122124000579.

Krasnopolsky:2020:RPB

[Kra20] B. Krasnopolsky. Revisiting performance of BiCGStab methods for solving systems with multiple right-hand sides. Computers and Mathematics with Applications, 79 (9):2574–2597, May 1, 2020. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122119305589.

Kravetc:2023:BSS

[Kra23] Tatiana Kravetc. Blending spline surfaces over polygon mesh and their application to isogeometric analysis. Computers and Mathematics with Applications, 149(??): 84–98, November 1, 2023. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122123003735.

Kazemian:2021:EGS

[KRESA21] Yousef Kazemian, Saman Rashidi, Javad Abolfazli Esfahani, and Omid Samimi-Abianeh. Effects of grains shapes of porous media on combustion onset — a numerical simulation using lattice Boltzmann method. Computers and Mathematics with Applications, 81(??): 547–561, January 1, 2021. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122119305048.

Krivovichev:2020:APM

[Kri20] Gerasim V. Krivovichev. Analysis of the parametric models of passive scalar transport used in the lattice Boltzmann method. Computers and Mathematics with Applications, 79(5): 1503–1524, March 1, 2020. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122119304584.

Kadri:2024:EAN

[KRO24] Tlili Kadri, Mohamed Rahmeni, and Khaled Omrani. Efficient and accurate numerical methods for nonlinear strongly damped wave equation in 2 + 1 dimensions. Computers and Mathematics with Applications, 160 (??):148–160, April 15, 2024. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122124000713.

Krzyzanowski:2022:APM

[Krz22] Piotr Krzyżanowski. Analysis of a preconditioner for a membrane diffusion problem with the Kedem–Katchalsky transmission condition. Computers and Mathematics with Applications, 122(??):28–34, September 15, 2022. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122122002917.

Kabre:2021:PSA

[KS21] Julienne Kabre and Qin Sheng. A preservative splitting approximation of the solution of a variable coefficient quenching problem. Computers and Mathematics with Applications, 100 (??):62–73, October 15, 2021. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122121003151.

Kumar:2025:NVE

[KS25] Sarvesh Kumar and Devika Shylaja. Nonconforming virtual element method for an incompressible miscible displacement problem in porous media. Computers and Mathematics with Applications, 183(??):153– 179, April 1, 2025. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122125000288.

Karatson:2025:QNI

[KSB25] János Karátson, Stanislav Sysala, and Michal Béreš. Quasi-Newton iterative solution approaches for nonsmooth elliptic operators with applications to elastoplasticity. Computers and Mathematics with Applications, 178(??):61–80, January 15, 2025. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122124005194.

Khara:2025:STF

[KSD^{+}25] Biswajit Khara, Kumar Saurabh, Robert Dyja, Anupam Sharma, and Baskar Ganapathysubramanian. Spacetime finite element analysis of the advection–diffusion equation using galerkin/ least-square stabilization. Computers and Mathematics with Applications, 185 (??):52–75, May 1, 2025. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122125000793.

Kumar:2023:CAW

[KSJ23] Naresh Kumar, Jasbir Singh, and Ram Jiwari. Convergence analysis of weak Galerkin finite element method for semilinear parabolic convection dominated diffusion equations on polygonal meshes. Computers and Mathematics with Applications, 145(??):141–158, September 1, 2023. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122123002638.

Kumar:2020:WAS

[KSS^{+}20] Yashveer Kumar, Somveer Singh, Nikhil Srivastava, Aman Singh, and Vineet Kumar Singh. Wavelet approximation scheme for distributed order fractional differential equations. Computers and Mathematics with Applications, 80(8):1985– 2017, October 15, 2020. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122120303230.

K:2022:HOF

[KSS22] Aarthika K., Ram Shiromani, and V. Shanthi. A higher-order finite difference method for twodimensional singularly perturbed reaction–diffusion with source-term-discontinuous problem. Computers and Mathematics with Applications, 118(??):56–73, July 15, 2022. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122122001729.

Kumar:2023:WBC

[KSSS23] Yashveer Kumar, Nikhil Srivastava, Aman Singh, and Vineet Kumar Singh. Wavelets based computational algorithms for multidimensional distributed order fractional differential equations with nonlinear source term. Computers and Mathematics with Applications, 132(??):73–103, February 15, 2023. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122122005028.

Klinkovsky:2023:LBM

[KTFI23] Jakub Klinkovský, Andrew C. Trautz, Radek Fučík, and Tissa H. Illangasekare. Lattice Boltzmann method-based efficient GPU simulator for vapor transport in the boundary layer over a moist soil: Development and experimental validation. Computers and Mathematics with Applications, 138 (??):65–87, May 15, 2023. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122123000809.

Keppens:2021:MAP

[KTXP21] Rony Keppens, Jannis Teunissen, Chun Xia, and Oliver Porth. MPI-AMRVAC: a parallel, grid-adaptive PDE toolkit. Computers and Mathematics with Applications, 81(??):316–333, January 1, 2021. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122120301279.

Kim:2023:LCI

[KTY23] Junseok Kim, Zhijun Tan, and Junxiang Yang. Linear and conservative IMEX Runge–Kutta finite difference schemes with provable energy stability for the Cahn–Hilliard model in arbitrary domains. Computers and Mathematics with Applications, 143(??):133– 150, August 1, 2023. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122123002006.

Kumar:2024:CAV

[KTY24] Naresh Kumar, Jai Tushar, and J. Y. Yuan. Convergence analysis of virtual element method for the electric interface model on polygonal meshes with small edges. Computers and Mathematics with Applications, 159 (??):254–266, April 1, 2024. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122124000725.

Ku:2024:ALS

[Ku24] JaEun Ku. Adaptive leastsquares finite element methods: Guaranteed upper bounds and convergence in L_{2} norm of the dual variable. Computers and Mathematics with Applications, 162 (??):52–60, May 15, 2024. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122124001019.

Karasozen:2020:ROO

[KUK20] Bülent Karasözen, Murat Uzunca, and Tugba Küçükseyhan. Reduced order optimal control of the convective FitzHugh– Nagumo equations. Computers and Mathematics with Applications, 79(4):982–995, February 15, 2020. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122119304067.

Korotov:2020:MAC

[KV20] Sergey Korotov and Jon Eivind Vatne. The minimum angle condition for d-simplices. Computers and Mathematics with Applications, 80(2): 367–370, July 2020. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122119302834.

Kapl:2022:ISS

[KV22] Mario Kapl and Vito Vitrih. C^{1} isogeometric spline space for trilinearly parameterized multi-patch volumes. Computers and Mathematics with Applications, 117 (??):53–68, July 1, 2022. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S089812212200164X.

Kunisch:2023:OPF

[KVV23] Karl Kunisch and Donato Vásquez-Varas. Optimal polynomial feedback laws for finite horizon control problems. Computers and Mathematics with Applications, 148(??):113–125, October 15, 2023. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122123003358.

Kim:2020:ABM

[KW20] Hyea Hyun Kim and Junxian Wang. An adaptive BDDC method enhanced with prior selected primal constraints. Computers and Mathematics with Applications, 80(8): 1928–1943, October 15, 2020. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122120303126.

Kramer:2020:LBS

[KWK^{+}20] Andreas Krämer, Dominik Wilde, Knut Küllmer, Dirk Reith, Holger Foysi, and Wolfgang Joppich. Lattice Boltzmann simulations on irregular grids: Introduction of the NATriuM library. Computers and Mathematics with Applications, 79(1):34–54, January 2020.

CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122118306382.

Kirby:2024:EDT

[KWK24] Robert C. Kirby, Xiaoyu Wei, and Andreas Klöckner. Exact domain truncation for the Morse–Ingard equations. Computers and Mathematics with Applications, 162 (??):22–32, May 15, 2024. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122124000919.

Kong:2024:PSO

[KWLW24] Linghua Kong, Yexiang Wu, Zhiqiang Liu, and Ping Wang. Pointwise second order convergence of structure-preserving scheme for the triple-coupled nonlinear Schrödinger equations. Computers and Mathematics with Applications, 154(??): 91–102, January 15, 2024. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122123004923.

Kadapa:2022:CAA

[KWM22] Chennakesava Kadapa, Xinyu Wang, and Yue Mei. A comprehensive assessment of accuracy of adaptive integration of cut cells for laminar fluid-structure interaction problems. Computers and Mathematics with Applications, 122(??):1–18, September 15, 2022. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122122002887.

Kummer:2021:BPM

[KWS21] Florian Kummer, Jens Weber, and Martin Smuda. BoSSS: a package for multigrid extended discontinuous Galerkin methods. Computers and Mathematics with Applications, 81(??):237– 257, January 1, 2021. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122120301917.

Kim:2020:PFR

[KYY^{+}20] Hyundong Kim, Ana Yun, Sungha Yoon, Chaeyoung Lee, Jintae Park, and Junseok Kim. Pattern formation in reaction–diffusion systems on evolving surfaces. Computers and Mathematics with Applications, 80(9):2019– 2028, November 1, 2020. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122120303400.

Luthi:2023:ASP

[LAB23] C. Lüthi, M. Afrasiabi, and M. Bambach. An adaptive smoothed particle hydrodynamics (SPH) scheme for efficient melt pool simulations in additive manufacturing. Computers and Mathematics with Applications, 139(??):7–27, June 1, 2023. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122123000925.

Liao:2024:CTI

[LAL^{+}24] Li-Dan Liao, Xia Ai, Rui-Xia Li, Di-Gen Li, and Wei Xu. Comparisons of two iteration methods for time-harmonic parabolic optimal control problems. Computers and Mathematics with Applications, 159 (??):44–59, April 1, 2024. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122124000257.

Lopez-Agredo:2023:NAM

[LARGVR23] Jorge L. López-Agredo, Diego A. Rueda-Gómez, and Élder J. Villamizar-Roa. Numerical analysis of a mathematical model describing the evolution of hypoxic glioma cells. Computers and Mathematics with Applications, 131(??):138–157, February 1, 2023. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122122005120.

Li:2022:NET

[LAX22] Minghui Li, Mejdi Azaiez, and Chuanju Xu. New efficient time-stepping schemes for the anisotropic phasefield dendritic crystal growth model. Computers and Mathematics with Applications, 109(??):204–215, March 1, 2022. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122122000244.

Li:2025:EDG

[LaZMY25] Zhaohua Li, Guang an Zou, Lina Ma, and Xiaofeng Yang. An efficient discontinuous Galerkin method for the hydro-dynamically coupled phase-field vesicle membrane model. Computers and Mathematics with Applications, 181(??):259– 286, March 1, 2025. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S089812212500001X.

Li:2024:EUE

[LaZZ24] Mengmeng Li, Guang an Zou, and Min Zhang. An efficient unconditional energystable finite element method for the electro-hydrodynamic equations. Computers and Mathematics with Applications, 176(??):447–468, December 15, 2024. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122124004942.

Los:2020:FIS

[LBPC20] M. Loś, P. Behnoudfar, M. Paszyński, and V. M. Calo. Fast isogeometric solvers for hyperbolic wave propagation problems. Computers and Mathematics with Applications, 80 (1):109–120, July 1, 2020. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122120300997.

Lee:2022:FIM

[LBTVG22] Jeonghun J. Lee, Tan Bui-Thanh, Umberto Villa, and Omar Ghattas. Forward and inverse modeling of fault transmissibility in subsurface flows. Computers and Mathematics with Applications, 128(??):354–367, December 15, 2022. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122122003935.

Liu:2021:TMM

[LC21] Yang Liu and Ai-Jie Cheng. Two meshless methods for Dirichlet boundary optimal control problem governed by elliptic PDEs. Computers and Mathematics with Applications, 82(??):113–129, January 15, 2021. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122120304387.

Liu:2022:ATLb

[LC22] Min Liu and Zhiqiang Cai. Adaptive two-layer ReLU neural network: II. Ritz approximation to elliptic PDEs. Computers and Mathematics with Applications, 113 (??):103–116, May 1, 2022. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122122000931.

Li:2023:ECD

[LC23a] Qing Li and Huanzhen Chen. An efficient compact difference-proper orthogonal decomposition algorithm for fractional viscoelastic plate vibration model. Computers and Mathematics with Applications, 151(??):190–214, December 1, 2023. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122123004091.

Lin:2023:DFE

[LC23b] Jitong Lin and Xuesong Chen. A diagonal finite element-projection-proximal gradient algorithm for elliptic optimal control problem. Computers and Mathematics with Applications, 148(??): 256–268, October 15, 2023. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122123003528.

Liu:2022:ATLa

[LCC22] Min Liu, Zhiqiang Cai, and Jingshuang Chen. Adaptive two-layer ReLU neural network: I. Best least-squares approximation. Computers and Mathematics with Applications, 113(??):34– 44, May 1, 2022. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122122000876.

Liu:2024:VSP

[LCC^{+}24] Youqiong Liu, Li Cai, Yaping Chen, Pengfei Ma, and Qian Zhong. Variable separated physics-informed neural networks based on adaptive weighted loss functions for blood flow model. Computers and Mathematics with Applications, 153(??):108– 122, January 1, 2024. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122123005096.

Lim:2023:DDT

[LCE23] Wei Xian Lim, Wai Lee Chan, and Basman Elhadidi. A domain decomposition technique for acoustic wave interactions and scattering. Computers and Mathematics with Applications, 152(??): 129–142, December 15, 2023. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122123004571.

Liu:2023:EAB

[LCH23] Ying Liu, Yanping Chen, and Yunqing Huang. Efficient algorithm based on two-grid method for semiconductor device problem. Computers and Mathematics with Applications, 144(??): 221–228, August 15, 2023. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122123002341.

Li:2020:CTV

[LCN20] Peng Li, Wengu Chen, and Michael K. Ng. Compressive total variation for image reconstruction and restoration. Computers and Mathematics with Applications, 80(5): 874–893, September 1, 2020. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122120302042.

Lin:2021:MFS

[LCT21] Bing-Ham Lin, Bang-Fuh Chen, and Chia-Cheng Tsai. Method of fundamental solutions on simulating sloshing liquids in a 2D tank. Computers and Mathematics with Applications, 88 (??):52–69, April 15, 2021. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122119303803.

Lu:2022:LTC

[LCT22] J. B. Lu, Z. X. Chen, and Z. F. Tian. A less timeconsuming upwind compact difference method with adjusted dissipation property for solving the unsteady incompressible Navier–Stokes equations. Computers and Mathematics with Applications, 126(??):149–171, November 15, 2022. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122122003911.

Li:2022:SPG

[LCW22] Shengyue Li, Wanrong Cao, and Yibo Wang. On spectral Petrov–Galerkin method for solving optimal control problem governed by a two-sided fractional diffusion equation. Computers and Mathematics with Applications, 107(??): 104–116, February 1, 2022. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122121004594.

Liu:2020:NMP

[LCX^{+}20] Li-Min Liu, Ying-Ying Cui, Jie Xu, Chao Li, and Qing-Hui Gao. The non-Markovian property of q-Gaussian process. Computers and Mathematics with Applications, 79(6):1802– 1812, March 15, 2020. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S089812211930495X.

Lee:2023:ITO

[LCY23] Kyungrok Lee, Jung-Il Choi, and Jungho Yoon. Improving the third-order WENO schemes by using exponential polynomial space with a locally optimized shape parameter. Computers and Mathematics with Applications, 149(??):24–37, November 1, 2023. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122123003590.

Liu:2020:VEM

[LD20] Wei Liu and Huoyuan Duan.

A virtual element method for stochastic Stokes equations. Computers and Mathematics with Applications, 79(6): 1813–1832, March 15, 2020. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122119304948.

Li:2021:DMC

[LD21] Jiaqi Li and Leszek Demkowicz. An L^{p} -DPG method for the convection–diffusion problem. Computers and Mathematics with Applications, 95(??):172–185, August 1, 2021. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122120303205.

Li:2022:LPF

[LD22] Qingtao Li and Guangzhi Du. Local and parallel finite element methods based on two-grid discretizations for a non-stationary coupled Stokes–Darcy model. Computers and Mathematics with Applications, 113 (??):254–269, May 1, 2022. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122122001195.

Luo:2023:LPD

[LD23] Longshan Luo and Cheng Dong. A linearity-preserving diamond scheme with extended least square interpolation for the heterogeneous and anisotropic diffusion problems on polyhedral meshes. Computers and Mathematics with Applications, 144(??):182– 197, August 15, 2023. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122123002262.

Li:2024:DMP

[LD24] Jiaqi Li and Leszek Demkowicz. A DPG method for planar div-curl problems. Computers and Mathematics with Applications, 159 (??):31–43, April 1, 2024. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122124000373.

Liu:2025:NAD

[LD25] Dong Liu and Weihua Deng. Numerical analysis of a discontinuous Galerkin method for time-fractional Navier– Stokes–Fokker–Planck equations with weakly singular solutions. Computers and Mathematics with Applications, 195(??):280–295, October 1, 2025. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122125003207.

Lin:2024:SSA

[LDH24] Xuelei Lin, Jiamei Dong, and Sean Hon. A single-sided allat-once preconditioning for linear system from a nonlocal evolutionary equation with weakly singular kernels. Computers and Mathematics with Applications, 169(??): 1–16, September 1, 2024. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122124002633.

Li:2025:SDM

[LDM25] Mingxia Li, Qianqian Ding, and Shipeng Mao. A streamline diffusion method for nonstationary incompressible magnetohydrodynamics system with variable density. Computers and Mathematics with Applications, 193(??): 1–19, September 1, 2025. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S089812212500224X.

Lopes:2021:CDG

[LdO21] N. D. Lopes and H. B.

de Oliveira. Continuous/ discontinuous Galerkin approximations for a fourthorder nonlinear problem. Computers and Mathematics with Applications, 97(??): 122–152, September 1, 2021. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122121002352.

Lazaridis:2024:PAI

[LDT24] Dimitrios S. Lazaridis, Konstantinos A. Draziotis, and Nikolaos L. Tsitsas. A parallel algorithm for the inversion of matrices with simultaneously diagonalizable blocks. Computers and Mathematics with Applications, 174(??): 340–351, November 15, 2024. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122124004206.

Li:2024:PIN

[LDW^{+}24] Xi’an Li, Jiaxin Deng, Jinran Wu, Shaotong Zhang, Weide Li, and You-Gan Wang. Physical informed neural networks with soft and hard boundary constraints for solving advection-diffusion equations using Fourier expansions. Computers and Mathematics with Applications, 159(??):60–75, April 1, 2024. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122124000348.

Le:2024:NCU

[LE24] Thanh Trung Le and Raluca Eftimie. Numerical challenges for the understanding of localised solutions with different symmetries in nonlocal hyperbolic systems. Computers and Mathematics with Applications, 169(??): 112–131, September 1, 2024. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S089812212400275X.

Lee:2020:NSE

[Lee20]  Dongsun Lee.  The nu-

merical solutions for the energy-dissipative and massconservative Allen–Cahn equation. Computers and Mathematics with Applications, 80(1):263–284, July 1, 2020. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122120301474.

Lee:2021:PGF

[Lee21] David Lee. Petrov–Galerkin flux upwinding for mixed mimetic spectral elements, and its application to geophysical flow problems. Computers and Mathematics with Applications, 89 (??):68–77, May 1, 2021. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122121000651.

Lee:2024:PEE

[Lee24] Jeonghun J. Lee. A posteriori error estimates of Darcy flows with Robin-type jump interface conditions. Computers and Mathematics with Applications, 174(??):417– 430, November 15, 2024. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122124004401.

Legrain:2021:NNM

[Leg21]  Grégory Legrain.  Non-

negative moment fitting quadrature rules for fictitious domain methods. Computers and Mathematics with Applications, 99(??):270–291, October 1, 2021. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122121002820.

Lackova:2024:ENM

[LF24a] Katarína Lacková and Peter Frolkovič. Efficient numerical methods for models of evolving interfaces enhanced with a small curvature term. Computers and Mathematics with Applications, 158(??):

244–258, March 15, 2024. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122124000440.

Li:2024:UEE

[LF24b] Lingling Li and Yayun Fu.

Unconditional error estimate of linearly-implicit and energy-preserving schemes for nonlocal wave equations. Computers and Mathematics with Applications, 176(??): 492–509, December 15, 2024. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122124004930.

Lacroix:2024:CSI

[LFF^{+}24] Martin Lacroix, Simon Février, Eduardo Fernández, Luc Papeleux, Romain Boman, and Jean-Philippe Ponthot. A comparative study of interpolation algorithms on non-matching meshes for PFEM-FEM fluid-structure interactions. Computers and Mathematics with Applications, 155(??):51–65, February 1, 2024. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122123005485.

Li:2023:RCL

[LFGZ23] Junpu Li, Zhuojia Fu, Yan Gu, and Lan Zhang. Rapid calculation of largescale acoustic scattering from complex targets by a duallevel fast direct solver. Computers and Mathematics with Applications, 130(??):1–9, January 15, 2023. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122122004631.

Liu:2020:TMF

[LFY^{+}20] Yang Liu, Enyu Fan, Baoli Yin, Hong Li, and Jinfeng Wang. TT-M finite element algorithm for a two-dimensional space fractional Gray–Scott model. Computers and Mathematics with Applications, 80(7): 1793–1809, October 1, 2020. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122120303187.

Liu:2021:NST

[LFY^{+}21] Yan-Cheng Liu, Chia-Ming Fan, Weichung Yeih, Cheng-Yu Ku, and Chiung-Lin Chu. Numerical solutions of two-dimensional Laplace and biharmonic equations by the localized Trefftz method. Computers and Mathematics with Applications, 88(??): 120–134, April 15, 2021. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122120303874.

Lee:2023:IPE

[LG23] Jeonghun J. Lee and Omar Ghattas. Interior overpenalized enriched Galerkin methods for second order elliptic equations. Computers and Mathematics with Applications, 152(??):102– 111, December 15, 2023. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122123004406.

Li:2020:ASS

[LGL20] Xiaolei Li, Bin Guo, and Menglan Liao. Asymptotic stability of solutions to quasilinear hyperbolic equations with variable sources. Computers and Mathematics with Applications, 79(4):1012– 1022, February 15, 2020. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122119304134.

Liu:2024:FSE

[LGL24] Hua-Yu Liu, Xiao-Wei Gao, and Jun Lv. The fundamental solution element method based on irregular polygonal meshes. Computers and Mathematics with Applications, 174(??):481– 494, November 15, 2024.

CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122124004541.

Li:2021:RAT

[LGQZ21] Junpu Li, Yan Gu, Qing-Hua Qin, and Lan Zhang. The rapid assessment for three-dimensional potential model of large-scale particle system by a modified multilevel fast multipole algorithm. Computers and Mathematics with Applications, 89 (??):127–138, May 1, 2021. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122121000717.

Liu:2023:ESE

[LGZC23] Hua-Yu Liu, Xiao-Wei Gao, Gui-Yong Zhang, and Miao Cui. Explicit spectral element collocation method for nonlinear transient heat transfer. Computers and Mathematics with Applications, 143(??):289–302, August 1, 2023. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122123002225.

Li:2023:FOD

[LH23a] Chengxue Li and Chuanjiang He. Fractional-order diffusion coupled with integerorder diffusion for multiplicative noise removal. Computers and Mathematics with Applications, 136 (??):34–43, April 15, 2023. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122123000494.

Liu:2023:SGD

[LH23b] Shuaijun Liu and Pengzhan Huang. A sparse grad– div stabilized algorithm for the incompressible magnetohydrodynamics equations. Computers and Mathematics with Applications, 138 (??):106–119, May 15, 2023. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122123001281.

Lampropoulos:2024:IIV

[LH24] Dimitrios S. Lampropoulos and Maria Hadjinicolaou. Investigating the impact of vessel geometry on cerebral aneurysm formation using multi-phase blood flow models. Computers and Mathematics with Applications, 176(??):257–269, December 15, 2024. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122124004887.

Lee:2022:ABN

[LHC^{+}22] Chak Shing Lee, François P. Hamon, Nicola Castelletto, Panayot S. Vassilevski, and Joshua A. White. An aggregation-based nonlinear multigrid solver for twophase flow and transport in porous media. Computers and Mathematics with Applications, 113(??):282– 299, May 1, 2022. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122122001171.

Li:2023:APM

[LHC23] Qian Li, Yuxiao Hu, and Yang Cao. Adaptive Perona– Malik model based on dynamical threshold for image multi-noise removal with details preservation. Computers and Mathematics with Applications, 137(??):28– 43, May 1, 2023. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122123000718.

Li:2024:SEO

[LHF24] Yang Li, Chaolang Hu, and Minfu Feng. On stabilized equal-order virtual element methods for the Navier– Stokes equations on polygonal meshes. Computers and Mathematics with Applications, 154(??):267–286, January 15, 2024. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122123005357.

Li:2022:EAO

[LHH22] Xiaodong Li, Steven Hulshoff, and Stefan Hickel. An enhanced algorithm for online proper orthogonal decomposition and its parallelization for unsteady simulations. Computers and Mathematics with Applications, 126(??):43–59, November 15, 2022. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122122003807.

Le:2024:ACN

[LHN24] Anh Ha Le, Toan T. Huynh, and Quan M. Nguyen. Analysis of a Crank–Nicolson finite difference scheme for (2 + 1)D perturbed nonlinear Schrödinger equations with saturable nonlinearity. Computers and Mathematics with Applications, 167(??): 160–177, August 1, 2024. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S089812212400227X.

Lv:2025:NSP

[LHQW25] Qianyu Lv, Bing He, Chunyan Qin, and Binghai Wen. Numerical simulation for pulmonary airway reopening in alveolar duct by lattice Boltzmann method. Computers and Mathematics with Applications, 190(??):206– 218, July 15, 2025. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122125001956.

Liu:2024:DFE

[LHSF24] Yaru Liu, Yinnian He, Dongwoo Sheen, and Xinlong Feng. A difference finite element method based on the conforming P_{1}(x, y) × Q_{1}(z, s) element for the 4D Poisson equation. Computers and Mathematics with Applications, 174(??):18–30, November 15, 2024. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122124003717.

Liu:2025:OBP

[LHTZ25] Chen Liu, Jingwei Hu, William T. Taitano, and Xiangxiong Zhang. An optimization-based positivitypreserving limiter in semiimplicit discontinuous Galerkin schemes solving Fokker– Planck equations. Computers and Mathematics with Applications, 192(??):54–71, August 15, 2025. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122125002068.

Li:2020:NRD

[Li20]  Chenglin Li.  A nonlo-

cal reaction–diffusion prey– predator model with free boundary. Computers and Mathematics with Applications, 79(2):378–390, January 15, 2020. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122119303426.

Li:2021:IBF

[Li21] Ming-Jian Li. Interaction between free surface flow and moving bodies with a dynamic mesh and interface geometric reconstruction approach. Computers and Mathematics with Applications, 81(??):649–663, January 1, 2021. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122120300456.

Li:2024:SAF

[Li24a] Jichun Li. Superconvergence analysis of finite element approximations to Maxwell’s equations in both metamaterials and PMLs. Computers and Mathematics with Applications, 172(??):111–124, October 15, 2024. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122124003407.

Li:2024:EAL

[Li24b] Yuan Li. Error analysis of the linearized Crank– Nicolson FEM for the incompressible vector potential magnetohydrodynamic system. Computers and Mathematics with Applications, 157 (??):65–73, March 1, 2024. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122123005825.

Li:2025:NMR

[Li25] Xinyan Li. A numerical method for reconstructing the potential in fractional Calderón problem with a single measurement. Computers and Mathematics with Applications, 183(??):256– 270, April 1, 2025. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122125000719.

Lin:2024:FVR

[Lin24] Zeng Lin. Finite volume reproducing kernel particle method. Computers and Mathematics with Applications, 167(??):150–159, August 1, 2024. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122124002153.

Liu:2020:MHM

[Liu20]  Tao Liu.  A multigrid-

homotopy method for nonlinear inverse problems. Computers and Mathematics with Applications, 79(6):1706– 1717, March 15, 2020. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122119304791.

Liu:2025:STF

[Liu25a] Jinghong Liu. Superconvergence of triquadratic finite elements for the second-order elliptic equation with variable coefficients. Computers and Mathematics with Applications, 181(??):44– 54, March 1, 2025. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122125000033.

Liu:2025:FMT

[Liu25b] Yi Liu. Fast method and theoretical study of magnetohydrodynamic flow and heat transfer for fractional Maxwell fluids. Computers and Mathematics with Applications, 196(??):355–375, October 15, 2025. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122125003554.

Liang:2024:LUE

[LJ24] Yihong Liang and Hongen Jia. Linear and unconditionally energy stable schemes for the modified phase field crystal equation. Computers and Mathematics with Applications, 153(??):197– 212, January 1, 2024. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122123004960.

Liu:2025:SVA

[LJY^{+}25] Jun Liu, Chenxi Ji, Wenbin Ye, Lei Gan, Lei Qin, Quansheng Zang, and Haibo Wang. Static and vibration analyses of laminated conical shells under various boundary conditions using a modified scaled boundary finite element method. Computers and Mathematics with Applications, 177(??): 147–166, January 1, 2025. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122124005224.

Lee:2020:VMP

[LK20] Philku Lee and Seongjai Kim. A variable-θ method for parabolic problems of nonsmooth data. Computers and Mathematics with Applications, 79(4):962–981, February 15, 2020. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122119304031.

Lee:2024:OSM

[LKC^{+}24] Gyeonggyu Lee, Soobin Kwak, Yongho Choi, Seunggyu Lee, Seungyoon Kang, Seokjun Ham, and Junseok Kim. An operator splitting method for the Cahn–Hilliard equation on nonuniform grids. Computers and Mathematics with Applications, 167(??):207– 216, August 1, 2024. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122124002402.

Li:2020:BLF

[LL20a] Botong Li and Fawang Liu.

Boundary layer flows of viscoelastic fluids over a nonuniform permeable surface. Computers and Mathematics with Applications, 79(8): 2376–2387, April 15, 2020. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122119305346.

Li:2020:CVB

[LL20b] Xiaolin Li and Shuling Li. A complex variable boundary point interpolation method for the nonlinear Signorini problem. Computers and Mathematics with Applications, 79(12):3297– 3309, June 15, 2020. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122120300559.

Lee:2021:MAA

[LL21a] Young-Ju Lee and Hengguang Li. The macroelement analysis for axisymmetric Stokes equations. Computers and Mathematics with Applications, 101(??):1–22, November 1, 2021. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122121003266.

Li:2021:LEF

[LL21b] Xiaolin Li and Shuling Li. A linearized elementfree Galerkin method for the complex Ginzburg– Landau equation. Computers and Mathematics with Applications, 90(??):135– 147, May 15, 2021. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122121001024.

Li:2022:OEA

[LL22a] Chenyang Li and Yuan Li.

Optimal L^{2} error analysis of first-order Euler linearized finite element scheme for the 2D magnetohydrodynamics system with variable density. Computers and Mathematics with Applications, 128(??): 96–107, December 15, 2022. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122122004369.

Li:2022:MHM

[LL22b] Meng Li and Xianbing Luo.

An MLMCE-HDG method for the convection diffusion equation with random diffusivity. Computers and Mathematics with Applications, 127(??):127–143, December 1, 2022. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122122004187.

Liang:2022:NNF

[LL22c] Ke Liang and Zheng Li.

A novel nonlinear FE perturbation method and its application to stacking sequence optimization for snap-through response of cylindrical shell panel. Computers and Mathematics with Applications, 112(??):154– 166, April 15, 2022. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122122000864.

Li:2023:MGA

[LL23] Xiaolin Li and Shuling Li. Meshless Galerkin analysis of the generalized Stokes problem. Computers and Mathematics with Applications, 144(??):164–181, August 15, 2023. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122123002328.

Liu:2024:OLT

[LL24a] Demin Liu and Youlei Liang. One-level and two-level operator splitting methods for the unsteady incompressible micropolar fluid equations with double diffusion convection. Computers and Mathematics with Applications, 170(??): 56–83, September 15, 2024. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122124002839.

Liu:2024:CAF

[LL24b] Xing Liu and Hui Li. Convergence analysis of a fully discrete scheme for diffusionwave equation forced by tempered fractional Brownian motion. Computers and Mathematics with Applications, 169(??):39–55, September 1, 2024. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122124002724.

Lee:2025:UEG

[LL25a] Gyeonggyu Lee and Seunggyu Lee. Unconditionally energy gradient stable numerical scheme for Cahn– Hilliard equation with highorder degenerate mobility. Computers and Mathematics with Applications, 196(??): 263–287, October 15, 2025. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122125003001.

Lei:2025:UWU

[LL25b] Wei-Min Lei and Hou-Biao Li. U-WNO: U-Net enhanced wavelet neural operator for solving parametric partial differential equations. Computers and Mathematics with Applications, 194(??):272– 287, September 15, 2025. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S089812212500269X.

Li:2025:NSD

[LL25c] Zongcheng Li and Jin Li. Numerical solutions for (2 + 1)- dimensional ZK-MEW equation using linear barycentric rational collocation method. Computers and Mathematics with Applications, 193(??): 332–345, September 1, 2025. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122125002664.

Legrand:2023:CLO

[LLD23] C. Legrand, F. Luddens, and I. Danaila. On the convergence of a low order Lagrange finite element approach for natural convection problems. Computers and Mathematics with Applications, 147(??): 259–277, October 1, 2023. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122123003188.

Li:2025:CII

[LLG25] Gengchen Li, Hongwei Lin, and Depeng Gao. Composite iteration for isogeometric collocation method using LSPIA and Schulz iteration. Computers and Mathematics with Applications, 181 (??):111–132, March 1, 2025. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122124005789.

Liu:2023:NCF

[LLJ23] Yi Liu, Fawang Liu, and Xiaoyun Jiang. Numerical calculation and fast method for the magnetohydrodynamic flow and heat transfer of fractional Jeffrey fluid on a two-dimensional irregular convex domain. Computers and Mathematics with Applications, 151(??):473–490, December 1, 2023. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122123004613.

Liu:2024:AVT

[LLLL24] Xinyuan Liu, Nan Liu, Yang Liu, and Hong Li. Analysis of variable-time-step BDF2 combined with the fast two-grid finite element algorithm for the FitzHugh– Nagumo model. Computers and Mathematics with Applications, 170(??):186– 203, September 15, 2024. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122124003018.

Liu:2025:DAT

[LLMC25] Cong Liu, Kaifu Li, Shaosong Min, and Yingbin Chai. Dynamic analysis of the three-phase magneto-electro-elastic (MEE) structures with the overlapping triangular finite elements. Computers and Mathematics with Applications, 179(??):148– 177, February 1, 2025. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122124005236.

Laghrib:2024:EOM

[LLN24a] Amine Laghrib, Fakhr-Eddine Limami, and Abdeljalil Nachaoui. Enhanced osmosis model with bilateral total variation for effective shadow removal. Computers and Mathematics with Applications, 174(??):71–83, November 15, 2024. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122124003687.

Le:2024:CCM

[LLN24b] Huynh P. N. Le, Thuy T. Le, and Loc H. Nguyen. The Carleman convexification method for Hamilton– Jacobi equations. Computers and Mathematics with Applications, 159(??):173– 185, April 1, 2024. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S089812212400066X.

Lee:2025:NDD

[LLR25] Chang-Ock Lee, Youngkyu Lee, and Byungeun Ryoo. A nonoverlapping domain decomposition method for extreme learning machines: Elliptic problems. Computers and Mathematics with Applications, 189(??):109– 128, July 1, 2025. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122125001403.

Langer:2024:AFE

[LLSY24] Ulrich Langer, Richard Löscher, Olaf Steinbach, and Huidong Yang. An adaptive finite element method for distributed elliptic optimal control problems with variable energy regularization. Computers and Mathematics with Applications, 160 (??):1–14, April 15, 2024. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122124000452.

Langer:2025:RFE

[LLSY25] Ulrich Langer, Richard Löscher, Olaf Steinbach, and Huidong Yang. Robust finite element solvers for distributed hyperbolic optimal control problems. Computers and Mathematics with Applications, 180(??):166–190, February 15, 2025. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S089812212400556X.

Liu:2024:FPE

[LLU24] Tao Liu, Ting Li, and Malik Zaka Ullah. On five-point equidistant stencils based on Gaussian function with application in numerical multidimensional option pricing. Computers and Mathematics with Applications, 176(??): 35–45, December 15, 2024. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122124004127.

Li:2020:LDS

[LLW20] Fuzhi Li, Yangrong Li, and Renhai Wang. Limiting dynamics for stochastic reaction–diffusion equations on the Sobolev space with thin domains. Computers and Mathematics with Applications, 79(2):457–475, January 15, 2020. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122119303475.

Lin:2023:EEC

[LLW23a] Lei Lin, Junliang Lv, and Dan Wu. Error estimate of the cell-centered nonlinear positivity-preserving two-point flux approximation schemes. Computers and Mathematics with Applications, 137(??):1–13, May 1, 2023. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122123000743.

Liu:2023:TGM

[LLW^{+}23b] Qingfang Liu, Baotong Li, Yujie Wang, Zhiheng Wang, and Jiakun Zhao. A two-grid mixed finite element method of a phase field model for two-phase incompressible flows. Computers and Mathematics with Applications, 137(??):14–27, May 1, 2023. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122123000676.

Liu:2025:WPF

[LLWX25] Xiaojuan Liu, Maojun Li, Kun Wang, and Jiangming Xie. Well-posedness and finite element analysis for the elastic scattering problem with a modified DtN map. Computers and Mathematics with Applications, 177 (??):58–77, January 1, 2025. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122124005133.

Li:2023:NOM

[LLX23] Dinggen Li, Faqiang Li, and Bo Xu. Non-orthogonal multiple-relaxation-time lattice Boltzmann method for vorticity-streamfunction formulation. Computers and Mathematics with Applications, 152(??):308–316, December 15, 2023. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122123004728.

Li:2024:MRT

[LLX24] Dinggen Li, Faqiang Li, and Bo Xu. Multirelaxation-time lattice Boltzmann method for anisotropic convection–diffusion equation with divergence-free velocity field. Computers and Mathematics with Applications, 171(??):1–5, October 1, 2024. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122124003055.

Liu:2023:MMN

[LLXY23] Haowei Liu, Zhiyong Liu, Qiuyan Xu, and Jiye Yang. Meshfree methods for nonlinear equilibrium radiation diffusion equation with jump coefficient. Computers and Mathematics with Applications, 147(??):239– 258, October 1, 2023. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122123003206.

Liu:2025:MMN

[LLXY25] Haowei Liu, Zhiyong Liu, Qiuyan Xu, and Jiye Yang. Meshfree methods for nonlinear equilibrium radiation diffusion equation with interface and discontinuous coefficient. Computers and Mathematics with Applications, 178(??):108–135, January 15, 2025. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122124005297.

Liu:2023:NST

[LLXZ23] Jinghong Liu, Qiyong Li, Zhiguang Xiong, and Qiding Zhu. L^{∞}- and L^{2}- norms superconvergence of the tetrahedral quadratic finite element. Computers and Mathematics with Applications, 149(??):71–74, November 1, 2023. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122123003693.

Li:2025:MCM

[LLY25] Jian Li, Shaoxuan Li, and Jing Yue. MC-CDNNs: the Monte Carlo-coupled deep neural networks approach for stochastic dual-porosity-Stokes flow coupled model. Computers and Mathematics with Applications, 181 (??):1–20, March 1, 2025. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122124005765.

Li:2020:PWD

[LLZ20] Peng Li, Xiaoping Lu, and Song-Ping Zhu. Pricing weather derivatives with the market price of risk extracted from the utility indifference valuation. Computers and Mathematics with Applications, 79(12):3394– 3409, June 15, 2020. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122120300687.

Li:2021:STM

[LLZ21] Chunguang Li, Cuihua Li, and Hong Zheng. Subspace tracking method for non-smooth yield surface. Computers and Mathematics with Applications, 90 (??):125–134, May 15, 2021. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122121000869.

Liu:2023:GDF

[LLZ^{+}23] W. Liu, Z. J. Liu, Z. L.

Zhang, C. J. Teo, and C. Shu. Grad’s distribution function for 13 moments-based moment gas kinetic solver for steady and unsteady rarefied flows: Discrete and explicit forms. Computers and Mathematics with Applications, 137(??):112– 125, May 1, 2023. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122123000640.

Liu:2024:CNC

[LLZ^{+}24] Zhen-Yi Liu, Chong-Jun Li, Ying Zhang, Yan-Mei Jia, and Juan Chen. Construction of a new class of quadrilateral spline elements based on the scaled boundary coordinates. Computers and Mathematics with Applications, 156(??):129–156, February 15, 2024. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122124000014.

Li:2025:PVA

[LLZ^{+}25] C. Li, S. Li, Y. M. Zhang, J. W. Cai, and S. K. Lai. In-plane vibration analysis of elastically restrained FGM skew plates using variational differential quadrature method. Computers and Mathematics with Applications, 178(??):136–153, January 15, 2025. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122124005212.

Li:2024:RDE

[LLZL24] Huanrong Li, Yuejie Li, Yihui Zeng, and Zhendong Luo. A reduceddimension extrapolation twogrid Crank–Nicolson finite element method of unknown solution coefficient vectors for spatial fractional nonlinear Allen–Cahn equations. Computers and Mathematics with Applications, 167(??): 110–122, August 1, 2024. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S089812212400213X.

Lee:2024:LCP

[LM24] Seulip Lee and Lin Mu. A low-cost, penalty parameterfree, and pressure-robust enriched Galerkin method for the Stokes equations. Computers and Mathematics with Applications, 166 (??):51–64, July 15, 2024. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122124001883.

Lee:2024:CVU

[LMBB24] David Lee, Alberto F.

Martín, Christopher Bladwell, and Santiago Badia. A comparison of variational upwinding schemes for geophysical fluids, and their application to potential enstrophy conserving discretisations. Computers and Mathematics with Applications, 165 (??):150–162, July 1, 2024. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122124001755.

Li:2022:WBD

[LMD22] Maojun Li, Rushuang Mu, and Haiyun Dong. A well-balanced discontinuous Galerkin method for the shallow water flows on erodible bottom. Computers and Mathematics with Applications, 119(??):13–20, August 1, 2022. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122122002267.

Latt:2021:PPL

[LMK^{+}21] Jonas Latt, Orestis Malaspinas, Dimitrios Kontaxakis, Andrea Parmigiani, Daniel Lagrava, Federico Brogi, Mohamed Ben Belgacem, Yann Thorimbert, Sébastien Leclaire, Sha Li, Francesco Marson, Jonathan Lemus, Christos Kotsalos, Raphaël Conradin, Christophe Coreixas, Rémy Petkantchin, Franck Raynaud, Joël Beny, and Bastien Chopard. Palabos: Parallel lattice Boltzmann solver. Computers and Mathematics with Applications, 81(??):334–350, January 1, 2021. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122120301267.

Liang:2023:TMB

[LML23] Ke Liang, Jiaqi Mu, and Zheng Li. Thermal-mechanical buckling analysis and optimization of the stringer stiffened cylinder using smeared stiffener based reduced-order models. Computers and Mathematics with Applications, 143(??):108–118, August 1, 2023. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122123001840.

Liu:2023:CDT

[LMLZ23] Hongyan Liu, Yanying Ma, Hu Li, and Wei Zhang. Combination of discrete technique on graded meshes with barycentric rational interpolation for solving a class of time-dependent partial integro-differential equations with weakly singular kernels.

Computers and Mathematics with Applications, 141 (??):159–169, July 1, 2023. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122123001529.

Los:2020:IRM

[LMMMP20] M. Loś, J. Muñoz-Matute, I. Muga, and M. Paszyński. Isogeometric residual minimization method (iGRM) with direction splitting for non-stationary advection– diffusion problems. Computers and Mathematics with Applications, 79(2):213–229, January 15, 2020. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122119303268.

Los:2021:IRM

[ŁMMMP21] M. Łoś, I. Muga, J. Muñoz-Matute, and M. Paszyński. Isogeometric residual minimization (iGRM) for nonstationary Stokes and Navier– Stokes problems. Computers and Mathematics with Applications, 95(??):200–214, August 1, 2021. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122120304417.

LaRocca:2023:MMA

[LMP23] Michele La Rocca, Andrea Montessori, and Pietro Prestininzi. A mesoscopic model approach for polar fluid flow. Computers and Mathematics with Applications, 151(??): 326–334, December 1, 2023. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S089812212300439X.

Loli:2020:ESS

[LMST20] Gabriele Loli, Monica Montardini, Giancarlo Sangalli, and Mattia Tani. An efficient solver for space–time isogeometric Galerkin methods for parabolic problems. Computers and Mathematics with Applications, 80(11): 2586–2603, December 2020. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122120303709.

Liu:2023:FEC

[LMZ23] Ziqi Liu, Shuai Miao, and Zhimin Zhang. A family of edge-centered finite volume schemes for heterogeneous and anisotropic diffusion problems on unstructured meshes. Computers and Mathematics with Applications, 146(??):165–175, September 15, 2023. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122123002821.

Liu:2021:PPE

[LN21] Ying Liu and Yufeng Nie.

A priori and a posteriori error estimates of the weak Galerkin finite element method for parabolic problems. Computers and Mathematics with Applications, 99 (??):73–83, October 1, 2021. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122121002868.

Lu:2022:ROF

[LN22] Jiashu Lu and Yufeng Nie.

A reduced-order fast reproducing kernel collocation method for nonlocal models with inhomogeneous volume constraints. Computers and Mathematics with Applications, 121(??):52–61, September 1, 2022. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S089812212200270X.

Lawrence:2024:NSM

[LNF24] Andrew P. Lawrence, Morten E.

Nielsen, and Bengt Fornberg. Node subsampling for multilevel meshfree elliptic PDE solvers. Computers and Mathematics with Applications, 164(??):79– 94, June 15, 2024. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122124001305.

Le:2024:TDR

[LNNP24] Thuy T. Le, Linh V. Nguyen, Loc H. Nguyen, and Hyunha Park. The time dimensional reduction method to determine the initial conditions without the knowledge of damping coefficients. Computers and Mathematics with Applications, 166 (??):77–90, July 15, 2024. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122124001706.

Le:2022:CBN

[LNT22] Thuy T. Le, Loc H. Nguyen, and Hung V. Tran. A Carleman-based numerical method for quasilinear elliptic equations with overdetermined boundary data and applications. Computers and Mathematics with Applications, 125(??):13–24, November 1, 2022. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122122003522.

Lobo:2021:FSF

[Lob21]  Diogo Lobo.  Fast sta-

ble finite difference schemes for nonlinear cross-diffusion. Computers and Mathematics with Applications, 101(??): 23–37, November 1, 2021. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122121002492.

Lohmann:2021:AFC

[Loh21] Christoph Lohmann. An algebraic flux correction scheme facilitating the use of Newton-like solution strategies. Computers and Mathematics with Applications, 84(??):56–76, February 15, 2021. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122120304727.

Lopez:2020:PAC

[Lop20]  Luciano Lopez.  Pref-

ace: Advanced computational methods for PDEs. Computers and Mathematics with Applications, 79 (7):1889, April 2020. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122120300699.

Lopez:2022:E

[Lop22] Luciano Lopez. Editorial.

Computers and Mathematics with Applications, 116 (??):1, June 15, 2022. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122122002292.

Llorente:2020:CIR

[LP20] Víctor J. Llorente and Antonio Pascau. Compact integration rules as a quadrature method with some applications. Computers and Mathematics with Applications, 79(5):1241–1265, March 1, 2020. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122119304420.

Lee:2021:VFS

[LP21] Chaemin Lee and Jongho Park. A variational framework for the strain-smoothed element method. Computers and Mathematics with Applications, 94(??):76–93, July 15, 2021. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122121001723.

Lopez:2022:STD

[LP22a] Luciano Lopez and Sabrina Francesca Pellegrino. A space-time discretization of a nonlinear peridynamic model on a 2D lamina. Computers and Mathematics with Applications, 116(??):161– 175, June 15, 2022. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122121002674.

Lotfi:2022:AAG

[LP22b] Yasaman Lotfi and Kourosh Parand. Anti-aliasing of gray-scale/color/outline images: Looking through the lens of numerical approaches for PDE-based models. Computers and Mathematics with Applications, 113(??):130– 147, May 1, 2022. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S089812212200102X.

Lotfi:2022:EID

[LP22c] Yasaman Lotfi and Kourosh Parand. Efficient image denoising technique using the meshless method: Investigation of operator splitting RBF collocation method for two anisotropic diffusionbased PDEs. Computers and Mathematics with Applications, 113(??):315– 331, May 1, 2022. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122122000980.

Lee:2023:PFE

[LP23] Chaemin Lee and Jongho Park. Preconditioning for finite element methods with strain smoothing. Computers and Mathematics with Applications, 130(??):41–57, January 15, 2023. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122122004813.

Lee:2022:SRS

[LPK22] Philku Lee, George V. Popescu, and Seongjai Kim. Stable rotational symmetric schemes for nonlinear reaction-diffusion equations. Computers and Mathematics with Applications, 109 (??):191–203, March 1, 2022. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122122000335.

Li:2025:LBB

[LpWW^{+}25] Jia-Qing Li, Ya ping Wang, Zhi-Hui Wang, Tian-Pei He, Xiao-Dong Niu, Adnan Khan, and Hiroshi Yamaguchi. A lattice Boltzmann-based diffuse interface immersed-boundary scheme for gas-liquid-solid multiphase flow. Computers and Mathematics with Applications, 199(??):325–342, December 1, 2025. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122125004341.

Li:2020:NTG

[LQ20] Wenbin Li and Jianliang Qian. Newton-type Gauss– Seidel Lax–Friedrichs highorder fast sweeping methods for solving generalized eikonal equations at largescale discretization. Computers and Mathematics with Applications, 79(4):1222– 1239, February 15, 2020. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122119304341.

Liu:2024:ASE

[LQ24] Liya Liu and Xiaolong Qin. An accelerated stochastic extragradient-like algorithm with new stepsize rules for stochastic variational inequalities. Computers and Mathematics with Applications, 163(??):117– 135, June 1, 2024. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122124001329.

Liu:2020:ASS

[LQCM20] Guodong Liu, Haokun Qi, Zhengbo Chang, and Xinzhu Meng. Asymptotic stability of a stochastic May mutualism system. Computers and Mathematics with Applications, 79(3):735–745, February 1, 2020. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122119303748.

Liu:2025:SLR

[LQF25] Yajun Liu, Yuanyang Qiao, and Xinlong Feng. A semi-Lagrangian radial basis function partition of unity closest point method for advection–diffusion equations on surfaces. Computers and Mathematics with Applications, 177(??):100– 114, January 1, 2025. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S089812212400511X.

Lei:2023:IDC

[LQLF23] Weidong Lei, Xiaofei Qin, Hongjun Li, and Youhua Fan. An iterative decomposition coupling algorithm between TD-FEM and TD-BEM with independent spatial discretization on the interface. Computers and Mathematics with Applications, 148(??):151–161, October 15, 2023. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122123003450.

Lin:2022:LSB

[LQW22] Ji Lin, Lin Qiu, and Fajie Wang. Localized singular boundary method for the simulation of large-scale problems of elliptic operators in complex geometries. Computers and Mathematics with Applications, 105 (??):94–106, January 1, 2022. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122121004156.

Liu:2022:UON

[LQY22] Nan Liu, Hongyu Qin, and Yin Yang. Unconditionally optimal H ^{1}-norm error estimates of a fast and linearized Galerkin method for nonlinear subdiffusion equations. Computers and Mathematics with Applications, 107 (??):70–81, February 1, 2022. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122121004454.

Li:2023:NSP

[LQYJ23] Qiang Li, Fangcao Qu, Jinyun Yuan, and Tao Jiang. Numerical simulation of phase field model for spherulite growth of semicrystalline polymers using FD-FV-LB method. Computers and Mathematics with Applications, 142(??):293– 306, July 15, 2023. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122123001013.

Liu:2023:MPI

[LQZ23] Chaoyu Liu, Zhonghua Qiao, and Qian Zhang. Multiphase image segmentation by the Allen–Cahn Chan– Vese model. Computers and Mathematics with Applications, 141(??):207– 220, July 1, 2023. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122122005272.

Li:2025:TOS

[LQZ25] Jiajie Li, Meizhi Qian, and Shengfeng Zhu. Topology optimization of Stokes eigenvalues by a level set method. Computers and Mathematics with Applications, 188 (??):50–71, June 15, 2025. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122125001075.

Lukes:2022:HLD

[LR22] Vladimír Lukeš and Eduard Rohan. Homogenization of large deforming fluidsaturated porous structures.

Computers and Mathematics with Applications, 110 (??):40–63, March 15, 2022. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122122000505.

Liang:2023:NLF

[LR23] Hao Liang and Hongxing Rui. The nonconforming locking-free virtual element method for the Biot’s consolidation model in poroelasticity. Computers and Mathematics with Applications, 148(??):269–281, October 15, 2023. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122123003504.

Liu:2023:HPS

[LRZ23] Yang Liu, Maohua Ran, and Li Zhang. Hamiltonianpreserving schemes for the two-dimensional fractional nonlinear Schrödinger wave equations. Computers and Mathematics with Applications, 150(??):54–69, November 15, 2023. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122123003966.

Liang:2020:AEI

[LS20] Ke Liang and Qin Sun. An accurate and efficient implementation of initial geometrical imperfections in the predictor–corrector reducedorder modeling method. Computers and Mathematics with Applications, 79(12): 3429–3446, June 15, 2020. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122120300663.

Liu:2021:IMF

[LS21] Zhixin Liu and Shicang Song. An isoparametric mixed finite element method for approximating a class of fourth-order elliptic problem. Computers and Mathematics with Applications, 96 (??):77–94, August 15, 2021. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122121001978.

Laymuns:2022:CFE

[LS22a] Genaro Laymuns and Manuel A. Sánchez. Corrected finite element methods on unfitted meshes for Stokes moving interface problem. Computers and Mathematics with Applications, 108(??):159–174, February 15, 2022. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122121004582.

Liu:2022:HNC

[LS22b] Xinwu Liu and Ting Sun.

Hybrid non-convex regularizers model for removing multiplicative noise. Computers and Mathematics with Applications, 126(??):182– 195, November 15, 2022. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122122003923.

Liu:2023:SEA

[LS23] Qian Liu and Dongyang Shi.

Superconvergence error analysis of an efficient mixed finite element method for time-dependent natural convection problem. Computers and Mathematics with Applications, 131(??):68–81, February 1, 2023. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122122005016.

Liang:2025:USA

[LS25a] Tiantian Liang and Dongyang Shi. Unconditional superconvergence analysis of a novel energy dissipation nonconforming Crank–Nicolson FEM for Sobolev equations with high order Burgers’ type nonlinearity. Computers and Mathematics with Applications, 180(??):86–101, February 15, 2025. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122124005558.

Lin:2025:FSB

[LS25b] Chuandong Lin and Chang Shu. A flux solver based on discrete Boltzmann method for compressible flows with nonequilibrium effects. Computers and Mathematics with Applications, 195(??):14–27, October 1, 2025. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122125002810.

Liang:2025:PRL

[LSF25] Sisi Liang, Haiyan Su, and Xinlong Feng. A pressurerobust lowest-order virtual element method for the natural convection problem on general mesh. Computers and Mathematics with Applications, 197(??):295–316, November 1, 2025. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122125003815.

Liu:2023:ASW

[LSGN23] Ying Liu, Xiaoqin Shen, Zhen Guan, and Yufeng Nie. The adaptive SAV weak Galerkin finite element method for the Allen– Cahn equation. Computers and Mathematics with Applications, 151(??):449–460, December 1, 2023. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122123004704.

Li:2021:NAS

[LSH21] Huanrong Li, Zhengyuan Song, and Junzhao Hu. Numerical analysis of a secondorder IPDGFE method for the Allen–Cahn equation and the curvature-driven geometric flow. Computers and Mathematics with Applications, 86(??):49–62, March 15, 2021. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122121000249.

Lopez-Salas:2024:PPI

[LSPRV24] José Germán López-Salas, Soledad Pérez-Rodríguez, and Carlos Vázquez. PDEs for pricing interest rate derivatives under the new generalized Forward Market Model (FMM). Computers and Mathematics with Applications, 169(??):88–98, September 1, 2024. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122124002785.

Lin:2024:NTM

[LSRL24] Ji Lin, Lianpeng Shi, Sergiy Reutskiy, and Jun Lu. Numerical treatment of multidimensional time-fractional Benjamin–Bona–Mahony– Burgers equations in arbitrary domains with a novel improvised RBF-based method. Computers and Mathematics with Applications, 167(??):178–198, August 1, 2024. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S089812212400230X.

Los:2023:SAS

[ŁSSP23] Marcin Łoś, Paulina Sepúlveda, Maciej Sikora, and Maciej Paszyński. Solver algorithm for stabilized space-time formulation of advection-dominated diffusion problem. Computers and Mathematics with Applications, 152(??):67–80, December 15, 2023. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122123004583.

Loli:2022:EEP

[LST22] Gabriele Loli, Giancarlo Sangalli, and Mattia Tani. Easy and efficient preconditioning of the isogeometric mass matrix. Computers and Mathematics with Applications, 116 (??):245–264, June 15, 2022. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122120304715.

Le:2021:EET

[LSV21] Van Chien Le, Marián Slodička, and Karel Van Bockstal. Error estimates for the time discretization of an electromagnetic contact problem with moving non-magnetic conductor. Computers and Mathematics with Applications, 87 (??):27–40, April 1, 2021. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122121000353.

Le:2024:NSS

[LSV24] Van Chien Le, Marián Slodička, and Karel Van Bockstal. A numerical scheme for solving an induction heating problem with moving non-magnetic conductor. Computers and Mathematics with Applications, 171(??):60–81, October 1, 2024. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122124003134.

Li:2020:USA

[LSW20] Meng Li, Dongyang Shi, and Junjun Wang. Unconditional superconvergence analysis of a linearized Crank–Nicolson Galerkin FEM for generalized Ginzburg–Landau equation. Computers and Mathematics with Applications, 79(8):2411–2425, April 15, 2020. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122119305413.

Lafontaine:2022:WEC

[LSW22] D. Lafontaine, E. A. Spence, and J. Wunsch. Wavenumberexplicit convergence of the hp -FEM for the full-space heterogeneous Helmholtz equation with smooth coefficients. Computers and Mathematics with Applications, 113 (??):59–69, May 1, 2022. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122122000918.

Liu:2021:VEM

[LSWY21] Yang Liu, Shi Shu, Huayi Wei, and Ying Yang. A virtual element method for the steady-state Poisson– Nernst–Planck equations on polygonal meshes. Computers and Mathematics with Applications, 102(??):95– 112, November 15, 2021. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122121003503.

Liu:2024:HOG

[LSY^{+}24] Yaguang Liu, Chang Shu, Peng Yu, Yangyang Liu, Hua Zhang, and Chun Lu. A highorder generalised differential quadrature element method for simulating 2D and 3D incompressible flows on unstructured meshes. Computers and Mathematics with Applications, 174(??):230– 247, November 15, 2024. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122124003870.

Liu:2024:VEM

[LSZ24a] Shuo Liu, Wanfang Shen, and Zhaojie Zhou. Virtual element method for semilinear elliptic Neumann boundary optimal control problem. Computers and Mathematics with Applications, 168(??): 218–238, August 15, 2024. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122124002980.

Loscher:2024:MHT

[LSZ24b] Richard Löscher, Olaf Steinbach, and Marco Zank. On a modified Hilbert transformation, the discrete inf-sup condition, and error estimates.

Computers and Mathematics with Applications, 171(??): 114–138, October 1, 2024. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122124003080.

Li:2023:TGA

[LT23] Kang Li and Zhijun Tan. Two-grid algorithms based on FEM for nonlinear timefractional wave equations with variable coefficient. Computers and Mathematics with Applications, 143(??): 119–132, August 1, 2023. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122123001839.

Liu:2024:CAM

[LTXZ24] Kai Liu, Ming Tang, Xiaoqing Xing, and Liuqiang Zhong. Convergence of adaptive mixed interior penalty discontinuous Galerkin methods for H(curl)-elliptic problems. Computers and Mathematics with Applications, 170 (??):84–94, September 15, 2024. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122124002888.

Liu:2025:SOU

[LTYZ25] Yiyi Liu, Xueqing Teng, Xiaoqiang Yan, and Hong Zhang. A second-order, unconditionally invariant-set-preserving scheme for the FitzHugh–Nagumo equation. Computers and Mathematics with Applications, 189 (??):161–175, July 1, 2025. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122125001579.

Ly:2025:HAP

[LVDDNT25] Duy-Khuong Ly, Huy-Cuong Vu-Do, Trung-Hau Dang, and T. Nguyen-Thoi. Hybrid active-passive control of multiphase composite shells through high-order Chebyshev frequency response modeling. Computers and Mathematics with Applications, 199(??):184–202, December 1, 2025. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122125003918.

Liu:2024:MUF

[LWFS24] Wei Liu, Zhifeng Wang, Gexian Fan, and Yingxue Song. Modified upwind finite volume scheme with second-order Lagrange multiplier method for dimensionally reduced transport model in intersecting fractured porous media. Computers and Mathematics with Applications, 175(??):202– 223, December 1, 2024. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122124004346.

Li:2025:GAP

[LWG^{+}25a] Wensheng Li, Chuncheng Wang, Hanting Guan, Jian Wang, Jie Yang, Chao Zhang, and Dacheng Tao. Generative adversarial physicsinformed neural networks for solving forward and inverse problem with small labeled samples. Computers and Mathematics with Applications, 183(??):98– 120, April 1, 2025. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S089812212500032X.

Li:2025:AAP

[LWG^{+}25b] Wensheng Li, Hao Wang, Hanting Guan, Ruifeng Zhou, Chao Zhang, and Dacheng Tao. AdaPW: an adaptive point-weighting method for training physicsinformed neural networks. Computers and Mathematics with Applications, 198(??): 255–273, November 15, 2025. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122125003591.

Li:2020:ADG

[LWL^{+}20] Chuan Li, Zhihan Wei, Guangqing Long, Cameron Campbell, Stacy Ashlyn, and Shan Zhao. Alternating direction ghost-fluid methods for solving the heat equation with interfaces. Computers and Mathematics with Applications, 80(5):714–732, September 1, 2020. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122120301747.

Li:2021:GVC

[LWL21] Sheng-Kun Li, Mao-Xiao Wang, and Gang Liu. A global variant of the COCR method for the complex symmetric Sylvester matrix equation ax + xb = C. Computers and Mathematics with Applications, 94 (??):104–113, July 15, 2021. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122121001735.

Li:2025:TFO

[LWL25] Yiqun Li, Hong Wang, and Wuchen Li. A time-fractional optimal transport: Formulation and algorithm. Computers and Mathematics with Applications, 180(??):261– 278, February 15, 2025. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122125000094.

Los:2023:FPI

[ŁWP^{+}23] Marcin Łoś, Maciej Woźniak, Keshav Pingali, Luis Emilio Garcia Castillo, Julen Alvarez-Arramberri, David Pardo, and Maciej Paszyński. Fast parallel IGA-ADS solver for time-dependent Maxwell’s equations. Computers and Mathematics with Applications, 151(??):36–49, December 1, 2023. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122123004200.

Li:2021:NAU

[LWSZ21] Huanrong Li, Dongmei Wang, Zhengyuan Song, and Fuchen Zhang. Numerical analysis of an unconditionally energy-stable reduced-order finite element method for the Allen–Cahn phase field model. Computers and Mathematics with Applications, 96(??):67–76, August 15, 2021. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122121001942.

Li:2024:CEW

[LWW24] Dan Li, Chunmei Wang, and Junping Wang. Curved elements in weak Galerkin finite element methods. Computers and Mathematics with Applications, 153(??):20–32, January 1, 2024. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122123005011.

Li:2020:ATD

[LWX20] Binjie Li, Tao Wang, and Xiaoping Xie. Analysis of a temporal discretization for a semilinear fractional diffusion equation. Computers and Mathematics with Applications, 80(10):2115– 2134, November 15, 2020. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122120303643.

Li:2021:ALS

[LWX21] Binjie Li, Tao Wang, and Xiaoping Xie. Analysis of the L1 scheme for fractional wave equations with nonsmooth data. Computers and Mathematics with Applications, 90(??):1–12, May 15, 2021. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122121000742.

Li:2022:DDM

[LWZ22] Sicheng Li, Shun Wang, and Xiufen Zou. Data-driven mathematical modeling and quantitative analysis of cell dynamics in the tumor microenvironment. Computers and Mathematics with Applications, 113(??):300– 314, May 1, 2022. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122122000967.

Liang:2022:FMM

[LX22] Jiangli Liang and Shuhuang Xiang. A fast multipole method for Fredholm integral equations of the second kind with general kernel K(x, y) = K(x − y). Computers and Mathematics with Applications, 118 (??):237–247, July 15, 2022. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122122002012.

Li:2024:CNA

[LX24a] Xindong Li and Wenwen Xu. Conservative numerical algorithm for simulating thermoelectrical semiconductor device with unconditional optimal convergence analysis. Computers and Mathematics with Applications, 176(??): 1–11, December 15, 2024. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S089812212400405X.

Linss:2024:BNE

[LX24b] Torsten Linß and Christos Xenophontos. Balanced norm estimates for rp-finite element methods applied to singularly perturbed fourth order boundary value problems. Computers and Mathematics with Applications, 164 (??):95–103, June 15, 2024. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122124001433.

Li:2024:SDC

[LXG^{+}24] Haichao Li, Jiawei Xu, Qingtao Gong, Yao Teng, Fuzhen Pang, and Ling Zhang. A study on the dynamic characteristics of the stiffened coupled plate with the effect of the dynamic vibration absorbers. Computers and Mathematics with Applications, 168(??):120– 132, August 15, 2024. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122124001913.

Ling:2024:NAH

[LXH24] Min Ling, Wenqiang Xiao, and Weimin Han. Numerical analysis of a historydependent mixed hemivariationalvariational inequality in contact problems. Computers and Mathematics with Applications, 166(??):65– 76, July 15, 2024. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122124001895.

Li:2021:TIE

[LXJ^{+}21] Hanqing Li, Bonan Xu, Hanhui Jin, Haiou Wang, and Jianren Fan. Two improved electronegativity equalization methods for charge distribution in large scale nonuniform system. Computers and Mathematics with Applications, 81(??):693–701, January 1, 2021. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122119305838.

Liu:2025:HOM

[LXL25] Guoliang Liu, Wenwen Xu, and Xindong Li. Higher order multipoint flux mixed finite element methods for parabolic equation. Computers and Mathematics with Applications, 189(??):144– 160, July 1, 2025. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122125001580.

Li:2024:UJR

[LXP^{+}24] Haichao Li, Jiawei Xu, Fuzhen Pang, Cong Gao, and Jiajun Zheng. A unified Jacobi–Ritz-spectral BEM for vibro-acoustic behavior of spherical shell. Computers and Mathematics with Applications, 176(??):415– 431, December 15, 2024. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122124004747.

Li:2022:SOP

[LXRQ22] Yi Li, Dandan Xue, Yao Rong, and Yi Qin. A second order partitioned method with grad-div stabilization for the non-stationary dualporosity-Stokes model. Computers and Mathematics with Applications, 124(??):111– 128, October 15, 2022. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122122003455.

Li:2023:LSS

[LXY23] Binjie Li, Xiaoping Xie, and Yubin Yan. L1 scheme for solving an inverse problem subject to a fractional diffusion equation. Computers and Mathematics with Applications, 134(??):112– 123, March 15, 2023. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122123000081.

Li:2023:CAS

[LXZ23] Shishun Li, Lin Xie, and Lingling Zhou. Convergence analysis of spacetime domain decomposition method for parabolic equations. Computers and Mathematics with Applications, 139 (??):209–215, June 1, 2023. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122122003327.

Liu:2024:PEE

[LXZZ24] Chunmei Liu, Yingying Xie, Liuqiang Zhong, and Liping Zhou. A posteriori error estimate of a weak Galerkin finite element method for solving linear elasticity problems. Computers and Mathematics with Applications, 173(??): 47–59, November 1, 2024. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S089812212400333X.

Liu:2021:MFE

[LY21] Xinfei Liu and Xiaoyuan Yang. Mixed finite element method for the nonlinear time-fractional stochastic fourth-order reaction– diffusion equation. Computers and Mathematics with Applications, 84(??):39–55, February 15, 2021. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122120304661.

Li:2023:ATS

[LY23] Huanrong Li and Rushuang Yang. Analysis of two spectral Galerkin approximation schemes for solving the perturbed FitzHugh–Nagumo neuron model. Computers and Mathematics with Applications, 143(??):1–9, August 1, 2023. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S089812212300175X.

Li:2024:UBC

[LY24] Changfeng Li and Yirang Yuan. Upwind blockcentered multistep differences for semiconductor device problem with heat and magnetic influences. Computers and Mathematics with Applications, 174(??):55–70, November 15, 2024. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122124003651.

Li:2025:ASG

[LY25] Huanrong Li and Rushuang Yang. Analysis of spectral Galerkin method with higher order time discretization for the nonlinear stochastic Fisher’s type equation driven by multiplicative noise. Computers and Mathematics with Applications, 190 (??):90–102, July 15, 2025. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122125001610.

Li:2020:ERP

[LYAZ20] Jian Li, Min Yao, Md. Abdullah Al Mahbub, and Haibiao Zheng. The efficient rotational pressurecorrection schemes for the coupling Stokes/Darcy problem. Computers and Mathematics with Applications, 79(2):337–353, January 15, 2020. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122119303360.

Liang:2024:ROM

[LYH24] Ke Liang, Zhen Yin, and Qiuyang Hao. A reducedorder method for geometrically nonlinear analysis of the wing-upper-skin panels in the presence of buckling. Computers and Mathematics with Applications, 170(??): 42–55, September 15, 2024. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122124002827.

Liu:2022:ADI

[LYL^{+}22] Yanqin Liu, Xiuling Yin, Fawang Liu, Xiaoyi Xin, Yanfeng Shen, and Libo Feng. An alternating direction implicit Legendre spectral method for simulating a 2D multi-term time-fractional Oldroyd-b fluid type diffusion equation. Computers and Mathematics with Applications, 113 (??):160–173, May 1, 2022. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122122001043.

Lyu:2025:CAV

[LYL25] Tong Lyu, Xingyang Ye, and Xiaoyue Liu. Convergence analysis of variable-time-step BDF2/spectral approximations for optimal control problems governed by linear reaction–diffusion equations. Computers and Mathematics with Applications, 189(??):48–69, July 1, 2025. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S089812212500118X.

Liu:2025:TDF

[LYLH25] Fuhao Liu, Wei Yang, Jichun Li, and Yunqing Huang. A time-domain finite element method for Kerr and Raman type nonlinear hyperbolic metamaterials with application for enhanced thirdharmonic generation. Computers and Mathematics with Applications, 198(??):178– 195, November 15, 2025. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122125003529.

Lee:2022:PFC

[LYP^{+}22] Chaeyoung Lee, Sungha Yoon, Jintae Park, Hyundong Kim, Yibao Li, Darae Jeong, Sangkwon Kim, Soobin Kwak, and Junseok Kim. Phase-field computations of anisotropic ice crystal growth on a spherical surface. Computers and Mathematics with Applications, 125(??):25–33, November 1, 2022. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122122003613.

Lin:2021:MRB

[LYRW21] Ji Lin, Hao Yu, Sergiy Reutskiy, and Yuan Wang. A meshless radial basis function based method for modeling dual-phase-lag heat transfer in irregular domains. Computers and Mathematics with Applications, 85 (??):1–17, March 1, 2021. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122121000158.

Lu:2021:COP

[LYXC21] J. Lu, C. Yu, W. Xu, and C. Chiu. Characteristic orthogonal polynomials-Ritz method for vibration behavior of functionally graded piezoelectric plates using FSDT. Computers and Mathematics with Applications, 98(??):157–168, September 15, 2021. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122121002698.

Liu:2023:GNA

[LYZ23] X. Liu, Z. W. Yang, and Y. M. Zeng. Global numerical analysis of an improved IMEX numerical scheme for a reaction diffusion SIS model in advective heterogeneous environments. Computers and Mathematics with Applications, 144(??):264– 273, August 15, 2023. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122123002705.

Li:2020:PIM

[LZ20a] Rui-Xia Li and Guo-Feng Zhang. Preconditioned iterative methods for the convective FitzHugh–Nagumo equations. Computers and Mathematics with Applications, 80(12):2915–2924, December 15, 2020. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122120304181.

Li:2020:ESC

[LZ20b] Xin Li and Luming Zhang. An efficient spectral-collocation difference method for twodimensional Schrödinger equation with Neumann boundary conditions. Computers and Mathematics with Applications, 79(8):2322– 2335, April 15, 2020. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122119305371.

Li:2020:OSS

[LZ20c] Xinfu Li and Junying Zhao. Orbital stability of standing waves for Schrödinger type equations with slowly decaying linear potential. Computers and Mathematics with Applications, 79(2): 303–316, January 15, 2020. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122119303335.

Lin:2020:NPC

[LZ20d] Sha Lin and Song-Ping Zhu.

Numerically pricing convertible bonds under stochastic volatility or stochastic interest rate with an ADI-based predictor–corrector scheme. Computers and Mathematics with Applications, 79(5): 1393–1419, March 1, 2020. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122119304444.

Liu:2020:NRB

[LZ20e] Ling Liu and Jiashan Zheng.

A new result for boundedness in the quasilinear parabolicparabolic Keller–Segel model (with logistic source). Computers and Mathematics with Applications, 79(4):1208– 1221, February 15, 2020. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122119304328.

Li:2021:PEE

[LZ21] Yuwen Li and Ludmil Zikatanov. A posteriori error estimates of finite element methods by preconditioning. Computers and Mathematics with Applications, 91 (??):192–201, June 1, 2021. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122120303072.

Li:2022:NFD

[LZ22] Liang Li and Jun Zhu. A new finite difference mapped unequal-sized WENO scheme for Hamilton–Jacobi equations. Computers and Mathematics with Applications, 119 (??):68–78, August 1, 2022. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122122002334.

Liu:2023:SPP

[LZ23] Xiaowei Liu and Jin Zhang. Supercloseness and postprocessing of finite element method in a balanced norm for singularly perturbed reaction–diffusion equation. Computers and Mathematics with Applications, 142(??): 1–8, July 15, 2023. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122123001311.

Li:2025:NIL

[LZ25a] Ziyan Li and Shun Zhang. Non-intrusive least-squares functional a posteriori error estimator: Linear and nonlinear problems with plain convergence. Computers and Mathematics with Applications, 191(??):275–295, August 1, 2025. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122125002111.

Liang:2025:LSV

[LZ25b] Yuxiang Liang and Shun Zhang. Least-squares versus partial least-squares finite element methods: Robust a priori and a posteriori error estimates of augmented mixed finite element methods. Computers and Mathematics with Applications, 184 (??):45–64, April 15, 2025. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122125000641.

Lu:2025:MMG

[LZ25c] Kang-Ya Lu and Xiao-Yun Zhang. Multigrid method with greedy partial block Jacobi smoother for solving two-dimensional spacefractional diffusion equations. Computers and Mathematics with Applications, 191(??):245–254, August 1, 2025. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S089812212500210X.

Li:2023:NTD

[LZA23] Jichun Li, Li Zhu, and Todd Arbogast. A new time-domain finite element method for simulating surface plasmon polaritons on graphene sheets. Computers and Mathematics with Applications, 142(??):268– 282, July 15, 2023. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S089812212300202X.

Lu:2025:LBI

[LZB25] Jinhua Lu, Song Zhao, and Pierre Boivin. A lattice-Boltzmann inspired finite volume solver for compressible flows. Computers and Mathematics with Applications, 187(??):50– 71, June 1, 2025. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122125001026.

Li:2024:DGM

[LZBC24] Rui Li, Mingbo Zhou, Linlin Bu, and Zhangxin Chen. Discontinuous Galerkin method for the coupled dual-porosity-Brinkman model. Computers and Mathematics with Applications, 163(??):84– 102, June 1, 2024. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S089812212400124X.

Li:2023:NGK

[LZF23] Weidong Li, Jinshan Zhao, and Ming Fang. A new gas kinetic BGK scheme based on the characteristic solution of the BGK model equation for viscous flows. Computers and Mathematics with Applications, 146(??):225– 240, September 15, 2023. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S089812212300295X.

Liu:2022:CSC

[LZFZ22] Zhongyun Liu, Fang Zhang, Carla Ferreira, and Yulin Zhang. On circulant and skew-circulant splitting algorithms for (continuous) Sylvester equations. Computers and Mathematics with Applications, 109(??):30– 43, March 1, 2022. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122122000347.

Liu:2023:TMB

[LZG^{+}23] Lufeng Liu, Xuan Zhou, Shaodong Guo, Yibing Chen, and Haibing Zhou. Topology modification based rezoning strategies for the arbitrary Lagrangian–Eulerian method on unstructured hexahedral meshes. Computers and Mathematics with Applications, 145(??):175–201, September 1, 2023. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122123002651.

Li:2021:PSB

[LZJC21] Chong-Jun Li, Ying Zhang, Yan-Mei Jia, and Juan Chen. The polygonal scaled boundary thin plate element based on the discrete Kirchhoff theory. Computers and Mathematics with Applications, 97(??):223–236, September 1, 2021. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122121002376.

Liu:2021:RPI

[LZL^{+}21] Qingxia Liu, Pinghui Zhuang, Fawang Liu, Minling Zheng, and Shanzhen Chen. Radial point interpolation collocation method based approximation for 2D fractional equation models. Computers and Mathematics with Applications, 97(??):153–161, September 1, 2021. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122121001875.

Li:2024:DSF

[LZL24] Jian Li, Xinyue Zhang, and Ruixia Li. A decoupled stabilized finite element method for nonstationary stochastic shale oil model based on superhydrophobic material modification. Computers and Mathematics with Applications, 172(??):94–110, October 15, 2024. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122124003389.

Liu:2025:MAO

[LZL^{+}25] Lin Liu, Sen Zhang, Lang Liu, Chiyu Xie, Lixiu Wang, Yijia Liu, Jing Zhu, and Libo Feng. The mechanism analysis of the oscillatory flow of Oldroyd-B fluid in a tube with a right triangular cross-section. Computers and Mathematics with Applications, 178(??):179–196, January 15, 2025. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S089812212400539X.

Liu:2025:IBK

[LZLN25] L. Liu, L. L. Zhang, M. Lei, and R. P. Niu. The improved boundary knot method with fictitious points for solving high-order Helmholtztype PDEs. Computers and Mathematics with Applications, 191(??):36–47, August 1, 2025. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122125001622.

Liu:2023:SAP

[LZS23] Nabing Liu, Lin Zhu, and Qin Sheng. A semiadaptive preservative scheme for a fractional quenching convective–diffusion problem. Computers and Mathematics with Applications, 151(??):288–299, December 1, 2023. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122123004285.

Liu:2025:USF

[LZS25] Hui Liu, Peizhi Zhao, and Tao Song. U-shaped factorized Fourier neural operator for solving partial differential equations. Computers and Mathematics with Applications, 196(??):233–245, October 15, 2025. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122125003013.

Liao:2021:CPI

[LZW21a] Li-Dan Liao, Guo-Feng Zhang, and Xiang Wang. Corrigendum to “Preconditioned iterative method for nonsymmetric saddle point linear systems” [Comput.

Math. Appl. 98 (2021) 69– 80]. Computers and Mathematics with Applications, 99(??):1, October 1, 2021. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122121002807.

See [LZW21b].

Liao:2021:PIM

[LZW21b] Li-Dan Liao, Guo-Feng Zhang, and Xiang Wang. Preconditioned iterative method for nonsymmetric saddle point linear systems. Computers and Mathematics with Applications, 98(??):69–80, September 15, 2021. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122121002583.

See corrigendum [LZW21a].

Li:2024:NLA

[LZY24a] Caojie Li, Haixiang Zhang, and Xuehua Yang. A new linearized ADI compact difference method on graded meshes for a nonlinear 2D and 3D PIDE with a WSK. Computers and Mathematics with Applications, 176(??): 349–370, December 15, 2024. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122124004978.

Liu:2024:VDG

[LZY24b] Kai Liu, Mingqian Zhang, and Xiong You. A variant of the discrete gradient method for the solution of the semilinear wave equation under different boundary conditions. Computers and Mathematics with Applications, 158(??):199– 218, March 15, 2024. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122124000464.

Mei:2022:EES

[MA22] Yanhua Mei and Rong An. Error estimates of secondorder BDF Galerkin finite element methods for a coupled nonlinear Schrödinger system. Computers and Mathematics with Applications, 122(??):117–125, September 15, 2022. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122122003091.

Ma:2025:EAM

[Ma25a] Huanhuan Ma. Error analysis of the moving least square material point method for large deformation problems. Computers and Mathematics with Applications, 193(??): 315–331, September 1, 2025. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122125002743.

Ma:2025:ESM

[Ma25b] Suna Ma. Efficient spectral method for the fourth order elliptic equation with a variable coefficient on the unit disc. Computers and Mathematics with Applications, 181 (??):340–350, March 1, 2025. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S089812212500029X.

Mallik:2022:GOP

[Mal22] Gouranga Mallik. Goaloriented a posteriori error estimation for the biharmonic problem based on an equilibrated moment tensor. Computers and Mathematics with Applications, 117 (??):312–325, July 1, 2022. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122122001778.

Manohar:2024:EAF

[Man24] Ram Manohar. Error analysis for finite element approximation of parabolic Neumann boundary control problems. Computers and Mathematics with Applications, 158(??):102– 117, March 15, 2024. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S089812212400004X.

Mascotto:2023:RSV

[Mas23] Lorenzo Mascotto. The role of stabilization in the virtual element method: a survey. Computers and Mathematics with Applications, 151(??): 244–251, December 1, 2023. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122123004303.

Mashayekhi:2020:HTE

[MAT^{+}20] Ramin Mashayekhi, Hossein Arasteh, Davood Toghraie, S. Hossein Motaharpour, Amir Keshmiri, and Masoud Afrand. Heat transfer enhancement of water– Al_{2}O_{3} nanofluid in an oval channel equipped with two rows of twisted conical strip inserts in various directions: a two-phase approach. Computers and Mathematics with Applications, 79(8): 2203–2215, April 15, 2020. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S089812211930519X.

Marquardt:2021:NEE

[MAT^{+}21] Jan E. Marquardt, Carsten-Rene Arlt, Robin Trunk, Matthias Franzreb, and Mathias J. Krause. Numerical and experimental examination of the retention of magnetic nanoparticles in magnetic chromatography. Computers and Mathematics with Applications, 89(??):34–43, May 1, 2021. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122121000523.

Mirfatah:2021:SIB

[MB21] S. M. Mirfatah and B. Boroomand. On the simulation of image-based cellular materials in a meshless style. Computers and Mathematics with Applications, 100(??): 99–125, October 15, 2021. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122121003138.

Magueresse:2024:AQN

[MB24] Alexandre Magueresse and Santiago Badia. Adaptive quadratures for nonlinear approximation of lowdimensional PDEs using smooth neural networks. Computers and Mathematics with Applications, 162 (??):1–21, May 15, 2024. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122124000877.

Mallik:2024:LPS

[MBG24] Gouranga Mallik, Rahul Biswas, and Thirupathi Gudi. A local projection stabilised HHO method for the Oseen problem. Computers and Mathematics with Applications, 176(??):202– 220, December 15, 2024. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122124004735.

Moutahir:2024:MSR

[MBSEA24] Fatima-Ezzahrae Moutahir, Youssef Belhamadia, Mohammed Seaid, and Mofdi El-Amrani. Modelling and simulation of radiative heat transfer in non-grey absorbing and emitting media under phase change. Computers and Mathematics with Applications, 176(??):432– 446, December 15, 2024. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122124004966.

Ma:2024:EFD

[MC24] Suna Ma and Hu Chen. Efficient finite difference/ spectral approximation for the time-fractional diffusion equation with an inverse square potential on the unit ball. Computers and Mathematics with Applications, 167(??):232– 238, August 1, 2024. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122124002323.

Mazzanti:2025:PVC

[MC25] Federico Mazzanti and Philip Cardiff. Performance of a vertex-centred block-coupled finite volume methodology for small-strain static elastoplasticity. Computers and Mathematics with Applications, 195(??):212–238, October 1, 2025. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122125003098.

Ma:2023:CAE

[MCD23] Huanhuan Ma, Jingrun Chen, and Jiansong Deng. Convergence analysis of the element-free Galerkin method for plate bending problem. Computers and Mathematics with Applications, 136(??):71– 79, April 15, 2023. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122123000524.

Ma:2024:EAE

[MCD24] Huanhuan Ma, Jingrun Chen, and Jiansong Deng.

Error analysis of the elementfree Galerkin method for a nonlinear plate problem. Computers and Mathematics with Applications, 163 (??):56–65, June 1, 2024. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122124001275.

Melchers:2023:CNC

[MCK^{+}23] Hugo Melchers, Daan Crommelin, Barry Koren, Vlado Menkovski, and Benjamin Sanderse. Comparison of neural closure models for discretised PDEs. Computers and Mathematics with Applications, 143(??):94–107, August 1, 2023. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122123001736.

Munar:2024:RBP

[MCV24] Mauricio Munar, Andrea Cangiani, and Iván Velásquez. Residual-based a posteriori error estimation for mixed virtual element methods. Computers and Mathematics with Applications, 166 (??):182–197, July 15, 2024. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122124002232.

Mao:2021:PEE

[MCW21] Wenting Mao, Yanping Chen, and Huasheng Wang. A-posteriori error estimations of the GJF-Petrov– Galerkin methods for fractional differential equations. Computers and Mathematics with Applications, 90 (??):159–170, May 15, 2021. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122121000961.

Miller:2022:VMM

[MCW22] Edward A. Miller, Xi Chen, and David M. Williams. Versatile mixed methods for non-isothermal incompressible flows. Computers and Mathematics with Applications, 125(??):150–175, November 1, 2022. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122122003704.

Ma:2025:NEN

[MCW25] Guodong Ma, Jinru Chen, and Feng Wang. A Nitsche’s extended nonconforming virtual element method for biharmonic PDEs involving interfaces. Computers and Mathematics with Applications, 192(??):134–154, August 15, 2025. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122125002147.

Ma:2020:MSA

[MCZL20] Ji Ma, Wen Chen, Chuanzeng Zhang, and Ji Lin. Meshless simulation of antiplane crack problems by the method of fundamental solutions using the crack Green’s function. Computers and Mathematics with Applications, 79(5):1543– 1560, March 1, 2020. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122119304663.

Migorski:2021:NCE

[MD21] Stanisław Migórski and Sylwia Dudek. A new class of elliptic quasi-variational-hemivariational inequalities for fluid flow with mixed boundary conditions. Computers and Mathematics with Applications, 100(??):51–61, October 15, 2021. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S089812212100314X.

Meddahi:2023:DMS

[Med23]  Salim Meddahi.  A DG

method for a stress formulation of the elasticity eigenproblem with strongly imposed symmetry. Computers and Mathematics with Applications, 135(??):19– 30, April 1, 2023. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122123000287.

Meng:2022:MMM

[Men22]  Zhaoliang Meng.  Modi-

fied Morley method by free formula for fourth order elliptic singular perturbation problems. Computers and Mathematics with Applications, 126(??):212– 225, November 15, 2022. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122122003960.

Moretto:2025:NMM

[MFF25] Daniele Moretto, Andrea Franceschini, and Massimiliano Ferronato. A novel mortar method integration using radial basis functions. Computers and Mathematics with Applications, 198(??): 38–58, November 15, 2025. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122125003359.

Ma:2023:RAF

[MFZX23] Li Ma, Hongfei Fu, Bingyin Zhang, and Shusen Xie. L1-robust analysis of a fourth-order block-centered finite difference method for two-dimensional variablecoefficient time-fractional reaction-diffusion equations. Computers and Mathematics with Applications, 148(??): 211–227, October 15, 2023. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122123003589.

Ma:2025:HOB

[MG25] Tingfu Ma and Yongbin Ge. A high-order blended compact difference (BCD) scheme for two-dimensional steady incompressible viscous flows. Computers and Mathematics with Applications, 194(??):288–315, September 15, 2025. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122125002718.

Mohanty:2022:HPI

[MGA22] R. K. Mohanty, Bishnu Pada Ghosh, and Urvashi Arora. High precision implicit method for 3D quasilinear hyperbolic equations on a dissimilar domain: Application to 3D telegraphic equation. Computers and Mathematics with Applications, 122(??): 93–116, September 15, 2022. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122122002930.

Majumder:2022:REP

[MGBN22] Sambit Majumder, Arnab Ghosh, Dipankar Narayan Basu, and Ganesh Natarajan. Re-examining the partially saturated-cells method for incompressible flows with stationary and moving bodies. Computers and Mathematics with Applications, 110 (??):19–39, March 15, 2022. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122122000499.

Ma:2023:SFW

[MGD23a] Jie Ma, Fuzheng Gao, and Ning Du. A stabilizer-free weak Galerkin finite element method with Alikhanov formula on nonuniform mesh for a linear reaction-subdiffusion problem. Computers and Mathematics with Applications, 148(??):180–187, October 15, 2023. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122123003474.

Marsic:2023:TON

[MGD23b] Nicolas Marsic, Christophe Geuzaine, and Herbert De Gersem. Transmission operators for the non-overlapping Schwarz method for solving Helmholtz problems in rectangular cavities. Computers and Mathematics with Applications, 138(??):37– 64, May 15, 2023. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122123000780.

Ma:2022:VPP

[MH22] Huimin Ma and Pengzhan Huang. A vector penaltyprojection approach for the time-dependent incompressible magnetohydrodynamics flows. Computers and Mathematics with Applications, 120 (??):28–44, August 15, 2022. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122122002644.

Mostajeran:2023:RBF

[MH23] F. Mostajeran and S. M. Hosseini. Radial basis function neural network (RBFNN) approximation of Cauchy inverse problems of the Laplace equation. Computers and Mathematics with Applications, 141(??):129– 144, July 1, 2023. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122123001694.

Mirzahashemi:2025:NSM

[MH25] Raziyeh Mirzahashemi and Mohammad Heydari. Numerical solutions of multidimensional nonlinear hyperbolic partial integrodifferential equations using a redefined cubic B-spline-based differential quadrature method. Computers and Mathematics with Applications, 198(??):214–238, November 15, 2025. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122125003505.

Marino:2020:EPK

[MHHR20] Enzo Marino, Seyed Farhad Hosseini, Ali Hashemian, and Alessandro Reali. Effects of parameterization and knot placement techniques on primal and mixed isogeometric collocation formulations of spatial sheardeformable beams with varying curvature and torsion. Computers and Mathematics with Applications, 80(11): 2563–2585, December 2020. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122120302558.

Meissner:2025:DLS

[MHKB25] Teddy Meissner, Edward Huynh, Paul Kuberry, and Pavel Bochev. A deep least-squares method for the Stokes equations. Computers and Mathematics with Applications, 196(??):1–12, October 15, 2025. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122125002858.

Maros:2020:FEM

[MI20] Gábor Maros and Ferenc Izsák. Finite element methods for fractional-order diffusion problems with optimal convergence order. Computers and Mathematics with Applications, 80(10):2105– 2114, November 15, 2020. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S089812212030362X.

Maeyama:2021:TCF

[MIOK21] Hirotaka Maeyama, Taro Imamura, Jun Osaka, and Naoki Kurimoto. Turbulent channel flow simulations using the lattice Boltzmann method with nearwall modeling on a nonbody-fitted Cartesian grid. Computers and Mathematics with Applications, 93 (??):20–31, July 1, 2021. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122121001292.

Mityushev:2022:EPTb

[Mit22] Vladimir Mityushev. Effective properties of twodimensional dispersed composites. Part II. Revision of self-consistent methods. Computers and Mathematics with Applications, 121(??): 74–84, September 1, 2022. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122122002863.

Mierke:2020:EAC

[MJR20] D. Mierke, C. F. Janßen, and T. Rung. An efficient algorithm for the calculation of sub-grid distances for higher-order LBM boundary conditions in a GPU simulation environment. Computers and Mathematics with Applications, 79(1):66–87, January 2020. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122118302207.

Majee:2020:DCN

[MJRM20] Sudeb Majee, Subit K.

Jain, Rajendra K. Ray, and Ananta K. Majee. On the development of a coupled nonlinear telegraph-diffusion model for image restoration. Computers and Mathematics with Applications, 80(7): 1745–1766, October 1, 2020. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122120303175.

Medeiros:2024:BCO

[MJVH24] R. Medeiros, E. Jané, F. Varas, and M. Higuera. Battery cell optimisation using time- and parameteradaptive reduced order models. Computers and Mathematics with Applications, 161 (??):137–154, May 1, 2024. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122124000890.

Mellouli:2020:FVA

[MJWD20] H. Mellouli, H. Jrad, M. Wali, and F. Dammak. Free vibration analysis of FG–CNTRC shell structures using the meshfree radial point interpolation method. Computers and Mathematics with Applications, 79(11):3160– 3178, June 1, 2020. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122120300316.

Matthaiou:2025:BPP

[MJZ25] Marina Matthaiou, Volker John, and Marwa Zainelabdeen. Bound-preserving physics-informed neural networks for steady-state convection– diffusion–reaction problems.

Computers and Mathematics with Applications, 199(??): 167–183, December 1, 2025. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122125003803.

Min:2020:DNB

[MK20] Youngho Min and Seungil Kim. Dirichlet-to-Neumann boundary conditions for multiple scattering in waveguides. Computers and Mathematics with Applications, 79(6):1661–1686, March 15, 2020. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122119304821.

Mittal:2020:HHA

[MKAM20] H. V. R. Mittal, Jiten C.

Kalita, and Qasem M. Al-Mdallal. A hybrid ψ–v HOC approach for surface tension driven flows in level set framework. Computers and Mathematics with Applications, 79(8):2350– 2375, April 15, 2020. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122119305358.

Mahajan:2024:FEA

[MKM24a] Sumit Mahajan, Arbaz Khan, and Manil T. Mohan. Finite element approximation for a delayed generalized Burgers–Huxley equation with weakly singular kernels: Part I wellposedness, regularity and conforming approximation. Computers and Mathematics with Applications, 174(??): 261–286, November 15, 2024. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122124004036.

Mellouli:2024:EBP

[MKM^{+}24b] H. Mellouli, S. Kalleli, H. Mallek, L. Ben Said, B. Ayadi, and F. Dammak. Electromechanical behavior of piezolaminated shell structures with imperfect functionally graded porous materials using an improved solid-shell element. Computers and Mathematics with Applications, 155(??):1–13, February 1, 2024. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122123005394.

Mishra:2023:SOB

[MKP23] Soumyarani Mishra, Morrakot Khebchareon, and Ambit K. Pany. Second order backward difference scheme combined with finite element method for a 2D Sobolev equation with Burgers’ type non-linearity. Computers and Mathematics with Applications, 141 (??):170–190, July 1, 2023. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122123001700.

Mousavi:2021:LES

[MKS^{+}21] Seyed Mahmood Mousavi, Reza Kamali, Freshteh Sotoudeh, Nader Karimi, and Danial Khojasteh. Large eddy simulation of pseudo shock structure in a convergent– long divergent duct. Computers and Mathematics with Applications, 81(??):823– 837, January 1, 2021. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122119305061.

Ming:2023:HTA

[MLHS23] Chunying Ming, Kexin Liu, Kelu Han, and Xinhui Si. Heat transfer analysis of Carreau fluid over a rotating disk with generalized thermal conductivity. Computers and Mathematics with Applications, 144(??):141– 149, August 15, 2023. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S089812212300233X.

McCullough:2020:ILN

[MLJ^{+}20] J. W. S. McCullough, C. R. Leonardi, B. D. Jones, S. M. Aminossadati, and J. R. Williams. Investigation of local and non-local lattice Boltzmann models for transient heat transfer between non-stationary, disparate media. Computers and Mathematics with Applications, 79 (1):174–194, January 2020. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122118300348.

Morgan:2025:HOI

[MLL25] Ian T. Morgan, Youzuo Lin, and Songting Luo. Highorder implicit Runge–Kutta Fourier pseudospectral methods for wave equations. Computers and Mathematics with Applications, 192(??):1–13, August 15, 2025. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122125002032.

Ming:2024:FLG

[MLLL24] Wanyuan Ming, Mengting Li, Yu Lu, and Meng Li. A fast linearized Galerkin finite element method for the nonlinear multi-term time fractional wave equation. Computers and Mathematics with Applications, 157 (??):27–48, March 1, 2024.

CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122123005813.

Meng:2022:NAF

[MLS22] Yahui Meng, Botong Li, and Xinhui Si. Numerical analysis of fractional viscoelastic fluid problem solved by finite difference scheme. Computers and Mathematics with Applications, 113 (??):225–242, May 1, 2022. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S089812212200116X.

Meyer:2021:PDM

[MM21] Marcela Molina Meyer and Frank Richard Prieto Medina. Polar differentiation matrices for the Laplace equation in the disk under nonhomogeneous Dirichlet, Neumann and Robin boundary conditions and the biharmonic equation under nonhomogeneous Dirichlet conditions. Computers and Mathematics with Applications, 89(??):1–19, May 1, 2021. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S089812212100047X.

McLean:2020:RTT

[MMAK20] William McLean, Kassem Mustapha, Raed Ali, and Omar M. Knio. Regularity theory for time-fractional advection–diffusion–reaction equations. Computers and Mathematics with Applications, 79(4):947–961, February 15, 2020. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122119304055.

See erratum [MMAK21].

McLean:2021:ERT

[MMAK21] William McLean, Kassem Mustapha, Raed Ali, and Omar M. Knio. Erratum to “Regularity theory for time-fractional advectiondiffusion-reaction equations” [Comput. Math. Appl. 79 (2020) 947–961]. Computers and Mathematics with Applications, 85(??):82–83, March 1, 2021. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122121000183.

See [MMAK20].

Malheiro:2021:PSA

[MMC21] M. T. Malheiro, Gaspar J. Machado, and Stéphane Clain. An a posteriori strategy for adaptive schemes in time for one-dimensional advection-diffusion transport equations. Computers and Mathematics with Applications, 103(??):65–81, December 1, 2021. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S089812212100376X.

MaraBeh:2025:PCV

[MMGS25] Raed Ali Mara’Beh, J. M.

Mantas, P. González, and Raymond J. Spiteri. Performance comparison of variable-stepsize IMEX SBDF methods on advection-diffusion-reaction models. Computers and Mathematics with Applications, 190(??):41– 56, July 15, 2025. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122125001476.

Maity:2021:PDF

[MMN21] Ruma Rani Maity, Apala Majumdar, and Neela Nataraj. Parameter dependent finite element analysis for ferronematics solutions. Computers and Mathematics with Applications, 103(??):127–155, December 1, 2021. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122121003813.

Matusik:2024:FTA

[MMN24] Radosław Matusik, Anna Michalak, and Andrzej Nowakowski. Fixed-time anti-synchronization for reaction– diffusion neural networks. Computers and Mathematics with Applications, 173(??): 19–32, November 1, 2024. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122124003365.

Mehar:2021:TBS

[MMP21] Kulmani Mehar, Pradeep Kumar Mishra, and Subrata Kumar Panda. Thermal buckling strength of smart nanotube-reinforced doubly curved hybrid composite panels. Computers and Mathematics with Applications, 90(??):13–24, May 15, 2021. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S089812212100078X.

Mascotto:2020:FBM

[MMPR20] Lorenzo Mascotto, Jens M. Melenk, Ilaria Perugia, and Alexander Rieder. FEM– BEM mortar coupling for the Helmholtz problem in three dimensions. Computers and Mathematics with Applications, 80(11):2351– 2378, December 2020. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122120301619.

Metcalfe:2021:QOT

[MN21] Stephen Metcalfe and Siva Nadarajah. A quasi-optimal test norm for a DPG discretization of the convectiondiffusion equation. Computers and Mathematics with Applications, 94(??):60– 75, July 15, 2021. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122121001607.

Mishra:2023:LPS

[MN23a] Sudheer Mishra and E. Natarajan. Local projection stabilization virtual element method for the convection– diffusion equation with nonlinear reaction term. Computers and Mathematics with Applications, 152(??):181– 198, December 15, 2023. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S089812212300473X.

Mohanty:2023:NPC

[MN23b] R. K. Mohanty and Niranjan.

Nine-point compact sixthorder approximation for twodimensional nonlinear elliptic partial differential equations: Application to biand tri-harmonic boundary value problems. Computers and Mathematics with Applications, 152(??):239– 249, December 15, 2023. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122123004765.

Mutombo:2024:GHK

[MN24] Franck Kalala Mutombo and Alice Nanyanzi. Generalised heat kernel invariants of a graph and application to object clustering. Computers and Mathematics with Applications, 167(??): 264–285, August 1, 2024. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122124002141.

Motamedi:2023:NTB

[MNB23] Ali Reza Motamedi, Nima Noormohammadi, and Bijan Boroomand. A novel Trefftzbased meshfree method for free vibration and buckling analysis of thin arbitrarily shaped laminated composite and isotropic plates. Computers and Mathematics with Applications, 145(??): 318–340, September 1, 2023. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122123002948.

Mishra:2023:PCN

[MP23] Bhaskar Mishra and Jörg Peters. Polyhedral control-net splines for analysis. Computers and Mathematics with Applications, 151(??):215– 221, December 1, 2023. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122123004261.

Mallesham:2025:VIP

[MPVA25] Harish Nagula Mallesham, Kamana Porwal, Jan Valdman, and Sanjib Kumar Acharya. Vectorized implementation of primal hybrid FEM in MATLAB. Computers and Mathematics with Applications, 180(??): 144–165, February 15, 2025. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122124005625.

Mityushev:2022:EPTa

[MR22] Vladimir Mityushev and Natalia Rylko. Effective properties of twodimensional dispersed composites. Part I. Schwarz’s alternating method. Computers and Mathematics with Applications, 111(??):50– 60, April 1, 2022. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122122000621.

Munoz:2024:TFD

[MR24] S. Muñoz and M.-L. Rapún. Towards flaw detection in welding joints via multifrequency topological derivative methods. Computers and Mathematics with Applications, 161(??):121– 136, May 1, 2024. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122124000841.

Madureira:2024:CNW

[MRAC24] Rodrigo L. R. Madureira, Mauro A. Rincon, Ricardo F. Apolaya, and Bruno A. Carmo. Control of a nonlinear wave equation with a dynamic boundary condition. Computers and Mathematics with Applications, 176(??): 140–149, December 15, 2024. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122124004437.

Martinez:2020:HOF

[MRN^{+}20] Abel Martínez, Luis Ramírez, Xesús Nogueira, Sofiane Khelladi, and Fermín Navarrina. A high-order finite volume method with improved isotherms reconstruction for the computation of multiphase flows using the Navier– Stokes-Korteweg equations. Computers and Mathematics with Applications, 79(3): 673–696, February 1, 2020. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122119303736.

Moayeri:2020:DBR

[MRP20] M. M. Moayeri, J. A. Rad, and K. Parand. Dynamical behavior of reaction– diffusion neural networks and their synchronization arising in modeling epileptic seizure: a numerical simulation study. Computers and Mathematics with Applications, 80(8): 1887–1927, October 15, 2020. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S089812212030328X.

Mendes:2021:FVS

[MRPK21] Julien Mendes, Antonio Russo, Sergio P. Perez, and Serafim Kalliadasis. A finite-volume scheme for gradient-flow equations with non-homogeneous diffusion. Computers and Mathematics with Applications, 89 (??):150–162, May 1, 2021. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122121000407.

Maljaars:2021:LPL

[MRS21] Jakob M. Maljaars, Chris N. Richardson, and Nathan Sime. LEoPart: a particle library for FEniCS. Computers and Mathematics with Applications, 81(??): 289–315, January 1, 2021. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S089812212030170X.

Ma:2024:NST

[MrZQ24] Rui Ma, Jian rong Zhu, and Cheng Qiu. A new scheme for two-way, nesting, quadrilateral grid in an estuarine model. Computers and Mathematics with Applications, 172(??):152–167, October 15, 2024. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122124003420.

Masoudi:2020:EPD

[MS20a] Mohsen Masoudi and Davod Khojasteh Salkuyeh. An extension of the positive-definite and skew-Hermitian splitting method for preconditioning of generalized saddle point problems. Computers and Mathematics with Applications, 79(8):2304– 2321, April 15, 2020. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122119305267.

Munar:2020:PEA

[MS20b] Mauricio Munar and Filánder A.

Sequeira. A posteriori error analysis of a mixed virtual element method for a nonlinear Brinkman model of porous media flow. Computers and Mathematics with Applications, 80(5):1240– 1259, September 1, 2020. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122120302546.

Mora:2022:VEM

[MS22] David Mora and Alberth Silgado. A C^{1} virtual element method for the stationary quasi-geostrophic equations of the ocean. Computers and Mathematics with Applications, 116(??):212– 228, June 15, 2022. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S089812212100208X.

Monsuur:2023:PFU

[MS23] Harald Monsuur and Rob Stevenson. A pollutionfree ultra-weak FOSLS discretization of the Helmholtz equation. Computers and Mathematics with Applications, 148(??):241–255, October 15, 2023. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122123003516.

Mucke:2023:MCG

[MSBO23] Nikolaj T. Mücke, Benjamin Sanderse, Sander M. Bohté, and Cornelis W. Oosterlee. Markov chain generative adversarial neural networks for solving Bayesian inverse problems in physics applications. Computers and Mathematics with Applications, 147(??):278–299, October 1, 2023. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122123003280.

Make:2022:SBS

[MSHB22] Michel Make, Thomas Spenke, Norbert Hosters, and Marek Behr. Spline-based spacetime finite element approach for fluid-structure interaction problems with a focus on fully enclosed domains. Computers and Mathematics with Applications, 114 (??):210–224, May 15, 2022. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122122001250.

Montardini:2025:LRS

[MST25] Monica Montardini, Giancarlo Sangalli, and Mattia Tani. A low-rank solver for conforming multipatch Isogeometric Analysis. Computers and Mathematics with Applications, 189 (??):86–108, July 1, 2025. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122125001385.

Mezzan:2025:MFE

[MT25] Youness Mezzan and Moulay Hicham Tber. A mixed finite element method for pricing American options and Greeks in the Heston model. Computers and Mathematics with Applications, 193(??):20–33, September 1, 2025. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122125002275.

Mojarrad:2023:NVS

[MVHÖ23] Fatemeh Nassajian Mojarrad, Maria Han Veiga, Jan S. Hesthaven, and Philipp Öffner. A new variable shape parameter strategy for RBF approximation using neural networks. Computers and Mathematics with Applications, 143(??):151– 168, August 1, 2023. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122123002031.

McKibben:2020:CSO

[MW20] Mark A. McKibben and Micah Webster. A class of second-order McKean– Vlasov stochastic evolution equations driven by fractional Brownian motion and Poisson jumps. Computers and Mathematics with Applications, 79(2):391–406, January 15, 2020. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122119303517.

Mohebujjaman:2022:EAP

[MWRA22] Muhammad Mohebujjaman, Hongwei Wang, Leo G. Rebholz, and Md. Abdullah Al Mahbub. An efficient algorithm for parameterized magnetohydrodynamic flow ensembles simulation. Computers and Mathematics with Applications, 112(??):167– 180, April 15, 2022. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122122000773.

Ma:2021:NOB

[MWY21] Yu Ma, Yahui Wang, and Junhe Yang. ntkFoam: an OpenFOAM based neutron transport kinetics solver for nuclear reactor simulation. Computers and Mathematics with Applications, 81(??): 512–531, January 1, 2021. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122119304651.

Miao:2023:IFC

[MWY23] Shuai Miao, Jiming Wu, and Yanzhong Yao. An interpolation-free cell-centered discretization of the heterogeneous and anisotropic diffusion problems on polygonal meshes. Computers and Mathematics with Applications, 130(??):105–118, January 15, 2023. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122122004928.

Ma:2022:HOI

[MWYY22] Chao Ma, Jie Wu, Haichuan Yu, and Liming Yang. A high-order implicit-explicit flux reconstruction lattice Boltzmann method for viscous incompressible flows. Computers and Mathematics with Applications, 105 (??):13–28, January 1, 2022. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122121004168.

Ma:2024:HPM

[MXLY24] Nan Ma, Qiuyan Xu, Zhiyong Liu, and Jiye Yang. High-precision meshfree methods for solving 3D radiation diffusion problem in irregular spaces. Computers and Mathematics with Applications, 169(??):163–176, September 1, 2024. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122124002815.

Ma:2023:SNM

[MZ23a] Xiaoqi Ma and Jin Zhang. Supercloseness of the NIPG method for a singularly perturbed convection diffusion problem on Shishkin mesh in 2D. Computers and Mathematics with Applications, 152(??):81–90, December 15, 2023. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122123004522.

Miazio:2023:PBL

[MZ23b] Ł. Miazio and G. Zboiński. A posteriori boundary layer detection and resolution in hpq-adaptive finite element methods for 3D-based hierarchical plate and shell models. Computers and Mathematics with Applications, 129 (??):50–68, January 1, 2023. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122122004734.

Mohammadian:2025:NSN

[MZCK25] Saeed Mohammadian, Zuduo Zheng, Shaoshuai Chu, and Alexander Kurganov. Numerical study of the nonconservative NET-RAT traffic flow model by pathconservative central-upwind schemes. Computers and Mathematics with Applications, 179(??):212–228, February 1, 2025. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122124005613.

Ma:2025:EPS

[MZGM25] Jiangming Ma, Xiang Zhang, Di Gao, and Jianjun Ma. An energy-preserving scheme for coupled fractional Gross– Pitaevskii equations based on energy discretization. Computers and Mathematics with Applications, 194 (??):335–361, September 15, 2025. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122125002706.

Miao:2024:AFS

[MZHJ24] Zhen Miao, Ren-Hao Zhang, Wei-Wei Han, and Yao-Lin Jiang. Analysis of a fractional-step parareal algorithm for the incompressible Navier–Stokes equations. Computers and Mathematics with Applications, 161 (??):78–89, May 1, 2024. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122124000804.

Miao:2021:GUT

[MZM21] Shu-Xin Miao, Jing Zhang, and Lingsheng Meng. A general Uzawa-type method for a class of 2 × 2 block structure linear system. Computers and Mathematics with Applications, 98(??):181– 200, September 15, 2021. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122121002704.

Ma:2023:CIM

[MZWD23] Fugui Ma, Lijing Zhao, Yejuan Wang, and Weihua Deng. The contour integral method for Feynman– Kac equation with two internal states. Computers and Mathematics with Applications, 151(??):80–100, December 1, 2023. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122123004224.

Mu:2025:HOG

[MZZ^{+}25] Junlei Mu, Congshan Zhuo, Qingdian Zhang, Sha Liu, and Chengwen Zhong. Highorder gas-kinetic scheme with TENO class reconstruction for the Euler and Navier–Stokes equations. Computers and Mathematics with Applications, 179(??): 126–147, February 1, 2025. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122124005418.

Nived:2022:AHO

[NAE22] M. R. Nived, Sai Saketha Chandra Athkuri, and Vinayak Eswaran. On the application of higher-order Backward Difference (BDF) methods for computing turbulent flows. Computers and Mathematics with Applications, 117 (??):299–311, July 1, 2022. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122122001948.

Naga:2024:IDA

[Nag24] Taha H. A. Naga. Improved dynamic analysis for plane elasticity problems with a boundary meshfree method and distributed sources. Computers and Mathematics with Applications, 163 (??):146–164, June 1, 2024. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122124001263.

Nadir-Alexandre:2020:HNC

[NAS20] Messai Nadir-Alexandre and Pernet Sébastien. hp nonconforming a priori error analysis of an interior penalty discontinuous Galerkin BEM for the Helmholtz equation. Computers and Mathematics with Applications, 80(12):2644– 2675, December 15, 2020. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122120304156.

Nastorg:2024:IGS

[NBF^{+}24] Matthieu Nastorg, Michele-Alessandro Bucci, Thibault Faney, Jean-Marc Gratien, Guillaume Charpiat, and Marc Schoenauer. An implicit GNN solver for Poisson-like problems. Computers and Mathematics with Applications, 176(??):270– 288, December 15, 2024. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122124004851.

Ngo:2020:MLA

[NC20] Long Cu Ngo and Hyoung Gwon Choi. A multilevel adaptive mesh refinement for an integrated finite element/level set formulation to simulate multiphase flows with surface tension. Computers and Mathematics with Applications, 79(4): 908–933, February 15, 2020. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122119304043.

Noorizadegan:2024:ETR

[NCCD24] A. Noorizadegan, C.-S.

Chen, R. Cavoretto, and A. De Rossi. Efficient truncated randomized SVD for mesh-free kernel methods. Computers and Mathematics with Applications, 164 (??):12–20, June 15, 2024. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122124001287.

Nino-Celis:2021:CPF

[NCRGVR21] Viviana Niño-Celis, Diego A.

Rueda-Gómez, and Élder J. Villamizar-Roa. Convergence and positivity of finite element methods for a haptotaxis model of tumoral invasion. Computers and Mathematics with Applications, 89(??):20–33, May 1, 2021. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122121000493.

Nie:2021:LDG

[ND21] Daxin Nie and Weihua Deng. Local discontinuous Galerkin method for the fractional diffusion equation with integral fractional Laplacian. Computers and Mathematics with Applications, 104(??): 44–49, December 15, 2021. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122121004041.

Nowak:2024:FTP

[NDB^{+}24] Marcin Nowak, Eduardo Divo, Tomasz Borkowski, Ewelina Marciniak, Marek Rojczyk, and Ryszard Białecki. Flow through a prosthetic mechanical aortic valve: Numerical model and experimental study. Computers and Mathematics with Applications, 175(??):184–201, December 1, 2024. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S089812212400419X.

Ngo:2022:HOL

[NDC22] Long Cu Ngo, Quang-Ngoc Dinh, and Hyoung Gwon Choi. High-order level set reinitialization for multiphase flow simulations based on unstructured grids. Computers and Mathematics with Applications, 120(??):60–77, August 15, 2022. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S089812212200267X.

Ndengna:2025:HOP

[Nde25] Arno Roland Ngatcha Ndengna.

A high order Path-Conservative Central-Upwind Arbitrary DERivative (PCCU-ADER) method for a generalized high order sediment transport model. Computers and Mathematics with Applications, 198(??):131–177, November 15, 2025. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122125003414.

Nasresfahani:2022:NSO

[NE22] F. Nasresfahani and M. R.

Eslahchi. Numerical solution of optimal control of atherosclerosis using direct and indirect methods with shooting/collocation approach. Computers and Mathematics with Applications, 126(??):60–76, November 15, 2022. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S089812212200373X.

Nolasco:2022:ENM

[NERB^{+}22] Irving E. Reyna Nolasco, Aimad Er-Raiy, Radouan Boukharfane, Anwar A. Aldhafeeri, Lisandro Dalcin, and Matteo Parsani. Eigenanalysis and non-modal analysis of collocated discontinuous Galerkin discretizations with the summationby-parts property. Computers and Mathematics with Applications, 124(??):196– 217, October 15, 2022. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S089812212200325X.

Namaki:2023:UPI

[NES23] Neda Namaki, M. R. Eslahchi, and Rezvan Salehi. The use of physics-informed neural network approach to image restoration via nonlinear PDE tools. Computers and Mathematics with Applications, 152(??):355– 363, December 15, 2023. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122123004431.

Nielsen:2025:HON

[NF25] Morten E. Nielsen and Bengt Fornberg. High-order numerical method for solving elliptic partial differential equations on unfitted node sets. Computers and Mathematics with Applications, 193(??): 103–116, September 1, 2025.

CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122125002287.

Nogueira:2024:MLB

[NFFR^{+}24] Xesús Nogueira, Javier Fernández-Fidalgo, Lucía Ramos, Iván Couceiro, and Luis Ramírez. Machine learning-based WENO5 scheme. Computers and Mathematics with Applications, 168 (??):84–99, August 15, 2024. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122124002505.

Nie:2024:MFV

[NFS24] Cunyun Nie, Jianglin Fang, and Shi Shu. A monotone finite volume element scheme for diffusion equations on arbitrary polygonal grids. Computers and Mathematics with Applications, 153(??): 225–236, January 1, 2024. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122123005278.

Ngondiep:2025:EHO

[Ngo25] Eric Ngondiep. An efficient high-order weak Galerkin finite element approach for Sobolev equation with variable matrix coefficients. Computers and Mathematics with Applications, 180(??): 279–298, February 15, 2025. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122125000124.

Nguyen:2020:NAD

[Ngu20]  Loc Hoang Nguyen.  A

new algorithm to determine the creation or depletion term of parabolic equations from boundary measurements. Computers and Mathematics with Applications, 80(10):2135– 2149, November 15, 2020. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122120303667.

Nataj:2024:AAN

[NH24] Sarah Nataj and Yunhui He. Anderson acceleration for nonlinear PDEs discretized by space–time spectral methods. Computers and Mathematics with Applications, 167(??):199– 206, August 1, 2024. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122124002116.

Niu:2022:ASA

[NHCJ22] Qiang Niu, Size Hou, Yang Cao, and Yanfei Jing. Algebraic spectral analysis of the DSSR preconditioner. Computers and Mathematics with Applications, 125(??): 80–89, November 1, 2022. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122122003686.

Niemi:2023:APP

[Nie23] Antti H. Niemi. Approximation properties of primal discontinuous Petrov– Galerkin method on general quadrilateral meshes. Computers and Mathematics with Applications, 151(??): 300–303, December 1, 2023. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122123004005.

Nguyen:2020:LSB

[NIT^{+}20] Truong Nguyen, Hiroshi Isakari, Toru Takahashi, Kentaro Yaji, Masato Yoshino, and Toshiro Matsumoto. Level-set based topology optimization of transient flow using lattice Boltzmann method considering an oscillating flow condition. Computers and Mathematics with Applications, 80(1):82–108, July 1, 2020. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122120301000.

Nordbotten:2025:TPS

[NK25] Jan Martin Nordbotten and Eirik Keilegavlen. Two-point stress approximation: a simple and robust finite volume method for linearized (poro- )elasticity and Stokes flow. Computers and Mathematics with Applications, 197(??): 259–294, November 1, 2025. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122125003244.

Nouri:2025:OTN

[NKPA^{+}25] Madyen Nouri, Mohammad-Javad Kazemzadeh-Parsi, Amine Ammar, Francisco Chinesta, Julien Artozoul, and Aude Caillaud. Optimal transport in non-convex geometries and its application in shrinkage porosity prediction. Computers and Mathematics with Applications, 193(??):222–240, September 1, 2025. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122125002573.

Najafian:2023:NSD

[NM23a] M. Najafian and S. Mortazavi. A numerical study of drop evaporation at high density ratios using Front-Tracking method. Computers and Mathematics with Applications, 134(??):1–15, March 15, 2023. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122122005247.

Ni:2023:OCS

[NM23b] Dongdong Ni and Wanbiao Ma. Optimal control of a size-structured unicellular algae model in harvesting and sewage treatment. Computers and Mathematics with Applications, 145(??): 240–259, September 1, 2023. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122123002699.

Ma:2024:MDS

[nMjWlS24] Ya nan Ma, Tian jun Wang, and You lin Shang. A multidomain spectral-Galerkin method for the Neumann problem on quadrilaterals. Computers and Mathematics with Applications, 156(??): 180–194, February 15, 2024. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122123005941.

Niu:2023:LTC

[NML23] Cuixia Niu, Heping Ma, and Dong Liang. Legendretau Chebyshev collocation spectral element method for Maxwell’s equations with material interfaces of two dimensional transverse magnetic mode. Computers and Mathematics with Applications, 147(??):222–238, October 1, 2023. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122123003176.

Nguyen:2025:ERA

[NNH^{+}25] Trung-Hieu Nguyen, Truong-Thang Nguyen, Duc-Minh Hoang, Viet-Hung Dang, and Xuan-Dat Pham. Efficient reliability analysis method for non-linear truss structures using machine learningbased uncertainty quantification. Computers and Mathematics with Applications, 182 (??):66–83, March 15, 2025. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122125000148.

Nguyen:2022:FEL

[NNNT22] Son H. Nguyen, Tan N. Nguyen, and Trung Nguyen-Thoi. A finite element level-set method for stressbased topology optimization of plate structures. Computers and Mathematics with Applications, 115 (??):26–40, June 1, 2022. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S089812212200150X.

Nguyen:2022:CBC

[NNT22] Dinh-Liem Nguyen, Loc H.

Nguyen, and Trung Truong. The Carleman-based contraction principle to reconstruct the potential of nonlinear hyperbolic equations. Computers and Mathematics with Applications, 128(??): 239–248, December 15, 2022. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122122004527.

Nguyen:2025:MMT

[NNTCB25] Minh Ngoc Nguyen, Nhon Nguyen-Thanh, Shunhua Chen, and Tinh Quoc Bui. Multi-material topology optimization of thermoelastic structures by an ordered SIMP-based phase field model. Computers and Mathematics with Applications, 186(??):84–100, May 15, 2025. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122125001002.

Nguyen:2024:AFV

[NP24] Son H. Nguyen and Duc-Huynh Phan. Analysis of free vibration in thinwalled plates using an enhanced polygonal plate element with selective interpolation approach. Computers and Mathematics with Applications, 156(??):42–55, February 15, 2024. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122123005643.

Novak:2022:SFC

[NR22] Andrej Novak and Nora Reinić. Shock filter as the classifier for image inpainting problem using the Cahn– Hilliard equation. Computers and Mathematics with Applications, 123(??):105– 114, October 1, 2022. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122122003121.

Naga:2025:CSI

[NR25a] Taha H. A. Naga and Youssef F. Rashed. Computation of singular integrals in variational boundary element method. Computers and Mathematics with Applications, 188(??):244– 263, June 15, 2025. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122125001300.

Normington:2025:DCF

[NR25b] Hywel Normington and Michele Ruggeri. A decoupled, convergent and fully linear algorithm for the Landau–Lifshitz–Gilbert equation with magnetoelastic effects. Computers and Mathematics with Applications, 187(??):1–29, June 1, 2025. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122125001038.

Nanda:2022:ITP

[NRV22a] P. Nanda, G. M. M. Reddy, and M. Vynnycky. Inverse two-phase nonlinear Stefan and Cauchy–Stefan problems: a phase-wise approach. Computers and Mathematics with Applications, 123(??): 216–226, October 1, 2022. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122122003297.

Nasser:2022:CCH

[NRV22b] Mohamed M. S. Nasser, Oona Rainio, and Matti Vuorinen. Condenser capacity and hyperbolic perimeter. Computers and Mathematics with Applications, 105 (??):54–74, January 1, 2022. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122121004132.

Nicaise:2020:SAP

[NS20] Serge Nicaise and Claire Scheid. Stability and asymptotic properties of a linearized hydrodynamic medium model for dispersive media in nanophotonics. Computers and Mathematics with Applications, 79(12): 3462–3494, June 15, 2020. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122120300675.

Nhangumbe:2024:NSO

[NS24] Clarinda Nhangumbe and Ercília Sousa. Numerical solutions of an option pricing rainfall weather derivatives model. Computers and Mathematics with Applications, 153(??):43–55, January 1, 2024. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122123005023.

Nandi:2021:ASI

[NSMM21] Ashish Kumar Nandi, Vaibhav Shekhar, Nachiketa Mishra, and Debasisha Mishra. Alternating stationary iterative methods based on double splittings. Computers and Mathematics with Applications, 89 (??):87–98, May 1, 2021. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122121000638.

Ngan:2025:HRO

[NSMR25] Valentin Nkana Ngan, Giovanni Stabile, Andrea Mola, and Gianluigi Rozza. A hybrid reduced-order model for segregated fluid-structure interaction solvers in an ALE approach at high Reynolds number. Computers and Mathematics with Applications, 180(??):299–321, February 15, 2025. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122125000045.

Nuca:2024:SSC

[NSRI24] Roberto Nuca, Erlend Storvik, Florin A. Radu, and Matteo Icardi. Splitting schemes for coupled differential equations: Block Schur-based approaches & partial Jacobi approximation. Computers and Mathematics with Applications, 161(??):190– 201, May 1, 2024. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122124000865.

Nakov:2021:CEF

[NT21] S. Nakov and I. Toulopoulos.

Convergence estimates of finite elements for a class of quasilinear elliptic problems. Computers and Mathematics with Applications, 104(??): 87–112, December 15, 2021. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122121003977.

Nguyen:2025:MCS

[NTP25] Van-Chinh Nguyen, Van-Ke Tran, and Van-Vinh Pham. Modified couple stress-based finite element model for analysis of partially supported 2D-FG sandwich microbeams with variable material length scale parameters. Computers and Mathematics with Applications, 198(??):106– 130, November 15, 2025. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122125003566.

Nannen:2024:KES

[NW24a] Lothar Nannen and Markus Wess. A Krylov eigenvalue solver based on filtered time domain solutions. Computers and Mathematics with Applications, 176(??):179– 188, December 15, 2024. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122124004498.

Nigam:2024:CFE

[NW24b] Nilima Nigam and David M.

Williams. Conforming finite element function spaces in four dimensions, part I: Foundational principles and the tesseract. Computers and Mathematics with Applications, 166(??):198– 223, July 15, 2024. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122124002074.

Nie:2022:MFV

[NY22] Cunyun Nie and Haiyuan Yu.

A monotone finite volume element scheme for diffusion equations on triangular grids. Computers and Mathematics with Applications, 105 (??):1–12, January 1, 2022. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122121004089.

Nan:2024:EBT

[NZ24] Caixia Nan and Qian Zhang.

Elastic bending total variation model for image inpainting with operator splitting method. Computers and Mathematics with Applications, 176(??):150–164, December 15, 2024. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122124004322.

Obalalu:2022:CSM

[OAA^{+}22] Adebowale Martins Obalalu, Adebayo Olusegun Ajala, Akintayo Oladimeji Akindele, Olayinka Akeem Oladapo, Olajide Olatunbosun Akintayo, and Oluwatosin Muinat Jimoh. Computational study of magneto-convective non-Newtonian nanofluid slip flow over a stretching/ shrinking sheet embedded in a porous medium. Computers and Mathematics with Applications, 119(??):319– 326, August 1, 2022. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122122002206.

Oliari:2024:CNS

[OAD24] Victor B. Oliari, Ricardo J. Hancco Ancori, and Philippe R. B. Devloo. A comparative numerical study of finite element methods resulting in mass conservation for Poisson’s problem: Primal hybrid, mixed and their hybridized formulations. Computers and Mathematics with Applications, 170(??):213–227, September 15, 2024. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122124002955.

Oulmelk:2024:ADM

[OAH^{+}24] A. Oulmelk, L. Afraites, A. Hadri, M. A. Zaky, and A. S. Hendy. Alternating direction multiplier method to estimate an unknown source term in the timefractional diffusion equation. Computers and Mathematics with Applications, 156(??): 195–206, February 15, 2024. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122123005837.

Ortega:2021:ETI

[OCCM21] E. Ortega, E. Castillo, R. C.

Cabrales, and N. O. Moraga. Effect of time integration scheme in the numerical approximation of thermally coupled problems: From first to third order. Computers and Mathematics with Applications, 99(??): 345–360, October 1, 2021. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122121003096.

Oleksy:2022:IMF

[OCRN22] Marta Oleksy, Witold Cecot, Waldemar Rachowicz, and Michał Nessel. An improved multiscale FEM for the free vibrations of heterogeneous solids. Computers and Mathematics with Applications, 110(??):110– 122, March 15, 2022. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122122000529.

Obbadi:2024:ISA

[OEASY24] Anouar Obbadi, Mofdi El-Amrani, Mohammed Seaid, and Driss Yakoubi. An improved splitting algorithm for unsteady generalized Newtonian fluid flow problems with natural boundary conditions. Computers and Mathematics with Applications, 167(??):92–109, August 1, 2024. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122124002165.

Otte:2020:SAC

[OF20] Philipp Otte and Martin Frank. A structured approach to the construction of stable linear Lattice Boltzmann collision operator. Computers and Mathematics with Applications, 79(5): 1447–1460, March 1, 2020. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122119304572.

Okubo:2021:TOA

[OKSS21] Carlos M. Okubo, César Y.

Kiyono, Luís F. N. Sá, and Emílio C. N. Silva. Topology optimization applied to 3D rotor flow path design based on the continuous adjoint approach. Computers and Mathematics with Applications, 96(??):16–30, August 15, 2021. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122121001863.

Omella:2024:ADL

[OP24] Ángel J. Omella and David Pardo. r-adaptive deep learning method for solving partial differential equations. Computers and Mathematics with Applications, 153 (??):33–42, January 1, 2024. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122123004959.

Ovall:2022:QID

[OR22] Jeffrey S. Ovall and Samuel E.

Reynolds. Quadrature for implicitly-defined finite element functions on curvilinear polygons. Computers and Mathematics with Applications, 107(??):1–16, February 1, 2022. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122121004296.

Orlando:2023:FMA

[Orl23] Giuseppe Orlando. A filtering monotonization approach for DG discretizations of hyperbolic problems. Computers and Mathematics with Applications, 129(??): 113–125, January 1, 2023. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122122004801.

Oruc:2020:TMM

[Oru20] Ömer Oruç. Two meshless methods based on local radial basis function and barycentric rational interpolation for solving 2D viscoelastic wave equation. Computers and Mathematics with Applications, 79(12): 3272–3288, June 15, 2020. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S089812212030050X.

Oruc:2022:ACM

[Oru22] Ömer Oruç. An accurate computational method for two-dimensional (2D) fractional Rayleigh–Stokes problem for a heated generalized second grade fluid via linear barycentric interpolation method. Computers and Mathematics with Applications, 118(??):120– 131, July 15, 2022. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122122001997.

Oruc:2025:UPA

[Oru25]  Ömer Oruç.  The use

of polynomial-augmented RBF collocation method with ghost points for plane elastostatic equations of anisotropic functionally graded materials. Computers and Mathematics with Applications, 184(??):116–133, April 15, 2025. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122125000720.

Opschoor:2024:DRN

[OS24] Joost A. A. Opschoor and Christoph Schwab. Deep ReLU networks and highorder finite element methods II: Chebyšev emulation. Computers and Mathematics with Applications, 169(??): 142–162, September 1, 2024. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122124002761.

Obalalu:2023:AHW

[OSO^{+}23] Adebowale Martins Obalalu, Sulyman Olakunle Salawu, Olalekan Adebayo Olayemi, Olusegun Adebayo Ajala, and Kazeem Issa. Analysis of hydromagnetic Williamson fluid flow over an inclined stretching sheet with Hall current using Galerkin Weighted Residual Method. Computers and Mathematics with Applications, 146(??): 22–32, September 15, 2023. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122123002730.

Oyarzua:2022:AUM

[OSZ22] Ricardo Oyarzúa, Manuel Solano, and Paulo Zúñiga. Analysis of an unfitted mixed finite element method for a class of quasi-Newtonian Stokes flow. Computers and Mathematics with Applications, 114(??):225– 243, May 15, 2022. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122122001365.

Oliver:2024:DGM

[OTC24] Dylan J. Oliver, Ian W. Turner, and Elliot J. Carr. Dual-grid mapping method for the advection-diffusion-reaction equation in a heterogeneous medium. Computers and Mathematics with Applications, 154(??):78–90, January 15, 2024. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122123005102.

Olshanskii:2021:FEM

[OXY21] Maxim Olshanskii, Xianmin Xu, and Vladimir Yushutin. A finite element method for Allen–Cahn equation on deforming surface. Computers and Mathematics with Applications, 90(??):148– 158, May 15, 2021. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122121000936.

Pacheco:2024:TSL

[Pac24] Douglas R. Q. Pacheco. On the temporal stability of least-squares methods for linear hyperbolic problems. Computers and Mathematics with Applications, 168 (??):33–38, August 15, 2024. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122124002426.

Park:2022:ASM

[Par22a]  Jongho Park.  Additive

Schwarz methods for convex optimization with backtracking. Computers and Mathematics with Applications, 113 (??):332–344, May 1, 2022.

CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122122001237.

Park:2022:INI

[Par22b] Won-Kwang Park. Investigation of a non-iterative technique based on topological derivatives for fast localization of small conductivity inclusions. Computers and Mathematics with Applications, 120(??):45–59, August 15, 2022. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122122002681.

Park:2022:SIO

[Par22c] Won-Kwang Park. Shape identification of open soundhard arcs without priori information in limited-view inverse scattering problem. Computers and Mathematics with Applications, 128(??): 55–68, December 15, 2022. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122122004321.

Park:2024:AMT

[Par24a] Won-Kwang Park. Application of MUSIC-type imaging for anomaly detection without background information. Computers and Mathematics with Applications, 172(??): 202–215, October 15, 2024. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122124003699.

Park:2024:UNI

[Par24b] Won-Kwang Park. Unveiling novel insights into Kirchhoff migration for a fast and effective object detection from experimental Fresnel dataset. Computers and Mathematics with Applications, 175(??): 345–355, December 1, 2024. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122124004620.

Peterson:2024:OBP

[PBR24] Kara Peterson, Pavel Bochev, and Denis Ridzal. Optimizationbased, property-preserving algorithm for passive tracer transport. Computers and Mathematics with Applications, 159(??):267–286, April 1, 2024. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122124000762.

Pan:2021:EMP

[PCC21] Xiaomin Pan, Soomin Chun, and Jung-Il Choi. Efficient monolithic projection-based method for chemotaxisdriven bioconvection problems. Computers and Mathematics with Applications, 84(??):166–184, February 15, 2021. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122121000079.

Pham:2025:NCF

[PCL25] Quoc Hoa Pham and Thanh Cuong-Le. A novel Chebyshev Finite Element approach for bending and free vibration analysis of tridirectional functionally graded porous nanoshell. Computers and Mathematics with Applications, 198(??):350– 376, November 15, 2025. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122125004286.

Peng:2025:SIL

[PCLZ25] Shuang Peng, Songze Chen, Hong Liang, and Chuang Zhang. Semi-implicit Lax– Wendroff kinetic scheme for multi-scale phonon transport. Computers and Mathematics with Applications, 187(??):72–84, June 1, 2025. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122125001130.

Park:2020:HCN

[PCM20] Moongyu Park, John H.

Cushman, and Michael J. Mueterthies. Homogenization of carbon-nanotube brush electrodes. Computers and Mathematics with Applications, 79(4):1177–1187, February 15, 2020. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122119304250.

Petrides:2021:AMS

[PD21] Socratis Petrides and Leszek Demkowicz. An adaptive multigrid solver for DPG methods with applications in linear acoustics and electromagnetics. Computers and Mathematics with Applications, 87 (??):12–26, April 1, 2021. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122121000274.

Pradhan:2025:FOT

[PD25] Gouranga Pradhan and Bhupen Deka. Fourth order time-stepping VEM for the 2D acoustic wave equations. Computers and Mathematics with Applications, 188 (??):115–136, June 15, 2025. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122125001257.

Pezzano:2022:GCA

[PDB22] Stefano Pezzano, Régis Duvigneau, and Mickaël Binois. Geometrically consistent aerodynamic optimization using an isogeometric Discontinuous Galerkin method. Computers and Mathematics with Applications, 128(??):368–381, December 15, 2022. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122122004606.

Pellegrino:2023:FCS

[Pel23] Sabrina F. Pellegrino. A filtered Chebyshev spectral method for conservation laws on network. Computers and Mathematics with Applications, 151(??):418–433, December 1, 2023. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122123004649.

Poveda:2023:CCG

[PFCZ23] Leonardo A. Poveda, Shubin Fu, Eric T. Chung, and Lina Zhao. Convergence of the CEM-GMsFEM for compressible flow in highly heterogeneous media. Computers and Mathematics with Applications, 151(??):153– 163, December 1, 2023. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122123004182.

Pal:2021:PEA

[PG21] Birupaksha Pal and Sashikumaar Ganesan. An a priori error analysis for a projection based variational multiscale finite element method for Oseen problems in a timedependent domain. Computers and Mathematics with Applications, 82(??):130– 147, January 15, 2021. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122120304375.

Pal:2025:CAN

[PG25] Arnab Pal and Thirupathi Gudi. Convergence of an AFEM for a non-local nonlinear problem of Kirchhoff type. Computers and Mathematics with Applications, 189 (??):176–193, July 1, 2025. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122125001737.

Pasquali:2020:NWT

[PGK20] Andrea Pasquali, Martin Geier, and Manfred Krafczyk. Near-wall treatment for the simulation of turbulent flow by the cumulant lattice Boltzmann method. Computers and Mathematics with Applications, 79(1):195–212, January 2020. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122117307484.

Pepona:2023:ECL

[PGR23] Marianna Pepona, John Gounley, and Amanda Randles. Effect of constitutive law on the erythrocyte membrane response to large strains. Computers and Mathematics with Applications, 132(??):145–160, February 15, 2023. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122122005107.

Peng:2023:QAM

[PH23] Qiyao Peng and Sander C. Hille. Quality of approximating a mass-emitting object by a point source in a diffusion model. Computers and Mathematics with Applications, 151(??):491–507, December 1, 2023. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S089812212300487X.

Pang:2025:STP

[PH25] Chen Hui Pang and Viet Ha Hoang. Sparse tensor product finite elements for two scale elliptic and parabolic equations with discontinuous coefficients. Computers and Mathematics with Applications, 179(??):17–40, February 1, 2025. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122124005157.

Pei:2021:LSO

[PHL21] Shuaichao Pei, Yanren Hou, and Qi Li. A linearly secondorder, unconditionally energy stable scheme and its error estimates for the modified phase field crystal equation. Computers and Mathematics with Applications, 103(??): 104–126, December 1, 2021. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122121003837.

Park:2023:IWG

[PK23] Hyeokjoo Park and Do Y.

Kwak. An immersed weak Galerkin method for elliptic interface problems on polygonal meshes. Computers and Mathematics with Applications, 147(??):185– 201, October 1, 2023. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S089812212300319X.

Priyanka:2025:CAF

[PK25] Priyanka and Sunil Kumar. Convergence analysis of a fast ADI compact finite difference method for two-dimensional semi-linear time-fractional reaction-diffusion equations with weak initial singularity. Computers and Mathematics with Applications, 199(??):1–21, December 1, 2025. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122125003621.

Pany:2021:NNE

[PKP21] Ambit Kumar Pany, Morrakot Khebchareon, and Amiya K. Pani. Negative norm estimates and superconvergence results in Galerkin method for strongly nonlinear parabolic problems. Computers and Mathematics with Applications, 99 (??):26–36, October 1, 2021. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122121002777.

Pasquale:2024:MPP

[PKPD^{+}24] Angelo Pasquale, Mohammad-Javad Kazemzadeh-Parsi, Daniele Di Lorenzo, Victor Champaney, Amine Ammar, and Francisco Chinesta. Modular parametric PGD enabling online solution of partial differential equations. Computers and Mathematics with Applications, 176(??): 244–256, December 15, 2024. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122124004863.

Peng:2024:CPI

[PL24] Keyi Peng and Jian Li. The coupled physical-informed neural networks for the two phase magnetohydrodynamic flows. Computers and Mathematics with Applications, 166(??):118– 128, July 15, 2024. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122124001767.

Peng:2025:ESB

[PL25] Gang Peng and Yuan Li.

An energy stable boundpreserving finite volume scheme for the Allen–Cahn equation based on operator splitting method. Computers and Mathematics with Applications, 178(??):47–60, January 15, 2025. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122124005108.

Plociniczak:2022:EGS

[Pło22]  Łukasz Płociniczak.  Er-

ror of the Galerkin scheme for a semilinear subdiffusion equation with timedependent coefficients and nonsmooth data. Computers and Mathematics with Applications, 127(??):181–191, December 1, 2022. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S089812212200414X.

Pellerin:2021:IBL

[PLR21] Nicolas Pellerin, Sébastien Leclaire, and Marcelo Reggio. An interpolation-based lattice Boltzmann method for non-conforming orthogonal meshes. Computers and Mathematics with Applications, 100(??):152–166, October 15, 2021. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122121003229.

Pleshkevich:2021:MLP

[PLVL21] Alexander Pleshkevich, Vadim Lisitsa, Dmitry Vishnevsky, and Vadim Levchenko. A multi-level parallel algorithm for seismic imaging based on one-way wave equation migration. Computers and Mathematics with Applications, 97(??):344–354, September 1, 2021. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122121002455.

Pei:2025:SAE

[PLZZ25] Lifang Pei, Ruixue Li, Jiwei Zhang, and Yanmin Zhao. Superconvergnce analysis of an energy-stable implicit scheme with variable time steps and anisotropic spatial nonconforming finite elements for the nonlinear Sobolev equations. Computers and Mathematics with Applications, 186 (??):37–52, May 15, 2025. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122125000951.

Pandit:2023:TMF

[PM23] Swapan K. Pandit and Rupchand Malo. Thermogravitational magnetonanofluid flow in a square cavity with isotropic heat flow field and viscous dissipation: a compact approach. Computers and Mathematics with Applications, 143(??):417– 434, August 1, 2023. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122123002249.

Pang:2020:IRA

[PMLC20] Zhi-Feng Pang, Ge Meng, Hui Li, and Ke Chen. Image restoration via the adaptive T V^{p} regularization. Computers and Mathematics with Applications, 80(5):569–587, September 1, 2020. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122120301838.

Pallikarakis:2024:AML

[PN24] Nikolaos Pallikarakis and Andreas Ntargaras. Application of machine learning regression models to inverse eigenvalue problems. Computers and Mathematics with Applications, 154(??): 162–174, January 15, 2024. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S089812212300545X.

Pham:2024:SVA

[PND^{+}24] Quoc Hoa Pham, Thien-Anh Nguyen, Ngoc-Tu Do, Van Ke Tran, and Minh-Nhan Nguyen. Static and vibration analyses of functionally graded porous shell structures by using an averaged edge/node-based smoothed MITC3 element. Computers and Mathematics with Applications, 153 (??):56–70, January 1, 2024. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S089812212300490X.

Prusak:2023:OBD

[PNT^{+}23] Ivan Prusak, Monica Nonino, Davide Torlo, Francesco Ballarin, and Gianluigi Rozza. An optimisationbased domain-decomposition reduced order model for the incompressible Navier– Stokes equations. Computers and Mathematics with Applications, 151(??):172–189, December 1, 2023. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122123004248.

Ponce:2020:TDF

[Pon20] Rodrigo Ponce. Time discretization of fractional subdiffusion equations via fractional resolvent operators. Computers and Mathematics with Applications, 80 (4):69–92, August 15, 2020. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122120301711.

Phan:2024:GEE

[PP24] J. D. Phan and A.-V. Phan.

A generalized energy eigenvalue problem for effectively solving the confined electron states in quantum semiconductor structures via boundary integral analysis. Computers and Mathematics with Applications, 170(??):228– 236, September 15, 2024. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122124003110.

Pang:2021:CBA

[PQSM21] Hong-Kui Pang, Hai-Hua Qin, Hai-Wei Sun, and Ting-Ting Ma. Circulant-based approximate inverse preconditioners for a class of fractional diffusion equations. Computers and Mathematics with Applications, 85 (??):18–29, March 1, 2021. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122121000171.

Picarelli:2020:DBP

[PR20] Athena Picarelli and Christoph Reisinger. Duality-based a posteriori error estimates for some approximation schemes for optimal investment problems. Computers and Mathematics with Applications, 79 (7):2099–2118, April 2020. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122119305681.

Pasquale:2023:TMD

[PRN^{+}23] Angelo Pasquale, Sebastian Rodriguez, Khanh Nguyen, Amine Ammar, and Francisco Chinesta. A time multiscale decomposition in cyclic elasto-plasticity. Computers and Mathematics with Applications, 149(??):75–83, November 1, 2023. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122123003553.

Piccialli:2024:P

[PRS24] Francesco Piccialli, Gianluigi Rozza, and Giovanni Stabile. Preface. Computers and Mathematics with Applications, 155(??): 14, February 1, 2024. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122123005497.

Palii:2020:CDP

[PS20] Olena Palii and Matthias Schlottbom. On a convergent DSA preconditioned source iteration for a DGFEM method for radiative transfer. Computers and Mathematics with Applications, 79(12):3366–3377, June 15, 2020. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0898122120300638.

Polz:2021:STD

[PS21] D. Pölz and M. Schanz. On the space-time discretization of variational retarded potential boundary integral equations. Computers and Mathematics with Applications, 99(??):195–210, October 1, 2021. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122121002881.

Pacheco:2022:IHO

[PS22] Douglas R. Q. Pacheco and Olaf Steinbach. On the initial higher-order pressure convergence in equal-order finite element discretizations of the Stokes system. Computers and Mathematics with Applications, 109(??):140– 145, March 1, 2022. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122122000293.

Pendar:2025:LPG

[PS25a] Masoud Pendar and Kamal Shanazari. Local Petrov– Galerkin meshfree method based on radial point interpolation for the numerical solution of 2D linear hyperbolic equations with variable coefficients. Computers and Mathematics with Applications, 190(??):16– 40, July 15, 2025. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122125001269.

Pouso:2025:ADS

[PS25b] Óscar López Pouso and Javier Segura. Analysis of difference schemes for the Fokker–Planck angular diffusion operator. Computers and Mathematics with Applications, 181(??):84– 110, March 1, 2025. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122125000057.

Paszynski:2022:SCD

[PSDP22] Maciej Paszyński, Leszek Siwik, Witold Dzwinel, and Keshav Pingali. Supermodeling, a convergent data assimilation meta-procedure used in simulation of tumor progression. Computers and Mathematics with Applications, 113(??):214– 224, May 1, 2022. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122122001158.

Putri:2024:DGA

[PSI^{+}24] Endah R. M. Putri, Muhammad L. Shahab, Mohammad Iqbal, Imam Mukhlash, Amirul Hakam, Lutfi Mardianto, and Hadi Susanto. A deep-genetic algorithm (deep-GA) approach for high-dimensional nonlinear parabolic partial differential equations. Computers and Mathematics with Applications, 154(??):120–127, January 15, 2024. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122123005126.

Perugia:2020:TPT

[PSSW20] Ilaria Perugia, Joachim Schöberl, Paul Stocker, and Christoph Wintersteiger. Tent pitching and Trefftz– DG method for the acoustic wave equation. Computers and Mathematics with Applications, 79(10):2987– 3000, May 15, 2020. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122120300134.

Plociniczak:2024:FDG

[PT24] Łukasz Płociniczak and Kacper Taźbierski. Fully discrete Galerkin scheme for a semilinear subdiffusion equation with nonsmooth data and time-dependent coefficient. Computers and Mathematics with Applications, 165 (??):217–223, July 1, 2024. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122124001974.

Ptaszny:2021:FMB

[Pta21] Jacek Ptaszny. A fast multipole BEM with higher-order elements for 3-D composite materials. Computers and Mathematics with Applications, 82(??):148–160, January 15, 2021. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122120304363.

Param:2023:TDD

[PTM^{+}23] H. Khalili Param, H. Tofighian, M. Mokhlesabadi, M. Nabaei, and A. Farnoud. Targeted drug delivery during radioembolization in a comprehensive hepatic artery system: a computational study. Computers and Mathematics with Applications, 135 (??):193–205, April 1, 2023. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122123000688.

Prusak:2024:OBD

[PTNR24] Ivan Prusak, Davide Torlo, Monica Nonino, and Gianluigi Rozza. An optimisation– based domain–decomposition reduced order model for parameter–dependent non– stationary fluid dynamics problems. Computers and Mathematics with Applications, 166(??):253– 268, July 15, 2024. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122124002098.

Puppi:2025:IDS

[Pup25] Riccardo Puppi. Isogeometric discretizations of the Stokes problem on trimmed geometries. Computers and Mathematics with Applications, 195(??):376–395, October 1, 2025. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122125002779.

Pauly:2020:PFT

[PV20] Dirk Pauly and Jan Valdman. Poincaré–Friedrichs type constants for operators involving grad, curl, and div: Theory and numerical experiments. Computers and Mathematics with Applications, 79(11):3027– 3067, June 1, 2020. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122120300110.

Panasenko:2022:MAP

[PV22] Grigory Panasenko and Marie-Claude Viallon. Method of asymptotic partial decomposition with discontinuous junctions. Computers and Mathematics with Applications, 105(??):75–93, January 1, 2022. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122121004144.

Pabisz:2025:CEI

[PVC^{+}25] Magdalena Pabisz, Askold Vilkha, Dominika Ciupek, Maciej Woźniak, and Maciej Paszyński. Concurrency of the exponential integrators for the glioblastoma brain tumor simulations. Computers and Mathematics with Applications, 196 (??):13–29, October 15, 2025. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122125002846.

Pan:2025:SUO

[PWBS25] Jiangong Pan, Wei Wan, Chenlong Bao, and Zuoqiang Shi. Solving unbalanced optimal transport on point cloud by tangent radial basis function method. Computers and Mathematics with Applications, 195(??): 161–176, October 1, 2025. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122125003050.

Pan:2023:DEG

[PWC^{+}23] Qingquan Pan, Lianjie Wang, Yun Cai, Xiaojing Liu, and Jinbiao Xiong. Densityextrapolation global variance reduction (DeGVR) method for large-scale radiation field calculation. Computers and Mathematics with Applications, 143(??):10–22, August 1, 2023. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122123001608.

Peng:2024:LBS

[PWJ^{+}24] Cheng Peng, Lian-Ping Wang, Li Ji, Songying Chen, and Zuchao Zhu. Lattice Boltzmann simulations of homogeneous shear turbulence laden with finite-size particles. Computers and Mathematics with Applications, 154 (??):65–77, January 15, 2024. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122123005060.

Partl:2023:RFD

[PWMC23] Ondřej Pártl, Ulrich Wilbrandt, Joaquín Mura, and Alfonso Caiazzo. Reconstruction of flow domain boundaries from velocity data via multistep optimization of distributed resistance. Computers and Mathematics with Applications, 129(??):11–33, January 1, 2023. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S089812212200462X.

Peng:2023:DNS

[PWWJ23] Cheng Peng, Xinnan Wu, Lian-Ping Wang, and Li Ji. Direct numerical simulation of homogeneous shear turbulence subject to a shear periodic boundary with the lattice Boltzmann method. Computers and Mathematics with Applications, 146(??):192–199, September 15, 2023. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122123002833.

Pan:2023:FSP

[PXLY23] Yunming Pan, Qiuyan Xu, Zhiyong Liu, and Jiye Yang. Fast successive permutation iterative algorithms for solving convection diffusion equation. Computers and Mathematics with Applications, 152(??):268–293, December 15, 2023. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122123004807.

Pei:2023:DNV

[PZL23] Lifang Pei, Chaofeng Zhang, and Meng Li. Dissipative nonconforming virtual element method for the fourth order nonlinear extended Fisher–Kolmogorov equation. Computers and Mathematics with Applications, 152(??):28–45, December 15, 2023. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122123004534.

Pu:2023:INE

[PZZ23] T. M. Pu, Y. Zhang, and C. H. Zhou. Implementation of a nonequilibrium analytical wall model in the local domainfree-discretization method for large eddy simulation. Computers and Mathematics with Applications, 133 (??):155–173, March 1, 2023. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122123000123.

Qin:2024:LGS

[QC24] Yonghui Qin and Yifan Cao. Legendre Galerkin spectral collocation least squares method for the Darcy flow in homogeneous medium and non-homogeneous medium. Computers and Mathematics with Applications, 166 (??):24–36, July 15, 2024. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122124001779.

Quiroz:2020:HHO

[QD20] Daniel Castanon Quiroz and Daniele A. Di Pietro. A hybrid high-order method for the incompressible Navier– Stokes problem robust for large irrotational body forces. Computers and Mathematics with Applications, 79(9): 2655–2677, May 1, 2020. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122119305632.

Qian:2023:PPN

[QD23] Xu Qian and Jian Dong.

Positivity-preserving nonstaggered central difference schemes solving the twolayer open channel flows. Computers and Mathematics with Applications, 148(??): 162–179, October 15, 2023. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122123003462.

Qu:2020:AAM

[QFL20] Wenzhen Qu, Chia-Ming Fan, and Xiaolin Li. Analysis of an augmented moving least squares approximation and the associated localized method of fundamental solutions. Computers and Mathematics with Applications, 80(1):13–30, July 1, 2020. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122120300821.

Qin:2023:PPT

[QFL23] Dan Qin, Kai Fu, and Dong Liang. Positivity preserving temporal second-order spatial fourth-order conservative characteristic methods for convection dominated diffusion equations. Computers and Mathematics with Applications, 149(??): 190–202, November 1, 2023. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122123003772.

Qiao:2022:FBA

[QGQ22] Leijie Qiao, Jing Guo, and Wenlin Qiu. Fast BDF2 ADI methods for the multi-dimensional tempered fractional integrodifferential equation of parabolic type. Computers and Mathematics with Applications, 123 (??):89–104, October 1, 2022.

CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122122003339.

Qi:2022:EAF

[QH22] Longzhao Qi and Yanren Hou. Error analysis of first- and second-order linear, unconditionally energystable schemes for the Swift– Hohenberg equation. Computers and Mathematics with Applications, 127(??):192– 212, December 1, 2022. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122122004308.

Qiu:2020:LBS

[QhCZ^{+}20] Ruo-Fan Qiu, Huan huan Che, Tao Zhou, Jian-Feng Zhu, and Yan-Cheng You. Lattice Boltzmann simulation for unsteady shock wave/boundary layer interaction in a shock tube. Computers and Mathematics with Applications, 80(10):2241– 2257, November 15, 2020. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122120302789.

Qin:2020:NAT

[QHHW20] Yi Qin, Yanren Hou, Pengzhan Huang, and Yongshuai Wang. Numerical analysis of two grad–div stabilization methods for the time-dependent Stokes/Darcy model. Computers and Mathematics with Applications, 79(3):817–832, February 1, 2020. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122119303888.

Qiu:2021:MLM

[Qiu21] Hailong Qiu. Multi-level mixed finite element algorithms for the stationary incompressible magnetohydrodynamics equations. Computers and Mathematics with Applications, 86 (??):33–48, March 15, 2021. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122121000213.

Qing:2024:AMG

[QL24] Lanyu Qing and Xiaolin Li. Analysis of a meshless generalized finite difference method for the time-fractional diffusionwave equation. Computers and Mathematics with Applications, 172(??):134–151, October 15, 2024. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122124003481.

Quaegebeur:2020:ICE

[QN20] Samuel Quaegebeur and Siva Nadarajah. Insights on the coercivity of the ESFR methods for elliptic problems. Computers and Mathematics with Applications, 80(9):2029–2044, November 1, 2020. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122120303515.

Quenjel:2024:ADS

[QP24] El-Houssaine Quenjel and Patrick Perré. Application of the 2D diamond scheme to predict the macroscopic thermal conductivity of a three-phase composite with local anisotropy directions. Computers and Mathematics with Applications, 163 (??):226–233, June 1, 2024. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122124001470.

Qiao:2021:SOA

[QQX21] Leijie Qiao, Wenlin Qiu, and Da Xu. A second-order ADI difference scheme based on non-uniform meshes for the three-dimensional nonlocal evolution problem. Computers and Mathematics with Applications, 102(??):137– 145, November 15, 2021. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S089812212100362X.

Qiu:2023:FEM

[QQY^{+}23] Meilan Qiu, Fang Qing, Xijun Yu, Jiangyong Hou, Dewang Li, and Xiaolong Zhao. Finite element method for the stationary dualporosity Navier–Stokes system with Beavers–Joseph interface conditions. Computers and Mathematics with Applications, 141(??):221– 241, July 1, 2023. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122123000160.

Qiu:2020:ASI

[QS20] Weifeng Qiu and Ke Shi. Analysis of a semi-implicit structure-preserving finite element method for the nonstationary incompressible magnetohydrodynamics equations. Computers and Mathematics with Applications, 80(10):2150–2161, November 15, 2020. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122120303539.

Qin:2023:MMB

[QSL23] Qiushuo Qin, Lina Song, and Fan Liu. A meshless method based on the generalized finite difference method for three-dimensional elliptic interface problems. Computers and Mathematics with Applications, 131(??):26–34, February 1, 2023. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122122004898.

Qi:2023:GWG

[QSW23] Wenya Qi, Padmanabhan Seshaiyer, and Junping Wang. Generalized weak Galerkin methods for Stokes equations. Computers and Mathematics with Applications, 134(??):181–193, March 15, 2023. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122123000147.

Querales:2022:MAA

[QV22] J. Querales and P. Venegas. Mixed approximation of the axisymmetric acoustic eigenvalue problem. Computers and Mathematics with Applications, 108(??):1–11, February 15, 2022. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122121004478.

Qin:2023:SOA

[QWC^{+}23] Yi Qin, Yang Wang, Lele Chen, Yi Li, and Jian Li. A second-order adaptive time filter algorithm with different subdomain variable time steps for the evolutionary Stokes/Darcy model. Computers and Mathematics with Applications, 150(??):170– 195, November 15, 2023. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122123004145.

Qiu:2025:PIR

[QWLQ25] Lin Qiu, Fajie Wang, Yingjie Liang, and Qing-Hua Qin. Physics-informed radial basis function network based on Hausdorff fractal distance for solving Hausdorff derivative elliptic problems. Computers and Mathematics with Applications, 183 (??):271–286, April 1, 2025. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122125000653.

Qiao:2023:SDS

[QX23] Yu Qiao and Xiangtuan Xiong. Simultaneous determination of the spacedependent source and initial value for a two-dimensional heat conduction equation. Computers and Mathematics with Applications, 147 (??):25–37, October 1, 2023. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122123003036.

Qiu:2020:ADI

[QXCG20] Wenlin Qiu, Da Xu, Haifan Chen, and Jing Guo. An alternating direction implicit Galerkin finite element method for the distributedorder time-fractional mobile– immobile equation in two dimensions. Computers and Mathematics with Applications, 80(12):3156–3172, December 15, 2020. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122120304284.

Qian:2023:SEE

[QZH23] Yanxia Qian, Yongchao Zhang, and Yunqing Huang. Stability and error estimates of GPAV-based unconditionally energy-stable scheme for phase field crystal equation. Computers and Mathematics with Applications, 151(??): 461–472, December 1, 2023. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122123004777.

Ravindran:2020:ELT

[Rav20] S. S. Ravindran. An efficient long-time stable secondorder accurate time-stepping scheme for evolutionary magneto-micropolar flows. Computers and Mathematics with Applications, 79(5): 1302–1320, March 1, 2020. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122119304390.

RambalducciDalla:2024:OEI

[RBSC24] Carlos Eduardo Rambalducci Dalla, Wellington Betencurte da Silva, Julio Cesar Sampaio Dutra, and Marcelo José Colaço. Online estimation of inlet contaminant concentration using Eulerian–Lagrangian method of fundamental solutions and Bayesian inference. Computers and Mathematics with Applications, 164 (??):131–138, June 15, 2024. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122124001780.

Ren:2025:OEE

[RC25] Jinmiao Ren and Xia Cui. Optimal error estimates of a decoupled, linear, unconditionally stable and charge-conservative finite element scheme for a thermally coupled incompressible IMHD system. Computers and Mathematics with Applications, 193(??):253– 278, September 1, 2025. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122125002627.

Ren:2023:IMM

[RCSM23] Jinlian Ren, David Culp, Brandon Smith, and Xia Ma. An investigation of the multi-mode Richtmyer– Meshkov instability at a gas/ HE interface using Pagosa. Computers and Mathematics with Applications, 139 (??):136–151, June 1, 2023. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122122001900.

Raeisi:2024:RPU

[RDF24] Banafsheh Raeisi, Mohammadreza Ahmadi Darani, and Mojtaba Fardi. The RBF partition of unity method for a 2D timefractional parabolic equation. Computers and Mathematics with Applications, 166 (??):237–252, July 15, 2024. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122124002244.

Ranocha:2020:FDE

[RDP20] Hendrik Ranocha, Lisandro Dalcin, and Matteo Parsani. Fully discrete explicit locally entropy-stable schemes for the compressible Euler and Navier– Stokes equations. Computers and Mathematics with Applications, 80(5):1343–1359, September 1, 2020. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122120302650.

DeRango:2021:OSE

[RFG^{+}21] Alessio De Rango, Luca Furnari, Andrea Giordano, Alfonso Senatore, Donato D’Ambrosio, William Spataro, Salvatore Straface, and Giuseppe Mendicino. Open-CAL system extension and application to the threedimensional Richards equation for unsaturated flow. Computers and Mathematics with Applications, 81(??): 133–158, January 1, 2021. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122120302248.

Rana:2021:NSC

[RG21] Puneet Rana and Gaurav Gupta. Numerical and sensitivity computations of threedimensional flow and heat transfer of nanoliquid over a wedge using modified Buongiorno model. Computers and Mathematics with Applications, 101(??):51–62, November 1, 2021. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122121003370.

Riya:2021:EAD

[RGL21] Riya, Bhupendra Gupta, and Subir Singh Lamba. An efficient anisotropic diffusion model for image denoising with edge preservation. Computers and Mathematics with Applications, 93 (??):106–119, July 1, 2021. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122121001103.

Ruhela:2022:NNC

[RGL22] Riya Ruhela, Bhupendra Gupta, and Subir Singh Lamba. A new nonconvex low rank minimization model to decompose an image into cartoon and texture components. Computers and Mathematics with Applications, 123(??):1–12, October 1, 2022. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122122003108.

Rutkowski:2020:OLO

[RGS^{+}20] Mariusz Rutkowski, Wojciech Gryglas, Jacek Szumbarski, Christopher Leonardi, and Lukasz Laniewski-Wołłk. Open-loop optimal control of a flapping wing using an adjoint Lattice Boltzmann method. Computers and Mathematics with Applications, 79(12):3547– 3569, June 15, 2020. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122120300870.

Roy:2024:LSA

[RGS24] Monisha Roy, Sukhendu Ghosh, and G. P. Raja Sekhar. Linear stability analysis of a Couette–Poiseuille flow: a fluid layer overlying an anisotropic and inhomogeneous porous layer. Computers and Mathematics with Applications, 173(??): 102–113, November 1, 2024. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S089812212400347X.

Roberts:2021:TSD

[RH21] Nathan V. Roberts and Stefan Henneking. Timestepping DPG formulations for the heat equation. Computers and Mathematics with Applications, 95(??):242– 255, August 1, 2021. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122120302467.

Rupp:2021:SEG

[RHA21] Andreas Rupp, Moritz Hauck, and Vadym Aizinger. A subcell-enriched Galerkin method for advection problems. Computers and Mathematics with Applications, 93 (??):120–129, July 1, 2021. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122121001425.

Rabago:2024:RAD

[RHA^{+}24] J. F. T. Rabago, A. Hadri, L. Afraites, A. S. Hendy, and M. A. Zaky. A robust alternating direction method of multipliers numerical scheme for solving geometric inverse problems in a shape optimization setting. Computers and Mathematics with Applications, 175(??): 19–32, December 1, 2024. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122124004012.

Rasin:2025:NDM

[RHA^{+}25] N. Rasin, H. Hirol, A. R. Zainal Abidin, M. H. Mokhtaram, M. A. Mohd Noor, and A. Y. Mohd Yassin. NURBS-Divergence-Meshless (NDM) method. Computers and Mathematics with Applications, 196(??):151–171, October 15, 2025. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122125003025.

Roper:2025:MLE

[RHB25] Taylor Roper, Harri Hakula, and Troy Butler. A machinelearning enabled framework for quantifying uncertainties in parameters of computational models. Computers and Mathematics with Applications, 182(??):184– 212, March 15, 2025. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122125000367.

Reuter:2021:FOU

[RHR^{+}21] Balthasar Reuter, Hennes Hajduk, Andreas Rupp, Florian Frank, Vadym Aizinger, and Peter Knabner. FES-TUNG 1.0: Overview, usage, and example applications of the MATLAB/GNU Octave toolbox for discontinuous Galerkin methods. Computers and Mathematics with Applications, 81 (??):3–41, January 1, 2021. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122120303254.

Roul:2025:NEA

[RKC25] P. Roul, S. N. Khandagale, and Jianxiong Cao. H^{1}-norm error analysis of an ADI compact finite difference method for a two-dimensional timefractional reaction–diffusion equation with variable coefficients. Computers and Mathematics with Applications, 194(??):135–157, September 15, 2025. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122125002548.

Roul:2025:GMT

[RKK25] Pradip Roul, Trishna Kumari, and Sameer N. Khandagale. A graded mesh technique for numerical approximation of a multi-term Caputo time-fractional Fokker– Planck equation in 2D space. Computers and Mathematics with Applications, 178(??): 92–107, January 15, 2025. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122124005285.

Rejwer-Kosinska:2022:ATK

[RKRRL22] Ewa Rejwer-Kosińska, Liliana Rybarska-Rusinek, and Aleksandr Linkov. On accuracy of translations by kernel independent fast multipole methods. Computers and Mathematics with Applications, 124(??):227–240, October 15, 2022. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122122003534.

Reutskiy:2020:RBT

[RL20] Sergiy Reutskiy and Ji Lin. A RBF-based technique for 3D convection–diffusion– reaction problems in an anisotropic inhomogeneous medium. Computers and Mathematics with Applications, 79(6):1875–1888, March 15, 2020. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122119304997.

Ren:2023:PCP

[RL23] Yuhang Ren and Demin Liu. Pressure correction projection finite element method for the 2D/3D timedependent thermomicropolar fluid problem. Computers and Mathematics with Applications, 136(??):136– 150, April 15, 2023. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122123000652.

Rais:2024:CBF

[RL24] Rania Rais and Frédérique Le Louër. On the coupling between finite elements and integral representation for linear elastic waves scattering problems: Analysis and simulation. Computers and Mathematics with Applications, 175(??):87–106, December 1, 2024. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122124004000.

Romanus:2021:IBL

[RLF21] Rodrigo S. Romanus, Alan Lugarini, and Admilson T. Franco. An immersed boundary-lattice Boltzmann framework for fully resolved simulations of nonspherical particle settling in unbounded domain. Computers and Mathematics with Applications, 102(??):206– 219, November 15, 2021. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122121003722.

Russo:2024:P

[RLS24] Alessandro Russo, Luciano Lopez, and Valeria Simonicini. Preface. Computers and Mathematics with Applications, 166(??): 50, July 15, 2024. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122124002050.

Rong:2023:LBM

[RLSC23] Fumei Rong, Qianhuan Li, Baochang Shi, and Zhenhua Chai. A lattice Boltzmann model based on Cole– Hopf transformation for Ndimensional coupled Burgers’ equations. Computers and Mathematics with Applications, 134(??):101– 111, March 15, 2023. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122123000172.

Ramirez:2021:SAF

[RLVR21] C. Ramirez, A. Leroyer, M. Visonneau, and Y. Roux. Study of asynchronicity in the fluid-structure interaction domain. Computers and Mathematics with Applications, 93(??):156– 167, July 1, 2021. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S089812212100122X.

Roberts:2024:ITD

[RMBC24] Nathan V. Roberts, Sean T. Miller, Stephen D. Bond, and Eric C. Cyr. An implicit-in-time DPG formulation of the 1D1V Vlasov– Poisson equations. Computers and Mathematics with Applications, 154(??):103– 119, January 15, 2024. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122123005047.

Robert:2025:ASS

[RNC25] Maria Robert, Suresh Kumar Nadupuri, and Nagaiah Chamakuri. Analysis and simulation of sparse optimal control of the monodomain model. Computers and Mathematics with Applications, 184(??):29– 44, April 15, 2025. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122125000616.

Reisch:2024:ANI

[RNMÖX24] Cordula Reisch, Adrián Navas-Montilla, and Ilhan Özgen-Xian. Analytical and numerical insights into wildfire dynamics: Exploring the advection–diffusion– reaction model. Computers and Mathematics with Applications, 158(??):179– 198, March 15, 2024. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122124000385.

Rafsanjani:2022:DBM

[RNN22] Hossein Khodabakhshi Rafsanjani, Hossein Noori, and Nasibe Naseri. Diffusion based method for impulse noise removal using residual feedback. Computers and Mathematics with Applications, 107(??):45–56, February 1, 2022. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122121004569.

Robinson:2025:CTS

[Rob25] Neville I. Robinson. Conformal transformation solutions of the extended motz problem. Computers and Mathematics with Applications, 183 (??):200–213, April 1, 2025. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122125000598.

Roul:2020:FON

[Rou20] Pradip Roul. A fourth order numerical method based on B-spline functions for pricing Asian options. Computers and Mathematics with Applications, 80(3):504–521, August 1, 2020. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122120301309.

Roul:2022:HON

[RR22] Pradip Roul and Vikas Rohil. A high-order numerical scheme based on graded mesh and its analysis for the two-dimensional time-fractional convectiondiffusion equation. Computers and Mathematics with Applications, 126(??):1–13, November 15, 2022. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122122003790.

Rukavishnikov:2023:WFE

[RR23] Viktor A. Rukavishnikov and Elena I. Rukavishnikova. Weighted finite element method and body of optimal parameters for elasticity problem with singularity. Computers and Mathematics with Applications, 151(??): 408–417, December 1, 2023. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122123004686.

Ray:2021:AFE

[RS21] Tanushree Ray and Rajen Kumar Sinha. An adaptive finite element method for parabolic interface problems with nonzero flux jumps. Computers and Mathematics with Applications, 82(??):97–112, January 15, 2021. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122120304399.

Ramalho:2024:FVC

[RSAG^{+}24] Luís D. C. Ramalho, Isidro J. Sánchez-Arce, Diogo C. Gonçalves, Raul D. S. G. Campilho, and Jorge Belinha. Forced vibrations of a cantilever beam using radial point interpolation methods: a comparison study. Computers and Mathematics with Applications, 163 (??):14–26, June 1, 2024. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122124001111.

Rasmussen:2021:OPM

[RSB^{+}21] Atgeirr Flø Rasmussen, Tor Harald Sandve, Kai Bao, Andreas Lauser, Joakim Hove, Bård Skaflestad, Robert Klöfkorn, Markus Blatt, Alf Birger Rustad, Ove Sævareid, Knut-Andreas Lie, and Andreas Thune. The Open Porous Media Flow reservoir simulator. Computers and Mathematics with Applications, 81(??):159– 185, January 1, 2021. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122120302182.

Rojas:2022:PFF

[RST^{+}22] Roberto Rojas, Verónica Sotomayor, Tomohiro Takaki, Kosuke Hayashi, and Akio Tomiyama. A phase fieldfinite difference lattice Boltzmann method for modeling dendritic growth solidification in the presence of melt convection. Computers and Mathematics with Applications, 114(??):180– 187, May 15, 2022. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122122001407.

Ren:2020:ESG

[RTA20] Shixian Ren, Ting Tan, and Jing An. An efficient spectral-Galerkin approximation based on dimension reduction scheme for transmission eigenvalues in polar geometries. Computers and Mathematics with Applications, 80(5):940–955, September 1, 2020. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122120302339.

Rojas-Ulloa:2024:IMG

[RUMT^{+}24] Carlos Rojas-Ulloa, Hélène Morch, Víctor Tuninetti, Laurent Duchêne, and Anne Marie Habraken. Implementation of a modified Graham– Walles viscosity function within a Chaboche viscoplastic constitutive model. Computers and Mathematics with Applications, 155(??):165– 175, February 1, 2024. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122123005515.

Ren:2023:CND

[RWaZ23] Jinmiao Ren, Bo Wang, and Guang an Zou. A Crank–Nicolson discontinuous Galerkin pressureprojection method for the hydrodynamic and sediment transport model. Computers and Mathematics with Applications, 142(??):175– 197, July 15, 2023. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122123001748.

Ren:2024:PBP

[RWZG24] Baifen Ren, Bao-Shan Wang, Xiangxiong Zhang, and Zhen Gao. Positivity and bound preserving wellbalanced high order compact finite difference scheme for Ripa and pollutant transport model. Computers and Mathematics with Applications, 176(??):545–563, December 15, 2024. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122124005091.

Rezaei:2021:NPB

[RYAM21] M. Rezaei, A. R. Yazdanian, A. Ashrafi, and S. M. Mahmoudi. Numerical pricing based on fractional black–Scholes equation with time-dependent parameters under the CEV model: Double barrier options. Computers and Mathematics with Applications, 90 (??):104–111, May 15, 2021. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122121000900.

Reisinger:2021:PSP

[RZ21] Christoph Reisinger and Yufei Zhang. A penalty scheme and policy iteration for nonlocal HJB variational inequalities with monotone nonlinearities. Computers and Mathematics with Applications, 93(??):199– 213, July 1, 2021. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122121001437.

Rachowicz:2021:DPG

[RZC21] Waldemar Rachowicz, Adam Zdunek, and Witold Cecot. A discontinuous Petrov–Galerkin method for compressible Navier–Stokes equations in three dimensions. Computers and Mathematics with Applications, 102(??):113–136, November 15, 2021. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122121003588.

Rek:2021:MMR

[RZŠ21] Zlatko Rek, Rizwan Zahoor, and Božidar Šarler. Modified method of regularized sources for potential flow. Computers and Mathematics with Applications, 88(??): 110–119, April 15, 2021. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122120302443.

Ran:2023:NPT

[RZSL23] Guoquan Ran, Qinghe Zhang, Gaochuang Shi, and Xia Li. A new Poissontype equation applicable to the three-dimensional nonhydrostatic model in the framework of the discontinuous Galerkin method. Computers and Mathematics with Applications, 145(??): 13–23, September 1, 2023. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122123002481.

Sharma:2022:ENW

[SA22] Vatsalya Sharma and Ashwani Assam. Evaluation of novel wall function approach for supersonic flow problems involving separations induced by geometry and shocks. Computers and Mathematics with Applications, 123(??):75–88, October 1, 2022. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122122003145.

Szabo:2024:DPA

[SA24] B. Szabó and R. Actis. The demarcation problem in the applied sciences. Computers and Mathematics with Applications, 162(??):206– 214, May 15, 2024. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122124001317.

Sajjadi:2021:EIH

[SADW21] H. Sajjadi, M. Atashafrooz, A. Amiri Delouei, and Y. Wang. The effect of indoor heating system location on particle deposition and convection heat transfer: DMRT-LBM. Computers and Mathematics with Applications, 86(??):90–105, March 15, 2021. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122121000262.

Szabo:2020:VPF

[SAR20] Barna Szabó, Ricardo Actis, and David Rusk. Validation of a predictor of fatigue failure in the high-cycle range. Computers and Mathematics with Applications, 80(11): 2451–2461, December 2020. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122120303114.

Sauter:2023:ISC

[Sau23] Stefan Sauter. The inf–sup constant for hp-Crouzeix– Raviart triangular elements. Computers and Mathematics with Applications, 149(??): 49–70, November 1, 2023. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122123003619.

Salazar:2022:SSA

[SAVHB22] Fabián Mauricio Vélez Salazar, Iván David Patiño Arcila, Carlos Alberto Ruiz Villa, and Abraham Hernández-Blanquisett. In-silico study about the influence of electroporation parameters on the chemotherapeutic drug transport in cancer tissues using the meshless method of approximate particular solutions. Computers and Mathematics with Applications, 125(??):116–135, November 1, 2022. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122122003546.

Sadeghi:2021:MST

[SB21] Sadegh Sadeghi and Mehdi Bidabadi. Mathematical study of threshold of thermal-diffusive instability in counter-flow nonpremixed biomass-fueled flames considering effective parameters. Computers and Mathematics with Applications, 81(??):602–617, January 1, 2021. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122119303177.

Sahu:2024:IDI

[SB24a] Abhishek Sahu and Shubhankar Bhowmick. Investigation of diabatic imparity involving asymmetric convection in two-dimensional longitudinal fins using lattice Boltzmann solver. Computers and Mathematics with Applications, 162(??):79– 93, May 15, 2024. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122124001044.

Santana:2024:EAN

[SB24b] Claudio Marcio Santana and Jose Eduardo Mautone Barros. Experimental analysis and numerical simulation of ignition delay time of diesel fuel using a shock tube. Computers and Mathematics with Applications, 158 (??):74–84, March 15, 2024. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122123005916.

Santra:2025:EAA

[SB25] Sudarshan Santra and Ratikanta Behera. Enhancing accuracy with an adaptive discretization for the nonlocal integro-partial differential equations involving initial time singularities. Computers and Mathematics with Applications, 192(??):212– 239, August 15, 2025. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122125002238.

Siddiqa:2021:ETR

[SBH^{+}21] Sadia Siddiqa, Naheed Begum, Md. Anwar Hossain, M. N. Abrar, Rama Subba Reddy Gorla, and Qasem Al-Mdallal. Effect of thermal radiation on conjugate natural convection flow of a micropolar fluid along a vertical surface. Computers and Mathematics with Applications, 83(??):74–83, February 1, 2021. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122120300249.

Sorgente:2022:PMQ

[SBMS22] T. Sorgente, S. Biasotti, G. Manzini, and M. Spagnuolo. Polyhedral mesh quality indicator for the Virtual Element Method. Computers and Mathematics with Applications, 114(??):151– 160, May 15, 2022. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122122001390.

Storvik:2023:RSS

[SBNR23] Erlend Storvik, Jakub Wiktor Both, Jan Martin Nordbotten, and Florin Adrian Radu. A robust solution strategy for the Cahn– Larché equations. Computers and Mathematics with Applications, 136(??):112– 126, April 15, 2023. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122123000536.

Sun:2024:NAC

[SBS24] Xiaoshu Sun, Timo Betcke, and Alexander Strohmaier. Numerical aspects of Casimir energy computation in acoustic scattering. Computers and Mathematics with Applications, 170(??):260–274, September 15, 2024. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122124002931.

Singh:2025:CGL

[SBY^{+}25] Chetan Singh, Deepak Bhoriya, Anshu Yadav, Harish Kumar, and Dinshaw S. Balsara. Chew, Goldberger & Low equations: Eigensystem analysis and applications to one-dimensional test problems. Computers and Mathematics with Applications, 188 (??):195–220, June 15, 2025. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122125001543.

Safari:2021:NAS

[SC21a] Farzaneh Safari and Wen Chen. Numerical approximations for space–time fractional Burgers’ equations via a new semi-analytical method. Computers and Mathematics with Applications, 96(??):55–66, August 15, 2021. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122121001012.

Sun:2021:NIR

[SC21b] Xiang Sun and Jung-Il Choi.

Non-intrusive reduced-order modeling for uncertainty quantification of space– time-dependent parameterized problems. Computers and Mathematics with Applications, 87(??):50–64, April 1, 2021. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122121000250.

Shang:2021:DUG

[SCCS21] Jinlong Shang, Zhenhua Chai, Xinmeng Chen, and Baochang Shi. Discrete unified gas kinetic scheme for incompressible Navier– Stokes equations. Computers and Mathematics with Applications, 97(??):45–60, September 1, 2021. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122121001991.

Schoof:2023:SDC

[SCD23] R. Schoof, G. F. Castelli, and W. Dörfler. Simulation of the deformation for cycling chemo-mechanically coupled battery active particles with mechanical constraints. Computers and Mathematics with Applications, 149(??):135–149, November 1, 2023. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122123003723.

Senesi:2025:EFI

[SCD25] Marco Desiderio Loffredo Senesi, Alessandro Coclite, and Michele Dassisti. On the effect of force isotropy in a multiphase cascaded lattice Boltzmann scheme with entropic stabilization. Computers and Mathematics with Applications, 194(??):214– 230, September 15, 2025. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122125002603.

Shukla:2021:HOD

[SCdH21] Khemraj Shukla, Jesse Chan, and Maarten V. de Hoop. A high order discontinuous Galerkin method for the symmetric form of the anisotropic viscoelastic wave equation. Computers and Mathematics with Applications, 99(??):113–132, October 1, 2021. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S089812212100287X.

Shi:2023:BPI

[SCG23] Chengxin Shi, Hao Cheng, and Xiaoxiao Geng. The backward problem for an inhomogeneous time-fractional diffusion-wave equation in an axis-symmetric cylinder. Computers and Mathematics with Applications, 137 (??):44–60, May 1, 2023. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122123000731.

Schaback:2021:ATV

[Sch21] Robert Schaback. An approximation theorist’s view on solving operator equations — with special attention to Trefftz, MFS, MPS, and DRM methods. Computers and Mathematics with Applications, 88 (??):70–77, April 15, 2021. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122119304523.

Schroder:2024:BFM

[Sch24] Andreas Schröder. C^{k} basis functions on meshes with hanging nodes. Computers and Mathematics with Applications, 166(??):169– 181, July 15, 2024. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122124001792.

Chen:2023:NES

[sCpLYhG23] Shu sheng Chen, Jin ping Li, Wu Yuan, and Zheng hong Gao. Nonlinear entropy stable Riemann solver with heuristic logarithmic pressure augmentation for supersonic and hypersonic flows. Computers and Mathematics with Applications, 141 (??):33–41, July 1, 2023. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122123001256.

Seshadri:2021:RSI

[SD21] Pradeep Kumar Seshadri and Ashoke De. A robust sharp interface based immersed boundary framework for moving body problems with applications to laminar incompressible flows. Computers and Mathematics with Applications, 83(??): 24–56, February 1, 2021. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S089812211930553X.

Safari:2025:SAC

[SD25a] Farzaneh Safari and Yanjun Duan. A semi-analytic collocation technique for solving 3D anomalous non-linear thermal conduction problem associated with the Caputo fractional derivative. Computers and Mathematics with Applications, 178 (??):81–91, January 15, 2025. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122124005261.

Sankhla:2025:STD

[SD25b] Kunj Sankhla and Sandipan Kumar Das. Solution of three-dimensional steady state Stokes flow problems with exit boundary conditions by the boundary element method. Computers and Mathematics with Applications, 195(??):1–13, October 1, 2025. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122125002767.

Shi:2020:CNR

[SDG20] Kehan Shi, Gang Dong, and Zhichang Guo. Cauchy noise removal by nonlinear diffusion equations. Computers and Mathematics with Applications, 80(9):2090–2103, November 1, 2020. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122120303412.

Sahabandu:2022:SAM

[SDJ^{+}22] C. W. Sahabandu, M. Dewasurendra, Z. A. M. S. Juman, K. Vajravelu, and Ali J. Chamkha. Semi-analytical method for propagation of harmonic waves in nonlinear magneto-thermo-elasticity. Computers and Mathematics with Applications, 105(??): 107–111, January 1, 2022. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S089812212100417X.

Salvador:2021:NIR

[SDM21] M. Salvador, L. Dedè, and A. Manzoni. Non intrusive reduced order modeling of parametrized PDEs by kernel POD and neural networks. Computers and Mathematics with Applications, 104(??): 1–13, December 15, 2021. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122121003928.

Sebestova:2024:RBP

[Šeb24] Ivana Šebestová. Residualbased a posteriori error estimation for E-based formulation of a time-dependent eddy current problem. Computers and Mathematics with Applications, 174(??):204– 218, November 15, 2024. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122124003730.

Sun:2025:TSF

[SF25] Kexin Sun and Minfu Feng. Time-space fractional anisotropic diffusion equations for multiplicative noise removal. Computers and Mathematics with Applications, 184(??):1–28, April 15, 2025. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122125000604.

Shafei:2021:NTV

[SFR21] Erfan Shafei, Shirko Faroughi, and Timon Rabczuk. Nonlinear transient vibration of viscoelastic plates: a NURBS-based isogeometric HSDT approach. Computers and Mathematics with Applications, 84(??):1–15, February 15, 2021. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122120304685.

Sahoo:2023:RUC

[SG23] Sanjay Ku Sahoo and Vikas Gupta. A robust uniformly convergent finite difference scheme for the timefractional singularly perturbed convection-diffusion problem. Computers and Mathematics with Applications, 137(??):126–146, May 1, 2023. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122123000755.

Scheuer:2020:MEP

[SGAC^{+}20] A. Scheuer, G. Grégoire, E. Abisset-Chavanne, F. Chinesta, and R. Keunings. Modelling the effect of particle inertia on the orientation kinematics of fibres and spheroids immersed in a simple shear flow. Computers and Mathematics with Applications, 79(3):539–554, February 1, 2020. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122118307375.

Saurabh:2021:ISL

[SGF^{+}21] Kumar Saurabh, Boshun Gao, Milinda Fernando, Songzhe Xu, Makrand A. Khanwale, Biswajit Khara, Ming-Chen Hsu, Adarsh Krishnamurthy, Hari Sundar, and Baskar Ganapathysubramanian. Industrial scale large eddy simulations with adaptive octree meshes using immersogeometric analysis. Computers and Mathematics with Applications, 97(??): 28–44, September 1, 2021.

CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122121002145.

Safa:2020:SPD

[SGJJ20] Raziyeh Safa, Ataallah Soltani Goharrizi, Saeed Jafari, and Ebrahim Jahanshahi Javaran. Simulation of particles dissolution in the shear flow: a combined concentration lattice Boltzmann and smoothed profile approach. Computers and Mathematics with Applications, 79(3): 603–622, February 1, 2020. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122119303530.

Sluzalec:2023:QOH

[SGR^{+}23] Tomasz Służalec, Rafał Grzeszczuk, Sergio Rojas, Witold Dzwinel, and Maciej Paszyński. Quasi-optimal hpfinite element refinements towards singularities via deep neural network prediction. Computers and Mathematics with Applications, 142 (??):157–174, July 15, 2023. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122123001591.

Storti:2020:MFC

[SGSD20] Bruno Storti, Luciano Garelli, Mario Storti, and Jorge D’Elía. A matrix-free Chimera approach based on Dirichlet–Dirichlet coupling for domain composition purposes. Computers and Mathematics with Applications, 79(12):3310– 3330, June 15, 2020. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122120300468.

Scabbia:2023:MIP

[SGZ^{+}23] Francesco Scabbia, Claudia Gasparrini, Mirco Zaccariotto, Ugo Galvanetto, Adam Larios, and Florin Bobaru. Moving interfaces in peridynamic diffusion models and the influence of discontinuous initial conditions: Numerical stability and convergence. Computers and Mathematics with Applications, 151(??):384–396, December 1, 2023. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122123004637.

Song:2022:TPS

[SH22] Shengzhong Song and Zhengda Huang. A two-parameter shift-splitting preconditioner for saddle point problems. Computers and Mathematics with Applications, 124 (??):7–20, October 15, 2022.

CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122122003388.

Shen:2023:ESO

[SHD23] Xiaojuan Shen, Yunqing Huang, and Xiaojing Dong. An effective second-order scheme for the nonstationary incompressible magnetohydrodynamics equations. Computers and Mathematics with Applications, 139 (??):184–208, June 1, 2023. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122122002905.

Shi:2024:ATV

[Shi24] Kehan Shi. Adaptive total variational regularization of Gaussian denoisers for multiplicative noise removal. Computers and Mathematics with Applications, 168(??): 207–217, August 15, 2024. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122124002852.

Sayyar:2021:HOS

[SHM21] Golsa Sayyar, Seyed Mohammad Hosseini, and Farinaz Mostajeran. A highorder scheme for time-space fractional diffusion equations with Caputo-Riesz derivatives. Computers and Mathematics with Applications, 104(??):34–43, December 15, 2021. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S089812212100393X.

Shi:2020:APS

[SHQL20] Yuying Shi, Zhimei Huo, Jing Qin, and Yilin Li. Automatic prior shape selection for image edge detection with modified Mumford– Shah model. Computers and Mathematics with Applications, 79(6):1644–1660, March 15, 2020. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122119304778.

Sarker:2025:MMM

[SHR25] Abu Naser Sarker, Ronald D. Haynes, and Michael Robertson. A moving mesh method for pitting corrosion of heterogeneous materials. Computers and Mathematics with Applications, 191(??):48–59, August 1, 2025. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S089812212500149X.

Sun:2024:SPH

[SHT^{+}24] Zewei Sun, Qingzhi Hou, Arris S. Tijsseling, Jijian Lian, and Jianguo Wei. Smoothed particle hydrodynamics with diffusive flux for advection–diffusion equation with discontinuities. Computers and Mathematics with Applications, 160 (??):70–85, April 15, 2024. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122124000518.

Shao:2022:CVM

[SHWC22] Yuanzhen Shao, Elizabeth Hawkins, Kai Wang, and Zhan Chen. A constrained variational model of biomolecular solvation and its numerical implementation. Computers and Mathematics with Applications, 107 (??):17–28, February 1, 2022. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122121004430.

Sun:2024:SFS

[SHXZ24] Zheng Sun, Yunjun Hua, Yunqing Xu, and Xiaomin Zhou. Simulation of fluidstructure interaction using the density smoothing Bspline material point method with a contact approach. Computers and Mathematics with Applications, 176(??): 525–544, December 15, 2024. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122124004991.

Suliman:2025:SEN

[SIAA25] Muhammad Suliman, Muhammad Ibrahim, Ebrahem A. Algehyne, and Vakkar Ali. A study of an efficient numerical method for solving the generalized fractional reaction-diffusion model involving a distributed-order operator along with stability analysis. Computers and Mathematics with Applications, 180(??):61–75, February 15, 2025. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122124005510.

Shi:2021:SAM

[SJ21] Dongyang Shi and Xu Jia. Superconvergence analysis of a mixed finite element approximation for the nonlinear fourth-order Rosenau-RLW equation. Computers and Mathematics with Applications, 98(??):169–180, September 15, 2021. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122121002686.

Shi:2023:SAF

[SJ23] Dongyang Shi and Peng Jian. Superconvergence analysis of finite element method for nonlinear semiconductor device problem. Computers and Mathematics with Applications, 138 (??):1–22, May 15, 2023. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122123000664.

Sha:2024:VRM

[SJ24] Bojiao Sha and Zupeng Jia. A very robust MMALE method based on a novel VoF method for two-dimensional compressible fluid flows. Computers and Mathematics with Applications, 160 (??):86–107, April 15, 2024. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122124000415.

Singh:2024:ILW

[SJM24] Bhagwan Singh, Komal Jangid, and Santwana Mukhopadhyay. Implementation of Legendre wavelet method for the size dependent bending analysis of nano beam resonator under nonlocal strain gradient theory. Computers and Mathematics with Applications, 153(??):94–107, January 1, 2024. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122123004893.

Sun:2022:CSB

[SJZ22] Qihang Sun, Bingquan Ji, and Luming Zhang. A convex splitting BDF2 method with variable time-steps for the extended Fisher– Kolmogorov equation. Computers and Mathematics with Applications, 114(??):73– 82, May 15, 2022. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122122001006.

Shi:2024:FMC

[SJZ24] Shan Shi, Xiaoyun Jiang, and Hui Zhang. A fast method and convergence analysis for the MHD flow model of generalized secondgrade fluid. Computers and Mathematics with Applications, 171(??):175–187, October 1, 2024. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122124003213.

Shin:2021:TSS

[SK21] Jaemin Shin and Mi-Young Kim. P_{0} time/space subcell limiting DG-DGLM method for hyperbolic systems of conservation laws. Computers and Mathematics with Applications, 94(??):114– 135, July 15, 2021. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122121001711.

Sato:2023:DSP

[SK23] Kenta Sato and Shunichi Koshimura. Development of a single-phase freesurface flow model with the improved lattice kinetic scheme. Computers and Mathematics with Applications, 145(??):275–288, September 1, 2023. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122123002766.

Singh:2025:CNC

[SK25] Harpal Singh and Arbaz Khan. Conforming/nonconforming mixed finite element methods for optimal control of velocity–vorticity– pressure formulation for the Oseen problem with variable viscosity. Computers and Mathematics with Applications, 198(??):59–92, November 15, 2025. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S089812212500330X.

Sivalingam:2024:CNN

[SKG24] S. M. Sivalingam, Pushpendra Kumar, and V. Govindaraj. A Chebyshev neural network-based numerical scheme to solve distributedorder fractional differential equations. Computers and Mathematics with Applications, 164(??):150– 165, June 15, 2024. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122124001639.

Sun:2025:SAA

[SKY25] Keyue Sun, Xiangjie Kong, and Junxiang Yang. Semianalytical algorithm for quasicrystal patterns. Computers and Mathematics with Applications, 180(??):130– 143, February 15, 2025. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122124005601.

Shi:2020:SAN

[SL20] Dongyang Shi and Qian Liu. Superconvergent analysis of a nonconforming mixed finite element method for non-stationary conduction– convection problem. Computers and Mathematics with Applications, 79(2):230–243, January 15, 2020. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122119303232.

Shi:2021:NAS

[SL21] Xiangyu Shi and Linzhang Lu. A new approach of superconvergence analysis of nonconforming Wilson finite element for semilinear parabolic problem. Computers and Mathematics with Applications, 94 (??):28–37, July 15, 2021. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122121001620.

Shi:2024:OEE

[SL24] Dongyang Shi and Conggang Liang. Optimal error estimates of nonconforming Wilson element for nonlinear nonlocal parabolic problems. Computers and Mathematics with Applications, 163 (??):234–241, June 1, 2024. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122124001809.

Sun:2024:DOV

[SLC^{+}24] Qizheng Sun, Xiaojing Liu, Xiang Chai, Hui He, and Tengfei Zhang. A discreteordinates variational nodal method for heterogeneous neutron Boltzmann transport problems. Computers and Mathematics with Applications, 170(??):142– 160, September 15, 2024. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122124003006.

Safari:2025:LFC

[SLD25] Farzaneh Safari, Ji Lin, and Yanjun Duan. A localized Fourier collocation method for the numerical solution of nonlinear fractional Fisher– Kolmogorov equation. Computers and Mathematics with Applications, 193(??):241– 252, September 1, 2025. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122125002652.

Song:2020:GFD

[SLGF20] Lina Song, Po-Wei Li, Yan Gu, and Chia-Ming Fan. Generalized finite difference method for solving stationary 2D and 3D Stokes equations with a mixed boundary condition. Computers and Mathematics with Applications, 80(6):1726–1743, September 15, 2020. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122120303102.

Shi:2023:HOM

[SLM23] Zhengguang Shi, Pin Lyu, and Jingtang Ma. Highorder methods for the option pricing under multivariate rough volatility models. Computers and Mathematics with Applications, 139 (??):173–183, June 1, 2023. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122122002656.

Shi:2024:SAN

[SLT24] Dongyang Shi, Minghao Li, and Qili Tang. Superconvergence analysis of a new stabilized nonconforming finite element method for the Stokes equations. Computers and Mathematics with Applications, 169(??):253– 259, September 1, 2024. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122124003201.

Szczucka-Lasota:2021:IMF

[SLUW^{+}21]  Bozena  Szczucka-Lasota,

Anita Uscilowska, Tomasz Wegrzyn, Zbigniew Stanik, and Jan Piwnik. Implementation of the Method of Fundamental Solutions for the correction of parameters of the thermal HM spraying process. Computers and Mathematics with Applications, 88(??):2–15, April 15, 2021. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122118306916.

Shi:2022:NSE

[SLW22] Xiangyu Shi, Linzhang Lu, and Haijie Wang. New superconvergence estimates of FEM for time-dependent Joule heating problem. Computers and Mathematics with Applications, 111(??):91– 97, April 1, 2022. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122122000682.

Shi:2023:SPS

[SLX^{+}23] Yu Shi, Yang Liu, Junhua Xue, Pengxiang Zhao, and Shugang Li. Study on particles sedimentation in porous media with the immersed boundary-lattice Boltzmann flux solver. Computers and Mathematics with Applications, 129(??):1–10, January 1, 2023. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122122004758.

Sun:2022:FSP

[SLYL22] Ziyu Sun, Yang Liu, Baoli Yin, and Hong Li. Fast structure-preserving difference algorithm for 2D nonlinear space-fractional wave models. Computers and Mathematics with Applications, 123(??):40–58, October 1, 2022. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S089812212200311X.

Saffarian:2022:RPO

[SM22a] Marziyeh Saffarian and Akbar Mohebbi. Reduced proper orthogonal decomposition spectral element method for the solution of 2D multi-term time fractional mixed diffusion and diffusion–wave equations in linear and nonlinear modes. Computers and Mathematics with Applications, 117 (??):127–154, July 1, 2022. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122122000797.

Shao:2022:NGV

[SM22b] Xin-Hui Shao and Hui-Nan Meng. A new generalized variant of the deteriorated PSS preconditioner for nonsymmetric saddle point problems. Computers and Mathematics with Applications, 118 (??):208–213, July 15, 2022. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122122002231.

Sana:2023:DNN

[SM23] Soura Sana and Bankim C. Mandal. Dirichlet–Neumann and Neumann–Neumann waveform relaxation algorithms for heterogeneous sub-diffusion and diffusionwave equations. Computers and Mathematics with Applications, 150(??):102– 124, November 15, 2023. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S089812212300398X.

Santra:2023:HOA

[SMDC23] Sudarshan Santra, Jugal Mohapatra, Pratibhamoy Das, and Debajyoti Choudhuri. Higher order approximations for fractional order integroparabolic partial differential equations on an adaptive mesh with error analysis. Computers and Mathematics with Applications, 150(??): 87–101, November 15, 2023. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122123003930.

Sashko:2025:PFL

[SMŁWL25] Dmytro Sashko, Travis R. Mitchell, Łukasz Łaniewski-Wołłk, and Christopher R. Leonardi. Phase field lattice Boltzmann method for liquid–gas flows in complex geometries with efficient and consistent wetting boundary treatment. Computers and Mathematics with Applications, 186(??):101– 129, May 15, 2025. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122125001099.

Sa:2024:TOD

[SMR^{+}24a] Luís F. N. Sá, Felipe Silva Maffei, Lucas N. B. S. Ribeiro, Julio Romano Meneghini, and Emílio Carlos Nelli Silva. Topology optimization design of labyrinth sealtype devices considering subsonic compressible turbulent flow conditions. Computers and Mathematics with Applications, 176(??):306– 323, December 15, 2024. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122124004711.

Singh:2024:NAS

[SMR24b] Ankit Singh, Vikas Maurya, and Manoj K. Rajpoot. Numerical analysis and simulation of European options under liquidity shocks: a coupled semilinear system approach with new IMEX methods. Computers and Mathematics with Applications, 165 (??):28–38, July 1, 2024. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122124001603.

Sheikhbahaei:2023:EPF

[SMS23a] Pouria Sheikhbahaei, Farshid Mossaiby, and Arman Shojaei. An efficient peridynamic framework based on the arc-length method for fracture modeling of brittle and quasi-brittle problems with snapping instabilities. Computers and Mathematics with Applications, 136 (??):165–190, April 15, 2023. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122123000792.

Stokke:2023:ASS

[SMS^{+}23b] Jakob S. Stokke, Koondanibha Mitra, Erlend Storvik, Jakub W. Both, and Florin A. Radu. An adaptive solution strategy for Richards’ equation. Computers and Mathematics with Applications, 152(??):155–167, December 15, 2023. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122123004674.

Santos:2021:CAN

[SMV^{+}21] Isaac P. Santos, Sandra M. C. Malta, Andrea M. P. Valli, Lucia Catabriga, and Regina C. Almeida. Convergence analysis of a new dynamic diffusion method. Computers and Mathematics with Applications, 98(??): 1–9, September 15, 2021. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122121002509.

Saidoglu:2020:MAR

[SN20a] Nataliya Y. Saidoglu and Alexander I. Nosich. Method of analytical regularization in the analysis of axially symmetric excitation of imperfect circular disk antennas. Computers and Mathematics with Applications, 79(10): 2872–2884, May 15, 2020. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122119305887.

Singh:2020:NSS

[SN20b] Maneesh Kumar Singh and Srinivasan Natesan. Numerical solution of 2D singularly perturbed reaction–diffusion system with multiple scales. Computers and Mathematics with Applications, 80 (4):36–53, August 15, 2020. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122120301759.

Sulman:2021:DDP

[SNHH21] M. H. M. Sulman, T. B. Nguyen, R. D. Haynes, and Weizhang Huang. Domain decomposition parabolic Monge– Ampère approach for fast generation of adaptive moving meshes. Computers and Mathematics with Applications, 84(??):97–111, February 15, 2021. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122120304697.

Sky:2025:FTS

[SNHZ25] Adam Sky, Michael Neunteufel, Jack S. Hale, and Andreas Zilian. Formulae and transformations for simplicial tensorial finite elements via polytopal templates. Computers and Mathematics with Applications, 195(??): 322–348, October 1, 2025. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122125003177.

Solovyev:2023:NIW

[SNL23] Sergey Solovyev, Mikhail Novikov, and Vadim Lisitsa. A numerical investigation of wave-induced fluid flows in anisotropic fractured porous media. Computers and Mathematics with Applications, 140(??):78– 88, June 15, 2023. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122123001098.

Saad:2025:SNP

[SOPH25] Khaled M. Saad, Kolade M.

Owolabi, Edson Pindza, and Waleed M. Hamanah. Simulating natural patterns via space-time accurate barycentric collocation methods for nonlinear PDEs. Computers and Mathematics with Applications, 199(??):378–402, December 1, 2025. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S089812212500433X.

Sharma:2020:MNF

[SP20] Nitin Sharma and Subrata Kumar Panda. Multiphysical numerical (FE–BE) solution of sound radiation responses of laminated sandwich shell panel including curvature effect. Computers and Mathematics with Applications, 80(5):1221–1239, September 1, 2020. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122120302595.

Shah:2023:DFE

[SPC^{+}23] Murad Ali Shah, Kejia Pan, Rui Chen, Zhengru Zhang, and Dongdong He. A decoupled finite element scheme for simulating the dynamics of red blood cells in an L-shaped cavity. Computers and Mathematics with Applications, 140(??):169– 182, June 15, 2023. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122123001414.

Su:2021:WDG

[SPZL21] Guanting Su, Tianyu Pan, Mengzong Zheng, and Qiushi Li. A well-defined grid line-based immersed boundary method for efficient and accurate simulations of incompressible flow. Computers and Mathematics with Applications, 89(??):99– 115, May 1, 2021. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122121000559.

She:2023:FTI

[SQ23] Zi-Hang She and Li-Min Qiu. Fast TTTS iteration methods for implicit Runge– Kutta temporal discretization of Riesz space fractional advection-diffusion equations. Computers and Mathematics with Applications, 141(??):42–53, July 1, 2023. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122123001074.

Shi:2024:USAb

[SQ24] Dongyang Shi and Zhenqi Qi. Unconditional superconvergence analysis of an energy-preserving finite element scheme for nonlinear BBM equation. Computers and Mathematics with Applications, 153(??):172– 185, January 1, 2024. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122123005230.

Sun:2022:ILE

[SQB22] Di Sun, Feng Qu, and Junqiang Bai. An implicit large eddy simulation method based on all-speed schemes. Computers and Mathematics with Applications, 114 (??):1–20, May 15, 2022. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122122001274.

Safari:2023:TDM

[SQC23] Farzaneh Safari, Tong Qingshan, and Wen Chen. Time discretization for modeling migration of groundwater contaminant in the presence of micro-organisms via a semi-analytic method. Computers and Mathematics with Applications, 151(??):397– 407, December 1, 2023. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122123004625.

Sun:2024:TDE

[SQGZ24] Wenxiang Sun, Wenzhen Qu, Yan Gu, and Shengdong Zhao. Three-dimensional elastodynamic analysis employing the generalized finite difference method with arbitrary-order accuracy. Computers and Mathematics with Applications, 174(??): 151–166, November 15, 2024. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122124003845.

She:2021:SCF

[SQL21] Zi-Hang She, Hai-Dong Qu, and Xuan Liu. Stability and convergence of finite difference method for twosided space-fractional diffusion equations. Computers and Mathematics with Applications, 89(??):78–86, May 1, 2021. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122121000663.

Sun:2021:IGM

[SQLZ21] Di Sun, Feng Qu, Qingsong Liu, and Jiaxiang Zhong. Improvement of the genuinely multidimensional ME-AUSMPW scheme for subsonic flows. Computers and Mathematics with Applications, 96(??):1–15, August 15, 2021. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122121001917.

Sang:2020:CBS

[SR20] Yanbin Sang and Yan Ren.

A critical p-biharmonic system with negative exponents. Computers and Mathematics with Applications, 79(5): 1335–1361, March 1, 2020. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122119304353.

Song:2021:ROF

[SR21] Junpeng Song and Hongxing Rui. A reduced-order finite element method based on POD for the incompressible miscible displacement problem. Computers and Mathematics with Applications, 98(??):99–117, September 15, 2021. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122121002601.

Song:2022:CSM

[SR22] Junpeng Song and Hongxing Rui. A combined stabilized mixed finite element and discontinuous Galerkin method for coupled Stokes and Darcy flows with transport. Computers and Mathematics with Applications, 120(??):92– 104, August 15, 2022. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122122002693.

Sass:2023:ARE

[SR23] Hauke Sass and Arnold Reusken. An accurate and robust Eulerian finite element method for partial differential equations on evolving surfaces. Computers and Mathematics with Applications, 146(??):253–270, September 15, 2023. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122123002924.

Song:2024:ROS

[SR24] Junpeng Song and Hongxing Rui. A reduced-order Schwarz domain decomposition method based on POD for the convectiondiffusion equation. Computers and Mathematics with Applications, 160(??):60– 69, April 15, 2024. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122124000555.

Siddique:2020:EMF

[SRA20] J. I. Siddique, U. A. Raja, and A. Ahmed. Effects of magnetic field on porosity and solid deformation for radial fluid flow through deformable porous shells. Computers and Mathematics with Applications, 80(5):1104– 1116, September 1, 2020. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122120302509.

Shi:2025:SSQ

[SRD25] Jinsong Shi, Kaysar Rahman, and Jiawen Deng. Split-step quintic uniform algebraic trigonometric tension B-spline collocation method for cubic Ginzburg– Landau equations. Computers and Mathematics with Applications, 196(??):395– 414, October 15, 2025. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122125003311.

Soutter:2021:MMD

[SRP21] Emile Soutter, Jacques Rappaz, and Marco Picasso. A mixture model for a dilute dispersion of gas in a liquid flow: Mathematical analysis and application to aluminium electrolysis. Computers and Mathematics with Applications, 90 (??):55–65, May 15, 2021. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122121000754.

Shukla:2020:MAC

[SS20] Vivekanand Shukla and Jeeoot Singh. Modeling and analysis of cross-ply and angle-ply laminated plates under patch loads using RBF based meshfree method and new HSDT. Computers and Mathematics with Applications, 79(8): 2240–2257, April 15, 2020. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122119305218.

Sun:2021:GPD

[SS21] Yizhi Sun and Zhilin Sun. Generating probability distributions on intervals and spheres with application to finite element method. Computers and Mathematics with Applications, 95(??): 282–295, August 1, 2021. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122120304193.

Shen:2022:BST

[SS22] Qin-Qin Shen and Quan Shi. Block symmetrictriangular preconditioners for generalized saddle point linear systems from piezoelectric equations. Computers and Mathematics with Applications, 119(??):100– 117, August 1, 2022. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122122002413.

Sun:2024:GPD

[SS24] Yizhi Sun and Zhilin Sun.

Generating probability distributions on intervals and spheres: Convex decomposition. Computers and Mathematics with Applications, 154 (??):12–23, January 15, 2024. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122123004984.

Shiromani:2023:HOH

[SSD23] Ram Shiromani, Vembu Shanthi, and Pratibhamoy Das. A higher order hybridnumerical approximation for a class of singularly perturbed two-dimensional convection–diffusion elliptic problem with non-smooth convection and source terms. Computers and Mathematics with Applications, 142 (??):9–30, July 15, 2023.

CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122123001402.

Silva:2020:GSC

[SSdA20] Edcarlos D. Silva, Maxwell L. Silva, and J. C. de Albuquerque. Ground states for a class of critical quasilinear coupled superlinear elliptic systems. Computers and Mathematics with Applications, 79(4):889–907, February 15, 2020. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122119303943.

Snyder:2023:NCS

[SSK^{+}23] William Snyder, Alex Santiago Anaya, Justin Krometis, Traian Iliescu, and Raffaella De Vita. A numerical comparison of simplified Galerkin and machine learning reduced order models for vaginal deformations. Computers and Mathematics with Applications, 152(??):168– 180, December 15, 2023. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122123004650.

Shao:2020:CMA

[SSW20] Wenting Shao, Qiwei Sheng, and Cheng Wang. A cascadic multigrid asymptoticpreserving discrete ordinate discontinuous streamline diffusion method for radiative transfer equations with diffusive scalings. Computers and Mathematics with Applications, 80(6):1650–1667, September 15, 2020. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122120303084.

Salhi:2022:WPN

[SSY22] Loubna Salhi, Mohammed Seaid, and Driss Yakoubi. Well-posedness and numerical approximation of steady convection-diffusion-reaction problems in porous media. Computers and Mathematics with Applications, 124(??): 129–148, October 15, 2022. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122122003443.

Sharafian:2020:FSM

[SSZ20] Amin Sharafian, Alireza Sharifi, and Weidong Zhang. Fractional sliding mode based on RBF neural network observer: Application to HIV infection mathematical model. Computers and Mathematics with Applications, 79(11):3179– 3188, June 1, 2020. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122120300304.

Sajavicius:2020:INB

[ST20a] Svajūnas Sajavičius and Thomas Takacs. Imposing nonlocal boundary conditions in Galerkintype methods based on non-interpolatory functions. Computers and Mathematics with Applications, 80(12): 2877–2895, December 15, 2020. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0898122120303783.

Sun:2020:HOU

[ST20b] Y. X. Sun and Z. F. Tian. High-order upwind compact finite-difference lattice Boltzmann method for viscous incompressible flows. Computers and Mathematics with Applications, 80(7): 1858–1872, October 1, 2020. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122120303217.

Su:2022:PPF

[ST22] Shuai Su and Huazhong Tang. A positivity-preserving and free energy dissipative hybrid scheme for the Poisson–Nernst–Planck equations on polygonal and polyhedral meshes. Computers and Mathematics with Applications, 108(??):33–48, February 15, 2022. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122121004600.

Slobodkins:2023:NEA

[ST23] Arkadijs Slobodkins and Johannes Tausch. A node elimination algorithm for cubature of high-dimensional polytopes. Computers and Mathematics with Applications, 144(??):229–236, August 15, 2023. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122123002468.

Seyednazari:2021:ESE

[STM21] Salar Seyednazari, Mehdi Tatari, and Davoud Mirzaei. Error and stability estimates of a least-squares variational kernel-based method for second order elliptic PDEs. Computers and Mathematics with Applications, 103 (??):1–11, December 1, 2021. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122121003734.

Storn:2020:RDP

[Sto20] Johannes Storn. On a relation of discontinuous Petrov– Galerkin and least-squares finite element methods. Computers and Mathematics with Applications, 79(12):3588– 3611, June 15, 2020. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122120300857.

Sekine:2022:INE

[STO22] Kouta Sekine, Kazuaki Tanaka, and Shin’ichi Oishi. Inverse norm estimation of perturbed Laplace operators and corresponding eigenvalue problems. Computers and Mathematics with Applications, 106(??):18–26, January 15, 2022. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122121004284.

Song:2023:SSS

[SU23] Yanlai Song and Malik Zaka Ullah. On a sparse and stable solver on graded meshes for solving high-dimensional parabolic pricing PDEs. Computers and Mathematics with Applications, 143(??): 224–233, August 1, 2023. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122123002079.

Schafelner:2021:NRA

[SV21] Andreas Schafelner and Panayot S. Vassilevski. Numerical results for adaptive (negative norm) constrained first order system least squares formulations. Computers and Mathematics with Applications, 95(??): 256–270, August 1, 2021. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122120303394.

Stevenson:2020:UPP

[SvV20] Rob Stevenson and Raymond van Venetië. Uniform preconditioners for problems of positive order. Computers and Mathematics with Applications, 79(12):3516– 3530, June 15, 2020. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122120300766.

Seibel:2020:RBE

[SW20a] Daniel Seibel and Steffen Weißer. Recovery-based error estimators for the VEM and BEM-based FEM. Computers and Mathematics with Applications, 80(9):2073– 2089, November 1, 2020. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122120303540.

Sun:2020:EIM

[SW20b] Zhen-Wei Sun and Li Wang. ERHSS iteration method for PDE optimal control problem. Computers and Mathematics with Applications, 79(3):656–672, February 1, 2020. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122119303633.

Stevenson:2021:MRS

[SW21] Rob Stevenson and Jan Westerdiep. Minimal residual space-time discretizations of parabolic equations: Asymmetric spatial operators. Computers and Mathematics with Applications, 101(??): 107–118, November 1, 2021. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122121003485.

Shen:2024:LSF

[SW24a] Mengxia Shen and Haiyong Wang. A least-squares Fourier frame method for nonlocal diffusion models on arbitrary domains. Computers and Mathematics with Applications, 175(??):433– 446, December 1, 2024. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S089812212400467X.

Sun:2024:LRN

[SW24b] Jingbo Sun and Fei Wang.

Local randomized neural networks with discontinuous Galerkin methods for diffusive-viscous wave equation. Computers and Mathematics with Applications, 154(??):128–137, January 15, 2024. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122123005199.

Sun:2024:DMT

[SWC24] Yao Sun, Pan Wang, and Bo Chen. A decomposition method for two and three dimensional fluid-solid interaction scattering problem. Computers and Mathematics with Applications, 165 (??):106–117, July 1, 2024. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S089812212400172X.

Si:2022:MCP

[SWW22] Zhiyong Si, Qing Wang, and Yunxia Wang. A modified characteristic projection finite element method for the Kelvin–Voigt viscoelastic fluid equations.

Computers and Mathematics with Applications, 109 (??):44–57, March 1, 2022. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122122000323.

Shen:2023:SFS

[SWWZ23] Xiaoqin Shen, Lixiao Wu, Juan Wen, and Juan Zhang. SAV Fourier-spectral method for diffuse-interface tumorgrowth model. Computers and Mathematics with Applications, 140(??):250– 259, June 15, 2023. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122122004497.

Sun:2023:NTO

[SWZ23] Qihang Sun, Jindi Wang, and Luming Zhang. A new third-order energy stable technique and error estimate for the extended Fisher– Kolmogorov equation. Computers and Mathematics with Applications, 142(??):198– 207, July 15, 2023. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S089812212300144X.

Sun:2023:TAW

[SWZC23] Tingting Sun, Peng Wang, Guanjun Zhang, and Yingbin Chai. Transient analyses of wave propagations in nonhomogeneous media employing the novel finite element method with the appropriate enrichment function. Computers and Mathematics with Applications, 129 (??):90–112, January 1, 2023. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S089812212200428X.

Shi:2025:NED

[SX25] Dongyang Shi and Xuemiao Xu. A new energy dissipationpreserving Crank–Nicolson type nonconforming FEM for damped wave equation with cubic nonlinearity. Computers and Mathematics with Applications, 190(??):170– 184, July 15, 2025. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122125001798.

Shi:2024:CDB

[SXCB24] Ziyu Shi, Yu Xiang, Jie Chen, and Yingchao Bao. A coupled double boundary Burton–Miller method without hypersingular integral. Computers and Mathematics with Applications, 167 (??):1–11, August 1, 2024. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122124002062.

Song:2024:UES

[SXKL24] Xin Song, Qing Xia, Junseok Kim, and Yibao Li. An unconditional energy stable data assimilation scheme for Navier–Stokes–Cahn– Hilliard equations with local discretized observed data. Computers and Mathematics with Applications, 164 (??):21–33, June 15, 2024. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122124001251.

Sun:2021:STB

[SXLZ21] HongGuang Sun, Yi Xu, Ji Lin, and Yuhui Zhang. A space-time backward substitution method for three-dimensional advectiondiffusion equations. Computers and Mathematics with Applications, 97(??):77–85, September 1, 2021. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S089812212100211X.

Su:2025:MPP

[SY25] Sheng Su and Junxiang Yang. Maximum principlepreserving, unconditionally energy-stable, and convergent method with secondorder accuracy for the phasefield model of image inpainting. Computers and Mathematics with Applications, 183 (??):32–45, April 1, 2025. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122125000380.

Si:2024:AMP

[SYHN24] Zhanfei Si, Tiantang Yu, Hirshikesh, and Sundararajan Natarajan. An adaptive multi-patch isogeometric phase-field model for dynamic brittle fracture. Computers and Mathematics with Applications, 153(??):1–19, January 1, 2024. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122123004947.

Shao:2023:SFI

[SYZ23] Baiqiang Shao, Haijian Yang, and Hong-Jie Zhao. Scalable fully implicit methods for subsurface flows in porous media with fractional derivative. Computers and Mathematics with Applications, 134(??):55–65, March 15, 2023. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122123000020.

Soleymani:2021:RFS

[SZ21] Fazlollah Soleymani and Shengfeng Zhu. RBF-FD solution for a financial partialintegro differential equation utilizing the generalized multiquadric function. Computers and Mathematics with Applications, 82(??): 161–178, January 15, 2021. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122120304351.

Shi:2022:LCN

[SZ22a] Dongyang Shi and Houchao Zhang. A linearized conservative nonconforming FEM for nonlinear Klein–Gordon– Schrödinger equations. Computers and Mathematics with Applications, 106(??):57–73, January 15, 2022. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122121004326.

Shi:2022:UOE

[SZ22b] Dongyang Shi and Houchao Zhang. Unconditional optimal error estimates and superconvergence analysis of energy-preserving FEM for general nonlinear Schrödinger equation with wave operator. Computers and Mathematics with Applications, 128(??):79–95, December 15, 2022. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122122004333.

Sheng:2024:ABF

[SZ24a] Ying Sheng and Tie Zhang.

An analysis of the bilinear finite volume method for the singularly-perturbed convection-diffusion problems on Shishkin mesh. Computers and Mathematics with Applications, 174(??):84–96, November 15, 2024. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122124003833.

Shi:2024:USAa

[SZ24b] Dongyang Shi and Lingen Zhang. Unconditional superconvergence analysis of modified finite difference streamlined diffusion method for nonlinear convectiondominated diffusion equation. Computers and Mathematics with Applications, 153 (??):81–93, January 1, 2024. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122123005114.

Sun:2024:SBM

[SZCW24] Linlin Sun, Qing Zhang, Zhikang Chen, and Xing Wei. A singular boundary method for transient coupled dynamic thermoelastic analysis. Computers and Mathematics with Applications, 158(??):259– 274, March 15, 2024. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122124000567.

Shao:2020:GTT

[SZS20] Xin-Hui Shao, Zhen-Duo Zhang, and Hai-Long Shen. A generalization of trigonometric transform splitting methods for spatial fractional diffusion equations. Computers and Mathematics with Applications, 79(6): 1845–1856, March 15, 2020. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122119304924.

Tang:2022:EAS

[TA22] Zheqian Tang and Rong An. Error analysis of the secondorder BDF finite element scheme for the thermally coupled incompressible magnetohydrodynamic system. Computers and Mathematics with Applications, 118 (??):110–119, July 15, 2022. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122122002024.

Tadi:2023:NMI

[Tad23]  M. Tadi.  A numer-

ical method for inverse Helmholtz problem based on approximate inverse of a matrix. Computers and Mathematics with Applications, 150(??):125–131, November 15, 2023. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122123004054.

Tan:2024:ARN

[Tan24] Zhijun Tan. An α-robust and new two-grid nonuniform L2-1_{σ} FEM for nonlinear timefractional diffusion equation. Computers and Mathematics with Applications, 174(??): 530–552, November 15, 2024. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122124004668.

Tan:2025:UES

[Tan25] Zengqiang Tan. Unconditionally energy stable and second-order accurate oneparameter ESAV schemes with non-uniform time stepsizes for the functionalized Cahn–Hilliard equation. Computers and Mathematics with Applications, 182(??): 163–183, March 15, 2025.

CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122125000318.

Tausch:2022:AQR

[Tau22] Johannes Tausch. Adaptive quadrature rules for Galerkin BEM. Computers and Mathematics with Applications, 113(??):270– 281, May 1, 2022. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122122001213.

Tyagi:2020:RBS

[TBO20] Mohit Tyagi, Paul E. Barbone, and Assad A. Oberai. Residual-based stabilized formulation for the solution of inverse elliptic partial differential equations. Computers and Mathematics with Applications, 80(5): 822–836, September 1, 2020. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122120301632.

Toutlini:2025:SIS

[TBS25] Nour-Eddine Toutlini, Abdelaziz Beljadid, and Azzeddine Soulaïmani. A semiimplicit second-order temporal scheme for solving the pressure head-based form of Richards’ and advectiondispersion equations. Computers and Mathematics with Applications, 187(??):106– 131, June 1, 2025. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122125001063.

Tan:2025:NPM

[TC25] Jianguo Tan and Jiling Cao.

Nonlinear PDE model for pricing European options with transaction costs under the 3/2 non-affine stochastic volatility model. Computers and Mathematics with Applications, 196(??):246–262, October 15, 2025. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122125003037.

Tan:2022:SAB

[TCA22a] Ting Tan, Waixiang Cao, and Jing An. Spectral approximation based on a mixed scheme and its error estimates for transmission eigenvalue problems. Computers and Mathematics with Applications, 111 (??):20–33, April 1, 2022. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122122000669.

Torregrosa:2022:HTB

[TCA^{+}22b] Sergio Torregrosa, Victor Champaney, Amine Ammar, Vincent Herbert, and Francisco Chinesta. Hybrid twins based on optimal transport. Computers and Mathematics with Applications, 127(??): 12–24, December 1, 2022. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122122004060.

Tian:2023:TGD

[TCW23] Zhikun Tian, Yanping Chen, and Jianyun Wang. A twogrid discretization method for nonlinear Schrödinger equation by mixed finite element methods. Computers and Mathematics with Applications, 130(??):10–20, January 15, 2023. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122122004783.

Tang:2023:PIN

[TFWX23] Siping Tang, Xinlong Feng, Wei Wu, and Hui Xu. Physics-informed neural networks combined with polynomial interpolation to solve nonlinear partial differential equations. Computers and Mathematics with Applications, 132(??):48–62, February 15, 2023. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122122005090.

Tinoco-Guerrero:2023:MFD

[TGARGT^{+}23] Gerardo Tinoco-Guerrero, Heriberto Arias-Rojas, José Alberto Guzmán-Torres, Ricardo Román-Gutiérrez, and José Gerardo Tinoco-Ruiz. A meshless finite difference scheme applied to the numerical solution of wave equation in highly irregular space regions. Computers and Mathematics with Applications, 136(??):25– 33, April 15, 2023. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122123000482.

Tinoco-Guerrero:2025:MMG

[TGDMGT^{+}25] Gerardo Tinoco-Guerrero, Francisco Javier Domínguez-Mota, José Alberto Guzmán-Torres, Gabriela Pedraza-Jiménez, and José Gerardo Tinoco-Ruiz. mGFD: a meshless generalized finite difference method. Computers and Mathematics with Applications, 195(??):396– 418, October 1, 2025. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122125003232.

Tse:2021:RST

[TGL21] Louis K. S. Tse, Yu Guan, and Larry K. B. Li. RANS simulations of terrain-disrupted turbulent airflow at Hong Kong International Airport. Computers and Mathematics with Applications, 81(??): 737–758, January 1, 2021. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122119305619.

Trong:2021:BPT

[TH21] Dang Duc Trong and Dinh Nguyen Duy Hai. Backward problem for timespace fractional diffusion equations in Hilbert scales. Computers and Mathematics with Applications, 93 (??):253–264, July 1, 2021. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122121001589.

Tang:2023:FAB

[TH23] Shi-Ping Tang and Yu-Mei Huang. A fast ADI based matrix splitting preconditioning method for the high dimensional space fractional diffusion equations in conservative form. Computers and Mathematics with Applications, 144(??):210– 220, August 15, 2023. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122123002316.

Tang:2025:PCD

[TH25] Zhenjie Tang and Li He.

Physics-constrained deep kernel learning for inverse problems with noisy data. Computers and Mathematics with Applications, 196(??): 135–150, October 15, 2025. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122125003104.

Takacs:2020:SDM

[THF20] Bálint Takács, Róbert Horváth, and István Faragó. Space dependent models for studying the spread of some diseases. Computers and Mathematics with Applications, 80(2):395–404, July 2020. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122119303396.

Tan:2024:SCM

[TjW24] Jia Tan and Tian jun Wang.

Spectral collocation method coupled with domain decomposition for exterior problems of the Fisher equation. Computers and Mathematics with Applications, 175(??): 289–312, December 1, 2024. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122124004462.

Tian:2022:RAU

[TJYL22] Chenlei Tian, Ting Jin, Xiangfeng Yang, and Qinyu Liu. Reliability analysis of the uncertain heat conduction model. Computers and Mathematics with Applications, 119(??):131– 140, August 1, 2022. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122122002255.

Tanwar:2020:XFE

[TK20] Sonam Tanwar and Preety Kumari. X-FEM: an efficient algorithm for simulation of phase transition during prostate cryosurgery. Computers and Mathematics with Applications, 79(8): 2119–2132, April 15, 2020. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122119304389.

Tushar:2022:VEM

[TKK22] Jai Tushar, Anil Kumar, and Sarvesh Kumar. Virtual element methods for general linear elliptic interface problems on polygonal meshes with small edges. Computers and Mathematics with Applications, 122(??):

61–75, September 15, 2022. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122122003066.

Tushar:2023:MVE

[TKK23] Jai Tushar, Anil Kumar, and Sarvesh Kumar. Mixed virtual element methods for optimal control of Darcy flow. Computers and Mathematics with Applications, 140 (??):134–153, June 15, 2023. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S089812212300158X.

Tisha:2025:DAA

[TKMK25] Mayesha Sharmim Tisha, Md. Kamrujjaman, Muhammad Mohebujjaman, and Taufiquar Khan. Decoupled algorithms and analyses for an advection–reaction– diffusion model with stocking and harvesting. Computers and Mathematics with Applications, 189(??):24– 47, July 1, 2025. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122125001208.

Titakis:2024:CSN

[TKT^{+}24] George Titakis, Iasson Karafyllis, Dionysios Theodosis, Ioannis Papamichail, and Markos Papageorgiou. A comparative study of numerical methods for approximating the solutions of a macroscopic automatedvehicle traffic flow model. Computers and Mathematics with Applications, 176(??): 469–490, December 15, 2024. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S089812212400498X.

Tang:2022:NHE

[TL22] Shujiang Tang and Mingjun Li. A novel high efficiency adaptive mapped WENO scheme. Computers and Mathematics with Applications, 124(??):149–162, October 15, 2022. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122122003431.

Tian:2024:NBB

[TLJ24] Hao Tian, Junke Lu, and Lili Ju. A novel bond-based nonlocal diffusion model with matrix-valued coefficients in non-divergence form and its collocation discretization. Computers and Mathematics with Applications, 173(??): 33–46, November 1, 2024. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122124003432.

Tang:2020:LOL

[TLZF20] Xialan Tang, Zhibin Liu, Baiju Zhang, and Minfu Feng. A low-order lockingfree virtual element for linear elasticity problems. Computers and Mathematics with Applications, 80(5):1260– 1274, September 1, 2020. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122120301863.

Tavares:2023:SLB

[TMML23] H. S. Tavares, B. Magacho, L. Moriconi, and J. B. R. Loureiro. A simplified lattice Boltzmann implementation of the quasi-static approximation in pipe flows under the presence of non-uniform magnetic fields. Computers and Mathematics with Applications, 148(??):93–112, October 15, 2023. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122123003322.

Tanbay:2020:FMS

[TO20] T. Tanbay and B. Ozgener.

Fully meshless solution of the one-dimensional multigroup neutron transport equation with the radial basis function collocation method.

Computers and Mathematics with Applications, 79(5): 1266–1286, March 1, 2020. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122119304419.

Toprakseven:2022:WGF

[Top22]  Şuayip Toprakseven.  A

weak Galerkin finite element method on temporal graded meshes for the multiterm time fractional diffusion equations. Computers and Mathematics with Applications, 128(??):108– 120, December 15, 2022. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122122004357.

Torsu:2020:VIM

[Tor20] Prosper Torsu. On variational iterative methods for semilinear problems. Computers and Mathematics with Applications, 80(5):1164– 1175, September 1, 2020. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122120302571.

Tiwari:2024:FRC

[TPA24] Awanish Kumar Tiwari, Ambuj Pandey, and Akash Anand. A fast, rapidly convergent Nyström solver for electromagnetic scattering from orthotropic media. Computers and Mathematics with Applications, 154 (??):58–64, January 15, 2024. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S089812212300500X.

Triet:2020:ASB

[TPOT20] Nguyen Anh Triet, Nguyen Duc Phuong, Donal O’Regan, and Nguyen Huy Tuan. Approximate solution of the backward problem for Kirchhoff’s model of parabolic type with discrete random noise. Computers and Mathematics with Applications, 80 (3):453–470, August 1, 2020. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S089812212030119X.

Tang:2023:AOS

[TQX23] Bo Tang, Leijie Qiao, and Da Xu. An ADI orthogonal spline collocation method for a new two-dimensional distributed-order fractional integro-differential equation. Computers and Mathematics with Applications, 132(??): 104–118, February 15, 2023. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122122005077.

Tian:2023:HOE

[TRL23] Zhihui Tian, Maohua Ran, and Yang Liu. Higher-order energy-preserving difference scheme for the fourth-order nonlinear strain wave equation. Computers and Mathematics with Applications, 135 (??):124–133, April 1, 2023. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122123000329.

Trent:2023:UIP

[TRSM23] Stephen Trent, Jamil Renno, Sadok Sassi, and M. Shadi Mohamed. Using image processing techniques in computational mechanics. Computers and Mathematics with Applications, 136(??):1–24, April 15, 2023. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S089812212200493X.

Tao:2021:GSM

[TS21] Zhen-Zhen Tao and Bing Sun. Galerkin spectral method for a fourth-order optimal control problem with H ^{1}-norm state constraint. Computers and Mathematics with Applications, 97(??): 1–17, September 1, 2021. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122121002091.

Tao:2024:SMP

[TS24] Yutian Tao and Eftychios Sifakis. A symmetric multigrid-preconditioned Krylov subspace solver for Stokes equations. Computers and Mathematics with Applications, 172(??):168–180, October 15, 2024. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122124003729.

Tian:2021:NSN

[TSARR21] Zhe Tian, Amin Shahsavar, Abdullah A. A. A. Al-Rashed, and Sara Rostami. Numerical simulation of nanofluid convective heat transfer in an oblique cavity with conductive edges equipped with a constant temperature heat source: Entropy production analysis. Computers and Mathematics with Applications, 81(??): 725–736, January 1, 2021. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122119305656.

Theodosiou:2024:EPI

[TSS24] Eleni Theodosiou, Jochen Schütz, and David Seal. An explicitness-preserving IMEX-split multiderivative method. Computers and Mathematics with Applications, 158(??):139–149, March 15, 2024. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S089812212400021X.

Tang:2025:SDD

[TSW^{+}25] Xiaosong Tang, Jiaxin Shen, Xinchang Wang, Zhaoyun Zeng, and Jingwen Zhu. Spatiotemporal dynamics deduced by nonlocal delay competition in a diffusive Lotka– Volterra population model. Computers and Mathematics with Applications, 196(??): 172–182, October 15, 2025. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122125002974.

Takacs:2025:MRI

[TT25] Stefan Takacs and Stefan Tyoler. Multi-resolution isogeometric analysis — efficient adaptivity utilizing the multi-patch structure. Computers and Mathematics with Applications, 179(??): 103–125, February 1, 2025. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122124005388.

Tuan:2020:REE

[TTKV20] Nguyen Huy Tuan, Bui Le Trong Thanh, Mokhtar Kirane, and Phan Thi Khanh Van. Regularization and error estimate for an initial inverse nonlocal diffusion problem. Computers and Mathematics with Applications, 79(12):3331– 3352, June 15, 2020. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122120300547.

Turk:2020:MSE

[Tür20] Önder Türk. An MHD Stokes eigenvalue problem and its approximation by a spectral collocation method. Computers and Mathematics with Applications, 80(9):2045–2056, November 1, 2020. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122120303527.

Trong:2021:TDS

[TVKN21] Dang Duc Trong, Tran Quoc Viet, Vo Dang Khoa, and Nguyen Thi Hong Nhung. A two-dimensional sideways problem with random discrete data. Computers and Mathematics with Applications, 86(??):16–32, March 15, 2021. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122121000237.

Tannahill:2023:MAI

[TW23] Connor Tannahill and Justin Wan. MM-ADMM: Implicit integration of MM-PDEs in parallel. Computers and Mathematics with Applications, 141(??):67– 79, July 1, 2023. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S089812212300113X.

Tao:2022:IRS

[TWH^{+}22] Shi Tao, Liang Wang, Qing He, Jiechao Chen, and Jiahong Luo. Interface-resolved simulation of solid-liquid phase change with thermal convection using discrete unified gas kinetic scheme. Computers and Mathematics with Applications, 125(??): 213–221, November 1, 2022. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122122003716.

Tian:2020:EOP

[TZ20] Yingxu Tian and Haoyan Zhang. European option pricing under stochastic volatility jump-diffusion models with transaction cost. Computers and Mathematics with Applications, 79(9): 2722–2741, May 1, 2020. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122119305590.

Tu:2021:BAA

[TZ21] Xuemin Tu and Jinjin Zhang. BDDC algorithms for advection-diffusion problems with HDG discretizations. Computers and Mathematics with Applications, 101(??): 74–106, November 1, 2021. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122121003400.

Tan:2025:PPP

[TZ25] Yan Tan and Jun Zhu. Purepositivity-preserving methods with an optimal sufficient CFL number for fifthorder MR-WENO schemes on structured meshes. Computers and Mathematics with Applications, 177(??):1–22, January 1, 2025. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122124005066.

Tang:2023:MWG

[TZX23] Ming Tang, Liuqiang Zhong, and Yingying Xie. A modified weak Galerkin method for H (curl)-elliptic problem.

Computers and Mathematics with Applications, 139 (??):224–229, June 1, 2023. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122122003984.

Uriarte:2025:OVP

[UBP^{+}25] Carlos Uriarte, Manuela Bastidas, David Pardo, Jamie M. Taylor, and Sergio Rojas. Optimizing variational physics-informed neural networks using least squares. Computers and Mathematics with Applications, 185(??):76–93, May 1, 2025. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122125000811.

Usman:2024:RIS

[UHL^{+}24] Muhammad Usman, Muhammad Hamid, Dianchen Lu, Zhengdi Zhang, and Wojciech Sumelka. Robust iterative spectral algorithms for smooth solutions of timefractional nonlinear diffusion problems and convergence analysis. Computers and Mathematics with Applications, 175(??):487–508, December 1, 2024. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122124004607.

Usman:2024:NSS

[UHLZ24] Muhammad Usman, Muhammad Hamid, Dianchen Lu, and Zhengdi Zhang. Nonsmooth solutions of timefractional Allen–Cahn problems via novel operational matrix based semi-spectral method with convergence analysis. Computers and Mathematics with Applications, 159(??):122–141, April 1, 2024. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122124000506.

Uy:2023:OIR

[UHP23] Wayne Isaac Tan Uy, Dirk Hartmann, and Benjamin Peherstorfer. Operator inference with roll outs for learning reduced models from scarce and low-quality data. Computers and Mathematics with Applications, 145(??): 224–239, September 1, 2023. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S089812212300264X.

Ullah:2020:RFS

[Ull20]  Malik Zaka Ullah.  An

RBF–FD sparse scheme to simulate high-dimensional Black–Scholes partial differential equations. Computers and Mathematics with Applications, 79(2):426–439, January 15, 2020. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122119303499.

Umbricht:2024:RTE

[UR24] Guillermo Federico Umbricht and Diana Rubio. Regularization techniques for estimating the space-dependent source in an n-dimensional linear parabolic equation using space-dependent noisy data. Computers and Mathematics with Applications, 172 (??):47–69, October 15, 2024. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122124003353.

Utyuzhnikov:2020:UIB

[US20] S. Utyuzhnikov and N. Smirnova. Unsteady interface boundary conditions for nearwall turbulence modeling. Computers and Mathematics with Applications, 79(5): 1483–1502, March 1, 2020. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122119304614.

Ullah:2021:PLS

[UuIUK21] Baseer Ullah, Siraj ul Islam, Zahur Ullah, and Wajid Khan. A parametrized level set based topology optimization method for analysing thermal problems. Computers and Mathematics with Applications, 99(??):99–112, October 1, 2021. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122121002819.

Vabishchevich:2022:NSH

[Vab22] P. N. Vabishchevich. Numerical solution of the heat conduction problem with memory. Computers and Mathematics with Applications, 118 (??):230–236, July 15, 2022. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122122002139.

Vabishchevich:2025:DTS

[Vab25] P. N. Vabishchevich. Decoupling technology for systems of evolutionary equations. Computers and Mathematics with Applications, 191(??): 105–128, August 1, 2025. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122125001750.

Voet:2022:IMA

[VADG22] Yannis Voet, Guillaume Anciaux, Simone Deparis, and Paola Gervasio. The IN-TERNODES method for applications in contact mechanics and dedicated preconditioning techniques. Computers and Mathematics with Applications, 127(??):48–64, December 1, 2022. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122122003996.

Vahabi:2022:ETD

[Vah22] Mohammad Vahabi. The effect of thixotropy on deformation of a single droplet under simple shear flow. Computers and Mathematics with Applications, 117 (??):206–215, July 1, 2022. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122122001791.

Vakilha:2025:FES

[Vak25] Mehran Vakilha. A fully explicit SPH method for modeling 2-D incompressible twophase fluid–structure interaction with modified periodic and open boundary conditions. Computers and Mathematics with Applications, 197(??):43–70, November 1, 2025. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S089812212500327X.

Valseth:2024:AVS

[VBDR24] Eirik Valseth, Pouria Behnoudfar, Clint Dawson, and Albert Romkes. Automatic variationally stable analysis for finite element computations: Transient convectiondiffusion problems. Computers and Mathematics with Applications, 157(??):15– 26, March 1, 2024. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122123005795.

Venegas:2022:KFC

[VGA^{+}22] P. Venegas, D. Gómez, M. Arrinda, M. Oyarbide, H. Macicior, and A. Bermúdez. Kalman filter and classical Preisach hysteresis model applied to the state of charge battery estimation. Computers and Mathematics with Applications, 118 (??):74–84, July 15, 2022. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122122001961.

Visinoni:2024:FTD

[Vis24] Michele Visinoni. A family of three-dimensional virtual elements for Hellinger– Reissner elasticity problems. Computers and Mathematics with Applications, 155(??): 97–109, February 1, 2024. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122123005370.

Vengatesan:2022:NNI

[VJ22] S. Vengatesan and P. V. Jeyakarthikeyan. A new numerical integration over arbitrary quadrilateral element based on smoothed strain energy and Richardson extrapolation. Computers and Mathematics with Applications, 123(??):136– 159, October 1, 2022. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122122003248.

Vlogman:2024:ECL

[VJ24] Tristan G. Vlogman and Kartik Jain. Efficient coupled lattice Boltzmann and Discrete Element Method simulations of small particles in complex geometries. Computers and Mathematics with Applications, 175(??): 313–329, December 1, 2024. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122124004474.

Vogl:2023:MRA

[VJH23] Christopher J. Vogl, Ilon Joseph, and Milan Holec. Mesh refinement for anisotropic diffusion in magnetized plasmas. Computers and Mathematics with Applications, 145(??):159–174, September 1, 2023. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122123002717.

Vienne:2021:RFD

[VL21] Lucien Vienne and Emmanuel Lévêque. Recursive finite-difference lattice Boltzmann schemes. Computers and Mathematics with Applications, 96(??):95–108, August 15, 2021. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122121001966.

Vodstrcil:2024:FBS

[VLD^{+}24] Petr Vodstrčil, Dalibor Lukáš, Zdeněk Dostál, Marie Sadowská, David Horák, Oldřich Vlach, Jiří Bouchala, and Jakub Kružík. On favorable bounds on the spectrum of discretized Steklov– Poincaré operator and applications to domain decomposition methods in 2D. Computers and Mathematics with Applications, 167 (??):12–20, August 1, 2024. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122124001986.

Valseth:2020:GOE

[VR20] Eirik Valseth and Albert Romkes. Goal-oriented error estimation for the automatic variationally stable FE method for convectiondominated diffusion problems. Computers and Mathematics with Applications, 80(12):3027–3043, December 15, 2020. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122120304211.

Vrionis:2021:TOF

[VSG21] P. Y. Vrionis, K. D. Samouchos, and K. C. Giannakoglou. Topology optimization in fluid mechanics using continuous adjoint and the cut-cell method. Computers and Mathematics with Applications, 97(??): 286–297, September 1, 2021. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122121002406.

Vasilyeva:2021:MLA

[VT21] Maria Vasilyeva and Aleksey Tyrylgin. Machine learning for accelerating macroscopic parameters prediction for poroelasticity problem in stochastic media. Computers and Mathematics with Applications, 84(??): 185–202, February 15, 2021.

CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122121000067.

Wout:2022:BIF

[vtWHG^{+}22] Elwin van ’t Wout, Seyyed R.

Haqshenas, Pierre Gélat, Timo Betcke, and Nader Saffari. Boundary integral formulations for acoustic modelling of high-contrast media. Computers and Mathematics with Applications, 105(??): 136–149, January 1, 2022. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122121004259.

Wang:2024:TGM

[Wan24]  Keyan Wang.  Two-grid

methods for nonlinear pseudoparabolic integro-differential equations by finite element method. Computers and Mathematics with Applications, 168(??):174–189, August 15, 2024. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S089812212400258X.

Wang:2025:HOD

[Wan25] Qiu-Ya Wang. High order difference schemes for nonlinear Riesz space variableorder fractional diffusion equations. Computers and Mathematics with Applications, 188(??):221–243, June 15, 2025. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122125001555.

Wu:2023:CRO

[WARX23] Liang Wu, Mejdi Azaïez, Tomás Chacón Rebollo, and Chuanju Xu. Certified reduced order method for the parametrized Allen– Cahn equation. Computers and Mathematics with Applications, 134(??):167– 180, March 15, 2023. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122123000068.

Wathen:2022:SOP

[Wat22] Andy Wathen. Some observations on preconditioning for non-self-adjoint and time-dependent problems. Computers and Mathematics with Applications, 116 (??):176–180, June 15, 2022. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122121002388.

Wazwaz:2020:DTD

[Waz20] Abdul-Majid Wazwaz. A (2 + 1)-dimensional timedependent Date–Jimbo–Kashiwara– Miwa equation: Painlevé integrability and multiple soliton solutions. Computers and Mathematics with Applications, 79(4):1145–1149, February 15, 2020. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122119304286.

Wang:2023:UES

[WaZWZ23] Meiting Wang, Guang an Zou, Bo Wang, and Wenju Zhao. Unconditionally energy-stable discontinuous Galerkin method for the chemo-repulsion-Navier– Stokes system. Computers and Mathematics with Applications, 150(??):132– 155, November 15, 2023. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122123003978.

Wang:2023:PEE

[WBY23] Shixi Wang, Hai Bi, and Yidu Yang. The a posteriori error estimates and an adaptive algorithm of the FEM for transmission eigenvalues for anisotropic media. Computers and Mathematics with Applications, 150(??): 156–169, November 15, 2023. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122123004030.

Wang:2021:PEE

[WC21] Yue Wang and Hu Chen. Pointwise error estimate of an alternating direction implicit difference scheme for two-dimensional timefractional diffusion equation. Computers and Mathematics with Applications, 99(??): 155–161, October 1, 2021. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122121002960.

Wang:2025:HPN

[WCC25] Jing Wang, Xuejuan Chen, and Jinghua Chen. A highprecision numerical method based on spectral deferred correction for solving the time-fractional Allen–Cahn equation. Computers and Mathematics with Applications, 180(??):1–27, February 15, 2025. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122124005376.

Wang:2025:MNE

[WCCC25] Nan Wang, Hanyu Chu, Jinru Chen, and Ying Cai. Mixed Nitsche extended finite element method for solving three-dimensional H(curl)-elliptic interface problems. Computers and Mathematics with Applications, 199(??):22–44, December 1, 2025. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122125003645.

Wang:2020:TGE

[WCH20] Yang Wang, Yanping Chen, and Yunqing Huang. A twogrid Eulerian–Lagrangian localized adjoint method to miscible displacement problems with dispersion term. Computers and Mathematics with Applications, 80 (4):54–68, August 15, 2020. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122120301450.

Wen:2023:AGD

[WCH23] Jing Wen, Zhangxing Chen, and Yinnian He. The analogue of grad-div stabilization in DG method for a coupled Stokes and Darcy problem. Computers and Mathematics with Applications, 149(??):99–114, November 1, 2023. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122123003747.

Wang:2020:EMS

[WCL20] Dan Wang, C. S. Chen, and Wen Li. An efficient MAPS for solving fourth order partial differential equations using trigonometric functions. Computers and Mathematics with Applications, 79(4): 934–946, February 15, 2020. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122119303955.

Wang:2025:XDS

[WCLQ25] Tao Wang, Yanping Chen, Xiangquan Li, and Fangfang Qin. XFEM for the diffusion and sub-diffusion problems in a non-convex domain. Computers and Mathematics with Applications, 196(??): 415–432, October 15, 2025. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122125003608.

Wang:2024:LEF

[WCS24] Chao Wang, Jintao Cui, and Zhengjia Sun. A linear edge finite element method for quad-curl problem. Computers and Mathematics with Applications, 176(??):99– 108, December 15, 2024. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122124004267.

Wang:2020:SFP

[WCW20a] Hanquan Wang, Ronghua Cheng, and Xinming Wu. A splitting Fourier pseudospectral method for Vlasov– Poisson–Fokker–Planck system. Computers and Mathematics with Applications, 79(6):1742–1758, March 15, 2020. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122119304845.

Wu:2020:TDA

[WCW20b] Chih-Ping Wu, Yen-Jung Chen, and Yung-Ming Wang. Three-dimensional asymptotic nonlocal elasticity theory for the free vibration analysis of embedded singlewalled carbon nanotubes. Computers and Mathematics with Applications, 80 (1):161–182, July 1, 2020. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122120301048.

Wang:2022:MVE

[WCWZ22] Feng Wang, Mingchao Cai, Gang Wang, and Yuping Zeng. A mixed virtual element method for Biot’s consolidation model. Computers and Mathematics with Applications, 126(??):31–42, November 15, 2022. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122122003789.

Wang:2023:HOA

[WCWZ23] Jingying Wang, Chen Cui, Zhifeng Weng, and Shuying Zhai. A high order accurate numerical algorithm for the space-fractional Swift– Hohenberg equation. Computers and Mathematics with Applications, 139(??):216– 223, June 1, 2023. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122122003947.

Wu:2024:DHM

[WCY^{+}24] Yao Wu, Zhenhua Chai, Xiaolei Yuan, Xiuya Guo, and Baochang Shi. A discrete Hermite moments based multiple-relaxation-time lattice Boltzmann model for convection-diffusion equations. Computers and Mathematics with Applications, 156(??):218–238, February 15, 2024. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122124000099.

Wang:2024:ECI

[WCZ24] Yong Wang, Jun Cao, and Chengwen Zhong. Enhancement of coupled immersed boundary–finite volume lattice Boltzmann method (IB–FVLBM) using least– square aided “ghost–cell” techniques. Computers and Mathematics with Applications, 159(??):219–239, April 1, 2024. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122124000476.

Wu:2020:SVI

[WCZD20] Jiayang Wu, Yongguang Cheng, Chunze Zhang, and Wei Diao. Simulating vortex induced vibration of an impulsively started flexible filament by an implicit IB–LB coupling scheme. Computers and Mathematics with Applications, 79 (1):159–173, January 2020. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122117305497.

Wang:2020:MET

[WDGC20] Hui Wang, Jinqiao Duan, Xianguo Geng, and Ying Chao. Mean escape time for randomly switching narrow gates in a steady flow. Computers and Mathematics with Applications, 79(10): 2795–2804, May 15, 2020. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122119305693.

Wang:2023:AFD

[WDGS23] Zhuo Wang, Lin Du, Feng Gao, and Xiaofeng Sun. Adaptive forcing distance in an immersed boundary method for internal flow simulation at high Reynolds numbers. Computers and Mathematics with Applications, 140(??):44– 63, June 15, 2023. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122123001153.

Wang:2024:SHO

[WDYZ24] Jihong Wang, Yana Di, Jerry Zhijian Yang, and Jiwei Zhang. A stable and highorder numerical scheme with discrete DtN-type artificial boundary conditions for a 2D peridynamic diffusion model. Computers and Mathematics with Applications, 158 (??):1–12, March 15, 2024. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122124000063.

Wang:2021:NLP

[WDZ21] Xinhui Wang, Guangzhi Du, and Liyun Zuo. A novel local and parallel finite element method for the mixed Navier–Stokes–Darcy problem. Computers and Mathematics with Applications, 90 (??):73–79, May 15, 2021. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122121000948.

Wang:2024:EMG

[WF24] Lu Wang and Minfu Feng.

On the enriched mixed Galerkin method with θ scheme for the elastic wave equation. Computers and Mathematics with Applications, 157(??):124–146, March 1, 2024. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122123005783.

Wang:2025:SWG

[WF25] Lu Wang and Minfu Feng.

The simplified weak Galerkin method with θ scheme and its reduced-order model for the elastodynamic problem on polygonal mesh. Computers and Mathematics with Applications, 178(??):19–46, January 15, 2025. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122124005200.

Wang:2020:DPG

[WFH^{+}20] Nan Wang, Mingfa Fei, Chengming Huang, Guoyu Zhang, and Meng Li. Dissipationpreserving Galerkin–Legendre spectral methods for twodimensional fractional nonlinear wave equations. Computers and Mathematics with Applications, 80(5):617–635, September 1, 2020. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122120301607.

Williams:2020:FSO

[WFMD20] David M. Williams, Cory V. Frontin, Edward A. Miller, and David L. Darmofal. A family of symmetric, optimized quadrature rules for pentatopes. Computers and Mathematics with Applications, 80(5):1405–1420, September 1, 2020. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122120302704.

Wang:2024:TOS

[WFW^{+}24] Chunxiu Wang, Lixue Fang, Xuan Wang, Huanlin Zhou, and Xin Guo. Topology optimization of steady Navier– Stokes flow using moving morphable void method. Computers and Mathematics with Applications, 161 (??):212–224, May 1, 2024. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S089812212400110X.

Wang:2024:ASP

[WFX24] Jie Wang, Xinlong Feng, and Hui Xu. Adaptive sampling points based multi-scale residual network for solving partial differential equations. Computers and Mathematics with Applications, 169(??): 223–236, September 1, 2024. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122124002979.

Wang:2023:SLO

[WGC23] Yue Wang, Fuzheng Gao, and Jintao Cui. Superconvergence of the lowestorder weak Galerkin finite element method for elliptic interface problems. Computers and Mathematics with Applications, 149(??):203–210, November 1, 2023. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S089812212300408X.

Wang:2024:NSN

[WGeL24] Shuaikang Wang, Yongbin Ge, and Sheng en Liu. Numerical solutions of the nonlinear wave equations with energy-preserving sixthorder finite difference schemes. Computers and Mathematics with Applications, 168(??):

100–119, August 15, 2024. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122124002475.

Wen:2020:SCF

[WH20] Jing Wen and Yinnian He. A strongly conservative finite element method for the coupled Stokes– Biot model. Computers and Mathematics with Applications, 80(5):1421–1442, September 1, 2020. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122120302674.

Wang:2021:FES

[WH21] Jun-Ya Wang and Qiong-Ao Huang. A family of effective structure-preserving schemes with second-order accuracy for the undamped sine–Gordon equation. Computers and Mathematics with Applications, 90(??):38– 45, May 15, 2021. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122121000778.

Wang:2025:PSB

[WH25] Yu Wang and Chuanjiang He. A PDE system based on background estimation for document binarization with non-uniform background. Computers and Mathematics with Applications, 194(??):158–176, September 15, 2025. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122125002561.

Wells:2024:VBM

[WHC24] H. Wells, M. E. Hubbard, and A. Cangiani. A velocitybased moving mesh virtual element method. Computers and Mathematics with Applications, 155(??):110– 125, February 1, 2024. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122123005552.

Wen:2022:SIU

[WHH22] Juan Wen, Yinnian He, and Ya-Ling He. Semiimplicit, unconditionally energy stable, stabilized finite element method based on multiscale enrichment for the Cahn–Hilliard–Navier– Stokes phase-field model. Computers and Mathematics with Applications, 126(??): 172–181, November 15, 2022. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S089812212200390X.

Wang:2022:LTO

[WHT22] Wansheng Wang, Yi Huang, and Jiao Tang. Lie– Trotter operator splitting spectral method for linear semiclassical fractional Schrödinger equation. Computers and Mathematics with Applications, 113(??):117– 129, May 1, 2022. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122122001018.

Wieners:2023:STD

[Wie23] Christian Wieners. A spacetime discontinuous Galerkin discretization for the linear transport equation. Computers and Mathematics with Applications, 152(??):294– 307, December 15, 2023. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122123004789.

Wang:2025:SDG

[WJ25] Deng Wang and Zupeng Jia. A staggered discontinuous Galerkin method for solving S_{N} transport equation on arbitrary polygonal grids. Computers and Mathematics with Applications, 182(??): 102–121, March 15, 2025. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122125000252.

Wang:2025:FMS

[WJC^{+}25] Lixiu Wang, Lueling Jia, Zijian Cao, Huiyuan Li, and Zhimin Zhang. Fast Maxwell solvers based on exact discrete eigen-decompositions I. Two-dimensional case. Computers and Mathematics with Applications, 193(??):346– 363, September 1, 2025. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122125002688.

Wang:2024:UWD

[WJZ^{+}24] Haijin Wang, Lulu Jiang, Qiang Zhang, Yuan Xu, and Xiaobin Shi. Ultraweak discontinuous Galerkin method with IMEX-BDF time marching for two dimensional convection–diffusion problems. Computers and Mathematics with Applications, 176(??):77–90, December 15, 2024. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122124004188.

Wu:2021:HOE

[WKC21] Xinhui Wu, Ethan J. Kubatko, and Jesse Chan. Highorder entropy stable discontinuous Galerkin methods for the shallow water equations: Curved triangular meshes and GPU acceleration. Computers and Mathematics with Applications, 82(??):179–199, January 15, 2021. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122120304314.

Wiegold:2021:NSV

[WKGH21] T. Wiegold, S. Klinge, R. P. Gilbert, and G. A. Holzapfel. Numerical simulation of the viral entry into a cell driven by receptor diffusion. Computers and Mathematics with Applications, 84(??):224– 243, February 15, 2021. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S089812212030482X.

Wu:2020:SFO

[WL20] Dong-Lun Wu and Fengying Li. Solutions for fourthorder Kirchhoff type elliptic equations involving concaveconvex nonlinearities in RN. Computers and Mathematics with Applications, 79(2): 489–499, January 15, 2020. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122119303451.

Wang:2022:SRN

[WL22] Junjun Wang and Meng Li. Superconvergence results for nonlinear Klein–Gordon– Schrödinger equation with backward differential formula finite element method. Computers and Mathematics with Applications, 118 (??):214–229, July 15, 2022. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122122001985.

Wang:2023:AST

[WL23] Yating Wang and Wing Tat Leung. An adaptive space and time method in partially explicit splitting scheme for multiscale flow problems. Computers and Mathematics with Applications, 144(??): 100–123, August 15, 2023. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122123002444.

Wang:2025:SEN

[WL25] Junjie Wang and Hongmin Li. Some effective numerical schemes for solving Klein–Gordon–Schrödinger system in 2D based on IEQ method. Computers and Mathematics with Applications, 194(??):231–256, September 15, 2025. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122125002615.

Wang:2024:TNL

[WLC24] Lele Wang, Xin Liao, and Can Chen. Two novel linearized energy-conserving finite element schemes for nonlinear regularized long wave equation. Computers and Mathematics with Applications, 174(??):361– 378, November 15, 2024. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122124004395.

Wen:2022:NSS

[WLD22] Weibin Wen, Tianhao Liu, and Shengyu Duan. A novel sub-step explicit time integration method based on cubic B-spline interpolation for linear and nonlinear dynamics. Computers and Mathematics with Applications, 127(??):154–180, December 1, 2022. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122122004175.

Wu:2024:FIM

[WLH24] Yawen Wu, Liming Ling, and Yubin Huang. A failureinformed multi-stage training algorithm for threecomponent nonlinear Schrödinger equation. Computers and Mathematics with Applications, 169(??):132–141, September 1, 2024. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122124002803.

Wang:2021:SAM

[WLJ21] Junjun Wang, Meng Li, and Mengping Jiang. Superconvergence analysis of a MFEM for BBM equation with a stable scheme. Computers and Mathematics with Applications, 93 (??):168–177, July 1, 2021. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122121001565.

Wu:2025:LCA

[WLJF25] Jilian Wu, Ning Li, Mengru Jiang, and Xinlong Feng. Low-cost adaptive pressure-correction projection method for the 2D/3D time-dependent natural convection problems. Computers and Mathematics with Applications, 194(??):86–109, September 15, 2025. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122125002536.

Wang:2023:UCS

[WLL23] Dan Wang, Meng Li, and Yu Lu. Unconditionally convergent and superconvergent analysis of secondorder weighted IMEX FEMs for nonlinear Ginzburg– Landau equation. Computers and Mathematics with Applications, 146(??):84–105, September 15, 2023. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122123002857.

Wu:2022:PDD

[WLQ22] Dan Wu, Junliang Lv, and Hao Qian. Parallel domain decomposition schemes based on finite volume element discretization for nonsteady-state diffusion equations on distorted meshes. Computers and Mathematics with Applications, 112(??):97– 115, April 15, 2022. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122122000840.

Wei:2022:GSA

[WLW^{+}22] Yabing Wei, Shujuan Lü, Fenling Wang, F. Liu, and Yanmin Zhao. Global superconvergence analysis of nonconforming finite element method for time fractional reaction-diffusion problem with anisotropic data. Computers and Mathematics with Applications, 119(??):159– 173, August 1, 2022. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122122002553.

Wang:2025:NAL

[WLW^{+}25] Kai Wang, Yubo Li, Anping Wu, Xiaoxin Yin, Feng Liu, and Xinlong Feng. Numerical analysis of a linear generalized finite difference scheme for Burgers’ equation. Computers and Mathematics with Applications, 193(??):117– 131, September 1, 2025. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122125002378.

Wang:2020:ADG

[WLX20] Donghuan Wang, Yeda Lian, and Hong Xiao. Application of discontinuous Galerkin method in supersonic and hypersonic gas flows. Computers and Mathematics with Applications, 80 (1):227–246, July 1, 2020. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122120301176.

Wang:2022:DGM

[WLX22] Tao Wang, Binjie Li, and Xiaoping Xie. Discontinuous Galerkin method for a distributed optimal control problem governed by a time fractional diffusion equation. Computers and Mathematics with Applications, 128(??): 1–11, December 15, 2022. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122122004023.

Wang:2022:INS

[WLZZ22] Yong Wang, Sha Liu, Congshan Zhuo, and Chengwen Zhong. Investigation of nonlinear squeezefilm damping involving rarefied gas effect in microelectro-mechanical systems. Computers and Mathematics with Applications, 114 (??):188–209, May 15, 2022. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122122001420.

Wang:2023:RRU

[WM23] Xu Wang and Weiyin Ma. Rational reparameterization of unstructured quadrilateral meshes for isogeometric analysis with optimal convergence. Computers and Mathematics with Applications, 151(??):304–325, December 1, 2023. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122123004418.

Wang:2021:FVE

[WML21] Yue Wang, Xiangyun Meng, and Yonghai Li. The finite volume element method on the Shishkin mesh for a singularly perturbed reaction– diffusion problem. Computers and Mathematics with Applications, 84(??):112– 127, February 15, 2021. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122120304739.

Wang:2025:MCC

[WMMH25] Xiaorong Wang, Xuerui Mao, Shipeng Mao, and Xiaoming He. A mass and charge conservative fully discrete scheme for a 3D diffuse interface model of the two-phase inductionless MHD flows. Computers and Mathematics with Applications, 182(??):139– 162, March 15, 2025. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122125000264.

Williams:2024:CFE

[WN24] David M. Williams and Nilima Nigam. Conforming finite element function spaces in four dimensions, part II: the pentatope and tetrahedral prism. Computers and Mathematics with Applications, 167(??):21–53, August 1, 2024. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122124002086.

Wang:2025:AMP

[WNF^{+}25] Lin Wang, Sundararajan Natarajan, Weihua Fang, Zhanfei Si, and Tiantang Yu. Adaptive 3D multi-patch isogeometric analysis with truncated hierarchical NURBS for complex elasticity. Computers and Mathematics with Applications, 192(??):104– 133, August 15, 2025. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122125002081.

Wolfmayr:2020:GLB

[Wol20] Monika Wolfmayr. Guaranteed lower bounds for cost functionals of timeperiodic parabolic optimization problems. Computers and Mathematics with Applications, 80(5):1050–1072, September 1, 2020. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122120301681.

Wang:2025:FMC

[WP25] Weizhu Wang and Stefanos-Aldo Papanicolopulos. Finding minimal cubature rules for finite elements passing the patch test. Computers and Mathematics with Applications, 179(??):1–16, February 1, 2025. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122124005273.

Wu:2022:LIS

[WPLW22] Pinxia Wu, Kejia Pan, Weiwei Ling, and Qihong Wu. A linearly implicit scheme and fast multigrid solver for 3D Fitzhugh–Nagumo equation. Computers and Mathematics with Applications, 117 (??):257–270, July 1, 2022. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122122001821.

Wang:2024:RTD

[WQ24] Yan Wang and Zhi Qian.

Regularizing a two-dimensional time-fractional inverse heat conduction problem by a fractional Landweber iteration method. Computers and Mathematics with Applications, 164(??):104– 115, June 15, 2024. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122124001524.

Wang:2025:TDG

[WQS25] Songhao Wang, Zhenghua Qian, and Yan Shang. Threedimensional geometrically nonlinear analysis of functionally graded microshell structures using corotational finite element method based on modified couple stress theory. Computers and Mathematics with Applications, 190(??):122– 153, July 15, 2025. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122125001774.

Wang:2023:SDM

[WR23] Xue Wang and Hongxing Rui. A semi-decoupled MAC scheme for the coupled fluidporoelastic material interaction. Computers and Mathematics with Applications, 139 (??):118–135, June 1, 2023. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122123001384.

Wittschieber:2023:MBA

[WRMB23] Stefan Wittschieber, Ajay Rangarajan, Georg May, and Marek Behr. Metric-based anisotropic mesh adaptation for viscoelastic flows.

Computers and Mathematics with Applications, 151(??): 67–79, December 1, 2023. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122123004170.

Wei:2020:SAT

[WS20a] Yifan Wei and Dongyang Shi. Superconvergence analysis of a two-grid method for nonlinear hyperbolic equations. Computers and Mathematics with Applications, 79(10):2846–2855, May 15, 2020. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122119305875.

Wen:2020:SMS

[WS20b] Jiaqiang Wen and Yufeng Shi. Symmetrical martingale solutions of backward doubly stochastic Volterra integral equations. Computers and Mathematics with Applications, 79(5):1435– 1446, March 1, 2020. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122119304535.

Wang:2021:ARF

[WS21a] Lin Wang and Martin Stynes. An α-robust finite difference method for a time-fractional radially symmetric diffusion problem. Computers and Mathematics with Applications, 97(??): 386–393, September 1, 2021. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122121002480.

Winter:2021:NSF

[WS21b] Ondřej Winter and Petr Sváček. On numerical simulation of flexibly supported airfoil in interaction with incompressible fluid flow using laminar–turbulence transition model. Computers and Mathematics with Applications, 83(??):57–73, February 1, 2021. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122119305966.

Wu:2025:QUU

[WS25] Yanmi Wu and Dongyang Shi. Quasi-uniform unconditional superconvergent error estimates of FEMs for the time-dependent singularly perturbed Bi-wave problem modeling d-wave superconductors. Computers and Mathematics with Applications, 191(??):24–35, August 1, 2025. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122125001646.

Wang:2024:SNE

[WSCP24] Junjun Wang, Dongyang Shi, Lina Cao, and Jiaxuan Pei. Superconvergence of a new energy dissipation finite element scheme for nonlinear Schrödinger equation with wave operator. Computers and Mathematics with Applications, 161(??):202– 211, May 1, 2024. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122124001093.

Wang:2021:NSF

[WSGT21] Le Wang, Junwei Su, Zhaolin Gu, and Liyu Tang. Numerical study on flow field and pollutant dispersion in an ideal street canyon within a real tree model at different wind velocities. Computers and Mathematics with Applications, 81(??):679–692, January 1, 2021. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122119306005.

Wodecki:2023:NOD

[WSO^{+}23] Aleš Wodecki, Pavel Strachota, Tomáš Oberhuber, Kateřina Škardová, Monika Balázsová, and Michal Bohatý. Numerical optimization of the Dirichlet boundary condition in the phase field model with an application to pure substance solidification. Computers and Mathematics with Applications, 145(??):90–105, September 1, 2023. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122123002584.

Wu:2022:NFD

[WSS22] Shuhao Wu, Yongli Song, and Qingyan Shi. Normal forms of double Hopf bifurcation for a reaction– diffusion system with delay and nonlocal spatial average and applications. Computers and Mathematics with Applications, 119(??):174– 192, August 1, 2022. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122122002541.

Wang:2025:BBN

[WT25] Xiaofang Wang and Hao Tian. A bond-based nonlocal anisotropic diffusion model with variable matrix-valued coefficients and its asymptotically compatible meshfree discretization. Computers and Mathematics with Applications, 199(??):148–166, December 1, 2025. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122125004018.

Wu:2020:RWD

[WTQ^{+}20] Xiao-Yu Wu, Bo Tian, Qi-Xing Qu, Yu-Qiang Yuan, and Xia-Xia Du. Rogue waves for a (2+1)-dimensional Gross–Pitaevskii equation with time-varying trapping potential in the Bose– Einstein condensate. Computers and Mathematics with Applications, 79(4):1023– 1030, February 15, 2020. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122119304122.

Wang:2020:LML

[WTSZ20] Meng Wang, Bo Tian, Yan Sun, and Ze Zhang. Lump, mixed lump-stripe and rogue wave-stripe solutions of a (3 + 1)-dimensional nonlinear wave equation for a liquid with gas bubbles. Computers and Mathematics with Applications, 79(3):576–587, February 1, 2020. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S089812211930344X.

Wang:2024:DLM

[WTSZ24] Sicong Wang, Bin Teng, Yufeng Shi, and Qingfeng Zhu. A deep learning method for solving multi-dimensional coupled forward–backward doubly SDEs. Computers and Mathematics with Applications, 169(??):260–272, September 1, 2024. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122124003158.

Wang:2020:PDW

[WW20a] Chunmei Wang and Junping Wang. Primal-dual weak Galerkin finite element methods for elliptic Cauchy problems. Computers and Mathematics with Applications, 79(3):746–763, February 1, 2020. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122119303839.

Wang:2020:LSV

[WW20b] Ying Wang and Gang Wang. A least-squares virtual element method for secondorder elliptic problems. Computers and Mathematics with Applications, 80(8):1873– 1886, October 15, 2020. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S089812212030331X.

Wang:2024:DGM

[WW24] Jindong Wang and Shuonan Wu. Discontinuous Galerkin methods for magnetic advection-diffusion problems. Computers and Mathematics with Applications, 174(??):43–54, November 15, 2024. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S089812212400381X.

Wu:2025:CAF

[WW25] Peixuan Wu and Yang Wang.

Convergence analysis of a fast virtual element method for coupled Darcy flows in fractured porous media on general polygonal meshes. Computers and Mathematics with Applications, 198(??): 1–23, November 15, 2025. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122125003293.

Wang:2024:FPP

[WWC^{+}24] Yujie Wang, Shengquan Wang, Yong Cai, Guidong Wang, and Guangyao Li. Fully parallel and pipelined sparse direct solver for large symmetric indefinite finite element problems. Computers and Mathematics with Applications, 175(??):447–469, December 1, 2024. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122124004589.

Wei:2023:AVE

[WWCJ23] Huayi Wei, Xin Wang, Chunyu Chen, and Kai Jiang. An adaptive virtual element method for the polymeric self-consistent field theory. Computers and Mathematics with Applications, 141(??):242– 254, July 1, 2023. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S089812212300072X.

Wang:2025:MDA

[WWG^{+}25] Jian Wang, Yu Wang, Shanshan Ge, Ziwei Han, and Maodong Pan. Morphological dynamics analysis on 3D surface using the Gray– Scott model. Computers and Mathematics with Applications, 199(??):242–259, December 1, 2025. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122125004043.

Wu:2023:NET

[WWH23] Shang Wu, Xue Wang, and Jianhua Huang. Numerical ergodicity of two dimensional stochastic Navier– Stokes equations with Gaussian noise. Computers and Mathematics with Applications, 141(??):15–32, July 1, 2023. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122123001244.

Wang:2022:RWG

[WWL22] Zhuoran Wang, Ruishu Wang, and Jiangguo Liu. Robust weak Galerkin finite element solvers for Stokes flow based on a lifting operator. Computers and Mathematics with Applications, 125(??): 90–100, November 1, 2022. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122122003698.

Wang:2024:ILC

[WWLG24] Yingying Wang, Wei Wu, Xuyang Lou, and Daniel Görges. Iterative learning control of an Euler–Bernoulli beam with time-varying boundary disturbance. Computers and Mathematics with Applications, 162(??):145– 154, May 15, 2024. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122124001068.

Wang:2025:AFE

[WWP^{+}25] Shuai Wang, Xinnan Wu, Cheng Peng, Songying Chen, and Hao Liu. Analysis on the force evaluation by the momentum exchange method and a localized refilling scheme for the lattice Boltzmann method. Computers and Mathematics with Applications, 183(??):180– 199, April 1, 2025. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122125000343.

Wang:2020:SDH

[WWQ20] Zhengdao Wang, Yikun Wei, and Yuehong Qian. A simple direct heating thermal immersed boundarylattice Boltzmann method for its application in incompressible flow. Computers and Mathematics with Applications, 80(6):1633–1649, September 15, 2020. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122120303096.

Wang:2023:PRV

[WWS23] Ying Wang, Gang Wang, and Yue Shen. A pressure-robust virtual element method for the Navier–Stokes problem on polygonal mesh. Computers and Mathematics with Applications, 131(??):124– 137, February 1, 2023. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S089812212200520X.

Wang:2021:AVE

[WWW21] Ying Wang, Gang Wang, and Feng Wang. An adaptive virtual element method for incompressible flow. Computers and Mathematics with Applications, 101(??):63–73, November 1, 2021. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122121003394.

Wang:2025:ADA

[WWY25] Wei Wang, Tiancheng Wang, and Wei Yang. Adaptive design and adaptive finite element verification of bifunctional layer-wise cloak metamaterials for thermal and electrical insulation. Computers and Mathematics with Applications, 197(??):183– 199, November 1, 2025. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122125003499.

Wang:2025:FFO

[WWYL25] Shenpei Wang, Tao Wang, Shuyu Yue, and Hairong Lian. A fast fourth-order scheme and its extrapolations for two-dimensional space fractional diffusion equations. Computers and Mathematics with Applications, 199(??):260–272, December 1, 2025. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122125004080.

Wang:2024:OSS

[WWZ^{+}24] Weiping Wang, Qingshan Wang, Rui Zhong, Xianjie Shi, and Liming Chen. Optimization of stacking sequence for quadrilateral laminated composite plates with curved edges based on kriging. Computers and Mathematics with Applications, 159 (??):142–154, April 1, 2024. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122124000403.

Wang:2025:SOM

[WWZQ25] Yan Wang, Haifeng Wang, Hong Zhang, and Xu Qian. A second-order maximum bound principle-preserving exponential Runge–Kutta scheme for the convective Allen–Cahn equation. Computers and Mathematics with Applications, 193(??):297– 314, September 1, 2025. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122125002755.

Wang:2025:RDM

[WWZX25] Hailong Wang, Liang Wang, Guoqing Zhu, and Chunguang Xiong. Rot-div mixed finite element method of two dimensional Hodge Laplacian problem. Computers and Mathematics with Applications, 184(??):134– 152, April 15, 2025. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122125000586.

Wang:2022:HAW

[WX22] An-Yang Wang and Hang Xu. Highly accurate wavelet-homotopy solutions for mixed convection hybrid nanofluid flow in an inclined square lid-driven cavity. Computers and Mathematics with Applications, 108(??): 88–108, February 15, 2022. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122122000049.

Wang:2025:NAS

[WXCL25] Wensi Wang, Hailing Xuan, Xiaoliang Cheng, and Kewei Liang. Numerical analysis and simulation of a quasistatic frictional bilateral contact problem with damage, long-term memory and wear. Computers and Mathematics with Applications, 177(??):130– 146, January 1, 2025. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122124005182.

Wang:2022:EMB

[WXF22] Yan Wang, Xufeng Xiao, and Xinlong Feng. An efficient maximum bound principle preserving p-adaptive operator-splitting method for three-dimensional phase field shape transformation model. Computers and Mathematics with Applications, 120(??):78–91, August 15, 2022. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122122002607.

Wang:2025:IGS

[WXH^{+}25] Ziping Wang, Guangqing Xia, Yajie Han, Chang Lu, Lin Zhang, and Gang Xu. An innovative Galerkin scheme based on anisotropic trilinear immersed finite elements for the magnetized plasma diffusion problem with plasma sheath interface. Computers and Mathematics with Applications, 192(??):155– 171, August 15, 2025. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122125002214.

Wang:2021:BIF

[WXZW21] Quanxiang Wang, Jianqiang Xie, Zhiyue Zhang, and Liqun Wang. Bilinear immersed finite volume element method for solving matrix coefficient elliptic interface problems with nonhomogeneous jump conditions. Computers and Mathematics with Applications, 86 (??):1–15, March 15, 2021. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122121000018.

Wang:2020:AFE

[WYH20] Hao Wang, Wei Yang, and Yunqing Huang. Adaptive finite element method for the sound wave problems in two kinds of media. Computers and Mathematics with Applications, 79(3): 789–801, February 1, 2020. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122119303815.

Wang:2024:BAO

[WYHQ24] Ruru Wang, Yubin Yan, A. S. Hendy, and Leijie Qiao. BDF2 ADI orthogonal spline collocation method for the fractional integro-differential equations of parabolic type in three dimensions. Computers and Mathematics with Applications, 155(??):126– 141, February 1, 2024. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122123005576.

Wang:2025:MSE

[WYLL25] Jiarui Wang, Yining Yang, Hong Li, and Yang Liu. Mixed spectral element method combined with secondorder time stepping schemes for a two-dimensional nonlinear fourth-order fractional diffusion equation. Computers and Mathematics with Applications, 188(??): 1–18, June 15, 2025. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122125001105.

Wang:2023:HET

[WYP23] Shuman Wang, Junxiang Yang, and Xiaomin Pan. Highly efficient time-marching method with enhanced energy consistency for the L^{2}- gradient flow based twophase incompressible fluid system. Computers and Mathematics with Applications, 139(??):68–90, June 1, 2023. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.

com/science/article/pii/
S0898122123001037.

Wang:2020:OWT

[WYS^{+}20] Shuoliang Wang, Chunlei Yu, Guoqiang Sang, Rongze Yu, and Feng Cheng. An oilwater two-phase reservoir numerical simulation coupled with dynamic capillary force based on the fullimplicit method. Computers and Mathematics with Applications, 79(9):2527– 2549, May 1, 2020. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122119305462.

Wu:2023:HEV

[WYT23] Jingwen Wu, Junxiang Yang, and Zhijun Tan. Highly efficient variant of SAV approach for the incompressible multi-component phasefield fluid models. Computers and Mathematics with Applications, 145(??):24–40, September 1, 2023. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122123002560.

Wang:2024:UAL

[WYW^{+}24] Yan Wang, Yining Yang, Jinfeng Wang, Hong Li, and Yang Liu. Unconditional analysis of the linearized second-order time-stepping scheme combined with a mixed element method for a nonlinear time fractional fourth-order wave equation. Computers and Mathematics with Applications, 157 (??):74–91, March 1, 2024. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122123005801.

Wang:2021:PSM

[WZ21] Yongshuai Wang and Haibiao Zheng. A partitioned scheme with multipletime-step technique for the nonstationary dual-porosity-Stokes problem. Computers and Mathematics with Applications, 93(??):265– 288, July 1, 2021. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122121001590.

Wang:2022:SOC

[WZ22] Yuan-Ming Wang and Zi-Yun Zheng. A secondorder L2-1_{σ} Crank–Nicolson difference method for twodimensional time-fractional wave equations with variable coefficients. Computers and Mathematics with Applications, 118(??):183– 207, July 15, 2022. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122122002127.

Wang:2024:TAN

[WZ24] Yiyang Wang and Zhongguo Zhou. Theoretical analysis and numerical scheme of local conservative characteristic finite difference for 2-d advection diffusion equations. Computers and Mathematics with Applications, 175(??): 255–275, December 1, 2024. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122124004413.

Wang:2025:TGA

[WZ25] Jianyun Wang and Zixin Zhong. Two-grid algorithm of the BDF2 finite element method for the two-dimensional linear Schrödinger equation. Computers and Mathematics with Applications, 189(??):194– 207, July 1, 2025. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122125001592.

Wang:2025:NHS

[WZA25] Wei Wang, Jihui Zheng, and Jing An. A novel Hermite spectral element method for elliptic interface eigenvalue problems. Computers and Mathematics with Applications, 198(??):293– 320, November 15, 2025. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122125004109.

Wang:2020:SFVa

[WZC^{+}20a] Yong Wang, Chengwen Zhong, Jun Cao, Congshan Zhuo, and Sha Liu. A simplified finite volume lattice Boltzmann method for simulations of fluid flows from laminar to turbulent regime, Part I: Numerical framework and its application to laminar flow simulation. Computers and Mathematics with Applications, 79(5): 1590–1618, March 1, 2020. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122119304675.

Wang:2020:SFVb

[WZC^{+}20b] Yong Wang, Chengwen Zhong, Jun Cao, Congshan Zhuo, and Sha Liu. A simplified finite volume lattice Boltzmann method for simulations of fluid flows from laminar to turbulent regime, Part II: Extension towards turbulent flow simulation. Computers and Mathematics with Applications, 79(8): 2133–2152, April 15, 2020. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122119305036.

Wu:2020:DES

[WZCS20] Yao Wu, Yong Zhao, Zhenhua Chai, and Baochang Shi. Discrete effects on some boundary schemes of multiple-relaxation-time lattice Boltzmann model for convection–diffusion equations. Computers and Mathematics with Applications, 80 (3):531–551, August 1, 2020. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122120301437.

Wang:2024:FEM

[WZD24] Shuo Wang, Xiangcheng Zheng, and Ning Du. Finite element method for an optimal control problem governed by a time fractional wave equation. Computers and Mathematics with Applications, 164 (??):45–66, June 15, 2024. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122124001482.

Wu:2025:PIM

[WZL25] Tianhao Wu, Changxiang Zhao, and Xuewen Liu. IS-PINN: Integrating mathematical principles to solve inhomogeneous wave equations via physics-informed neural networks. Computers and Mathematics with Applications, 193(??):177– 204, September 1, 2025. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122125002512.

Wang:2024:PFE

[WZS24a] Guoliang Wang, Bo Zheng, and Yueqiang Shang. A parallel finite element postprocessing algorithm for the damped Stokes equations. Computers and Mathematics with Applications, 153 (??):71–80, January 1, 2024. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122123004935.

Wen:2024:GFB

[WZS^{+}24b] J. C. Wen, Y. R. Zhou, J. Sladek, V. Sladek, and P. H. Wen. Galerkin finite block method in solid mechanics. Computers and Mathematics with Applications, 155(??):66–79, February 1, 2024. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122123005254.

Wang:2023:DPD

[WZT23] Kang Wang, Lei Zhang, and Shaoqiang Tang. Discovery of PDEs driven by data with sharp gradient or discontinuity. Computers and Mathematics with Applications, 140 (??):33–43, June 15, 2023. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122123001165.

Wang:2023:HFO

[WZTZ23] Zhenming Wang, Jun Zhu, Linlin Tian, and Ning Zhao. Hybrid fifth-order unequalsized weighted essentially non-oscillatory scheme for shallow water equations. Computers and Mathematics with Applications, 150(??): 1–14, November 15, 2023. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122123003887.

Wu:2021:CTI

[WZXN21] Shuang Wu, Jie Zhang, Qi Xiao, and Ming-Jiu Ni. Comparison of two interfacial flow solvers: Specific case of a single droplet impacting onto a deep pool. Computers and Mathematics with Applications, 81(??): 664–678, January 1, 2021. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122120300237.

Wang:2022:DSO

[WZY22] Ziqiang Wang, Jun Zhang, and Xiaofeng Yang. Decoupled, second-order accurate in time and unconditionally energy stable scheme for a hydrodynamically coupled ternary Cahn–Hilliard phase-field model of triblock copolymer melts. Computers and Mathematics with Applications, 124(??):241–257, October 15, 2022. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122122003728.

Wang:2023:DDV

[WZY23] Li Wang, Zijian Zhou, and Zhenya Yan. Data-driven vortex solitons and parameter discovery of 2D generalized nonlinear Schrödinger equations with a PT - symmetric optical lattice. Computers and Mathematics with Applications, 140 (??):17–23, June 15, 2023. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122123001116.

Wang:2020:NFO

[WZZ20] Zhenming Wang, Jun Zhu, and Ning Zhao. A new fifthorder finite difference wellbalanced multi-resolution WENO scheme for solving shallow water equations. Computers and Mathematics with Applications, 80(5):1387–1404, September 1, 2020. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122120302698.

Xi:2021:MPB

[XCFC21] Qiang Xi, C. S. Chen, Zhuojia Fu, and Eva Comino. The MAPS with polynomial basis functions for solving axisymmetric time-fractional equations. Computers and Mathematics with Applications, 88 (??):78–90, April 15, 2021. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122119305474.

Xiao:2024:NAV

[XCLX24] Qichang Xiao, Xiaoliang Cheng, Kewei Liang, and Hailing Xuan. Numerical analysis of a variationalhemivariational inequality governed by the Stokes equations. Computers and Mathematics with Applications, 159(??):1–10, April 1, 2024. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122124000233.

Xie:2025:SPS

[XDL^{+}25] Na Xie, Jian Dong, Min Liu, Zhen Li, and Dingfang Li. A structure-preserving surface reconstruction scheme for the movable bottom topography Saint-Venant–Exner model with wet-dry interfaces. Computers and Mathematics with Applications, 198(??): 321–349, November 15, 2025. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122125004134.

Xenophontos:2020:IAS

[Xen20]  C. Xenophontos.  Isoge-

ometric analysis for singularly perturbed highorder, two-point boundary value problems of reaction– diffusion type. Computers and Mathematics with Applications, 80(11):2340– 2350, December 2020. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122120302157.

Xu:2023:TNC

[XF23] Zhuangzhi Xu and Yayun Fu. Two novel conservative exponential relaxation methods for the space-fractional nonlinear Schrödinger equation. Computers and Mathematics with Applications, 142 (??):97–106, July 15, 2023. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122123001451.

Xie:2022:HHO

[XFLZ22] Chun-Mei Xie, Min-Fu Feng, Yan Luo, and Li Zhang. A hybrid high-order method for Sobolev equation with convection-dominated term. Computers and Mathematics with Applications, 118 (??):85–94, July 15, 2022. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122122001730.

Xiao:2023:ENS

[XFS23] Xufeng Xiao, Xinlong Feng, and Zuoqiang Shi. Efficient numerical simulation of Cahn–Hilliard type models by a dimension splitting method. Computers and Mathematics with Applications, 136(??):54– 70, April 15, 2023. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122123000500.

Xu:2022:NLC

[XFX22] Wen-Zhi Xu, Zhuo-Jia Fu, and Qiang Xi. A novel localized collocation solver based on a radial Trefftz basis for thermal conduction analysis in FGMs with exponential variations. Computers and Mathematics with Applications, 117 (??):24–38, July 1, 2022. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122122001560.

Xi:2024:FPU

[XFX^{+}24] Qiang Xi, Zhuojia Fu, Wenzhi Xu, Mi-An Xue, Youssef F. Rashed, and Jinhai Zheng. FEM-PIKFNN for underwater acoustic propagation induced by structural vibrations in different ocean environments. Computers and Mathematics with Applications, 176(??): 46–54, December 15, 2024. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122124004152.

Xu:2021:NSB

[XGL21] Yupeng Xu, Xi Gao, and Tingwen Li. Numerical study of the bi-disperse particles segregation inside a spherical tumbler with Discrete Element Method (DEM). Computers and Mathematics with Applications, 81(??): 588–601, January 1, 2021. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122119303621.

Xu:2023:HOA

[XHH^{+}23] Xingchun Xu, Yanwei Hu, Yurong He, Jiecai Han, and Jiaqi Zhu. High-order analysis of lattice Boltzmann models for the conservative Allen–Cahn equation. Computers and Mathematics with Applications, 146(??):106– 125, September 15, 2023. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122123002869.

Xue:2021:ALP

[XHL21] Dandan Xue, Yanren Hou, and Yi Li. Analysis of the local and parallel spacetime algorithm for the heat equation. Computers and Mathematics with Applications, 100(??):167–181, October 15, 2021. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122121003357.

Xue:2024:SOT

[XHL24] Dandan Xue, Yanren Hou, and Yi Li. A secondorder time parallel decoupled algorithm for the Stokes/ Darcy model. Computers and Mathematics with Applications, 161(??):13– 31, May 1, 2024. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122124000786.

Xu:2020:NDD

[XHM20] Fei Xu, Qiumei Huang, and Hongkun Ma. A novel domain decomposition framework for the ground state solution of Bose–Einstein condensates. Computers and Mathematics with Applications, 80(5):1287–1300, September 1, 2020. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122120302637.

Xie:2021:CGB

[XHY21] Yaning Xie, Zhongyi Huang, and Wenjun Ying. A Cartesian grid based tailored finite point method for reactiondiffusion equation on complex domains. Computers and Mathematics with Applications, 97(??):298–313, September 1, 2021. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122121002005.

Xiao:2021:IAW

[XJ21] Huiwen Xiao and Yee-Chung Jin. Improvement and application of weakly compressible moving particle semiimplicit method with kernelsmoothing algorithm. Computers and Mathematics with Applications, 99(??):37–51, October 1, 2021. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122121002789.

Xue:2020:IAG

[XJY^{+}20] Yaqiang Xue, Guoyong Jin, Tiangui Ye, Kangkang Shi, Saifeng Zhong, and Chuanmeng Yang. Isogeometric analysis for geometric modelling and acoustic attenuation performances of reactive mufflers. Computers and Mathematics with Applications, 79(12):3447– 3461, June 15, 2020. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122120300651.

Xia:2022:UES

[XKXL22] Qing Xia, Junseok Kim, Binhu Xia, and Yibao Li. An unconditionally energy stable method for binary incompressible heat conductive fluids based on the phase–field model. Computers and Mathematics with Applications, 123(??):26–39, October 1, 2022. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122122003133.

Xie:2020:NPN

[XL20] Xian Xie and Hou-Biao Li. A note on preconditioning for the 3 × 3 block saddle point problem. Computers and Mathematics with Applications, 79(12):3289– 3296, June 15, 2020. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S089812212030047X.

Xia:2022:EIN

[XL22] Boyang Xia and Jun Li. An efficient implementation of nodal discontinuous Galerkin lattice Boltzmann method and validation for direct numerical simulation of turbulent flows. Computers and Mathematics with Applications, 117 (??):284–298, July 1, 2022. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S089812212200178X.

Xu:2024:ADB

[XL24] Qiuyan Xu and Zhiyong Liu. Approximation of the derivatives beyond Taylor expansion. Computers and Mathematics with Applications, 158 (??):46–55, March 15, 2024. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122124000117.

Xu:2025:KFS

[XL25a] Qiuyan Xu and Zhiyong Liu. A kind of fast successive permutation iterative algorithms with the relaxation factor for nonlinear radiation diffusion problem. Computers and Mathematics with Applications, 187 (??):132–149, June 1, 2025. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122125001270.

Xu:2025:DTA

[XL25b] Sishang Xu and Wanshan Li.

Development and theoretical analysis of the energy stable nonsymmetric time-domain schemes for the 2D transverse electric mode of nonlinear Maxwell’s equations. Computers and Mathematics with Applications, 196(??): 433–457, October 15, 2025. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122125003517.

Xu:2020:SFL

[XLCP20] Yiran Xu, Jingye Li, Xiaohong Chen, and Guofei Pang. Solving fractional Laplacian visco-acoustic wave equations on complex-geometry domains using Grünwaldformula based radial basis collocation method. Computers and Mathematics with Applications, 79(8):2153– 2167, April 15, 2020. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122119305024.

Xu:2020:FDF

[XLLA20] Tao Xu, Fawang Liu, Shujuan Lü, and Vo V. Anh. Finite difference/finite element method for two-dimensional time–space fractional Bloch– Torrey equations with variable coefficients on irregular convex domains. Computers and Mathematics with Applications, 80(12):3173– 3192, December 15, 2020. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122120304326.

Xu:2025:SPB

[XLLR25] Yitong Xu, Ji Lin, Jun Lu, and Sergiy Reutskiy. Simulation of plate bending vibration problems by the meshless backward substitution method. Computers and Mathematics with Applications, 194(??):202– 213, September 15, 2025. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122125002585.

Xu:2022:EMM

[XLS22] Yuan Xu, Siu-Long Lei, and Hai-Wei Sun. An efficient multigrid method with preconditioned smoother for two-dimensional anisotropic space-fractional diffusion equations. Computers and Mathematics with Applications, 124(??):218–226, October 15, 2022. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122122003509.

Xi:2024:FEC

[XLS24] Yingxia Xi, Junshan Lin, and Jiguang Sun. A finite element contour integral method for computing the resonances of metallic grating structures with subwavelength holes. Computers and Mathematics with Applications, 170(??):161– 171, September 15, 2024. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122124002906.

Xu:2025:PFF

[XLSF25] Yuan Xu, Siu-Long Lei, Hai-Wei Sun, and Zhi-Wei Fang. A preconditioned fast finite volume method for two-dimensional conservative space-fractional diffusion equations on nonuniform meshes. Computers and Mathematics with Applications, 196(??):30–48, October 15, 2025. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S089812212500286X.

Xu:2025:KBC

[XLYY25] Wenwen Xu, Siqing Li, Junhong Yue, and Qi Ye. Kernelbased collocation methods with TBC for the elastic wave scattering by obstacles. Computers and Mathematics with Applications, 197(??): 200–216, November 1, 2025. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122125003542.

Xu:2021:NSW

[XLZ^{+}21] Fei Xu, Shuang Liang, Yaning Zhang, Bingxi Li, and Yiran Hu. Numerical study of water–air distribution in unsaturated soil by using lattice Boltzmann method. Computers and Mathematics with Applications, 81(??): 573–587, January 1, 2021. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122119304109.

Xu:2021:GPM

[XMQ^{+}21] Nengxiong Xu, Gang Mei, Jiayu Qin, Yazhe Li, and Liangliang Xu. GeoMFree^{3D}: a package of meshfree local Radial Point Interpolation Method (RPIM) for geomechanics. Computers and Mathematics with Applications, 81(??):113–132, January 1, 2021. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S089812212030242X.

Xu:2022:UOE

[XP22] Chao Xu and Lifang Pei.

Unconditional optimal error estimates of a modified finite element fully discrete scheme for the complex Ginzburg–Landau equation. Computers and Mathematics with Applications, 115(??):1–13, June 1, 2022. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122122001511.

Xuan:2025:AST

[XRS25] Anqing Xuan, Ziyan Ren, and Lian Shen. Accelerating simulations of turbulent flows over waves leveraging GPU parallelization. Computers and Mathematics with Applications, 183 (??):1–19, April 1, 2025. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122125000379.

Xu:2025:SMB

[XS25] Zhaodong Xu and Zhiqiang Sheng. Subspace method based on neural networks for solving the partial differential equation. Computers and Mathematics with Applications, 195(??):109– 138, October 1, 2025. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122125002822.

Xie:2024:NQF

[XSL24a] Huazhao Xie, Dongyang Shi, and Qian Liu. Nonconforming quadrilateral finite element analysis for the nonlinear Ginzburg– Landau equation. Computers and Mathematics with Applications, 171(??):139– 153, October 1, 2024. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122124003195.

Xu:2024:APU

[XSL24b] Xiaojing Xu, Wenjun Sun, and Qi Li. The asymptotic preserving unified gas kinetic scheme for the multi-scale kinetic SIR epidemic model. Computers and Mathematics with Applications, 174(??): 298–324, November 15, 2024.

CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122124004309.

Xu:2023:NMM

[XSZ^{+}23] Yi Xu, HongGuang Sun, Yuhui Zhang, Hai-Wei Sun, and Ji Lin. A novel meshless method based on RBF for solving variable-order time fractional advection– diffusion–reaction equation in linear or nonlinear systems. Computers and Mathematics with Applications, 142 (??):107–120, July 15, 2023. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122123001517.

Xie:2021:FSN

[XTZ^{+}21] Wenjia Xie, Zhengyu Tian, Ye Zhang, Hang Yu, and Weijie Ren. Further studies on numerical instabilities of Godunov-type schemes for strong shocks. Computers and Mathematics with Applications, 102(??):65–86, November 15, 2021. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122121003564.

Xu:2024:ECL

[XTZZ24] Yunjie Xu, Linlin Tian, Chunling Zhu, and Ning Zhao. Enhanced cascaded lattice Boltzmann model for multiphase flow simulations at large density ratio. Computers and Mathematics with Applications, 159 (??):92–101, April 1, 2024. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S089812212400035X.

Xu:2024:ECF

[Xu24] Qinwu Xu. An efficient computational framework for data assimilation of fractional-order dynamical system with sparse observations. Computers and Mathematics with Applications, 176(??):12–28, December 15, 2024. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122124004139.

Xiao:2024:BUT

[XW24a] Xiao-Yong Xiao and Cha-Sheng Wang. A block upper triangular preconditioner with two parameters for saddle-point problems. Computers and Mathematics with Applications, 160 (??):15–29, April 15, 2024. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122124000543.

Xu:2024:RBD

[XW24b] Hao Xu and Zeng-Qi Wang.

Rotated block diagonal preconditioners for Navier– Stokes control problems. Computers and Mathematics with Applications, 154 (??):1–11, January 15, 2024. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122123005072.

Xie:2023:EDS

[XWXF23] Na Xie, Yan Wang, Xufeng Xiao, and Xinlong Feng. An efficient dimension splitting p-adaptive method for the binary fluid surfactant phase field model. Computers and Mathematics with Applications, 142(??):84– 96, July 15, 2023. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122123001499.

Xu:2021:DAO

[XXAD21] Fei Xu, Qingang Xiong, Vadym Aizinger, and Guillaume Ducrozet. Development and application of open-source software for problems with numerical PDEs. Computers and Mathematics with Applications, 81 (??):1–2, January 1, 2021. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122120304557.

Xiong:2023:PSE

[XXLS23] Chunguang Xiong, Manting Xie, Fusheng Luo, and Hongling Su. A posteriori and superconvergence error analysis for finite element approximation of the Steklov eigenvalue problem. Computers and Mathematics with Applications, 144 (??):90–99, August 15, 2023. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122123002298.

Xu:2020:MET

[XXY20] Lin Xu, Shaosheng Xu, and Dingjun Yao. Maximizing expected terminal utility of an insurer with high gain tax by investment and reinsurance. Computers and Mathematics with Applications, 79(3):716–734, February 1, 2020. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S089812211930375X.

Xie:2023:EPF

[XXYL23] Wenxuan Xie, Qing Xia, Qian Yu, and Yibao Li. An effective phase field method for topology optimization without the curvature effects. Computers and Mathematics with Applications, 146(??):200–212, September 15, 2023. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122123002894.

Xiao:2024:LSF

[XYS^{+}24] Y. Xiao, L. M. Yang, C. Shu, H. Dong, Y. J. Du, and Y. X. Song. Least-square finite difference-based physicsinformed neural network for steady incompressible flows. Computers and Mathematics with Applications, 175(??): 33–48, December 1, 2024. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122124004024.

Xie:2021:TOI

[XYT^{+}21] Suqiong Xie, Kentaro Yaji, Toru Takahashi, Hiroshi Isakari, Masato Yoshino, and Toshiro Matsumoto. Topology optimization for incompressible viscous fluid flow using the lattice kinetic scheme. Computers and Mathematics with Applications, 97(??):251–266, September 1, 2021. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122121002248.

Xie:2023:EHO

[XZ23] Jianqiang Xie and Zhiyue Zhang. Efficient high-order physical property-preserving difference methods for nonlinear fourth-order wave equation with damping. Computers and Mathematics with Applications, 142 (??):64–83, July 15, 2023. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122123001475.

Xing:2024:HOG

[XZ24] Yanan Xing and Haibiao Zheng. A high order generalized finite difference method for solving the anisotropic elliptic interface problem in static and moving systems. Computers and Mathematics with Applications, 166 (??):1–23, July 15, 2024. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122124001871.

Xiang:2025:PTS

[XZ25] Yuting Xiang and Haibiao Zheng. Partitioned time stepping method for time-fractional Stokes–Darcy model with the Beavers– Joseph–Saffman interface conditions. Computers and Mathematics with Applications, 178(??):154–178, January 15, 2025. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122124005303.

Xu:2023:IBC

[XZS^{+}23] Xiaofeng Xu, Lian Zhang, Yin Shi, Long-Qing Chen, and Jinchao Xu. Integral boundary conditions in phase field models. Computers and Mathematics with Applications, 135(??): 1–5, April 1, 2023. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122122004941.

Xu:2023:PPF

[XZSY23] Jinjing Xu, Fei Zhao, Zhiqiang Sheng, and Guangwei Yuan. A positivitypreserving finite volume scheme for multi-group neutron diffusion kinetics equations with delayed neutrons. Computers and Mathematics with Applications, 129 (??):69–89, January 1, 2023. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122122004710.

Xie:2023:PEE

[XZT23] Yingying Xie, Liuqiang Zhong, and Ming Tang. A posteriori error estimate for a modified weak Galerkin method of 2D H(curl) elliptic problems. Computers and Mathematics with Applications, 142(??):307– 321, July 15, 2023. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122123001712.

Yazdi:2021:EAA

[YA21] Saeed Reza Sabbagh Yazdi and Tayebeh Amiri. An efficient automatic adaptive algorithm for cohesive crack propagation modeling of concrete structures using matrix-free unstructured Galerkin finite volume method. Computers and Mathematics with Applications, 97(??):237–250, September 1, 2021. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S089812212100242X.

Yan:2023:TLF

[YA23] Lexiang Yan and Rong An. Two-level finite element methods for the steady bio-convection flows problem. Computers and Mathematics with Applications, 146(??):126–136, September 15, 2023. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122123002870.

Yousuf:2025:FSM

[YAA25] M. Yousuf, M. Alshayqi, and S. S. Alzahrani. Fourier spectral methods based on restricted Padé approximations for space fractional reaction–diffusion systems. Computers and Mathematics with Applications, 181 (??):228–247, March 1, 2025. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122124005753.

Yang:2020:FNS

[Yan20] Xi Yang. A fast null-space method for the unsteady Stokes equations. Computers and Mathematics with Applications, 80(5):1459–1477, September 1, 2020. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122120302728.

Yang:2024:SPF

[Yan24]  Junxiang Yang.  Surface

phase-field surfactant fluid model and its practical closest point type finite difference computation. Computers and Mathematics with Applications, 154(??):24–38, January 15, 2024. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122123005217.

Yan:2025:EGR

[Yan25a]  Chunyu Yan.  Exploring

Gaussian radial basis function integrals for weight generation with application in financial option pricing. Computers and Mathematics with Applications, 181 (??):71–83, March 1, 2025. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122124005741.

Yang:2025:NEA

[Yan25b] Huaijun Yang. A new error analysis of a linearized BDF2 Galerkin scheme for Schrödinger equation with cubic nonlinearity. Computers and Mathematics with Applications, 188(??):83– 96, June 15, 2025. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S089812212500121X.

Yang:2024:ANS

[YCCW24] Zhengya Yang, Xuejuan Chen, Yanping Chen, and Jing Wang. Accurate numerical simulations for fractional diffusion equations using spectral deferred correction methods. Computers and Mathematics with Applications, 153(??):123– 129, January 1, 2024. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122123004911.

Yuan:2025:ESF

[YCGL25] Xiaodong Yuan, Aimin Chen, Rui Guo, and Maojun Li. An energy stable finite element method for the nonlocal electron heat transport model. Computers and Mathematics with Applications, 177(??):23–40, January 1, 2025. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122124005078.

You:2020:EBS

[YCL20] Xiangyu You, Yingbin Chai, and Wei Li. Edged-based smoothed point interpolation method for acoustic radiation with perfectly matched layer. Computers and Mathematics with Applications, 80(6):1596–1618, September 15, 2020. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S089812212030287X.

Yuan:2020:GLB

[YCWS20] Xiaolei Yuan, Zhenhua Chai, Huili Wang, and Baochang Shi. A generalized lattice Boltzmann model for fluid flow system and its application in two-phase flows. Computers and Mathematics with Applications, 79(6): 1759–1780, March 15, 2020. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122119304961.

Yang:2021:FHT

[YCZ^{+}21] Weidong Yang, Xuehui Chen, Xinru Zhang, Liancun Zheng, and Fawang Liu. Flow and heat transfer of viscoelastic fluid with a novel space distributedorder constitution relationship. Computers and Mathematics with Applications, 94 (??):94–103, July 15, 2021. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122121001632.

Yadav:2021:HMM

[YD21] K. S. Yadav and D. C. Dalal. The heterogeneous multiscale method to study particle size and partitioning effects in drug delivery. Computers and Mathematics with Applications, 92 (??):134–148, June 15, 2021. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122121001000.

Yadav:2023:MCS

[YD23] K. S. Yadav and D. C.

Dalal. A multiscale computational study of the effects of fluid flow and drug metabolism on drug delivery. Computers and Mathematics with Applications, 130 (??):58–68, January 15, 2023. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122122004825.

Yang:2023:NNB

[YDY^{+}23] Qihong Yang, Yangtao Deng, Yu Yang, Qiaolin He, and Shiquan Zhang. Neural networks based on power method and inverse power method for solving linear eigenvalue problems. Computers and Mathematics with Applications, 147(??):14–24, October 1, 2023. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122123003085.

Yemm:2024:EHH

[Yem24] Liam Yemm. An enriched hybrid high-order method for the Stokes problem with application to flow around submerged cylinders. Computers and Mathematics with Applications, 171(??):100– 113, October 1, 2024. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122124003122.

Yazidi:2021:SSI

[YEO21] Youness El Yazidi, Abdellatif Ellabib, and Youssef Ouakrim. On the stabilization of singular identification problem of an unknown discontinuous diffusion parameter in elliptic equation. Computers and Mathematics with Applications, 97(??): 267–285, September 1, 2021. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122121002443.

Yousuf:2020:HOT

[YFK20] M. Yousuf, K. M. Furati, and A. Q. M. Khaliq. Highorder time-stepping methods for two-dimensional Riesz fractional nonlinear reaction–diffusion equations. Computers and Mathematics with Applications, 80 (1):204–226, July 1, 2020. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122120301140.

Yousuf:2023:HFO

[YFK23] M. Yousuf, K. M. Furati, and A. Q. M. Khaliq. A hybrid fourth order time stepping method for space distributed order nonlinear reaction–diffusion equations. Computers and Mathematics with Applications, 151(??): 116–126, December 1, 2023. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122123004169.

Yang:2022:ADS

[YGZ^{+}22] Jiabao Yang, Zhichang Guo, Dazhi Zhang, Boying Wu, and Shan Du. An anisotropic diffusion system with nonlinear time-delay structure tensor for image enhancement and segmentation. Computers and Mathematics with Applications, 107(??):29–44, February 1, 2022. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122121004314.

Yang:2021:DGM

[YH21a] Di Yang and Yinnian He.

A discontinuous Galerkin method by patch reconstruction for convection-diffusion-reaction problems over polytopic meshes. Computers and Mathematics with Applications, 97(??):175–206, September 1, 2021. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122121002364.

Yuan:2021:PWD

[YH21b] Long Yuan and Qiya Hu. Plane wave discontinuous Galerkin methods for the Helmholtz equation and Maxwell equations in anisotropic media. Computers and Mathematics with Applications, 97(??):355–374, September 1, 2021. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122121002467.

Yin:2024:EFO

[YHF24] Fengli Yin, Dongdong Hu, and Yayun Fu. An efficient fourth-order structurepreserving scheme for the nonlocal Klein–Gordon–Schrödinger system. Computers and Mathematics with Applications, 167(??):123–133, August 1, 2024. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122124002268.

Yang:2025:WPS

[YHL25] Yi Yang, Jin Huang, and Hu Li. Well-posedness and strong convergence analysis of stochastic space-time fractional wave problems driven by fractional Brownian sheet. Computers and Mathematics with Applications, 194(??):177–201, September 15, 2025. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122125002597.

Yi:2022:EGM

[YHLA22] Son-Young Yi, Xiaozhe Hu, Sanghyun Lee, and James H. Adler. An enriched Galerkin method for the Stokes equations. Computers and Mathematics with Applications, 120(??):115– 131, August 15, 2022. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122122002632.

Yang:2022:ACP

[YHZ22] Di Yang, Yinnian He, and Yarong Zhang. Analysis and computation of a pressurerobust method for the rotation form of the incompressible Navier–Stokes equations with high-order finite elements. Computers and Mathematics with Applications, 112(??):1–22, April 15, 2022. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122122000815.

Yang:2024:SEA

[YJ24] Huaijun Yang and Xu Jia.

Superconvergence error analysis of Ciarlet–Raviart mixed finite element method for damped Boussinesq equation. Computers and Mathematics with Applications, 154(??):199–211, January 15, 2024. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122123005448.

Yang:2022:FDT

[YJL22] Jing-Yu Yang, Yao-Lin Jiang, and Jun Li. A fully discrete two-grid method for the diffusive Peterlin viscoelastic model. Computers and Mathematics with Applications, 119(??):118– 130, August 1, 2022. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122122002218.

Yu:2020:IAC

[YJZ20] Ying-Ying Yu, Ye Ji, and Chun-Gang Zhu. An improved algorithm for checking the injectivity of 2D toric surface patches. Computers and Mathematics with Applications, 79(10):2973– 2986, May 15, 2020. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S089812212030002X.

Yang:2020:PFM

[YK20] Junxiang Yang and Junseok Kim. A phase-field method for two-phase fluid flow in arbitrary domains. Computers and Mathematics with Applications, 79(6): 1857–1874, March 15, 2020. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122119304973.

Yang:2024:IAP

[YK24] Hee Jun Yang and Hyea Hyun Kim. Iterative algorithms for partitioned neural network approximation to partial differential equations. Computers and Mathematics with Applications, 170(??):237– 259, September 15, 2024. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122124003079.

Yang:2025:UMP

[YKK25] Junxiang Yang, Sangkwon Kim, and Junseok Kim. Unconditionally maximum principle preserving scheme for the Allen–Cahn equation with a logarithmic free energy. Computers and Mathematics with Applications, 195(??):366–375, October 1, 2025. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122125003219.

Yavich:2021:COT

[YKMZ21] Nikolay Yavich, Nikolay Khokhlov, Mikhail Malovichko, and Michael S. Zhdanov. Contraction operator transformation for the complex heterogeneous Helmholtz equation. Computers and Mathematics with Applications, 86(??):63–72, March 15, 2021. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122121000341.

Yosibash:2023:FSM

[YKNS23] Zohar Yosibash, Yekutiel Katz, Trabelsi Nir, and Amir Sternheim. Femurs segmentation by machine learning from CT scans combined with autonomous finite elements in orthopedic and endocrinology applications. Computers and Mathematics with Applications, 152(??): 16–27, December 15, 2023. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122123004297.

Yang:2020:NBT

[YL20a] Dong-Sheng Yang and Jing Ling. A novel boundarytype element-free method for 3D thermal analysis in inhomogeneous media with variable thermal source. Computers and Mathematics with Applications, 80(12):3123– 3136, December 15, 2020. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122120304260.

Yazdani:2020:NSP

[YL20b] Shima Yazdani and Esmail Lakzian. Numerical simulation and passive control of condensing flow through turbine blade by NVD Method using Eulerian–Lagrangian model. Computers and Mathematics with Applications, 80 (1):140–160, July 1, 2020. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S089812212030105X.

Yang:2022:NAF

[YL22] Xiaozhong Yang and Xinlong Liu. Numerical analysis of fourth-order compact difference scheme for inhomogeneous time-fractional Burgers–Huxley equation. Computers and Mathematics with Applications, 125 (??):1–12, November 1, 2022. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122122003510.

Yang:2025:PTP

[YL25] Hong Yang and Cheng-Xue Lao. Preconditioning techniques and parameterized regular splitting iteration methods for the Riesz distributed-order space-fractional diffusion equations with variable coefficients. Computers and Mathematics with Applications, 194(??):1–17, September 15, 2025. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122125002482.

Young:2021:TSM

[YLCC21] D. L. Young, Shin-Ruei Lin, Chuin-Shan Chen, and C. S. Chen. Two-step MPS-MFS ghost point method for solving partial differential equations. Computers and Mathematics with Applications, 94 (??):38–46, July 15, 2021. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122121001279.

Ye:2020:NSM

[YLFL20] Wenbin Ye, Jun Liu, Hongyuan Fang, and Gao Lin. Numerical solutions for magnetoelectro-elastic laminated plates resting on Winkler foundation or elastic half-space. Computers and Mathematics with Applications, 79(8): 2388–2410, April 15, 2020. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122119305401.

Yao:2025:EPR

[YLFZ25] Changhui Yao, Lei Li, Huijun Fan, and Yanmin Zhao. Energy-preserving RERK-FEM for the regularized logarithmic Schrödinger equation. Computers and Mathematics with Applications, 180(??):76–85, February 15, 2025. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122124005546.

Yang:2024:FDK

[YLL24] Jiye Yang, Yuqing Li, and Zhiyong Liu. A finite difference/Kansa method for the two-dimensional time and space fractional Bloch– Torrey equation. Computers and Mathematics with Applications, 156(??):1–15, February 15, 2024. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122123005539.

Yang:2022:OIR

[YLQ22] Zhi Yang, Xiaochuan Luo, and Jinwei Qiao. Optimal investigation of the reheating furnace based on a 3D numerical model with the formation of both oxide scale layer and skid marks. Computers and Mathematics with Applications, 128(??): 12–20, December 15, 2022. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122122003959.

Yong:2024:DMR

[YLSY24] Jinjun Yong, Xianbing Luo, Shuyu Sun, and Changlun Ye. Deep mixed residual method for solving PDE-constrained optimization problems. Computers and Mathematics with Applications, 176(??):510–524, December 15, 2024. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S089812212400508X.

Yang:2022:NBT

[YLW^{+}22] Dong-Sheng Yang, Jing Ling, Xiao-Bin Wang, Chi Pan, and Kun-Peng Qin. A new boundary-type meshfree method with RBFI for 2D thermo-elastic problems. Computers and Mathematics with Applications, 111 (??):81–90, April 1, 2022.

CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122122000633.

Yang:2023:GST

[YLW23] Peng Yang, Yonghai Li, and Xiang Wang. Global L^{2} superconvergence of the tetrahedral quadratic finite element. Computers and Mathematics with Applications, 133(??):104– 123, March 1, 2023. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S089812212300007X.

Yu:2020:EBC

[YLWZ20] Yang Yu, Xiaochuan Luo, Yuan Wang, and Huaxi (Yulin) Zhang. Estimation of boundary condition of twodimensional nonlinear PDE with application to continuous casting. Computers and Mathematics with Applications, 80(12):3082–3097, December 15, 2020. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122120304235.

Yang:2020:SST

[YLY20] X. L. Yang, Y. Liu, and L. Yang. A shear stress transport incorporated elliptic blending turbulence model applied to near-wall, separated and impinging jet flows and heat transfer. Computers and Mathematics with Applications, 79(12):3257– 3271, June 15, 2020. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122120300493.

Yao:2020:SAT

[YLZ20a] C. H. Yao, F. R. Li, and Y. M. Zhao. Superconvergence analysis of two-grid FEM for Maxwell’s equations with a thermal effect. Computers and Mathematics with Applications, 79(12): 3378–3393, June 15, 2020. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122120300626.

Ye:2020:BSF

[YLZ20b] Hailong Ye, Qiang Liu, and Mingjun Zhou. An L_{∞} bound for solutions of a fractional Cahn–Hilliard equation. Computers and Mathematics with Applications, 79(12):3353–3365, June 15, 2020. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122120300535.

Ye:2024:HON

[YLZ^{+}24] Xiao Ye, Jun Liu, Bingyin Zhang, Hongfei Fu, and Yue Liu. High order numerical methods based on quadratic spline collocation method and averaged L1 scheme for the variable-order time fractional mobile/immobile diffusion equation. Computers and Mathematics with Applications, 171(??):82–99, October 1, 2024. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122124003092.

Ye:2020:IBB

[YLZL20] Wenbin Ye, Jun Liu, Quansheng Zang, and Gao Lin. Investigation of bending behavior for laminated composite magneto-electro-elastic cylindrical shells subjected to mechanical or electric/ magnetic loads. Computers and Mathematics with Applications, 80(7):1839–1857, October 1, 2020. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122120303229.

Yu:2021:MUM

[YM21] Xiaojuan Yu and Changfeng Ma. The modified Uzawa methods for solving singular linear systems. Computers and Mathematics with Applications, 104(??):71–86, December 15, 2021. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122121004065.

Yu:2023:DLS

[YM23] Xuhong Yu and Qingrui Mao. Diagonalized Legendre spectral method for secondorder eigenvalue problems. Computers and Mathematics with Applications, 143(??): 269–276, August 1, 2023. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122123002274.

Yang:2025:BBL

[YMDW25] Zhiwei Yang, Jie Ma, Ning Du, and Hong Wang. A bond-based linear peridynamic model for viscoelastic materials and its efficient collocation method. Computers and Mathematics with Applications, 183 (??):121–136, April 1, 2025. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122125000331.

Yang:2025:SNS

[YMTP25] Yin Yang, Sayyed Ehsan Monabbati, Emran Tohidi, and Atena Pasban. Spatiotemporal numerical simulation of breast cancer tumors in one-dimensional nonlinear moving boundary models via temporal-spatial spectral collocation method. Computers and Mathematics with Applications, 187 (??):30–49, June 1, 2025. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122125001014.

Yosibash:2020:AFA

[YMTS20] Zohar Yosibash, Kent Myers, Nir Trabelsi, and Amir Sternheim. Autonomous FEs (AFE) — a stride toward personalized medicine. Computers and Mathematics with Applications, 80(11): 2417–2432, December 2020. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122120301164.

Yeoh:2020:LBH

[YOF20] Chin Vern Yeoh, Ean Hin Ooi, and Ji Jinn Foo. Lattice-Boltzmann hydrodynamics of single-square-grid generated turbulence — a partial entropic stabilisation approach. Computers and Mathematics with Applications, 80(5):1301–1326, September 1, 2020. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122120302625.

Yan:2021:NDI

[YPC^{+}21] Zhen-Guo Yan, Yu Pan, Giacomo Castiglioni, Koen Hillewaert, Joaquim Peiró, David Moxey, and Spencer J. Sherwin. Nektar++: Design and implementation of an implicit, spectral/hp element, compressible flow solver using a Jacobian-free Newton Krylov approach. Computers and Mathematics with Applications, 81(??): 351–372, January 1, 2021. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122120301073.

Yue:2021:MRT

[YPZ^{+}21] Xiaoqiang Yue, Kejia Pan, Jie Zhou, Zhifeng Weng, Shi Shu, and Juan Tang. A multigrid-reduction-in-time solver with a new two-level convergence for unsteady fractional Laplacian problems. Computers and Mathematics with Applications, 89(??):57–67, May 1, 2021. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122121000687.

Ye:2022:AAS

[YQZL22] Ximeng Ye, Guoliang Qin, Ya Zhuo, and Miaomiao Li. Analysis/application of stabilization by the overintegration technique in CBS-SEM for incompressible flow. Computers and Mathematics with Applications, 117(??):1–13, July 1, 2022. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122122001547.

Yang:2021:EAD

[YQZT21] Xuehua Yang, Wenlin Qiu, Haixiang Zhang, and Liang Tang. An efficient alternating direction implicit finite difference scheme for the three-dimensional timefractional telegraph equation. Computers and Mathematics with Applications, 102(??):233–247, November 15, 2021. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122121003758.

Yadav:2025:UAT

[YRC25] Vivek S. Yadav, Nishant Ranwan, and Nagaiah Chamakuri. Understanding avascular tumor growth and drug interactions through numerical analysis: a finite element method approach. Computers and Mathematics with Applications, 181 (??):55–70, March 1, 2025. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S089812212400573X.

Yang:2023:SOT

[YS23] Caijie Yang and Tongjun Sun. Second-order time discretization for reaction coefficient estimation of bilinear parabolic optimization problem with Neumann boundary conditions. Computers and Mathematics with Applications, 140(??):211– 224, June 15, 2023. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122123001542.

Yadav:2024:SCC

[YSB^{+}24] Sonia Yadav, Sukhjit Singh, R. P. Badoni, Ajay Kumar, and Mehakpreet Singh. Semilocal convergence of Chebyshev Kurchatov type methods for non-differentiable operators. Computers and Mathematics with Applications, 170(??):275–281, September 15, 2024. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122124003031.

Yang:2020:HOT

[YSGD20] Zhiqiang Yang, Yi Sun, Tianyu Guan, and Hao Dong. A high-order threescale approach for predicting thermo-mechanical properties of porous materials with interior surface radiation. Computers and Mathematics with Applications, 79 (9):2742–2770, May 1, 2020. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S089812211930567X.

Shen:2022:SOE

[ySRC22] Jin ye Shen, Jincheng Ren, and Shanzhen Chen. A second-order energy stable and nonuniform timestepping scheme for time fractional Burgers’ equation. Computers and Mathematics with Applications, 123(??): 227–240, October 1, 2022. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122122003261.

Yoshioka:2022:HJB

[YT22] Hidekazu Yoshioka and Motoh Tsujimura. Hamilton– Jacobi–Bellman–Isaacs equation for rational inattention in the long-run management of river environments under uncertainty. Computers and Mathematics with Applications, 112 (??):23–54, April 15, 2022. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122122000761.

Yoshioka:2023:LEA

[YTA23] Hidekazu Yoshioka, Tomomi Tanaka, and Futoshi Aranishi. Limit equations of adaptive Erlangization and their application to environmental management. Computers and Mathematics with Applications, 146(??):271– 293, September 15, 2023. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122123002973.

Yoshioka:2021:HFP

[YTH^{+}21] Hidekazu Yoshioka, Motoh Tsujimura, Kunihiko Hamagami, Yuta Yaegashi, and Yumi Yoshioka. HJB and Fokker–Planck equations for river environmental management based on stochastic impulse control with discrete and random observation. Computers and Mathematics with Applications, 96(??): 131–154, August 15, 2021. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122121001954.

Yang:2021:HOT

[YTK21] Junxiang Yang, Zhijun Tan, and Junseok Kim. High-order time-accurate, efficient, and structurepreserving numerical methods for the conservative Swift–Hohenberg model. Computers and Mathematics with Applications, 102(??):160– 174, November 15, 2021. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122121003709.

Yoshioka:2022:MCI

[YTT^{+}22] Hidekazu Yoshioka, Motoh Tsujimura, Tomohiro Tanaka, Yumi Yoshioka, and Ayumi Hashiguchi. Modeling and computation of an integral operator Riccati equation for an infinitedimensional stochastic differential equation governing streamflow discharge. Computers and Mathematics with Applications, 126(??):115– 148, November 15, 2022. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122122003893.

Yoshioka:2024:NAE

[YTY24] Hidekazu Yoshioka, Motoh Tsujimura, and Yumi Yoshioka. Numerical analysis of an extended mean field game for harvesting common fishery resource. Computers and Mathematics with Applications, 165 (??):88–105, July 1, 2024. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122124001615.

Yu:2023:LOL

[Yu23] Yue Yu. A lowest-order locking-free nonconforming virtual element method based on the reduced integration technique for linear elasticity problems. Computers and Mathematics with Applications, 135(??):157– 170, April 1, 2023. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122123000366.

Yuan:2023:GST

[Yua23] Long Yuan. Global spacetime Trefftz DG schemes for the time-dependent linear wave equation. Computers and Mathematics with Applications, 131(??):82–102, February 1, 2023. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122122004977.

Yuan:2024:P

[Yua24]  Jinyun Yuan.  Preface.

Computers and Mathematics with Applications, 156 (??):28, February 15, 2024. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S089812212300559X.

Yue:2020:AVV

[YW20] Gaocheng Yue and Jintao Wang. Attractors of the velocity-vorticity-Voigt model of the 3D Navier– Stokes equations with damping. Computers and Mathematics with Applications, 80 (3):434–452, August 1, 2020. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122120301218.

Yu:2024:ELL

[YW24] Xuhong Yu and Xiaoyang Wang. Efficient Legendre-Laguerre spectral element methods for problems on unbounded domains with diagonalization technique. Computers and Mathematics with Applications, 159 (??):11–20, April 1, 2024. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122124000361.

Yao:2021:GFE

[YWL21] Xin Yao, Yihe Wang, and Jianxing Leng. A general finite element method: Extension of variational analysis for nonlinear heat conduction with temperaturedependent properties and boundary conditions, and its implementation as local refinement. Computers and Mathematics with Applications, 100(??):11–29, October 15, 2021. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122121003163.

Yang:2023:SAL

[YWL23] Huaijun Yang, Lele Wang, and Xin Liao. Superconvergence analysis of a linearized energy-conservative Galerkin method for the nonlinear Schrödinger equation with wave operator. Computers and Mathematics with Applications, 133 (??):142–154, March 1, 2023. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122123000305.

Yue:2021:LNS

[YWLF21] Xingxing Yue, Fajie Wang, Po-Wei Li, and Chia-Ming Fan. Local non-singular knot method for large-scale computation of acoustic problems in complicated geometries. Computers and Mathematics with Applications, 84(??):128–143, February 15, 2021. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122120304843.

Yu:2024:OCT

[YWPS24] Yang Yu, Yu Wang, Xinfu Pang, and Liangliang Sun. Optimal control of threedimensional unsteady partial differential equations with convection term in continuous casting. Computers and Mathematics with Applications, 157(??):49– 64, March 1, 2024. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S089812212300576X.

Yang:2022:SPF

[YWT22] Junxiang Yang, Jian Wang, and Zhijun Tan. A simple and practical finite difference method for the phasefield crystal model with a strong nonlinear vacancy potential on 3D surfaces. Computers and Mathematics with Applications, 121(??):131– 144, September 1, 2022. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122122002929.

Yao:2021:LFF

[YWZ21] C. H. Yao, Z. Y. Wang, and Y. M. Zhao. A leap-frog finite element method for wave propagation of Maxwell– Schrödinger equations with nonlocal effect in metamaterials. Computers and Mathematics with Applications, 90 (??):25–37, May 15, 2021. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122121000675.

Yuan:2023:PFB

[YWZ^{+}23] Xiaolei Yuan, Yao Wu, Chunhua Zhang, Zhenhua Chai, and Baochang Shi. A phase-field-based multiplerelaxation-time lattice Boltzmann method for incompressible multiphase flows with density and viscosity contrasts. Computers and Mathematics with Applications, 144(??):237–256, August 15, 2023. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122123002377.

Yu:2021:HBW

[YX21] Qiang Yu and Hang Xu. A homotopy-based wavelet approach for large deflection of a circular plate on nonlinear foundations with parameterized boundaries. Computers and Mathematics with Applications, 90 (??):80–95, May 15, 2021. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122121000894.

Yaghobipour:2020:SQS

[YY20a] S. Yaghobipour and M. Yarahmadi. Solving quantum stochastic LQR optimal control problem in Fock space and its application in finance. Computers and Mathematics with Applications, 79(10): 2832–2845, May 15, 2020. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122119305826.

Yoshioka:2020:RSC

[YY20b] Hidekazu Yoshioka and Yumi Yoshioka. Regime-switching constrained viscosity solutions approach for controlling dam–reservoir systems. Computers and Mathematics with Applications, 80(9):2057–2072, November 1, 2020. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122120303618.

Yadav:2023:SFC

[YY23] Pramod Kumar Yadav and Nitisha Yadav. A study on the flow of couple stress fluid in a porous curved channel. Computers and Mathematics with Applications, 152(??): 1–15, December 15, 2023. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122123004510.

Yao:2025:MSC

[YY25] Jinran Yao and Zhengwei Yin. Mean-square contractivity and convergence rate of stochastic theta schemes for non-autonomous SDEs with non-globally Lipschitz continuous coefficients. Computers and Mathematics with Applications, 193(??):157– 176, September 1, 2025. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122125002408.

Yu:2020:FVS

[YYSL20] Yao Yu, Guangwei Yuan, Zhiqiang Sheng, and Yonghai Li. The finite volume scheme preserving maximum principle for diffusion equations with discontinuous coefficient. Computers and Mathematics with Applications, 79(8):2168– 2188, April 15, 2020. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122119305012.

Yoshioka:2020:ACD

[YYY^{+}20] Hidekazu Yoshioka, Yumi Yoshioka, Yuta Yaegashi, Tomomi Tanaka, Masahiro Horinouchi, and Futoshi Aranishi. Analysis and computation of a discrete costly observation model for growth estimation and management of biological resources. Computers and Mathematics with Applications, 79(4):1072– 1093, February 15, 2020. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122119304146.

Yodono:2022:TOE

[YYY^{+}22] Takashi Yodono, Kentaro Yaji, Takayuki Yamada, Kozo Furuta, Kazuhiro Izui, and Shinji Nishiwaki. Topology optimization for the elastic field using the lattice Boltzmann method. Computers and Mathematics with Applications, 110(??):123– 134, March 15, 2022. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122122000463.

Yang:2022:SHE

[YYZ22] Xiaochen Yang, Zhanwen Yang, and Chiping Zhang. Stochastic heat equation: Numerical positivity and almost surely exponential stability. Computers and Mathematics with Applications, 119(??):312–318, August 1, 2022. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122122002243.

Yang:2021:NID

[YZ21] Jian-Qing Yang and Bing-Xin Zhao. Numerical investigation of double-diffusive convection in rectangular cavities with different aspect ratio I: High-accuracy numerical method. Computers and Mathematics with Applications, 94(??):155– 169, July 15, 2021. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122121001759.

Yang:2022:NIE

[YZJ^{+}22] Qin Yang, Ang Zhang, Bin Jiang, Liang Gao, Zhipeng Guo, Jiangfeng Song, Shoumei Xiong, and Fusheng Pan. Numerical investigation of eutectic growth dynamics under convection by 3D phasefield method. Computers and Mathematics with Applications, 114(??):83– 94, May 15, 2022. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122122001146.

Yan:2020:NSU

[YZL20] Dong Yan, Song-Ping Zhu, and Xiaoping Lu. A numerical study of the utilityindifference approach for pricing American options. Computers and Mathematics with Applications, 80(5): 894–905, September 1, 2020. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122120302054.

Yang:2022:TGM

[YZN22] Jiming Yang, Jing Zhou, and Cunyun Nie. A twogrid method for discontinuous Galerkin approximations to compressible miscible displacement problems. Computers and Mathematics with Applications, 115 (??):57–67, June 1, 2022. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122121004570.

Yang:2021:OSF

[YZT21] Xuehua Yang, Haixiang Zhang, and Jie Tang. The OSC solver for the fourthorder sub-diffusion equation with weakly singular solutions. Computers and Mathematics with Applications, 82 (??):1–12, January 15, 2021. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122120304430.

Yue:2022:APF

[YZW^{+}22] Qiang Yue, Wei Zhou, Qiao Wang, Y. T. Feng, Gang Ma, and Xiaolin Chang. An adaptive phase-field model based on bilinear elements for tensile-compressive-shear fracture. Computers and Mathematics with Applications, 105(??):112–135, January 1, 2022. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122121004077.

Ye:2020:LFW

[YZZ20] Xiu Ye, Shangyou Zhang, and Zhimin Zhang. A locking-free weak Galerkin finite element method for Reissner–Mindlin plate on polygonal meshes. Computers and Mathematics with Applications, 80(5):906–916, September 1, 2020. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122120302194.

Zhao:2023:TDQ

[ZA23] Yong Zhi Zhao and Zhi Yong Ai. The transformed differential quadrature method for solving time-dependent partial differential equations: Framework and examples. Computers and Mathematics with Applications, 140 (??):183–194, June 15, 2023.

CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122123001505.

Zhang:2023:LEE

[ZAH23] Dan Zhang, Na An, and Chaobao Huang. Local error estimates of the fourthorder compact difference scheme for a time-fractional diffusion-wave equation. Computers and Mathematics with Applications, 142(??):283– 292, July 15, 2023. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122123002080.

Zygiridis:2023:FDT

[ZAPK23] Theodoros T. Zygiridis, Stamatios A. Amanatiadis, Aristeides D. Papadopoulos, and Nikolaos V. Kantartzis. A finite-difference time-domain method for Lorentz dispersive media with reduced errors within arbitrary frequency bands. Computers and Mathematics with Applications, 137(??):102– 111, May 1, 2023. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122123000779.

Zamora:2024:EFV

[ZBCS24] Esteban A. Zamora, Laura Battaglia, Marcela Cruchaga, and Mario Storti. Embedded finite volume technique for fluid/rigid-body interaction problems. Computers and Mathematics with Applications, 156(??):157–166, February 15, 2024. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122123005989.

Zarareh:2022:ISA

[ZBKC22] Amin Zarareh, Stephen B. Burnside, Sorush Khajepor, and Baixin Chen. Improving the staircase approximation for wettability implementation of phase-field model: Part 2 — three-component permeation. Computers and Mathematics with Applications, 109(??):100–124, March 1, 2022. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122122000050.

Zhang:2022:GCS

[ZC22] Zemian Zhang and Xuesong Chen. A global convergent semi-smooth Newton method for semi-linear elliptic optimal control problem. Computers and Mathematics with Applications, 120 (??):15–27, August 15, 2022. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122122002589.

Zhang:2023:EEF

[ZC23] Jiehua Zhang and Zhongying Chen. L^{2} error estimates for a family of cubic finite volume methods on triangular meshes. Computers and Mathematics with Applications, 143(??):189– 223, August 1, 2023. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122123001815.

Zheng:2023:NSM

[ZCA23] Jihui Zheng, Waixiang Cao, and Jing An. A novel spectral method and error estimates for fourth-order problems with mixed boundary conditions in a cylindrical domain. Computers and Mathematics with Applications, 152(??):221–238, December 15, 2023. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122123004716.

Zeng:2020:RPS

[ZCL20] Jiaoyan Zeng, Yanping Chen, and Guichang Liu. Rough polyharmonic splines method for optimal control problem governed by parabolic systems with rough coefficient. Computers and Mathematics with Applications, 80(1):121–139, July 1, 2020. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122120300985.

Zheng:2024:EVE

[ZCW24] Xianyan Zheng, Jinru Chen, and Feng Wang. An extended virtual element method for elliptic interface problems. Computers and Mathematics with Applications, 156(??):87–102, February 15, 2024. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122123005680.

Zhang:2020:NGP

[ZCXG20] Jia-Le Zhang, Hong-Quan Chen, Sheng-Guan Xu, and Huan-Qin Gao. A novel GPU-parallelized meshless method for solving compressible turbulent flows. Computers and Mathematics with Applications, 80(12):2738– 2763, December 15, 2020. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122120303849.

Zhan:2022:LLB

[ZCY22] Ningyu Zhan, Rongqian Chen, and Yancheng You. Linear lattice Boltzmann flux solver for simulating acoustic propagation. Computers and Mathematics with Applications, 114(??):21– 40, May 15, 2022. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122122001249.

Zhan:2024:LLB

[ZCYL24] Ningyu Zhan, Rongqian Chen, Yancheng You, and Zelun Lin. Linear lattice Boltzmann flux solver with compact third-order finite volume method for acoustic propagation simulations on three-dimensional hybrid unstructured grids. Computers and Mathematics with Applications, 156(??):103–120, February 15, 2024. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122123005527.

Zhao:2021:DPR

[ZCZ^{+}21] Wenke Zhao, Xin Chen, Yaning Zhang, Wentao Su, Fei Xu, and Bingxi Li. Deicing performances of a road unit driven by a hydronic heating system in severely cold regions of China. Computers and Mathematics with Applications, 81(??):838–850, January 1, 2021. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S089812211930505X.

Zghal:2021:VCP

[ZD21] S. Zghal and F. Dammak. Vibration characteristics of plates and shells with functionally graded pores imperfections using an enhanced finite shell element. Computers and Mathematics with Applications, 99 (??):52–72, October 1, 2021. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122121002856.

Zhang:2022:DUE

[ZD22] Xiaodi Zhang and Qianqian Ding. A decoupled, unconditionally energy stable and charge-conservative finite element method for inductionless magnetohydrodynamic equations. Computers and Mathematics with Applications, 127(??):80–96, December 1, 2022. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122122004035.

Zhang:2022:PRP

[ZDLC22] Yongou Zhang, Sina Dang, Wei Li, and Yingbin Chai. Performance of the radial point interpolation method (RPIM) with implicit time integration scheme for transient wave propagation dynamics. Computers and Mathematics with Applications, 114(??):95–111, May 15, 2022. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122122001201.

Zhou:2022:LBM

[ZDLL22] Xun Zhou, Bo Dong, Weizhong Li, and Kunfeng Liang. Lattice Boltzmann modeling of primary liquid film breakup at the prefilmer edge: Effects of surface wettability. Computers and Mathematics with Applications, 109 (??):146–157, March 1, 2022. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S089812212200030X.

Zhang:2022:DQV

[ZDT^{+}22] Shun Hu Zhang, Lei Deng, Wen Hao Tian, Li Zhi Che, and Yan Li. Deduction of a quadratic velocity field and its application to rolling force of extra-thick plate. Computers and Mathematics with Applications, 109 (??):58–73, March 1, 2022. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122122000311.

Zdunek:2021:TFM

[Zdu21] Adam Zdunek. Tests with FALKSOL — a massively parallel multi-level domain decomposing direct solver. Computers and Mathematics with Applications, 97(??): 207–222, September 1, 2021. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S089812212100239X.

Zhang:2020:NSD

[ZDW^{+}20] Ziying Zhang, Jun Du, Zhengying Wei, Zhen Wang, Haoqiang Zhang, Minghui Li, and Yiping Tang. Numerical simulation of dynamic seeding of mesenchymal stem cells in pore structure. Computers and Mathematics with Applications, 79(1):88–99, January 2020. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122118300439.

Zhang:2023:MME

[ZDW23] Yuhao Zhang, Haiyun Dong, and Kun Wang. Mass, momentum and energy identicalrelation-preserving scheme for the Navier–Stokes equations with variable density. Computers and Mathematics with Applications, 137 (??):73–92, May 1, 2023. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122123000627.

Zhang:2022:NIR

[ZDWR22] Wen Zhang, Zirong Ding, Zewen Wang, and Zhousheng Ruan. Numerical inversion of reaction parameter for a time-fractional diffusion equation by Legendre spectral collocation and mollification method. Computers and Mathematics with Applications, 128(??):188– 197, December 15, 2022. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122122004515.

Zhang:2022:VEM

[ZF22] Yadong Zhang and Minfu Feng. The virtual element method for the time fractional convection diffusion reaction equation with non-smooth data. Computers and Mathematics with Applications, 110(??):1–18, March 15, 2022. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122122000487.

Zhang:2024:MVE

[ZF24] Xi Zhang and Minfu Feng.

A mixed virtual element method for the two-dimensional Navier–Stokes equations in stream-function formulation. Computers and Mathematics with Applications, 175(??): 119–137, December 1, 2024. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122124004292.

Zhang:2022:GLS

[ZFL22] Yanming Zhang, Yan Fan, and Yu Li. General linear and spectral Galerkin methods for the nonlinear two-sided space distributedorder diffusion equation. Computers and Mathematics with Applications, 113(??): 1–12, May 1, 2022. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122122001055.

Zhu:2024:EAC

[ZFQ24] Danchen Zhu, Xinlong Feng, and Lingzhi Qian. Error analysis of Crank–Nicolsonleapfrog scheme for the two-phase Cahn–Hilliard– Navier–Stokes incompressible flows. Computers and Mathematics with Applications, 172(??):78–93, October 15, 2024. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122124003328.

Zhang:2020:ITG

[ZFZ20] Weifeng Zhang, Ronghong Fan, and Liuqiang Zhong. Iterative two-grid methods for semilinear elliptic equations. Computers and Mathematics with Applications, 80 (3):522–530, August 1, 2020. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122120301449.

Zhang:2024:TGW

[ZGC24] Jianghong Zhang, Fuzheng Gao, and Jintao Cui. Twogrid weak Galerkin finite element method for nonlinear parabolic equations. Computers and Mathematics with Applications, 175(??): 356–365, December 1, 2024. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122124004504.

Zhang:2025:SDL

[ZGG^{+}25] Zhengfang Zhang, Lulu Guo, Xiangjing Gao, Weifeng Chen, and Xiaoliang Cheng. Shape derivative of the Laplacian eigenvalue problem. Computers and Mathematics with Applications, 199(??):127–147, December 1, 2025. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122125003943.

Zhao:2023:BPL

[ZGL23] Yong-Liang Zhao, Xian-Ming Gu, and Hu Li. A bilateral preconditioning for an L2-type all-at-once system from time-space non-local evolution equations with a weakly singular kernel. Computers and Mathematics with Applications, 148(??):200– 210, October 15, 2023. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S089812212300353X.

Zhang:2023:TCD

[ZGW23] Chunhua Zhang, Zhaoli Guo, and Lian-Ping Wang. A thermodynamically consistent diffuse interface model for multi-component two-phase flow with partial miscibility. Computers and Mathematics with Applications, 150(??): 22–36, November 15, 2023. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122123003917.

Zhang:2020:NDG

[ZH20] Jiansong Zhang and Huiran Han. A new discontinuous Galerkin mixed finite element method for compressible miscible displacement problem. Computers and Mathematics with Applications, 80(6):1714–1725, September 15, 2020. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122120303151.

Zhang:2023:MMM

[ZH23a] Xuqing Zhang and Jiayu Han. Multigrid methods for the modified elastic transmission eigenvalue problem. Computers and Mathematics with Applications, 142 (??):31–47, July 15, 2023. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122123001426.

Zhang:2023:FEM

[ZH23b] Yarong Zhang and Yinnian He. Finite element method for semi-linear elliptical equation with a 2m − 1 polynomial nonlinearity. Computers and Mathematics with Applications, 139 (??):91–98, June 1, 2023. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122123001025.

Zhou:2023:TGA

[ZH23c] Yingcong Zhou and Tianliang Hou. Two-grid algorithm of lumped mass finite element approximation for Allen–Cahn equations. Computers and Mathematics with Applications, 152(??): 46–55, December 15, 2023. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122123004601.

Zhang:2021:ASM

[Zha21] Yinglong J. Zhang. Assessment of subgrid method in a finite-volume model. Computers and Mathematics with Applications, 81(??): 220–236, January 1, 2021. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122120301929.

Zhang:2022:FQF

[Zha22a]  Jiehua Zhang.  A fam-

ily of quadratic finite volume method for solving the Stokes equation. Computers and Mathematics with Applications, 117(??):155– 186, July 1, 2022. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122122001705.

Zhang:2022:BGP

[Zha22b]  Shun Zhang.  Battling

Gibbs phenomenon: On finite element approximations of discontinuous solutions of PDEs. Computers and Mathematics with Applications, 122(??):35–47, September 15, 2022. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122122002978.

Zhang:2022:HOI

[Zha22c] Xiaodi Zhang. High order interface-penalty finite element methods for elasticity interface problems in 3D. Computers and Mathematics with Applications, 114 (??):161–170, May 15, 2022. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122122001419.

Zhao:2022:GFD

[Zha22d]  Jia Zhao.  A general

framework to derive linear, decoupled and energystable schemes for reversibleirreversible thermodynamically consistent models. Computers and Mathematics with Applications, 110 (??):91–109, March 15, 2022. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122121004466.

Zhao:2022:BCL

[Zha22e] Zhenyu Zhao. Boundary condition limitation in an inverse source problem and its overcoming. Computers and Mathematics with Applications, 111(??):124– 133, April 1, 2022. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122122000694.

Zhang:2023:SPC

[Zha23] Shun Zhang. A simple proof of coerciveness of first-order system least-squares methods for general second-order elliptic PDEs. Computers and Mathematics with Applications, 130(??):98–104, January 15, 2023. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122122004904.

Zhao:2024:PFM

[Zha24] Lina Zhao. A parameterfree mixed formulation for the Stokes equations and linear elasticity with strongly symmetric stress. Computers and Mathematics with Applications, 155(??):35–50, February 1, 2024. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122123005436.

Zheng:2022:DDB

[Zhe22] QingQing Zheng. Domain decomposition based preconditioner combined local low-rank approximation with global corrections ^{1}. Computers and Mathematics with Applications, 114 (??):41–46, May 15, 2022. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S089812212200089X.

Zhao:2021:CAN

[ZHH21] Xinyue Evelyn Zhao, Wenrui Hao, and Bei Hu. Convergence analysis of neural networks for solving a free boundary problem. Computers and Mathematics with Applications, 93 (??):144–155, July 1, 2021. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122121001139.

Zhang:2023:DIL

[ZHHL23] Shi-Ting Zhang, Yang Hu, Qiang He, and Qian-Ping Li. A diffuse interface–lattice Boltzmann model for conjugate heat transfer with imperfect interface. Computers and Mathematics with Applications, 151(??):134–152, December 1, 2023. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122123004194.

Zhou:2023:UPS

[ZHPW23] Zhongguo Zhou, Tongtong Hang, Hao Pan, and Yan Wang. The upwind PPM scheme and analysis for solving two-sided space-fractional advection– diffusion equations in three dimension. Computers and Mathematics with Applications, 150(??):70–86, November 15, 2023. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122123003905.

Zhu:2020:NIM

[ZHW20] Wenxing Zhu, Pengzhan Huang, and Kun Wang. Newton iterative method based on finite element discretization for the stationary Darcy–Brinkman equations. Computers and Mathematics with Applications, 80(12):3098–3122, December 15, 2020. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122120304223.

Zhang:2021:NTG

[ZHYL21] Jiansong Zhang, Huiran Han, Yun Yu, and Jun Liu. A new two-grid mixed finite element analysis of semi-linear reaction–diffusion equation. Computers and Mathematics with Applications, 92 (??):172–179, June 15, 2021. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122121001140.

Zhang:2022:FPM

[ZHZ^{+}22] Fan Zhang, Can Huang, Huashan Zhang, Tiegang Liu, and Moubin Liu. A finite particle method based on a Riemann solver for modeling incompressible flows. Computers and Mathematics with Applications, 124 (??):74–88, October 15, 2022. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122122003376.

Zhao:2023:UES

[ZHZ23] Guomei Zhao, Shuaifei Hu, and Peicheng Zhu. Unconditional energy stability and temporal convergence of first-order numerical scheme for the square phase-field crystal model. Computers and Mathematics with Applications, 143(??): 318–326, August 1, 2023. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122123002213.

Zhang:2025:PBW

[ZJ25] Ren-Hao Zhang and Yao-Lin Jiang. Parareal-based waveform relaxation method for semi-linear parabolic equation. Computers and Mathematics with Applications, 195 (??):52–67, October 1, 2025. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122125002731.

Zhao:2021:NSC

[ZJCG21] Xiao Zhao, Hui Jin, Yunan Chen, and Zhiwei Ge. Numerical study of H _{2}, CH _{4}, CO, O _{2} and CO _{2} diffusion in water near the critical point with molecular dynamics simulation. Computers and Mathematics with Applications, 81(??): 759–771, January 1, 2021. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122119305450.

Zhang:2020:OEE

[ZJJ20] Hui Zhang, Junqing Jia, and Xiaoyun Jiang. An optimal error estimate for the two-dimensional nonlinear time fractional advection– diffusion equation with smooth and non-smooth solutions. Computers and Mathematics with Applications, 79(10): 2819–2831, May 15, 2020. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122119305735.

Zhang:2024:MPP

[ZJZ24] Yi Zhang, Ye Ji, and Chun-Gang Zhu. Multi-patch parameterization method for isogeometric analysis using singular structure of crossfield. Computers and Mathematics with Applications, 162 (??):61–78, May 15, 2024. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122124001007.

Zimmerman:2021:MFE

[ZK21] Alexander G. Zimmerman and Julia Kowalski. Mixed finite elements for convectioncoupled phase-change in enthalpy form: Open software verified and applied to 2D benchmarks. Computers and Mathematics with Applications, 84(??):77–96, February 15, 2021. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122120304338.

Zarareh:2021:ISA

[ZKBC21] Amin Zarareh, Sorush Khajepor, Stephen B. Burnside, and Baixin Chen. Improving the staircase approximation for wettability implementation of phase-field model: Part 1 — static contact angle. Computers and Mathematics with Applications, 98(??):218–238, September 15, 2021. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122121002765.

Zarth:2021:TSO

[ZKTK21] Asher Zarth, Fabian Klemens, Gudrun Thäter, and Mathias J. Krause. Towards shape optimisation of fluid flows using lattice Boltzmann methods and automatic differentiation. Computers and Mathematics with Applications, 90 (??):46–54, May 15, 2021. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S089812212100064X.

Zhang:2020:VMI

[ZL20a] Tao Zhang and Xiaolin Li. Variational multiscale interpolating element-free Galerkin method for the nonlinear Darcy–Forchheimer model. Computers and Mathematics with Applications, 79(2):363–377, January 15, 2020. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122119303414.

Zhang:2020:NNB

[ZL20b] Xingyu Zhang and Hua Li.

Neural-network-based control of discrete-phase concentration in a gas-particle corner flow with optimal energy consumption. Computers and Mathematics with Applications, 80(5):1360–1374, September 1, 2020. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122120302686.

Zhang:2022:USS

[ZL22a] Jin Zhang and Xiaowei Liu. Uniform stability of the SUPG method for the evolutionary convection– diffusion–reaction equation. Computers and Mathematics with Applications, 124 (??):1–6, October 15, 2022. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122122003340.

Zheng:2022:TAG

[ZL22b] Zhiyin Zheng and Xiaolin Li. Theoretical analysis of the generalized finite difference method. Computers and Mathematics with Applications, 120(??):1–14, August 15, 2022. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122122002620.

Zhang:2023:CCP

[ZL23a] Wen Zhang and Jian Li. CPINNs: a coupled physicsinformed neural networks for the closed-loop geothermal system. Computers and Mathematics with Applications, 132(??):161–179, February 15, 2023. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122123000032.

Zhang:2023:RPI

[ZL23b] Wen Zhang and Jian Li. The robust physics-informed neural networks for a typical fourth-order phase field model. Computers and Mathematics with Applications, 140(??):64–77, June 15, 2023. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0898122123001104.

Zhao:2023:NST

[ZL23c] Zhihui Zhao and Hong Li. Numerical study of twodimensional Burgers’ equation by using a continuous Galerkin method. Computers and Mathematics with Applications, 149(??):38–48, November 1, 2023. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122123003759.

Zhang:2025:TGM

[ZL25] Jiansong Zhang and Yanyu Liu. Two-grid mixed finite element analysis of semilinear second order hyperbolic problem. Computers and Mathematics with Applications, 189(??):70– 85, July 1, 2025. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122125001373.

Zhou:2023:EPE

[ZLD^{+}23] Xun Zhou, Miaomiao Liu, Bo Dong, Weizhong Li, and Kunfeng Liang. Effects of prefilmer edge configuration on primary liquid film breakup: a lattice Boltzmann study. Computers and Mathematics with Applications, 146(??):33–44, September 15, 2023. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122123002754.

Zhao:2022:OSF

[ZLF^{+}22] Han Zhao, Xiangbei Liu, Andrew H. Fletcher, Ru Xiang, John T. Hwang, and David Kamensky. An opensource framework for coupling non-matching isogeometric shells with application to aerospace structures. Computers and Mathematics with Applications, 111 (??):109–123, April 1, 2022. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122122000657.

Zhao:2022:IFI

[ZLH22] Wei Zhao, Min Lei, and Yiu-Chung Hon. An improved finite integration method for nonlocal nonlinear Schrödinger equations. Computers and Mathematics with Applications, 113 (??):24–33, May 1, 2022. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122122000906.

Zhang:2024:PPE

[ZLHC24] Jun Zhang, Zijiang Luo, Jiayu Han, and Hu Chen. Apriori and a-posteriori error estimates for discontinuous Galerkin method of the Maxwell eigenvalue problem. Computers and Mathematics with Applications, 176(??): 190–201, December 15, 2024. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122124004693.

Zhang:2022:RSA

[ZLJ^{+}22] Ying Zhang, Chong-Jun Li, Yan-Mei Jia, Zhen-Yi Liu, and Juan Chen. Revised SBFEM on arbitrary polygons and faceted polyhedrons with the second order completeness by elimination of bubble functions. Computers and Mathematics with Applications, 109 (??):74–99, March 1, 2022. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122122000128.

Zheng:2020:FDS

[ZLJT20] Rumeng Zheng, Fawang Liu, Xiaoyun Jiang, and Ian W. Turner. Finite difference/ spectral methods for the two-dimensional distributedorder time-fractional cable equation. Computers and Mathematics with Applications, 80(6):1523–1537, September 15, 2020. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122120302662.

Zhao:2020:ACG

[ZLL20] Zhihui Zhao, Hong Li, and Yang Liu. Analysis of a continuous Galerkin method with mesh modification for two-dimensional telegraph equation. Computers and Mathematics with Applications, 79(3):588–602, February 1, 2020. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122119303438.

Zhao:2023:HOC

[ZLL23] Xuan Zhao, Ziyan Li, and Xiaoli Li. High order compact finite difference methods for non-Fickian flows in porous media. Computers and Mathematics with Applications, 136(??):95– 111, April 15, 2023. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122123000706.

Zhang:2020:LAS

[ZLPR20] Qifeng Zhang, Xiaoman Lin, Kejia Pan, and Yunzhu Ren. Linearized ADI schemes for two-dimensional space-fractional nonlinear Ginzburg–Landau equation. Computers and Mathematics with Applications, 80 (5):1201–1220, September 1, 2020. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122120302492.

Zhang:2023:GMS

[ZLQ23] Tianwei Zhang, Yuntao Liu, and Huizhen Qu. Global mean-square exponential stability and random periodicity of discrete-time stochastic inertial neural networks with discrete spatial diffusions and Dirichlet boundary condition. Computers and Mathematics with Applications, 141(??):116– 128, July 1, 2023. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122123001487.

Liang:2024:MIB

[zLqLZ^{+}24] Wen zhi Liang, Pei qing Liu, Jin Zhang, Shu tong Yang, and Qiu lin Qu. A mathematical interpolation bounce back wall modeled lattice Boltzmann method based on hierarchical Cartesian mesh applied to 30P30N airfoil aeroacoustics simulation. Computers and Mathematics with Applications, 158 (??):21–35, March 15, 2024. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122124000087.

Zhang:2021:FVM

[ZLT^{+}21] Mengchen Zhang, Fawang Liu, Ian W. Turner, Vo V. Anh, and Libo Feng. A finite volume method for the two-dimensional time and space variable-order fractional Bloch-Torrey equation with variable coefficients on irregular domains. Computers and Mathematics with Applications, 98(??): 81–98, September 15, 2021. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122121002510.

Zhou:2023:EMD

[ZLW23] Zhongguo Zhou, Dong Liang, and Yaushu Wong. An efficient MMOCAA-DDM for solving advection diffusion equations. Computers and Mathematics with Applications, 131(??):35–53, February 1, 2023. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122122004990.

Zang:2020:IBE

[ZLYL20] Quansheng Zang, Jun Liu, Wenbin Ye, and Gao Lin. Isogeometric boundary element for analyzing steadystate heat conduction problems under spatially varying conductivity and internal heat source. Computers and Mathematics with Applications, 80(7):1767–1792, October 1, 2020. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122120303163.

Zhao:2025:MAB

[ZLZ^{+}25a] Jingui Zhao, Guirong Liu, Jinhui Zhao, Gang Wang, Zhonghu Wang, and Zirui Li. Modal analysis of biological structures based on the smoothed finite element methods. Computers and Mathematics with Applications, 191(??):188– 227, August 1, 2025. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122125002020.

Zhao:2025:HOL

[ZLZ25b] Lei Zhao, Dong Liang, and Zhiyue Zhang. High-order linearly implicit energystable schemes for twodimensional Navier–Stokes equations in vorticity-velocity formulation. Computers and Mathematics with Applications, 196(??):338–354, October 15, 2025. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122125003372.

Zaky:2020:MDS

[ZM20] Mahmoud A. Zaky and J. Tenreiro Machado. Multidimensional spectral tau methods for distributedorder fractional diffusion equations. Computers and Mathematics with Applications, 79(2):476–488, January 15, 2020. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122119303463.

Zhou:2021:LON

[ZMS21] Xinchen Zhou, Zhaoliang Meng, and Jing Su. Loworder nonconforming brick elements for the 3D Brinkman model. Computers and Mathematics with Applications, 98(??):201–217, September 15, 2021. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122121002728.

Zhu:2024:LRA

[ZMZW24] Yujun Zhu, Ju Ming, Jie Zhu, and Zhongming Wang. Low rank approximation method for perturbed linear systems with applications to elliptic type stochastic PDEs. Computers and Mathematics with Applications, 173(??): 60–73, November 1, 2024. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122124003456.

Zhang:2024:VIM

[ZNC24] Zhizhuo Zhang, Xiaobing Nie, and Jinde Cao. Variational inequalities of multilayer elastic contact systems with interlayer friction: Existence and uniqueness of solution and convergence of numerical solution. Computers and Mathematics with Applications, 174(??): 248–260, November 15, 2024. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122124003973.

Zheng:2024:LMS

[ZNH24] Zhoushun Zheng, Xinyue Ni, and Jilong He. Lagrange multiplier structurepreserving algorithm for time-fractional Allen–Cahn equation. Computers and Mathematics with Applications, 164(??):67–78, June 15, 2024. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122124001445.

Zhou:2023:SSN

[ZNMS23] Xinchen Zhou, Hexin Niu, Zhaoliang Meng, and Jing Su. Superconvergence of some nonconforming brick elements for the 3D Stokes problem. Computers and Mathematics with Applications, 151(??):50–66, December 1, 2023. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122123004066.

Zhang:2021:UES

[ZO21] Chenhui Zhang and Jie Ouyang. Unconditionally energy stable second-order numerical schemes for the functionalized Cahn–Hilliard gradient flow equation based on the SAV approach. Computers and Mathematics with Applications, 84(??):16–38, February 15, 2021. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122120304624.

Zafeiris:2024:OIA

[ZP24] Spiros Zafeiris and George Papadakis. An overset interpolation algorithm for multiphase flows using 3D multiblock polyhedral meshes. Computers and Mathematics with Applications, 161 (??):155–173, May 1, 2024. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122124000993.

Zhao:2020:SDG

[ZPC20] Lina Zhao, Eun-Jae Park, and Eric T. Chung. Staggered discontinuous Galerkin methods for the Helmholtz equation with large wave number. Computers and Mathematics with Applications, 80(12):2676–2690, December 15, 2020. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122120303813.

Zhao:2022:PRS

[ZPC22] Lina Zhao, Eun-Jae Park, and Eric Chung. A pressure robust staggered discontinuous Galerkin method for the Stokes equations. Computers and Mathematics with Applications, 128(??): 163–179, December 15, 2022. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122122004424.

Zhao:2022:SCC

[ZPK22] Lina Zhao, Eun-Jae Park, and Wonjong Kim. A staggered cell-centered DG method for the biharmonic Steklov problem on polygonal meshes: a priori and a posteriori analysis. Computers and Mathematics with Applications, 117(??):216– 228, July 1, 2022. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122122001742.

Zhou:2025:NCG

[ZQ25] Jintao Zhou and Guanghui Qing. Non-conforming generalized mixed element methods based on the volume coordinate system. Computers and Mathematics with Applications, 182(??):236– 255, March 15, 2025. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122125000355.

Zhang:2020:DGM

[ZQM20] Yongchao Zhang, Yanxia Qian, and Liquan Mei. Discontinuous Galerkin methods for the Stokes equations with nonlinear damping term on general meshes. Computers and Mathematics with Applications, 79(8): 2258–2275, April 15, 2020. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S089812211930522X.

Zhou:2024:REA

[ZQS24] Yongtao Zhou, Hongyu Qin, and Martin Stynes. An α-robust error analysis of the L̄1 scheme for timefractional integro–differential initial-boundary value problems. Computers and Mathematics with Applications, 162 (??):196–205, May 15, 2024. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122124001226.

Zhang:2023:HMD

[ZQY^{+}23] Jiansong Zhang, Rong Qin, Yun Yu, Jiang Zhu, and Yue Yu. Hybrid mixed discontinuous Galerkin finite element method for incompressible wormhole propagation problem. Computers and Mathematics with Applications, 138 (??):23–36, May 15, 2023. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122123000895.

Zhu:2020:SHO

[ZQZ20] Hongqiang Zhu, Jianxian Qiu, and Jun Zhu. A simple, high-order and compact WENO limiter for RKDG method. Computers and Mathematics with Applications, 79(2):317–336, January 15, 2020. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122119303372.

Zhang:2022:GMF

[ZR22] Jing Zhang and Hongxing Rui. Galerkin method for the fully coupled quasi-static thermo-poroelastic problem. Computers and Mathematics with Applications, 118 (??):95–109, July 15, 2022.

CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122122001754.

Zdunek:2023:WMD

[ZR23] Adam Zdunek and Waldemar Rachowicz. To ℘ or not to p — the mixed displacement–pressure p, versus the higher order ℘ displacement finite element formulation, for nearly incompressible linear elasticity. Computers and Mathematics with Applications, 148(??): 313–337, October 15, 2023. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S089812212300370X.

Zhang:2022:STB

[ZRL^{+}22] Yuhui Zhang, Timon Rabczuk, Jun Lu, Shifa Lin, and Ji Lin. Space-time backward substitution method for nonlinear transient heat conduction problems in functionally graded materials. Computers and Mathematics with Applications, 124(??): 98–110, October 15, 2022. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122122003467.

Zheng:2021:LPF

[ZS21] Bo Zheng and Yueqiang Shang. Local and parallel finite element algorithms based on domain decomposition for the 2D/3D Stokes equations with damping. Computers and Mathematics with Applications, 103(??): 82–103, December 1, 2021. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122121003825.

Zhang:2023:DUE

[ZS23a] Xiaodi Zhang and Haiyan Su. A decoupled, unconditionally energy-stable and structure-preserving finite element scheme for the incompressible MHD equations with magnetic-electric formulation. Computers and Mathematics with Applications, 146(??):45–59, September 15, 2023. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S089812212300281X.

Zheng:2023:PFE

[ZS23b] Bo Zheng and Yueqiang Shang. A parallel finite element method based on fully overlapping domain decomposition for the steadystate Smagorinsky model. Computers and Mathematics with Applications, 147 (??):76–91, October 1, 2023. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122123003048.

Zhou:2024:NFF

[ZS24] Yanhui Zhou and Shuai Su. A novel family of Q_{1}-finite volume element schemes on quadrilateral meshes. Computers and Mathematics with Applications, 172(??):216– 240, October 15, 2024. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122124003742.

Zeng:2025:SBK

[ZS25] Min-Li Zeng and Martin Stoll. A splittingbased KPIK method for eddy current optimal control problems in an all-at-once approach. Computers and Mathematics with Applications, 190(??):1–15, July 15, 2025. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122125001397.

Zhang:2023:SOR

[ZSF23] Zhe Zhang, Haiyan Su, and Xinlong Feng. Second-order rotational velocity correction projection finite element method for unsteady MHD coupled heat equation. Computers and Mathematics with Applications, 144(??): 306–322, August 15, 2023. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122123002742.

Zhang:2024:ODD

[ZSGT24] Jinjin Zhang, Yanru Su, Xinfeng Gao, and Xuemin Tu. Overlapping domain decomposition methods for finite volume discretizations. Computers and Mathematics with Applications, 174(??):510– 529, November 15, 2024. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122124004619.

Zeidan:2021:PSI

[ZSH^{+}21] Dia Zeidan, Theodore E.

Simos, Charis Harley, Ashoke De, and Eric Goncalves. Preface to special issue of selected papers from the 13th International Symposium on Numerical Analysis of Fluid Flow, Heat and Mass Transfer — Numerical Fluids 2018. Computers and Mathematics with Applications, 83 (??):1–3, February 1, 2021. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122121000080.

Zhang:2023:MSM

[ZSK23] Shengliang Zhang, Zhengjie Sun, and Alpesh Kumar. Meshless symplectic and multi-symplectic scheme for the coupled nonlinear Schrödinger system based on local RBF approximation. Computers and Mathematics with Applications, 134(??):16–32, March 15, 2023. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122122005260.

Zhou:2022:CSM

[ZSL22] Rui-Rui Zhou, Ya-Song Sun, and Ben-Wen Li. The collocation spectral method with domain decomposition for radiative heat transfer in two-dimensional enclosures. Computers and Mathematics with Applications, 123(??): 204–215, October 1, 2022. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122122003285.

Zhou:2020:ECF

[ZSPW20] Zhongguo Zhou, Xinxin Sun, Hao Pan, and Yan Wang. An efficient characteristic finite difference S-DDM scheme for convection– diffusion equations. Computers and Mathematics with Applications, 80(12):3044– 3065, December 15, 2020.

CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122120304259.

Zhou:2022:TLP

[ZST22] Xianghai Zhou, Haiyan Su, and Bo Tang. Twolevel Picard coupling correction finite element method based on charge-conservation for stationary inductionless magnetohydrodynamic equations. Computers and Mathematics with Applications, 115(??):41–56, June 1, 2022. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122122001535.

Zhao:2023:TDP

[ZSW23] Meiling Zhao, Jieyu Shi, and Liqun Wang. A three-dimensional Petrov– Galerkin finite element interface method for solving inhomogeneous anisotropic Maxwell’s equations in irregular regions. Computers and Mathematics with Applications, 152(??):364– 377, December 15, 2023. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122123004881.

Zhou:2022:FVS

[ZSY22] Huifang Zhou, Zhiqiang Sheng, and Guangwei Yuan. A finite volume scheme preserving the invariant region property for the coupled system of FitzHugh-Nagumo equations on distorted meshes. Computers and Mathematics with Applications, 117(??):39– 52, July 1, 2022. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122122001663.

Zhang:2022:FDA

[ZTL22] Lei Zhang, Shaoqiang Tang, and Baiyili Liu. Finite difference approach for multiscale computations of atomic chain at finite temperature. Computers and Mathematics with Applications, 110 (??):77–90, March 15, 2022. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122122000475.

Zhang:2022:WBS

[ZTW22] Shengyuan Zhang, Jun Tang, and Huiying Wu. Wetting boundary schemes in modified phase-field lattice Boltzmann method for binary fluids with large density ratios. Computers and Mathematics with Applications, 113 (??):243–253, May 1, 2022. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122122001134.

Zandbergen:2025:IPT

[ZvNH25] Anouk Zandbergen, Tycho van Noorden, and Alexander Heinlein. Improving pseudo-time stepping convergence for CFD simulations with neural networks. Computers and Mathematics with Applications, 196 (??):64–83, October 15, 2025. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122125002871.

Zhou:2020:FQF

[ZW20a] Yanhui Zhou and Jiming Wu.

A family of quadratic finite volume element schemes over triangular meshes for elliptic equations. Computers and Mathematics with Applications, 79(9):2473– 2491, May 1, 2020. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122119305504.

Zou:2020:RSD

[ZW20b] Chengye Zou and Xingyuan Wang. Robust stability of delayed Markovian switching genetic regulatory networks with reaction– diffusion terms. Computers and Mathematics with Applications, 79(4):1150–1164, February 15, 2020. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122119304274.

Zak:2021:SBF

[ŻW21a] A. Żak and W. Waszkowiak. A spline-based FE approach to modelling of high frequency dynamics of 1-D structures. Computers and Mathematics with Applications, 104(??):14–33, December 15, 2021. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122121003941.

Zhou:2021:HOL

[ZW21b] Yanhui Zhou and Jiming Wu. High order locally conservative finite element solutions for anisotropic diffusion problems in two dimensions. Computers and Mathematics with Applications, 92(??):1–12, June 15, 2021. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122121000973.

Zhang:2025:TDL

[ZW25] Houchao Zhang and Junjun Wang. Two decoupled and linearized structurepreserving FEMs for the Klein–Gordon–Schrödinger equations with damping mechanism. Computers and Mathematics with Applications, 195(??):68–92, October 1, 2025. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122125002780.

Zhao:2020:BTR

[ZWCS20] Yong Zhao, Yao Wu, Zhenhua Chai, and Baochang Shi. A block triplerelaxation-time lattice Boltzmann model for nonlinear anisotropic convection– diffusion equations. Computers and Mathematics with Applications, 79(9):2550– 2573, May 1, 2020. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122119305516.

Zhai:2024:EDM

[ZWMF24] Shuying Zhai, Zhifeng Weng, Yuchang Mo, and Xinlong Feng. Energy dissipation and maximum bound principle preserving scheme for solving a nonlocal ternary Allen–Cahn model. Computers and Mathematics with Applications, 155(??):150– 164, February 1, 2024. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122123005564.

Zapletal:2021:SAI

[ZWOM21] Jan Zapletal, Raphael Watschinger, Günther Of, and Michal Merta. Semi-analytic integration for a parallel space-time boundary element method modelling the heat equation. Computers and Mathematics with Applications, 103(??):156–170, December 1, 2021. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122121003795.

Zhou:2022:ILB

[ZWS22] Min Zhou, JinRong Wang, and Dong Shen. Iterative learning based consensus control for distributed parameter type multi-agent differential inclusion systems with time-delay. Computers and Mathematics with Applications, 127(??):25–47, December 1, 2022. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S089812212200400X.

Zhang:2023:HED

[ZWT23] Xin Zhang, Jingwen Wu, and Zhijun Tan. Highly efficient, decoupled and unconditionally stable numerical schemes for a modified phase-field crystal model with a strong nonlinear vacancy potential. Computers and Mathematics with Applications, 132(??): 119–134, February 15, 2023. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122122005119.

Zhou:2023:NDO

[ZWW^{+}23] Xingyu Zhou, Haoping Wang, Ke Wu, Yang Tian, and Gang Zheng. Nonlinear disturbance observerbased robust predefined time tracking and vibration suppression control for the rigid-flexible coupled robotic mechanisms with large beam-deformations. Computers and Mathematics with Applications, 148 (??):1–25, October 15, 2023. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122123003334.

Zheng:2024:APS

[ZWX24] Fenglian Zheng, Yan Wang, and Xufeng Xiao. Accurate and parallel simulation of the anisotropic dendrite crystal growth by Lagrangian data assimilation with directional operator splitting. Computers and Mathematics with Applications, 175(??): 416–432, December 1, 2024. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122124004644.

Zhou:2023:DNN

[ZWY23] Zijian Zhou, Li Wang, and Zhenya Yan. Deep neural networks learning forward and inverse problems of two-dimensional nonlinear wave equations with rational solitons. Computers and Mathematics with Applications, 151(??):164–171, December 1, 2023. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122123004388.

Zeng:2024:NFI

[ZWZD24] Fang Zeng, Jiajia Wang, Shuang Zhou, and Haiyun Dong. Near-field imaging method for interior inverse elastic scattering problem. Computers and Mathematics with Applications, 168 (??):10–21, August 15, 2024. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122124001950.

Zhou:2020:ADI

[ZX20] Jun Zhou and Da Xu. Alternating direction implicit difference scheme for the multiterm time-fractional integrodifferential equation with a weakly singular kernel. Computers and Mathematics with Applications, 79(2): 244–255, January 15, 2020. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S089812211930330X.

Zhu:2022:SPS

[ZX22] Peng Zhu and Shenglan Xie. Supercloseness and postprocessing of stabilizerfree weak Galerkin finite element approximations for parabolic problems. Computers and Mathematics with Applications, 119(??):79–88, August 1, 2022. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122122002322.

Zhang:2024:NAV

[ZX24] Jianhua Zhang and Guo Xu.

A novel α-absolute value preconditioner for all-at-once systems from heat equations. Computers and Mathematics with Applications, 165 (??):196–204, July 1, 2024. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122124001962.

Zhong:2024:TAP

[ZXG^{+}24] Yudong Zhong, Guizhong Xie, Hongrui Geng, Junjian Hou, Dengfeng Zhao, and Wenbin He. Thermal analysis for plate structures using a transformation BEM based on complex poles. Computers and Mathematics with Applications, 161(??):32–42, May 1, 2024. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122124000798.

Zhang:2022:NIN

[ZXGF22] Tao Zhang, Hui Xu, Lei Guo, and Xinlong Feng. A non-intrusive neural network model order reduction algorithm for parameterized parabolic PDEs. Computers and Mathematics with Applications, 119 (??):59–67, August 1, 2022. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122122002280.

Zhou:2022:ARE

[ZXQQ22] Jun Zhou, Da Xu, Wenlin Qiu, and Leijie Qiao. An accurate, robust, and efficient weak Galerkin finite element scheme with graded meshes for the time-fractional quasilinear diffusion equation. Computers and Mathematics with Applications, 124(??): 188–195, October 15, 2022. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S089812212200342X.

Zhang:2023:HOS

[ZXXX23] Weijie Zhang, Yulong Xing, Yinhua Xia, and Yan Xu. High order structurepreserving arbitrary Lagrangian– Eulerian discontinuous Galerkin methods for the Euler equations under gravitational fields. Computers and Mathematics with Applications, 146(??):339–359, September 15, 2023. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122123003292.

Zhao:2025:NMS

[ZXZ25] Yedan Zhao, Yinghong Xu, and Lipu Zhang. Numerical methods for solving the inverse problem of 1D and 2D PT-symmetric potentials in the NLSE. Computers and Mathematics with Applications, 183(??):137– 152, April 1, 2025. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122125000306.

Zhao:2020:PGF

[ZXZF20] Shubo Zhao, Xufeng Xiao, Jianping Zhao, and Xinlong Feng. A Petrov–Galerkin finite element method for simulating chemotaxis models on stationary surfaces. Computers and Mathematics with Applications, 79(11): 3189–3205, June 1, 2020. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122120300353.

Zhang:2021:GLE

[ZY21a] Yu Zhang and Yidu Yang. Guaranteed lower eigenvalue bounds for two spectral problems arising in fluid mechanics. Computers and Mathematics with Applications, 90 (??):66–72, May 15, 2021. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122121000730.

Zhu:2021:MDC

[ZY21b] Qiming Zhu and Jinhui Yan. A moving-domain CFD solver in FEniCS with applications to tidal turbine simulations in turbulent flows. Computers and Mathematics with Applications, 81(??): 532–546, January 1, 2021. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122119303906.

Zhang:2023:AUM

[ZY23] Biao Zhang and Yin Yang. An adaptive unconditional maximum principle preserving and energy stability scheme for the space fractional Allen–Cahn equation. Computers and Mathematics with Applications, 139 (??):28–37, June 1, 2023. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122123000810.

Zhang:2024:IDR

[ZYC^{+}24] Senlin Zhang, Bo Yu, Leilei Chen, Haojie Lian, and Stephane P. A. Bordas. Isogeometric dual reciprocity BEM for solving timedomain acoustic wave problems. Computers and Mathematics with Applications, 160 (??):125–141, April 15, 2024. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122124000683.

Zhang:2024:NAT

[ZYD24] Yunzhang Zhang, Xinghui Yong, and Xiaogang Du. Numerical analysis of time filter method for the stabilized incompressible diffusive Peterlin viscoelastic fluid model. Computers and Mathematics with Applications, 168(??): 239–253, August 15, 2024. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S089812212400302X.

Zhang:2024:MMS

[ZYM24] Lei Zhang, Enjian Yang, and Yulin Miao. Mesoscopic modeling of silver particles melting process on silicon wafer in inkjet-printed solar cell metallization. Computers and Mathematics with Applications, 156(??):56–63, February 15, 2024. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122123005679.

Zhang:2024:NSM

[ZYQ^{+}24] Ang Zhang, Minghang Yang, Lang Qin, Jing Cheng, Yuchen Tang, Jinglian Du, Wenbo Yu, Zhihua Dong, Feng Liu, Bin Jiang, and Fusheng Pan. Numerical study of magnesium dendrite microstructure under convection: Change of dendrite symmetry. Computers and Mathematics with Applications, 176(??):289– 305, December 15, 2024. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122124004875.

Zhao:2022:ILB

[ZYS22] Xiang Zhao, Liming Yang, and Chang Shu. An implicit lattice Boltzmann flux solver for simulation of compressible flows. Computers and Mathematics with Applications, 107(??):82–94, February 1, 2022. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S089812212100448X.

Zhang:2022:REA

[ZYTX22] Haixiang Zhang, Xuehua Yang, Qiong Tang, and Da Xu. A robust error analysis of the OSC method for a multi-term fourth-order subdiffusion equation. Computers and Mathematics with Applications, 109(??):180– 190, March 1, 2022. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122122000074.

Zhu:2020:PDP

[ZYW20] Jun-Li Zhu, Ai-Li Yang, and Yu-Jiang Wu. A parameterized deteriorated PSS preconditioner and its optimization for nonsymmetric saddle point problems. Computers and Mathematics with Applications, 79(5): 1420–1434, March 1, 2020. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122119304511.

Zhao:2020:LRS

[ZZ20a] Jianxi Zhao and Lina Zhao. Low-rank and sparse matrices fitting algorithm for lowrank representation. Computers and Mathematics with Applications, 79(2):407–425, January 15, 2020. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122119303505.

Zhao:2020:SVT

[ZZ20b] Jin Zhao and Zhimin Zhang. Seven-velocity three-dimensional vectorial lattice Boltzmann method including various types of approximations to the pressure and twoparameterized second-order boundary treatments. Computers and Mathematics with Applications, 80(12):2764– 2779, December 15, 2020. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122120303886.

Zhang:2021:FII

[ZZ21] Min Zhang and Guo-Feng Zhang. Fast image inpainting strategy based on the space-fractional modified Cahn–Hilliard equations. Computers and Mathematics with Applications, 102(??): 1–14, November 15, 2021. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122121003539.

Zhang:2023:SFE

[ZZ23a] Chunxiao Zhang and Jin Zhang. Supercloseness of finite element method for a singularly perturbed convection-diffusion problem on bakhvalov-type triangular meshes. Computers and Mathematics with Applications, 150(??):243–253, November 15, 2023. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122123004133.

Zhang:2023:MIM

[ZZ23b] Jianjun Zhang and Xi Zhang.

A modulus iteration method for non-negatively constrained TV image restoration. Computers and Mathematics with Applications, 148(??):62–69, October 15, 2023. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122123003309.

Zhang:2024:UCF

[ZZ24a] Chunxiao Zhang and Jin Zhang. Uniform convergence of finite element method on Bakhvalov-type triangular mesh for a singularly perturbed problem featuring exponential layers. Computers and Mathematics with Applications, 160(??):171– 180, April 15, 2024. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122124000749.

Zhang:2024:NEA

[ZZ24b] Xiaodi Zhang and Xianghai Zhou. New error analysis of charge-conservative finite element methods for stationary inductionless MHD equations. Computers and Mathematics with Applications, 157 (??):147–158, March 1, 2024. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122123005977.

Zhou:2024:NSS

[ZZ24c] Shiping Zhou and Yanzhi Zhang. A novel and simple spectral method for nonlocal PDEs with the fractional Laplacian. Computers and Mathematics with Applications, 168(??):133– 147, August 15, 2024. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122124002621.

Zhang:2025:LWF

[ZZ25a] Tao Zhang and Xiaofeng Zhou. A local weak form of the generalized finite difference method (GFDM) with control volume in heat conduction problems. Computers and Mathematics with Applications, 199(??):203– 224, December 1, 2025. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122125003979.

Zhao:2025:UCF

[ZZ25b] Xianyang Zhao and Jin Zhang. Uniform convergence of finite element method on Vulanović–Bakhvalov mesh for singularly perturbed convection–diffusion equation in 2D. Computers and Mathematics with Applications, 188(??):183– 194, June 15, 2025. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S089812212500152X.

Zhou:2020:IDS

[ZZB20] Yongtao Zhou, Chengjian Zhang, and Luigi Brugnano. An implicit difference scheme with the KPS preconditioner for two-dimensional time– space fractional convection– diffusion equations. Computers and Mathematics with Applications, 80(1):31– 42, July 1, 2020. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S089812212030081X.

Zhang:2020:LFN

[ZZCM20] Jingyuan Zhang, Chenguang Zhou, Yanzhao Cao, and A. J. Meir. A locking free numerical approximation for quasilinear poroelasticity problems. Computers and Mathematics with Applications, 80(6):1538–1554, September 15, 2020. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122120302777.

Zhu:2021:ESO

[ZZFL21] Chen Zhu, Bingyin Zhang, Hongfei Fu, and Jun Liu. Efficient second-order ADI difference schemes for threedimensional Riesz spacefractional diffusion equations. Computers and Mathematics with Applications, 98(??):24–39, September 15, 2021. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122121002546.

Zhang:2023:PEE

[ZZH23] Xindan Zhang, Jianping Zhao, and Yanren Hou. A priori error estimates of Crank–Nicolson finite element method for parabolic optimal control problems.

Computers and Mathematics with Applications, 144(??): 274–289, August 15, 2023. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122123002687.

Zhang:2022:TLR

[ZZL22] Lei Zhang, Guo-Feng Zhang, and Zhao-Zheng Liang. Tensorized low-rank circulant preconditioners for multilevel Toeplitz linear systems from high-dimensional fractional Riesz equations. Computers and Mathematics with Applications, 110 (??):64–76, March 15, 2022. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122122000037.

Zhang:2023:MBN

[ZZL23] Limei Zhang, Hong Zheng, and Feng Liu. MLSbased numerical manifold method for steady-state three-dimensional heat conduction problems of functionally graded materials. Computers and Mathematics with Applications, 144(??): 124–140, August 15, 2023. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122123002353.

Zhang:2023:PFO

[ZZS23] Lu Zhang, Qifeng Zhang, and Hai-Wei Sun. Preconditioned fourth-order exponential integrator for two-dimensional nonlinear fractional Ginzburg–Landau equation. Computers and Mathematics with Applications, 150(??):211–228, November 15, 2023. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122123004121.

Zhou:2024:LPF

[ZZS24] Xianghai Zhou, Xiaodi Zhang, and Haiyan Su. Local and parallel finite element algorithms based on charge-conservation approximation for the stationary inductionless magnetohydrodynamic problem. Computers and Mathematics with Applications, 154(??):224– 235, January 15, 2024. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122123005473.

Zhang:2020:NUT

[ZZW20] Jun Zhang, Jia Zhao, and JinRong Wang. A nonuniform time-stepping convex splitting scheme for the time-fractional Cahn– Hilliard equation. Computers and Mathematics with Applications, 80(5):837–850, September 1, 2020. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S089812212030184X.

Zhou:2023:PFV

[ZZW23] Yanhui Zhou, Yanlong Zhang, and Jiming Wu. A polygonal finite volume element method for anisotropic diffusion problems. Computers and Mathematics with Applications, 140(??):225– 236, June 15, 2023. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122123001670.

Zhang:2025:UES

[ZZW25] Dan Zhang, YuXing Zhang, and Bo Wang. Unconditionally energy-stable discontinuous Galerkin method for the dynamics model of protein folding. Computers and Mathematics with Applications, 192(??):14–36, August 15, 2025. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122125001920.

Zhao:2024:HOA

[ZZY^{+}24] Xiaolong Zhao, Shijun Zou, Xijun Yu, Dongyang Shi, and Shicang Song. A highorder arbitrary Lagrangian– Eulerian discontinuous Galerkin method for compressible flows in two-dimensional Cartesian and cylindrical coordinates. Computers and Mathematics with Applications, 170(??):95–115, September 15, 2024. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S089812212400289X.

Zhang:2024:PDA

[ZZZ24] Xuetian Zhang, Chunrui Zhang, and Yazhuo Zhang. Pattern dynamics analysis of a time–space discrete FitzHugh–Nagumo (FHN) model based on coupled map lattices. Computers and Mathematics with Applications, 157(??):92–123, March 1, 2024. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122123005928.

Zhang:2025:APN

[ZZZ25] Xiaodi Zhang, Meiying Zhang, and Xianghai Zhou. Analysis of the Picard– Newton finite element iteration for the stationary incompressible inductionless MHD equations. Computers and Mathematics with Applications, 182(??):46–65, March 15, 2025. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122125000161.

Zhai:2024:NBP

[ZZZC24] Qinglan Zhai, Lin Zheng, Song Zheng, and Hutao Cui. A neutrally buoyant particle motion in a double-lid-driven square cavity. Computers and Mathematics with Applications, 161(??):51–62, May 1, 2024. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122124000737.

Zhao:2024:HTV

[ZZZD24] Jikun Zhao, Fengchen Zhou, Bei Zhang, and Xiaojing Dong. The Hermite-type virtual element method for second order problem. Computers and Mathematics with Applications, 172(??):70–77, October 15, 2024. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122124003341.